A Complete Bibliography of Lecture Notes in Mathematics (1985–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.04

Title word cross-reference

(1 < p ≤ ∞) [LS87f]. (2) [HR88a]. (2m − 2) [KL88]. (A_{0}, A_{1})_{θ1} [Xu87a]. (α, β) [Pie88a, Fin88a]. (d ≥ 1) [Wsc85a]. (λ)^{κ} [DM85b]. (Z/2)^{k} [Car86b]. (nα) [Sch85h]. (ϕ^{4})_{2} [BM89c]. (τ − λ)u = f [Wei87r]. (x, t) [Lum87, Lum89]. (X_{1} − X_{3}, X_{2} − X_{3}) [SW87]. 0 [Caz88, Kas86, Pro87]. 0 < p < 1 [Cle87]. 1 [Bak85a, DD85, Drm87, Eli88, FT88a, Gek86d, HN88, Kos86a, LT89, Pet89a, Pro87, Tan87, vdG89]. 1/4 [KS86e]. 1 ≤ q < 2 [Gue86]. 2 [BPPS87, Cam88, Cat85a, ES87b, Gan85e, Gol86a, HRL89g, Hei85, Hua86, Kan89, KB86, Li86, LT89, Mil87b, Mur85a, Qui85b, SP89, Shi85, Spe86, Wal85b, Wan86]. 2m − 2 [Kos88b]. 2m − 3 [Kos88b]. 2m − 4 [Kos88b]. 2 × 2 [Vog88]. 3 [Aso89, BPPS87, BW85c, BG88b, Che86d, Fis86, Gab85, Gu87a, HLM85b, Kam89b, Kir89c, Lev85c, Mil85c, Néd86, Pet86, Ron86, Sch85b, ST88, Tur88b, Wan86, Wen85, tDP89, vdW86]. 4 [Bau88a, Don85a, FKV88, Kha88, Kir89l, SS86, Seg85b, Wal85b]. 5 [Ito89, Kir89e, SV85]. 5(4) [Cas86]. 5819539783680 [KSX87]. 6 [PH89, Žub88].

7 [Bry85, DM89b]. [−1, 1] [LS87f]. [0, ∞) [Lub85]. [λ]^{<κ} [Mat89a]. _{1}^{0} [Kuč85]. _{2}^{1} [Gri85a]. _{3} [D’O85]. _{m}^{⟨k⟩} → ⊖_{m−1} [Sto85e]. _{R} [Kec88a]. A [GSZ87, Fer88a]. A − B = C [Fre89]. A[k] [Sto85h]. A_{∗}^{⟨k⟩} [Sto85c]. A_{1}(C) [Ale86]. A_{n} [Con86, Shi86c, Shi86g, Shi86m, Shi86a, Shi86t, Shi86r, Shi86u, Wie86, Shi86n]. a_{n} → −1/4 [Jac86a]. a_{n} → ∞ [JJW87, SW86b]. ℵ_{0} [Bus88b]. ℵ_{1} [Kec88c]. α [DM85c, GN88, Ski86]. B [CAXD86g, Hej88]. μ̄ [Lev88]. ¯∂ [BS87b, Ran87]. ¯∂_{b} [Chr89, Mai88]. ¯∂u = f [Ama87]. BC_{p} [Deb87]. β [DM85c]. BG [DW88]. mod2 [KI87]. modp [McC86a]. bP_{m} [Sto85a]. BS^{3} [DMW87]. BZ/pZ [DW88]. C [BDT87b, Gek86g, Shi86h, San87]. C(S × T) [LC85h]. C∗ [Ara85, AB85, BDE85, Bra86g, Eva85a, Eva85b, Goo85, HO85, Lud88, Wan87c, Wan87g, dlH85]. C ∗ (G) [Pat88]. c^{∗} [Gro86b]. C^{0} [Shi87a]. C^{2} [Lev85b, Rea86]. C^{∞} [Lun86a, Miz88, jQxQ88]. C^{k} [lS87d]. C^{n} [CS87b, Rea86, Sae87]. C^{ω} [Shi87b, Shi87f]. C^{r} [Kif88a, Shi87c]. C_{0} [Tim88, AOR88]. C_{∞} [Sch85e]. c_{n+p}/c_{n} → c [vdB87c]. C_{o} [Sam86, AS85f]. C_{o}(X) [Are86a, Gre86b, Gre86d, NS86c]. cat^{n} [Gil87]. χ → +∞

[IN86k, IN86j]. χ → −∞ [IN86j]. CP^{2} ♯^{9}(−CP^{2}) [Kir89b]. CP^{n} [BDK87]. D [CL85, Wsc85a]. D(f) [Pre88m]. D_{2}(k) [Cau86b]. D_{n} [Con86]. δ [Coo89]. Diff^{Ω}(X × R^{+}, relX × [0]) [CM89a]. E [Sla85a, Con89]. E^{3} [Yan89]. E_{1} [Sto85h]. E_{2} [CGMM87]. E_{6} [Don85b]. E_{7} [Don85c]. E_{8} [Don85d]. E_{np}(W) [LS88o]. ℓ [JM85]. ℓ^{P} [Nei85, KK85]. ℓ_{1} [ADR85]. ℓ_{1}^{n} [MS86h]. ℓ_{p}^{m} [MS86h]. ℓ_{p}^{m} → ℓ_{q}^{m} [Mak88]. ℓ_{p}^{n} [BT87b, MS86d]. ℓ_{r}^{n} [Sch87d]. ϵ [LM89b, Sch89c]. η [GS89, Jin88a]. F [CL85, LM88a, Met85, Raw85, Shi88b, Tay88, Fut88c, Fut88b, Fut88d, Fut88e, Fut88h]. f^{(k)}(σ; τ_{i}) [Wid85a]. F^{∞} [BR88]. F_{4} [Don85e]. F_{p} [CC88]. F_{q} [Lus87]. G [Ass89, CC85a, Kat86b, Kaw89, KK89b, Liu88, Pat88, Ric87, DR89b, MR85a, PM89]. G^{1}(G) [Rei89e, Rei89j]. G_{0} [Lor88]. G_{2} [Bry85, Don85f]. G_{n,k} [Lev89c, Lev89e, Lev89l]. galL/K [Tho86b]. Γ [Fra86, Koź86]. Γ_{0} [Shi89]. Γ_{2} ≥ 0 [Bak85c, Bak85b]. Γ_{o}(N) [Kno85]. GL(n, F_{P}) [Woo85]. GL_{2} [Kai87c, Kai87b]. GL_{n} [Kai87a, Kai87c]. G × GL(n) [GPS87a]. H [Fin88c, Kir89q, Mül87k, AD85b, Fou88]. H(curl) [Néd86]. H(div) [Néd86]. H ∗ (BG; Z_{p}) [DLS85]. H^{1} [Fer86, Jan89, MS89c, Oro89]. H^{3} [Ma89]. H^{P} [Cle87, Fef88, Ill88]. H′ [Jac85c]. H_{0} × H_{1} [Ped85]. H_{2}(A, ∂A) = 0 [Sch85b]. H_{∞} [Bru86a, Bru86b, Bru86c, BMMS86, May86a, McC86b, Ste86a]. H_{∞}^{d} [McC86c]. H_{n} [Ste86a]. H_{O}^{∞}(D) [AAGV85]. Ĵ [Ran86]. Hilb^{3}P^{2} [EL88]. Hom_{u_{p}}(, H^{∗}V) [Zar85]. h\underline{o} lim F [Woj87]. I∗ [Wt87d, Wt87a, Wt87b]. II_{1} [Cho85a]. ℑJ [CP89c]. ∫_{0}^{∞} (sin(\frac{π}{4} + θ)e^{−θy} + sin(\frac{π}{4} − θ)e^{θy})π(x, y) dy = √2coshθx [McG88b]. ∫_{X} |f|^{2λ}□ [Bar86b]. IP^{3} [BE87]. IP_{(2)}^{m} [Le 86a]. IR^{3} [Oza85]. IR^{n} [MTJ87]. J [MT87, Sto85e]. J∗ [Mes85]. K [AS85h, Cle88, FT87a, Fre87, Gai88a, Hah86, Heg88, Hod89, JV85, Kuh89, McC85b, Mur85b, PW85b, Per85, Phi87b, Phi87c, Phi87f, Ric86, Vog85, Wal85a, McC86a, Nat89b]. K(a_{n}/1) [Jac86a, JT86, Jac86b, JJW87, SW86b]. K(π, l) [Dav85b]. K + K [Fre87]. K3 [Bar85c, Mar89b, Hoy87, PS89a]. k = n + j + 1/2 [Wei86d]. k[X] [Hel89].

K^{+} − p [AC87]. K_{−I} [MR89a]. K_{3} [Sil89f]. K_{4} [Con89]. κ [She86i]. KK [Phi87e]. K ⊂ Q(μ_{n}) [Tho86b]. L [And87a, CS85a, Dur87, Ebe86, GPSR87, GPS87a, HS85, Mor86f, PSR87, Ral87c, Sun86, Ral87d]. L(n) [Kuh87]. L^{1} [Di 86, Krá89, Ros88a]. L^{1}(G) [Joh88a]. L^{2} [And88, BW85a, CKS87a, Jaj88, Mül87a, Mül87h, Mül88, Pat87, Sul86]. L^{2}(Γ_{n}/G) [Wei86c]. L^{2}(R^{n}) [DR87]. l^{∞} [AS85f]. L^{∞}(G) [Joh88a]. L^{P} [Cie88, Cec85, Gue85, Nag87, Sil88b]. L^{q} [Gue86]. L_{1} [FFR88, LC85a]. L_{2} [Jaj89a, Wei87a]. L_{N}^{∞} [Bou85b]. L_{p} [AA89, Bas87, BKT89, GZ89, LS87f, LS88p, Mas88b]. L_{P} (μ) [Ill88]. L_{p}(X) [HLR85]. L_{s}^{p}(R^{2}) [Pon88]. Λ [Sin87]. λ_{1} [BBD88]. LC^{n} [Dra87b]. {t_{n}}_{n=1}^{∞} [CS89e]. l ≤ p < +∞ [Gue85]. M [BT86a, Coo86b, GH85a, Mei86b, She86m, Fin88a, KL88, Kos88b, Odi85, Sch87c]. M(ρ, s) [Wei86e]. M(T) [Que87a]. map_{∗}(BS^{3}, BS^{3}) [DM87b]. A^{n} [Pop87b]. C [Lev89a, Pou88]. C^{2} [Dem86b]. C^{N} [Kor87a, Cou89, D’A87a, Oka85]. C^{ν} [Éca85]. C^{P} [Szp86, Szp86]. G_{a} [Pop87b]. P [HS86]. P^{1} [SS86]. P^{3} [CO89, Idà89, Rao87]. P_{2} [Dre87, HH86]. R^{2} [Lev85c]. R^{3} [BC86b, Dum85b, Tro88, Cam88]. R^{4} [Lev85c, Lev85e]. R^{d} [Let89, Wsc85d]. R^{k} [Mil87a]. R^{n} [Els85c, Gor88]. RP^{3} [Kha88]. Z [AR87, CC88]. Z/p [Lew88b]. Z_{2} [Mas86b]. E [AA86]. R^{\} [Bou87d, Bou87e]. μ [Sch88k, LS87b]. N [Aga86a, BT86a, LS88l, Bar86e, BCZ87, Cho89a, Düm85a, GLM85, Gil87, MS86d, MP89b, Shi86b, Shi86n]. n > 2 [Shi86c, Wei87d]. n^{th} [BDT87a]. N_{1} [Sch89d]. N_{λ} [Shi86p, Shi86q]. no(M) [She86m]. n × n [For87]. O [Szp86]. O(n^{3}L) [Roo89]. ω^{2}S^{2n+1} [CM87d]. ω_{1} [Har87b]. ω_{4}^{so} = Z [Kir89a]. ω_{4}^{spin} = Z [Kir89a]. ωU(n) [Cra87]. P [AK87, Sch89g, BB88b, Dra87a, Ele87, GZ89, Gou88c, KQ88, Mad89, Not88, RW85, Shi86v, Vig89, Wil86a]. p = 2 [LS88f]. p > 0 [Smi86c]. P^{2} [Mar86c, ES87b, JM86a]. P^{3} [APS89, Her89]. P^{4} [vGW89]. P^{n} [Kle89]. p′ [Wil86a]. P^{v} [Sch89e]. p_{1} [Lev85h]. p_{1}(M) = 3σ(M) [Kir89a]. P_{κ}λ [CP89a, Zwi89]. ∂ [Sto85e]. ∂_{b} [Ber87c]. ∂: a_{m}^{⟨k⟩} → ⊖_{m−1} [Sto85a]. p ≥ 1 [BT87b]. ϕ [FJ88]. Π [Gri85a, Kuč85]. π_{∗} [Sto85n]. π_{∗}(ωS) ⊗ Q [Hal86, FHTT86]. Π_{3,2} [LV86a]. Π_{m,2} [Le 86b]. pq [Bar87c]. q = b + 1 [Gan85b]. Q_{m} [tC87]. QX [McC86a]. R [Fun88, Njå88]. R^{2} [Bar85d, LX89]. R^{4} [Kir89h]. R^{6} [Kir89l]. R^{d} [KM88]. R^{n+k} [Lev89f]. R_{K} [Han87f, Han87g]. R_{n} [Wan87a]. RΓ [Lüc89a]. rmGL_{2} [Kai87c]. rmGL_{n} [Kai87c]. RO(G) [Lew88b]. S [Abe87g, Abe87h, Abe87c, Ler86, Seg85b, AM88c, Kaw89]. S + 1 [Dra85]. S^{1} [Dyl88, Hat86, Jaw86, Kam89a, LX89, Pup88]. S^{2} × S^{2} [Kob87]. S^{3} [Jaw86]. S^{k} [Bré89b]. S^{n−1} [MS86k]. S_{n} [Wan87a]. S_{p,q} [CAXD86u]. σ^{−1}(λ) [Shi86k, Shi86a, Shi86t, Shi86q]. Σ_{1} [SW89b]. Σ_{2} [Cho85b, Cho89b]. SK_{1} [Son85]. SL(2, C) [Aso89]. SL(2, C) [Bar85b]. SL(3, p^{n}) [cY86]. sl(n, D) [Sel88g]. Spec(R) [Mum88a]. Sq^{i} [Woo87b]. SU(3, p^{n}) [cY86].

∑_{n=0}^{∞} c_{n} \frac{z^{n}}{n!} , ∑_{n=0}^{∞} c_{n}z^{n}and ∑_{n=0c_{n}n!z^{n}} [vdB87c, vdB87b]. ∑_{n≤NB_{2}(\{nα\})} [Sch87i]. T

[CS89b, CS89c, Cat85b, JW89, Val86a]. T 1 [FHJW89]. T^{K} [Cat85b]. T_{0} [Wei87g, Wei87q, Wei87t]. T_{2} [He89]. T_{Lie}^{3} [Kir89b]. T_{M} [CS89f]. T_{m,n} [Wer87b]. τ [Dec88, Shi86l]. θ [CDMS89]. Ã_{n} [Shi86m]. Ã_{n−1} [Shi86c]. top^{N} [HJ87a]. U [Gou88b]. U(n) [Cra87]. U(p, q)/U(p) × U(q) [Ank88]. u′ = AK ∗ u + BM ∗ u [Pug86]. U_{n,R}(x) [LS88g]. υ_{p} [KMM87]. UV^{k} [Fer87]. V [Ala88]. v + 2 [Sch89e]. V_{a} [BP88b]. v_{p} [PX87]. ϱ [Wei87f]. W [CAXD86g]. W (4, 2, 2) [Gol86a]. W (m: n) [Wil89]. W ∗ [Pas89]. w^{∗} [Gro86d, Gro86b, Gro86e]. W^{+} [MvR88]. X^{a} + Y^{b} + Z^{c} = 0 [Bar86d]. X_{1} [SW87]. x_{1}^{5} + x_{2}^{11} [LP88f]. X_{2} [SW87]. X_{3} [SW87]. X_{p}^{m} [Sch87d]. X_{pq}^{μ}(ω, ℓ) [Shi88a]. Ξ [Fis87c]. ξ(w, k) [Shi86p]. Z [Woj87]. Z^{d} [Sul86]. Z_{15} [Kor88]. Z_{∞} [Poh89a]. ζ(3) [PS89a].

* [LL85, Phi87c, Lev89a]. *-action [Lev89a]. *-algebras [Phi87c].

- [Con89, DM85c, Vog88, Bar85c, Kaw89]. -3 [Per85]. -actions [Car86b, Hat86, Kat86b, Ped85]. -adequate [Fou88]. -adic [Gou88c, KQ88, Dur87]. -adique [Vig89]. -algebra [Sin87]. -algebras [Gro86b, Gro86e, Eva85a, Wan87g, AB85, BDE85, Bra86g, Eva85b, Goo85, Gro86d, HO85, Mes85, Pas89, Wan87c, dlH85]. -algèbres [GSZ87]. -amenability [JV85]. -analytic [Pou88]. -and [Gro86b]. -anneaux [CL85, Val86a]. -Arithmetic [Abe87c, Abe87g, Abe87h]. -block [HN88, Wil86a]. -blocks [Shi86v]. -bound [lS87d]. -bounded [Nag87]. -Calderón [CDMS89]. -classes [Kuč85]. -cobordism [AM88c, Kaw89, Kir89q]. -codimensional [DD85]. -cohomology [Pat87]. -collection [SW89b]. -compacta [Dra87b]. -complete [Bus88b]. -complexes [Qui85b]. -connected [DM89b, FT88a]. -connectedness [Li86]. -convergence [Fra86]. -Convexity [Ros88a]. -core [Wil86a]. -cube [Gil87]. -D [BPPS87]. -decay [Ski86]. -degree [Dyl88]. -degrees [Odi85]. -dense [Lud88]. -dérivations [Ler86]. -Dimensional [MS86d, Eli88, Gu87a, MP89b, Seg85b, Tan87, Ric86]. -dimensions [BCZ87]. -dissipative [Sch87c]. -distance [Fer88a]. -dual [San87]. -dynamical [Ara85]. -endomorphism [Bar85d]. -equivalent [Fer87]. -factors [Ral87d, Cho85a]. -fold [BW85c]. -folds [Pet86]. -Folgen [Pro87]. -Fraction [Coo89, AK87, JW89]. -fractions [Ran86]. -free [Nat89b, Con89]. -freeness [She86i, Phi87b, Phi87f]. -function [Jin88a]. -Functions [GPSR87, Ral87c, And87a, CC85a, GPS87a, HS85, Mor86f, PSR87, Sun86]. -funktion [Sch85e]. -Gleichverteilung [BDT87b]. -gonal [Sch89g]. -graded [AR87, Lew88b]. -graph [Fin88c]. -group [Gil87]. -groupoids [Cho89a]. -groups [Hej88, CS85a]. -handlebodies [KB86]. -Heights [Mur85a]. -Index [Mül87h, Mül88, Mül87a]. -infinitely [GH85a]. -injectives [LS87b]. -invariant [Shi86l]. -invariants [Lev88]. -kernels [GLM85]. -knots [Kan89]. -Komponente [Sto85a, Sto85e]. -konkavem [Kos86a]. -level [Düm85a]. -Lévy [Mad89]. -like [Har87b]. -local [Hua86]. -logic [Gri85a]. -manifold [Bau88a, ST88]. -Manifolds [Don85a, Kir89c, Lev85c, Sch85b, Gab85, HLM85b, KL88, Kam89b, Kir89e, Kir89l, Kos88b, Mil85c, Tur88b, Žub88, Kam89a]. -maps [Dyl88]. -martingales [BW85a]. -matrix [Seg85b]. -measure [Wt87d, Wt87a, Wt87b]. -modules [Han87g, Lüc89a]. -Moduli [Dec88]. -Neumann [BS87b, Ran87]. -normal [tC87]. -norms [Kif88a]. -nuclear [Ele87]. -PADF [Fin88a]. -paramètres [Wsc85a]. -part [Pro87]. -pinched [KS86e]. -polynomial [LM88a]. -power [Shi86v]. -products [GS89]. -property [Shi88b]. -rank [Gol86a]. -recursion [DM85c, Sla85a]. -Regularized [LM89b]. -semi-parfaits [CL85]. -semigroups [He89, Tim88]. -sequences [Pro87]. -sets [Liu88]. -signature [KK89b]. -singularities [jQxQ88]. -solutions [Wei87a]. -space [Wen85]. -spaces [Ass89, JM85, Mil87b, AA89, Jaj89a]. -sphere [Aso89, BG88b, FKV88]. -stable [GZ89, RW85, Ric87]. -strictly [KMM87]. -strongly [GN88]. -structures [PM89, Raw85]. -subgroups [Ebe86]. -summable [BB88b]. -surfaces [PS89a]. -symmetry [Pup88]. -tables [Coo86b]. -Teilfolgen [Pro87]. -Term [Sto85h]. -th [Bar86e, Fin88a]. -théorie [Gai88a]. -theories [D’O85, Kuh89]. -theory [McC86a, FT87a, Hah86, Heg88, Hod89, McC85b, Mur85b, PW85b, Phi87c, Phi87e, Vog85, Wal85a]. -toral [Not88]. -torsion [Cat85a]. -transform [FJ88]. -trees [Sch89d]. -twistor [Raw85]. -valued [Bré89b]. -variétés [Ala88]. -version [LC85a].

13th [Ste89b]. 14th [GPR87]. 16th [CQ85, Rou88]. 17th [SSS86]. 1956 [Hop89a]. 1982 [Ław85, LS86a]. 1982/83 [JY85]. 1982/84 [Ław85]. 1983 [ACSC85, AIJH85, AMSV85, Bre85, CAWXD85, Di 85, DH85b, GWW85, LDS86, RLQ85, Rol85]. 1983/1984 [LDS86]. 1983/84 [AY85]. 1984 [AvW85, ABS86, AH85a, All85a, AW86b, BDH^{+}85, BBT85, BDM^{+}85, Com85, CDG86, DGM86a, DGM86b, DFZ85, EMS85, FGHS85, Giu85, Gra85a, Hen86b, Hen86c, HSS85, KZ85b, KS85a, LRT85, RR85, Ros86a, SJ85, Smi85b, Tau86c, Tua86, Tur85]. 1984/85 [AY86]. 1985 [ABS87, Bal86b, BS86, CDV87, CKGP87, DT87, FP86, FO86, FHMM86, GPR87, GPS87b, GZ87b, Gra86b, GBB87, HT86, Han87j, Hey86, IH86, JP86, KPZ87, LL86, LP86, NFM86, Nor86a, SSS86, Thr86, VZ86, vO86]. 1986 [CGW87, Tal86a]. 1st [ABS86, Giu85, LP86, Tal86a].

2nd [ACSC85, ABS87, FP86, Gra85a, HT86, NFM86].

36ème [Mal85c]. 37ème [Mal86b]. 3rd [Bre85, Ros86a].

4th [All85a].

6th [Di 85].

82 [Hla85d]. 84 [CCCN85]. 85 [CCCN87]. 86 [Hla87b]. 8th [KZ85b].

9th [KPZ87].

A-posteriori [Ban86]. A-Priori [DZ85]. A. [ABB85, Par86c]. A.C.C. [EK88b]. Aarhus [Bal86b]. Abbildungen [Kos86a]. Abdelkader [Dem87a]. abelian [And86a, Asa86, Bou88d, CGC88, HS85, She86i, Sil89b, Asa89, Atz88, Bar89a, Bar87c, CT88, Goo85, IMW88, Jag85, Kat89a, Lau87b, LM85, Mas87, Möh87, Muk89, SC85, Sch88b, Sch89e, Tre87, Wu86, vdG89]. abéliennes [Gil89b]. abelschen [Fal85]. Abhyankar [tDP89]. Ableitungen [Wei86a]. Ablowitz [IN86j]. above [KS86c, LW88]. absence [Pic85, Xu87b]. absolue [Yor85a]. Absolute [Cha85a, Jan86a, KS86b, Shi85, Bru87a, Jur89, Ōsh87, Str89a]. absolutely [Lub85, Wei87i]. absorption [Arv85, vW85]. absorptive [Bar85d]. Abstract [Fra86, AT86, Bai89, KÜ88, Lab86a, Lun86a, Mil88a, Oll85d, Sch89b, She87c]. acceleration [JJW87, Sko87, VSP89]. account [Lew88a]. accuracy [Kor87c]. accurate [Spe86]. across [Bel89b]. Actes [FHMM86]. Action [Ale86, CAXD86a, BN85, Eck87, Ker86, LP88g, Lev89a, Ocn85g, Ocn85h, Ocn85i, Woo87b]. Actions [Bra86g, OD85a, Ocn85a, AG89, AP85b, Aso89, Ass85, AV85, Ass86a, Bea88b, BR88, Car86b, Cha88, FH89, Goo85, GM87b, Hat86, Hat89, Jaw86, JV85, Kam85, Kam89b, Kat86b, Lev89f, Lew88b, LV85, Mil87a, MS85, Ped85, Pop87b, Ulr88d]. acts [FK88, Lex88]. acyclic [Koź86]. AD [Kec88a, Jac88a, Kec88a]. Adams [Sto85g, Sto85h, Bru86a, CGMM87, Dav87d, DLS85, Sch87e, Sto85b]. Adams-Spektralsequenz [Sto85g, Sto85h, Sto85b]. adapted [CH85]. adaptive [Fuc86, Hah89, Mét86, Ban86]. ADD [Ste85c]. addenda [Jos87a]. Addendum [San86, dlP89]. Additional [Bra86f, MS86c, Kar89, Par86d, Pie89b]. Additive [FT87a, Los85b, Nat89a, All85b, All89, ESS85, EN87, FS88b]. additively [HvL88]. additives [Bia89]. adelic [Rei89a]. adequate [Fou88]. Adiabatic [CLS85]. adic [Dur87, Gou88c, KQ88]. adique [Vig89]. adjoint [BDT87a, Bri86, GG85, Wei87g, Wei87j, Wei87q, Wei87t, Wid85f, Wid85g]. adjointness [Ngu85]. adjunction [Som86]. admettent [Reu86a]. Admissibility [Sle87]. Admissible [MC89, Shi86b]. Advanced [PlT88a, Sol86]. aerodynamics [BPPS87]. AF [Phi87a]. Affine [Li89a, Shi86j, Shi87c, Shi87b, Tit86, BK88a, Bar85b, DM89a, Lor86, Mum88u, RS85a, Shi86c, Shi86g, Shi86m, Shi86n, Shi86t, Shi86u, Sim89, Smi86c, Wu86]. Affinsphären [Sch85k]. after [CAXD86h]. again [BR86, Kne89c]. Age [Bos86]. Aggregate [Kiw85a]. aggregation [Man86a]. Ahlfors [Hay88]. aided [Sho87]. aircraft [Jam85]. Albanese [Hai87a]. aléatoire [Geb86]. Aléatoires [Ibr85, Wsc85e, Alt89, Bou86a, Car86d, Fer89]. aleph [Ara88]. aleph-nought-continuous [Ara88]. Alexander [Lis87]. Algebra [Com85, AH86, BR86, Bor87a, Bou87c, Bru87a, Bry89a, Car86b, Cib86, CD87, Don85i, Don85j, Don85m, Dra85, Dra88, FFR88, Gla85, GvW89, HL88a, HR86, Hüb89b, Hud89, Jaj88, Jaj89b, Jan87, Lam88, Lud88, Lus87, Mc 87, Mer86, Mic89b, MvR88, Mum88v, Que87a, Qui85a, Rei89e, Rei89j, Ric87, Roo88, Ros85a, Ser89, She86o, Sin87, Sta88b, Str89a, Ton87b, Van86a, Var89, VP88, Wil89, Roo86b]. Algebraic [BS89a, BV89, Bui86c, Don85n, Gra86b, Hah86, Hud88, Kow85, Mor86a, Net88a, Roo86b, Sch88c, Sil89a, Wal85a, AM88b, AM88e, AM88g, BK88a, Beĭ87a, BS89b, BS89d, Cat88, CGC88, Che86c, D’A85a, Dic85, DR86, Dub88, Ger89, GM87b, HRL89d, HRL89e, HRL89h, HH85, Kol86, Mor88, MS86u, Mor86b, Mum88q, Mur85b, OP89, Pat87, PW85b, Ran85, Sch88e, Sch88g, Shi86c, Sil89c, Sil89d, Spr86, SY89, Swe85, Van86a, Vog85, Web88, tW87a, vdE87, vW86, CAWXD85, RLQ85, Smi85b]. Algebraically [LP87b, Lüc89b, Lud88]. Algebras [AMSV85, BV88a, Die86c, DGM86a, DGM86b, Dür86c, Eva85a, Jar85e, Wan87g, vEdG85a, Abe87b, AP85a, AC85, AL86, AB85, Bae86, Bea86a, BMS87, BGO^{+}89, Ber86i, Bou85a, BDE85, Bra86g, BP88d, Cec88, Cec89, Ced88, Con86, DH86, DV86, Dic88, Die86a, Die86b, Die86d, DR88a, Dop85, DLS86, Dür86b, EK88a, Eld89, Est86, Eva85b, FT87b, FdlP86, Fis86, FR85, FGL86, Gei86, GL87, Gla89a, Goo85, Gro86b, Gro86c, Gro86d, Gro86e, HO85, Hap86, Hap89, Hen89a, Hen89b, HS87d, Iar85, Jaj85a, Jaj85b, Jaj85e, Jar85c, Jar85d, Jež85, Jøn85, Kai87c, KN89, Kis85, Kos86b, KŠ88c, Kra88c, Kum85, Lan85a, LL85, Le 88a, Len86, Lor86, Łuc89f, Mad85, McN85, MMT87, Mer86, Mes85, Ocn85a, Ocn85k, Pas89, Pet89h, Phi87a, Phi87c, Phi87i, Qua88b, Reg88, Rei87b, Rei89e]. algebras [Row89, Sel86, Sel88c, Sel88e, Sel88d, Sel88h, Sel88i, Sel88j, Sel89a, SW89a, She86e, She86l, Shi86e, Smi86b, Smi87, TW86, Tac86, Vos86b, Wan87c, Zel88, dS87, dP86, dlH85]. Algèbre [Ser89, Fer86, Lev85a, Mal88]. Algèbres [AS89b, Boh88, Kai87a, Kai87c, AL88b, Car89a, Cau86a, Cle88, Cru87, Deh89, GSZ87, Reu86a]. algébrique [Gui88a, Spr85]. algébriques [Ler86]. Algebroid [Sto87b]. algorithm [Arb89, BSK87b, Cha86a, GMW87, LC85c, LC85b, LC85h, LC85a, Noc86, Shi86h]. Algorithmen [KPT87]. algorithmic [Che85, GKL89]. Algorithms [FdlP86, Ste86b, BT86b, BS89e, Gol86b, Jam86, Kan87, KM86b, Lal88, Mét86, MW86, Ste85a, Via89]. Alle [Los85a]. allocation [DM87a]. alloys [Eng86]. Almansi [Ieş87e]. Almost [AOP86, AR87, BL89b, BK88b, DP85, Gue86, Hat86, Jaj85a, KRT86, LB86, MS86e, MS86d, Rog87, AU86, Ber88c, BG88b, BKS87, CG86a, KS86e, Mis85, MB85, Ray88, Sch86k, She87b]. along [Dau88f, Lau87a, Lin86b, Sch89e]. alternating [LC85c]. alternative [Anc87, EK88b, Wei87r]. altruistic [Kim86]. amenability [JV85]. Amenable [Ger88a, Ocn85b, LV85, Ocn85a, Oll85a, Oll85e, Pat88]. American [Di 85]. among [Dav87a, Tho85g]. amorphous [Ker85]. Ampère [Che88b, Fuk86b, Tru86, Won87]. amplifier [IN86c]. amplitude [AC87, Bac89, Yaj85b]. amplitude- [Yaj85b]. analog [Ani86, Kag89a]. analogue [Die88a, GZ86b]. Analogues [KL87b, Sun87, VJ85]. Analyse [Bon89, KPT87, Yan87]. Analysis [AS85i, BGV86, BKK87, CDV87, DF86b, FGHS85, Gra86b, Hen86b, Hla85d, Hla87b, LP86, LMZ86j, NR85c, Poh89e, jQxQ88, Rao89, Sch86i, SS85b, Tur85, vdB87f, AC87, Bea88a, Bri86, Brü85, Bur86, Cha86c, Dau87, Dia88, Eke86b, Fef88, Frö86, GMSW86, Hel88e, HM88a, IN86e, JMN86, Lin86b, Mac88b, Maw88, MS86b, NR85a, NR85b, Oll85h, Pie88b, Pie89a, Poh89f, Pre88l, Rav85, Rei88d, Rei89c, Sch85c, Sol86, Tay86, Üst87, Wai89, Wak87, Zol86]. Analytic [AK87, Jos88, Kan86a, Mor86a, Mur88a, Wih85, AFP87, Bal86c, Bin89a, BR85c, CS88a, DGR87, Éca85, Fer85c, God87a, GR88, GR85k, Hög89, JTW89, Jar89, Jon89, KZ87, Kal86b, Lau87a, Lie87, Pau86, Paw85, Pou88, Riv88, Rod87b, Rod87c, Shi88d, Ton87b, Wal87, Wie87, vEKM89, Thr86]. Analytical [Ado87, Mor86b, Nov88e, AD85a, Cha86c]. Analyticity [vEdG85b, Bou87c, vEdG85a, vEdG85d]. analytioues [Ofm86a]. Analytiques [Zin89, Ble86, Cou89, Del85, Den86, Joy86, Ofm86b, Ofm86c, Ofm86e, Szp86, Var86a]. Analytischer [Hla85b, Pro85]. Anfangswertproblem [Vog87a]. anharmonic [IN86n]. Anisotropic [Ieş87a]. anneau [Adj87, Kra87b]. Anneaux [Rou87, Alf86, Cau88, CL85, Ler86, Les88, Val86a]. Année [Mal85c, Mal86b]. Années [LDS86]. annulateurs [Kra87b]. Anosov [Ghy88, Lon85]. antichains [Hig85]. anticipative [DZ89b]. Antisymétriques [Yan88a]. Antoine [Bor86]. Antwerp [vO86]. Anwendungen [Col86]. Apéry [PS89a]. appear [Tho86b]. Appendix [Bér86g, BV88b, Els85a, Fie88a, Her89, HS86, Kog86a, LeP89a, LC85e, LC85d, Nat85b, Nic89, Wei87c, Wei87d, Wei87b, Wei87e]. Application [Aga85a, Jeu85, Lan88a, Ara87, BH86b, Bri87, Deh89, El 87, Fan88, FT88b, Gon88, Gri85a, Grü88b, Gyö85, Hei86, MK88, MN88a, McC89, MP88, Mou88, NR85e, Pic88, Pic87, Rof85, Sil89a, iS88, Vog88, dBJ87]. Applications [AvW85, Aga85d, Bla89, BGV86, BDM^{+}85, Bro89, BMMS86, CC85a, CC85b, Clo85a, DT87, Dia88, Dic85, Dür86a, Fie88b, Gan85a, GPS87b, HK85, IH86, LP88a, Léa88, Mar87, MS86c, MS86s, Mur85b, Pav87a, Pav87c, Pue88, RR85, Rum89a, Shi86d, Str88b, TW86, Tău86b, Tau86c, VZ86, Win85b, ABC^{+}85, AM88a, Bér85a, Bér86b, Bou88c, BLR87, BJO89b, Brü87c, Cha86c, Che87, tC87, CC87b, CC88, CC89b, Col86, DT89, Don85h, DW87, FH89, FJ88, Fri88c, Fri89, Fun86, Gau85, GP87b, Gyö89, Had86, Hal87, Hil89, HPS87a, Hud86, Ill86, JY85, JL85, Kal86b, Kat88b, Kif88a, Kon85a, Kon85d, Kon85f, Kor87a, KP86l, Lan86, LR89b, Lev89k, tLqL87, Mac88b, Mad89, Man86b, May86b, Min87, Mod87, Pik86b, Sil88b, Rod87e, Shi86v, Sun86, Tie88, Üst87, Van87g, Var88a]. applications [Vie85, Vig86, qW88, tW87a, Yan87, Att86, Bar89b, OD85a]. applied [LV89, McF88]. Approach [Fis87a, Mil85d, Sch86i, AD85b, BR85b, BG89b, Bor88b, Bou88a, CQ85, DM87a, DR86, FG88, FS89a, Fle85b, Fuk86b, Fut88a, Gri86c, Hej85, Ker85, Lew86a, Lic85, LM88b, MC88, Mil85e, Sén86, Tro89a, Via89, Wüs87a, Wüs87b, Zam88a]. approaches [BRA86e, CM87a, Sta87d]. Approche [Nic87, AD85b, Ber85b]. approximants [AK87, Dra87d, Dra88, Kov87, McC89, Njå89a, Sta87d]. Approximate [BLR87, Bin86, Kod86, Roo89, Sai86]. Approximately [HR88a]. Approximation [BLL89, CB85, DZ89a, GPS87b, Gon88, HW88, LC85f, LV85, LS88c, LS88d, Oza88, Pic88, Rof85, Shi87d, Voj87a, Wsc85a, vdB87a, BG88a, Blu85a, BM89c, Bre88, Bre86c, Che85, Col86, Csö89, DP88, Dun88, GR88, Ill88, KN88a, KK89a, KL87b, Kop89, Kor87c, Kuc87, KS87, LS87c, LM89b, LS88n, Mak88, Mas88b, Mat86b, Mou85, Njå86, Pap88a, Rec87, Saf88, Sch86q, Sch86r, Sha87, Sho85, SW89c, SS85e, dB85]. Approximational [Yom88]. Approximations [Lew87, Voj87e, CC85a, CC85b, Dav85b, Dun87, Gau87, GRT88, JM89a, LR88a, Voj87d]. Approximationstheorie [Col86]. April [ACSC85, ABS87, BBT85, EMS85, vO86]. Arbeitstagung [HSS85]. arbitrary [BDG88, Dzi88, Sch86m, Wan86]. arbre [Akk88]. arc [HSW85]. Archimedean [Rum89e]. Arcs [Fit87a, GLM85]. area [ABL88, Bar85d, Cat85b]. Arguesian [Day85]. arising [Ben85b, Bri85b, Bum85, Eke88, GL87, Len86, Que87c, Rei85]. Arithmetic [Abe87c, COT89, Fal87, Mor86a, SY89, Abe87g, Abe87h, Bak87a, CC89b, Clo85a, Clo85b, Har87a, Har87b, Mor86b, PW85a, Rub89, SS89b]. arithmetical [Bou85b, Pix85]. arithmetically [CO89, MR89c]. Arithmétique [CTKS87]. Arithmetische [Fal85]. Arnold [FZ85]. Arnold-conjecture [FZ85]. arrêtées [Yor85b, Yor85d]. arrêtés [SY87]. arrays [Bin86]. article [Ati85a, Gal89]. Artin [Bae86, Cae86, Gei86, Mer86]. Artinian [FGL86]. Askey [Koo88, Mag88]. ASL [BV88p]. Aspects [ACSC85, FHMM86, Kes86, NS89b, Tor87g, BN88, Che85, Coh87a, D’A85a, Elw88a, Gli88, Hen85, Hop86, Lyc86, Rod87c, Seg85b, Sim85a, Sim85d, Tak88a, Win85c]. Aspherical [DH89]. Asplund [Phe89a, Phe89h, Phe89j]. assisted [CC87b]. Associated [ILV88, Mag88, And87a, Ant85, Ask89, Aur89, AV88b, BN85, CM88a, Cho87b, Coo86b, Fis87c, Fun86, GLM85, Gru86a, Hoy85, Hus89, Jos87a, Jos87b, LS87f, McC89, Mic89b, Pan85a, Que87d, Shi86n, Sie85, Wih85, vdB87c, vdB87b, vdBL88]. associative [AA89, Dic88]. associativity [IM89]. associées [Azn87, Dao89]. assumptions [Bin86, Rug89, Wei87k]. asymmetric [Ito89, Laz88]. Asymptotic [Bax86, BM85, Ens88, IN86b, IN86a, Jaj89a, LN87, Mar89c, Mei88, Mil88a, MS86f, Sai86, Sch86i, Sch86j, Sch86k, Sén86, Wid85b, vdB87c, vdB87b, vdB87f, Ark89, BG88a, Ens86, Hay88, Jen86, Kor89, NS89a, Rie87, SS89b, Van87f, IN86e]. Asymptotically [Hal87, Van87a, Kas88, RZ87]. Asymptotics [CM88a, CKS87b, FP85, Gor86, Gre86b, Gre86c, Gro86b, HS87e, Neu86a, Van87b, AO85, Ges86b, HL88b, IK88, IN86i, IN86j, IN86l, IN86k, LYZ89, LS87f, LS88o, LS88p, LS88l, Sch87l, Wat88, vdB87e]. asymptotique [Bia89, Mou88]. asymptotiques [BM87]. Asymptotische [Kir85b, Kir87b, KP87b]. Atiyah [Sil88a]. atlases [Hop89b]. atom [TL86]. Atomic [For88b, Kaw87, FG88, ST89b]. attached [Tit85]. attaining [Car85a]. attractor [Tak85]. Attractors [GZ87b, Hir85, CD85, EB88, GZ87a, Rös85c]. au-dessus [Dem86b]. Aubin [Fut88a]. auflösbarer [Sch87f]. Auftritt [Hla85b]. Aug [AMSV85, Gra85a, Han87j, Roo86b, SSS86, Tua86, Tur85, VZ86]. August [Di 85, DGM86a, DGM86b, DH85b, GZ87b, Ros86a]. Auslander [Bjö89a, Bre86a, Die87, Eks89, IT86]. Austria [LRT85]. Auswertung [KPS87]. automata [Gil88, Que87b, Que87h, Que87i]. Automated [McF88]. Automorphe [Wei86b]. Automorphic [GPSR87, Wei86b]. automorphism [BM89a, BDK87, GK87b, Ito89, Lau87b, Mei86b, SVS87]. Automorphismen [Mat87e]. Automorphismes [Ber85e, Cou89]. Automorphisms [Gro87b, Gup89a, Iva88, LP88b, Abe87a, Boy88, BBGM89, FM88, Hei87, LV85, Mat87e, Pra85, Qua88b, Wan89]. autonomous [BF88a, DG86a, Poh89a, Sev86e]. Autour [Bou87a]. aux [Azn87, BH86b, Bou86b, Sil88b, Vig86]. Average [Nov88a]. Averaging [AO88, MPS87, PPP87, RT89]. Aviemore [Thr86]. Axiom [Mum88r, Wat87]. Axiomatic [Lai88, Mas85, Wt87d]. axiomatizability [Cor85]. axiomatization [Pre85]. axon [Bri85b, Gri85b]. ayant [Let89]. Azéma [Éme89]. Azumaya [Dus88].

B [Sta87b]. B. [Mey88c]. B.K.W. [Hel88b]. Back [Går87a, Hel88a, Hel86a, Kir89d, Lüc89c, MS86g, PlT88b, Phe89b, Pik86a, Poh89b, Sim85b, Sio86a, Tim86a, Yaj85a, Zha89a]. background [Mor86c]. bad [Lec86a]. Baer [Esp88]. baffled [Fil85]. Baire [LMZ86a]. baker [Sta86b, AP88, Wüs87a, Wüs87b]. Balanced [Jaw86, BT86a]. Balayage [Han88, Wer87a]. ball [Bou85a, Bru87a, MP89b, MY87b, Tho87]. ball-algebras [Bou85a]. balls [CS87b, Gro87a, Led86a, Pin86]. Banach [BS86, BDH^{+}85, FO86, FHMM86, Ara87, Are86b, ADR85, Aro88a, BN85, BT88b, BP88c, Bou85a, Bou88b, Bur86, CS88c, DL86a, DR88a, DG86a, Ele87, Est86, FL86b, Gre86c, Gre86e, Gri86c, Hei86, Jar85c, Jar85e, Joh88b, KS85a, Lar88, LC85d, Lin88b, Łuc89f, NS86a, NS86b, NS86d, Nei88, Neu86a, Phe89c, Pis86, Pra88, Rea88, Rei88b, Ros85a, Rub86, Sch85a, Sch88a, SS85b, SS85e, SS85f, SS85d, ST89n, Yuk86]. Bar [BDM^{+}85]. Bar-le-Duc [BDM^{+}85]. Barcelona [CAWXD85, CKGP87]. Barge [CM88c]. Barlow [Dav87c]. barren [HMW85]. Base [LMZ86b]. based [Bak87a, Kes88b, Kus86a, Lev88, McN85, Obr85, Vir88, tW87a, dG85]. Basener [Ton87a]. bases [Bri85a, CS89k, EN87, Gol88, Sio86b]. Basic [AS85c, Fis87b, GNOY88, Gro86c, Kne89a, NS86a, NS86b, NS86c, Pet89b, Phi87b, SS85a, VJ85, Win85a, CS89a, DK85a, ST89c]. Basics [Imk88a]. basin [GNOY88]. basiques [DF86a]. basis [BL89b, CS89g, CS89j, Clo85a, LS87e]. Bäumen [Pro85]. Baxter [Mad89]. Bayesian [Pic89b]. Bazilevic [Aha87]. be [Aul85, Nag87]. beaded [Kre86a]. Beeson [Kob89]. beginning [Shi86p, Zwi89]. Behavior [Hin87, Bou89b, Daw86a, Gil88, GL85, LN87, Mar89c, Mei88, Mil88a, Ros86c, Vuo88a]. behaviour [Aki89, Bax86, BM85, Fac89, Fer88c, Lig85a, LMZ86c, Pas89, Sev86a, vdB87b]. Beijing [Tua86]. Beiträge [Now85]. Bell [Pet89c]. Bellman [Da 85, Rof85]. below [LS86b]. Beltrami [Dzi88, iS88]. belts [ABC^{+}85]. Bemerkung [Hla85a, Hla85c, Tic85]. Bemerkungen [Tic87, Tur87]. Bénard [GR85e]. Bennequin [Eli88]. Beppo [Oht88]. Berechnung [Sto85i]. Berezin [Ne’85]. Berezinian [RM87]. Bergman [Bel87c, Dwi87, Gra87, Lig85b]. Berlin [FGHS85]. Bernoulli [BR88, Fei89, MM85]. Bernstein [Bar87b, BG89c, KT85a, LS88e]. Berry [Pop87a, Sla85b]. Besov [AFP88, Boh88, CF88, Mur87]. Bessel [Cle88]. Besson [Bér86g]. Best [CHM89, Kov87, BJ86, Kik89, Saf88]. Bestimmung [Pro85]. Beta [Ask89]. better [Gau87]. better-than-uniform [Gau87]. Betti [Roo88, Sim87, vS87b]. between [AM88g, And87a, AH86, Beĭ87a, Blu86, Bra87, BH89d, CS89e, Cat85b, CS87b, DZ87, Dyl88, Fre89, GR85c, Hel88i, Ino89, Kan86a, Kan88b, Kir87a, LS87e, Mil88c, OR87a, OW87, Pic89b, PX87, Pro89, SP89, Ton87a, vEdG85b]. Beyond [BE87]. Bézout [Oka85]. Bi [IN85]. Bi-orthogonal [IN85]. bialgebras [Dür86e]. Bianchi [Ves85]. bias [Rie87, RZ87]. bias-robust [RZ87]. BiBoS [ABS86, ABS87]. BiBoS-Symposium [ABS86]. bicanoniques [Gan85a]. bicentralizer [Wri89a, Wri89b]. bidualisant [Lev85a]. Bielefeld [ABS86, ABS87]. Bifurcation [Ant85, Bar89d, JMN86, Kog86b, vdM85c, BT88b, Cat85b, Chi85, Fom88b, Gus87, Kuč86a, Pie89d, Pie89e, vdM85b]. Bifurcations [BBT85, Sad86, Bor88a, KH86, Mir85, SC85]. Bifurcationss [BW88]. biharmonic [Bir85, Ste85b]. Biholomorphic [GK87a, Bar87e]. Bijective [Que87b, Rog86]. Bilinear [AS85d, Dol86, TK88]. Billiards [KSLP86, GK88b, Str86d]. Bimeasures [Rao89]. bimodule [BN85]. bimodules [Hil85b]. binôme [Far88]. binaire [Las87]. binaires [Dix85]. Binären [Pro85]. binary [HS89a, Hel89, Joh89b, Kun87, Pet89a]. Binomial [Ste89a]. Biological [Tau86c]. biology [Ede86, Tău86b, Tău86a]. Birkhoff [BF88a, Reu86b]. Birth [Had86, Bos86]. bisectional [Mok86]. Bismut [Kus86a]. bivariate [Obr85, OR87a]. blade [Alt88]. Blaschke [Sat86]. Block [CS89a, Cas86, FS85b, HN88, Wil86a]. blocking [CS89b]. Blocks [Fei86b, Shi86i, Shi86v]. Blow [LS88a, Beb86, Har88b]. blow-up [Beb86]. blow-ups [Har88b]. Blumberg [LMZ86a]. BMO [Bou87a, Yor85d]. Bo [Dav87d, Mah87]. Bochner [Cle87, Dem86a, Hel86b, LT87, LC85e]. bodies [Bou87d, BL88, HN86, Ieş87d, KM87, MTJ87, vdM85h]. body [Bal86c, Ens85, Kai87c]. Bois [Hla85a, Hla85c]. Bologna [FO86]. Boltzmann [FA89, Fun86]. Bonn [HSS85]. bonne [Lev89a]. Boolean [Die86a, Die86b, Die86c, Die86d, Jež85, She86e, She86l, She86o]. bordism [BD89b]. Bordismusgruppen [Sto85c]. bords [Pol86]. Borel [LSR88]. boréliens [Wsc85d]. bornage [Sil88b]. bornés [AA86, JM86b, Par86c, Sam86]. Borromean [HLM88]. Borsuk [ČW87, DKL88, FH89, Nak89]. boson [Fan89, vdBLP88, vdBL88]. Bott [Mit87]. bottles [Bâr88c]. bottom [Hel88b]. boule [AAGV85]. bound [BB88a, Che87, HK89, lS87d, Sla85b]. boundaries [Bre86c, GNOY88, Maz86, McG88a, Sle87, Sto85l, Ton87a]. Boundary [Aki89, Bar87e, BB86, Bru87a, Fil85, Krá88a, Kuf86, Lie87, LMZ86c, MPST85, Maz86, Rei87a, Roz87, Vuo88a, Waa89, Aga86b, AL88a, Bel87b, Bem88, BBGM89, Caf89, ČZ86, Col86, CL89b, Dau88i, Dau88h, EM86, Erk87, Fab88, FB86, Fla87, Fri86, Fu88, Gol87a, Gon88, Gra87, Gre86a, Gru86a, Gru87, Grü88b, GF87, HW88, Her88c, HSW85, Kes88a, Kod86, Kra88b, Laz88, tLqL87, Li88, Lig85a, Mar88a, MP88, Nit86, PS87a, Paw85, Rac88, Rea86, Rod87d, Sch85c, Sch86c, Šed86, Shi88d, Sho85, ST88, SU87, Tim88, Ton87b, Tru86, Wit88a, Woe86, Wu88a, YM89, Zvy88]. boundary-value [Wit88a, Zvy88]. Bounded [CS89b, Wid85b, Alb85, Alb88, AV88a, AS85l, Ber88b, For88b, Fuk86a, Haa85, Har88a, Hud89, Kaz89, Möh87, Nag87, PW85a, Tim88, Vol88b, Vol88e, dSCV86]. Boundedly [AM88c, AM88b, AM88f]. Boundedness [AS85m, Mur88b, Wei87e]. Bounding [Kir89e]. boundness [CC87a]. Bounds [Agm85, AP86, BG88a, Bor85, CHM89, DGR87, GA86, HGYR89, Nov88a, Nov88b, Nov88c, iS88]. Bourgain [Sch87a]. bouts [Tho88]. box [CM87b]. box-scheme [CM87b]. bracket [Kra88c]. braid [Jia89]. Braids [HjJ89]. Branched [Sko87, BW85c, Eel85, KS87, Ril89, Sch89e]. Branching [Kog86d, Kog86c, Gor86, Göt86, Iva86, MRC89]. Brauer [Bea86a, Brz85, Mic86, Yui85]. breakdown [Pic85]. Breathers [Smi89]. Brelot [Cho89c]. Bremen [AW86b]. bridge [Bra87]. bridges [Kre86c]. brief [Bér86a, Bor86]. Brillouin [JMN86]. Brno [VZ86]. Brown [Whi89, Anc87]. Brownian [Dav87b, Bas87, CL89b, DT88b, Elw88b, Gar86a, Kuo86, Kuo87, Kuo88, Kuo89, Kus86a, Lew86a, Lin86b, MRR88, Pin86, Tan87, Tan88b]. Brownien [BLY87, Le 88b]. Brownienne [CMJ85]. Browniens [Jeu85, Bia89]. Brunk [Hei85]. Brunswick [RLQ85]. Bu [Mah87]. Buchsbaum [MR89c]. budworm [GS85b]. Buffalo [LS86a]. buildings [Coh86, Ron86, Ros86a]. Bulgaria [KPZ87]. Bundle [Kiw85b, Aur89, Dem88, Hul85, Rec85, SS86, Sch86d]. bundles [AG89, AHKHK89, AS85b, App88, BT87a, ES87b, HH86, HS86, Kar89, Koc88, LR89a, Mar86c, Nar87, PM89, Rao87]. Burgers [DL89]. Burnside [Kom89b, LMM86]. Bus [KM86b]. Busemann [Shi86w]. Busteni [AMSV85]. bvp [Tan88a].

C [Bau85, Phi87c, AA86]. C*PCP [Sch88j]. C.I.M.E [Bre85, FP86, Giu85, Gra85a, LP86, Ros86a, Tal86a]. C.R. [Mez86]. Cônes [Fd89]. Calcul [FHMM86, JY85, AY87, Bar86d, Lav87, Léa88, Mar88c, Mey86b]. calculability [Wt87e]. calculating [GA86, HN86]. Calculation [Ros85b, Lav87]. Calculations [CGMM87, BA89, Bra86a, Dav87d, Jam85, Jam86]. Calculs [Yan88a, Yan88b]. calculus [AP85a, Bah88, BL89a, Bry89b, Cla86, Fri85b, Gal85, Itō88, Kuo88, Kus86a, Mas88a, Nor86b, Nua86, Nua88, NZ89, Oco88, PlT88c, Par87a, Par86b, Par86d, Par88b, Pic88, Qua88a, Str86a, Swe89, Üst87, Üst88, Vir88, Yas89]. Calderón [CDMS89, Mur88b]. Canada [DGM86a, DGM86b, Rol85]. Cancellation [Cae88, Men88a, Azn88]. Canonical [Kun87, Lev85b, Aoy86, BV88d, Car86e, Coe87, Gra86a, Kon85j, Kra88c, Mak88, Mei86a, Rum89b]. canonique [Bar86d, Las87]. canoniques [Gan85a]. Cantor [Tak88b]. capacities [Kor87a, Mur88a]. capacity [Ada88, Jon89, Lon88, PV88, Rum89d, Rum89e, Rum89f, Tak88b, Vuo88c]. capillarity [Fin88b]. capillary [CS86]. capillo [Bol88]. capillo-viscoelastic [Bol88]. Cappell [Mil85a]. Caracas [Di 85]. caracterisation [Akk88, Del89, DS86b, Éme88]. caractéristique [Rou87]. caractéristiques [Aud87]. Carathéodory [JNT86, JNT87]. Cardinal [MN88b]. cardinals [Fra89]. Carleman [Ken89]. Carleson [Ber87c]. Carlo [Nov88c]. carré [Mey86c]. carrier [Gaj86]. Cartan [Blu86, Mas85, Rob86, tW87b, cY86]. Cartesian [Frö86]. cas [Mey89e, Sal89]. cascade [BFKP86]. case [Ark89, CW89, CMM89, Fac89, GR85a, GR85b, Gul86b, Hel88b, Kog86f, LeP89a, LS88f, Már89a, Obr89, Phi87d, Ral87f, Rei89b, Rei89a, Rum89e, Rum89f, Sch86o, Sch86p, Siu85a, Siu87, SS88d, Tim88, Wid85a, Wid85f, Wid85g, dBJ87]. Casson [Kir89f]. Castigliano [Har85b]. Categorical [Dür86b, KP86a, CP87, Dür86e, GM88, KP86b, Mak85, Maw88, She87b]. catégorie [Sch88k]. Categories [Die86a, Die86b, Die86c, Die86d, ART88, Bar88b, CGC88, Die88b, DLS86, FK88, Frö86, Hap87, HV88, Rog87, AS89b, LS86a]. Category [AM88d, Ani89, Beĭ87c, HL88a, HR86, HK87a, LM86e, May86c, Mon87, Mum88b]. catenarian [BF85]. catenary [Die88a]. Cauchy [ABB86, Bli86, CM87c, DS86b, DST89, Gon88, IN86i, IN86l, Jon89, KN88b, tLmC89, Lie88, Mat85, Miz88, Mur88c]. Cauchy-type [Gon88]. Cauchysche [Vog87a]. causal [Ceg89]. Cayley [Bro89, Bus88a]. Cecchini [Acc89]. Cech [Kee87, Wat87]. cell [Tău86b]. Cells [Shi86j, Shi86l, Shi86k, Shi86r, Shi86u]. censored [HKW89]. Center [CCCN85, CCCN87, Roo89]. centered [DL87]. Centers [Le 88a, BP88a, Jar89]. Central [AB89, LC85g, Ale85, Bin86, DV86, Dür85b, DG86b, FQ85, FTO87, GZ86a, GZ89, He89, Kon85h, Kon85i, Kru85, Led86a, Mas86a, Sch86h, Sze87, Yuk86, vW86]. central-server [Sze87]. Centraliseurs [Cau88]. Centralizers [Har87e]. centre [Reu86a]. Centro [Bre85, FP86, Giu85, Gra85a, LP86, Ros86a, Tal86a]. centroidal [PS89b]. cercle [Pic89a]. cerclés [Sam86]. Certain [EK88a, Shi86j, AEE85, AOR88, BC86c, Bel87a, Bra86c, CRT89, DvdB87, EN86, Gel86, GK87b, Gre86a, GGL87, Gup89a, He89, Hei87, Jur89, KS86a, Kus88, LW88, LM87, LS88c, LS88d, LN87, PPP87, Rau85, Rei89d, Rob86, Sar85, Sch89e, Sel86, SW86b, Sul86, SY89, UW88, Wu88b, Xin89, Zvy88, dSCV86, vS87b]. certaine [Str85b]. certaines [Bia89, DvdB87, GV87, ZM86]. certains [Adi87, Loe86, MVW87f, Ofm86e, Sel87, Yor85d, dC86]. Chaînes [Kra87b]. Chain [Bru88]. chains [Acc89]. chaleur [Ank88, Aro88b]. chamber [Mei86b, Tim86b]. champ [Lav85, Mey86c]. Change [LM86c, Hah89]. Changes [Går87b]. channels [Bar88a]. Chaos [HM88c, Mis85, Cas85a, Kah88, LM89a, Mey88e, Nua89, Str87b, Yan87]. chaotic [Bar85d, Cas85b, Cat85b]. chaotique [Mey89e]. Chapter [Sev86b, Sel88e, Wod87]. character [Bry89b, Fut88c, Fut88b, Fut88d, Fut88e, Fut88h, Sch87e]. Characterisation [BH89b, Bre86a, MRR88, Sch87b]. Characteristic [CCE86, Kif86, AG89, Alu88, Asa85, Bor87b, CAXD86c, Caz88, Che86d, Eck87, FT88a, Kam89a, Kar89, KS86c, LS86b, Lev89i, Mei87, Nak89, Smi86c, Swe85, Swe89, Vir88, Wu88a]. characteristical [SS88d]. characteristics [BBP88, Bor88b, Bor89, jQxQ88]. Characterization [Are86a, Are86b, Din85b, DST89, Gro86d, NS86d, BDK87, Brü87b, Fer88c, Hua86, IT86, Kow85, Mas85, OD85b, Obr85, PR89a, Pau86, PS89b, Plu85, DS89]. Characterizations [GK87b, BV89, Coh86, Kur87, Nev87]. characterized [Shi86l]. Characterizing [SW87]. characters [CAXD86p, CS85b, FT86b, Res89, Rey86b, Sch88c, Sch88e, Sch88g]. Charakteristische [KP87b]. Charleston [Com85]. chasing [Aga86b]. Chebyshev [Dun87, Kov87, VSP89]. chemical [Sta86a]. Chemins [GB86]. cherchant [Éme88]. Chern [Sch87e, Siu87, sT89a]. chi [NV89]. chi-square [NV89]. Chile [dPR86]. chirality [RD85]. chirotopes [BS89d]. choc [Mét89]. Choice [HPS87b]. Choosing [Kei87]. Choquet [Ada88, Kes88a, Kes88b, Kes88c, Ros88b, Win85a, Win85b, Win85c]. Chow [Bro89, Fab89, Gai88a, RX88]. Christoffel [GLM85, LN87]. Church [MO88]. circadian [Car85c]. Circle [Hat89, ÉY87, Kam89b, Wei87m]. circles [Fri88a, Wan89]. circuit [Pas86]. circular [JT86]. City [CCCN85, CCCN87]. class [BBP88, BW85c, BV88d, Coh85, Con86, Cor85, DR87, Di 86, Dud85, Fit87b, Fuk88a, FM88, GL87, GM85b, GQS85, Gre86f, Hei87, Hua88, Ill89, Iwa87, Jur85, KN89, Kaz89, LB86, Mil88b, Oli85, PS88a, Pat88, PS88b, jQxQ88, Rea88, RC85, Sal85b, Shi86a, Siu87, SW86b, ST89f, sT89a, Wan88, Wie86, Wu88a, Zha88, vS87b]. classe [Gér86a, Pic86, Str85b, Var86a]. Classes [AH85b, Aud87, Nyi89, AG89, ASH85, Asa85, BGO^{+}89, BP88b, Bor87b, Car89a, DH88, Fuk88b, Gel86, HLM85b, Kai87c, Kar89, Kaw86a, Ker86, Kuč85, LP87b, Ler85, Lev89i, Miz88, MVW87f, OR87a, She87c, She87d, Shi86r, Sie85, Spr86, sT89a, Wu87a, Zap88]. classic [Aga85b]. Classical [ACSC85, AA85, Gra86a, GRT88, Mor86c, Tru86, BC86f, BCSSC85, CB87, Coh87a, CBD87, Don85g, Fut88c, Hah86, Hej85, Hel88e, Hör87, KP86b, KK89a, tLmC89, NR85b, PSR87, Rog87, Tro89a, Yaj85b, Zam88a]. Classification [Bal87, Gan85b, Kir89g, Neh87a, She86a, She86b, She87c, Tur88a, Žub88, Bau85, Bue87, Éca85, Her87a, Kor88, Let89, Mes85, Mic86, Ocn85e, Rao89, Sil89b, SS85f, Tur88b, Fom88b]. Classifications [Neh87b]. Classifying [Lev85c, She86c, BD89b, DZ87, Heg88, JM89a]. classiques [Mar88c]. Classnumbers [DV86]. Clausthal [DH85b]. Clifford [BW85a, Deh89, LYZ89, Sel89a]. cliques [LN89]. Clones [ACN85]. Closed [Gro85, McC85a, Mum88c, Shi88a, ART88, CCS^{+}87, Els85b, Els85d, Frö86, FK87, Kat88a, Koz85, LP87b, NR85d, Pre88a, Res89, Sil88a, Sze87]. closed-loop [NR85d]. Closure [GR85k, Kos88c, Bec88, Han87c]. clouds [ABC^{+}85]. Cluster [Tho85a]. clustering [BE86, CG86b]. CM [Sch88f, Sch87j]. CM-representation [Sch87j]. Co [ABL88, Day85]. Co-area [ABL88]. co-ordinization [Day85]. coadjointe [Lic86]. coalgebras [Sch86h]. cobordism [AM88c, DM89b, Kaw89, Kir89q, Rav88]. cochain [AP85b]. Cocycles [Hel86c, AJL89, Hel86d, Mey88a]. codes [KS88a, KL87a, OD85b, Sha89]. Codimension [Ghy88, LT89, BW88, CB89, Ein88, Eld89, Fab89, IMW88, Mag86]. codimension-one [CB89]. codimension-two [Fab89]. codimensional [DD85]. codimensionels [Mal86a]. codimensions [Reg88]. coding [Kec88b, She86d, She86g]. coefficient [GvW89, LN88, Sch86o, Sch86p]. coefficients [AK87, Bih85, BH89c, Bre87, Brü85, CM88a, DK89, Går87f, Gar88b, Giu88, Ich87, IN86h, IN86m, Lie87, Lis87, LB86, Mur85c, Ngu85, Ste89a, Van87a, Wei87c, Wei87h, Dao89]. coextensions [DK88a]. Cogroups [BH89a]. Cohen [BEH87, Dow89, Haj87, Han87c, Her87b, Rei86]. Cohérence [Hir89b, Gla89c]. Coherent [Gla89a, Mum88d, Dau87, Gla89d, Lis87, Mum88t, Ros87b]. cohérents [Gai88a]. Cohomological [Sta88a, Dra87a, RS87]. cohomologie [Gui88a, Rab87, Sch88k]. cohomologies [CKS87a, Her87a]. cohomologiques [Var86a]. Cohomology [CP88, Le 86a, Ocn85c, Ass86a, BC88, Bor87a, CGC88, Dic88, Dur87, ES87a, Far89, FFR88, FT88a, GM85b, Gur86, Hap89, Har88c, Hat86, Hüb89b, Jan87, Kne89b, KI87, Lan88b, Lew88b, LAB88, Lis87, Pat87, Rob87, SZ87, Wei89a]. Coifman [Mey85a, Roc88]. coincide [Ced88]. Coincidences [Hus89, Bor89]. coincides [Il’85]. coins [Mou85]. coinvariants [Mùi85]. collapse [Fra89, IN86p]. collection [Cho89b, SW89b]. collectionwise [NP89a]. Collective [Pre87, Liv87]. College [AH85a, Ser89, Ser89]. colligations [DLdS87]. collocation [Rei87a]. Colloque [LL86]. Cologne [HT86]. Coloring [Glo87]. Columbia [KS85a]. combination [Bra86b]. combinatoire [Ber85b, LL86, Lon86, Str85a, LL86]. Combinatorial [Bak87a, Bau88b, Ber86i, BS89b, Bro86, Law89, LV86b, dSCV85, Bok88, Bre86a, DH85a, Dre85, Nat89a, Pri89, Vie85]. Combinatorics [Str86b, Kat86b, OD85a]. combustion [Emb87]. comme [Bou87a]. Commentary [Ati85a, MT87]. comments [CS89i, Mor86e, Par89, Üst88]. common [Hög89]. Commutants [Cau86a]. commutatifs [Cau86b, GG84, Gal89, Rou87]. commutation [FL87]. Commutative [Jaj85f, Okn88a, App88, Blo86, BH89b, BP88d, Car86b, Cec85, Con85a, Dia88, Dra85, Dra88, Dus88, Est86, GvW89, Haj87, Jaj89a, Jar85c, Küm88, Lev85a, Lon86, Phi87d, Roo88, Smi86b, VP88, vW86, CDV87]. commutativity [GT85a]. Commutator [Jen86, Mur88b]. commutators [D’A87b, Lon89]. commutes [CFdG87]. Commuting [Liv87, Mac87, Mat85, Sch85i]. Como [Bre85, Gra85a, LP86, Ros86a]. compacité [Str85c]. Compact [Abe87a, Abe87c, Dow89, Mok86, Ral87f, Sch87c, Abe87f, AV85, BN85, CMS89, Cle87, GH85a, GK87b, Hu87, Ill89, JM89a, JM85, Kan86a, Kan86b, Kaw89, Kes88a, KS86f, KI87, MN88d, Mil87a, Muk89, NR85f, Nyi89, Pat88, Pou88, SS87, Van87d]. compacta [Dra87a, Dra87b, Fer87]. compacte [Var86b]. compactes [Var86a]. Compactification [Don89a, Fuk86a, Kee87, Neč86, RD85].

Compactifications [Ike88]. compactifying [MS85]. Compactness [Bel87b, Urb86, DM86, Din85b, Sei87, TT88]. compacts [Ofm86e, Szp86]. Comparaison [Fer85a, Zah89, Sel87]. comparative [VSP89]. Comparison [Fin88b, Gra85b, CD87, SH86, Zah89]. compatible [Neh87d]. Compensation [dC85a, Imk88c]. complêts [Rou87]. Complement [She87e, Sco87, Zha89b]. complementary [NR85f]. Complemented [Als88, AC85, CS89k, MS89c]. complementing [EH88a]. Complements [BT87b, Bui86a, Gek86a, Sel88b, GM89, Oko86a]. complet [Bou87a]. Complete [BC86a, Bra86c, CAXD86b, Caz88, GP87a, HH86, Hul85, HS86, Ma89, Mum88e, Str86b, TK88, And88, BC86e, Bus88b, Die85a, Dwi85, Ebe87c, Hüb89f, Jin88b, Kuč85, MD89, Net88b, PS85, Pra85, Rog89, Shi86w, Sil88a, Xia89, Zhe87]. Completely [Dho85, Dwi85, FQ85, Haa85, Rum85, Rum86]. Completeness [Faj85, MN88c, Jen86]. complètes [Bea86b, Ber85e]. completion [Don89a]. Completions [DK85b]. Complex [Edg85, Gre87, Gru87, Mum88f, Rav88, Sch86d, ST89d, Won87, And87a, Bea88a, BD89b, Bla87, BD88, Coe87, Dup88, Dus88, Fuk86b, Gen87, Hat86, KP86c, Kat89a, KS86f, Kra87a, ŁR85, LL89, Lew88b, Mas86b, Mes85, OR87b, Oko86a, Ost85, RTVW88, Roc88, Sch88b, Sch86q, Siu85a, SS88c, SS88d, iS88, Wal85b, Zal88, Gra86b]. complex-valued [Ost85]. Complexe [Lev85a, Bak89, BM87, Ofm86e]. Complexes [Nor86a, Ass85, Ass86b, Col88b, Dao89, FZ87, HZ87b, HL86, Kne89c, Koź86, Mah87, Qui85b, Roo86a, Var86a, Vig86, Wei85b, dlP87b, dlP89]. complexite [Pol86]. complexity [Yam89, Yom88]. component [Csö89, LF89, LP88c]. Components [SV85, DM87b, Fre85, Iar85, Kag89a, Kle87, Ung86, Vos86a]. Comportement [Bia89]. composite [Ble85a, EP87, Nic89]. composition [BSK87b]. Compound [KS85c, vEdG85c]. Compressed [Iar85]. compressible [Jam86, Neu86c, Val86b, Bei88]. computable [OP89]. Computation [CAXD86c, Ebe87a, Eng87b, Wei87f, IW86, JJW86, Wt87d]. Computational [CC89b, Pet89d, Yam89, Lab86b]. Computations [Gal85, Lac86, AM88h, BPPS87, EL88]. Computer [CC87b, Ede86, Yas88, CC88, DT88b, Hof85, Mc 87, Sho87]. computer-aided [Sho87]. Computing [Kem87, RX88, BSK87b, HS89b, Rei87a]. concatenation [LT88]. Concavity [Oko86a]. concentrations [Ste85a]. Concepts [Nei88, Gla89b, Pet89b, Wt87e, Win85a]. Conceptual [Ber87b]. conceptualization [GM88]. concernant [Aub86, Dix85]. Concerning [Mat89a, Joh86, Koz85, Rei88b, Res89, Sch85a, Sch88a, Sch89c, Sta86b, Šve86]. concise [Tău86a]. Conclusions [Pie89b]. concomitants [LP87a]. Concrete [BB87]. condition [Abe87d, Abe87f, Bak85b, Bjö89a, Bou87c, CAXD86p, Eks89, Gol89, HJ85, Krá89, Led86a, Mai88, Ple87, Wil87]. Conditional [LeP89b, Pet88, Bin89a, Jaj85b, Num89, Plu85]. conditionally [Hol88]. conditionnelles [Bou87a]. Conditions [Alf86, Aul85, CAXD86d, AL88c, Bar89c, Ben85a, Cam86, CAXD86f, CAXD86k, CAXD86m, CAXD86l, CAXD86r, CAXD86s, D’A87a, FB86, Gia88a, Gol87a, GF87, HRL89g, Her88c, Kas88, Kuč86a, Nag87, Rip87, Sch86m, Tsc85, Tvr86]. conductivity [Car89c]. cone [Il’85]. Cones [Fd89]. Conference [All85a, AMSV85, BS86, BDH^{+}85, CDV87, CAWXD85, Com85, CGW87, DT87, DGM86a, DGM86b, DH85b, DFZ85, EMS85, FO86, FGHS85, GPR87, Gra86b, GWW85, HT86, Hey86, IH86, KS85a, RLQ85, RR85, Rol85, Roo86b, SJ85, Smi85b, Tau86c, VZ86, vO86]. Confidence [Oko85]. configuration [BC88]. Configurations [BM89b, Pie89g, Str86b, WKK86]. Conformal [Kam89b, Vuo88b, Fer88b, KP86l, Pap85]. conformosymplectiques [Cru85b]. Congruence [Jon88, KV88, Pix85, Stu87, Tsc85]. Congruences [Kom89b, Kom89a, AK88, DK88a, Lan88d]. conical [CD87]. Conics [CAXD86f, CAXD86e, CAXD86h, CAXD86b, CAXD86d, CAXD86t, CAXD86u, Kat88b]. conjecture [BE87, BH86a, Bou89a, BH87, CCS85, CC85b, Dem87b, FZ85, KL87b, Lan85b, May86d, MVW87e, Mül87f, Not88, Res89, Sat86, Sch85b, Sch87h, Sch89g, Ste87a, Str88a, Voj87b, Wen85, Sch89e]. conjectured [BH89d]. conjectures [Bre86b, Dix85, Hel87, Mic86, Sel88a, Sta86b]. Conjugacy [AH85b, Spr86, ASH85, HN88]. conjugaison [MVW87f]. conjugate [Bra86b, Noc86]. Conley [Cha85b]. Connected [Fre85, GM86, Ass89, Con86, DM89b, Dub85, FT88a, HN89, Kaw88, Mas89, Pup88, Rod87b, Siu85b, Wt87f, Žub88]. connectedness [Li86]. connection [GR85c, IN86o, IN86j, IN86l, IN86k, KV89, KMR89, Mic89b, Pro89]. Connections [AMSV85, Blu85c, FB86, Blu86, Dav87a, Mil88c]. connectives [LE85]. Connectivity [LMZ86d]. Connes [Phi87f, Wri89a]. conormal [Sch87l]. Consequences [MS86k, Voj87b, NS86e, Van87h]. conservation [Che88a, God87a, God87b, Hof87, Key87, KK89a, Roe87]. conservative [DiP87, Tad87]. Considerations [GMSW86, FS88b, Sel88f]. Consistency [Kob89, She89a]. consistent [Bar89c, Már89a]. constant [AO85, Bou89b, Che86a, Che89, DD85, Dav87b, Del86, DK88b, DW87, Elb85, Går87f, Gud87, Gut87, Ma89, Wu87a, dC89]. Constante [Cru85a]. constants [Cha85a, LC85i]. Constrained [Sta88b, Kiw85c, Kiw85f]. constraints [Gea89, Gro86a, Kog86h, Lic85, Shn86, Vin85]. construct [McN85]. constructed [Sel88j]. constructible [Kec88c]. constructibles [Hir89a]. Constructing [Gro87c, Lon85, Lec86a]. Construction [Aga85b, Aga85c, Daw86a, Fri85b, Fun86, Hel88b, MZ85, Mey89a, Sch85j, Sel88c, Sel88e, Sel88d, Sto85f, Wu87c, Bra86d, CS89d, CS89h, Kes88b, LN89, Lev89l, MR87, PW86, Ron86, Sho87, ZM86]. Constructions [Cru85b, GPSR87, Mei86b, Gla89g, Sto85k, Woo87a]. Constructive [GWW85, Hop86, tW87c, Brü87b, OD85b, Rob89, tW87a]. constructs [Her88b]. Contact [Arn88a, BV89, CH85, GPT89, Lic86, Str86c]. contained [CAXD86u]. containing [ADR85, Nat89b]. contamination [Saf88]. continua [Gul88a, Hus87]. Continuation [BM86, Ken89, KH86, Ros87a, Ros87b]. Continued [Lan86, Thr86, Thr89, AJBL89, AK87, BJ86, BA85, DGR87, Hov89, Jac86a, JT86, Jac86b, JJW87, JTW89, JNT87, KL86, Kuc87, KS87, McC89, Njå89a, NW89, Njå89b, Par87b, Sko87, SW86b, Waa86, Waa87, dBJ87]. continuées [Die85b]. continues [Mey88d, Pas87b, Yor85b, Yor85d]. continuité [BH86b, Yor85a]. Continuity [CM87a, Dol86, Tho85b, BB86, Bru87a, Jan86a, Jur89, KS86b, Ōsh87, PV88, Por85, Rei89b, Rei89a, Sch88a, Wat87, Wer87b]. Continuous [Arv85, Her88a, Ara88, ADR85, BD86b, Coh88a, EK88a, FR86, Ger88b, Hol89a, Kik89, Kum85, MRC89, Tim88, Wei87i, vEdG85a]. continus [Del88]. Continuum [LS86a, Nol86, SS89a, Wil86b]. contrôle [Hau85]. contracting [Abe87a]. Contraction [Jaj88]. contractions [Bou87a, Jaj89b, Wak87]. Contribution [CCS85, KZ87, Mül87m]. contributions [JTW89, LR88b, Kur87]. Control [Kog86b, NR85c, Bar85a, Ben85b, But88, CM89b, Col88e, EM86, Fla87, Fle85b, GESEZ85, GQS85, Hop87, Hop86, Kog86d, Kog86c, Kog86g, Kog86h, Lio85, Lio89, MR85b, MW86, Mos85, NR85d, NR85e, OD85a, Rof85, San86, Tvr86, Vin85]. controlled [AM88b, AM88c, AM88f, DM87a]. convection [BLL89, Fit87b]. conventions [Imk88d, Win85g]. Convergence [AS85e, Avr87, Bol88, Cav86, Dal86, Dra88, HRL89a, HRL89b, HRL89c, HN89, JJW87, Jaj85b, MN88d, Pic87, Wie87, AS85h, Att86, BKS87, Din85b, DR88b, Edg85, Els85a, FS85b, Fra86, GM87a, GA86, Göt86, Gri86a, Gro86g, Hoh87, Jac86a, JT86, Jac86b, Jaj85a, Kal85, KPA89, KM88, Kru85, LM86b, Mas86a, Mét87, Njå89b, Obr89, Ome89, Par87b, Pic88, SW86a, Sch86s, Sko87, SW86b, Sta87d, Sul86, Žen86, Att86]. Convergence-of-types [HN89]. Convergences [Pas89, Kru89, Mos88]. convergent [AS85l]. Convex [Koe88, Phe89c, AOR88, Bal88, Bou87d, Bou87e, Che88b, Dal86, EM88, Eke86b, GH85a, Gol86b, Jar89, Kes88a, Kiw85a, Kiw85c, KM87, Led86a, Lon88, Mas88a, MTJ87, Phe89g]. Convexe [Cho87a]. convexité [Loe86]. Convexity [Kaw85d, Ros88a, Ben85a, BB88c, Oko86a, RS85a]. Convoluteurs [Del88, Der88, Der89]. Convolution [BL89a, Muk89, BH89b, Haz89, McG88b, MS87b, Ost85]. Convolutors [BS87c, MS87a]. Conway [Van85]. coordinates [LeP89b, Wag87]. Coordinatization [Neh87c]. coordinatized [Lev85b]. copies [Aga86a]. core [Nag88, Wil86a]. corner [Blu85b]. corner-regions [Blu85b]. corners [Dob85]. corps [Cau86b, Cau88, Kai87c, Van86b, Vig89]. correction [Hem86, Spe86]. Correspondance [AAGV85, MVW87a, Vig89]. correspondence [Hej85, Voj87c]. Correspondences [Gob88, BB86]. corresponding [Kan88b, Kis89]. Corrigenda [Jos87a]. Cosmology [Ker87]. Cosserat [Ieş87d]. cost [Kog86g, MR85b]. cosurfaces [AHKH86]. cosymplectic [CG86a]. Cotype [MS86t, Xu87a, Mil87b]. Coulomb [GR87a]. Countable [Har87b, Pre88a, Dow89, Oll85e, Sch87j, She86j]. counterexample [Jin88b, Wen85]. Counting [Fri88a, Ges86a, PW85a, Yam89]. coupled [Cha86a, KM86b, San86]. coupling [AO85, AFL89]. couplings [Cec89]. Courants [Pas87a]. courbe [GS87]. courbes [Gan85g, Hir87, Hir89c]. courbure [Aub86, Bak87b]. cours [Ser89]. course [Coh87a]. courses [Ser89]. covariance [EK88a]. Covariant [Car89c, Isr89, AHKHK89]. covered [MY87b]. Covering [Dvi85, HJ87a, Kar85, ACM85, Ceg89, Fre85, KM87]. coverings [BW85c, Caz88, CP87, DLS86, Far87b, Har87a]. Covers [Pas86, ABC^{+}85, HN88, Ril89, Sch89e]. Cowling [Mey85a]. Coxeter [Shi86e]. CR [Bej85, Chr89, Dwi87]. CR-submanifolds [Bej85]. cracks [Kre85, Sch85f]. Cramér [Kag89a, Ros87a, Šig87]. cranks [Mur88a]. Critèare [Sel87]. critères [Str85c]. criteria [BS89a, HSG88, HPS87a, Kon85a, Mos88, Sle87, Wei87m]. criterion [AD85a, Gar88b, Koe88, Meh85, OP89]. Critical [Bah88, CG86b, Maw86, Nel85, Shi86w, vdB87b, BN88, CW89, Fom88b, GPT89, PlT88a, PlT88d, Sjö87, qW88]. critiques [MK88]. Crossed [CW85]. crossing [qW89]. crossratio [Thr89]. crystals [ABL88, HKL88]. cube [Gil87, ST89n]. cubes [Bal89]. cubic [Pet89a, Gui88b]. cubics [Alu88, EPS87, KSX87, KS88b, MD89, Pie85]. cubique [Gai88a, Gui88a]. Cubiques [CTKS87, GP87b, Gui88b, MPS86, Pue88]. Cugat [CKGP87]. currents [Tay88]. Curvature [Asa85, GPT89, MR86, Bal85, Bal86a, BC86a, BV89, Che86a, Che89, DD85, Die88a, Drd85, Ebe85, Ebe86, Gul88b, Gut87, Kas88, Ma89, Mok86, Nis86, Pan86, Sch89f, Wu87a, dC89, SSS86]. curvatures [CP89b, Che89, Fuk86a, lS87d]. Curve [Len86, AHKH86, Cil87, Els85b, Els85d, LP88g, Sab85, Sch87b, Tro88, Web88]. curved [BG88b, RVV85, Xia87]. Curves [Bar89a, Gek86b, Smi87, BE87, BT87a, Cat85a, CO89, Chr87, Dub85, ES87a, Fab89, Fre89, GL87, Gek86g, GP87a, GR85k, Gud88, Har88c, HH85, Hul85, Kle87, Kor88, Lan88d, Lon88, Mar89b, MW89, MS89b, MR89c, Nar87, Pet86, Reg87, Rub89, Sab85, Sar85, Sch89g, Sil89i, SZ87, Str87a, VS87a, Wai89, Yui85, Zag85]. cusp [Mül87d, Mül87i]. cuspidal [MD89]. Cusps [Mül87b, Mül87h, Gek86e, Mül87g]. cut [EEK88]. cutsets [Con89]. CW [Ass85, HL86]. cycle [Sev86a]. Cycles [CAXD86g, Azn87, Beĭ87a, CAXD86q, Cib86, GMM88, Mur85b, Ofm86b, Ofm86c, Ofm86a, Ofm86e, Sev86c, Sil89a, Sze87, dP86]. Cyclic [BP88d, FT87b, Fri88b, Kar89, Mar86a, Sch89e, VP88, Bry87b, CRT89, Chr87, Fei86b, Ino89, JM86a, Ril89, Roo88, Wun87, tD89a]. cyclicity [Jos87a, Jos87b, Rou88]. Cylinder [Sio86b]. cylinders [Ieş87e]. cylindrical [CD87, Ele87, Sch89a]. cylindriques [Lum87, Lum89, NÜ89]. Czechoslovakia [VZ86].

d [Mey89a, BPPS87, Néd86, Spe86, vdW86]. D.C [GZ87b]. d’absolue [BH86b]. d’additivité [Hud86]. DAEs [Gea89]. d’Alembert [Kôn85k]. d’Algèbre [Mal85c, Mal86b]. d’algèbres [AL88b, Cru85b, Mal86a]. damped [Avr87, Eng87a, ENS87]. d’Analyse [LDS86]. d’annulation [Azn88, Dem87a]. dans [Alf86, AH89, Bak87b, Bak88, BCZ86, BH86b, Cau86b, Cau88, DZ89a, DvdB87, Die85b, Fer85a, FHMM86, Krá89, LT87, Loe86, Lum89, MVW87f, Mou85, Van86b, Wsc85d, Yp89]. d’applications [Vig86]. d’après [Mey85a, Mey88b]. Darmois [KZ85a]. d’arrêt [MM86]. Darstellungsformen [tDL85, tDL86]. data [BO85, BH89c, CAXD86m, FNT87, God87a, IN86m, IN86j, IN86k, Joh89a, Lie88, Lio86, Tho86a, Tit85]. databases [Plo88]. death [Bos86, Had86]. Decay [Don88c, Agm85, Don89b, Hel88c, Ski86]. Decker [Mac88a]. decomposability [Haz89]. decomposable [She86k]. Decomposition [KS88a, MP88, Sch86l, ST89e, BW85a, BV88m, CM88c, Dav87a, DB87, DZ89b, GP85, Haa85, He89, Kan87, Kre85, Ros87a, ST89b, Szé86]. decompositions [FG88, For88b, Sch87a, Wan86]. decreasing [Bre87, IN86j, MMT87, Ric86]. deep [CS86, ZG88]. Defect [Hem86, Fei86b, Shi86v, Spe86]. Deficiency [Wei87g]. Definable [PR85c, SS88a]. defined [Bou87f, CS88a, CQ85, Cho89a, Lüc89b, Lüc89e, Tak88b]. Definiert [Los85b]. defining [EPS87]. Definite [Bre86b, FP85, Hel89, Hol88]. Definition [Mum88g, Mum88h, Mum88k, dBJ87]. Definitions [Han87a, CD85, DK85a, Dud85, Poh89f, SS85a, Win85g]. definitizing [Ber88a]. defoliation [GS85b]. deform [Kan89]. deform-spun [Kan89]. Deformation [CP89b, IN86f, Knö87, KLR85, Reg87, Nis86]. Deformations [Ław85, Pou88, Ran89, IN86h, IN86g, MR86, Yan89]. Degenerate [Els85b, Che88b, Fu88, Hel88b, Kal85, Kru89, Kus86b, Li88, MR85b, Sjö87]. degenerating [Els85c]. degeneration [CAXD86i]. degenerations [Net88b, vdG89]. Degree [Kus88, Bro87, BT87c, Cam88, Dyl88, Kha88, Kor88, Kun87, LW89, Lüc88, PH89, SP89, Tro88]. Degrees [Kni87, AS85a, Bla85b, FS85a, Hom85, Odi85, Reu86b, SS89a, SS88a]. Dehn [Hej88]. d’Eilenberg [FT88b]. d’Einstein [Aub89]. Delay [Smí86a, DL86a, GL85, KN88b]. deletion [Kiw85g]. Deligne [GP87b, Tie88]. delooping [PW85b, Sch87e]. demiconvergence [FS85b]. Démonstration [Kôn85k, Deb89]. d’énergie [God87b]. Denjoy [Tho85c]. Denmark [Han87j]. dénombrables [GG84, Gal89]. dénombrement [Arq86]. dense [Lud88, ST89f]. d’ensembles [Wsc85d]. Densitè [Léa87, Bou87b]. densities [Yan88c]. Density [Kat88a, LMZ86e, Ark89, Bou87c, CMM89, Cie88, Cra86a, Düm85a, Gar88c, Joh86, Kir85a]. denting [RS85b]. dependence [BDD89, BD86b, Sch89c, Wih86]. dependent [DG86b, For85, Göt86, Jan86a, Lic85, SH86]. depending [Imk88a]. depth [LS87a]. d’équations [Lap88]. derivability [BH89c]. derivates [Tho85g]. Derivation [Wid85c]. Derivations [Bra86g, BV88c, Joh88a, Zin89, Bra86i, Mic89b, Ler86]. derivative [AL88c, DK88b, Sch89a]. derivatives [D’A87b, Hel87, Jan89, Oro89, Pet89g, Waa86]. derived [Beĭ87c, HR86, Mùi85, Thr89]. derivées [Bou86b, AS89b]. Descent [BP88c, Kiw85e, Bui86b, Sal85b]. Descente [Gai88a, Gui88a]. described [MS86u]. describing [HGYR89, vW85]. Description [HRL89d, Beb86, LV86a, Shi86g, Woe86]. descriptions [And87b, Shi86c]. descriptive [Gri85a, Kre86c]. désingularisation [Den86].

Desingularization [Tor87a]. Désintegration [Kah88]. dessus [Dem86b]. destruction [Sap86]. d’Été [AIJH85, Hen86c]. determinacy [LSR88]. determinant [Adj87]. determinantal [PW86]. determination [RS87, Tak85]. determine [GNOY88]. determined [BT87b, FKZ86, Ill89, Ost85]. deterministes [Hmi89]. Deterministic [Nov88b, AGHK^{+}86, Kir86, Rof85, Tir88]. detonation [Ros89a]. Deux [Hud86, Gan85g, Las86a]. developing [Ede86]. developmental [Ede86]. developments [Bér86a, Fef88, Hay88, Kno89, Sch85d, Tri86, Roo86b]. Développement [Mou88, Yan87, LM89a, Nua89]. Développements [BM87]. deviation [BCSSC85, Koi87, Kus86a, Lew88a, OR87a, Rie87]. deviations [BCSSC85, Fle85b, Jai85, Var88a, dA85b, vdBLP88, DD86]. devices [BCM86]. d’existence [Ras87]. d’extrema [Wsc85b]. Dga [Wt87c]. d’Halphen [Ell89]. d’Hermite [Str85a]. d’holonômte [FT86a]. d’hypergroupes [Gal86]. d’hypersurfaces [Bea86b]. diagonal [Coh87c, Dav85b, Fre88]. Diagonales [CTKS87]. Diagonals [Kum85]. diagonaux [GB86]. diagram [Pas86]. Diagrams [Ros86a, Ebe87a, Ebe87b, Far87a, KP86e]. Diameter [Mil87a, Bou85b]. Diameters [CHS85, Fuk86a]. dianalytic [Bâr88c]. dichotomy [KS89]. Dicriticalness [Can88]. dictionary [AH86]. d’idéaux [Kra87b]. dielectric [Del86]. diffeomorphic [FKV88]. Diffeomorphieklassifikation [Sto85p]. diffeomorphisms [Bax86, Éca85, Kir89q, PS88b, Sev86d, Sev86f, Sev86g]. Difference [Ler89, Bel89b, Bri85b, Dho85, Kir87a, MN88a, NR85a, NR85b, NR85d, NR85e, OW87, PR89b, Sch86c, Sho87, Tad87, VSP89]. differencing [Roe87]. different [CM87a, Sta87d]. Differentiability [Kal86a, Phe89e]. differentiable [Che86b, cH89a, Lev89b, Lev89j, MC88]. Differential [Bui86b, Bui86c, Bur89, CS85b, CH89b, DT87, Els85h, FGHS85, GZ87b, Grü88a, GBB87, Hüb89a, Hüb89h, Ker85, Kol86, Mul86, Pav87c, PM89, SJ85, Smi86c, Smi86b, VZ86, Wei87h, Wei87u, Wil89, AEE85, Ado87, Arn88a, Bai89, Bar88a, Bar86e, BDT87a, Bel87a, Bel89b, BD86a, BF88a, Ber87b, BG89b, Bih85, Böh86, BH89c, Bou87g, BDE85, Bri85b, BDG88, Bui86d, BHT89, CV87, CFM86, CQ85, CM89b, Chu88a, CC85b, CC89b, Col88e, Col88g, Con85a, Ehr85, Els85c, Els85e, Els85f, Els85g, Eng87b, Fen89, Fri89, Gar88c, GL85, Gre87, Gri86c, Gro87d, Gru87, Gud87, Gyö89, HRL89d, HRL89e, HRL89h, HL88a, Hel87, Hen86a, Hsi87, Hüb89e, Ich87, IN86h, IN86m, Iwa87, Jac88c, JSC87, KN88a, Kaz87, Koc86, Kre86b, Kuc86b, Kuo89, KJ86, Kus86b, Lab86a, LS88b]. differential [Lax87, LV86b, Lie87, Lit88, Lyc86, Mac87, Már89a, Maw86, MV88, Mic89b, MH88a, MH88b, MH88c, MS86u, Mur89, Mur87, Nag87, Net88a, Neu86b, Oco87, Olv87, Pav87a, Pic88, Pic87, Rei85, Roz87, Sch88h, Sch87l, Sev86e, Smi87, SS88c, SS88d, Šve86, Swe85, Tah86, Tal86b, Tre87, Wal86a, Wei87j, Wei87r, tW87a, Zha88, ZG88, vW85, CGW87, DH85b, FO86, GPR87, GBB87, Han87j, Hop89a, NFM86, Wei86g]. differential-algebraic [HRL89d, HRL89e, HRL89h, MS86u]. differential-difference [Bri85b]. Differentiale [Sto85g]. Differentialgleichungen [Vog87a].

Differentialoperatoren [Wei86g]. differentials [BV88j]. différentiel [Deb87]. différentielles [BH86b, Lap88]. differentiels [And87b, Lev89a, Cau88, Mal88]. Diffraction [Ros89a, Mei86a]. diffractive [Lau87a, Wu88a]. diffuse [Går87h]. diffusing [LS85b]. Diffusion [FP86, Kim86, UW88, ABGS85, AHK85, Bak89, BLL89, ČZ86, DD86, FB86, Fit87b, Föl88, Her86, Ich86, IK88, Kot89, Kuč86a, LM89a, LN88, Lum89, MPS87, Mar87, McG88b, Nic87, Nua89, Ōsh87, Par86a, Sal85c, Sch85j, Sta86a, Tan87, Tan88b]. Diffusions [BÉ85, Dür85b, Bak88, Car86a, CE88, DT88a, Elw88a, EH88b, Hud88, Hud89, Kuo88, MR85b, Nel85, Ros86b, Sul86, Üst88, ZM86]. digit [KPT85]. digraphs [Ito89]. Digression [Sil89b, Kir89c]. dilation [Bal86c, FA89, Pau86]. dilation-analytic [Bal86c, Pau86]. Dilations [BL85]. Diliberto [LC85h, LC85a]. Dimension [Gro88, Hes86, Mum88i, Bak85a, Bou87d, Bum85, Col88a, Deb89, Dra87a, Gei86, GNOY88, GGL87, Gup89b, Gur86, Kah88, Kör87b, Koź86, Kré88, Le 86c, Mañ88, Mil85d, Net88a, Okn88a, PV88, Sta87a, Ste89b, Szp86, Tak85, Tak88b, Van87i, Wal85b, tDP89]. Dimensional [DGM86a, MS86c, MS86d, MS86f, MS86i, MS86q, Tor87a, AJL89, Alb88, AK87, Ara87, Asa89, BW85b, BCSSC87, Bor88a, Bor89, Bou88b, BKT89, Brü87b, Can85, CRD87, CH85, CG86b, Dau88g, Die88a, Din85a, Dra87a, Eli88, Eng86, Eva85a, FFR88, GS85a, GL87, Ges86b, Gol89, Gu87a, HN86, Hap89, Har87d, Hid89, Hig87, IN86p, IN86q, Jen86, JLS87, KP87a, KP86e, Kos85, KŠ88c, Kuc87, Kuo88, tLcZ88, MS89a, Mes85, Mil85d, MP89b, Mir85, PlT88m, Pan85a, Pap88b, PS88b, Rei87b, Ric86, Sch87a, Seg85b, Sel86, Sel88i, Smi87, Tan87, Tan88b, Ton87b, VG88, Vos86b, Whi86, Zab87, dS87]. dimensionality [GZ87a, Gri86b]. dimensionally [OP89]. dimensionnelle [Ras87]. dimensions [BM88, BCZ87, Car88b, GR87a, HS89b, Kos88b, Lio89, Man85, McC85c, NR88, Sto87a]. dimodule [Bea86a]. Dini [Cau86a]. d’intégrales [Loe87]. d’intéraction [Cru87]. d’intersections [Bea86b]. Diophantine [CC85a, CC89b, Kop89, Voj87d, Voj87e]. Dirac [Ati85b, Bac89, BCSSC87, Duv87, Lav85, Mül87a, Wei87b, Wit88b, vEdG85e]. Direct [Bér86g, BFKP86, IN86b, IN86a]. directe [Sch85j]. directes [Hir89a]. directions [Kiw85f, Kra87a]. Directly [Bus88b]. directs [Zay87]. Direkten [Hla87a]. Dirichlet [Ber87d, BH86b, Bou87a, BH87, Fla87, FNT87, Kno85, Kop89, LMZ86h, Mou85, Sim89, SS88b, Zha88, dC88]. disc [Tro88]. Discerning [AP88]. disconjugate [Tre86]. discontinuities [Wag87]. discontinuous [Ros89c]. discoveries [Ber85d]. discrepancy [Mha87]. Discrete [Dau87, GM87a, NR85c, BN85, Bou87f, BRA86e, Coh88a, Die85a, Gyö85, Hir85, NR85d, Ocn85a, Vas86, dlH85, DZ89a, Mey89e, Rou87]. discrete-state [BRA86e]. discrete-time [Hir85]. discreteness [BM88]. discretization [AMM86, Pic87, Rek86]. discrets [GG84, Gal89, Yan88b]. discriminant [Pet89a]. Discussion [Fie88c, BC86d]. disjunctive [KS88a]. disk [LS87c, Pet89c]. Diskrepanz [Sch85h]. Diskrepanzbegriff [Fle85a]. diskrete [Wei86c]. Diskreten [Gol87c, KT85b, Tur87]. disks [Wal87]. dislocations [Koc86]. disordered [Ben86b]. displacement [Ste85a, lY87]. disque [Dem86b]. dissipation [Hof87, Tad87]. Dissipations [Bra86h, Bra86g]. dissipative [CLS85, Sch87c]. Distal [IMW88]. distance [Fer88a, LS87e, Rum89b, ST89n]. distances [Koi87]. Distinguished [Wan87d]. distortion [AVV88, Fer88b]. Distributed [Lio86, Fla87, KL87c, Mel89]. Distribution [Hel86d, Voj87e, Afa85, All89, BDG88, Dra87c, DST89, FJ88, Ger88c, GLR86, Has85, Hu88, Il’85, Kal85, Koi87, Kor87c, Plu85, Pom85, Ros87a, SW87, Siu86, Sto86, Tan87, Tur89, Van87h]. Distributions [Kon85g, San87, Ark89, Bin89a, Bos86, Cas85a, Dud85, Gaj86, Göt86, Jan86a, Kag89a, Kon85d, Kor89, Kré88, Kre86c, KT87, LR88a, Mar85, MY87a, Muk89, NS87, Nik87, NV89, NÜ89, Obr85, OR87a, Obr89, Oco87, OW87, SW87, Sch87g, ST89h, Yan87]. distributive [AC85, Dwi85]. divergence [Zha88]. diviseurs [Car89a]. divisibility [And87a, FQ85]. divisible [Csö89, DM89a, GH85a, Il’85, LeP89a, Ste89a]. divisibles [GG84, Gal89]. Division [BCZ86, Che86c, Kai87a, Kai87c, Le 88a, Van86a, Yge87]. divisor [BV88d, DH88]. Divisors [Wei87v, Kan85, Pir89, SV85]. do [Kaw89, Ste85c, vK87]. d’occupation [Bou87b]. dogleg [WKK86]. Dolbeault [Rod87a]. domain [Alb88, BF85, GR85b, Gen87, HSW85, Kan87, Kri85f, Mat85, Sad86, She86h, Vas86]. Domaine [Cho87a, Yp89]. Domaines [AA86, Lum87, Lum89, Sam86]. Domains [Wid85b, Blu85a, Bus88a, Che88b, Dau88g, Dau88i, Dau88h, Dob85, Eng89, Fab88, Fef88, FS87, For88b, GR85a, GR85b, GR85f, GR85g, GR85i, GP85, GK87a, GK87b, Har88a, Her88c, Hol89c, Lon88, Maz86, Mei88, MH88b]. domination [Ass89]. Donsker [Dud85]. dont [Bou86a]. Doob [BW85a]. Doolittle [KP86e]. d’opérateurs [JM86b, Par86c, Gér86a]. d’origine [Bou86b]. d’Ornstein [Jac85b]. Double [Per85, BR86, BBP88, Gra89, Sad86, WKK86]. Doubly [Ito89]. dozen [Sta86b]. Drinfeld [Gek86c, Gek86b, Gek86d]. Dual [Don86c, MW86, Dup88, ILV88, Ne’87, Res89, Sán85a, San87, Sim85d]. duales [MVW87c]. dualité [Hir89b, Ofm86d]. Duality [Bea88b, KP86d, KP86f, LÖ86, ST89g, Att86, BPSTJ89, Cam86, Des85, Går87b, HK89, HS87b, HV88, KP86i, LR89b, LM86d, Lis87, May86d]. duals [RS85b]. Dubinin [Bae87]. Dubins [Hei86]. Dubreil [Mal85c, Mal86b]. Duc [BDM^{+}85]. due [Vse85]. Duflo [Jos87a, Jos87b]. d’un [Akk88, AY87, AH89, Cho87a, Kra87b, Léa87, Mou88, Tho88]. Dundee [SJ85]. d’une [AY89, El 87, FT88b, Fer86, Gan85e, Geb86, GS87, Gil89b, GV87, Las87, LM89a, Lev89a, MK88, Nua89, Par86a, Sch85j, Str88a, Tho88, Var86b, Var86a, Yoe85, Yor85c]. Dunford [Dre86]. Dupin [CC89a]. Durch [Los85b]. Durchschnittlichen [Kir85b, Kir87b]. Dvoretzky [ABB85, Sch89c]. Dynamic [DFZ85, Sch88i, Vin85, Bar85a]. Dynamical [Lin88a, Oll85a, Oll85b, Que87d, Que87c, AHK85, Ara85, Col88a, Fel88, Hir85, Lic85, Oll85d, Que87e, Que87j, Str86d, BBT85]. dynamically [Tak88b]. Dynamics [Bre85, DK88b, DK88c, GR85h, Fri86, GP85, Hol89b, IN86c, KH86, MMA86, Shn86, Sho87, Wan87b]. Dynkin [Ebe87a, Ebe87b, Hap86]. Dyson [Sch86e].

E. [Die85b, Sta87b]. each [CFdG87, Koo85]. ebenen [Sch85k]. École [AIJH85, Hen86c]. economics [Eke88, KMR89]. economy [DM87a]. Edgar [Whi89]. edges [Dau88f]. effect [Jen86]. Effective [Wt87d, AS85b, Bar88b, Del86, MW89]. effectively [CK88]. effects [Ros89a]. efficiency [HGYR89, Pet89d]. Efficient [CRW89, Gol86b, BT86b, Cha86a]. égalité [Pas87a, Lin86a]. EHP [MT87]. Ehrhard [Bor86]. Eigenfunction [GG85]. eigenfunctions [Agm85, BDG88, Don88c, Don89b, Hel88c, Hin87, IN86n, Rei87a]. eigenvalue [Bar86a, Don86b, Gos88, Mut86, Sad86, Vol88c, Vol88b, Vol88d]. Eigenvalues [Ati85b, Que87e, Zal88, AB87, Bes88, FP85, GA86, Kir87a, Pav85, Rei87a, Ura86]. Eighth [SJ85]. Einbettungsprobleme [Mat87b]. Einem [Tic85]. Einheitskreis [Hla85b]. Einige [KPT87, Col86, Now85, Tic87, Tur87]. Einstein [Bry87a, Fut88g, KS86f, Sán85a, Siu87]. Eisenstein [BH89d, Kri85a, Mül87c, Ral87h, Wei86h]. Eisenstein- [Kri85a]. Eisensteinreihen [Wei86o, Wei86a, Wei86h, Wei86i, Wei86l, Wei86f]. elastic [AC87, Ieş87d, Kre85]. élasticité [Ras87]. Elasticity [GR85h, Bol88, EEŚ85, HSW85, Maz86]. elastodynamics [GR85i]. elastoplasticity [CL87]. elastostatics [GR85g]. Electric [Del86, Ben85c, Hel88h]. electro [MP88]. electro- [MP88]. electromagnetic [IN86c]. électromagnétique [Lav85]. Electrons [Ben85c]. Electrostatic [HvR88]. element [AMM86, BO85, BPPS87, CFdG87, ČZ86, Dob85, Els85d, FR86, GP85, MPST85, NW86, Néd86, Nit86, Sch85c, Shi86n, Shi86p, Ste85a, Tho86a, Whi86, Žen86]. élémentaire [Deb89]. elementare [Sto85i]. Elementary [PlT88d, Shi89, Cha85b, Deb89, Lew86a, PR85c, She87c]. Elemente [Sto85j]. Elements [Dzi88, CMS89, Kei87, Kří86, Sch85f, Sel88h, Shi86q, Tit87, Mey86a, Mey87a, Mey89b, Mey89c, Yan88a, Yan88b]. elimination [CLS85, Van86b]. elle [Lin86a]. ellipsoids [MP89b]. Elliptic [CC88, Els85c, Hoy89, Lan88b, Lan88c, Sch88d, BM86, Blu85a, Bra86d, Bri86, Dob85, Dur85a, Els85a, FM89, FP85, Fre88, Fre89, FK87, Gau87, Gos88, Gri86c, HW88, Hei88a, Hof85, Hoy87, Kus86b, Lan88a, Lan88d, MW89, Mei88, Ngu85, Nir88, Rhe89, Rub89, Sch85c, Sil89g, Sim86, Tan88a, VSP89, Vas86, Wat88, Whi86, Wil88, Yan88e, YM89, Zag88, Žen86, Zha88]. ellipticity [Tay88]. elliptiques [Dao89, Och88]. Embedded [Hof85]. Embedding [CS89c, DM89a, Hus87, MS86h, Sch87d, BC85, Dav87a, Haw87, JLS87, Mak88, Wil87]. Embeddings [FS85a, AS88, Bri85a, Can85, HR88b]. emergence [Sch86m]. emission [vW85]. empirical [DL86b, Dud85, GZ86a, Mas86a]. employing [vdW86]. Enchev [Mey88b]. Enclosing [Sch86c]. enclosure [Col86]. enclosure-solution [Col86]. Encounter [ACSC85]. end [PR85c]. endlicher [Sch87f]. endo [She86e]. endo-rigid [She86e]. endomorphic [CFdG87]. Endomorphism [FK88, Aoy86, Bar85d, GG86, Par88a]. endomorphisms [HS88, Jež85, Mir85]. ends [CD87, Xia89]. Énergie [Bou87b]. Energy [Cas85a, GR85a, GR85b, GR85h, PS87c, ABGS85, CK85, Fom88b, GR85g, GR85i, Rab88, Rof85, vdM85a]. energy-momentum [vdM85a]. engineering [Sta86a]. England [Tur85]. enhanced [NR88]. Enlacements [Rum86]. Eno [Har87d]. Enriques [Bar85c]. ensemble [AH89, Kon85j]. Ensembles [Cho87a, Wsc85e, Yp89, Ble86, Daw86b, Szp86]. enter [AAGV85, vK87]. entire [Fis87c, HS87b, Lub85, MMT87]. entrance [Tim88]. entrant [BW85b]. Entropic [BB85]. Entropie [Yor85c]. Entropy [KSLP86, Neč86, Oll85d, Oll85c, Bes85, BPSTJ89, KS86c, LS86b, Lin88a, Tad87, Yuk86]. enumerable [AS85a, Cho89b, Joc85]. enumerating [Pie85]. Enumeration [KL88, Ker86, dSCV86, Bre86b]. Enumerative [CAXD86h, KS88b, CAXD86a, LL86]. Enveloping [EN88, FT87b]. enveloppantes [AL88b, Mal86a]. environment [Gre86f, Tan88b]. Epi [Att86]. Epi-convergence [Att86]. epidemic [Kuu86]. epimorphism [LAB88]. Equadiff [VZ86]. equal [Mor86d, Reu86b]. equality [Gro86a]. equation [Alb85, Aro86, AD85b, BS87b, Bér85a, Bér86d, BC85, BCSSC87, Böh86, Bom87, BH89c, Bri85b, CV87, CW89, Chr89, Da 85, DL89, DG86a, For88a, FA89, Fun86, Går87e, HSW85, IN86i, IN86n, IN86o, IN86j, IN86l, IN86k, Kuo87, Lab86a, Lio89, MMA86, Mat85, McG88b, MRC89, Mül87d, PS88c, Pug86, Rei85, Rof85, Ros89c, Sch86j, Sch86k, Sch86l, Sch86m, Sch86o, Sch86p, Smi89, Vog87b, Vog89, Wei87r, Won87, Zam88a, ZG88, vW85, Aki89, Hau85, Lav85, Ama87, AD85b]. Equations [CAXD86i, DT87, Els85h, FO86, GZ87b, GBB87, IN86f, LRT85, Pav87c, SJ85, VZ86, AEE85, AT86, Ado87, Alb88, ABHK85, AOP86, Arn88a, Aro88a, Avr87, Axe86, Bai89, Bar88a, Bar86e, Bel87a, Bel89b, BD86a, Bem88, BLL89, BF88a, Ben89, Ber87b, Ber86j, BG89b, Bih85, BDD89, Bol88, Bra86c, BHT89, CM87b, Che88b, CL89a, CQ85, Cho87b, CM89b, CC85b, CC89b, Coh89, Col88e, Col88g, Coo89, CK88, DL86a, DG89, DF86b, Di 86, DZ85, Don86b, Dub85, Duv87, Ehr85, Els85a, Els85d, Emb87, EN86, Eng87a, ENS87, Eng89, Eng87b, FR86, Fen89, FB86, Fil85, FKZ86, Fri89, Gar88b, GZ87a, Gli86, Gol87a, GL85, GS86, Gri86c, Grü88a, Gud87, Gyö89, Hab87, HRL89h, HT87, HK87b, Hel87, Hem86, Hen86a, HSG88, Her86, Hör87]. equations [Hua88, Ich87, IK88, IN86h, IN86g, IN86e, IN86m, Jac88c, Jia89, Joh89a, Kač86, KPA89, KN88a, Kaz87, Kno85, Kre86b, Kuc86b, Kuo89, KJ86, Kus86b, Lan88a, LF89, LS88b, Lax87, LV86b, Li88, tLmC89, Lie87, Lio89, Lit88, LS88i, Lun86a, Lyc86, MPS87, Már89a, MN88a, Maw86, Maz86, Men88b, MV88, Mil88a, MH88a, MH88b, MH88c, MH88d, Miz88, MS86u, Mul86, NR85a, NR85b, NR85d, NR85e, Net88a, Neu86b, Neu86c, Nir88, Oco87, Oha86, Olv87, Pav87a, Pic88, Pic87, Pik86b, Pon88, Pov86, Rod87b, Roz87, Sán85a, Sch86c, SH86, Sev86e, Sim86, Spe86, SS88c, SS88d, Šve86, Tah86, Tal86b, Tre86, Tre87, Tru86, Val86b, VSP89, Wal86a, Wan87b, Wil88, Zha88, Bei88, BH86b, Bou86b, God87b, MS87b, Mey89a, Lum89, Mey89d]. equatorial [Car88b]. equilibria [Cav86]. equilibrium [Bra86a, DL88, Le 86c, LR88a, MC88, Pie89f, Sch88i, Sev86c, Sev86f, WKK86]. equimeasurable [Sch86a]. Equimodular [KL86]. equipped [ŁR85]. equitype [Wan86]. Equivalence [GS85a, Oll85e, CAXD86r, CAXD86s, DR89b, HS88, Le 89, Pic89b, Shi86a, Shi86r]. equivalences [Bou88d, Ino89, Šve86]. equivalent [BK88b, CK88, Fer87, Kec88a, Mer86, Tsc85]. équivalents [Yp89]. Equivariant [Lev89b, LM86d, LMS86c, LM86g, LM86f, MR89a, Phi87c, Smi85a, Ulr88a, qW88, And86b, Ass86b, BR89, BF88b, Dre85, FZ87, GZ86b, Ill86, Ill89, Ino89, Kaw86a, Kom89a, Lev89f, LMM86, LM86e, Lew88b, Lüc88, Mil87a, Phi87e, Ulr88b]. équivariants [Och88]. Erdos [KL87b, Tab86]. Ergodic [Ara85, AMSV85, Hur88, Jaj85c, Kan86b, Lev85i, Oll85f, Oll85g, Bou88a, EN88, Mar88b, Mod86]. Ergodicity [Don88b, Cho85a, EN88, Kan88a, Wal86c]. Erice [CDG86]. Erlang [Kal85]. Errata [LMZ86f]. Erratum [dlP89]. Error [Bre88, Nov88c, Tho86a, Axe86, Ban86, BG88a, HRL89f, Nov88a, Nov88b, Sch85c]. errors [CB85]. Erzeugendensystemen [Mat87d]. escape [Gor88]. espace [DZ89a, Lum89, Mar85, Mey86b, MY87a, NÜ89, Ofm86e, Sam86, San87, Wu87c, Wu87d]. espace-temps [Lum89]. Espaces [FHMM86, MVW87b, dC85b, Adi87, Ank88, BCZ86, BH86b, Deb87, DvdB87, Fer85a, Lic86, Loe86, Mar85, Pra88, Sch88k, Yor85d]. espèces [Joy86]. espérances [Bou87a]. essai [Fug86]. Esseen [Pop87a, Sla85b]. Essential [Wan89, Wei87i, dS87, iS88, Wei87a]. est [Bou86a, Lin86a]. est-elle [Lin86a]. estimate [Bre88, Che87, DZ85, KS86e, Oka85]. Estimates [BCSSC85, BG89c, BLM89, Kir87a, LC85i, Mur87, OR87a, Axe86, BCM88, Ban86, BB88a, Bli88, Gar88b, Gru86a, Jen87, Jut89, KV85, Ken89, Kus86b, MW89, Sch86j, Sch86k, Sch86l, Šig87, Tho86a]. Estimation [Mit86, Cie88, Hah89, HRL89f, Hoh87, KS86c, Kik89, KMM87, Koi87, LS86b, Már89a, Pic89b, Sze87]. Estimations [Ama87, DD86, Kod86, RV89]. estimator [RZ87]. estimators [Nik87]. Estivo [Bre85, FP86, Giu85, Gra85a, LP86, Ros86a, Tal86a]. eta [Gil89a]. Etale [LP87a]. étude [Jeu85, Lon86, Adi87, AY89, Bak85a, Bak85b, Bak87b, Dao89, Gil89b, Reu86a]. Euclidean [AHKHK89, BL89b, GM89, Hab87, Hoh87, KS86a, LN89, MS86e, MS86d, MS86q, Ulr88a]. Euler [BPPS87, Bra86c, CM87b, Eck87, FT88a, Hem86, Pon88, Spe86, Vir88]. Eulerian [Wag87]. European [HT86]. Evaluating [Mil85b]. evaluation [Bre86b, LM88a, Loe87]. évanescents [Azn87]. Even [Lub85, Gil89a, Kun87]. events [Alo89]. ever [Rei89f]. Everywhere [Yan88e, Kec85, Mis85]. Evolution [Gol87a, Pav87c, Aro88a, Axe86, Ber86j, BR85b, CL89a, Cho87b, DG89, EN86, Kač86, KPA89, Lun86a, Mil88a, Oha86, Pav87b, Pie88b, Pov86, Rek86]. evolutions [App85, Ens88]. evolving [Gre86f]. exact [CGC88, Mc 87, Mdz87]. example [CG86a, Don85h, Joh86, Mil85a, Sch88a, Sco87, Sel88g, Ton89]. Examples [BM88, CAXD86j, CAXD86k, DD85, GKL89, HRL89e, Kan89, Koz85, Pre88b, BC86e, BC86f, Bry87b, CFdG87, DH88, Eke86a, Kiw85i, LS87c, LMZ86l, Pav85, Sch88i, vEdG85d, vdB87e, Gal86]. Exceptional [Mur89, Gut88, LN89, Sel88c, Sel88d, Sel88i, Sel88j]. Exchange [GR85c]. Exchangeability [Ald85]. exchangeable [Gue86, Jan86a]. Excursions [McG88a, Pit87]. exemple [Dem86b, Hu88]. Exemples [Cau86b, Gal86, JY85]. Exercises [BS87a]. exhibiting [Wat88]. exist [Bin89a]. Existence [CV87, Car86a, Dun87, Eng87a, GZ87a, KS86d, Mas88a, She86e, Sta87c, Wsc85b, Wu88b, Alb85, Alb88, Ale85, Alo89, AU86, Bac89, Bar89c, BS89e, Che88c, Chu88a, Hof87, Koź86, LM89b, Mey86c, Mor89, NR85f, Neu86c, OP89, Sch89b, She86d]. exists [She86h]. Exit [Zab87, KM86a, Pin86]. Exotic [Kir89h]. expanding [Col88a, Kif88b]. expansion [Kus86a, Pro89, Vol88f, Vol88e, Vol88g, Wid85e, Wid85f, Wid85g, Wih85]. Expansions [Gek86e, Wid85b, AK87, GG85, JW89, NW89, Sch87l, Ste85b, Wid85h, vdB87c, vdB87b, vdB87g]. Expectation [Cho87b, Pet88]. expectations [Acc89, Jaj85b]. expert [GQS85]. Explicit [EL88, GPSR87, Ber87b, Cas86, LV86a]. Explizite [Sch87i]. Explosion [Ich86]. exponent [Böh86, Led86b, Wih86]. Exponential [Haj89, Agm85, Dür86d, HS87a, Jut89, Koi87, MP85, NV89, OR87a, Ski86]. exponentially [IN86j]. exponentielles [Let89]. Exponents [AW86b, Ani89, AKO86, AW86a, CCE86, Car86f, Cra86b, DL88, Eng86, Kif86, Kir86, Lac86, LS86b, Wih85]. exposant [Bou86a]. exposition [Har87c]. expository [Lew88a]. expression [IN86k]. expressions [GV87, LT88, Wei87j, vdB87c, vdB87b]. ext [BR86, Sto85i]. ext-algebra [BR86]. Ext-Gruppen [Sto85i]. Extendability [Ste87b, Rea86]. Extended [May86a, Din85b, Dud85, Han85, LMS86b, MS86b, Mül89, Rog85, Rug89]. Extendibility [Bel87c]. Extending [MRR88]. extensible [Len87]. Extension [LMZ86g, AS88, BCM88, Bou87f, CS88c, D’O85, DLdS87, Gur85, Har87d, HMW85, Siu86, Wie87]. Extensions [KN89, KV88, Pre88c, ST89h, BGS89, EK88a, EM85, FS85a, Gla89h, GR86, GK88a, Gro87b, KP86h, Kuč85, Lja88, McC85b, MPS86, PR85c, Ste85c, Vie85, Wei87g, Wei87q, Wei87t, Wu86, dS87]. extent [GS86]. exterior [GR85b, GR85f, Gru86a, HSW85, Mei88]. External [vdB87d, Ben85c]. Extraction [Ray88]. Extrema [Sie85, Sim85c]. Extremal [Kas86, AVV88, Fut88g, Gu87a, Kes88a, LS87f, LS88e, LS88k, LS88p, LS88l, LS88r, Oht88]. Extremals [Kog86b]. Extreme [Rac85, Csö89, Obr89, Ome89, Pan85b]. extreme-sum [Csö89].

F [tW87a]. facteurs [Zay87]. Factor [AB85, DM86, HN88]. Factoriality [Han87b]. Factoring [Nei85, WS87]. factorised [Sto87c]. Factorization [MN88a, Rob86]. Factorizations [Ger89, BHS88, HK87a]. factorized [Amb89]. factors [Cho85a, CHS85, Ral87d]. facts [Fis87b, Lin88b]. faible [Krá89]. Failure [Cai85]. faisceaux [Dre87, Gai88a]. Faithful [Bar87c]. false [EP87]. Families [MS86c, MS86m, Sev86c, Bel87b, tC87, DT89, FGL86, HH86, HPS87f, HS86, Kle89, Mas87, Neh87d, Spe88, Veg85, Ves87, Wal87]. famille [Gil89b]. familles [Bea86b, Let89]. family [BJ86, JW89, Rao87]. Fano [Con85b]. Far [GF87]. Fast [Los85a, vdW86, GMW87]. faster [Haj89]. fat [AP88]. Fatou [ABB85]. Feasible [Kiw85c, Kiw85f]. February [CGW87]. feedbacks [GQS85]. Fejér [Rei87a]. Feller [Kis89]. Fermi [GR86, GR89, PS89a]. fermion [RD85, Str86a]. feuille [Tho88]. feuilletage [Tho88]. feuilletages [Leh86]. few [Jac88c, Mil88c]. Feynman [AHKHK89, CS88a, HM88c, Lav87]. Fiber [Yag87]. fiberings [Hoy89]. Fibonacci [Men89]. Fibration [vdM85a]. fibrations [Gan85b, Gan85c, Gan85e, Wt87f, Yag87]. fibre [Dem86b, DR89b, Jaw86, Koc88, vS87b, Dem86b]. fibred [Mod87]. fibrées [Gan85g]. fibres [BM87, Mum88j, Adi87]. field [Ben85c, BGR85, Bli86, Bra87, D’A85a, Del86, Dop85, GS85a, GZ86b, GF87, KS85b, Sta88b, Tor87c, Yui85, dM88]. Fields [Bui86c, CDG86, Mat86d, Mor86a, Mum88k, Tor87a, AHKH86, BW88, BM89c, Bri85a, BM85, Cam88, Car88a, CCS^{+}87, Cho89a, Coh85, Coh88b, DV86, Dum85b, Fei89, Föl88, Gor86, Gos89, Got88, Hab87, Heg88, Hel88g, Hri87, LS87a, LN89, Mat86c, Mil88c, Mor86b, Nel88, Pan85a, Pre85, Ran88, Rei89f, Rei89g, Row89, SB88, Sew87, Sha87, ST88, SY89, Tău86b, Xu87b]. Figiel [CS89d]. filter [CMM89, Kak89]. filtered [AV88b, DB87, Eks89, LV89, McC85b, SZ87, vdE87]. filtering [CM87a, Coh88a, Mét86, Pic87, PS88c, Üst88]. filters [OP89]. filtration [AH89, CMJ85, Fu88, Jac85b, Mit87, Par86a, Sch88k, Yoe85, Yor85c]. Filtrations [Abe87b, DR89a, Don85n, Bou87a, JY85, Kni86, Pro89, Yor85a]. filtré [Adj87]. filtres [DR88b]. Filtrierung [Sto85j]. final [Ber87a, CS89i]. findable [CK88]. finding [BS89f]. Fine [Fug88a, LMZ86j, LMZ86i, Mey85b, LMZ86d, LMZ86h, LMZ86l, LMZ86m, LMZ86o]. Finely [LMZ86k, Fug88b]. finement [Adj87, Fug86]. Finetti [Fan88]. fini [Zay87]. finie [Pas87a]. Finite [Abe87c, Asc86, Aus85, BO85, ByC86, ČZ86, D’A87a, DGM86a, Dzi88, GP85, KP86e, Mad85, MS86c, MS86f, MS86i, MS86q, Rei87b, Sch87j, Vos86b, AMM86, AO85, Ade89, AV85, Ass86a, Ass89, BC86a, Bar89a, Bar89b, Bau88a, Ben86a, BPPS87, Bon85, Bou88b, BKT89, Bre86a, Bue86, Bus88b, CRW89, Cil87, CK88, Cur86, Dob85, DLS86, Dra87a, Els85d, FR86, Fis86, FS88a, GL87, GZ87a, Gul88b, Hap89, Har87e, Heg88, IT86, IMW88, Jag85, JLS87, Ker86, Kha89, KR85, Kos86b, KŠ88c, LM86a, LS87a, LP87b, LS88h, MO88, McN85, Mes85, Mic86, Mil85d, MY87b, Mum88s, NW86, Néd86, Nit86, Njå88, OP89, Pan85a, Pas86, Ran88, Rob86, Rog86, Rou88, Sch87a, Sch85c, Sch85f, She86k, SW89b, Smi87, Ste85a, SY89, Tho86a, Tho86b, Tim86b]. finite [Wan89, Whi86, Yaj85b, Yui85, Zal88, Žen86]. finite-dimensional [Hap89]. finite-dimensionality [GZ87a]. Finite-type [D’A87a]. finitely [AL86, Ber88a, CK88, Hem85, McN85]. finiteness [And86b, Mil85b]. finitude [Dem87a]. fins [de 89]. Fintushel [NZ85]. First [AV88a, AB87, BT88a, CAXD86n, CGC88, Går87d, GMM88, Gos88, Gur86, Had85, KV85, Kes86, LT88, Mut86, NV89, Pik86b, Siu87, Tre87, cY86]. first-kind [Pik86b]. first-order [LT88]. fit [NV89]. fitting [Zap88]. Five [Che85, CAXD86f]. fixe [Cou89]. Fixed [BGS89, FZ85, BG89b, Bor89, ČW87, Fad89, Fel88, Hög89, Kom89a, KS85c, Not88, Rad87, Ulr88b]. fixed-point [Not88]. fixes [Vig86]. Fixpunktmengen [tDL85, tDL86]. Flat [FGL86, BK88a, Can85, Rea86, Wai89, Xia89]. floating [Mal85b]. Flour [AIJH85, Hen86c, AIJH85]. flous [MM86]. flow [And86a, BC85, CHM89, DL88, Don88b, Jam85, Jam86, LF89, tLcZ88, Neč86, dLQ89, vdW86]. Flows [CE88, Bax86, BA89, BFKP86, CCE86, Car86f, Cha88, Fuc86, GR85d, Ghy88, Gus87, HN86, Har86, HR88a, IMW88, Kan88a, Kif88a, Le 86c, Nit86]. ≪produits [dC85a]. ^{≪}Supersymétriques^{≫} [Yan88a]. Fluid [Bre85, GR85h, Car89c, Fei86a, Fri86, GP85, Hol89b, Isr89, Neu86c]. fluids [Val86b]. Focal [PlT88e]. Fock [Le 88c, Mey86b, Mey87b, Par86d, Par88b, dC85b]. fold [BW85c]. folds [Pet86, SS86]. Folgen [Los85b, Los85a, Sch85h, Gol87b, Pro87]. foliated [Blu85c, Mil85c]. foliation [Can88]. Foliations [Gut88, Wan87g, Wan87e, AV88a, CB89, GMM88, Hur88, KT85a, KRT86, Mos89, NR85f, Net88a, Pou88, Wan87c]. foncteur [Zar85]. Foncteurs [Joy86]. fonction [El 87, Let89]. fonctionnelles [Arq86, Bia89]. fonctionnels [Bou86b]. Fonctions [Bar86c, Cle88, DF86a, Fug86, Roo86a, Adi87, AAGV85, Bad85, BD89a, BCZ86, Dao89, DR88b, Fer89, Hir89a, Le 88b, Nor86a]. fondamentaux [Bar87f]. forced [Car89b]. forcing [Bla89, MR87]. form [Che87, D’A87b, Don85a, FK87, Ger88c, KS86a, Mas85, McG88b, Mor86d, NR85a, NR85b, Yan88e, Zha88]. Formal [Lev89c, Bak87a, CC88, Kob89, Lau86, LP88b, LP88h, Meh85, OD85a, Poh89f, Rei85, Sch85i, Wie87, Yui88]. formalism [vEdG85e]. Formalité [FT88b]. Formalizations [Rau85]. Formally [Wei87j]. Formation [ABHK85]. Forme [Las87, Alt89, Bar86d, Bou87a, GV87]. Formel [Sch87i]. formels [Cau88]. Formen [Wei86b]. Formes [Cou89, Bar86c, Dix85, Mar85, Mar88c, Rab87]. Forms [Hüb89h, Kri85d, SW89a, And88, Asa85, Asa89, Blo85, BWW85, Bon85, Bro89, CS88b, DD85, Elw88b, Fal87, FNT87, Gek86h, Gou88c, Haa87, Hel89, Hüb89a, Hüb89e, Kag85, KZ87, Kir89m, Kri85c, Kun87, Kur87, Lan88b, LM86b, Mei87, Mic89b, Mod87, Mur85c, PlT88m, Pet89a, Rei89d, Rei89f, Sch88f, Sch88h, Sch85i, Sel88c, Sel88j, Sev86f, Stu87, TK88, Tir88, Wei86b, Wei86n, Wil89, Wüs87a, Wüs87b, Zag85, Zhe87, tD89a, tD89b, vdM85a, vdM85f]. Formula [Fis87a, AHKHK89, BG89d, CAXD86l, CAXD86n, CAXD86q, Cas86, Coh88a, DR87, DZ89b, Fis87d, Grü88b, Ill86, Jaj89a, KÜ88, Kuu88, LS88e, Mac88a, Mas89, Mei87, Par88b, Paw85, Ral87f, Ral87g, Wei85a, Yan88c]. formulae [IN86j, IN86l, IN86k, Lev85h, Lev89i, Ral87b, SS89b, Wat88]. formulas [Col88b, Dür86d, GLM85, Hes86, Hüb89e, Kat88b, MM85, Mul86, Pix85, Sai86, Van87f, Wid85i]. formulation [Ben89, BHT89, Cha86c, HSW85].

Formule [Far88]. Formules [Wsc85c, Yge87, Sil88b, Roo86a]. forte [Alt89, Pol86]. forts [dC88]. forward [AC87]. found [Ber85d]. Foundations [Mc 87, Rum89c, AM88d, CO86, HPS87e, Ran85]. Four [MC88, Win85c, vdB87e, Can85, Kir88, LW89]. Fourier [Bin89a, BB88b, Bre87, Bur86, CS88c, DB87, Dia88, Fef88, Fuj87, Koo85, Kuo87, LM86b, Lus87, Mur85c]. Fourier-analytic [Bin89a]. Fourth [DGM86a, DGM86b, Hen86b]. fractal [HS89b]. fraction [AJBL89, AK87, BA85, DGR87, JW89, Lan86, NW89, Thr89, Coo89]. Fractions [Die85b, Thr86, BJ86, Cau88, Hov89, Jac86a, JT86, Jac86b, JJW87, JTW89, JNT86, JNT87, KL86, Kuc87, KS87, McC89, Njå89a, Njå89b, Par87b, Pas87b, Ran86, Sko87, SW86b, Waa86, Waa87, dBJ87]. fracture [DDL85, MPST85]. frame [Li89b]. framed [Kir89k]. framework [Ler89]. France [ACSC85, Ser89, Ser89]. Fréchet [Fer85a, HS87b, Oha86]. Fredholm [Dau88a, PlT88j, Pik86b]. Free [AM88e, Bem88, Caf89, Car86b, Fou88, Fri86, Li88, Nit86, AP85b, App85, BT86a, Bou87d, Bre86c, Car88a, Cas86, Che88c, Coh87b, Con89, CK88, Fre85, Grü88b, Gup89a, Han87g, HLM85b, Kam89a, tLqL87, LT88, MR87, Möh87, Nat89b, Roo88, Ros89c, Sta87c, Wal87, Wit88a, Woe86]. Freedman [Hei86, Kir89i]. Freedom [MR87, Sén86]. freeness [Phi87b, Phi87f, She86i]. Freidlin [BCM88]. French [ACSC85, AA86, Adi87, Adj87, Akk88, Ala88, AIJH85, Ale86, AAGV85, Alf86, Alt89, Ama87, And87b, Ank88, Aro88b, Arq86, Asa89, AS89b, Aub86, Aub89, AL88b, Aud87, AY87, AY89, AH89, AD85b, Azn87, Azn88, Bad85, Bak85c, Bak85a, Bak85b, Bak87b, Bak88, Bak89, Bar86d, Bar86c, BM87, Bar87b, Bar87f, Bea86b, BD89a, Ben87, Ber85b, Ber85e, Ber87d, BLY87, Bia89, Big85, Ble86, Boh88, BCZ86, Bon89, Bou86a, BH86b, BL86, Bou87a, Bou87b, BJO89b, Bou86b, Bri87, Car89a, Car86d, Cau86a, Cau86b, Cau88, CMJ85, Cho87a, Cho89c, Cle88, CTKS87, CL85, Cou89, Cru85b, Cru85a, Cru87, DZ89a, Dao89, Deb89, Deb87, Deh89, Del88, Del85, Del89, DvdB87, Dem86b, Dem86a, Dem87b, Dem87a, Den86, Der88, Der89, DF86a, Die85b, Dix85, DD86]. French [Dre87, DR88b, DS86b, El 87, Ell89, Éme88, FL86a, Far88, FT86a, FT88b, Fer85a, Fer89, Fer86, Fd89, FL87, Fug86, Gai88a, Gai88b, GG84, Gal86, Gal89, Gan85a, Gan85b, Gan85c, Gan85d, Gan85e, Gan85f, Geb86, Gér86a, GS87, Gil89b, God87b, GSZ87, GV87, GB86, GP87b, Gui88a, Gui88b, Han89, Hau85, Hen86c, Hir87, Hir89b, Hir89a, Hir89c, Hmi87, HM88c, HM88b, Hu88, Hub86, Hud86, Ibr85, Jac85b, Jac85c, Jac85a, JM89b, Jeu85, JM86b, Joy86, Kah88, Kai87a, Kai87c, Kai87b, Kôn85k, Krá89, Kra87b, Kré88, Lap88, LS85a, Las86a, Las87, LT85, LT87, Lav85, Lav87, Le 88b, Le 87, Léa87, Léa88, LM89a, LDS86, Ler86, Les88, Let89, Lev85a, Lev89a, Lic86, Lim85, Loe86, Loe87, Lon86, Lum87, Lum89, Maa87, Maa89, Mag86, Mal85c, Mal86a]. French [Mal86b, Mal88, MK88, Mar85, Mar88c, MM86, MS87b, Mét89, MZ85, Mey85b, Mey85a, Mey86c, Mey86a, Mey86b, Mey87a, MY87a, Mey88b, Mey88e, Mey88d, Mey88c, Mey89e, Mey89a, Mey89b, Mey89c, Mey89d, Mez86, MPS86, MVW87f, MVW87a, MVW87b, MVW87c, MVW87e, MVW87d, Mou85, Mou88, Nic87, NÜ89, Nua89, Och88, Ofm86b, Ofm86c, Ofm86a, Ofm86d, Ofm86e, Oua88, Par86a, Pas87a, Pas87b, Pic86, Pic89a, PSS86, Pra88, Pue88, Rab87, Ras87, Sil88b, Reu86a, Roo86a, Rot89, Rou87, Sal89, Sam86, San87, Sch85j, Sch88k, Sel87, Ser89, SY87, Spr85, Str85a, Str85c, Str85b, Str88a, Szp86, Tho88, Tit86, Tor87g, Val86a, Van86b, Var86b, Var86a, Vig89, Vig86, Wsc85a, Wsc85b, Wsc85c, Wsc85d, Wu87c, Wu87d, Yan87, Yan88a, Yan88b, Yge87, Yp89, Yoe85, Yor85c, Yor85a, Yor85b, Yor85d, Zah89]. French [Zar85, Zay87, ZM86, Zin89, dC85a, dC85b, dC85c, dC86, dC88, de 89, Lin86a, dlP87b, dlP87a, dlP89]. frequencies [Coo86a]. Freud [KL87b]. FRG [DH85b]. Frobenius [Lor88, Sco87]. Front [Går87c, Hel88d, Hel86e, Jac88b, Kir89j, Lüc89d, Mic89a, MS86j, Mum88l, Mum88m, Mum88n, Mum88o, Nel89, PlT88f, Phe89d, Poh89c, Rei88c, Ros89b, Tro89b, Bri85b, Che88a, Går87k, Lau87a]. Frontière [Cho87a, Krá89, Sal89, Yp89]. fronts [Eli88, Går87h]. Fubinu [Pat87]. Fuchs [Tre87]. Fuchsian [Els85e, FS88a, Tah86]. fuchsiens [Ber85e]. fuel [Beb86]. Fulkerson [Joh89b]. full [Cho85a, Don86a]. Fuller [Fen89]. Fully [Nir88, Lio89, Lun86a]. Function [Bui86c, Fis87a, MN88e, AVV88, Ark89, Bel87c, BCZ87, BT86b, Boj88b, Bru89, CF88, DV86, Din85b, EB88, Fac89, FZ87, Fer88c, Fis87c, Fit87a, FF89, Gar88c, GM85a, Gos89, HS87a, HW88, Jin88a, LN87, Min87, Mül87l, Now87, Oka85, Oll85c, Oza88, Rod87c, Sch88d, Tho85h, Tri86, Wal88]. Functional [BKK87, Dub85, LRT85, BC85, BHT89, Col88g, Dra85, Fen89, GM86, Kno85, LE85, Mat85, Sch89f, Wu88b]. functionals [Bas87, CS88a, Cho87b, Giu88, HGYR89, KM88, Kuo86, Kuo87, Kuo89, Rac85]. Functions [Ber88b, Bre87, Dür86c, GPSR87, HPS87b, Ral87c, ST89i, Tho85f, vEdG85g, Aha87, All85b, All89, Anc88, And87a, AFP87, AFP88, ADR85, Att86, Bel89a, BS87c, BB88b, Bor85, Bru87a, BH89d, Bur89, CS88a, Cec85, Ceg87, tC87, CC85a, Cil87, CM87c, Cle87, Coh87c, DGR87, DP85, Fel88, Fer85b, Fer85c, FS87, Fug88b, Fuk88a, Gal85, Gek86h, Gek86j, GPS87a, GN88, GR88, Gon88, GR85j, HS89a, HS85, HS87b, Hel87, Hof85, Hol88, Ill88, IN86o, JJW86, Jan89, JW89, Jon89, Jos85a, Kal86a, Kal86b, Kas88, Kaw88, Kec85, Kiw85h, Kno89, Koe88, Kon85j, Kov87, Kri85e, Kuc87, KS87, Le 88a, Lew87, Li89a, Lon88, Lub85, LMZ86c, LMZ86k, Mac87, Man86b, Mat85, MN88b, MN88g, MMT87, MS87a, Mor86f, Mul86, Mül87c, Njå88, Oht88, Oro89, PV87, Phe89c]. functions [Phe89g, PSR87, Rea86, Rec87, Ric86, Rip87, Riv88, Rod87b, Ros87a, Rub86, Shi86w, Shi88d, SS89b, SS88a, Sta85, Ste85b, Ste87b, SW89c, ST89k, Sun86, sT89a, Tal86b, Tim88, Wal87, Wan88, Yan88c, Yom88, dG85, vdM85f]. functor [Mum88p]. Functors [KP86f, Dür86d, TW86, Tit85]. Fundamental [FKZ86, Gla89b, Sun88, Wei87k, Wt87e, AT86, BB87, Bau88a, Går87d, Gar88b, Gil87, Gol85, Hin85, Hol89c, IK88, Kri85f, Kus86b, Shi89, Xia87, Mat87c]. Fundamentalgruppen [Mat87c]. Fundamentals [Kiw85d, Wod87]. funktion [Sch85e]. Funktional [Vog87a]. Funktional-Differentialgleichungen [Vog87a].

Funktionaldifferentialgleichungen [KPT87]. Funktionalgleichungen [Hah85]. Funktionen [Drm87, Hla85b, Tic87, KPS87]. Further [JJW86, LS88g, LMZ86l, GL85]. future [Jam89, Kra87a].

G [Bér86g, CDG86]. Gâteaux [Phe89e]. Gabber [LV89]. GABs [Kan86c]. Galerkin [Els85a]. Gallardo [Gal89]. Galois [And87b, Bui86d, Des85, DLS86, Gou88a, Har87a, Kai87c, Len85, McC85d, Plo88, Sal85b, Tay85, Vig89]. Galois-quaternions [Vig89]. galoisiens [Kai87c]. Galoistheorie [Mat87a]. game [LT88, Zam88b]. games [Cav86, Ste88]. gamma [JJW86, Sch88d]. Gap [KS86a, HK89]. Garben [OS85]. Garsia [Bou89a]. gas [LF89, Sho87, Wan87b, vdBLP88, vdBL88]. gases [dLQ89]. gauche [Val86a]. gauches [Hir87, Hir89c, Van86b]. gauge [AHKH86, Bru87d, Gro86g, Hen85, Mc 87, Mod87]. Gauss [HK85, Riv88, Wan87a]. Gaussian [Fer85b, Jan86a, Kag89a, LM86b, Plu85, TC85]. gaussien [CMJ85, Bad85]. gaussiennes [Fer85a, Fer89]. Gebuhrer [Gal89]. Gel’fand [Aga86b]. genera [Lan88c]. General [Bra86i, Fie88d, HPS87a, LMZ86m, MS86d, Rug89, Sel88f, Sil89c, Bar89a, CK89, Cra86a, ESS85, Faj85, Fis87d, Går87b, Heg88, HJ87b, Kaw89, Kuu86, LL85, Ler89, MC88, Ngu85, Que87h, Que87i, Sch88i, Sch89g, Sze87, Tre86, Vie85, Wei87k]. générale [Alt89, AY87]. Généralisation [El 87, Sil88b]. generalisations [Pas87b]. généralisé [Jac85b, Yoe85]. generalised [dBJ87, Aki89]. généralisée [Far88]. généralisés [Str85a]. Generalities [Hel88e]. Generalization [CAXD86l, Ost85, Böh86, Bra88, Fut88b, Hol88, MP86, Phe89f, Pic89b, Ton87b]. generalizations [BA85, DST89, Rod87e, Zap88]. Generalized [Aba86, DG89, GT85a, Kan86c, Koc88, Pet89e, Sch87g, vEdG85g, Aga85c, Ass85, BLR87, CC87a, DP85, DR86, DR89b, EK88a, Had85, Kan85, KK89b, Kuo86, Kuo87, Kuo89, Kus86a, Maw88, MRC89, Mül87a, Par88b, Rod87b, Rod87c, RV89, Sch85f, Sch89e, Sel89a, She86c, Shi86l, Wil86a, Wro87, dG85]. generated [CK88, Hem85, Sjö87]. generates [Dum85b]. Generating [For86, Idà89, Ros89c]. Generation [Göt86, BM89a, Bar89a, EEK88, Sei89b]. Generation-dependent [Göt86]. Generators [AS85a, Kis89]. Generic [BV88e, BV88f, Cat85a, Fie88e, Fie88f, NR85g, Ste85c, Deb89, LP88c, Le 88a, MR87, Veg85]. Generically [BV88g]. Genericity [CAXD86m, Joc85, Fie88a]. génériques [Deb89, Hir87]. genetics [Don86c, Kim86]. Genre [GS87, Gan85e, Gan85g]. Genres [Och88]. genügen [Hah85]. genus [Bar89a, BBD88, Cil87, LW88, Sch89g, SV85, Zag88]. Geodesic [Don88b, DD85, Kan88a, NS86e, Pin86]. geodesics [Gen87, Gro85, McC85a]. géodésique [Loe86]. geodesiques [Vig86]. geodetic [Wit88a]. geographical [Nag86c]. geometri [Gro87d]. Geometric [Bér86b, BS89c, DH85b, Elw88a, GPR87, Mon87, Qui85a, RLQ85, Shi86f, AM88f, AM88g, Bér85a, BB88a, BS89a, BMMP88, Bou87g, CS85b, Die88a, Duv87, Grü88b, Ill86, Joh86, Jos88, KP87a, KP86l, MP86, Mil88d, Min87, NS86e, Nol86, Sim85c, SS85a, TT88, Tur88b, Wal85b]. Geometrical [Gli88, MS86t, Fom88b, Ker85, Lyc86, NR85f, Shi86g, FHMM86, Sto85k]. Geometrically [Lüc89e, FS88a]. géométrie [Aro88b, Cru85a, Dem86a, Pra88]. Geometries [SZ85, Wal85b, GI85, Li86, Pas86, tW87a]. Géométrique [Wsc85e, Bou86b]. géométriques [Léa88]. Geometrische [Sto85j, Sto85k]. Geometry [Bér86g, BKT89, CAWXD85, DH88, FGHS85, Gra86b, GBB87, Han87j, Hop89a, Lic85, MD89, NFM86, Ros86a, SS85b, Ara87, Bal88, BM88, BS87b, BD86a, Bér86a, Bér86h, Bis86, Bro86, BV89, Bue86, CC89a, Cha85b, Con85a, Daw86a, Dic85, DS89, Dub88, Går87e, Gre87, Haa87, Hsi87, Kaz87, KS88b, Koc86, KP86l, Lüc89a, Man85, Mur89, PlT88h, Pis86, RS85a, Sam85, She87a, Str88b, iS88, Sun86, Sun87, Swe85, VG88, Vuo88b, tW87a, AH85a, Ves85]. German [ACSC85, Bar87d, Bar85b, BDT87b, Buc85, Col86, Drm87, Fal85, Fle85a, Gol87b, Gol87c, Hah85, Has85, Hla85a, Hla85c, Hla85b, Hla87a, Kir85b, KT85b, KPT85, KPS87, Kir87b, KP87b, KPT87, Knö87, Kos86a, Los85b, Los85a, Mat87b, Mat87c, Mat87d, Mat87e, Mat87a, Now85, OS85, Pro85, Pro87, Sch87f, Sch85e, Sch85g, Sch85h, Sch87i, Sch85k, Sto85b, Sto85c, Sto85g, Sto85p, Sto85i, Sto85j, Sto85k, Sto85l, Sto85d, Sto85a, Sto85e, Sto85m, Sto85f, Sto85h, Sto85o, Tic85, Tic87, Tur87, Vog87a, Wei86b, Wei86a, Wei86g, Wei86h, Wei86i, Wei86d, Wei86c, Wei86j, Wei86e, Wei86l, Wei86f, Wei86m, Wei86n, tDL85, tDL86]. Germany [AvW85, ABS86, ABS87, EMS85, GWW85, RR85]. germs [Dum85b, Gag88, SB88]. Geronimus [ASV88]. geschlossener [Sto85p]. Gestalt [Kir85b, Kir87b]. Gevrey [Iwa87, Miz88]. Gevrey-hypoellipticity [Iwa87]. Gewisser [Kir85b, KPS87, Kir87b, Pro85, Sch85g]. Gihman [Sch87c]. Ginzburg [Fun88]. Giorgi [Wan88]. Girsanov [Jac85c, Pro89, Yoe85]. Gitterpunkte [Now85]. given [Bal86b, Bre85, FP86, Giu85, Gra85a, LS86a, LP86, Res89, Ros86a, Shi86v, Tal86a, Ves85, dPR86]. gives [DL87]. glace [Rot89]. glass [AH86]. Gleason [Łuc89f]. Gleichmässig [Drm87, Drm87]. Gleichverteilt [Los85a]. Gleichverteilte [Drm87, Los85a]. Gleichverteilter [Gol87b]. Gleichverteilung [BDT87b, Gol87c, Has85, Hla87a, KT85b, Los85b, Tur87]. Global [Aro88a, Bac89, BGV86, FGHS85, tLqL87, tLcZ88, tLmC89, Mor89, Owe86, Ral87a, Rum89d, Sta87a, Tak88a, Veg85, Aba86, DV86, EEŚ85, Fie88e, Fie88d, Går87d, GH85b, GLC86, Gu87a, Neu86b, Neu86c, SC85, FGHS85]. globale [Ras87]. globales [GV87]. globally [Hu87, Nag88]. glue [Beĭ87b]. Golitschek [LC85b]. Golod [Avr86]. gonal [Sch89g]. Gonality [MS89b]. good [Jac89]. goodness [NV89]. goodness-of-fit [NV89]. Gordon [Bac89, Sch86k, Smi89]. Gorenstein [Rog86, Wie86]. Goto [Aoy86]. Göttingen [Gra86b, Smi85b]. Gotzmann [Gre89]. governed [GESEZ85]. Graded [Bat87, Mic89b, Alm87, AR87, AV88b, Bry89a, Cae86, Cae88, Eks89, Lew88b, RM87]. gradient [Aba86, GZ86b, MW86, Noc86]. gradients [Bra86b, Veg85]. gradings [Bea88b]. Graduate [CCCN85, CCCN87]. graduées [Car89a, Cru85b, Cru87]. gradués [Alf86, Les88]. grande [Mag86]. grandes [DD86]. grands [Alt89]. graph [Fin88c, Kis85]. Graphenklassen [Kir87b, Kir85b]. graphics [Hof85]. graphs [Arb89, Bus88a, Con89, Ger88a, Mur88c, PS89b, She89a, Wan87h]. Grassmann [Str87a]. Grassmannian [Din85a, Hus89, Lev89c, Lev89d]. Grassmannians [lS89]. Grauert [Dem87b]. gravitation [Mc 87, Sew87]. gravity [Ne’87]. greatest [Fac89]. greedoids [GKL89]. Green [Dao89, Gar88c, JM86a, Kuu88, Lon88, Oza88, Sal89, Sch89g, Sel87, de 89]. Grenzefall [Wei86d]. grid [BM86, Fuc86, Hem86]. grids [Neh87a]. Griffiths [Bar89a]. Groningen [BBT85]. gros [Hir89b]. Grossissement [CMJ85, Jac85b, Jac85c, Par86a, Yor85a, Jeu85, Yoe85, Yor85c]. Grossissements [JY85]. Grothendieck [CC85b, Dav85a, HS87c]. ground [Gur85, TL86]. ground-state [TL86]. Group [Bra86g, FS88b, JV85, Koo88, AHKH86, AM88f, Bak87a, Bar89a, Bar87c, Bau88a, Bea86a, Bea88b, Bin86, Bin89a, BN85, BW85c, BDK87, BV88d, BH89d, CAXD86a, Ced88, DM89a, DM85a, Die85a, DR86, EK88a, Eck87, Fab89, FH89, FG88, FT87b, Fut88h, Gia87, Gil87, GvW89, GM85b, Goo85, Gur86, Haz89, HS87c, Hid89, HLM88, Hoh87, Hua86, Ill89, Joh89b, JM86a, Ker86, Kim86, Kos86b, Kri85f, LP87b, Lor88, Lud88, Mar86a, Mei86b, Mit87, Mor88, Oli85, Pan85a, Pri89, Rei89b, Rei89a, Rei89c, Rei89h, Reu86b, Rog88, SVS87, San85b, Shi86c, Shi86g, Shi86m, Shi86n, Shi86t, Shi86u, Tim86b, Tit85, Ulr88d, VG88, Wu86, Xia87, Yui85, Tua86]. group-valued [AHKH86]. Groupe [Car89a, Dre87, Bea86b, MPS86, MVW87d]. groupes [Ale86, And87b, Bak85a, Bak85b, Ber85e, Hmi89, Lum89, Maa89, MVW87f]. groupoids [Cho89a, Lja88]. Groups [Abe87c, ACM85, AH85b, CDV87, DGM86b, Don85n, Gal89, Hey86, JP86, Rhe89, Shi86j, Tit85, Abe87g, Abe87h, Abe87b, Ade89, Amb89, AM88e, AM88g, And87a, ASH85, Asc86, AV85, Ass86a, Atz88, BM89a, BSK87a, Ben86a, Ber86i, Bok88, Bru86b, Bue86, BBGM89, Buz88, CRT89, CRW89, CS85a, CFdG87, CGC88, Che86c, CC88, Cle87, Cur86, DM86, Don86a, Don85g, Dvi85, Ebe87e, Ebe87d, FS89a, FS89b, Fei86b, FKZ86, FS88a, FRN85, Frö85, GM87a, Ger88a, Gol85, GK87b, GLR86, Gup89a, Hah86, Hai87b, HJ89, Har87e, HN89, Heg88, Hei87, Hej85, Hej88, Hem85, Hol89c, Hsi87, cH89a, IMW88, Ino89, IKPR88, Ito89, Iva88, JM89a, Jag85, Jan86b, KP87a, Kan86b, Kar85, Kat89a, Kaw89, Kaw87, Kha89, Kol86, KI87, Kos86b, KŠ88c, Lau87b, LP87b, Len87, Lev85g]. groups [LM86c, LAB88, RX88, LV85, LM87, Lon89, MO88, Mar88b, MS86a, Mat87c, Men89, Mic86, Mil85b, Möh87, MY87b, Mur89, Neu89, Not88, Nov88e, OD85a, Ocn85a, Ocn85b, Oli85, Oll85a, Oll85e, PlT88l, Pat88, PSR87, Pir89, Pom87, Ral87h, Rei88d, Rei89h, Rob86, Rob87, Rob89, Ros89c, Ros87b, Sal85b, Seg85a, She86i, Shi86e, Shi86v, Sin87, Son85, Spr86, Sto87c, Sun88, Tho86b, Tit87, UW88, Wil87, cY86, Yui88, Zal88, Zap88, dlH85]. Growth [Ril89, Aha87, Bac86, Gia88a, Gor86, Kas88, Rip87, Ros86b]. Gruppen [Sch87f, Sto85i, Tic87]. Guanajuato [Hen86b]. Gudkov [KV88]. guide [HZ87a, Oco88].

H [Bau85, Whi89]. H. [AY87, Sta87b, Wen85]. Haagerup [Wri89b]. Haar [Szé86]. Hadamard [Aga85a, Aga85b, Aga85c, Aga85d]. hadron [Ne’87]. Hahn [FL86b, ILV88, Mag88]. Halasz [LR88b]. Halbgruppen [Los85b]. Half [Kri85d, Erk87, GR85a, Kec88a, LMS86b, Ost85, Sch86j, Sch86k, Sch86l, ST89i, Wei87c]. half-line [Ost85, Wei87c]. half-space [ST89i]. Half-Spaces [Kri85d]. Halphen [CAXD86h, CAXD86l, CAXD86n]. Hamilton [Eva87, Går87e, Rof85]. Hamiltonian [BGO^{+}89, Fom88a, GH85b, HS87e, Hol89b, Rab88, vdM85b, vdM85c, vdM85f]. Hamiltonians [Bal86c, Jen86, Lic85, Sta88b]. hamiltoniens [Lim85]. Handlebodies [Kir89k, Qui85b, KB86]. handles [Kir89f]. Hanf [She86f]. Hankel [CS88b]. Hardy [CF88, For88b, Kaw87, ST89j]. hare [Bel85a]. Harmonic [CDV87, Ess87, Hil85a, Jos88, Kas88, LMZ86n, Sal85a, Sam85, Wak87, Zhe87, Anc88, And88, Bur89, Bur87, CT88, Drd85, Elw88b, Ike88, Jen87, Jon89, Jos85b, Kal86b, Kaw88, Kes88b, Li89a, Mac88b, PS87a, Pie88b, Raw85, Rei88d, Rei89c, Ura86, Wai89, Woo87a, Xin87, Xin89, dG85, Giu85]. harmonically [Fil85]. harmonicité [Bak88]. harmonique [DZ89a]. harmonizable [Rao89]. Harnack [Wan88]. Hartogs [MS87b]. Hausdorff [Ada88, Ber88b, Bou88c, Bum85, Col88a, JS88, Kau87, Mañ88, Mum88r, PV88, Tak88b]. Hausdorffness [NP89a]. Haussdorf [Le 86c]. having [Coh88a, CK88, GMM88, Gul88a, Gut88, Joh89b, MO88]. hazard [Obr85]. heat [ABHK85, And86a, BS87b, Bér85a, Bér86d, Don86b, Kuo87, Mül87d, Mül87e, UW88, Wat88, Wid85e, Wid85h, Zam88a]. heavy [Str89b]. Hecke [Gou88b, HS85, Sch88c, Sch88e, Sch88g, Shi86e]. Heidelberg [AvW85, Tau86c]. Height [Beĭ87a, Tor86, dSCV86]. Heights [Mur85a, Voj87f, Mas87]. Heiles [RV89]. Heisenberg [DS86b]. Held [DH85b, AvW85, ACSC85, ABS86, ABS87, AH85a, All85a, AMSV85, AW86b, BS86, BDH^{+}85, BBT85, BDM^{+}85, Bre85, CDV87, CAWXD85, CKGP87, CCCN85, CCCN87, Com85, CDG86, CGW87, DT87, Di 85, DGM86a, DGM86b, DFZ85, EMS85, FP86, FO86, FGHS85, GPR87, GZ87b, Giu85, Gra85a, Gra86b, GWW85, GBB87, HT86, Han87j, Hen86b, Hey86, HSS85, IH86, JP86, KZ85b, KPZ87, KS85a, LL86, LS86a, LP86, LRT85, NFM86, RLQ85, RR85, Rol85, Roo86b, Ros86a, SSS86, SJ85, Smi85b, Tal86a, Tau86c, Thr86, Tua86, Tur85, VZ86, dPR86, vO86]. Helical [Sak86]. Henon [RV89]. hereditarily [NP89a, Nei85]. hereditarily- [Nei85]. Hereditary [GG86, Jøn85, Bae86, Len86, Mer86, PR89a, Sch86f]. Hereroclinic [BT88b]. Hermite [AK87, dSCV85]. Hermitian [CFM86, Kri85b, SVS87, Siu85b, Vol88c, Vol88f, Wid85a]. hermitienne [Bar86d, Dem86a]. hermitiennes [Bar86c]. hermitiens [MVW87b]. Heterogeneous [Ieş87b]. Heuristics [Gro86g]. Heyde [KZ85a]. Heyting [AC85]. Hida [Hu88]. hierarchies [Daw86b, Ler85]. hierarchy [Gab85, AD85a]. High [DL89, HL86, Ler85]. high/low [Ler85]. Higher [Dav85b, Gup89b, Hai87a, IM89, Kos85, Sim86, Asa89, CAXD86k, CAXD86l, GR87a, Hem86, NR88, Wil88]. highest [Sel88e]. Highly [Sto85l]. Highly-related [Sto85l]. Higman [Hua86]. Hilbert [AL86, Ani88, Bru89, CQ85, Cil87, Coh87c, Da 85, Die88b, DLdS87, Fri89, Gal85, Gri86a, Hir89c, Hud89, Iar85, Kle87, Lar88, Lis89, MS89b, PlT88j, Pov86, Rou88, Ste89b, Sul87, Szp86, TC85]. hilbertiennes [Mey88b]. Hilbertkörpern [Mat87b]. Hilfe [KPS87, Pro85]. Hirzebruch [Mül87f]. historical [Bra86j, Tău86a]. hitting [BT88a, CL89b, Pin86]. Hjelmslev [Mat86a, Mat86d]. Hochschild [Bry87b, Cib86, Hap89, Hüb89a, Hüb89b, Hüb89g]. Hochzusammenhängende [Sto85l]. hochzusammenhängender [Sto85p, Sto85o]. Hodge [And86a, Azn87, Bar87b, CS87a, CKGP87, CKS87b, GP87b, HZ87a, Hai87b, HZ87b, Har85a, Kas87, Loe87, Pue88, Sch89e, SZ87, Ste87a]. Hofen [LRT85]. hold [Kaw89]. holes [FTO87]. holomorphe [Dem86b, Wei86c]. holomorpher [Kos86a]. holomorphes [BCZ86, Fug86, Vig86]. Holomorphic [Rea86, Ves87, AV88a, BB86, Bel87b, BS87c, Cha88, FM88, GMM88, Lig85a, MS87a, Mok86, Pou88, Sal85a, lS89, Sta85, Ste87b, dM88]. holomorphy [Ara87]. holonomy [Bry85]. homeomorphic [FKV88]. homeomorphism [BC85]. homeomorphisms [Anc87, Bou87f, Kif86]. homogène [Akk88]. Homogeneous [Joh88b, Str87b, BHS88, Bou88b, Cam88, CG86a, DS89, EEK88, Hri87, Jac88c, Kam85, Kat86a, KL87a, KS86f, OR87b, Oko86b, Sch86g, She86j, Sim87, tW87b, Xin89]. homogènes [AA86, FL87, Lic86]. homogenization [Att86, FNT87, Mod86]. Homological [Bae86, BF88b, Don85i, Don85j, Haj87, Don85m, MS86b, Sta88b]. homologie [Gai88a]. homologies [Bac86]. homologique [Mal88]. Homology [Ade89, DK89, Kne89b, Mùi85, Roo88, Ste86a, Abe87i, Bau87, Bry87b, BP88d, Cib86, DW87, Far87a, FT87b, Gol88, Hüb89a, Hüb89b, Hüb89g, Kar89, LMM86, Lis87, Mar86a, Mdz87, Now87, Ril89, Sin87, SS88b, VP88, Wat87]. homomorphism [Hüb89d, Shi86v]. Homomorphisms [Che86c, PS85, Avr86]. homotopic [DMW87]. homotopie [Oua88]. Homotopies [AL88b, DK85c, Kne89c]. Homotopy [Ani89, Ass86a, Far87a, JM89a, LMS86c, Woo85, tW87c, Ani86, Ass89, AH86, Bau88a, Bau85, Bau88b, BM89d, BS89j, Bru86b, Bru86c, Coh87a, DR89b, DM87b, FZ87, Hai87b, HK87a, HS87d, Ill89, Ino89, IM89, Kos85, Lai86, Lev85g, MR85a, Mas86b, May86b, McC86b, Nov88e, Sch88h, Sie85, Wt87a]. homtervals [Pre88k]. Hopf [BJM87, SC85, Wen85, vdM85b, vdM85c]. Horizontal [Rec85]. horizontales [Bar86c]. Horrocks [Aur89]. horseshoes [PV88]. Howe [MVW87a, MVW87e]. Hua [HB89]. huge [Fra89]. hulls [Ale87, Wer87a]. Hunt [Ike88]. Hurwitz [ŁR85]. Hybrid [Sár89, BR85a]. hydrodynamic [Alb88]. hydrodynamical [ABHK85, Bel85a, Fun88, Ros86c]. hydrodynamics [Str89b, vW89]. hydrodynamique [BJO89b]. hydrogen [TL86]. hydromagnetic [GR85e]. Hyperbolic [Går87f, Rec87, BBP88, EB88, Eng89, FR86, Går87d, GESEZ85, GF87, Har88a, Hör87, KS86a, Key87, Las86b, Lax87, Ler89, tLqL87, Mar88a, Mar86b, Min87, Miz88, MS85, Pov86, RB87, Roe87, Tah86, Zeu86]. hyperbolicity [Anc88]. hyperboliques [God87b]. hypercartes [Arq86]. hypercontractives [BÉ85]. Hyperebenen [Wei86j]. hyperelliptic [BBGM89, Far87b]. hypergeometric [Koo88, VJ85]. hypergéométrique [Deb87]. hypergéométriques [DF86a]. hypergroup [BH89b]. hypergroupes [GG84, Gal89, Geb86]. hypergroups [Blo86, Gal86, KS89, Zeu86]. hyperharmonic [LMZ86k]. hyperoctahedral [Gia87]. Hyperpaths [NP89b]. hyperplane [SS86]. hyperplanes [Gre89, Sch89e, Voj87a, Wei86j]. Hyperrésolutions [GP87b, Gui88b, Gui88b, Pue88]. hypersurface [HS87c, LP88a, Mut86, Sim89, vS87b]. Hypersurfaces [dC89, CAXD86i, Che89, D’A87a, DT88b, HK85, Wu87a]. hypo [SW86a]. hypo-convergence [SW86a]. hypocontinuity [AS85d]. hypoelliptic [Hua88]. hypoellipticity [Iwa87, Mai88]. hypoelliptique [Aro88b]. hypothèse [Bak85c, Jac85c]. hypothesis [SS89a]. hysteresis [Kre86b].

I* [Wt87f, tW87b, tW87c]. I*-measure [tW87b, tW87c]. I*-measures [Wt87f]. I.I.D. [Fac89, Mae87, Ber88c]. icebergs [Swe89]. icosahedron [Coh85]. ideal [Car89c]. Ideals [BV88h, Mor86d, Bel89a, Bor87b, BV88g, BV88k, CP89a, DR89a, EEK88, Fra89, GGL87, Han87f, Han87i, Lud88, McC85d, MMT87, PW86, Row89, She86l, She86o, Shi88a, Sim87, Zwi89]. ideas [Ber87b]. Ideaux [Mal86a, Big85, Reu86a]. Idempotent [Blo86, HvL88, Nag88]. idempotents [Mar86a]. identical [Muk89]. identically [Mel89]. identification [KL87c, Tan88a]. identité [Bou87a, Dem86a]. identités [Str85a]. identities [And89, Bak87a, BA85, Bre89a, Gia87, NZ89, Thr89]. identity [And89, Wei87d]. if [Fre87]. IFR [Obr89]. IFR-distributions [Obr89]. ignition [Beb86]. ihre [Sto85l]. II [AvW85, ABS87, DGM86b, HT86, Hla87b, Thr86, Ves85, BGV86, BDE85, BKK87, DF86a, DH88, Die85a, Don88b, Fef88, Hal86, HL88b, Hla85c, IN86g, IN86n, IN86j, IN86k, Jac86a, Jar85b, Kai87c, KS86a, Kir87b, Kon85i, Lio89, LS88d, Mül89, OS85, PS88a, Sei89b, She86b, She87c, She87e, Tit87, Wüs87b, tDL86]. II- [Ves85]. III [CHS85, IN86g, IN86o, IN86l, Jar85c, Sat86]. IIMAS [Hen86b]. ill [Nat86]. ill-posed [Nat86]. Illustrative [vEdG85d]. ILU [vdW86]. ILU/SIP [vdW86]. Im [Hla85b]. Image [MT87, MK88, Var88b, Bri87, Var86b]. images [CFdG87, Hir89a]. imaginary [IN86k]. imaging [Ble85b, Grü86b]. Imbedding [AS85f, Bla87, Boj88a, Hu87, Mét87]. Immersing [Kir89l]. immersion [Car88b, CT89, Lev89e]. Immersions [BWW85, Giu85, Lev85c, Lev89f, Lev89g, BR89, Blo85, DD85, Eel85, KL88, Kos88b, Lev85e, Lev89b, Lev89d, Lev89j, LT89, LX89, Rec85, Sak86]. Immeubles [Tit86]. immigration [Göt86]. immune [Fri85a]. impaire [Deb89]. Implications [Abe87d]. implicit [Arn88a, DL87, Pic87]. implies [Aus85, NP89a]. imposed [CAXD86d]. Improper [CAXD86o]. improved [AP89]. Improving [CP87]. impulse [IN86c, Mos85]. inadmissible [Cho85b]. Incidence [Dür86c, Dür86b, Dür86e]. inclination [Fin88c]. inclined [Poz88]. inclinées [AS89b]. inclusion [PR89b, Ste86b]. inclusions [Gel88]. Incomplete [Net88b, Tie88, JJW86]. incompressible [HN86, Bei88]. incorporating [Kim86]. increasing [IN86k]. Indecomposable [Pan85a, Rey86b]. indecomposables [Vos86a]. indefinite [BT86c, Don85a, ES86, Hoy85]. Independence [DH88, Day85]. independent [CAXD86f, For88a, KZ87, Kon85b, Kor87c, Kru85, Mel89, Shc85, dA85b]. independent-oscillator [For88a]. Index [Brü87c, COT89, FH89, Mül87h, Mül88, BG89b, Bli86, Bli88, BH87, Cra86a, Fen89, Got88, Gul88b, GZ86b, HRL89a, HRL89b, HRL89c, HRL89g, Kas86, Kog86e, LYZ89, LS88m, Mül87a, Mül87m, Ulr88b, Van86a, Wit88b]. India [All85a]. indices [Bli86, Kom89a, Ran88, Wei87g]. Induced [Mac88b, DvdB87, Gro87b, Las89]. induction [Der89, MVW87a]. inductive [Poh89f]. induites [DvdB87]. Inégalité [Wu87d, Lin86a]. Inégalités [SY87, Yor85b, Yor85d, Bak85c]. inequalites [Nag86a]. Inequalities [AVV88, Min87, Tab86, Bar85a, Bas87, Bér85a, Ber87c, Boj88a, Bou87d, Bou88c, CM89b, CC85a, Dav87c, Fer85b, Ger86b, Jac89, Kan86a, Ken89, Lon88, LS88h, MTJ87, Mil88d, Pop87a, RW85, Wan88]. Inequality [MS86k, ABB85, BMMP88, Bou89b, Dav85a, ÉY87, Gor88, HB89, Hei86, Jac87, Kik89, MP89a, Mil85d, Mil87b, NS86e, ST89c, Wro87]. inertia [MP89b]. inertial [MPS87]. infections [Had86]. infimum [GM86]. infiniment [GG84, Gal89]. infinitary [Poh89a, Yas89]. Infinite [Amb89, BM89a, Dra87a, FQ85, Hid89, LS88h, Ara87, Axe86, BR88, Bor89, Brü87b, Buz88, CFdG87, DL86a, EH88a, FFR88, FGL86, Gre86f, KP87a, Kör87b, Kuh89, Kuo88, Lio89, MS89a, McC85c, Neu89, Sel86, Ton87b, Wil87, Zab87, dS87]. infinite- [Bor89]. Infinite-dimensional [Dra87a, KP87a, Sel86, Ton87b]. Infinite-finite [LS88h]. infinitely [Axe86, Csö89, DM89a, GH85a, Il’85, LeP89a, Wai89]. Infinitesimal [Yas89, Har85a, Kis89, Yan89, lY87]. infinity [Bah88]. influence [Car85c]. Information [DB87]. Ingham [BR85a]. inhomogeneous [Wei87r]. Initial [AL88a, Fu88, Mar88a, Rac88, ZG88, Alb85, BH89c, Emb87, Jac85c, Joh89a, Pon88, Sch89b, Sch86m, Ste86b, Yor85c]. Initial-boundary [AL88a]. Injection [DvdB87, DvdB87]. Injective [HPS87b]. injectives [LS87b]. Injectivité [Ofm86b, Ofm86c, Ofm86a]. Injectivity [Haa85, KS86e]. injury [SW89b]. Inner [Ral87b, Kur87]. innovation [PSS86]. Input [Bar88a, BO85, GMW87]. inserting [Shi86h]. inspired [FS88b]. instability [Bel85b, LB86]. instables [GSZ87]. instanton [BT87a]. Instantons [BM89d]. Institut [HSS85]. integers [All89, EH88a, Ger89, Nat89b, Tay85, WS87]. Integrability [Kon85a, Bra86c, CL89a, Kif88a, LV89, Rei88a]. Integrable [Asa89, CS88a, Car89b, Dho85, FG88, Fom88a, Fom88b, Sev86d]. Integral [CT88, GN88, Han87c, Kri85b, LS88i, Ran87, Sil89i, Asa85, AL88c, Bre86b, Bri85a, Fil85, Fuj87, Gli86, GMM88, Han87b, Her87b, Hoh87, HSW85, IN86e, JM86a, Jur85, Khi86, Kon85e, KJ86, LC85e, LM88b, Mad85, Maz86, McG88b, Mei87, MH88d, Mor86d, Ne’85, Oli85, Rod87b, Rog88, Rub86, Sch86a, SS88c, Vir88, Voj87f, ZG88]. integral-differential [ZG88]. intégrale [FL86a, HM88c, JM86b, Maa87, Ofm86e, Par86c, Roo86a]. intégrales [God87b, HM88b, dC85c]. Integrals [Kon85b, Ada88, AHKH86, AV88a, Ask89, Bra86c, Dav89, DZ89b, Dre86, Går87b, Går87h, Går87k, Gon88, Hai87b, HS85, Jon89, Jur89, Kno85, Nua86, Nua88, Orl87, Por85, RW85, SC85, Sch88b, Sch88d, ST89l, Tre87, Wal86b, dSCV85]. integrated [Ben89, Joh86]. Intégration [Pra88, AF88, EM86, FRN85, HK87b, Mas88b, Nol86, BM87, Ber87d, Kré88, Ofm86b, Ofm86c, Ofm86a]. intégraux [Bou86b]. Integrity [BV88i]. integro [Gar88c]. integro-differential [Gar88c]. integrodifferential [DL86a, Eng89, GS86, Pug86]. Interaction [Bea87, BM88, Hel88i, MR89b, Pap88a, Pic85]. Interactions [Roo86b, AGHK^{+}86]. interchanging [Shi86i]. interface [Blu85b, Tvr86]. interface-problems [Blu85b]. interior [Hua88, Kes88c]. intermediate [Rau85, Rei88b]. Intermittancy [Tak88a]. Intermittent [Gus87]. internal [Gab85]. International [BDH^{+}85, CDV87, CDG86, DGM86a, DGM86b, DH85b, GPR87, IH86, KZ85b, KPZ87, LRT85, NFM86, SSS86, Tua86, VZ86, vO86]. Internationale [Bre85, Giu85, Gra85a]. Internazionale [FP86, LP86, Ros86a, Tal86a]. interpolants [PPP87, Pie88a, Sta87c]. interpolating [Tho87]. Interpolation [AM88a, DP88, Hig85, Hla85b, KP86f, LS85a, RTVW88, Sae87, Bak89, Cai85, Cas86, DGM88, Dec88, EM88, Has85, KS87, Nei88, PX87, Roc88, ST89d]. Interpolationspolynome [Has85]. Interprétation [AY87, HvR88, Web88, dSCV85, vEdG85e]. interpretations [Koo88, Shi86f]. interpreted [Yas89]. Intersection [Kir89m, CAXD86o, CKS87a, Don85a, Gro87a, Gul86a, Hem85, Hul85]. intersections [Deb89, Ebe87c, Hüb89f, Net88b]. interval [BC85, CM88c]. intervals [Oko85]. intimate [Dav87a]. Into-isomorphisms [Jar85b]. intrinsèques [dlP87b, dlP89]. Intrinsic [Bor85, BR85b]. Introduction [Abe87e, Aga85e, AS85g, Bra86j, Col88c, Con85a, Dau88b, Die86e, Don85k, Ens86, Fie88g, Fis87e, Fut88f, Gla89c, Han87d, HPS87c, Hüb89c, Imk88b, IN86d, Jaj85d, KP86g, Kaw85a, Kir89n, Koc86, Kon85c, LP88d, Law86, Laz88, Lev85d, Lev89h, LS88j, MN88f, Nov88d, Ocn85d, Pet89f, Phi87d, Phi87e, Pie89c, Poh89d, Pre88d, Sch86n, Sev86b, Shi87e, Shi88c, SS85c, Swe85, Ulr88c, Van87c, Vol88a, Wan87f, Wei87l, Wei86k, Wid85d, Win85d, Wri89c, vdM85d, Hal89, Har85a, Hel88f, Isr89, Jai85, LC85j, May86c, MS85, Roa88, Sch85c, Sta88b, Wal86a]. Introductory [COT89, GR85d, Lan88c, Sel88g]. Invariance [DL86b, Kaw86a, Thr89]. Invariant [Cha88, Fom88a, KSLP86, Olv87, PS87b, Sul87, Als88, AFP87, BJM87, BV88f, Col88a, Cra86b, Fut88c, Gar88c, Gil89a, Gro87c, Gro85, HS87c, KP87a, KS86d, Kuc86b, Lun86b, Mañ88, Mil87a, NZ85, Pre88a, Rea88, RZ87, Sab85, Shi86l, Wan89, cY86]. Invarianten [Sto85j]. invariantes [San87]. Invariants [Bou86b, Ebe87c, Ocn85e, Pom87, BB88a, CS85b, Deb87, Dix85, For86, For87, Fut88b, Gan85c, Gan85d, Gea89, Gra87, Kem87, Kos85, Kos88a, LS87a, LP87a, Lev88, LM85, Loe86, Lüc89b, Lüc89e, PR85a, PM89]. Inverse [Bar86a, BDT87a, Bér86g, Ble85b, IN86e, Sch86i, Win85e, Win85f, Anc87, Ber86k, BFKP86, CRD87, FS88b, Men88b, Sai86, SU87, Tan88a, Tal86a]. inverses [CC87a]. inversion [Pik86b]. invertible [Cat85a, Khi86]. investigation [Pas89, Sho87]. inviscid [DL89, LF89]. involution [Gia87, Son85, Vog85]. Involutions [HR85, KB86, Jos87a, Jos87b, Mas86b]. Involutorial [Sel88h]. involving [AL88c, SS89b]. ion [Str89b]. Irees [JV85]. Irrational [Hov89]. irreducibile [Han87f]. Irreducible [Atz88, BSK87a, Mum88q, Hei87, Jež85, LM86a, Lud88, Nag88]. isentropic [Alb88, tLcZ88]. Ising [Eva85a, Hig87, IN86q]. Isoclinic [Gol86a]. isoenergetic [Fom88b]. Isolated [AU86, Sim85c, Aus85, Die87, LP88a, Rei86, qW88]. isolé [Tho88]. isometric [DD85, Rec85]. isometries [AA89, DLdS87, Kan86a, Ves87, Jar85c]. isometry [Gro85]. isometry-invariant [Gro85]. Isomonodromic [IN86f, IN86h, IN86g]. Isomorphic [Mil88d, ST89n]. Isomorphism [BGO^{+}89, AM88g, Gla85, Ill89, Ocn85h, San85b]. isomorphismproblem [Rog88]. Isomorphisms [CS89e, Han87e, Jar85b, Jar85a]. Isoparametric [PlT88g, Mut86, PlT88k, PH89]. Isoperimetric [Bér85a, Bér86c, Bér86d, Gul88a, MS86k, Bér86b, Lon88, Mil85d, NS86e]. isospectral [Bus88a, Dho85]. isotopies [HR85, MR89a]. Isotopy [Kor88]. Isotropic [Har86, MP89b, DL88, Sel88e]. Itô [AOP86, BW85a, Gli86, KÜ88, Par88b, Qua88a]. Italy [Bre85, DT87, DFZ85, FP86, Giu85, Gra85a, LP86, Ros86a, Tal86a]. Iterated [D’A87b, Hai87b, Har88b, Shi86i, Asa89, EB88, Fis87f, Fra89, HKW89, KK85, Kom89a]. Iteration [Coh89, Éca85, Kat88b, Ban86, ES86, Meh85, Mil85e, Nev89, Rei85, Sch85i, Tor87g, LRT85]. Iterations [Pra85, Pet89c, Pet89e]. Iterative [Pet89g, Man86a, Roa88]. Itineraries [Hel85]. IV [ESS85, Jar85d]. Iwasawa [FG86]. IX [Yan88a].

J. [Bau85, Par86c, Sch87a, tW87a]. Jacobi [Rof85, BG88a, Eva87, Går87e, Koo85, Str86b, Yui88]. Jacobian [CCS85]. Jacobsthal [Kat85b]. January [All85a]. Japan [SSS86]. JB [HO85]. JB-algebras [HO85]. jets [Ger88c]. jeux [Del85]. JH* [Sch88j]. Johnson [CS89d]. Joint [Kon85d, Kre86c, Aoy86]. Jones [qW89]. Jordan [Cle88, GLM85, GR85k]. Journé [Par86c]. Journées [FHMM86]. Jr. [Whi89]. Juin [Nor86a, FHMM86]. Jul [Han87j, Tur85]. Julius [JTW89]. July [BDH^{+}85, Bre85, Com85, FO86, Giu85, Gra86b, GBB87, Hen86b, IH86, JP86, RLQ85]. Jump [BCSSC87, CM88a, CM88b, LB86]. Jumps [Imk88c, GJN88, Had85, Kis86]. June [BS86, CDV87, CKGP87, FP86, FGHS85, GPR87, Giu85, Gra86b, GBB87, HSS85, KS85a, LL86, LP86, NFM86, RR85, Rol85, SJ85, Tal86a, Thr86]. k* [LP88g]. k*-action [LP88g]. K. [Kát85a]. K.M.S. [Cec85]. K.P. [AD85a]. Kac [AHKHK89, Sel89a, Tit85]. Kähler [Aub89, BK88a, BV88j, Bry87a, DP85, Eno88, Fut88g, KS86f, Mok86, Sam85, Siu85b, Siu87]. Kählerian [Bej85, Fut88b]. kählérienne [Var86b]. kähleriennes [Ala88]. Kalman [GMW87]. Kaluza [Ker87]. Kan [KP86h]. Kasparov [Hil85b]. Katata [SSS86]. Katsylo [Hir89c]. Kazhdan [Shi86d, Shi86f, Shi86j]. Kchinchine [BG89d]. KdV [Dho85]. KdV-type [Dho85]. Keller [CCS85]. Kepler [ABHK85, LS88a]. kernel [Bel87c, Fis87f, Gra87, LP88e, LM88b, Mül87e, TC85, Wat88, vEdG85c]. kernels [BH89b, Fal86, Fd89, GLM85, UW88]. Kervaire [BJM87]. killed [Kis89]. killing [Had86, Sán85a, Zhe87]. kind [Pik86b, Tre87]. kinds [Shi86s]. Kirby [Mil88b]. Kirwan [Sch87h]. Klasse [Gol87b, Vog87a]. Klassenzahlen [Mat87d]. Kleene [Bla85b, SS89a]. Klein [Bac89, Bâr88c, BBGM89, Caz88, Coh85, Ker87]. Kleinian [Hem85]. Klingenschen [Wei86a, Wei86i]. Kloosterman [PS87b]. knapsack [CHM89]. Knot [Rol85]. knots [Ble85a, HLM85a, Kan89, Lit85, LM85, Ril89, Van85]. knottings [FKV88]. known [Hej85, Sab85]. Kodaira [Dem86a, LP88e]. Kohomologie [Sto85d]. Kojima [Jin88a]. Kokern [Sto85e]. Kolmogorov [Bou89a, Sev86d, Sev86f]. Kolyvagin [Rub89]. Kompakte [Fle85a]. Kompaktifizierung [Fal85]. komplexen [Kos86a]. Komponente [Sto85a, Sto85e]. konkavem [Kos86a]. Konstruktion [Sto85m, Sto85f]. Konstruktionen [Sto85k]. Korovkin [Fer88a]. Korteweg [Sch86j, Sch86l, Sch86m, Sch86o, Sch86p]. Krein [DLdS87, Sch85a]. Krivine [MS86l]. Krull [Gei86, Okn88a]. Kumanogo [Fuj87]. Kummer [Hoy89]. Kunita [Lin86a]. Künneth [Mac88a]. Kunze [Geb86]. Kurvensingularitäten [Knö87]. Kushner [Ben89]. Kutta [Cas86, HRL89g, HRL89h, NS89b].

L [Par86c, Sil88b, D’O85]. L-I [Sil88b]. L. [Gal89]. l.f.t. [Thr89]. labour [DM87a]. labour-surplus [DM87a]. lacets [Sch88k]. lack [KR85, Sch87c]. lack-of-memory [KR85]. lacunary [Dra85]. lacunes [Ell89]. Lagrange [CK85, Lio86, Wei87d]. Lagrangian [Bru87d, Bry87a, GS85a, Wag87]. Lagrangians [Att86]. lagrangiennes [Aud87]. Laguerre [BDM^{+}85, Die85b, Mag88, dSCV85]. lamb [For85]. lambda [Jac89]. Lancaster [Tur85]. Łańcut [GPS87b]. Landau [Fun88, GZ87a]. Landweber [Zag88]. Langevin [For88a]. Längste [Pro87]. languages [KS88a, Yam89]. Laplace [All85b, CD87, Dau88c, SS88c, iS88]. Laplacian [Bér86e, Bes88, FTO87, Kuo86, Lin86b, Mül87i, Mut86, Oba89, Pav85, Rei87a, Sim89, Ura86]. Laplacians [Sun88]. Large [Bro87, BT87c, Gro86f, Hop89a, IN86i, Var88a, vdBLP88, Alb88, Bac89, BCSSC85, BBD88, ESN89, Fle85b, Hei85, Hei86, Hen89a, Hen89b, Hil89, HJ85, Jai85, Jaj85e, Jaj89b, Kus86a, Lew88a, Num89, Sch89g, WS87, dA85b]. Large-time [IN86i]. largest [LS88k, Tab86]. laser [IN86c]. later [Roo86b]. Latin [Di 85, Ges86a]. Lattice [BM89c, Bre89a, Com85, Gro86g, Bou87e, Esp88, Hai85, Hoy85, KN89, Lex88, LM87, Oll85i, Rog87, Var89, dLQ89]. Lattices [Gek86f, AC85, Are86b, BB87, Ben85a, Day85, Dwi85, Ebe87e, Ebe87d, Fre85, Gre86c, Gre86e, Hig85, NS86a, Nat85a, Rei88b, Rog86]. Laurent [McC85b]. law [BV88a, Car88b, HKW89, Hei85, Hei86, HJ85, KK85, LS85b, Liu88, Num89, Owe86, Ski86]. laws [Bak87a, Che88a, Csö89, God87a, Hen89a, Hen89b, Hof87, Jaj85e, Jaj89b, Jur85, Key87, KK89a, KMM87, KŠ88c, NS89a, Oko85, Roe87, Sch86e, Fer85b]. Layer [Fab88]. layers [Shi86o]. leading [CM88a]. leaf [Paw85]. Leafspaces [Rei88a]. least [GP85, GLC86]. leaves [AV88a, Gut88, dM88]. Lebesgue [Kon85e]. Leblond [Duv87]. Lecture [Gal89]. Lectures [Bal86b, Bar85c, Bre85, CT89, FP86, GZ86a, Giu85, Gra85a, Hop89a, Jos85b, LS86a, LP86, Lev85g, MR85c, Ros86a, Tal86a, Ves85, dPR86, Che85, MC88, Sel89b, dPR86]. Lefschetz [FF89, LYZ89]. Left [Shi86l, Ced88, DR89a, Sch86j, Shi86k]. leftward [Sch86k, Sch86l]. Legendre [Lan88d, Yui88]. Legendrian [Eli88]. Leja [Ple87]. Lemma [Hla85a, Hla85c, Tic85, AS85l, AS85k, Asa86, Cha88, Kas87, ST89n, Cha86b]. lemmas [Sei87, vdB87a]. lemniscate [MM85]. length [BT86a, Oht88, Sch86b]. lens [HR85, LT89]. Lenstra [LN89]. letters [BT86a]. leur [Car86d]. leurs [AL88b, Cau88]. Level [COT89, Kaw85d, AP85b, BRA86e, Cas85a, Düm85a, Kaw88, LW89, Lon88, LM89b, McC86b]. Level-Index [COT89]. levels [LW88]. Levi [Che87, D’A87b, Oht88]. Levy [Car86d, Duv87, Fal86, Oba89, BG89d, Mad89, MRR88, MS86c, MS86m]. Lewis [BF88a]. Liaison [Kle89]. Liapunov [CC87a]. Lie [CDV87, Abe87b, AHKH86, AL88b, BGO^{+}89, Ber86i, Bor87a, CC89a, Cle87, Cru85b, Cru87, DM86, Eld89, Far89, FFR88, FS89a, FS89b, FRN85, FR85, HN89, Hsi87, Ill89, JM89a, Jan87, Kaw89, KN89, Kaw87, KI87, KŠ88c, Lud88, Lus87, Mar88b, Mur89, OP89, Reu86a, Ric87, Ros85a, Sel86, SW89a, Str89a, Var89, VG88, Zal88, dS87]. Lie-groups [FRN85]. Lieear [Els85f]. lifespan [Hör87]. Lifschitz [GZ87a]. Lifshitz [Kir87a]. Lifting [Fut88h, Lev85e, HN86]. liftings [CS88b]. lifts [Rec85]. Liftungen [Wei86m]. light [Car85c, vW85]. Lignes [Sal89, MM86]. like [DD85, Erd85, Har87b, McC85c, Sel86]. limb [Ede86]. Limit [Ber85c, Hol89a, Iva86, Jaj85f, Mel89, Pan85b, Tak88b, Tan87, Tan88b, Wei87m, vdBL88, AFL89, ART88, Ale85, Aul85, Bin86, CBD87, Düm85a, Dür85b, DG86b, FQ85, FTO87, Fun88, GZ86a, GZ89, Gra86a, He89, HS87d, Hoh87, IN86q, Jac86a, JT86, Jur85, KL86, Kon85h, Kon85i, Kru85, Led86a, Mae87, Mal85a, Mas86a, PV88, Par87b, RS87, Ros86c, Sch86h, WKK86, Yaj85b, Yuk86, Bei88, dBJ87, vW86]. limit-closed [ART88]. limit-periodic [Par87b, dBJ87]. Limitations [Cas85b]. Limite [Jac85a, Mey88d]. Limites [Ibr85]. Limiting [Fac89]. limits [AB89, IW86, Poh89a, Win85e, Win85f]. Lindelöf [Nev89]. line [Arb89, BDT87a, Coh86, Col88b, Dav87d, Ges86a, Har87a, Hel86c, Lub85, Ost85, Rös85c, Wei87c]. line-graphs [Arb89]. linéaire [Cou89, Ras87, Van86b]. linéaires [Bon89, Mar85, Mar88c, Gér86a]. Linear [Els85g, Går87j, Hai85, Her87b, Jur89, Kog86e, MH88a, Rod87b, Tre86, ABHK85, Ale87, AOP86, AKO86, Ass86b, Bai89, Bar87d, Bar86e, Ber86k, Blu85a, Böh86, BM89b, Cam86, CV87, CS89b, Che86c, CC85b, CC89b, Coh88a, DL86a, DP88, Del86, DLdS87, Din85a, Dra85, Dum85b, ES86, EN86, FRN85, FV89, Går87i, GR85c, GR85i, GI85, Gil88, Gre89, Gro86a, Gro88, GR85j, Har88a, Heg88, HSG88, HM88a, Hör87, IN86h, IN86m, Joh89a, Kag85, KZ87, Khi86, Kog86d, Kog86c, Kří86, KJ86, Lai88, Lew88b, Lie87, LB86, MR85a, MP88, MH88c, Nag86a, NW86, Neu86b, OD85b, Pic87, Plu89, Roo89, Sán85a, Sap86, Sch87c, SP89, Tur89, Via89, Wih86, Wüs87a, Wüs87b, Yan88d, vEdG85a, vEdG85b]. linearen [Hah85]. Linearer [Vog87a]. Lineares [Bar87d]. linearity [Plu85, Vog87b]. linearizability [Gol89]. Linearized [DS86a, DF86b]. linearly [Sch86k]. lines [Idà89, Sch86j, Sch86k, Sch86l, Smi86c]. Link [Kir88, Kos85, Kos88a, LM88a, Vog88]. Linking [tD89a, Ped85, tD89b]. Links [Fre89, Dur87, Jin88a, Kir89k, Lev88, Lev89c, Mur85a, Tur88a, qW89]. Liouville [Brü85, Fom88a, Fom88b, Wei87e, Wei87i, Wei87o, Wei87p, Wu87b, Xin87]. Lipschitz [BCZ86, BH89c, Bou87f, Fab88, For88b, Gon88, Hil85b, JLS87, Mil87a]. lipschitzien [Yp89]. liquid [ABL88, HKL88]. lisse [GS87]. Littlewood [Bak85c, BL86, Bou89b, Mey85a]. Lizorkin [FHJW89]. Lloyd [Lac86]. loaded [Ieş87e]. loads [Pie89h]. Local [BS87a, CAXD86p, Col88d, DGM88, HRL89f, Lev89i, PlT88h, Par88b, Ral87d, Rum89e, Rum89f, Ser89, Tay88, Tho85d, Waa87, AP85a, Alb85, And87a, AH86, BR86, Ban86, Boj88b, Cha88, Col88i, Col88h, Col88j, Col88k, Dav87c, DL86b, Éca85, Fd89, Fie88f, FGL86, Frö85, GQS85, Gul86b, HLR85, Hua86, Jac87, KM86a, LYZ89, LP88c, LP87a, Le 87, Lec86a, Lev85h, Mil85e, Mil88c, Owe86, Pra85, Rei89b, Rei89f, Rei89g, Sal85c, Sch89b, Sil88a, Sio86b, Wsc85a]. locale [Bou87b, Ser89]. locales [JS88]. localic [IKPR88]. localisation [Sch86f]. Locality [BDE85, EEŚ85, Kiw85h]. localization [AP85b, vdE87]. Localizations [Kis89]. localized [War86]. Locally [Can85, DK85d, Die88b, Hru87, ART88, Die86d, DLS86, DW87, Har87e, Kan86b, Kat86a, LP87b, MR85a, Pat88, Rob87]. Location [WKK86, Lan86, Rie87, RZ87]. locaux [Ber87d, Fd89, Jeu85, Wsc85b]. Lochau [LRT85]. locus [BV88i, SV85]. Łódź [Ław85]. log [Mal85b]. logarithm [HKW89, KK85, Reu86b]. logarithmic [ÉY87, Fac89, Oka85]. logarithms [Pra85, Wüs87a, Wüs87b]. Logic [Di 85, Ceg89, GM88, Gri85a, Kat89b, Lam88, LT88, Mak85, She86f, Yas89]. logics [Rau85]. Loi [Str85c, Alt89, Car86d, DS86b]. Lois [GG84, Gal89, God87b]. Long [Rös85c, Ste88, Ens85, Ens88, IN86c, Pic85]. long-range [Ens88]. longest [Pro87]. Longtime [Fit87b]. looking [AH86]. Loop [Seg85a, Asa89, Coh87b, Kuh89, Mit87, NR85d, NR85e, Roo88, Wit88b]. loops [Eli88]. Lorentz [Lun86b]. Lorentzian [BEMG85]. Lösungseinschließung [Col86]. low [ABGS85, Clo85a, Düm85a, Ein88, JLS87]. low-basis [Clo85a]. Lower [Ass85, Phe89g, iS88, Che87, GA86]. lowest [Koź86]. Lozanovskii [Rei88b]. Lozi [Lev85i]. Luigi [Ves85]. Lukacs [Kag89b]. Luminy [CDV87]. Lund [Pee88]. Lusin [LMZ86o]. Lusternik [Mon87]. Lusztig [Shi86d, Shi86f, Shi86j]. Lyapounov [Bou86a, Cra86a, DL88]. Lyapunov [AKO86, AW86b, AW86a, Böh86, Car86f, Cra86b, Eng86, KS86c, Kir86, Lac86, LS86b, Mor89, Wih85, Wih86]. Lyngby [Han87j].

M [DHS85, Pop87a]. M. [Anc87]. Macaulay [BEH87, Gre89, Haj87, Her87b, Rei86]. Macauleyness [Han87c]. MacDonald [Bre86b]. Mach [Emb87]. machines [HZ85]. magnetic [Hel88g]. magneto [MP88]. magneto-static [MP88]. Magnetohydrodynamic [Bra86a]. magnitude [Fit87a]. Mahler [Bom87]. Main [Fie88h, Ocn85f, LS88q, Voj87b]. maisons [Del85]. Majoration [Lap88]. majorization [Ess87]. makes [Smí86a]. Malliavin [Itō88, Kus86a, Léa88, Nor86b, Nua86, NZ89, Üst87, Üst88, Wu87c]. mammalian [Car85c]. Mandelbrot [PR85b]. Manifold [HV88, Bau88a, CMS89, Che88c, CG86a, Dav87a, FG86, Gol89, IN86j, IN86l, IN86k, Kas88, Koz85, Mod87, PS87a, Siu85a, ST88]. Manifolds [Bal85, Don85a, KSLP86, Kir89c, Lev85c, Mül87g, Mül87h, Rol85, Sch85b, SSS86, Alm87, AS85b, AM88c, And88, AV85, Bal86a, BK88a, Bat87, BEMG85, Bej85, Bla87, Blu85c, Blu86, BS89f, Bry87a, Can85, CT88, Car86f, Cha88, CH85, Chr89, Cra86b, DH89, DP85, Don86b, Don88c, Drd85, DW87, Ebe85, Eel85, Elw88a, Fuk86a, Gab85, Gli86, GPT89, GR87b, Gro86f, Gul88b, Hai87a, Hat89, Hil85a, Hil85b, HLM85b, Jin88b, KL88, Kam89a, Kam89b, Kan86a, KS86d, Kat86b, Kir89e, Kir89l, KS86e, KS86f, Kos88b, KLR85, Kus88, LR89a, Lev85e, Lev85f, Lev89g, Lev89j, Lew86a, Li89b, MPS87, McC85a, Mil85c, Mok86, NR85f, Nis86, OR87b, Oko86b, Pan86, PS88b, Pup88, Sch87k, Shi87c, Shi87b, Shi87a, Shi87f, Siu85b, SS88c, SS88d, Sto85l, Thi86, Tur88b, Veg85, Xia87]. manifolds [Zhe87, Žub88]. Manin [Ati85a]. Mannigfaltigkeiten [Kos86a, Sto85l, Sto85p, Sto85o]. many [Ber88a, FTO87, Gel86, Kat89b, Kaw89, Oza88, Rhe89]. many-valued [Gel86, Kat89b]. map [BGS89, Fut88d, GR87b, KK89b, LP88e, Lev85i, PS87a, Wan87a]. mapped [Koo85]. mapping [BW85c, Böd87, Coe87, CK85, GM85b, HK85, KS85b, Pap85, Var88b]. Mappings [Blu86, Giu85, Pre88e, Bel87b, Bor89, CT88, CS87b, DP85, FQ85, Fer88b, Fer88c, Gel86, Hil85a, Hög89, Ker86, Li89b, Lig85a, Pan86, Pet89h, Pre88f, Vuo88d, vEdG85a, vEdG85b]. Maps [DZ87, DW88, KSLP86, Lev85c, AS85d, Bar87e, BG89a, BR89, BF88b, Bou87f, Bré89b, Bur87, ČW87, CM88c, DK88b, Dyl88, Erk88, FZ85, Frö86, Glo87, GK87a, Haa85, HN88, HJ87a, Hel85, Jaw86, Jia89, Jos85b, Jos88, KS86d, Kir88, Kom89a, Kos85, Kos88a, Lev85e, Lev85f, LX89, LC85g, Lon85, Mañ88, Mat89b, Phe89f, Ran89, Raw85, Sal85a, Sam85, lS89, Sie85, Spe88, Ulr88b, Ura86, Woj87, Woo87a, Xin87, Xin89, Yag87, vdM85a]. Marcel [Cho89c, Pee88]. March [ACSC85, DFZ85]. Marchaud [Wro87]. Marche [Geb86]. Marches [Ibr85]. Marginal [LR89b, Att86, HJ87b]. Marie [Mal85c, Mal86b]. Markoff [Haa87]. Markov [Acc89, AHKH86, Bes85, BK88b, DL86b, Fri85b, Fuk88b, Kis86, Küm88, Le 88c, Mal85a, Nag86a, ŌT87, PS87c, Rog89]. Markovian [AJL89]. marriages [HPS87d]. Marseille [ACSC85, CDV87]. Martin [CM88c, Sal89, Woe86]. Martinelli [LT87]. Martingale [Cha86c, Rub86, AY89, Bin86, DM87a, Dur85c, Edg85, Hei86, Jac87, JM86b, Par86c]. Martingales [Bur86, MS86n, Pic89a, Wal86b, BW85a, Bur89, Éme88, Éme89, FL86a, FS85b, Imk88c, Imk88e, Jaj85b, Kaz89, Kik89, Mey88d, Mey89d, PS88a, Sch86s, Yor85b, Yor85d]. Maryland [AH85a]. mass [BN88, Xu87b]. Massey [Eke86a, FFR88, Lau86]. matchings [Str86b]. Matematico [Bre85, Giu85, Gra85a, LP86, Ros86a, Tal86a]. materials [Ieş87a]. Matermatico [FP86]. Mathematical [BFKP86, BN88, Di 85, DH85b, DFZ85, Fei86a, GPR87, Tau86c, AL88a, Eke88, Gar88a, Grü86b, HKL88, KMR89, Man86b, Rug89, vEdG85e]. Mathematics [ACSC85, ABS86, ABS87, CB87, CDG86, GPS87b, Bra87]. Mathematik [HSS85]. Mathieu [IN86o]. Matric [Lau86]. Matrices [Aga85d, Que87f, Aga85a, Aga85b, Aga85c, Bou86a, DT89, For87, Kri85b, Le 89, Led86b, Rob86, Vog88, Vol88c, Vol88f, Zal88]. Matrix [AS85i, BHS88, Dav85a, Que87g, AS85c, For86, LP87a, Pet89h, Reg88, SZ85, Seg85b, Wal86c, Wei87f]. Matroids [BS89d]. Matscience [All85a]. Matter [Går87a, Går87c, Hel88a, Hel88d, Hel86a, Hel86e, Jac88b, Kir89d, Kir89j, Lüc89c, Lüc89d, Mic89a, MS86g, MS86j, Mum88l, Mum88m, Mum88n, Mum88o, Nel89, PlT88b, PlT88f, Phe89b, Phe89d, Pik86a, Poh89b, Poh89c, Rei88c, Ros89b, Sim85b, Sio86a, Tim86a, Tro89b, Yaj85a, Zha89a]. Maurey [MS86o]. Max [HSS85]. Max-Planck-Institut [HSS85]. maxima [Kru85, OW87]. Maximal [Que87h, VS87a, BE87, Bou87d, Bru89, BV88h, BV88k, Ced88, DV86, Des85, Jac87, Kha89, KI87, Li89a, Mai88, ST89k, Wei87n]. Maximale [Pol86]. maximaux [Cau86b]. Maximum [Kaw85b, Bar85a, Bou87a, Die88a, Gol86a, HS87a, Lum87]. May [KPZ87, LL86, LP86, Tal86a]. Mayer [CGC88]. Mazur [ST89n]. MD [He89]. MD-semigroups [He89]. Mean [Jut89, Pin86, Sch86s, Ber88b, Che89, Die88a, Már89a, Oba89, dC89]. means [BV89, Cle87, Din85b, GN88, Mur87]. Measure [Daw86a, Kuč85, Ale87, DL86b, EH88a, Ess87, Grü88b, Jon89, MRC89, Mét87, Qua88b, Ros86b, Tay86, Wt87d, Wt87a, Wt87b, tW87b, tW87c, Wt87e]. Measure-valued [Daw86a]. measurement [Hol89a]. measurements [Hol89a]. Measures [Gal89, Hey86, AHKH86, Blo86, Brü87b, CM87c, Col88a, Cra86b, DM86, Ele87, Fuk88b, Ger88b, Hei88a, Hel86f, HL88b, Jan86a, Jen87, Kan88b, Kan86b, Kau87, Kec88b, Kes88a, Kiw85h, LL89, Lóp88, Łuc89f, Ost85, Par88b, Que87f, Rac85, SW86a, Sch86e, Sio86b, Sto87a, Szé86, Wt87f, Win85f]. mechanical [ABGS85, tW87a]. Mechanics [Gut87, Kon85g, Oll85g, ABC^{+}85, Att86, BT88a, Bel85b, BB85, BCSSC85, Car86a, Car88a, Cas85a, Cas85b, DDL85, Fei86a, GR87b, Isr89, JL85, Lew88a, MPST85, Nel88, Nol86, Oll85b, Oll85i, TL86, WST88, Yas88, vEdG85e]. mechanism [Bel85a]. Medford [BDH^{+}85]. media [Car89c, Del86, Ieş87b, Pap88b]. medians [Urb86]. medical [Grü86b]. medium [Aro86, Tan87]. Meet [Han87f]. Meet-irreducibile [Han87f]. meeting [HSS85]. Mehrdimensionalen [Col86]. Mellin [BM87, Sch87l]. memory [KR85]. Menchoff [LMZ86o]. merit [BT86b]. Merkurjev [vdK85]. meromorphe [Yge87]. meromorphes [Pas87a, Adi87, Rab87]. meromorphic [Bry89a, tC87, GMM88, Hel87, Rec87]. mesh [Ban86, Gor88]. Mesure [Wsc85e]. Mesures [NÜ89, DZ89a, Fer85a, Kah88]. Metaplectic [Rei88d, BH89d, Rei89b, Rei89a]. métaplectiques [MVW87e]. Method [IN86f, Kač86, Aga86b, Bra86b, BT86c, CC87a, DF86b, DDL85, Els85a, Els85d, GMSW86, GA86, GLC86, Gyö85, Jar89, KP86c, KP86k, KL88, Koi87, Kos88b, Lec86a, Mar86b, MV86, Mur88c, Noc86, Rek86, RTVW88, RS87, Roo89, RZ86, Sch85c, Sla85a, SW89b, Spe86, Vas86]. Methode [Sto85i, Ber85b]. Methoden [Hla87a, Pro85]. Methods [AS85i, Bre85, DH85b, DFZ85, GR85h, GPR87, GWW85, HT86, HS89b, Kiw85e, Kiw85f, Kiw85g, Kiw85h, LMZ86j, WS87, Aba86, All85b, BO85, Bér86b, Bér86c, Bér86d, Ber86k, BPPS87, Ble85b, BS89b, BS89c, BS89h, BF88b, Bur87, BR85c, Car86e, Chu88b, DGR87, Dob85, Fab88, Far89, FW89, FV89, GR85a, GR85b, GR85g, GR85i, GP85, GT85b, Gri86a, HN86, HRL89g, HRL89h, Hen86a, Her86, Jen86, KP86b, Kiw85a, Kiw85b, Kiw85c, KH86, Lab86b, LMZ86q, Mae87, Man86a, MPST85, Mas88b, Mor89, NW86, Nir89, Nit86, NS89b, Nov88c, Pap85, Pet89d, Pet89g, Pis86, Rav85, Raw87, Sch86c, SH86, Str88b, Tho86a, Whi86, Wil88, Di 85]. Metric [Gro87d, Her88c, KV85, Yuk86, Bar89d, CK89, Ill88, JLS87, Kif86, Koi87, Min87, Pat87, Rad87, Rog89]. Metrics [Bry85, BK88a, CH85, Fut88g, GPT89, KS86f, KS85c, Nis86, Sch89f, Siu87]. Métriques [Aub89]. Metrische [BDT87b]. Mexico [Hen86b]. Meyer [Acc88, BW85a, Par86c, Str88a]. Mg [SH86]. Mg-methods [SH86]. Mhaskar [Sta87b]. Michell [Ieş87e]. micro [vdE87]. micro-localization [vdE87]. microlocale [Bon89]. Microlocalisation [Spr85]. Mills [Bou87g, Don89a, Hen85, Sán85a, lS87d, Xu87b]. Milman [Gor88, Sch85a]. Milnor [Ebe87d, NR88, vS87b]. minima [Col88d, Col88i, Col88k, Hel88b, Por85]. Minimal [BC86b, Bry87a, DK88a, Eel85, Giu85, Gul88b, Hom85, Hop89b, LC85k, Lud88, MM85, Mos89, Ros89c, ABL88, BC86a, BC86c, BC86e, Ber88a, Bor88a, Bue87, Ceg87, Cha85a, Eno88, EN87, Ess87, Hei88b, Hof85, Kaw86b, KS85c, Leh86, Mil87a, Mut86, Pas86, RVV85, She89b, Tro88, Wei87n, Wt87c, Xia89]. minimality [Mas88b]. minimization [Kiw85a, Kiw85c, Kiw85g]. minimizer [Wu88b]. minimizers [Gia86, Gia88b]. minimizing [Kiw85h, Tay88]. minimum [BN88, Fuk86b]. Minkowski [Bac89, BLM88, CK89, Gu87a, Wei85a]. minkowskiennes [Cru85b]. minorée [Bak87b]. minors [BV88a, BV88h, BV88k]. Miscellaneous [HPS87d, May86b, LC85d, Pre88l]. Miscellany [Don85l]. Missouri [KS85a]. Mixed [Néd86, Bry89b, HZ87a, Hai87b, HZ87b, Las86b, NW86, SZ87]. mixing [BR88, DL86b, GK88b, Kan86b, Qua88b]. mixtes [Azn87]. Mixtures [KT87, Res89]. MKdV [IN86i]. Mladineo [Lev89g]. Möbius [AFP87, AS88, DS89, Dür86c, HS87a]. modêle [Oua88]. Modal [GM88]. mode [IN86p, NS87]. Model [MMA86, Ocn85h, Ocn85i, AP89, Beb86, Bel85a, BGR85, Coh87b, CG86b, D’O85, Dic85, Dow89, Eva85a, For88a, Fun88, Gri85b, Hig87, IN86q, Kaw86b, Kei87, Kim86, Kot89, Lac86, Ocn85g, PR85b, Pic85, Rie87, Wt87c]. Modele [Leh86]. modelled [LR89a]. Modelling [NR85a, Bri85b, Smí86a]. Models [Ede86, GM88, KZ85b, KPZ87, She86a, Win85b, Cho89b, Din85a, Fan89, Har87b, Kif88b, Kni87, Nag86c, Ne’87, Ros86b, Sán85a, Sch88i, She86j, She86k, She89a]. modérée [Oua88]. Modern [Bor88b]. Modification [dBJ87, tD89b, CM88a, Las89]. modifications [CM88b]. Modified [CS89f, BJ86, Sch86l, Sch86p, Wit88a]. modulaire [Gil89b, Vig89]. Modular [Gek86b, Gek86g, Gek86h, Jan86b, Kno85, Kri85d, Kri85c, Kri85e, Kri85f, Woo85, Zag85, Cha86a, Coh87c, Coh89, Fal87, Gou88c, Hol89c, Hru87, Iva88, Lan88b, LW88, LW89, Mùi85, Mur85c, Sch88f, Stu87, Wei86n, Wei87v, Wei89a]. Modularer [KPS87]. modularity [Tsc85]. Module [Rum85, BV88j, Car85b, McC85d, She86h, Tay85]. Modules [Ben86a, Gur85, vdE87, AM88e, Aoy86, Bar87c, BP88c, BV88e, BEH87, Cae88, Cau86a, DvdB87, Des85, Dre87, Fis86, Gek86c, Gek86d, Han87g, HSG88, Her87b, Les88, LV89, Lis87, Lüc89a, Men88a, Mum88d, Mum88t, Reg87, Rei86, Sel86, Sel88e, Sel88d, Sel88h, Sel89a, Wun87, Zay87]. Modulformen [Wei86o, Wei86n]. Moduli [Bui86c, Cat88, Shi88d, Don89a, ES87b, Fab89, Fac89, Har88b, Har88c, Lau86, LP88b, LP88c, LP88f, LP88h, Mar86c, Wro87, Dec88]. Moduln [Knö87]. Modulo [Drm87, Dra87a, Mer86]. Modulräume [Kos86a]. Modulraums [Fal85]. Modulus [Vuo88c]. Moh [tDP89]. Moisezon [Pet86, AS85b]. mollified [Neu86c]. Moment [RW85, Ale85, AOP86, Bri87, Coo86b, Fut88d, Ran86]. moments [All85b, LeP89b, Plu85, Sze87, Urb86]. momentum [CK85, KS85b, vdM85a]. monômiaux [Les88]. Monadic [She86f, Cai85, MO88]. Monge [Che88b, Fuk86b, Tru86, Won87]. mono [Ras87]. mono-dimensionnelle [Ras87]. monodromie [Bea86b, Gai88b]. Monodromy [Bar86b, Ebe87e, IN86m, CKS87b, Ebe87d, IN86b, IN86a, IN86e, IN86j, IN86k]. monoid [CK88, McC85d, Okn88a]. monoids [Alm88, FK88, Lal88]. Monomial [Chr87]. Monotone [Phe89h, Axe86, BP88b, EM88, Hir85, Lub85, Phe89f, Pre88f]. Monotonicity [Gro87a, Tho85e, Grü88b]. Monte [Nov88c]. Montecatini [FP86, Giu85, Tal86a]. Montesinos [Tur88a]. Montréal [LL86]. Moody [Sel89a, Tit85]. Moore [EN88, FT88b, Sel88a, Smi85a]. Morava [Kuh89]. Morita [Hah86]. morphism [Mum88g, Mum88h, Mum88j]. morphisms [Bâr88c, Her88b, Mum88s, dP86]. Morse [Gag88, qW88]. Moser [Tab86]. most [Erd85, Fis86, RZ87]. motion [Bas87, Cas85b, CL89b, Dav87b, DT88b, Elw88b, Gar86a, Hoh87, Kuo88, Lew86a, MRR88, Neu86c, Pin86]. motions [Liv87]. motivated [Grü86b, Mor88]. Motivations [Bér86f, Fri88c]. Motives [Sch88e, Sch88f]. Mountain [Cha86b]. mouvement [Le 88b]. mouvements [Bia89]. movability [Yag87]. movable [IN86n, IN86o]. moving [ČZ86, Sch86j, Sch86k, Sch86l, Sto86]. Multi [BRA86e, AJL89, BM86, ČW87, Fuc86, GH85b, GMW87, Hem86, LF89, Les88, Njå86, Wak87]. multi-component [LF89]. multi-dimensional [AJL89]. multi-gradués [Les88]. multi-grid [BM86, Fuc86]. Multi-level [BRA86e]. multi-output [GMW87]. multi-parameter [Wak87]. multi-point [Njå86]. multi-valued [ČW87]. Multidimensional [EH88b, LeP89a, Sap88, Thi86, Ark89, Col86, Il’85, Par87b, Wal86b]. multidimensionnel [DD86]. multidimensionnelles [Mét89]. Multigrid [BT86c, HN86, HT86, Jam86, Vas86, AMM86, Ban86, Bra86a, Bra86b, KM86b, Man86a, MV86, RZ86, vdW86, HT86]. Multiparameter [Vol88c, Vol88b, Vol88d, Vol88f, Vol88e, Vol88g]. multiparticle [Ens86, Ens88]. multiphase [Grü88a]. Multiple [Col88b, Kos88a, Rog89, Kat88b, jQxQ88, WKK86, Wu88a]. Multiple-point [Col88b]. multiples [HM88b, dC85c]. multiplicatif [Kah88]. multiplication [Sch88b]. multiplications [Jar85a]. Multiplicative [GS89, Jos85a, dC85a]. multiplicatively [HvL88]. multiplicités [Fer86, Ser89]. Multiplicities [Sim87, Gul86a, Kuu88, RS87, Ser89, Shi88f, Wei87s]. Multiplicity [CAXD86q, Hel86f, Laz88, Que87i]. multiplicy [Bes88]. multiplier [Ale85, AS85l]. Multipliers [BB88b, GZ89, Lio86, ST89l]. Multipoint [Njå89a, Aga86b]. multisécantes [Hir87]. Multisoliton [Sch86o, Sch86p]. multistage [MS86u]. multitype [Iva86]. multivalued [Gel88]. multivariate [Kag89a, Wer87b]. multivariées [DF86a]. Mumford [Aur89, TT88]. munies [Lev89a]. Mutual [Qua88a]. my [Erd85]. myelinated [Bri85b, Gri85b].

N [Die85b, Sta87b]. Nagell [Han85]. Nagoya [IH86]. Nakano [Dem86a]. n’appartenant [Reu86a]. Nash [Shi87c, Shi87b, Shi87a, Shi87f]. Nasu [Boy88]. Natural [MN88g, CFM86, Han88, KPT85, Tim88]. naturelle [AH89]. naturelles [Let89]. Natürlicher [KPT85]. Navier [Bem88, Pon88, SH86, Val86b, Bei88]. Nazarova [BT86d]. near [Anc87, Gos88, Hel88b, Mas88b, Sab85, Sad86, Sev86a, Sev86c, Sev86f, ST88, WKK86]. near-minimality [Mas88b]. Nearby [BP88a]. Nearest [Lig86, Woe86]. nearly [BT86c, VJ85]. nearly-poised [VJ85]. nearness [NW89]. Necessary [HJ85, Mai88, Abe87d, Abe87f]. negative [Ber88a, Bre87, DL88, DR87, Gut87, Pan86, Sch86d]. negatively [BG88b]. negativity [Eno88]. Nehari [Elb85, Koe88]. neighborhoods [Lev85b, Pin86]. neighbour [Woe86]. neighbourhood [Ulr88a]. Nelson [Car86a, Gli88, MZ85]. Néron [Pir89]. nerve [Bri85b, Gri85b]. nests [Arv85, Lar88]. Netherlands [BBT85]. nets [EPS87]. network [CHM89, Sze87, Win89]. Neumann [DMR87, BS87b, Cec88, Cec89, FB86, Hen89a, Hen89b, Jaj85a, Jaj85b, Jaj85e, Jaj88, Jaj89b, Krá89, LN88, Łuc89f, Mül87i, Ocn85a, Qua88b, Ran87]. Neutral [Nag86c, BHT89]. neutral-type [BHT89]. Nevanlinna [KQ88, Sto87b, Voj87c]. Neveu [BMS87]. Newton [Aba86, Fin88a, HRL89i, Has85, LS85a]. Newton-type [Fin88a]. Newtonian [CE88]. Newton’sche [Has85]. Nicholson [Mul86]. Nicholson-type [Mul86]. Nicolai [GR87b]. Nielsen [Fel88, HPY89, Zha89b]. Nielsen-type [HPY89]. Nikodym [Cec88, Sch87a, GM85a, Phe89a, Sch85a]. nil [Zel88]. nilmanifolds [CP88]. nilpotence [LM87]. nilpotency [Zel88]. Nilpotent [Bor87b, Eck87, HN89, Len87, Lud88, Möh87, Rob87, UW88, Wei85b]. nilpotent-extensible [Len87]. nilpotente [Sch88k]. Niveau [Wsc85e]. No [Fie88i, HM87, Cib86, KK89a, Ron86]. Nodal [Hen86a, vGW89, Cat85a, KS88b]. Noether [CAXD86q, MP86]. Noetherian [AV88b, BR86, Bjö89a, Bro85, Eks89, FT88a, Haj87, Hod89, Lec86a, Okn88b, Sta87a]. noethérien [Kra87b]. noetheriennes [Cau86a, Alf86]. noise [LM88b, Sch89a]. Nombre [Gan85e]. nombres [Alt89, Lon86]. Non [And86a, Asa86, BG89a, Cam86, Cec85, Din85a, EB88, FKV88, Jaj85f, Kle87, KS86f, MH88b, Sel88i, Sev86e, She86h, vW86, ABHK85, Ale85, App88, Asa85, Asa89, AA89, Bal85, Bar89a, Bar87c, BM88, Bar86e, BF88a, Ber85e, Bon89, Bou88d, Bri86, BBGM89, CV87, Car85c, CGC88, CMS89, Coh88a, Con85a, DD85, DL87, DP88, Del86, Dem86b, Dia88, Dol86, DG86a, Dra85, Dra88, Eke86a, Eli88, FP85, Gér86a, GG85, GvW89, GK87b, Gro88, Had85, Haj87, Har86, Hel88b, Hör87, IMW88, Jaj89a, Jos85a, JM86b, Kan86a, Kes88a, Küm88, Lai88, Le 88c, Lev85a, Lon86, Lum87, Lum89, Mas88a, MP88, MVW87c, Mor86d, Muk89, Nag86a, OP89, Par86c, Pas86, Ras87, Rog87, Roo88, Rou87, Sán85a, Sap86, Sch89d, She87c, Sjö87, TC85, CDV87]. Non-abelian [And86a, Bou88d, CGC88, Bar87c]. Non-autonomous [Sev86e, DG86a]. non-bilinear [Dol86]. non-classical [Rog87]. Non-Commutative [Jaj85f, vW86, App88, Dia88, Dra85, Dra88, GvW89, Haj87, Jaj89a, Küm88, Lon86, CDV87]. non-compact [CMS89, GK87b, Kan86a]. non-cylindriques [Lum87, Lum89]. non-definite [FP85]. non-degenerate [Hel88b, Sjö87]. Non-diffeomorphic [FKV88]. non-discreteness [BM88]. non-existence [Ale85, OP89]. non-finite [Bar89a]. non-form [Mor86d]. non-Gaussian [TC85]. Non-homogeneous [KS86f]. non-identical [Muk89]. non-integral [Asa85]. Non-linear [Cam86, Din85a, Bar86e, Coh88a, DP88, Del86, Gro88, Hör87, Lai88, MP88, Nag86a, Sán85a, Sap86]. Non-orientable [BG89a, BBGM89]. Non-oscillation [MH88b]. non-oscillatory [DL87]. non-parametric [Car85c]. non-positive [Had85]. non-random [Har86]. non-rational [Roo88]. Non-reduced [Kle87, Sel88i]. non-regular [Kes88a]. non-self-adjoint [Bri86]. Non-self-similar [EB88]. non-special [Sch89d]. non-spherical [Pas86]. non-standard [Eli88]. non-vanishing [Jos85a]. non-zero [Eke86a]. nonabelian [Sco87]. Nonaffine [Shi87f]. Nonamphicheiral [Kha88]. Nonarchimedean [Rum89f]. Nonclausal [Nua88]. Noncommutative [Bla85a, Bra86k, War86, Wod87, Bar89c, BR85c, Cec88, Hol88]. noncommuting [Bra86c]. noncompact [FH89, Fal86, Jin88b, Win85a]. nonconnective [PW85b]. nonconvex [Kiw85f, Kiw85g, Kiw85h]. noncooperative [Zam88b]. nondegenerate [LX89]. Nondifferentiable [Kiw85e, Chi85]. Nonelliptic [Bea88a]. Nonequidimensional [Siu86]. Nonexistence [BC86c]. nonholonomic [Che88c, VG88]. Nonhomogeneous [Mar88a, Wan87b]. Nonintegrable [vdM85e]. noninvertible [dP86]. nonisentropic [Wan87b]. noniterated [Poh89f]. Nonlinear [DiP87, Eva87, Hei88a, Kog86g, KH86, Pav87b, Pav87c, Pav87d, Rod87c, SS85b, Ant85, BM86, Bea87, Bel87a, BLL89, Ber87b, Ber86j, Bih85, BT86b, Bor88b, CW89, ČZ86, CM87a, Cha86a, CL89a, CK89, DS86a, Dob85, Emb87, Fei86a, FA89, GR85c, GR85e, GP85, Gol87a, Gri86a, HRL89i, HT87, Her86, Kog86f, Ler89, tLmC89, Lio89, Lun86a, Mar86b, Maw86, NR85d, Nir88, Nir89, Pan85b, RR87, Ros89a, Sch89b, Sch86c, Sim86, Tad87, VSP89, qW88, ZG88, FP86, GZ87b]. nonlinearities [Laz88]. nonlinearity [Vog87b]. nonlocal [Tan88a]. nonnegative [Kas88, Mok86]. Nonparametric [Cie88]. nonparasite [Krb86]. nonpositive [Bal86a, Ebe85, Ebe86]. nonrecursive [Hom85]. nonreflexive [Rea88]. Nonresonance [Gos88]. nonself [Wid85f]. nonself-adjoint [Wid85f]. nonselfadjoint [Liv87]. nonsimply [Ass89]. nonsimply-connected [Ass89]. nonsingular [HS88]. nonsmooth [Tho86a]. Nonstandard [Lin86b, vdB87f, vdB87e]. nonstrictly [Key87]. nonsymmetric [ES86, Man86a]. nontrivial [Ges86b]. nonwandering [CM88c]. nonzero [Sch86o, Sch86p]. Nordic [Han87j]. norm [Cie88, Mur87, Roĭ86]. Normal [Bri85a, D’A85a, Erk87, Kat86b, Kru89, Sch85i, Sev86f, Tir88, Zap88, vdM85f, CM89a, CO89, tC87, Csö89, Ger88c, Hul85, Il’85, Jaj89a, Kor87c, Lau87b, LM87, Rey86b, She86l, vdM85a]. Normale [Ves85]. normales [Car89a]. Normality [Bui86d, Il’85, Ric87, BV88i]. normalization [Pan85b]. normalize [Lau87b]. normalized [Shi86q]. Normally [Ass86b]. Normed [MS86c, MS86e, MS86d, MS86f, MS86i, MS86q, MS86s, MS86t, CG89, JLS87, Mil85d, Mil85e, MP89b]. norms [AFP88, Dav85a, Kif88a, Mor86d]. Notation [LS88m, Ral87e, Han87a]. Notations [Gek86i, Imk88d, Shi88e, Win85g]. Note [Zag88, Acc88, Ada88, Chu88a, Gul86a, Hög89, Jac89, KN88a, Kha85, Kha89, Kif88a, Kif88b, Kis85, Mey88a, NZ85, Nev89, Rou88, San86, She86i, Sta87b, Waa86]. notebook [Ber85d]. notebooks [AJBL89]. Notes [Gal89, Hoy87, MS89a, Phe89i]. Notions [HL88a, KP86a]. nought [Ara88]. nouvel [Hu88]. Nouvelles [Hir89a]. Nov [Smi85b]. November [AW86b, GWW85, Hey86]. Noyau [Aro88b, Akk88, Ank88, Bar87f, Geb86, Hmi87]. Noyaux [Ama87, Del89, Fd89, Sel87, Yan88b, Yor85a, Zah89]. nth [GL85]. nuage [Rot89]. nuclear [Ele87, KPA89]. nuclearity [D’A85a, Sio86b]. nuclei [Zah89]. nul [Bou86a]. null [AO88, Sto85j]. Number [Mor86a, Poh89e, Sun87, ACM85, Bar87d, BB88a, Ber85d, Bum85, BR85c, CC87b, CC88, Emb87, Erd85, EP87, EN87, FS88b, FGL86, Joh86, KR85, KMR89, KSX87, Kor87c, Mal85a, Mat89b, Mel89, Mor86b, Nat89a, NR88, Oza85, Rav88, Sár89, Sar85, Shi86v, Sri85, qW89, Zha89b, Zvy88, All85a, CCCN85, CCCN87]. Numbers [She86f, Alu88, Ber89, BPSTJ89, CAXD86c, CAXD86o, Ehr85, HPY89, Hei85, Hei86, Hen89a, Hen89b, HJ85, Jaj85e, Jaj89b, Kam89a, Kar85, KS86c, KPT85, KM87, KK89b, Mak88, Nic89, Num89, PR85a, Ped85, Pom85, Reu86b, She86o, Sim87, vS87b]. Numerical [Bel87a, BA89, Car85a, CL87, FV89, Kiw85i, Nat86, Rek86, DDL85, Fei86a, GMSW86, GP85, Hen86a, IT86, IW86, Pap85, Tad87, Tak85, Bre85, Hen86b, Tur85]. Numerically [AS85b]. numériques [Gan85c, Gan85d]. Numerische [Sch85g]. Numerischer [KPS87]. Nyquist [Haj89].

O. [Dem87a, Gal89]. OATE [AMSV85]. ober [KP87b, KPT85, Sch85e, Sch85h]. Obersicht [Buc85]. Oberwolfach [CGW87, EMS85, GWW85, Hey86, RR85]. objects [NR85g, Shi85]. observables [BR85b, Ens86]. observation [Coh88a]. observations [Jan86a, Mil88c, PS88c]. obstacle [ABB86]. obstacles [Bir85, Oza88]. obstruction [And86b, Fut88e, FM88, Mil85b, NR85f]. Obstructions [Che88c, Chr87]. obtained [DGR87, Fra89, PPP87]. obtenue [Ofm86b, Ofm86c, Ofm86a]. occupations [Le 88c]. occurrence [KV85]. occurring [Pik86b]. Oct [DT87, LRT85]. October [AvW85, Bal86b, BDM^{+}85, CAWXD85, HT86]. Octobre [Nor86a]. odd [Deb89]. ODE [Sha89]. ODEs [Gea89]. Oeuvre [Cho89c, Die85b]. Off [Rog85]. Off-shell [Rog85]. Ogoma [Lec86b]. Ohsawa [Dem87a]. oil [Poz88]. oil-wells [Poz88]. Oka [Fer85c]. Old [Rob88, Wal88]. Oliver [Wei85a]. ondelettes [JM89b]. Ondes [Mét89]. One [Mül87h, Nag86b, Alb88, BW88, CRT89, CB89, DGM88, Dic88, Eld89, Els85c, Els85e, Els85f, Eng86, FRN85, GS85a, Ges86b, Jen86, Kal85, Kas86, tLcZ88, Lie87, Mir85, Mül87b, Mül87g, Pap88b, Pre88e, Rei85, Sch87g, Sel88d, Sel88i, Sta85, Tan88b, Tor87f]. one-dimensional [Alb88, Eng86, Ges86b, Jen86, tLcZ88, Mir85, Pap88b, Sel88i, Tan88b]. One-parameter [Nag86b, FRN85]. one-relator [CRT89, Dic88]. one-sided [Kas86, Lie87, Sch87g]. ones [KS85c]. only [Kor88, She86j]. onto [Jar85a, Koo85]. onto-isomorphisms [Jar85a]. Ootacamund [All85a]. Open [Kaw88, Tro88, Bla87, Gro85, Lig86, Lig85a, NR85e, Sac85, She86o, Shi86w, Siu85b, Sta85]. open-loop [NR85e]. Operad [LMS86a]. opérateur [Han89, Mey86c, Wu87c]. Operateurs [Lev89a, Dao89, Deb87, Mal88, dC86]. operation [BL85, FT86a]. operations [ACN85, Dwi85, FFR88, KI87, McC86c, Mùi85, Sch87e, Shi86i, Shi86o, Ste86a]. Operator [Wei86e, Wer87b, AGHK^{+}86, Ati85b, Bir85, Bor88a, Brü85, Dau88c, Dop85, Dzi88, FS89a, Fuk86b, FK87, Går87d, GESEZ85, Gar88c, Gro86c, Gru86a, Hil85b, Jar85d, LMZ86b, MB85, Nis86, RT89, RS85b, Sai86, iS88, Wat88, Wei87b, Wei87n, Wih85, Wit88b, AMSV85]. Operators [Bal86b, CGW87, Els85h, FS89b, Gra85a, Nag86b, Pav87c, Wei87c, Wei87u, Wid85b, vEdG85g, AB89, Agm85, AS85e, AB87, Bal86c, BDT87a, BCZ87, BBP88, Bor88b, BPSTJ89, BDE85, Bri86, Brü87c, BDG88, Car85a, Car86c, CFM86, CS89b, Chi85, CDMS89, CBD87, CH89b, DR87, DMR87, Dho85, Din85b, Dub88, Ele87, Els85b, Els85c, Els85e, Els85f, Els85g, Fuj87, Går87f, Går87g, GR85e, Gar86b, Gel88, GG85, Gil89a, Gou88b, Hel88g, Hel88h, Hin87, Hud89, Iwa87, Jaj89a, JM85, JSC87, Kha85, Khi86, Kir85a, Kir86, Kir87a, Kos88c, Krá88a, Kuo86, Liv87, Mac87, Mül87a, Mur87, Nag87, NS86a, Nei85, Ngu85, Pau86, Pav87b, Phe89f, Phe89h, jQxQ88, Que87k, Rei89h, Rei89g, Rei89j, Sch86a, Sch87c, Sch87l, Sim85a, Smi86c, Smi86b, Smi87, SS88c, Vol88b, Vol88d]. operators [Vol88e, Vol88g, Wei87e, Wei87h, Wei87i, Wei87o, Wei87p, Wei86g, Yan88d]. optical [Pik86b]. Optimal [Col88e, GESEZ85, Kog86b, Kog86h, Lio85, Bar85a, DM87a, Fla87, Hau85, Hop87, Hop86, Lio89, MM86, MR85c, MW86, Mos85, NR85d, NR85e, OD85b, Tvr86, Vin85]. optimality [Cam86]. Optimization [Col88f, GQS85, HM88a, Kiw85e, MS86u, Alt88, GJN88, Law89, LM89b, Phe89a, Zol86, CDG86]. oracles [HZ85, Yam89]. orbit [HS88, KK89b, LÖ86, MS86a, Shi86m]. Orbital [Orl87]. orbits [ABHK85, Bor87b, Jaw86, Kat88a, Mat89b, qW88]. Order [BB88c, HRL89g, Win85b, AEE85, Aro88a, BB87, Bar87c, Bar86e, BDT87a, Bih85, Böh86, BDG88, CAXD86k, CAXD86l, DR87, DL89, Des85, Dra85, ENS87, Går87d, Gil89a, Gos88, GL85, Gre86a, Hem86, JSC87, JM86a, Kar85, Kos85, Lab86a, Las86b, LS88b, LX89, Lio89, LT88, Mas85, MN88a, NV89, Nir88, Pan85b, Ril89, Sch85c, Spe86, Tur89, Wan86, Wan89, Wil88, Win85c]. ordered [Hir85]. ordering [Ler85]. Orders [DGM86b, Brz85, CW85, DV86, Frö85, HvL88, Rog86, Rum85, Rum86, Sal85b, VS87a, Wie86, RR85]. Ordinal [Poh89f, Poh89e, Poh89a]. ordinals [Bau89, Cho85b, Fri85a, Jac88a]. Ordinary [Els85h, Neu86b, Wei87u, Ado87, Ber87b, BDG88, CQ85, Col88e, Els85f, Els85g, IN86h, IN86m, KJ86, Lew88b, Lit88, MH88a, MH88c, Ste86b, LV86b, SJ85]. ordinization [Day85]. ordre [BD89a]. Ore [Bro85]. organisées [FHMM86]. orientable [BG89a, BBGM89]. oriented [Cib86, Str86b, Tur88a, dP86]. Orlicz [AS85j, MP85, Por85]. Ornstein [Had85]. Orthogonal [CM88b, Dra85, EM88, Hen89a, Lit88, Njå88, Van87b, Van87d, Van87e, AA85, Ask89, Bou89a, CM88a, Gau85, Ger88b, Hen89b, HL88b, IN85, Koo85, Koo88, Kra88b, Las89, Lev89f, LR88a, Lóp88, Mag88, MM85, Mas88b, MN88a, Nik87, Ral87h, Vie85]. orthogonal-series [Nik87]. Orthogonality [CCS^{+}87, Kow85, dB85]. Orthogonalpolynome [Hah85]. Orthogonaux [BDM^{+}85, AAGV85, Die85b, Mar88c]. Orthonormal [LS88f]. orthonormalization [Pap85]. orthosymplectiques [Cru85b, Cru87]. Oscillation [Bou86a, Wei87o, Wei87p, Ber88b, MH88b]. oscillations [Arn88a]. Oscillator [Ral87c, For85, For88a, IN86n]. oscillatory [Bar86e, DL87, DZ85, Eng87b, Går87b, Går87h, Går87k, GL85]. oscillatory-equations [DZ85]. other [Koo85, MN88c, Men89, Mil88c, Shi86r]. OTOP [Har87c]. Ottawa [DGM86a, DGM86b]. outfix [KS88a]. outgoing [Ens88]. output [Bar88a, GMW87]. ouvert [Mou85]. ouverts [Loe86, Ofm86e, de 89]. Oval [JT86]. overrings [Gla89g]. Overview [Kog86i, Hai85, Lan88c].

P. [Car86d, Par86c]. P.-A. [Par86c]. P.I. [Jøn85, PR89a]. PA [Kuč85]. packing [Bou87e, Tay86]. Padé [AK87, Bre88, CC85a, CC85b, Dra87d, Dra88, McC89, Njå89a, Njå86, Sta87d, Wer87b, dB85]. Padé-Operator [Wer87b]. Padé-type [Dra87d]. PADF [Fin88a]. Painlevé [IN86f, IN86g, IN86n, IN86o, IN86j, IN86l, IN86k]. paired [Går87h]. Paires [MVW87c]. pairing [Beĭ87a]. pairs [Bat87, Kog86c, ŁR85, Sch85i]. Paley [BL86, Mey85a, Bak85c, Bou89b]. palindroms [BT86a]. Paneitz [DHS85]. paper [HS86, Kát85a, Par86c]. parabola [Waa89]. Parabolic [DM86, AT86, Bel87a, Di 86, Gar88b, GS86, Lan88a, Li88, Mar89c, Mor89, NW86, SJ86, Tho86a, Tit87, Wan88]. paraboliques [Lum87]. paradox [Coh85]. Parallel [Kan87, App88, Gli86, MV86, NS89b, Pet89c, Xia89, lY87]. Parallelism [Bel89b, Sil88a]. Parameter [KP87b, Pro85, Wih86, Bar89d, CF88, Fla87, FRN85, Grü88a, Imk88a, Imk88f, KL87c, Nag86b, Sad86, Veg85, Wak87, Wal86b]. parameters [KMM87, Kör87b, Kre86c, MH88b, Oko85]. paramètre [DD86]. paramètres [Wsc85a]. Parametric [GJN88, Car85c]. Paramétrix [Rab87, FK87, Går87d]. Parametrization [IN86j]. Parametrized [Nak89, Str87a, Thi86]. parasitic [Had86]. parentheses [BT86a]. Pareto [Cec86]. parfaits [CL85]. Paris [JY85, Mal85c, Mal86b, HM87]. Park [AH85a]. Parry [Pre88g]. part [Don88b, Dum85b, Pro87, Sch86q, Sch86r, Jac86a, LeP89a, Lio89, Rau85]. Partial [DT87, Gek86j, Gia86, GZ87b, Kag85, Kaz87, Pav87c, SJ85, Ado87, Bel87a, Bel89a, Ber87b, Bre86b, CV87, CM89b, Ehr85, Eng87b, Gia88a, GMW87, Gyö89, HvL88, Hen86a, Her88b, Jac88c, Kre86b, Kuo89, Lie87, Lit88, Lja88, Már89a, MV88, Oco87, Pav87a, PS85, Roz87, Tah86, Tal86b, Tor87b, Waa86, Wal86a, Zha88]. partially [Grü88b]. particle [AGHK^{+}86, Dur86f, Gre86f, Lig86, Pre87, Rav85, Ros86c]. particles [LS85b]. Particular [Gla89d, tW87a]. particuliers [Sal89]. partie [Bak85a, Pas87a]. Partieller [Vog87a]. partielles [Bou86b]. parties [Deb87]. Partition [Clo85b, Bau89, Fri85a, She89a, Shi86n, Yor85c]. partitioning [Arb89, Vas86]. Partitions [Aga86a, Dür86d, ESN89, Kat85b, Sta86b, SS85d]. parts [ESN89]. Pass [Cha86b]. Passage [Kör87b, Had85, Kes86]. past [Jam89, Rea86]. Patch [Kne89c]. path [Car86a]. Pathfollowing [GJN88]. Pathology [Neu89]. paths [Bre89a, Kis86, Lin86b, MP85]. patrie [Bak85b]. patterning [GS85b]. Paul [Mal85c, Mal86b]. Paule [Mal85c, Mal86b]. PDE [Eva87, Kaw85d]. PDE’s [Roc88]. peculiar [Dub85, Dur86f]. penalty [Man86b]. 1984 [LDS86]. 83 [JY85]. 84 [AY85, Ław85]. 85 [AY86]. low [Ler85]. SIP [vdW86]. pencil [PS89a]. Penetration [Fit88]. Peñíscola [NFM86]. Percolation [Gri86b, Hig87, Kes86]. perfect [BV88g, Con89, LF89]. performance [Kog86e]. Périmètres [Wsc85d]. period [Kno89, Var88b]. Periodic [Bih85, Coo86a, DL86a, Gil88, Kre86b, Prü86, Rab88, Afa85, BF88b, Chu88a, Col88e, Dho85, Fen89, GU87b, HPY89, Hel88h, HS87e, HjJ89, Jac86a, KL86, KH86, Mat89b, Mil85b, MB85, Par87b, Son85, Van87a, Wei87c, Wei87h, dBJ87]. periodical [AK87]. periodicity [Sei89a]. périodiques [Lim85]. periods [Sch88g, Tre87]. permutation [Dvi85, Pir87]. permutational [Lon89]. Permutations [CS89g, Wan86]. Perron [JNT86, JNT87, KJ86]. Persistent [GMM88]. personal [cH89a]. perspective [cH89a]. Perturbation [Maa89, NR85c, Ben85b, BCM86, Gra86a, tLcZ88, NR85b, Yan88d]. Perturbations [Jar85e, Mez86, Bol88, Chi85, Lin86b, Sch87c, Sev86d, Tre86, Jar85d, Jar85a]. perturbative [Gyö85, Seg85b]. perturbed [HS87e, Men88b, NR85a, NR85d, Sch86c]. perverse [Beĭ87c, Beĭ87b]. petit [Léa87, MVW87d]. petits [Gan85c]. Pettis [AS85j, Dre86]. pf [Her88a]. phase [BR85b, MMA86, PR85b, Sch86o, Sch86p, Sle87]. phenomena [Pre87]. Phénomène [MS87b]. Phillips [AS85k]. Physical [Fri88c, Kon85f, Bér86f, Bra87, FKZ86, Gyö85]. Physics [ACSC85, ABS86, ABS87, DH85b, GPR87, GPS87b, LS86a, AL88a, GRT88, Mit86, Sta88b, Sta88a, Wil86b, vK87]. physiology [RB87]. Picard [Dre87, Hol89c, Nev89, Tre87]. Piecewise [Lev89j, Pre88f, Che86a, Lev89b, Lev89d, Maz86]. pilgerschritt [FR85]. pinched [KS86e]. pinching [Kat86a]. Pisier [MS86o]. Pitlochry [Thr86]. PL [DH89, Lev89c]. plan [Dem86b]. planaires [Arq86]. planar [Bus88a, Kre85]. Plancherel [PV87]. Planck [HSS85, Cru85a]. Plane [CO89, Gud88, LP88g, Str86d, Wan87g, Alu88, AS88, Cat85a, GS85a, Her88c, KS88b, Kor88, Lev85e, Lev85f, Lon88, MD89, Pic85, Ros89c, Sch87b, Sta86b, Wal86b, Wan87c, Wan87e, Web88]. planes [Hir87]. Planetesimal [DT88a]. plasma [AP89]. plasmas [Wei89b]. plate [Fil85, Sad86]. Plateau [Bor88a, Die88a]. pluriharmonic [DP85]. Plurisubharmonic [FS87]. plus [Mag86]. plusieures [Fug86]. Plusieurs [Nor86a, Bia89, LS85a, Roo86a]. PN [Lev85b]. Poincaré [Ani88, Ass86b, Dix85, EH88a, FT88a, FGL86, Gul86b, HV88, Hen85, Kas87, Kri85a, Les88, Mas85, Mc 87, Pan85a, Roo88, Le 86b, Asa86, Reu86b]. Poincaré-series [Kri85a]. Point [Coh86, AGHK^{+}86, Bér86f, BG89b, Bli86, ČW87, Col88b, CL89b, Dra87d, Dre85, Els85c, Els85e, Els85f, EM86, Fad89, Fel88, FZ85, FF89, He89, Hög89, KM88, Kat88b, Kes88a, Kiw85c, Kom89a, Kor88, Kos88a, Lab86a, Maw86, Njå86, Not88, PlT88a, PlT88d, Pap88a, Pre88e, Rad87, Tho88, Ulr88b, Wei87m, WST88, Yp89]. point-limit [Wei87m]. Point-line [Coh86]. pointclasses [Bec88]. pointées [Arq86]. Points [MM86, Arn88a, Bah88, Bor89, BGS89, BV88f, Cil87, Els85g, Fer88c, Fin88a, GH85b, Hah89, HPY89, KR85, Kog86d, Mag86, MK88, Mor86d, Mum88p, PlT88e, Rea86, Rei87a, Rog89, RS85b, Sar85, Sch88a, Shi86w, Sil89i, Sjö87, Str86c, Tim88, Vig86, Voj87f, Wai89, WKK86, Zag85]. pointwise [Bou88a, CM88c]. poised [VJ85]. Poisson [Akk88, BW85b, BGR85, FL86a, Fri88c, Hmi87, Kni86, Wu87c, Wu87d, dC85b]. polaires [Le 88b, Mag86]. Polar [SZ87, Web88]. polarities [Dol86]. polarized [CKS87b, Kas87]. polaron [AO85]. polemical [May86c]. poles [Bar86b, IN86n, IN86o, Kov87, Njå88, Sta87c]. Polstellen [Wei86l]. polycycle [Mou88]. polycyclic [Rob89]. polydiscs [LM86a]. Polyeder [Buc85]. polygonal [Blu85a]. polygons [Kan86c]. polyheadral [BS89f]. polyhedra [Buc85]. polyhedral [Dau88h, KP85]. polymer [Sto87a]. polynôme [Bar87b]. Polynômes [Str85a, Val86a, AAGV85, Big85, Cau88, Die85b, Las86a, Mar85, Mar88c, Tor87g, BDM^{+}85]. Polynomial [Ale87, ES86, Gia87, Gla89e, LS88n, Ple87, ASV88, Arb89, BF85, BDG88, Cie88, Hom85, KL87b, Koo85, LM88a, Lit88, Mas88b, McC89, Net88a, PPP87, Saf88, Wer87a, Woo87b, qW89]. Polynomials [Ger88b, Van87b, AA85, Ask89, BHS88, BG88a, Ber88a, BD88, CM88a, CM88b, Dra85, EM88, Esp88, Fri88b, Gau85, Gro87c, GR85k, Has85, HL88b, IN85, ILV88, JNT86, Koo88, Kow85, Kra88b, Lan88d, Las89, LS87c, LR88a, Lóp88, LS87f, LS88c, LS88d, LS88e, LS88f, LS88k, LS88p, LS88l, LS88r, Mag88, MM85, MN88a, NS85, Nak89, PH89, Pet89c, She87a, Shi86f, Van87d, Van87e, Van85, Vie85, Yui88, dG85, dSCV85]. polytopes [Dre85]. Poncelet [BT87a]. pont [BLY87]. Pontrjagin [Sto85f, Kaw86a, KK89b, Mei87]. Pontryagin [Lis87, Sto85f]. population [Don86c, Kim86, Smí86a]. porous [Aro86]. portrait [Fom88a]. posed [Nat86]. posets [BV88a]. position [MP89b]. Positive [Anc88, Don86b, FM89, Hel89, JW89, Nag86b, Pet89h, Swe89, Van85, Are86a, Are86b, AB87, Bal85, DD85, Dem88, FQ85, Gre86b, Gre86c, Gre86d, Gre86e, Gro86b, Gro86d, Gro86e, Gul88b, Had85, Hol88, Lac86, NS87, NS86a, Nis86, Pet89a, She89a, Siu87, Swe85, YM89]. Positively [RVV85, Xia87]. positivité [dC86]. Positivity [KN88b, Led86b]. possible [Koź86]. posteriori [Ban86, Kod86]. Postnikov [Ass85]. potato [Sch86b]. Potential [LMZ86j, Wil88, BH89b, BB88c, Chu88b, Fab88, Fug88a, GR87a, HN86, Han88, Har86, Hel88b, Krá88a, Lai88, Lan88a, LMZ86i, PS89a, Roa88, Ros85b, Sch86q, Sch86r, Sha87, vdW86]. potential-theoretic [Krá88a]. potentialistes [Tor87g]. potentials [AB87, CBD87, Cra86a, DMR87, Ens85, Ens88, FW89, GU87b, Hel88h, Kan88b, Kir86, LS88n, Yaj85b]. Potentiel [HM87, Bar87f, Geb86]. Potentiels [de 89, Del89, Maa87, dlP87b, dlP89]. Pour [Ibr85, Jac85a, Alt89, Bak85a, Bak85b, Bar86d, Bar87b, Ber87d, Bou86b, Cle88, Deb89, DvdB87, DD86, Fer86, Gai88a, Han89, Hir87, Le 88b, Lim85, Lon86, Lum87, SY87, Vig89, dC85b]. Power [Dür86c, McC86c, LR88b, Meh85, Rei85, Sch85i, She86m, Shi86v, Wie87]. power-sum [LR88b]. powered [Dun87]. Powers [ASH85, BV88k, Dem88, Gru87, LMS86b, May86a, MS86b, SS89b]. Poznań [JP86]. practical [Ste85a]. PRC [Pre85]. PRC-fields [Pre85]. pre [AS89b]. pre-inclinées [AS89b]. Preamble [May86c]. Precise [HGYR89, Kus86b]. preconditioners [Bra86d]. Precursors [CS89h]. predicate [GM88, She86b]. predicativity [Poh89a]. Predictable [KM86a]. preduals [Gro86e]. Prefix [BT86a]. Prefix-free [BT86a]. prehomogeneous [Che86d]. Preinjective [Ung86]. Préliminaires [Gan85f]. Preliminaries [BS89g, BV88l, Bui86e, Dau88d, Fut88i, Gla89f, Jar85f, KP86j, MS86i, MS86p, Mül87j, PlT88i, Rei89c, Rum89g, Shi87g, Sil89d, Ulr88d, vdM85g, Ral87e]. Preliminary [Har87d, Oll85h, BC86d]. Première [Bak85a]. premiers [Ler86, Mal86a]. preparation [Cha88]. Preprocessing [BS89h]. preprojective [Vos86a]. preprojectives [Mer86]. Presburger [Har87b]. prescheme [Mum88p]. preschemes [Mum88b, Mum88x]. prescribed [Die88a, Kni87, Kos88c, Rab88, Sel88e, Sta87c]. presence [GS85b]. present [Jam89]. Presentability [Abe87c, Abe87a, Abe87f]. presentable [ART88]. Presentation [HLM85b, Den86]. presentations [CRW89, MO88]. presented [AL86, Lal88]. preserving [CP89b, Jaw86, Sil88a, vdM85a]. presheaves [GM88]. presque [Lim85]. presque-périodiques [Lim85]. preuve [Dem87b]. prevarieties [Mum88h]. prévisible [AH89]. primal [Gol86b]. primality [Len85]. Primary [BV88m]. Prime [Ble85a, FT86b, Ber89, Kat85b, Ste89a]. Primitive [Row89, Bor87b]. Principal [BV88n, Pix85, Rey86b, AG89, CP89b, Frö85, Hri87, LR89a, Ma89, Wil86a]. principale [Pas87a]. principales [Hir87]. principe [Bou87a, Cru87, Hud86]. Principes [Lum87, Bar87f]. Principle [AS85m, Arv85, Bar85a, CS89b, Die88a, Fuk86b, Ich87, Lew88a, MPS87, Oll85c, Phe89j, Saf88, Sar85, Shn86]. principles [DL86b, Fin88b, Kaw85b]. Priori [DZ85, Lap88]. priority [Sla85a, SW89b]. probabiliste [AD85b, Kôn85k]. Probabilistic [Alo89, HZ85, Pis86, Van87f, Win85b, AD85b, Rad87, Zam88a]. probabilité [GG84]. Probabilités [AIJH85, AY85, AY86, AYM87, FHMM86, Hen86c, Yan88a, Gal89, Mey86a, Mey87a, Mey89b, Mey89c, Mey89d, Yan88b]. probabilities [DM89a, Mañ88, Pin86]. Probability [AvW85, AIJH85, Bis86, Chu88b, FHMM86, Gal89, Jaj85f, Kon85g, Oco87, dPR86, BL89a, Bin89a, Bin89b, Brü87b, Dia88, Fri89, KT87, Mey87b, Nik87, Par89, Pet88, Rac85, SW86a, BS86, BDH^{+}85, Hey86, LP86]. problem [AMM86, AFL89, ABB86, Ani86, BW85b, BS87b, Bla87, BFKP86, BPSTJ89, Bou89a, BLR87, BN88, CT88, Car86b, CAXD86a, ČZ86, CHM89, CG89, Die88a, Don88a, DLS85, Emb87, Fla87, GR85e, Gos88, GMW87, Gre86a, Gro85, Grü88b, Had85, Han87h, HT87, Her87a, HJ87b, Hop87, HSW85, Hu87, Ieş87c, Ieş87d, IN86b, IN86a, IN86e, IN86i, IN86l, Jin88b, KT85a, Kat86a, Kaw86b, Kes88b, Kes88c, KMR89, Kob87, Koi87, Lab86a, LS88a, tLcZ88, tLmC89, Lig85a, LN88, LMZ86h, Mar88a, McG88b, Men88b, Miz88, MY87b, NR85d, NR85e, Njå86, Pie85, Pie89a, Pie89d, Pie89e, Pie89f, Pon88, Prü86, Ran87, Rea88, Rod87d, Sad86, Sai86, San85b, Sei89b, Ste89b, SU87, Szé86, Tan88a, Wit88a, Wri89a, Wri89b, Wu88a, YM89, Zab87, Zha88, tDP89, vdG85, vdM85h, CQ85, Rou88]. Problème [Ben87, BJO89b, Krá89, Pol86, Aub86, Hau85, Let89, MM86, Mou85, Rot89, KPT85]. problèmes [Nic87]. Problems [Bér86g, Dop85, ESS85, EN87, FP86, Fer88b, KZ85b, KPZ87, Kan88b, Mal85b, Sch86i, Tal86a, Wai89, Aga86b, AL88a, ABHK85, Ant85, Arn88b, Att86, BO85, Bal89, Bar86a, Bau85, Bea88a, Bem88, Ben85b, Ber86k, Blu85b, Bra86d, BRA86e, BT86c, BCM86, Bro86, Brü85, But88, Caf89, Cec86, CC89b, Col86, Coo86a, Coo86b, Dau88i, Dau88h, DT89, Dia88, Dob85, Dur85a, Eke88, EM86, Erd85, Erk87, Fab88, Fei86a, FP85, Fle85b, FS88b, Fri86, Fu88, Gla85, GP85, Gri86a, Gri86c, Gru86a, Gru87, Grü86b, Gul88a, HRL89a, HRL89b, HRL89c, HRL89d, HW88, Har88a, Har87d, Hel87, HM88a, Ich86, Ieş87e, Joh89b, Kaw88, KN88b, KPT85, Kiw85c, Kiw85f, KL87b, Kod86, Kog86d, Kog86c, Kog86h, Kra88b, Kuč86a, Kuf86, Lan85a, Las86b, Laz88, LR89b, tLqL87]. problems [Li88, Lie88, Lig86, LM89b, Man86a, Mar86b, Mas88a, MPST85, Mas85, Mei86a, MR85b, MW86, MP88, Mos85, Nag86a, NW86, Nat89a, Nat86, Nir89, Nit86, NW89, PW85a, Pie89b, Pri89, Rac88, Ran86, Rek86, Roa88, Roz87, RM87, Sap86, Sár89, Sch85c, Sch89b, Sch86c, Sch85i, Šed86, Sho85, Sim85c, Sim89, Siu85b, Sko87, Smí86a, Sta86a, Sta85, Ste86b, Str88b, Šve86, Tal89, Thi86, Tho86a, Tro88, Tru86, Vas86, Vin85, Vol88c, Vol88b, Vol88d, qW88, Whi86, Won87, Žen86, ZG88, Zvy88, vW89]. procedure [Rug89]. procedures [Mil85e]. Proceedings [AvW85, ACSC85, ABS86, ABS87, AH85a, All85a, AMSV85, AW86b, BS86, BDH^{+}85, BBT85, BDM^{+}85, CDV87, CAWXD85, CKGP87, Com85, CDG86, CGW87, DT87, Di 85, DGM86a, DGM86b, DH85b, DFZ85, EMS85, FO86, FGHS85, GPR87, GPS87b, GZ87b, Gra86b, GWW85, GBB87, HT86, Han87j, Hen86b, Hey86, HSS85, IH86, JP86, KZ85b, KPZ87, KS85a, LL86, Ław85, LRT85, Mal85c, Mal86b, NFM86, RLQ85, RR85, Rol85, Roo86b, SSS86, SJ85, Smi85b, Tau86c, Thr86, Tua86, Tur85, VZ86, vO86, AY85, AY86]. process [Afa85, Kas86, LB86, MRC89, ŌT87, Sch89a, Tan88b]. Processes [ABS86, ABS87, IH86, Tau86c, ABGS85, AHK85, BL85, Ber85c, BR85b, BGR85, BCSSC87, Bos86, Daw86a, Don86c, DL86b, Ede86, EM85, Faj85, FS89a, FS89b, Föl88, Fri85b, Fri88c, Fuk88b, Fun86, GZ86a, Göt86, Gri86b, Gyö85, Hab88, Had86, Had85, He89, Hol89a, Ich86, Ike88, Imk88a, Imk88f, JL85, KV85, KM88, Kis86, Kis89, KÜ88, Küm88, Las89, Le 88c, LeP89a, MP85, Már89a, MC89, Mas86a, Mét87, MS86u, Nag86a, Ōsh87, Plu89, PS87c, Rao89, Rog89, Str86c, TC85, Tău86a, UW88, vK87, vdBL88]. processors [MV86]. Processus [Jac85a, AY87, Ber87d, BLY87, BL86, DD86, FL86a, Jac85b, Léa87, MZ85, Pic86, SY87, Wsc85a, dC85b, dC88]. Product [FRN85, LC85f, RS88, CW85, D’A85a, Eva85b, Fef88, Fuj87, GGL87, HJ89, HK87b, Ill86, Kur87, KJ86, Lev85b, LC85j, Mas89, Mor86f, Ral87b, Yui85]. Product-integration [FRN85]. production [Rof85]. Products [AH85b, Don85n, Mum88r, ASH85, CRT89, Eke86a, Ele87, FK88, GS89, HO85, HLM85b, JS88, KMR89, Lau86, LM86c, LMS86b, Liu88, Muk89, Phi87h, Que87g, Wal86c, Woe86, dSCV85, vEdG85c]. produit [Fer85a]. produits [Bou86a]. Produkt [Hla85b]. program [Bar87d]. Programm [Bar87d]. Programming [DFZ85, Bar85a, BT86b, FV89, GMSW86, Man86b, Roo89, Via89, Vin85]. programs [Gol86b, Jar89]. progress [Ani88, Cla86, Kra87a, Rog88]. progressing [BM88, MR89b]. projecteurs [Der88]. projectif [Ofm86e]. Projection [BL88, MS86r, Cha85a, Chr89, LC85i]. Projections [MS86q, Cha85a, Coh87c, FL87, LC85k, Lig85b, PPP87]. Projective [Han87g, Mat86a, Mat86d, Oli85, AM88e, BR89, BM89b, Cae88, CAXD86a, Cil87, Eke86a, ES87a, FV89, GL87, Gud88, Hat86, HR88b, Hop89b, Jac88a, Jag85, Lew88b, Mas86b, RS85a, SZ85, Smi86c, Som86, sT89a, Via89]. projectivity [DR89a]. Prologue [LMZ86p, vEdG85f]. prolongement [Yge87]. Proof [LE85, LS88o, LS88p, Pre88h, Wei87d, Wid85e, Wid85f, Wid85g, Anc87, Bin89a, Cha85b, Dav87d, DKL88, Deb89, EN88, ÉY87, Hai85, Lec86b, Pre88g, Sta87d, TT88, Wri89a]. Proofs [Dau88e, Alo89, Mur88b, Van87f]. Propagation [Går87j, Lau87a, Ant85, Går87i]. Proper [CS87b, Her87a, Kam85, MS86a, Mum88s, PlT88j, BB86, CAXD86a]. properness [Col88h, Col88j]. Properties [MS86t, NS87, PPP87, Rei89d, Sab85, AK87, Ara85, Avr87, Bae86, Bar86e, Bar85d, Bec88, Car86a, Coe87, CBD87, Dur86f, Dvi85, ESS85, GR85f, GKL89, GS89, GU87b, Haj87, Haw87, HPS87f, Kal86a, Kan86b, Kos88c, Lan88a, Lev85i, LS88g, LMZ86d, LMZ86m, MN88c, MW86, MvR88, NS89a, Obr85, Phi87b, Pir87, Plu89, Qua88b, Shi86r, Shi89, SS85a, Val86b, Waa87, Wei87k, Wei87s, Wer87b, Zwi89]. property [BF85, Ceg89, GM85a, Gru87, Har86, Hem85, Joh89b, KR85, Lon89, LMZ86o, Mad89, Met85, Okn88b, ŌT87, Phe89a, Rou88, Sch85a, Sch87a, Sch87c, Shi88b]. proportional [Mil85d]. proportionality [Sha89]. propos [Der89, MY87a, Str88a]. propres [Han89]. propriété [AH89, Bak88]. Propriétés [BH86b, Cho87a, Lim85, Var86a, Geb86, Zar85]. propulsive [Bel85a]. prorepresenting [LP88h]. proving [McF88, tW87a]. Proximinality [LC85l]. proximity [LC85g]. Prüfer [BF85]. Prym [Kan85]. Pseudo [Cho87a, Loe86, AC85, Blo85, DD85, EM85, Gru87, Iwa87, Lon85, MR89a, Mur87, Nag87, Rec85, Sch87l, CGW87]. pseudo-Anosov [Lon85]. pseudo-complemented [AC85]. Pseudo-Convexe [Cho87a]. Pseudo-convexité [Loe86]. pseudo-differential [Gru87, Iwa87, Mur87, Nag87, Sch87l, CGW87]. pseudo-isotopies [MR89a]. pseudo-processes [EM85]. pseudo-Riemannian [Blo85, DD85, Rec85]. pseudoconvex [AS85b, GK87b]. Pseudodifferential [Els85h, Går87g, Wid85b, DR87, Els85a, Els85b, Els85d, Går87d]. Pseudofunctors [Bou88d]. pseudogroups [Sil88a]. pseudovarieties [Alm88]. pullback [Got88]. pulses [BPW86]. punctual [Iar85]. Pure [Poh89e, IN86k]. purement [Mal86a].

Quadratic [Bon85, Giu88, HS89a, Imk88e, Kur87, CHM89, FT87b, GMSW86, Gol86b, Gul86b, Hel89, Kat85b, Kaz89, LM86b, Qua88a, Rei89d, Rei89f, Sel88c, Sel88j, Tre87]. quadratique [Let89]. quadrature [Riv88]. quadrics [BEH87, Deb89, EEK88, EPS87, Her87b]. quadriques [Deb89]. Qualitative [Fom88b, Her86, Val86b, Daw86a, Fer88a, Lan88a]. Quand [Lin86a, AD85b]. quantifier [Kak89]. quantifiers [Cai85, HZ85, She86c]. quantiques [Mey86a, Mey87a, Mey89b, Mey89c, Mey89d, Yan88a, Yan88b]. quantitatives [Léa88]. quantization [BM89c, Duv87, Gli88, LL85, Lic85]. quantized [Seg85b]. Quantum [AW89, AS89a, ABGS85, ACSC85, App88, Bra87, Ens85, Fri88c, Hud89, Kon85g, Par86b, Sew87, Str86c, Bar88a, BL89a, Bel85b, BR85b, BGR85, Bes85, BB85, BCSSC85, CLS85, Cas85a, Cas85b, D’A85a, Dop85, Ens86, Ens88, EH88b, For85, For88a, Fri85b, Fri89, FA89, GR87b, Hud88, JL85, Lin88a, Lun86b, PS88a, Par89, Pet88, Qua88a, RV89, vEdG85e, vW85, AvW85]. quartic [Gud88]. quartics [Yui88]. Quasi [App85, Cam88, Eva85b, LL85, LMZ86q, Mum88t, Paw85, AS88, CS88c, DR88a, Hei88b, Joh89a, Kes88c, Wan86, Yaj85b]. Quasi-*-algebras [LL85]. quasi-Banach [CS88c, DR88a]. quasi-classical [Yaj85b]. Quasi-coherent [Mum88t]. quasi-equitype [Wan86]. Quasi-free [App85]. Quasi-homogeneous [Cam88]. quasi-interior [Kes88c]. quasi-linear [Joh89a]. quasi-minimal [Hei88b]. quasi-Möbius [AS88]. Quasi-product [Eva85b]. Quasi-regular [Paw85]. Quasi-topological [LMZ86q]. Quasiclassical [IN86p]. Quasiconformal [AS88, Pan86, Fer88c]. Quasiconvexity [Gia88a]. quasigroups [PS85]. Quasihomogene [Bar85b]. quasihomogènes [Ber85e]. quasilinear [Alb85, Eng89, tLqL87, Mar88a, Mei88, SJ86, Yan88e, YM89, Zha88]. Quasimartingales [Mey88b]. quasimonotone [Mar89c]. quasinormed [RTVW88]. Quasiregular [Vuo88d]. quaternion [Hej85]. quaternionic [Sul87]. Quaternions [Kri85d, Vig89]. Quatre [And87b]. que [Mag86]. Québec [LL86]. queer [LS88b]. quelconque [Kré88, Lav85, SY87, Yor85b, Yor85d]. Quelques [Del85, Dix85, Geb86, Mey86b, Pas87b, Szp86, Vig86, Zar85, Mar85, Sal89]. questions [Pug86, Sac85]. queueing [Sze87]. qui [BLY87, Mar85, Reu86a]. Quillen [VP88]. quintic [Aur89]. quintics [vGW89]. quiver [Cib86, Die87, Hes86]. quivers [Ben86a, Bre86a, IT86, Sch86f]. Quotient [MS86d, Ced88, Chr87, Lic86, Wun87]. quotients [BKT89, Her88a, KI87, LM86a, Par88a, RS88, Web88].

R. [Lev89g, Sal89]. r.e [FS85a]. r.i [Ray88]. rôle [Yor85d]. Rademacher [CG89, LM86b, MS86r]. radiales [Deb87]. radially [Bea87, Eng87a]. radiation [ABC^{+}85, Fil85]. radical [BV88n, FHTT86, Hal86]. radius [Car85a, KS86e]. Radon [Brü87b, Cec88, GM85a, Ofm86e, Phe89a, Ric86, Sch85a, Sch87a, SS88c]. Radon-transform [Ric86]. Raising [Gar86b, Shi86o]. Rajeev [Mey88c]. Ramachandra [Kát85a]. Ramanujan [And89, AJBL89, Ber85d, Ber89, BA85, Bre89a, Erd89, Han85]. ramification [CW85, Voj87g]. ramifiées [MVW87c]. Ramsey [Hin89, MS86s, Sim85d]. Rand [Kos86a]. Ränder [Sto85l]. Random [Car86c, Föl88, Jur85, Kir85a, Mil85d, MS86q, PX87, Tău86b, AW89, Bor85, BM85, Coh88a, Del86, DG86b, Fac89, Fer85b, FNT87, Ger86b, GH85a, Gor88, Gor86, Gre86f, Han87h, Har86, Iva86, Jur89, KZ87, Kif86, Kir86, Kir87a, KMR89, Kon85b, Kor87c, Kor89, Kru85, Kru89, KS89, Lis89, Mae87, Már89a, McG88a, Mel89, Nel88, Pap88b, Pic85, SW87, Sch86g, Shc85, Ste85c, Sto87a, Sul86, Tay86, Wal86c, Woe86, Yam89, dA85b, Buc85]. Randwertaufgaben [Col86]. rang [Hir89a, MVW87d]. range [Ara88, Ens85, Ens88, LS88h, Pic85, Yaj85b]. Rank [Mül87h, AP86, BSK87a, BE87, DD85, ES87b, Gek86d, Gol86a, Ito89, Mül87b, Mül87g, Ron86, Sel88d, Shi86b, Str89a, vdG89]. rank- [ES87b, vdG89]. ranks [Möh87]. Rapid [Don89b]. rapidly [Bre87, Ric86]. rapport [Bar87f, PSS86]. rare [Alo89]. raréfaction [Mét89]. rate [GR88, Hoh87, Kal85, Kuu86, Obr85, Obr89, Ome89, Pic88]. Rates [Gri86a, Mas86a, Bac86]. ratio [AB87, CM88a, Kik89, Mil87b, Shc85]. Rational [BC88, Don85n, Erk88, KS87, LR88a, Pet86, Sil89e, tW87c, Ani86, AFP88, AH86, FT88a, GR88, Har88b, Ill88, IN86h, IN86m, Kaw86a, Kno89, Kov87, Mañ88, Njå88, Pie88a, Roo88, Sil89a, Sta87c, Str87a, GPS87b]. Rationalité [Hir89c]. rationality [ES87b, Mar86c]. rationals [Dun87]. rationnelle [FT86a]. rationnelles [Hir89c]. Ratner [FS88a]. Räume [Fle85a]. Räumen [Gol87c, KT85b]. rays [Lau87a]. re [BW85b]. re-entrant [BW85b]. reacting [BA89]. Reaction [Sta86a, FB86, Kot89, Kuč86a, MPS87, Mar87, RB87]. Reaction-diffusion [Sta86a, Kot89, Kuč86a, MPS87, Mar87]. reaction-hyperbolic [RB87]. reactions [Str89b]. Real [Est86, Hel87, LL89, LMZ86j, Sil89g, Sil89h, Sil89f, Tho85f, BSK87a, BR89, Bor85, BS89j, D’A87a, Dic85, Dub88, Gud88, Hop89b, IN86n, IN86j, IN86l, KP86k, Kec88c, LF89, Lub85, Mes85, Mur88c, Nei88, Phe89c, PX87, Pre88l, RW85, Sch86r, She87a, She89b, Sil89a, Sil89b, Sil89c, Sil89d, Sil89e, Wal85b]. real-valued [Bor85, IN86n, IN86j, IN86l]. realisation [GMW87]. realizability [BS89a]. reals [Ste85c]. Réarrangement [Bad85, Bou89a]. Rearrangements [Kaw85c, Kaw85d, Tal86b, CG89]. reasoning [Rey86a, vdB87e]. reciprocal [Fri88b]. reciprocals [Jos85a, LS87c]. reciprocity [Liu88, Lor88]. records [Nev87]. Recouvrement [Cho87a]. recovery [MR85c]. rectangles [Ges86a]. rectangular [Sad86]. Rectification [Gal89]. Recurrence [FNT87, PV87, HS88, MM85, ŌT87, PS89a, Van87a]. Recurrent [Sch86g]. recurring [Tur89]. Recursion [Cho85b, Sim85d, DM85c, NR85g, Sac85, Sla85a, EMS85]. Recursively [Cho89b, AS85a, Joc85]. Reduced [dlH85, Aba86, Kle87, Pie89a, Pou88, Sel88i]. reduces [Vog87b]. reducibilities [DM85c]. reducible [Rum85, Rum86]. Reducing [Gul86b]. réductifs [DvdB87]. Reduction [ABHK85, Pie89d, Sto87b, Wri89a]. Reductions [Don85o, Pre88i]. reductive [And87a, DvdB87, Jan86b, Tit87]. réductives [MVW87c]. réduites [Cou89]. Ree [Coh88b]. réelles [Bar86c, Reu86a]. reels [Pas87b]. reference [Kim86]. refinement [Ban86, Cae86, KZ85a, Kag89b]. Refinements [Pre88j]. reflected [BPW86, Tan87]. reflecting [CL89b]. Reflection [Sla85a, Kat89a, Sch86o, Sch86p, TW86]. reflectionless [Sch86q, Sch86r]. Reflections [Rob89]. reflective [ART88]. reflector [Ble85b]. Reflexive [Wun87, Lan85a]. régénératif [AH89]. regenerative [KV85, Sze87]. region [Noc86, Oza88]. regions [BJ86, Blu85b, FTO87, IW86, JT86, Jac86b, vdB87b]. register [Pre88b]. register-shifts [Pre88b]. réglée [GS87]. regression [Plu89]. Regular [Ced88, Dre85, Gen87, Hej88, Van87i, AK88, Ara88, APS89, BG89a, Brü87c, Bus88b, Dho85, Her89, Hru87, Ito89, Kes88a, Kni86, LV89, Men88a, Paw85, Wei87o, vdE87]. Régularisation [LT85]. Régularité [Fer89, Hmi87]. Regularity [Lie88, Seg85b, Xin89, Blu85b, Cra86a, DL89, Gia86, Gia88a, Gia88b, God87a, GS86, Gud87, Hua88, Krá88a, Las86b, LS88b, Lie87, Lig85b, Lun86a, NP89a, Sch88a, Sim86, SJ86, Tay88, Wu88b, Xin87, Yan88e]. Regularization [Ber86k, Gro86a]. Regularized [LM89b]. regulière [MK88, Gér86a]. regurarity [Cla86]. Reihen [KPS87, Sch85g]. Reinhardt [AA86]. Reiten [Bre86a, Die87, IT86]. Related [CDG86, Ald85, Bal89, Bér86e, BCSSC87, BLM89, CRW89, Dav85a, DM89b, Dho85, DR89a, FZ85, Fut88a, Gal85, Han85, Har86, Her88b, Hes86, KPT85, KL87b, Kob89, Kop89, Lew87, LN87, Mur89, Njå89a, PS89a, Qua88b, Rei89h, Sch89f, Siu85a, Sto85l, Sul86, Wai89, Won87]. relatif [Leh86]. relation [AD85a, DZ89b, Fre85, Kan88b, Lal88, LS87e, Mad85, Roĭ86, SP89, dP86]. Relationen [Sto85n]. Relations [Arq86, MS86t, Tho85g, AHKH86, ACN85, BB85, Clo85b, DLdS87, FT87b, GvW89, Gul86b, PV87, Tho85c, Ton87a]. relationship [RM87]. Relationships [And87a, BH89d]. Relative [HL88b, Tay85, Bor87a, GGL87, Hmi87, Sel88e, Tor86, Zha89b]. relatively [Gup89a, MR87]. relatives [Bla89]. Relativistic [Mül89, Rug89, Str89b, Wei89b, AP89, Hol89b, vW89]. relativity [CK89]. Relativized [LT88]. relator [CRT89, Dic88]. Relaxation [But88, Arn88a]. relaxed [Col88d, Ieş87c]. remainders [vdB87c, vdB87b]. remark [AHK85, BS87c, Fuj87, Kát85a, Kaz89, May86d, Par86c, PH89, Sch89c, Hla85c]. remarkable [MvR88]. Remarks [Bau89, Ber85d, Boj88b, Boj88a, Bou87f, DG86b, FR86, Fre88, HKL88, HS88, HT87, IKPR88, KS86e, Lex88, Mor86e, Pap88a, Phe89i, She86o, She86n, Sri85, SU87, Xia87, Bal88, Bou85a, Bra86f, Bra86j, Car89b, CKS87a, CW89, CP89c, Gia88b, Hop87, LM86b, Lev85h, Mil88b, Mil87b, NW89, Par86d, Sch85a, Sch86a, Ste87a, vEKM89]. remarquable [AY89]. Remarque [Mey85b, Bak85c, Mey88e, Nua89, Str85b, dC88]. Remarques [Big85, Mey86b]. removable [Kau87]. renewal [KV89]. renewals [Mal85a]. Renyi [KV89, Vse85]. repeated [Hej88, Hol89a]. repellent [Sto87a].

Repetitive [Hap87]. Report [Gos89, Sch88h, LW89]. represent [Kes88a]. representability [Her88b]. représentable [JM86b, Par86c]. Representation [Ben86a, BV88o, Die85a, HS87b, PS88a, Ral87c, Vos86b, Lic86, Maa87, Ark89, BC86f, Bon85, Dyl88, FdlP86, Fis86, Fug88b, FM88, GL87, Ill89, Jur85, Kni86, Le 88c, Len86, McC85c, Neu89, Pas87b, Pet89a, Rao89, Rog86, Ros88b, Rum86, Sch85d, Sch87j, Szé86, Wei87s, Wei87t, tD89a, tD89b, AH89, Car86d, FL86a, Mey89e, PSS86, Roo86a, DGM86a, DGM86b]. representation-finite [Fis86]. Representations [DGM86a, DGM86b, Don85n, Dür86c, Gol85, Sel88j, Woo85, Kai87c, Kai87b, MVW87e, MVW87d, Ade89, And87a, Atz88, BSK87a, BB87, Bar87a, Ber86i, Buz88, Cur86, DvdB87, FG88, Frö85, Gou88a, Ino89, Jag85, Jan86b, Kai87c, KŠ88c, Lai86, Lud88, Mac88b, MR85a, Mor88, Pan85a, Ran87, Reu86b, Ros87b, Rum85, Shi86e, Zal88, DvdB87]. Representing [Bry89a]. Reproducing [TC85]. reproduction [Dur86f, Kuu86]. répulsive [Cou89]. research [CO86]. réseaux [Nic87]. Residu [Ofm86d]. residual [GLC86]. residue [Hüb89d, VSP89, Wod87]. residues [Hüb89e, Ron86]. resistivity [Tan88a]. resolubles [Reu86a]. Resolution [Tor87c, LV86b, Mül87k, SZ87]. resolutions [Dra87b, PW86, Bou87a]. Résolvante [dlP89]. Résolvantes [dlP87b, dlP89, dlP87a]. resolvent [Fis87f, FK87]. resolvents [Wei87q]. Resonance [Ben86b, vdM85e]. resonances [Men88b, Mül88, Sjö87]. resonant [Sev86g]. resource [DM87a]. respect [Ada88, BH89c, Dra85, Ger88b, HL88b, Lev89f, Lóp88]. response [Del86]. ressemble [BLY87]. Restoration [MW86]. Restricted [Jan87, Fra89, Mat85, SW89a, vdM85h]. Restriction [Ken89, Bou87c]. Restrictions [Gre89, Ost85]. result [AEE85, BF88a, ESS85, Lec86b, Tie88]. résultante [Las86a]. resultants [Fri88b]. résultats [Del85, Dix85, Szp86]. Results [Lig86, AS85c, BCSSC85, BS89i, Bok88, Bre86b, Bre86c, CV87, CM89a, Da 85, Dau88a, Day85, Eke86b, EN87, Fab89, Fie88h, FZ85, Fut88a, GR89, GL85, Gre89, Gro86c, Har88a, Har87d, Hel88f, JJW86, Kob89, Kop89, Kří86, Lan85a, Las86b, LS88b, LC85d, LS88q, Mas88a, NS86a, NS86b, NS86c, NW86, Njå89b, Ocn85f, OR87b, PS87a, Pie89e, Pre88l, Rob88, Sil89c, Siu85a, SJ86, Vse85, Žen86, LM89b]. Retarded [Col88g]. retour [Mou88]. retournement [Par86a, Pic86]. retracts [Ulr88a]. reversal [Föl86]. Reversible [Sev86h, Le 88c, Sev86a, Sev86c, Sev86d, Sev86e, Sev86f, Sev86g]. reversibles [MZ85]. Review [dLQ89, Kra88b]. revisited [ABB85, Bro87, BT87c, Don85j, Ral87g, Str87b, Vog89]. Reymond [Hla85c, Hla85a, Hla85c]. RG [Bea86a]. RG-dimodule [Bea86a]. Rham [Wt87b, Asa89, Dup88, Gui88a]. rhythms [Car85c]. Riccati [Coo89, Vog87b, Vog89]. Ricci [Bak87b, BK88a]. Ricci-flat [BK88a]. Rice [Wsc85c]. Riemann [BLR87, BJO89b, Bry89a, BBD88, Erk88, Far87b, KK89a, tLcZ88, Rod87a, Rod87d, lS89]. Riemannian [SSS86, ABB86, And88, Bér86h, Blo85, CH85, DD85, Don86b, Don88c, Drd85, Fal86, Fuk86a, Gli86, GPT89, GR87b, Gro86f, Hil85a, Hu87, Kan86a, Koz85, Kus88, Lev89j, Nis86, PS88c, Rec85, Sch89f, Thi86, Zhe87]. riemannienne [Aro88b]. riemanniennes [Bak87b]. Riemenschneider [Dem87b]. Riesz [Bak85a, Bak85b, Bak87b, Cle87, Fug88b, Pee88, Que87g]. right [Ced88, DR89a, Fou88, Sch86j, Sch86k, Sch86l, Shi86l, Shi86k]. rigid [She86e]. rigidity [BV88c, FS88a, Kuh87]. Rigorous [Sch86e, GR89, Sén86]. Ring [BMMS86, Gla89g, Gla89h, BR86, Bru86a, Bru86b, Bru86c, Coh85, Dus88, For86, FS87, Gur85, JM86a, Kom89b, LMM86, LMS86a, May86a, McC86c, McC86b, Ste86a, Woo87b, vO86]. Rings [Par88a, Aoy86, Ara88, AR87, AV88b, Ben86a, Bjö89a, Bok88, BF85, Bro85, Bru88, Bus88b, Ced88, Che86c, DH86, Die85a, Eks89, Esp88, EK88b, Est88, GG86, Gla89d, Gla89e, Gul86b, Haj87, Her88a, HS87c, HLM88, Hod89, LS87a, Lec86a, LV89, Lor88, Mar86a, McC85b, McC85d, Meh85, Men88a, Okn88a, Okn88b, Oli85, PR89a, Pra85, Rei85, Rog88, San85b, Sch86f, Sta87a, Tay85, Van87i, War86, vdE87]. Ritt [tW87a]. rl [Shi86a]. rl-equivalence [Shi86a]. robust [RZ87, vdW86, Wei86n]. robustesse [Hub86]. robustness [HM88a, Hub86]. Roch [Rod87a]. Rochberg [Mey85a]. rod [Pie89f]. Rogers [And89, AJBL89, Bre89a]. Rohlin [KV88, Kir89o, Ocn85j]. Roiter [BT86d]. role [DLdS87, Kak89]. Romania [AMSV85]. Rome [DFZ85]. roof [HK89]. roof-duality [HK89]. roos [BH86a]. root [LS85b, LS88l, Pet89e, Sel88i]. Rosser [MO88]. Rotation [Mir85, Hid89, Joh86]. rotations [Dum85b]. rotator [Pic85]. Rothe [Kač86, Mar86b]. rough [BO85, Kan86a]. round [Pom85]. row [Fin88a]. rpg [AM88e]. RPS [Kob89]. Rudvalis [Hua86]. rule [Gar86b]. ruled [APS89, Her89, SS86, Sil89h]. rules [Kiw85g]. Runge [Cas86, HRL89g, HRL89h, NS89b]. Rutgers [RLQ85].

S [AA86, Sal89]. S.D.G. [Gag88]. Sacksteder [Hur88]. Saff [Sta87b]. Saint [AIJH85, Hen86c, AIJH85, Ieş87c, Ieş87d]. Saint-Flour [AIJH85, AIJH85]. Saint-Venant [Ieş87c, Ieş87d]. Salamanca [GPR87]. same [DD85]. sample [Car86a, MP85]. sans [Den86, Zay87]. Sant [CKGP87]. Santaló [MP89a]. Santiago [dPR86]. satellite [LM85]. satisfying [CAXD86f, Kra88b, Már89a]. Sato [Bar87b]. Saturated [Fra89, Kei87]. saturation [Phi87f]. Sätze [BDT87b]. Sauer [ST89n]. sauts [Léa87]. SCABs [Kan86c]. scalaire [Aub86]. Scalar [Gra87, Kre85, MH88c, MH88d, Mor86f, Car88a, FB86, Fun88, Gul88b, Sch89f]. scale [NR85e]. scales [Aro88a]. scaling [IN86q, Sch86e]. scattered [Kre85]. Scattering [Coo86a, Ges86b, Sch86i, AL88a, ABGS85, BDT87a, CRD87, Ens85, Ens86, HT87, JMN86, Men88b, Sai86, Yaj85b]. scenario [Col88h]. Schattengrenzen [Sch85k]. schéma [Szp86]. schémas [Hir89c]. Scheme [EEK88, CM87b, DL87, FR86, Fuc86, Iar85, Kle87, Koo88]. Scheme-theoretic [EEK88]. schemes [Brz85, DGM88, Har87d, Hem86, Ler89, MS89b, OD85a, Pic87, Roe87, Sho87, Tad87, Wu86]. Schloss [LRT85]. Schmidt [CC87a]. Schnirelmann [Mon87]. Schoenberg [Fal86]. School [AIJH85, CCCN85, CCCN87, CDG86, Han87j, Tur85, dPR86]. Schottky [Don88a, vdG85]. Schouten [Kra88c]. Schrödinger [Agm85, ABHK85, AGHK^{+}86, AB87, AD85b, Bal86b, Car86c, CW89, CBD87, DMR87, Gra85a, Gru86a, Han89, HT87, Hel88g, Hel88h, Hin87, Kir85a, Kir86, Kir87a, MB85, Sai86, Sim85a, Wih85, Zah89]. Schubert [Gal85, Hal89, KSX87]. Schur [AS85l, AS85k, DM85a, JNT86]. Schwach [Drm87]. Schwartz [Lud88]. Schwartz-algebra [Lud88]. Schwarz [BMS87]. Schwarzian [DK88b]. science [CC88]. scientific [Bor86]. scindes [Rou87]. Scotland [SJ85, Thr86]. Scott [Kni87]. Scuola [Ves85]. sea [Bel85a]. secant [Gri86a]. secants [Lon86]. Second [Spe86, Abe87i, Aro88a, BD89a, Ber85d, Bih85, Böh86, ENS87, Gos88, Hin85, JSC87, Lab86a, Las86b, Lau87a, LS88b, Lio89, Mas85, MN88a, Nir88, Owe86, Sch85c, Tur89]. second-order [Sch85c, Tur89]. Seconde [Bak85b]. section [SS86, Wei87c, Wei87d, Wei87b, Wei87e]. sectional [Bal86a]. Sections [Hir87, Bal89, BL89b, CO89]. sector [Ros85b]. see [Erd85]. Segal [Itō88, May86d, Rei89b, Rei89a, Rei89e, Rei89j]. Segur [IN86j]. Seifert [KLR85, LR89a]. Selberg [Fis87a, Fis87c, Fis87d]. selection [Kim86]. selections [Gel86]. Self [Des85, Ngu85, Sto87a, BDT87a, Bri86, EB88, GG85, GK87a, Sán85a, SY87, Wei87g, Wei87j, Wei87q, Wei87t, Wid85g]. self-adjoint [BDT87a, Wei87g, Wei87j, Wei87q, Wei87t, Wid85g]. Self-adjointness [Ngu85]. self-dual [Sán85a]. Self-duality [Des85]. self-maps [GK87a]. Self-repellent [Sto87a]. self-similaires [SY87]. selfadjoint [Chi85]. Selfinjective [Tac86, TW86]. selon [Car86d, Kah88]. Semi [Est88, GRT88, Hmi89, Wei87e, Ass85, Bak85a, Bak85b, CBD87, CL85, Dau88a, DM87a, Dre87, Har88a, Hel88e, Kam89a, Lum89, Maa89, Mar88c, San87]. Semi-boundedness [Wei87e]. Semi-Classical [GRT88, CBD87, Hel88e]. semi-classiques [Mar88c]. semi-Fredholm [Dau88a]. semi-free [Kam89a]. Semi-groupes [Hmi89, Bak85a, Bak85b, Lum89, Maa89]. semi-linear [Har88a]. semi-martingale [DM87a]. Semi-rings [Est88]. semi-simple [Ass85, San87]. semi-stables [Dre87]. Semialgebraic [DK85d, Kne89a]. Semiclassical [Sjö87]. semiconductor [BCM86, Sei89b]. semiconductors [Gaj86]. semicontinuous [Phe89g]. semidistributive [Nat85a]. Semifree [AV85]. Semigroup [DH86, Lja88, EN88, Eva87, FS88b, Haz89, Hög89, Mar87, Muk89, Okn88b]. semigroupes [Bak89]. Semigroups [GZ87b, Nag86b, Pav87c, AHK85, AK88, Are86a, Are86b, BL89a, BH89b, DK88a, DS86a, Est86, Fou88, Gob88, Gre86b, Gre86c, Gre86d, Gre86e, Gro86b, Gro86c, Gro86d, Gro86e, Hal87, He89, Jaj88, Jon88, Lal88, McF88, NS86b, NS86d, Nag88, Neu86a, Pav87d, Pir87, Tim88, Wak87, Yan88d]. semihereditary [GG86]. Semilinear [Oha86, Avr87, BM88, Di 86, ENS87, FM89, LN88, Mor89, Tho86a]. semimartingales [Str85c, Str85b]. semimodular [Jon88]. Séminaire [JY85, Nor86a, AY85, AY86, AYM87, HM87, LDS86, Mal85c, Mal86b]. Seminar [CCCN85, CCCN87, Hop89a, KZ85b, KPZ87, Thr86, Ves85, Ław85]. Seminarberichte [Hla85d, Hla87b]. semipositive [Siu85a]. Semiprime [EK88b, Sta87a]. semirings [HvL88]. Semisimple [Józ88, DM86, Jar85c, Kaw87, Lus87, Ric87]. Semmes [Roc88]. sense [Fom88a]. sentences [She86k]. separable [DH86]. séparables [Fer85a]. separably [CCS^{+}87]. separate [Njå89b]. Separation [Ben85a, Cam86, GG85, Ich87, Wei87b]. séparé [Hau85]. Sept [AMSV85, DT87, Gra85a, KZ85b, LRT85]. September [ABS86, CDG86, DH85b, Ros86a, Tau86c, Tua86]. sequence [Ber88a, Bru86a, CGC88, Dav87d, DL86b, FZ87, MT87, Pro87, Shi86p]. sequences [AO88, Anc87, Att86, AU86, AR87, CS89a, EH88a, ESS85, Gue85, Gue86, Hen89b, Led86b, MS89a, Mir85, Pro87, Que87d, Rog87, Shi89, Sil88a, Tho87, Tur89]. Sequential [EM86, GMSW86, MW86, Nyi89]. SERC [Tur85]. Series [Dür86c, Ste85b, Les88, AL86, Ani88, AS85l, BSK87b, BH89d, FGL86, Gre89, GP87a, Gul86b, Hen89a, Kno85, Kri85a, Kri85g, Meh85, Nik87, PX87, Ral87h, Rei85, Roo88, SP89, Sch85i, SS88b, Sol86, VJ85, Wei86h, Wid85c, Wie87, Le 86b, DF86a, Dix85]. server [Sze87]. sesquilinear [Kur87]. Session [Bre85, FP86, Giu85, Gra85a, LP86, Ros86a, Tal86a]. Set [HPS87e, Aul85, Fer85c, Fuk86a, Gel88, Gri85a, Kak89, Kni87, Lau87a, LN87, Mel89, PR85b, Pre88m, RC85, Shi86h, Shi86q, Tal89, Van87d]. set-theorists [Tal89]. set-valued [Fer85c]. Sets [Kaw85d, Kec85, ASV88, All89, Bal88, BP88a, Bou85b, Bou87e, Bou87f, Bro85, Bue87, Bum85, Cho89b, Cil87, Cor85, Dal86, Dau87, Går87k, Gau87, GH85a, GNOY88, Hom85, Joc85, Kau87, Kaw88, Liu88, Maw88, Mum88q, Pre88a, Rey86a, Sch86a, She86d, She86o, Tak88b, Tay86, Tho85a, Van87e, vEKM89, vdB87d]. setting [Han88]. seventh [Kor88]. several [DGM88, Els85g, Jaj89b, Kra87a, Roc88, SH86]. Severi [Brz85, DH88, Pir89]. sextics [Per85]. sexual [Dur86f]. Shabat [Sch86q, Sch86r]. Shadow [vdB87g]. shallow [DF86b]. Shaneson [Mil85a]. Shanghai [GBB87]. shape [Alt88, Bla85a, SS87, She87e, Yag87]. Sharp [Går87h, Las86b, ST89k]. sharper [Mac88a]. sharpness [LS88h]. Sheaflike [Dür86e]. Sheaves [BMS87, Die86c, DW87, Mum88u, Beĭ87c, Beĭ87b, Cat85a, Die86a, Die86b, Die86d, Eno88, Lis87]. sheet [Kus86a]. sheets [Gar88a]. Shelah [ST89n]. shell [Rog85]. shift [BK88b, For85, Mül87l]. shifts [Mey88a, Pre88b, Sch86o, Sch86p, Sul87]. Shilov [Ton87b]. Shimura [CS87a]. Shiro [Aoy86]. shock [Che88a]. shocks [DL87]. Short [IK88, Ens85]. shortest [NP89b, Shn86]. Shrinking [KL87a]. Sibony [Ton87a]. Sicily [CDG86]. side [Bou87g, Tvr86]. sided [BT86a, Kas86, Lie87, Par87a, Sch87g, Shi86u]. Sidon [Bou85b]. Siebenmann [Mil88b]. Siegel [Fal87, LW88, LW89, Ral87f, Ral87g, Wei87v, Wei89a, dM88]. Siegelschen [Wei86f]. sieve [All85b, DH85a, GR85j, Hil89]. Sigma [Din85a, Sán85a]. sign [Shi86b, Shi86s]. signal [Dau87]. signaling [Haj89]. Signature [Hil85b, KK89b]. significance [Joh86, Jos88]. similaires [SY87]. similar [EB88]. similarity [Grü88a, Lar88]. Simple [Die86c, FW89, MS86t, Pil87, Asc86, Ass85, Ass89, Bar87c, Boy88, Bue86, CRW89, Dem87b, DV86, Die86b, Die86d, ÉY87, Fis86, HR88a, HS87c, Ill89, Jac86b, Jež85, Mic86, Mur89, SB88, San87, SW89a, Var88b, dlH85]. simple-homotopy [Ill89]. simplex [Kes88a, Kes88b]. simplexes [Kes88c]. Simplices [Win85b, Win85e]. Simplicial [Kne89d, Dup88, DK89]. simplicity [Han87b]. Simplified [Nor86b, HRL89i]. Simply [Con86, Pup88, HN89, Žub88]. simply-connected [Žub88]. Sims [Hua86]. simulation [DT88b, Ede86, Fuc86, Kat89b]. simulations [dLQ89]. Simultaneous [CL89b, Le 89, dB85, AK87, Pet89d, Pie88a]. sine [Sch86k, Smi89]. sine-Gordon [Sch86k, Smi89]. single [GMW87, Lal88]. singleton [Aul85]. Singular [BCM86, Dav89, Els85h, Kon85h, Kon85i, NR85b, NR85c, ST89l, ABHK85, AV88a, Arn88a, Ben85b, Ber86j, Bli86, BT86c, Brü85, Brü87c, BV88i, Can88, DB87, DG86a, Els85e, Els85f, Els85g, Fin88a, Fu88, Ger88b, Gri86c, Kec85, Khi86, Kor88, MC89, Per85, Pou88, Pov86, She86m, SV85, Wei87p, ZG88]. singularitè [Gér86a]. Singularités [Mag86, Ber85e, Ble86, Bon89]. Singularities [Dau88f, Dur85a, Går87i, Går87j, KSLP86, Sch85f, SS88d, Whi86, Asa85, AU86, Bar87e, BW88, Bré89b, Chr87, Col88a, Con85b, Ebe87b, HS87c, KS86d, LP88a, LP88g, Len86, Lev85f, LV89, Pap85, jQxQ88, Ran88, RS88, Rei86, Rei87b, RZ86, Sch87b, Sim85c, SZ87, Ste85a, SS88c, Str86d, Var88b, Wun87, vS87b, vdE87, GWW85]. singularity [Aus85, Die87, Jan86a, KL88, Kos88b, MPST85, SB88, Web88]. Singularly [NR85d, NR85a]. singulier [Tho88]. singuliers [Gér86a]. Sinks [Pre88k]. site [Hir89b]. Sitges [CAWXD85]. six [Idà89]. size [Bru89]. skeleta [HL86]. skeletons [Dav85b]. Sketch [KSX87]. Skew [GR85e, Mat86d, Idà89, Lor88, McC85b, Sel88h]. Skew-symmetric [GR85e]. Skitovitch [KZ85a]. Skorohod [Ben87, Fit88, McC85c]. Skorohod-like [McC85c]. sleeping [Sap88]. slightly [BT86c]. slit [Ste85b]. slowly [MMT87, Sch86j, Sch86l]. Small [Mas87, Pav85, ACM85, Bue87, FTO87, Fit87a, FK88, Fre87, Joh89a, Kar85, Led86a, Wat88]. small-time [Wat88]. smash [LM86c, LMS86b]. SMF [FHMM86]. Smooth [AG89, Aso89, KSLP86, MR85a, SS85d, Alu88, Bea87, Che88c, CH89b, Dau88i, DH89, Eke86a, Frö86, Fuk88b, Hal87, Jar89, KS86d, Kle87, Lev89d, Lie88, Mas86b, Mat89b, Maz86, MS89b, Phe89j, Shi88d, Str86d]. Smoothing [HH85, KP85, Lev89k, Pic89b, VSP89, vdW86]. Smoothness [Che88a, SS85e, SS85f, Wro87]. Sobolev [Boh88, Boj88a, BG89c, CF88, ÉY87, Ger86b, Har85b, Ken89, Lew87, Mét87, Ngu85, Por85, Sch87k, Wu87d]. Sobolev-imbedding [Mét87]. Sobolev-type [Ngu85]. sofic [BK88b, HN88]. sojourns [Ber85c]. solenoids [Kee87]. solid [AGHK^{+}86, Beb86, Ben85c, Car89c, GRT88]. solids [Ker85]. Soliton [Sch86i]. solitons [Sch86m]. Soluble [Hei87, Len87, Neu89]. Solution [HRL89i, BO85, Bel87a, Ber87b, BFKP86, BH89c, BLR87, ČZ86, Col86, Går87d, Gar88b, GP85, HRL89e, Hem86, Hen86a, IN86b, IN86a, IN86i, KK89a, McG88b, SH86, VSP89, Vas86, Wri89b, vW85]. Solutions [Ama87, Joh89a, AEE85, AT86, Ado87, Aki89, Alb88, ABHK85, Bac89, Bar86e, Bih85, Bol88, Bri85b, CAXD86a, Che88a, Chu88a, Coo89, DL86a, DG89, DL89, Don86b, Eng87a, Eng87b, Eva87, Fen89, Fit87b, FM89, Fre89, GR85f, Gel88, God87a, GL85, Gri86c, Grü88a, Gul88a, GF87, HW88, Hel87, HS87e, Hör87, Hua88, HjJ89, IK88, IN86n, IN86o, IN86j, IN86l, IN86k, Jac88c, KPA89, Kod86, Krb86, Kre86b, KH86, Kuc86b, Kus86b, Lan88a, Lap88, tLqL87, tLmC89, Li89b, Lie87, Lie88, Lio85, Lio89, Lit88, Mar89c, Mc 87, Mei88, MV88, Mey89a, Mil88a, Mou85, Net88a, Neu86c, Oco87, PS88c, Prü86, Rab88, Sch87l, Sch86j, Sch86k, Sch86l, Sim86, Spe86, SS88c, SS88d, Ste86b, Val86b, Wag87, Wei87a, Wei87r, Yan88e, YM89, Zvy88]. solvability [Aro88a, Erk87, HSG88]. Solvable [Abe87c, Abe87h, GLR86, Gup89a, KŠ88c, MB85]. solve [GMW87]. solved [Erd85, Mei86a]. solver [AMM86, vdW86]. solving [Aga86b, DF86b, Jar89]. Some [AK88, AL88c, Bae89, Bal88, Ben85b, Bok88, Bou85a, Bry87b, BH89d, CFdG87, Car86e, Car89b, CS89i, CKS87a, CW89, Cec88, CM89a, Coe87, CP89c, Da 85, Dav87d, Day85, Dra87c, Dur86f, Eke88, Fab89, Fef88, Fri89, GR85f, Gau85, Gia88b, GR86, GL85, Grü86b, GU87b, HSG88, Hil89, Hop87, Jen87, JL85, Kal86b, Kop89, Kor87a, Lab86b, Lal88, Lan85a, LM86b, LS87c, Lev85h, Lev89k, Lev89l, Lin88b, Lon88, Mae87, Man86b, MW89, Mei86a, MR85b, Mha87, Mil88b, MS86k, MS86s, MS86t, Mil87b, Mor88, MvR88, MY87b, Mum88v, NS89a, NW86, Nat85a, NW89, Obr85, PS87a, Par86d, Par89, Pie88b, Plu89, Pre88l, Pri89, Pug86, Rad87, Rod87e, Rog88, Sac85, Sch85a, Sch86a, Shi86r, Shi86s, Sim85a, SS89b, Siu85a, Sri85, Sta85, SJ86]. Some [Ste87a, Str86c, Šve86, Tho86b, Üst87, Üst88, Van87g, Vir88, Vse85, Wei85b, Žen86, vW89, Ara85, Arn88b, BSK87a, Bar89c, Bar85d, Bér86a, BB88a, Ber86i, Ber86j, Ber85d, BG89b, Bih85, Bor85, BM85, CM87a, Cha86c, Cho87b, Coh87a, DST89, Ebe87b, Erd85, Fei86b, Fer85b, FP85, Fle85b, Gal85, Gre89, Hab87, Hay88, Hei85, Jos85a, Kat89a, Kob89, Kod86, Kon85a, KP85, Kus86b, LR89b, tLqL87, LR88b, Mar86c, Mei88, Men89, Mil88c, Mos85, Mul86, Njå89b, NZ89, Oco87, Pik86b, Pov86, Rog87, Shi86v, Sun86, TC85, dS87, CC89b, Col86, NS86e]. son [San87]. soniques [Mét89]. sont [Mar85]. sound [Fil85]. source [Roe87, WST88]. Sous [Cho87a, Aro88b, Bak85c, Bak85b, Bak88, Cau86b, Den86, GB86, Szp86]. sous-analytiques [Den86]. sous-corps [Cau86b]. sous-diagonaux [GB86]. Sous-Ensembles [Cho87a, Szp86]. sous-harmonicité [Bak88]. sous-riemannienne [Aro88b]. Space [CS89f, GP87a, MS86e, AHKHK89, Bac89, BE86, BM88, BR85b, BN85, Blo85, BWW85, Bou85a, BHT89, CS89a, CS89d, CS89e, CS89g, CS89j, CS89k, Che86d, Che89, Cil87, Coe87, Coh87b, CP89c, DD85, DM85b, ES87b, Erk87, Fab89, Föl86, Fri89, GR85a, Gri86a, Gu87a, Har88c, HH85, HM88a, Hud89, Itō88, KS86a, Kes88b, Kle87, KÜ88, Kow85, Lar88, Le 88c, Lis89, Łuc89f, McC86b, MN88e, Mey87b, MP89b, NR85a, Ne’87, PlT88m, Par86d, Par88b, PS88c, Qua88b, Rac85, Rec85, Rog89, Ros85a, Rub86, Sch89a, SZ85, Seg85b, Sel88i, Shi86m, Shi88a, ST89i, Sul87, TC85, Tim88, Wen85, Wit88b, tW87b, Wu87a, dM88]. space-level [McC86b]. space-like [DD85]. space-valued [Łuc89f]. Spaces [BS86, BDH^{+}85, Bel89a, DK85d, FO86, FHMM86, KS85a, Kri85d, LC85f, Mat86d, MS86c, MS86d, MS86f, MS86i, MS86q, MS86s, MS86t, Sch87k, SS85b, vEdG85g, AM88a, AM88b, Ani89, Ara87, AFP87, ADR85, Asa89, Ass89, AA89, Bar89d, BD89b, BS87c, BR89, BT88b, BB88b, Böd87, BC88, Boj88b, Bou88b, BL89b, BG89c, Brü87b, Bur86, BG89d, Cec85, Che86a, Che86b, CG89, CP87, CF88, CHS85, CD87, CS88c, Da 85, DL86a, DP88, DvdB87, DLdS87, Din85b, Don89a, DG86a, Dow89, DZ87, Ebe86, EH88a, Ele87, Eno88, Est88, Fal86, Far87a, FT88a, For88b, FJ88, FHJW89, Gar86a, GM85a, Gil87, Gri86c, HW88, Han88, HS87b, Hat86, HV88, Heg88, Hei85, Hei86, Hir85, HR85, Hop89b, Hri87, Hu87, Hus89, Ike88, JM89a, Jaj89a, JM85, JLS87]. spaces [Joh88b, KPA89, Kam85, Kat86a, Kaw87, Kif86, KK85, Kne89a, Kne89d, Kuc86b, Kuf86, Kuh89, KLR85, Kur87, Lar88, Led86a, Lew88b, LT89, LC85d, LC85j, LS87e, Lin88b, LÖ86, LMZ86a, LMZ86b, LMZ86n, MP85, Mar86c, Mas86b, Mat86a, Mat86e, MS86a, MS87a, Mil85d, Mil85e, Mil87b, MS85, Mur87, NS86b, NS86c, NS86d, Nei85, Nei88, Neu86a, Now87, Nyi89, Oha86, Oko86a, Orl87, Oro89, PlT88j, Phe89a, Phe89c, Phe89e, Phe89h, Phe89j, Pie89g, Pie89h, Pis86, PX87, Por85, Pou88, Pov86, Rad87, Raw85, Ray88, Rea88, Rei89f, Rei89g, RTVW88, Roo88, RS85b, Sch85a, Sch88a, Sch87d, Sch86g, Smi85a, ST89j, SS85e, SS85f, SS85d, Tri86, Vog85, Wal85a, Wal88, Win85f, Xin89, Xu87a, Yuk86, dG85, vEdG85a, vEdG85b, vEdG85c, vEdG85d]. Spain [BS86, CAWXD85, CKGP87, GPR87, NFM86]. spanned [CS89e]. spanning [Tro88]. Spatial [GS85b, Tau86c, Ges86b, Gor86, GS86, Kuu86, MPS87, Tău86a]. spdes [Par87a]. Special [AH85a, BT87a, HS85, Koo85, Neh87d, Ral87h, Sch89d, Wei87s, All89, Ebe87b, JW89, Kim86, Mul86, Pir89, Shi86s]. speciality [GP87a]. Specialization [Mum88w]. species [Lab86b]. Spectra [BMMS86, DR88a, Bru86a, Bru86b, Bru86c, Gro88, HS87e, LM86f, LMS86a, Lew86b, Mah87, May86a, McC86b, Ste86a]. Spectral [Bér86g, Bri86, Brü85, CBD87, GN86, Gre86d, Gre86e, Gro86e, Hel86h, Mül87h, Oro89, Que87i, Que87j, Que87k, Wei87u, Bér86a, Bro86, Bru86a, Dav87d, Est88, FZ87, Fer85c, GT85b, Grü88a, GU87b, Hej85, Hel86f, Hel86g, Kuc86b, Mül87k, Mül87l, NS87, Que87h, Sun87, Wei87f, Wei87s, Wei87t]. spectrale [FT88b, Nic87]. spectrales [Lim85]. spectre [Fer86]. Spectrum [RV89, Coe87, DM89b, DT89, EN86, Esp88, HR88a, Hin87, Koz85, Phi87f, iS88, Ton87b, Wei87i, Wei87a]. speed [Dav87b]. Spektralsequenz [Sto85g, Sto85h, Sto85b]. Spektrum [OS85, Wei86c]. Spektrums [Sto85m]. Spencer [LP88e]. Sphäre [BDT87b]. sphere [Aso89, Bry89a, BG88b, CC89a, Don88b, FKV88, HK85, Kir88, Mut86, VG88]. spheres [CGMM87, Dyl88, Ino89, Lev85g, NS86e, Ped85]. Spherical [Bar89b, AL88c, ByC86, DvdB87, Mac87, Pas86, Ron86, Sap86]. sphériques [DvdB87, San87]. Spin [Kir89p, Sel86, Bry85, Fan89, Tur88a]. spin-boson [Fan89]. Spin-like [Sel86]. spline [DGM88, Pic89b]. splines [Els85a]. split [AU86, AR87, Rog87]. Splitting [Hri87, MMT87, Woo87b, BEMG85, CM87d, Cra87, Giu88, Hul85, Ocn85i]. splittings [Ass86b, Böd87, LMM86]. Spontaneous [RD85]. sporadic [Fei86b, Kar85]. spread [Ros86b]. Springer [Ros87b]. spruce [GS85b]. spun [Kan89, Lit85]. Square [Jon89, BW85a, Le 89, LS85b, NV89, Sch86s]. squares [Ber88a, GP85, GLC86, NS85, She86n, Woo87b]. Squaring [KI87]. squeezing [AB89]. Stabile [Wei86o, Wei86m, Wei86n]. Stabilité [Ala88]. stabilities [GR85c]. Stability [Axe86, Bai89, Bel85b, CS86, Eno88, KR85, Kuč86a, She89b, Ura86, Zol86, AEE85, AOP86, CK89, DS86a, DT89, Gaj86, GR85c, GR85d, Gyö89, Hol89b, KN88b, PS87a, SB88, Šig87, Veg85, Jar85g, KZ85b, KPZ87]. stabilization [GR85e]. stabilized [AMM86]. Stable [Ara88, Böd87, Bry89b, LeP89a, Lev85c, LMS86c, Lis89, Mar86c, Ark89, APS89, Car86d, Cra87, ES87b, Fab89, GZ89, GM85b, Her89, HH86, HS86, Kas86, KMM87, LeP89b, Lev85e, Lev85f, LM86e, MC89, May86b, May86c, NS87, NS89a, Oko85, PS88b, Pic86, Ric87, RW85, Sch87g, Shc85]. stables [BL86, Dre87, Lap88]. Stallings [Sch85b]. Stampacchia [CDG86]. Standard [Gol88, Tor87d, Eli88, She86h]. Stanford [Hop89a]. star [Con89, LT88, Shi86i]. star-free [LT88]. Stark [Jen86]. starlike [GN88]. State [BHT89, BRA86e, CHS85, Gaj86, GRT88, GF87, HM88a, NR85a, TL86]. State-space [BHT89]. Statement [LS88q]. statements [Tor87c]. States [Han87i, AB85, Ben86b, BB88a, Bes85, Cra86a, D’A85a, Dau87, Eva85b, Fan89, Joh86, Kir85a]. static [MP88, Xu87b]. stationarity [Las89]. Stationary [Pit87, Bor85, Da 85, Fri85b, Küm88, Led86b, MS89a, Mat89a, She86g]. stationnaires [Fer89]. statistic [SW87, Sla85b]. Statistical [FHMM86, Kon85g, Oll85g, Oll85i, DL88, Le 86c, Lew88a, Mit86, Oll85b]. Statistics [BPW86, dPR86, Bar89c, Pan85b]. steady [DL87, Gaj86, GR85f, GF87, Hem86, Spe86]. steady-state [Gaj86]. Steenrod [Woo87b]. Stein [BL86, Dem86b, Geb86, LT87, Mey85a]. Steiner [BLM89, PS85]. step [EM86, Kni86]. steps [Bel89b]. Stern [NZ85]. Stetige [Tic87]. Steven [DHS85]. Stickelberger [McC85d]. Stiefel [McC85a]. Stieltjes [AC87, MH88d]. stiff [Axe86]. stimulated [JMN86]. Stirling [Reu86b]. Stochastic [AP85a, AF88, AHKH86, BD86a, Car88a, Cec89, Daw86b, Hab87, HK87b, KZ85b, KPZ87, Kuo89, MP85, Mod86, Nel88, Ros86b, Tău86a, WST88, ABC^{+}85, AGHK^{+}86, App85, App88, AKO86, BL85, Bar88a, BL89a, BT88a, Bax86, Ben85b, Ber85c, BR85b, Böh86, BM89c, BH89c, BRA86e, CV87, Car86a, CCE86, Car86f, Cho87b, CM89b, Cra86a, Cra86b, DL88, Daw86a, DG89, Faj85, FS89b, Fla87, Fle85b, Fri85b, Fri89, Fuk86b, Fun86, Gli86, Gli88, GQS85, Gyö89, Hab88, Har86, Hop87, Ich87, JL85, KPA89, KN88a, Kif88a, Kif88b, Kot89, Las89, Lio89, LB86, Mét86, MV88, Mos85, Nua86, Nua88, Oco87, Oco88, Par87a, Par86b, Par86d, Par88b, Pic88, Plu89, Pre87, Qua88a, RW85, Ros86c, Roz87, Str86a, Str86c, Tir88, TL86, Üst87]. stochastic [Wal86a, Wal86b, Wih86, Yas88, vK87, vdBL88, vW85, ACSC85, ABS86, ABS87, DT87, IH86, Tau86c]. stochastics [Le 86c, LR89b].

Stochastique [JY85, Ber87d, Hau85, JM86b, Kré88, Mey86b, Pra88, Yan87, Par86c]. stochastiques [BH86b, Lap88, dC85c]. Stockholm [Roo86b]. Stokes [AMM86, Bem88, Nit86, Paw85, Pon88, SH86, Val86b, Bei88]. Stone [Kee87]. Stong [Zag88]. stop [AS89a]. Stopping [Nag86a, Fit88]. straightening [BV88a]. strange [Pas89]. Strasbourg [FHMM86]. Strategies [Tor87a]. strategy [BT86b, IW86, Tor87b, Tor87f, WKK86]. Stratonovich [DZ89b, vW85]. Stratonovitch [HM88b]. Straus [LC85h, LC85a]. strength [Gud87, LSR88]. stress [Ste85a]. stretching [KL87a]. Strict [CAXD86r, CAXD86s, Kni86]. stricte [Del88]. Strictement [Cho87a]. strictly [Ark89, Ens88, Gol86b, KMM87, Sch87b]. Strings [Ne’87, Kre86a]. strip [Lac86]. Strong [Bau87, Cho85a, Col88i, DM85c, Hen89b, Jaj85e, Jaj85f, Jaj89b, LS87f, Coo86b, Hen89a, Phi87f, Ran86, Sch88a]. Strongly [ENS87, LV89, AS85b, Avr87, Axe86, Cho85b, Eng87a, GN88, Hir85, Oko85, Sim86]. strongly-damped [Avr87]. Structural [BDD89, HPS87f, SB88, Zwi89]. Structure [Bal86a, Ebe85, Sev86g, Wil86b, Arn88a, Bea86a, Bou85a, BV88p, CP89a, CKS87b, Che88c, Dav87a, DH89, Dum85b, Eli88, Fre87, Gel88, HZ87a, HLR85, Imk88e, JM86a, Kas87, LS87b, LP87a, McC85d, Mey89a, Mey89d, Mir85, Ne’87, Odi85, Pre88m, SS86, Som86, SZ87, Tay85]. Structures [CS87a, Azn87, Dür86b, Dür86e, Dwi87, Hai87b, Har85a, Joy86, Kar89, Kir89p, KL87a, KP86l, Law89, ŁR85, MR87, Mez86, MS85, PM89, Raw85, Ton87b, Tur88a, Wal85b]. Student [Sla85b]. Studies [BGV86]. Study [Hel88i, tW87c, All85b, Fei86a, FV89, GH85b, Gu87a, Kee87, Lan88a, VSP89]. Sturm [Brü85, Wei87e, Wei87i, Wei87o, Wei87p]. sub [Paw85]. sub-analytic [Paw85]. subalgebra [Eld89, GI85, Var89]. subanalytic [MS86a]. subcategories [ART88]. subcomplexes [Lev89e]. subdifferentials [Phe89h]. Subdirectly [Jež85, Nag88]. Subelliptic [JSC87]. subgradient [Kiw85a, Kiw85g, Kiw85h]. Subgroup [Haw87, Bea86a, BR88, Lau87b]. Subgroups [Phi87g, Asc86, CM89a, DM86, Don86a, Ebe86, FRN85, GGL87, Gup89b, Hua86, Kha89, Len87, LM87, Möh87, Rey86b, Rhe89, Sto87c, Tit87, Wil87]. Subharmonic [Ceg87, Rip87]. subharmonics [CC87a]. subideals [KN89]. submanifold [Wan87a]. Submanifolds [GM89, Hin85, ByC86, Bar89b, Bej85, Bry87a, CC89a, KS86a, PlT88g, PlT88h, PlT88j, PlT88k, RVV85, She89b]. submartingales [FS85b]. submersion [Rec85]. submodules [Hal89]. Subnormal [Sto87c]. Suboptimal [Els85a]. subordination [Bur89, Maa89, dlP87a, dlP89]. subpreschemes [Mum88c]. subrings [Bac86]. subschemes [Kle89]. subsemigroup [Res89]. Subsequences [CS89j, Ber88c, CS89e, Ray88]. subset [Mil87a, Njå88, Shi86t]. Subsets [Dun88, Kec88c, AOR88, EH88a, Mat89a, Maw88, MN88d, Mil87a]. subshifts [HN88]. subspace [CT88, Rea88, Vas86]. Subspaces [Bou85b, MS86e, MS86d, MS86q, Als88, BT87b, BKT89, CS89e, CS89k, Gor88, Mil85d, MS89c, Sul87]. substitution [Kas86, Que87e]. substitutions [Que87b, Que87c]. substratum [LP88h]. substructuring [Bra86d]. subsystem [CK88]. subtournament [Tab86]. Subvarieties [Oko86b, BM89b, Ric87]. subvariety [Siu86]. succeeding [Kir87a]. successes [Dra87c]. Successive [Coh87c]. Suchbaumen [KP87b]. Sudakov [MTJ87]. sue [Str85b]. sufficiency [Kag85]. sufficient [Bar89c, HJ85, Mai88, Nag87]. suite [FT88b, LP88c, LP88f]. suites [AAGV85, DR88b]. suivant [DR88b, Hir89c, Yan87]. sujet [Mey86b]. Sullivan [Oua88, Wt87b]. sum [Csö89, Fac89, HS87a, KPT85, Kor87c, Lac86, LR88b, Mas89]. summability [Mae87]. summable [BB88b]. summands [CG89]. summary [Buc85, GR85j, NZ89]. summation [Mel89]. Summer [AIJH85, Han87j, Tur85]. Sums [Jos85a, BLM88, CG89, GM86, HKW89, Haj89, Han85, Jut89, Kat85b, Kor89, Kru85, Kru89, NS85, OW87, PS87b, Tie88, dA85b]. Sumsets [Nat89b]. SUNY [LS86a]. super [Bry89a, Sel88d]. super-exceptional [Sel88d]. superalgebra [SVS87]. superalgebras [Józ88]. superchamp [Le 87]. Superconvergence [Kří86, NW86]. Supergauge [Alm87]. supergravities [CB87]. supergravity [Rog85]. superharmonic [Fug88b]. Superiore [Ves85]. Supermanifolds [Ne’85, Bru87d]. superperfect [Bla89]. Supersingular [Lan88d]. supersolvable [Var89]. superspace [Rog85]. superstable [Pil87]. supersymmetric [GR87b]. Supersymmetry [Hab88]. support [CMM89, DLS86]. support-finite [DLS86]. supports [Gyö89, Jur89, SS87]. sure [AOP86, LB86]. Surface [Jia89, BW85c, Gan85e, Gut87, Hul85, Oza85, SS86, Wan89, Wun87]. Surfaces [Aur89, BC86b, CTKS87, Gan85g, PlT88m, Wsc85e, ABL88, APS89, BC86a, BC86c, BC86e, Bar85c, Ber85e, BC88, BWW85, Bor88a, BBGM89, BBD88, Cat88, Caz88, Ceg87, CP89b, CQ85, Dav89, Die88a, Dzi88, Far87b, FKV88, Fom88b, Glo87, Gol85, Gra85b, Gu87a, Gul88b, Gut88, Har88b, Hei88b, Her89, Hof85, Hoy85, Hoy87, Hoy89, Kha88, KP85, Lev89a, Li89a, Ma89, Mar89b, Mil85c, MY87b, Per85, PS89a, Pir89, Reg87, SB88, Sch86d, lS89, Shi86w, Sil89a, Sil89b, Sil89c, Sil89g, Sil89h, Sil89e, Sil89f, Sil89i, Str88b, SY89, Tro88, Wai89, Xia89, Yan89, Zag85]. Surgering [Car88b]. Surgery [BD88, LR88c, Mas86b]. Surjectivity [Šed86]. surmartingale [Mey88d]. surplus [DM87a]. Survey [Car86f, COT89, Küm88, Nar87, AW86a, Bej85, Bér86a, Bor86, BR85c, Caf89, DGR87, Dic88, Fie88i, Hel87, JNT86, Key87, LS88b, Neu86b, Njå89b, Oli85, PW86, Sae87, San85b, Sch89b, Tău86a, Wih86]. Suslin [vdK85]. suspension [Coh87b, Gra89]. suspensions [BH89a, Ghy88]. Swan [Mil85b]. Switching [Sei89a]. Sylvester [And89]. symbol [FK87]. symbolic [Lav87]. symbolique [Lav87]. symbols [Sch87k]. Symétrie [Bar87b]. symétrique [San87]. symétriques [Ank88, Bak85a, Bak85b, Deb87, DvdB87, DF86a, Loe86]. Symmetric [Blo85, Gue85, Bou87d, Bou87e, BL89b, DvdB87, DZ89b, DLdS87, Ebe87e, Eng87a, Fal86, Fuk88b, GR85e, Hu87, Ich86, KRT86, LS87e, Liu88, Nag86a, Orl87, Ōsh87, ŌT87, PS88c, Ray88, Reu86b, Sev86a, Sev86c, Sin87, sT89a]. Symmetries [Lit85]. symmetrization [Bae87, Bae89, Mil88d]. symmetrizations [BLM88, BLM89]. symmetrizing [Lax87]. Symmetry [Ker86, Fie88i, Fie88j, FKZ86, GR85c, Kan89, Pic85, Pup88, Sap86]. symmetry-spun [Kan89]. symplectic [Arn88b, Cha85b, DP85, FZ85, Går87e, Hat89]. symplectique [Cru85a, MVW87d]. Symposium [ABS86, ABS87, BDM^{+}85, Di 85, GZ87b, GBB87, JP86, LRT85, NFM86, SSS86, Tua86]. Syntactical [Kat89b]. Synthèse [Adj87]. synthesis [Kuc86b, Oro89, Win89]. Synthetic [Rey86a, Koc86]. system [AEE85, Ber88c, Chu88a, Coh88a, Dur86f, Fla87, Fu88, GESEZ85, GH85b, GQS85, IN86b, IN86a, KK89a, Kob89, tLcZ88, Neu86c, Poh89a, Que87e, RB87, RV89, Sch86q, Sch86r, Shi89, Sho87, Wan87b, Wt87d]. Système [Deb87]. systèmes [MM86]. Systems [ACSC85, BBT85, CAXD86t, CAXD86u, CK88, Jac88c, Mod87, NR85c, Sev86h, Ara85, AKO86, Bac89, BM86, Bar88a, Ben86b, BDD89, Blu85a, Bou89a, BK88b, BV88n, Car86e, Car89b, CLS85, Cha86a, Che88a, Col88a, Cra86b, Dho85, DiP87, Düm85a, EB88, ES86, EN86, Ens85, Ens88, Fel88, Fit87b, FM89, Fom88b, Fre88, Ges86b, God87a, GMW87, Gre86f, GF87, HRL89e, HRL89g, HRL89i, HSG88, HM88a, HS87e, Hir85, Hof87, IN86h, IN86g, IN86e, IN86m, KM86a, Kog86e, Kog86g, KM86b, Koo85, KL87c, KH86, Kuč86a, Ler89, LV86b, tLqL87, Lic85, Lig86, Lin88a, Lio86, Liv87, LB86, Mar88a, Mar87, Mar89c, Mei88, Mei86b, Mor89, Oll85a, Oll85b, Oll85d, PS85, Pre87, Que87d, Que87c, Que87j, Rab88, Rof85, Ros89c, Ros86c, San86, Sch85i, Sei89a, Shn86]. systems [SJ86, Str86d, Šve86, Tho85d, Tim86b, Tir88, Tvr86, Wih86, Yan88e, Zab87, Zha88, vdM85e]. systems- [KH86]. Szego [Chr89, Dwi87, JNT86, Lóp88, Str85a, Wid85f, Wid85g, Wid85h].

T. [Dem87a]. table [Qua88a]. tableaux [GB86, dSCV86]. tables [Coo86b, Fin88a]. tails [Kir87a]. tall [Fac89]. Tame [Frö85, Hes86, Len86, Sch88h]. Tanaka [AY87, Par88b]. tangent [BW88, DK88c, Eno88, Ran88, Sch86d]. tangles [Ble85a]. Taniguchi [SSS86, Fuj87]. targets [Sto86]. Tauberian [Bin89b]. taxonomy [Mal85b]. Tchebycheff [dSCV85]. Technical [DH85b]. technique [CLS85]. techniques [CL87, Din85a, Hol89b, LS88a, Mét87, Roa88]. Teichmüller [Bra88, Iva88, Tro89a]. Teilfolgen [Los85a, Pro87]. Temperature [Fan89, For85]. Temperature-dependent [For85]. Temps [Ber87d, Le 87, Bou87b, Jeu85, Léa87, Lum89, Par86a, Pic86, SY87, Wsc85a, Yor85b, Yor85d]. Ten [Siu85b]. tend [AD85b]. tends [AD85b]. tense [GM88]. Tensor [Don85n, LC85f, Phi87h, Bry89b, Dem88, Ele87, HO85, Hin85, LC85j, vEdG85c]. Tensorial [Kan88a]. tent [AM88a]. Tents [Tho87]. Term [Sto85h, BG88a, Voj87g]. Terme [FP86, Tal86a]. terminal [Vin85]. terms [AC85, ABGS85, AHK85, Ass85, BR85b, CS88a, CGMM87, Fac89, Kat85b, Roe87, Sei89b]. ternary [Hoy85]. tessellations [Dre85]. test [Bar89c, NV89]. testing [Len85]. Tests [BE86, Col88j]. th [Bar86e, Fin88a, LS88l]. thêta [BD89a]. Thacker [DF86b]. Their [Aga85d, GWW85, IH86, Pas89, AHKH86, AK87, AMSV85, Asc86, BMMS86, Cra86a, Dau87, DLdS87, EM85, FKZ86, Gro86e, HRL89e, HPS87a, Imk88c, IN86o, IN86j, IN86k, KMM87, Lex88, LS88h, MP88, Ost85, RR85, Roo86b, RM87, Sel88h, Sev86f, Shi86e, Shi88f, Sto85l, VZ86, Whi86]. them [vEdG85b]. Theme [CS89l, ABB85]. Theorem [Hel86h, MS86l, MS86s, Mül87h, Reu86b, Aha87, Alb85, Ale85, AP85b, Anc87, AM88c, Bae87, BR85a, BEMG85, Bin86, Bor87a, Bra88, BT86d, Cae86, ČW87, Cha85b, Che86a, Clo85a, CM88c, DKL88, Deb89, Dud85, DG86b, Eli88, EN88, Fan88, FL86b, FTO87, FS88a, FHJW89, Fuj87, GZ86a, GZ89, Gud87, Hay88, HN89, Hel86b, Hel86g, Hoh87, Hur88, Jac86b, Kag89a, KV89, Kec88b, Kir89o, Kru85, KK89b, KS89, Kus88, LM86a, Led86a, Lev89e, LV89, Lis87, MP86, Mas86a, Mat86b, McF88, Oba89, Ocn85j, Ost85, Pic89b, Pre88g, Pro89, Rei89i, Ros88b, Ros87a, Sch87a, Sch89c, Sch86h, Shi87d, Šig87, Sim85d, Sta87b, SW89c, Tan88b, TT88, Vse85, Wei85a, Wt87b, tW87a, tW87b, Wüs87a, Wüs87b, Yan88d, Yuk86, vdK85, Kop89, Lóp88, MS86o, Waa89]. theorem-proving [tW87a]. Théorème [Yoe85, Deb89, Jac85c, Kôn85k, Mey88c, PSS86, Ras87]. Théorèmes [Azn88, Gai88b, Ibr85, Jac85a, Dem87a]. Theorems [CS89c, Cec88, Jaj85f, Pre88h, Azn88, Bea88b, Ber88b, Ber85c, Bin89b, Bou88a, Cae88, Cam86, D’O85, Dem88, Dür85b, Ein88, FH89, Faj85, FQ85, Fer88a, Gra85b, He89, Hof87, Hol89a, HPS87d, Iva86, Jaj85c, KZ85a, Kag89b, KS86a, Kat88a, Kaw89, Ken89, Kir89g, Kon85h, Kon85i, LR89b, LMZ86g, Mae87, Mal85a, Mel89, Mha87, MP88, Nak89, Neh87c, Oll85f, Pan85b, Rad87, Rei88b, Rob87, Rod87a, She89a, She87e, Siu85a, Spe88, Vol88f, Vol88e, Vol88g, Waa89, Wu87b, Xin87, vEdG85c, vdBL88, vW86, AS85j, Gyö89]. theoretic [AM88d, EEK88, Far89, FS88b, HPS87e, Koo88, Krá88a, Sim85d, Som86, Sun87]. theoretical [Ber85d, Rek86, Sta88b]. Theorie [Now85, BL86, Hub86, Adj87, AY87, Bar87f, Den86, Gai88a, GP87b, Jeu85, Kré88, Loe87, Mey85a, Pue88, HM87]. Theories [Tau86c, Alm87, Bau87, Bru87d, D’O85, GS85a, Hen85, Ker87, Kuh89, Lai88, Lex88, McC86c, Mod87, Pil87, RC85, Sch89b, She86j, She87b, Žen86]. theorists [Rav88, Tal89]. Theory [All85a, AMSV85, CAXD86h, CKGP87, CCCN85, CCCN87, Com85, DGM86a, DGM86b, Dub88, EMS85, Ieş87e, IN86f, LMS86c, LRT85, LC85f, LMZ86j, MS86c, MS86f, Mül87h, Oll85g, Poh89e, Rol85, She86a, Thr86, Tua86, Voj87e, Wei87u, Woo85, tW87c, vO86, And86a, Asa89, AH86, Bal87, Bar87d, Bar89d, BCZ87, Ber89, BGR85, Bin89b, Bla85a, BB88c, Bok88, BP88c, BF88b, Bor88b, Bor89, Bou87g, Bra87, Bra86i, BS89j, Bru86c, Brü87c, BV88f, BV88o, Bui86b, Bui86d, Bum85, BR85c, But88, CP89a, Car89c, Cav86, Cha86c, Che85, Cho85b, tC87, CB85, CC87b, CC88, Chu88b, Cla86, Coh87a, CT89, Col86, Col88f, Cor85, Cur86, D’A85a, Dav87a, Dic85, DM85c, DLdS87, Dop85, Dra87c, Dwi87, EN88, Ens85, Erd85, EN87, Eva87, FH89, Fad89, Faj85, Fal87, Far89]. theory [Fef88, FT87a, Fel88, Fer85c, Fie88e, Fie88d, Fie88f, FdlP86, Fom88b, FS88b, Fug88a, Gar88a, GL87, Ges86b, GR86, GR89, Gos89, Gra86a, GN86, Gre86d, Gre86e, Gri85a, Gri86c, Gro86e, Gro86g, Grü88b, Hah86, Hai85, Han88, HKL88, Heg88, Hin89, Hod89, Hol89b, Hub86, Hud88, Hur88, IN86b, IN86a, IN86e, JTW89, Jos88, KP87a, KZ87, Kak89, KV89, Ker85, KQ88, Kne89a, Kno89, Koc86, KS85b, Küm88, Lab86b, Lai86, Lan88a, Laz88, Len85, Len86, Lev89k, LMM86, LM86d, LR88b, LMZ86i, Lun86b, Lyc86, Mar88b, Mar87, MC88, Mas86b, Maw86, May86b, McC86a, McC86b, McC85b, Mc 87, McF88, McG88b, Mei86a, Mey87b, Mil85e, Mil88c, Min87, Mit86, Mod86, Mur85b, Nat89a, NR85g, Neu86b, Neu89, Nol86, Nov88e, OD85a, Olv87, Ome89, PlT88a, PlT88d, PW85b]. theory [Phi87c, Phi87e, PS87b, Plo88, Poh89f, Pri89, Que87j, Que87k, Ral87a, Ran87, Ran85, Rav88, Rei85, Rei89c, Rob89, Rod87c, Rof85, Rum89d, Rum89e, Rum89f, Rum86, Sac85, Sae87, Sán85a, Sár89, Sch88h, Sch85d, Sch89f, SS87, lS87d, She87e, Shi86d, SS89a, Sho85, Sim85a, Sim89, Siu85a, Siu86, Sla85a, SS88b, Sri85, Sto86, Sto87b, Str86c, Swe85, Tac86, Tim86b, Tor86, Tri86, Tro88, Tro89a, Vie85, Vog88, Vog85, Voj87c, Vos86b, Wal85a, qW88, Wei87o, Wei87p, Wil88, Win85a, tW87a, Yag87, Zam88b]. Thermodynamics [FL86b, CO86, Isr89, Mül89, Owe86, Rug89, Sew87]. thermoelasticity [Rac88]. thermostat [Prü86]. Theta [Kan85, Kri85g, BH89d, SP89, SV85]. Theta-series [Kri85g]. thetacharacteristics [Cat85a]. thinness [Ess87]. third [AEE85, Gre86a, LX89]. third-order [AEE85]. Thom [Sto85f, LM86f, Lew86b, Mah87, Sto85f]. Thomas [GR86, GR89]. Three [CRD87, NR85e, Sta87d, Tor87a, Bal86c, Bar89a, BW85b, Can85, CH85, Deb89, Eel85, Ens85, Fer88c, Ges86a, GPT89, Gol89, HN86, IN86p, Kre86c, Lev85e, Lev85f, PlT88m, Tor87e, Van87i, Veg85, VG88, Whi86, vdM85h]. three-body [Bal86c, Ens85]. Three-Dimensional [Tor87a, CRD87, BW85b, CH85, HN86, IN86p, VG88, Whi86]. three-line [Ges86a]. three-manifolds [Eel85, GPT89, Veg85]. Three-time-scale [NR85e]. three-web [Gol89]. threefold [Bar89a]. threefolds [Aur89, Con85b, Wei89a]. Thue [Bom87]. Tight [KP85]. tightness [Dow89]. tilting [Tac86]. Time [Föl86, Afa85, Axe86, BE86, Bar85a, Coo86a, Dav87c, DL86b, Ens88, Hir85, Hop86, IK88, IN86i, KV85, Lic85, NR85e, Ne’87, Rek86, Sch89a, SH86, Seg85b, Sol86, Wat88]. time-dependent [Lic85, SH86]. time-derivative [Sch89a]. times [AS89a, BT88a, Coh88a, Fit88, Jac87, JTW89, KM86a, Le 88c, Pin86, Sal85c]. Titchmarsh [Ost85]. Tits [Tim86b]. Toeplitz [BCZ87, CS88b, Kha85]. Tolerance [Sha89]. Tonelli [Cla86]. tools [Pas89]. top [CK85, Her88b]. Topics [Bér86h, Cur86, GT85b, Sol86, Zam88b, Ald85, Bae89, Bér86e, GR85d, Lew87, Mur89, Sch89f]. Topological [Bli88, Bor89, HN88, Kaw86a, Lyc86, PR85a, Ped85, Tal89, Win89, BS89i, Bor88b, Hüb89b, Hüb89d, LMZ86q, Sat86, Tur88b, Web88, Žub88, Mat87e]. Topologie [Mey85b, Del88]. topologies [LMZ86d, LMZ86e, LMZ86l, LMZ86m, LMZ86o, Mat86b, MN88e]. Topologische [Mat87e]. Topologischen [Tic87]. Topology [AH85a, AMSV85, Jam89, LMZ86j, PlT88k, RLQ85, Roo86b, SSS86, Smi85b, Whi89, AM88b, Arn88b, CC88, GM87b, Khi86, LMZ86i, Mor88, She86o]. Toponogov [BEMG85]. tops [Sap88]. toral [AG89, Not88, Str89a]. tordus [Val86a]. Torelli [Deb89]. tori [DS89, Fom88b, KI87, Sev86d, Sev86f]. Torsion [CS85a, Möh87, And86b, Cat85a, Des85, DR89b, Ill86, Lex88, Ran85, Tor86]. Torsion-free [Möh87]. torsions [Ino89]. Torsionsfreie [Knö87]. torsionsfreier [OS85]. torus [AP85b, AP86, GM87b, Mos89]. total [BC86a, Sch89f]. totally [DD85, NR85f, She89b]. Totient [Han85]. tournament [Tab86]. Trace [Fis87a, Gru86a, Hüb89e, Wid85i, Bar86c, EK88a, Fis87d, Fis87f, Kum85, dlH85]. tracée [GS87]. Traces [Hüb89f, Hüb89h, Hüb89g]. track [HJ89]. Trägheitsgruppen [Sto85o]. trait [Kim86]. trajectories [Poz88, Sev86a]. Trajectory [vEdG85g, vEdG85a, vEdG85b]. transcendence [Bro87, BT87c]. transcendency [Gos89]. Transcendental [Ehr85, Oka85]. transfer [LM86g, LR88c]. transform [Bru89, FR85, FJ88, Koo85, Mur88c, Ric86]. Transformation [LT87, PlT88l, Ade89, BM87, DB87, Hsi87, cH89a, MS86a, Ofm86b, Ofm86c, Ofm86a, ŌT87, Wag87, JP86]. Transformations [Bak85a, Bak85b, AP88, Bak87b, Dun87, Gro87c, HR88a, Kif88b, Qua88b, VJ85]. Transformée [Ofm86e]. transforms [All85b, BB88b, Dur85c, Kuo87, Lus87, Rub86, CM87c]. transient [BH89b, Gal86, Sei89b]. transients [Gal86]. transition [Acc89, Yan88c]. transitions [PR85b, Tor87d]. transitive [HR88a, Mei86b, Tab86]. transitivity [San86]. translates [LL89, MC89]. translation [Als88, AD85a, Gli86, MMA86, PR85a]. translation-invariant [Als88]. transmission [Gru87]. transmutation [Car86e]. Transonic [Jam85, BFKP86, Neč86, vdW86]. transport [App88]. transversal [LV85]. Transversality [Mil85c, Spe88]. transversally [KRT86]. Trapping [ABC^{+}85, CBD87]. Travelling [Bri85b]. treating [Rek86]. Treatment [BW85b, GWW85, HGYR89, Nat86, Pap85, RZ86, Ste85a, Whi86]. tree [Con86, RT89]. tree-class [Con86]. trees [MS85, Sch89d, Ung86, Wan87d, Wan87h]. Trento [DT87]. trialité [Cru87]. triangle [Rob88]. triangular [Kří86, Yan88e]. triangulated [Lev89g]. triangulation [MS86a]. triangulations [DK85b]. Tribus [FL87]. Triebel [FHJW89]. Trigonometric [NS85, Ber88c]. trimmed [HKW89]. triple [PS85]. tripotents [Neh87d]. tritangencies [Oza85]. trivial [DL88]. Triviality [Son85]. trois [Deb89]. Truncations [HZ87b]. trust [Noc86]. truth [CB85]. Tschirnhaus [Gro87c]. Tsirelson [CS89a, CS89d, CS89e, CS89f, CS89h, CS89g, CS89j, CS89k]. tubular [Che89, HR86, Pin86]. tumour [HGYR89]. tunnelling [WKK86]. Turan [LR88b]. turbine [Alt88]. turning [GH85b, Pre88e]. twist [Lit85]. twist-spun [Lit85]. Twisted [LMS86b, Don86a, EK88a, EPS87, KSX87, Pie85]. Twistor [Bur87, Raw87, Woo87a, Raw85]. twists [Hej88]. Two [BA85, CS87a, Dau88g, Dra87d, Eke86a, Eke86b, Fad89, Hof87, Imk88f, Nev87, Par87a, Sch88i, Sei87, dG85, vEKM89, Ani89, AK87, AB87, BSK87a, BM88, BT86a, BCSSC87, Bor88a, Coh89, CG86b, CL89b, Die88a, Din85a, Don88b, EM86, Ens85, Eva85a, Fab89, Fuk88b, Gol89, Har87d, Hei87, Hig87, IN86q, Kör87b, Kuc87, Lab86a, LW88, LW89, LM89b, MR89b, MH88b, NR85f, Net88a, Ōsh87, PS88b, Rei88b, Shi86u, Sto87a, Šve86, Tie88, Tor87e, Veg85, Whi86, Wu87a, Yui85]. two- [Ens85, Whi86]. Two-dimensional [Dau88g, AK87, Bor88a, Eva85a, Gol89, Har87d, Hig87, IN86q, Kuc87, PS88b]. two-level [LM89b]. Two-parameter [Imk88f, Veg85]. two-parameters [Kör87b]. Two-point [Dra87d, EM86]. Two-sided [Par87a, BT86a, Shi86u]. two-sphere [Don88b]. two-valued [Coh89]. Typ [Wei86i]. Type [MS86t, Tit86, tW87c, Ass89, Aus85, Bar86c, ByC86, Bar89b, Bau88a, Ben87, BF88a, Ber88b, Ber87c, Boh88, Bon85, BHT89, Cai85, Cau86a, CHS85, CS88c, D’A87a, Deb87, DMR87, Dho85, Dra87d, DS86b, DST89, Dwi87, DM87b, Fac89, Fal86, Fin88a, Fu88, Geb86, Gon88, Goo85, HPY89, Hei85, Hu87, Ill89, KL87b, Kob87, McC89, Mei86b, MTJ87, Mor89, Mul86, Ngu85, Pie88a, Que87h, Rog86, Ron86, Sch87j, Shi86c, Shi85, Tor87d, Tor87f, Tru86, Wei87e, Xin87, Zal88, Zay87, vdB87c, vdB87b, Phi87i]. Types [RC85, Tor87e, CCS^{+}87, Die85a, HN89, Hru87, IN86g, Kod86, Shi86b, Shi86s].

U.S [CKGP87]. Uhlenbeck [Had85, Jac85b]. Uhlenbeck-processes [Had85]. Ulam [FH89, Nak89]. Ultrafilters [Hin89, Bla85b, PR85c]. Ultraproduct [Ocn85k]. Ultraproducts [Mak85]. Ultraproduits [Zay87]. Ultrasingularities [RR87]. umbilical [NR85f]. umbrella [BW88]. UMD [Gar86a]. UMD-spaces [Gar86a]. Umkehrproblem [Mat87a]. Unary [Dwi85]. Unbounded [CS88a, vEdG85g, Eng89, FW89, GR85a, GR85b, GR85g, GR85i, Gau87, Hil85b, Ich87, LN87, MR85b, Pic85, Van87e, Vol88d, Vol88g]. unboundedness [Dur85c]. uncertainty [BB85, DB87, Sar85]. uncomplete [Lio86]. Unconditional [CS89k]. unconstrained [Kiw85a, Kiw85g]. uncorrelated [CMM89]. uncountable [Row89]. unequal [ESN89]. unicity [Dun88]. unidimensionnels [Reu86a]. unification [Ne’87]. unified [FG88, Via89]. Uniform [AS85m, CC87a, Gau87, GLR86, Kre86a, MR87, Bou87f, Din85b, Dun88, Has85, HJ85, Jaj85a, Ken89, LM86b, Ton87b, Tur89, vdB87b]. uniformity [EEŚ85]. Uniformization [KP86l, Kec88a]. uniformly [Led86a]. unilateral [Dal86, Kuč86a, Rip87]. unimodal [Hel85]. unimodular [Sch87b]. unions [Bar88b]. Unipotent [Bar87a, Tit87, HZ87a, Mül87m, Pom87]. Unique [Wal86c, Ken89, MS86a, dlH85]. uniquely [BT87b, FKZ86, Ost85]. Uniqueness [BD86b, PS88c, ABB86, Ber88a, CV87, DMW87, Gaj86, Kis89, Sta87c, Wri89a, Wu88b]. uniserial [She86h]. unit [Bus88b, CS89g, CS89j, Ino89, LS87c, MP89b, Mut86, Tho87]. unit-regular [Bus88b]. unitaires [MVW87f]. unitarity [Bar87a]. Unitary [DLdS87, FA89, AHKHK89, BSK87a, GvW89, Hid89, Kam89a, Que87k, Rei89j]. unité [AAGV85]. units [LN89]. unity [Lam88, SS85d]. univalence [Koe88]. univalent [Aha87, CM87c, Fit87a, Fuk88a]. Universal [AC85, HLM85a, Sch86f, She87d, Ale85, FT87b, Gla85, HKW89, HLM88, Com85]. universality [GK88a]. universe [LM86c]. universelles [Bea86b]. Université [JY85, LL86]. University [AH85a, CCCN85, CCCN87, DH85b, Hop89a, RLQ85]. Unobstructed [MR89c]. Unramified [Far87b]. unsolved [Mei86a]. Unstable [Lai86, CGMM87]. Untergruppenverbände [Sch87f]. Untersuchungen [Kir85b, Kir87b, KP87b]. Untwisted [Kan89]. update [Hin89]. upper [HGYR89, ST89i]. ups [Har88b]. Upwind [Roe87, Spe86]. Urysohn [Mil87b]. USA [BDH^{+}85, KS85a, RLQ85]. Usco [Phe89f]. use [Dau87, Spe86, Tay86]. useful [Lin88b]. Using [Kee87, All85b, BCSSC85, Bra86a, Mét87, Nev87, dLQ89]. USSR [KZ85b]. utilizing [BT86b]. Uzhgorod [KZ85b].

V [BDH^{+}85, Nor86a, Jar85g]. V. [Pop87a]. valeur [Pas87a]. Valeurs [Han89, Fer89]. valid [Axe86]. Valuation [Mat86b, Die85a, Rou87]. Valuations [Coh88b, Mat86c, Mat86d, Bru88, Rum85]. Value [Kon85j, Sto86, Voj87e, Aga86b, AL88a, BJ86, Col86, Dau88i, Dau88h, DB87, Dra87c, Emb87, EM86, Erk87, Fab88, Fla87, Gon88, Gre86a, Gru87, HW88, HSW85, Jut89, Kod86, Kra88b, Kuf86, Laz88, Mar88a, MP88, Oba89, Obr89, Ome89, Pon88, Rac88, Sch85c, Sch89b, Sch86c, Šed86, Sho85, Siu86, Ste86b, SU87, Tru86, Wit88a, Wu88a, YM89, ZG88, Zvy88]. value-distribution [Dra87c]. Valued [Mat86e, AHKH86, BL85, Bor85, Bou88c, Bré89b, ČW87, Coh89, Daw86a, Fer85c, Gel86, IN86n, IN86j, IN86l, Kal86a, Kal86b, Kat89b, Kog86e, Łuc89f, Mét87, Mod87, Ost85, Rub86, SW89c]. values [Alb85, CC85a, Fom88b, HS85, Hay88, KÜ88, Mas87]. Vancouver [Rol85]. Vanishing [Dem88, Ein88, Rob87, Bor87a, Che86a, Dav87d, Ebe87e, Jos85a, LM86a, Ocn85c, Siu85a, SZ87]. Varenna [LP86]. variable [CGC88, Gar88b, Mur88c, Rei85, Rey86a, Sta85]. Variables [Nor86a, Alt89, Car86d, DGM88, DG86b, Fac89, Fug86, Går87b, GG85, HS89a, KZ87, KMR89, Kon85b, Kor87c, Kra87a, Kru85, LS85a, LeP89a, Mae87, Mel89, Roc88, Roo86a, Shc85]. variance [Bou87a, Let89]. variant [Xu87b]. Variation [Tho85h, Alb85, El 87, Imk88e, Kas87, Kaz89, Nag86c, Qua88a]. Variational [CM89b, Dau88i, Dau88h, Nir89, RM87, Sch89f, Sho85, Bah88, Bar85a, Eke88, FW89, Mas85, Mos88, Nag86a, Oll85c, Phe89j, Por85, Sap86, Sim85c, Str88b, Thi86, Tro89a]. variationnelle [Éme88]. Variations [AJBL89, ABB85, CS89l, CKS87b, Cla86, HZ87a, Har85a, Lev89l, Mas88a, Oco88, PlT88c, Pic88]. Variationsrechnung [Hla87a, Tic85]. Varietäten [Bar85b, Fal85]. variété [Cou89, GS87, LT85, LT87, Sch85j, Var86b]. variétés [Ala88, Bak87b, Bak88, Dre87, Gil89b, Var86a]. Varieties [Bui86c, CS87a, EEK88, GI85, Mum88x, Ben86a, BS89d, CH89b, DH88, Ein88, Eke86a, HR88b, Jon88, Kan85, Kis85, LW88, Mas87, Mor86d, Mum88e, Mum88f, Mum88u, Nat85a, Pat87, Pix85, Rob88, Sch89e, Sil89a, Sil89b, Sil89d, Som86, Str87a, sT89a, Wu86, vdG89]. variety [Car85b, EPS87, LW89, Wei87v]. Various [Wan87h, HW88]. Varna [KPZ87]. varying [HL88b, Lóp88, jQxQ88]. Vector [Bou88c, Rei89f, ST88, Tor87a, App88, BW88, Bli86, BG89d, Cam88, CS89g, CS89j, Che86d, Dem88, Dum85b, ES87b, Got88, GZ86b, HH86, HS86, Kal86a, Kal86b, Kog86e, Kog86g, Mar86c, Mat86e, MH88a, Mod87, Nar87, Rao87, Rei89g, RW85, SB88, Sha87, SW89c, Tor87c, dM88]. Vector-valued [Bou88c, Kal86b]. vectorfield [Sev86c]. vectorfields [Bra86k, Bry89a, Sev86a, Sev86d, Sev86f, Sev86g]. vectoriel [Mar85]. vectorielles [Alt89, Fer89]. vectors [Jos87a, Jos87b, LeP89b, Lis89, Sán85a, SW87, dA85b]. Venant [Ieş87c, Ieş87d]. Venezuela [Di 85]. Ventsel [BCM88]. verification [KSX87]. Veronese [Bac86, Hul85]. Vers [Kai87c, AD85b]. Verschlingung [tDL85]. Verschlingungszahlen [tDL86]. version [BR85a, Cae86, LC85a, Str85a, ST89n, Wit88a]. versions [Owe86, Waa89]. versus [EEK88]. vertex [BW85b]. vertical [Ulr88a, Ulr88b]. Verwandte [KPT85]. VI [DS89, JY85]. via [And86a, Cha85a, FG88, Fis87a, Fra89, Gri86c, Gro87c, MR87, McC86b, Mor89, RS85a, Sch88h, Tur88a, tW87c]. vibrating [Fil85]. Vietoris [CGC88]. view [Bér86f, Dre85]. Viewing [Noc86]. viewpoint [KP86b]. views [Pie88b]. vignettes [Fad89]. VII [Gal89]. VIII [Hey86]. Virasoro [BMS87, Bry89a]. Virtual [Fie88j]. viscoelastic [Bol88]. Viscosity [Lio89, Eva87, KK89a, Lio85, Tad87]. viscous [GR85d, Neu86c]. VMO [MS89c]. Vogan [BSK87b]. voisinage [Tho88]. Vojta [Lan85b]. Vollständig [Gol87b]. Vollständige [Gol87c]. Volterra [MH88d]. Volume [Mil85e, AV88a, CMS89, GV87, Gro87a, LM86a, Mil87b]. Volumes [Bal89, Hol89c, Gra85b]. vom [Wei86i]. vortex [Gar88a, Gus87]. voter [CG86b]. Vries [Sch86j, Sch86l, Sch86m, Sch86o, Sch86p].

Wadge [LSR88]. Wahba [Pic89b]. waiting [Afa85]. waiting-time [Afa85]. walk [Han87h]. walks [AW89, Ger86b, Iva86, KS89, Sch86g, Sto87a, Sul86, Woe86]. wall [And86b, Kir89q]. warm [AP89]. Warsaw [Ław85]. Washington [GZ87b]. Watanabe [Lin86a]. water [CS86, DF86b, ZG88]. Wave [Bal86c, Går87j, Går87k, Alb85, Ant85, Avr87, Bri85b, Eli88, Eng87a, Går87i, IN86p, Joh89a, Lau87a, tLmC89, Men88b]. Waves [Pap88b, BM88, Bea87, CS86, Kre85, MR89b, RR87, Ros89a]. way [Shn86]. we-2 [Nag88]. Weak [Col88k, JS88, KPA89, KM88, Mét87, Tur89, AFL89, Bou88d, Din85b, God87a, Kru85, LS87e, MV88, Neu86c, Num89, Sch86s, Sim86]. Weak* [RS85b]. Weak*-denting [RS85b]. Weakley [ADR85]. Weakly [BR88, GK88b, AO88, Bue87, DG86b, GK87b, JM85, Kne89a, Ros89a]. web [Gol89]. webs [Gol86a]. Week [EMS85]. Weier [Ton89]. Weierstrass [BC86f, Mag86]. Weight [BS87a, CS87a, SP89]. Weighted [GR85g, GR85i, GR85h, LS88r, Oht88, Dud85, GL87, GR85k, HS87b, Kuf86, Lig85b, LS88c, LS88d, MMT87, Sch87k]. Weights [ST89m, KL87b, Lub85, LS87f, Sel88e, ST89e]. Weil [Frö85, Kai87c, Kai87b, Mor88, Rei89h, Rei89g, Rei89i, Rei89j]. Weingarten [PlT88m]. Weinstein [Aki89]. Weisner [Ber85b]. Weiss [Mey85a]. well [BT86a, Hej85]. well-balanced [BT86a]. well-known [Hej85]. wells [Hel88i, Poz88]. Welters [AD85a]. Wert [Sch85g]. West [AvW85, ABS86, ABS87, EMS85, GWW85, RR85]. Weyl [DR87, HB89, LL85, Ros87b, Shi86c, Shi86g, Shi86j, Shi86m, Shi86n, Shi86t, Shi86u, Wei87r]. where [Jac86a, JJW87, Tho86b, vdB87c]. which [BR88, BH89a, BT87b, CFdG87, CS89e, DL87, DH86, ESN89, Fit87a, Gor88, Il’85, Kes88a, Lau87b, McN85, Mer86, Pat88, SS86, Sch87c, Ste85c, Tho86b, dlH85]. White [Sch89a]. Whitehead [AM88e, AM88f, AM88g, And86b, Ani86, Bau85, DR86, DR89b, HJ89, Ill86, Ino89]. Whitney [BW88]. whole [Lub85]. whose [Ced88, Cib86, Dum85b]. Wichura [Dud85]. Wick^{≫} [dC85a]. width [Gro88]. widths [BG89c]. Wielandt [Sco87]. Wiener [Hla85d, Hla87b, Bir85, Föl86, Gar88b, HM88c, KÜ88, Kus88, LM89a, MY87a, Mey88e, MvR88, Mos88, NÜ89, Sch89a, Wsc85a, Yan87, dC85b]. Wigner [DMR87]. wild [Bae86, CW85]. Willmore [Kob87]. Wilson [Mag88]. winning [Tor87b, Tor87f]. Winter [dPR86]. within [Ros87a, Ste85a]. without [BBGM89, BD86b, Csö89, DH89, Pet89g, Pup88, Sán85a]. witnesses [EP87]. Witt [Reu86b]. Witten [Hel88f]. word [Ber87a, Gla85]. words [BT86a]. work [Aoy86, Bor86, Dic88, DHS85, Kir89i, Roc88, Rub89, Whi89]. works [tW87a]. Workshop [AvW85, AW86b, BBT85, CKGP87, Hen86b, LS86a, Sel89b, Thr86]. world [Bra87]. Worpitzky [JTW89, Waa89]. would [Erd85]. wreath [FK88]. wtt [FS85a]. wtt-degrees [FS85a].

X [Mey89b, Yan88b]. XI [Mey89c]. XIII [AIJH85]. XIV [Hen86c]. XIX [AY85]. XX [AY86]. XXI [AYM87].

Yamabe [Aub86, GM86, Jin88b]. Yang [Bou87g, Don89a, Hen85, Sán85a, lS87d, Xu87b]. Year [AH85a]. Yoneda [BR86]. Yor [Dav87c]. York [CCCN85, CCCN87, Hop89a]. Young [Ber88b, Bou88c, Gar86b, GB86, Tho85c, dSCV86].

Zadeh [Maw88]. Zahlen [KPT85]. Zahlentheoretische [Hla85d, Hla87b]. Zahlentheorie [Bar87d]. Zakai [Ben89, Hau85, Lio89]. Zakharov [Sch86q, Sch86r]. Zaragoza [BS86]. Zariski [Hir89b]. Zehnder [Cha85b]. Zero [Van87h, Wal87, dP86, AD85b, Eke86a, Emb87, GLC86, Lan86, PS89a, AD85b, Rou87]. Zero-free [Wal87]. Zeros [Shi88f, BDG88, DGR87, Hel87, HvR88, Kre86a, LS88k, LS88r, Mul86]. Zeta [Fis87a, Fel88, Fis87c, Gek86j, Le 88a]. zeta- [Le 88a]. Zeta-Function [Fis87a, Fis87c]. Ziffernsumme [KPT85]. Zolotarev [Pop87a]. zonal [Mac87]. zones [Sev86g]. zonoids [MP89b]. Zufällige [Buc85]. Zum [Hla85a, Hla85c]. Zur [Kir85b, Now85, Col86, Kir87b, Sto85i, Tur87, KPS87, Sto85p]. Zygmund [CDMS89].

References

Andrews:1985:COP

[AA85] George E. Andrews and Richard Askey. Classical orthogonal polynomials. Lecture Notes in Mathematics, 1171:36–62, 1985. CO-DEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076530/.

Abd-Alla:1986:DRB

[AA86] Mohamed Abd-Alla. Domaines de Reinhardt bornés homogènes de E(s,C). (French) []. Lecture Notes in Mathematics, 1198:1– 34, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077040/.

Ayupov:1989:INA

[AA89] Sh. A. Ayupov and R. Z. Abdullaev. On isometries of non associative L_{p}-spaces. Lecture Notes in Mathematics, 1396:99–106, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083547/.

Alfaro:1985:CES

[AAGV85] M. Alfaro, Ma. P. Alfaro, J. J. Guadalupe, and L. Vigil. Correspondance enter suites de polynômes orthogonaux et fonctions de la boule unité de H_{O}^{∞}(D). (French) []. Lecture Notes in Mathematics, 1171:158–163, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076540/.

Archbold:1985:FSA

[AB85] R. J. Archbold and C. J. K. Batty. Factor states on C∗-algebras. Lecture Notes in Mathematics, 1132:12–16, 1985. CODEN LN-MAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074875/.

Ashbaugh:1987:RFT

[AB87] Mark S. Ashbaugh and Rafael Benguria. On the ratio of the first two eigenvalues of Schrödinger operators with positive potentials. Lecture Notes in Mathematics, 1285:16–25, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080577/.

Accardi:1989:CLS

[AB89] L. Accardi and A. Bach. Central limits of squeezing operators. Lecture Notes in Mathematics, 1396:7–19, 1989. CODEN LN-MAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083541/.

Abadie:1986:GRG

[Aba86] J. Abadie. Generalized reduced gradient and global Newton methods. Lecture Notes in Mathematics, 1190:1–20, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076700/.

Austin:1985:FIR

[ABB85] D. Austin, A. Bellow, and N. Bouzar. The Fatou inequality revisited. — variations on a theme by A. Dvoretzky. Lecture Notes in Mathematics, 1153:17–39, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074944/.

Alexander:1986:CUR

[ABB86] Stephanie B. Alexander, I. David Berg, and Richard L. Bishop. Cauchy uniqueness in the Riemannian obstacle problem. Lecture Notes in Mathematics, 1209:1–7, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076617/.

Albeverio:1985:TSM

[ABC^{+}85] S. Albeverio, Ph. Blanchard, Ph. Combe, R. Rodriguez, M. Sirugue, and M. Sirugue-Collin. Trapping in stochastic mechanics and applications to covers of clouds and radiation belts. Lecture Notes in Mathematics, 1136:24–39, 1985. CODEN LN-MAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074456/.

Abels:1987:CPC

[Abe87a] Herbert Abels. Compact presentability and contracting automorphisms. Lecture Notes in Mathematics, 1261:15–26, 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079710/.

Abels:1987:FLA

[Abe87b] Herbert Abels. Filtrations of Lie algebras and groups. Lecture Notes in Mathematics, 1261:27–48, 1987. CODEN LN-MAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079711/.

Abels:1987:FPA

[Abe87c] Herbert Abels. Finite Presentability of S-Arithmetic Groups Compact Presentability of Solvable Groups, volume 1261 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 178 pp. LCCN QA3

.L28 no. 1261. URL http://link.springer.com/book/10.1007/
BFb0079708;   http://www.springerlink.com/content/978-
3-540-47198-1.

Abels:1987:INC

[Abe87d] Herbert Abels. Implications of the necessary condition. Lecture Notes in Mathematics, 1261:61–89, 1987. CODEN LN-MAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079713/.

Abels:1987:I

[Abe87e] Herbert Abels. Introduction. Lecture Notes in Mathematics, 1261: 1–14, 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079709/.

Abels:1987:NCC

[Abe87f] Herbert Abels. A necessary condition for compact presentability. Lecture Notes in Mathematics, 1261:49–60, 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079712/.

Abels:1987:AG

[Abe87g] Herbert Abels. S-arithmetic groups. Lecture Notes in Mathematics, 1261:123–145, 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079715/.

Abels:1987:ASG

[Abe87h] Herbert Abels. S-arithmetic solvable groups. Lecture Notes in Mathematics, 1261:146–168, 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079716/.

Abels:1987:SH

[Abe87i] Herbert Abels. The second homology. Lecture Notes in Mathematics, 1261:90–122, 1987. CODEN LNMAA2. ISBN 3-540-17975-5 (print), 3-540-47198-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079714/.

Albeverio:1985:QML

[ABGS85] S. Albeverio, Ph. Blanchard, F. Gesztesy, and L. Streit. Quantum mechanical low energy scattering in terms of diffusion processes. Lecture Notes in Mathematics, 1109:207–227, 1985. CO-DEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0101546/.

Albeverio:1985:RNL

[ABHK85] S. Albeverio, Ph. Blanchard, and R. Høegh-Krohn. Reduction of non linear problems to Schrödinger or heat equations: Formation of Kepler orbits, singular solutions for hydrodynamical equations. Lecture Notes in Mathematics, 1109:189–206, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101545/.

Almgren:1988:CAL

[ABL88] F. Almgren, W. Browder, and E. H. Lieb. Co-area, liquid crystals, and minimal surfaces. Lecture Notes in Mathematics, 1306: 1–22, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082921/.

Albeverio:1986:SPM

[ABS86] Sergio A. Albeverio, Philippe Blanchard, and Ludwig Streit, editors. Stochastic Processes — Mathematics and Physics: Proceedings of the 1st BiBoS-Symposium held in Bielefeld, West Germany, September 10–15, 1984, volume 1158 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1158. URL http://link.springer.com/book/10.1007/
BFb0080206;   http://www.springerlink.com/content/978-
3-540-39703-8.

Albeverio:1987:SPM

[ABS87] Sergio Albeverio, Philippe Blanchard, and Ludwig Streit, editors. Stochastic Processes — Mathematics and Physics II: Proceedings of the 2nd BiBoS Symposium held in Bielefeld, West Germany, April 15–19, 1985, volume 1250 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1250. URL http:

//link.springer.com/book/10.1007/BFb0077343;   http://
www.springerlink.com/content/978-3-540-47835-5.

Adams:1985:UTP

[AC85] M. E. Adams and D. M. Clark. Universal terms for pseudocomplemented distributive lattices and Heyting algebras. Lecture Notes in Mathematics, 1149:1–16, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098451/.

Antolin:1987:SAF

[AC87] J. Antolin and A. Cruz. A Stieltjes analysis of the K^{+} − p forward elastic amplitude. Lecture Notes in Mathematics, 1237:251– 269, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072469/.

Accardi:1988:NMN

[Acc88] Luigi Accardi. A note on Meyer’s note. Lecture Notes in Mathematics, 1303:1–5, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078050/.

Accardi:1989:CTE

[Acc89] Luigi Accardi. Cecchini’s transition expectations and Markov chains. Lecture Notes in Mathematics, 1396:1–6, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083540/.

Arad:1985:GSC

[ACM85] Zvi Arad, David Chillag, and Gadi Moran. Groups with a small covering number. Lecture Notes in Mathematics, 1112:222– 244, 1985. CODEN LNMAA2. ISBN 3-540-13916-8 (print), 3-540-39142-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072289/.

Andreka:1985:COR

[ACN85] H. Andréka, S. D. Comer, and I. Németi. Clones of operations on relations. Lecture Notes in Mathematics, 1149:17–21, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098452/.

Albeverio:1985:SAC

[ACSC85] Sergio Albeverio, Philippe Combe, and Madeleine Sirugue-Collin, editors. Stochastic Aspects of Classical and Quantum Systems: Proceedings of the 2nd French–German Encounter in Mathematics and Physics, held in Marseille, France, March 28–April 1, 1983, volume 1109 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1109. URL http:

//link.springer.com/book/10.1007/BFb0101532;   http://
www.springerlink.com/content/978-3-540-39138-8.

Arbarello:1985:ATC

[AD85a] Enrico Arbarello and Corrado De Concini. An analytical translation of a criterion of Welters and its relation with the K.P. Hierarchy. Lecture Notes in Mathematics, 1124:1–20, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074993/.

Azencott:1985:LSQ

[AD85b] R. Azencott and H. Doss. L’Équation de Schrödinger quand h tend vers zéro; une approche probabiliste. (French) [Schrödinger’s equation when h tends toward zero: a probabilistic approach]. Lecture Notes in Mathematics, 1109:1–17, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101533/.

Adams:1988:NCI

[Ada88] David R. Adams. A note on Choquet integrals with respect to Hausdorff capacity. Lecture Notes in Mathematics, 1302:115– 124, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078867/.

Adem:1989:HRF

[Ade89] Alejandro Adem. Homology representations of finite transformation groups. Lecture Notes in Mathematics, 1370:15–23, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085215/.

Adib:1987:EFM

[Adi87] Khalid Filali Adib. Étude des fonctions méromorphes sur certains espaces fibrés. (French) []. Lecture Notes in Mathematics, 1295:59–95, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081978/.

Adjamagbo:1987:STD

[Adj87] Kossivi Adjamagbo. Synthèse de la théorie du determinant sur un anneau finement filtré. (French) []. Lecture Notes in Mathematics, 1296:1–27, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078521/.

Adomian:1987:ASO

[Ado87] G. Adomian. Analytical solutions for ordinary and partial differential equations. Lecture Notes in Mathematics, 1285:1–15, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080576/.

Aron:1985:WCF

[ADR85] R. M. Aron, J. Diestel, and A. K. Rajappa. Weakley continuous functions on Banach spaces containing ℓ_{1}. Lecture Notes in Mathematics, 1166:1–3, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074685/.

Abou-El-Ela:1985:SRS

[AEE85] A. M. A. Abou-El-Ela. A stability result for the solutions of a certain system of third-order differential equations. Lecture Notes in Mathematics, 1151:1–8, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074709/.

Accardi:1988:SI

[AF88] Luigi Accardi and Franco Fagnola. Stochastic integration. Lecture Notes in Mathematics, 1303:6–19, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078051/.

Afanaseva:1985:PDW

[Afa85] L. G. Afanas’eva. On periodic distribution of waiting-time process. Lecture Notes in Mathematics, 1155:1–20, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074810/.

Accardi:1989:WCL

[AFL89] L. Accardi, A. Frigerio, and Yun Gang Lu. On the weak coupling limit problem. Lecture Notes in Mathematics, 1396:20–58, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083542/.

Arazy:1987:MIS

[AFP87] J. Arazy, S. Fisher, and J. Peetre. Möbius invariant spaces of analytic functions. Lecture Notes in Mathematics, 1275:10–22, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078341/.

Arazy:1988:BNR

[AFP88] J. Arazy, S. D. Fisher, and J. Peetre. Besov norms of rational functions. Lecture Notes in Mathematics, 1302:125–129, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078868/.

Alamo:1989:STA

[AG89] N. Alamo and F. Gómez. Smooth toral actions on principal bundles and characteristic classes. Lecture Notes in Mathematics, 1410:1–26, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086408/.

Agaian:1985:AHM

[Aga85a] S. S. Agaian. Application of Hadamard matrices. Lecture Notes in Mathematics, 1168:134–177, 1985. CODEN LNMAA2. ISBN 3-540-16056-6 (print), 3-540-39740-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101077/.

Agaian:1985:CCH

[Aga85b] S. S. Agaian. Construction of classic Hadamard matrices. Lecture Notes in Mathematics, 1168:11–102, 1985. CODEN LN-MAA2. ISBN 3-540-16056-6 (print), 3-540-39740-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101075/.

Agaian:1985:CGH

[Aga85c] S. S. Agaian. Construction of generalized Hadamard matrices. Lecture Notes in Mathematics, 1168:103–133, 1985. CODEN LN-MAA2. ISBN 3-540-16056-6 (print), 3-540-39740-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101076/.

Agaian:1985:HMT

[Aga85d] S. S. Agaian. Hadamard Matrices and Their Applications, volume 1168 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16056-6 (print), 3-540-39740-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 227 pp. LCCN QA3 .L28 no. 1168. URL http:

//link.springer.com/book/10.1007/BFb0101073;   http://
www.springerlink.com/content/978-3-540-39740-3.

Agaian:1985:I

[Aga85e] S. S. Agaian. Introduction. Lecture Notes in Mathematics, 1168: 1–10, 1985. CODEN LNMAA2. ISBN 3-540-16056-6 (print), 3-540-39740-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101074/.

Agarwal:1986:PC

[Aga86a] A. K. Agarwal. Partitions with ‘N copies of N’. Lecture Notes in Mathematics, 1234:1–4, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072504/.

Agarwal:1986:GMC

[Aga86b] R. P. Agarwal. On Gel’fand’s method of chasing for solving multipoint boundary value problems. Lecture Notes in Mathematics, 1192:267–274, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076080/.

Albeverio:1986:SOP

[AGHK^{+}86] S. Albeverio, F. Gesztesy, R. Høegh-Krohn, H. Holden, and W. Kirsch. The Schrödinger operator for a particle in a solid with deterministic and stochastic point interactions. Lecture Notes in Mathematics, 1218:1–38, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073042/.

Agmon:1985:BED

[Agm85] Shmuel Agmon. Bounds on exponential decay of eigenfunctions of Schrödinger operators. Lecture Notes in Mathematics, 1159: 1–38, 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0080331/;   http://link.springer.com/content/
pdf/bfm:978-3-540-39706-9/1.pdf.

Alexander:1985:GTP

[AH85a] James C. Alexander and John L. Harer, editors. Geometry and Topology: Proceedings of the Special Year held at the University of Maryland, College Park 1983–1984, volume 1167 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1167. URL http://link.springer.com/book/10.1007/
BFb0075212;   http://www.springerlink.com/content/978-
3-540-39738-0.

Arad:1985:PCC

[AH85b] Zvi Arad and Marcel Herzog. Products of Conjugacy Classes in Groups, volume 1112 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-13916-8 (print), 3-540-39142-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 244 pp. LCCN QA3 .L28 no. 1112. URL http:

//link.springer.com/book/10.1007/BFb0072284;   http://
www.springerlink.com/content/978-3-540-39142-5.

Avramov:1986:TLG

[AH86] Luchezar Avramov and Stephen Halperin. Through the looking glass: a dictionary between rational homotopy theory and local algebra. Lecture Notes in Mathematics, 1183:1–27, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075446/.

Azema:1989:PRP

[AH89] J. Azéma and K. Hamza. La propriété de représentation prévisible dans la filtration naturelle d’un ensemble régénératif. (French) []. Lecture Notes in Mathematics, 1372:131–138, 1989. CODEN LN-MAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083963/.

Aharonov:1987:BTG

[Aha87] Dov Aharonov. Bazilevič theorem and the growth of univalent functions. Lecture Notes in Mathematics, 1275:1–9, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078340/.

Albeverio:1985:RDS

[AHK85] Sergio Albeverio and Raphael Høegh-Krohn. A remark on dynamical semigroups in terms of diffusion processes. Lecture Notes in Mathematics, 1136:40–45, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074457/.

Albeverio:1986:SLG

[AHKH86] Sergio Albeverio, Raphael Høegh-Krohn, and Helge Holden. Stochastic Lie group-valued measures and their relations to stochastic curve integrals, gauge fields and Markov cosurfaces. Lecture Notes in Mathematics, 1158:1–24, 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080207/.

Albeverio:1989:CFK

[AHKHK89] S. Albeverio, R. Høegh-Krohn, H. Holden, and T. Kolsrud. A covariant Feynman–Kac formula for unitary bundles over Euclidean space. Lecture Notes in Mathematics, 1390:1–12, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083931/.

Aldous:1985:EEP

[AIJH85] D. J. (David J.) Aldous, I. A. (Ildar Abdulovich) Ibragimov, Jean Jacod, and Paul Louis Hennequin, editors. École d’Été de Probabilités de Saint-Flour XIII — 1983. (French) [Saint-Flour Summer School on Probability XIII — 1983], volume 1117 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CO-DEN LNMAA2. ISBN 3-540-15203-2 (print), 3-540-39316-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1117; QA3 .L28 no. 1117 QA273.18. URL http:

//link.springer.com/book/10.1007/BFb0099420;   http://
www.springerlink.com/content/978-3-540-39316-0.

Andrews:1989:VRR

[AJBL89] George E. Andrews, Lisa Jacobsen, Bruce C. Berndt, and Robert L. Lamphere. Variations on the Rogers–Ramanujan continued fraction in Ramanujan’s notebooks. Lecture Notes in Mathematics, 1395:73–83, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086400/.

Accardi:1989:MDM

[AJL89] L. Accardi, J.-L. Journé, and J. M. Lindsay. On multi-dimensional Markovian cocycles. Lecture Notes in Mathematics, 1396:59– 67, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083543/.

Aptekarev:1987:APT

[AK87] A. I. Aptekarev and V. A. Kalyagin. Analytic properties of twodimensional continued P -fraction expansions with periodical coefficients and their simultaneous Padé–Hermite approximants. Lecture Notes in Mathematics, 1237:145–160, 1987. CODEN LN-MAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072461/.

Alimpic:1988:SCR

[AK88] Branka P. Alimpić and Dragica N. Krgović. Some congruences on regular semigroups. Lecture Notes in Mathematics, 1320:1– 10, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083419/.

Akin:1989:BBS

[Aki89] Ömer Akin. Boundary behaviour of solutions to the Generalised Weinstein Equation. Lecture Notes in Mathematics, 1393:24– 38, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085769/.

Akkouchi:1988:CDN

[Akk88] Mohamed Akkouchi. Une caracterisation du noyau de Poisson d’un arbre homogène. (French) []. Lecture Notes in Mathematics, 1359:52–59, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086587/.

Arnold:1986:LEL

[AKO86] L. Arnold, W. Kliemann, and E. Oeljeklaus. Lyapunov exponents of linear stochastic systems. Lecture Notes in Mathematics, 1186:85–125, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076835/.

Anick:1986:HSF

[AL86] David Anick and Clas Löfwall. Hilbert series of finitely presented algebras. Lecture Notes in Mathematics, 1183:32–55, 1986. CO-DEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075448/.

Alber:1988:IBV

[AL88a] H.-D. Alber and R. Leis. Initial-boundary value and scattering problems in mathematical physics. Lecture Notes in Mathematics, 1357:23–60, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082861/.

Aubry:1988:HAL

[AL88b] Marc Aubry and Jean-Michel Lemaire. Homotopies d’algèbres de Lie et de leurs algèbres enveloppantes. (French) []. Lecture Notes in Mathematics, 1318:26–30, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077792/.

Aulaskari:1988:SIC

[AL88c] Rauno Aulaskari and Peter Lappan. Some integral conditions involving the spherical derivative. Lecture Notes in Mathematics, 1351:28–37, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081240/.

Alaoui:1988:SVK

[Ala88] Aziz El Kacimi Alaoui. Stabilité des V -variétés kähleriennes. (French) []. Lecture Notes in Mathematics, 1345:111–123, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081398/.

Alber:1985:LET

[Alb85] H.-D. Alber. A local existence theorem for the quasilinear wave equation with initial values of bounded variation. Lecture Notes in Mathematics, 1151:9–24, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074710/.

Alber:1988:ELS

[Alb88] H.-D. Alber. On the existence in the large of solutions to the one-dimensional, isentropic hydrodynamic equations in a bounded domain. Lecture Notes in Mathematics, 1357:1–22, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082860/.

Aldous:1985:ERT

[Ald85] David J. Aldous. Exchangeability and related topics. Lecture Notes in Mathematics, 1117:1–198, 1985. CODEN LN-MAA2. ISBN 3-540-15203-2 (print), 3-540-39316-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099421/.

Alexander:1985:NEU

[Ale85] Kenneth S. Alexander. The non-existence of a universal multiplier moment for the central limit theorem. Lecture Notes in Mathematics, 1153:15–16, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074943.

Alev:1986:AGF

[Ale86] J. Alev. Action de groupes sur A_{1}(C). (French) []. Lecture Notes in Mathematics, 1197:1–9, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076307/.

Alexander:1987:PHL

[Ale87] H. Alexander. Polynomial hulls and linear measure. Lecture Notes in Mathematics, 1276:1–11, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078949/.

Alfonsi:1986:CNA

[Alf86] B. Alfonsi. Conditions Noethériennes dans les anneaux gradués. (French) []. Lecture Notes in Mathematics, 1197:10–19, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076308/.

Alladi:1985:NTP

[All85a] Krishnaswami Alladi, editor. Number Theory: Proceedings of the 4th Matscience Conference held at Ootacamund, India, January 5–10, 1984, volume 1122 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1122. URL http:

//link.springer.com/book/10.1007/BFb0075746;   http://
www.springerlink.com/content/978-3-540-39642-0.

Alladi:1985:SMA

[All85b] Krishnaswami Alladi. A study of the moments of additive functions using Laplace transforms and sieve methods. Lecture Notes in Mathematics, 1122:1–37, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075747/.

Alladi:1989:DAF

[All89] Krishnaswami Alladi. The distribution of additive functions in special sets of integers. Lecture Notes in Mathematics, 1395:21– 63, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086398/.

Almorox:1987:STG

[Alm87] Antonio López Almorox. Supergauge theories in graded manifolds. Lecture Notes in Mathematics, 1251:114–136, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077318/.

Almeida:1988:PM

[Alm88] Jorge Almeida. On pseudovarieties of monoids. Lecture Notes in Mathematics, 1320:11–17, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083420/.

Alon:1989:PPE

[Alo89] Noga Alon. Probabilistic proofs of existence of rare events. Lecture Notes in Mathematics, 1376:186–201, 1989. CODEN LN-MAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090055/.

Alspach:1988:CTI

[Als88] Dale E. Alspach. Complemented translation-invariant subspaces. Lecture Notes in Mathematics, 1332:112–125, 1988. CODEN LN-MAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081614/.

Alt:1988:SOT

[Alt88] Hans Wilhelm Alt. On shape optimization of a turbine blade. Lecture Notes in Mathematics, 1357:61–95, 1988. CODEN LN-MAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082862/.

Alt:1989:FGL

[Alt89] Jean-Christian Alt. Une forme générale de la loi forte des grands nombres pour des variables aléatoires vectorielles. (French) []. Lecture Notes in Mathematics, 1391:1–15, 1989. CODEN LN-MAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083375/.

Aluffi:1988:CNS

[Alu88] Paolo Aluffi. The characteristic numbers of smooth plane cubics. Lecture Notes in Mathematics, 1311:1–8, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082905/.

Alvarez:1988:ITS

[AM88a] J. Alvarez and M. Milman. Interpolation of tent spaces and applications. Lecture Notes in Mathematics, 1302:11–21, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078860/.

Anderson:1988:ATB

[AM88b] Douglas R. Anderson and Hans J. Munkholm. The algebraic topology of boundedly controlled spaces. Lecture Notes in Mathematics, 1323:40–98, 1988. CODEN LNMAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079808/.

Anderson:1988:BCM

[AM88c] Douglas R. Anderson and Hans J. Munkholm. Boundedly controlled manifolds and the s-cobordism theorem. Lecture Notes in Mathematics, 1323:212–259, 1988. CODEN LNMAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079812/.

Anderson:1988:CTF

[AM88d] Douglas R. Anderson and Hans J. Munkholm. Category theoretic foundations. Lecture Notes in Mathematics, 1323:1–39, 1988. CODEN LNMAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079807/.

Anderson:1988:FPR

[AM88e] Douglas R. Anderson and Hans J. Munkholm. Free and projective rpg modules the algebraic Whitehead groups of rpg. Lecture Notes in Mathematics, 1323:123–175, 1988. CODEN LNMAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079810/.

Anderson:1988:GBC

[AM88f] Douglas R. Anderson and Hans J. Munkholm. The geometric, boundedly controlled Whitehead group. Lecture Notes in Mathematics, 1323:99–122, 1988. CODEN LNMAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079809/.

Anderson:1988:IBG

[AM88g] Douglas R. Anderson and Hans J. Munkholm. The isomorphism between the geometric and algebraic Whitehead groups. Lecture Notes in Mathematics, 1323:176–211, 1988. CODEN LN-MAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079811/.

Anderson:1988:TC

[AM88h] Douglas R. Anderson and Hans J. Munkholm. Toward computations. Lecture Notes in Mathematics, 1323:260–300, 1988. CODEN LNMAA2. ISBN 3-540-19397-9 (print), 3-540-39249-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079813/.

Amar:1987:EPN

[Ama87] Éric Amar. Estimations, par noyaux, des solutions de l’Équation ¯∂u = f. (French) []. Lecture Notes in Mathematics, 1276:12– 24, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078950/.

Amberg:1989:IFG

[Amb89] Bernhard Amberg. Infinite factorized groups. Lecture Notes in Mathematics, 1398:1–24, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086235/.

Abdalass:1986:MSS

[AMM86] E. M. Abdalass, J.-F. Maitre, and F. Musy. A multigrid solver for a stabilized finite element discretization of the Stokes problem. Lecture Notes in Mathematics, 1228:1–6, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072637/.

Araki:1985:OAT

[AMSV85] Huzihiro Araki, Calvin C. Moore, Şerban-Valentin Stratila, and Dan-Virgil Voiculescu, editors. Operator Algebras and their Connections with Topology and Ergodic Theory: Proceedings of the OATE Conference held in Buşteni, Romania, Aug. 29–Sept. 9, 1983, volume 1132 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1132. URL http:

//link.springer.com/book/10.1007/BFb0074873;   http://
www.springerlink.com/content/978-3-540-39514-0.

Ancel:1987:APM

[Anc87] Fredric D. Ancel. An alternative proof of M. Brown’s theorem on inverse sequences of near homeomorphisms. Lecture Notes in Mathematics, 1283:1–2, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081413.

Ancona:1988:PHF

[Anc88] Alano Ancona. Positive harmonic functions and hyperbolicity. Lecture Notes in Mathematics, 1344:1–23, 1988. CODEN LN-MAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103341/.

Andersson:1986:NAH

[And86a] Stig I. Andersson. Non-abelian Hodge theory via heat flow. Lecture Notes in Mathematics, 1209:8–36, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076618/.

Andrzejewski:1986:EWF

[And86b] Paweł Andrzejewski. The equivariant wall finiteness obstruction and Whitehead torsion. Lecture Notes in Mathematics, 1217:11– 25, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072812/.

Andler:1987:RDB

[And87a] Martin Andler. Relationships of divisibility between local Lfunctions associated to representations of complex reductive groups. Lecture Notes in Mathematics, 1243:1–14, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073015/.

Andre:1987:QDG

[And87b] Yves André. Quatre descriptions des groupes de Galois differentiels. (French) []. Lecture Notes in Mathematics, 1296:28–41, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078522/.

Anderson:1988:HFC

[And88] Michael T. Anderson. L^{2} harmonic forms on complete Riemannian manifolds. Lecture Notes in Mathematics, 1339:1–19, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083043/.

Andrews:1989:ISR

[And89] George E. Andrews. An identity of Sylvester and the Rogers– Ramanujan identities. Lecture Notes in Mathematics, 1395:64– 72, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086399/.

Anick:1986:RHA

[Ani86] David J. Anick. A rational homotopy analog of Whitehead’s problem. Lecture Notes in Mathematics, 1183:28–31, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075447/.

Anick:1988:RPH

[Ani88] David J. Anick. Recent progress in Hilbert and Poincaré series. Lecture Notes in Mathematics, 1318:1–25, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077791/.

Anick:1989:HES

[Ani89] David J. Anick. Homotopy exponents for spaces of category two. Lecture Notes in Mathematics, 1370:24–52, 1989. CODEN LN-MAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085216/.

Anker:1988:NCE

[Ank88] Jean-Philippe Anker. Le noyau de la chaleur sur les espaces symétriques U(p, q)/U(p) × U(q). (French) []. Lecture Notes in Mathematics, 1359:60–82, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086588/.

Antman:1985:BPA

[Ant85] Stuart S. Antman. Bifurcation problems associated with nonlinear wave propagation. Lecture Notes in Mathematics, 1151: 25–34, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074711/.

Accardi:1985:PAF

[AO85] Luigi Accardi and Stefano Olla. On the polaron asymptotics at finite coupling constant. Lecture Notes in Mathematics, 1136: 1–8, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074454/.

Alspach:1988:AWN

[AO88] D. Alspach and E. Odell. Averaging weakly null sequences. Lecture Notes in Mathematics, 1332:126–144, 1988. CODEN LN-MAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081615/.

Arnold:1986:ASM

[AOP86] L. Arnold, E. Oeljeklaus, and É. Pardoux. Almost sure and moment stability for linear Itô equations. Lecture Notes in Mathematics, 1186:129–159, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076837/.

Argyros:1988:CCS

[AOR88] S. Argyros, E. Odell, and H. Rosenthal. On certain convex subsets of c_{0}. Lecture Notes in Mathematics, 1332:80–111, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081613/.

Aoyama:1986:ERC

[Aoy86] Yōichi Aoyama. On endomorphism rings of canonical modules (joint work with Shiro Goto). Lecture Notes in Mathematics, 1183: 56–58, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075449/.

Accardi:1985:SCL

[AP85a] Luigi Accardi and K. R. Parthasarathy. Stochastic calculus on local algebras. Lecture Notes in Mathematics, 1136:9–23, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074455/.

Allday:1985:LTC

[AP85b] C. Allday and V. Puppe. On the localization theorem at the cochain level and free torus actions. Lecture Notes in Mathematics, 1172:1–16, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0074421/;   http://link.springer.com/content/
pdf/bfm:978-3-540-39745-8/1.pdf.

Allday:1986:BTR

[AP86] C. Allday and V. Puppe. Bounds on the torus rank. Lecture Notes in Mathematics, 1217:1–10, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072811/.

Alexander:1988:DFB

[AP88] J. C. Alexander and William Parry. Discerning fat Baker’s transformations. Lecture Notes in Mathematics, 1342:1–6, 1988. CO-DEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082820/.

Anile:1989:IRW

[AP89] A. M. Anile and S. Pennisi. An improved relativistic warm plasma model. Lecture Notes in Mathematics, 1385:238–256, 1989. CO-DEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084032/.

Applebaum:1985:QFS

[App85] David Applebaum. Quasi-free stochastic evolutions. Lecture Notes in Mathematics, 1136:46–56, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074458/.

Applebaum:1988:QSP

[App88] David Applebaum. Quantum stochastic parallel transport on non-commutative vector bundles. Lecture Notes in Mathematics, 1303:20–36, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078052/.

Arrondo:1989:RSR

[APS89] E. Arrondo, M. Pedreira, and I. Sols. On regular and stable ruled surfaces in P^{3}. Lecture Notes in Mathematics, 1389:1–15, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085919/.

Auslander:1987:ASS

[AR87] Maurice Auslander and Idun Reiten. Almost split sequences for Zgraded rings. Lecture Notes in Mathematics, 1273:232–243, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078847/.

Araki:1985:EPS

[Ara85] Huzihiro Araki. Ergodic properties of some C∗-dynamical systems. Lecture Notes in Mathematics, 1132:1–11, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074874/.

Arazy:1987:AID

[Ara87] Jonathan Arazy. An application of infinite dimensional holomorphy to the geometry of Banach spaces. Lecture Notes in Mathematics, 1267:122–150, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078141/.

Ara:1988:SRA

[Ara88] Pere Ara. Stable range of aleph-nought-continuous regular rings. Lecture Notes in Mathematics, 1328:1–7, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100909/.

Arbib:1989:PAP

[Arb89] Claudio Arbib. A polynomial algorithm for partitioning linegraphs. Lecture Notes in Mathematics, 1403:198–208, 1989. CO-DEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0083465/.

Arendt:1986:CPSa

[Are86a] Wolfgang Arendt. Characterization of positive semigroups on C_{o}(X). Lecture Notes in Mathematics, 1184:122–162, 1986. CO-DEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074928/.

Arendt:1986:CPSb

[Are86b] Wolfgang Arendt. Characterization of positive semigroups on Banach lattices. Lecture Notes in Mathematics, 1184:247–291, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074932/.

Arkhipov:1989:DFA

[Ark89] S. V. Arkhipov. The density function’s asymptotic representation in the case of multidimensional strictly stable distributions. Lecture Notes in Mathematics, 1412:1–21, 1989. CODEN LN-MAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084161/.

Arnold:1988:CSR

[Arn88a] V. I. Arnol’d. Contact structure, relaxation oscillations and singular points of implicit differential equations. Lecture Notes in Mathematics, 1334:173–179, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080428/.

Arnold:1988:SPS

[Arn88b] V. I. Arnol’d. On some problems in symplectic topology. Lecture Notes in Mathematics, 1346:1–5, 1988. CODEN LN-MAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082767/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-45958-3/
1.pdf.

Aronson:1986:PME

[Aro86] D. G. Aronson. The porous medium equation. Lecture Notes in Mathematics, 1224:1–46, 1986. CODEN LNMAA2. ISBN 3-540-17192-4 (print), 3-540-47352-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072687/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-47352-7/
1.pdf.

Arosio:1988:GSS

[Aro88a] Alberto Arosio. Global solvability of second order evolution equations in Banach scales. Lecture Notes in Mathematics, 1340:1– 10, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082880/.

Arous:1988:NCH

[Aro88b] Gérard Ben Arous. Noyau de la chaleur hypoelliptique et géométrie sous-riemannienne. (French) []. Lecture Notes in Mathematics, 1322:1–16, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077862/.

Arques:1986:RFD

[Arq86] Didier Arquès. Relations fonctionnelles et dénombrement des hypercartes planaires pointées. (French) []. Lecture Notes in Mathematics, 1234:5–26, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072505/.

Adamek:1988:ALC

[ART88] J. Adámek, J. Rosický, and V. Trnková. Are all limit-closed subcategories of locally presentable categories reflective? Lecture Notes in Mathematics, 1348:1–18, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081345/.

Arveson:1985:CNA

[Arv85] William Arveson. Continuous nests and the absorption principle. Lecture Notes in Mathematics, 1132:17–29, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074876/.

Ambos-Spies:1985:GRE

[AS85a] Klaus Ambos-Spies. Generators of the recursively enumerable degrees. Lecture Notes in Mathematics, 1141:1–28, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076212/.

Ancona:1985:NEB

[AS85b] Vincenzo Ancona and Alessandro Silva. Numerically effective bundles on Moišezon and strongly pseudoconvex manifolds. Lecture Notes in Mathematics, 1165:1–13, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076142/.

Antosik:1985:BMR

[AS85c] Piotr Antosik and Charles Swartz. Basic matrix results. Lecture Notes in Mathematics, 1113:7–10, 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072266/.

Antosik:1985:BMH

[AS85d] Piotr Antosik and Charles Swartz. Bilinear maps and hypocontinuity. Lecture Notes in Mathematics, 1113:45–57, 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072270/.

Antosik:1985:CO

[AS85e] Piotr Antosik and Charles Swartz. Convergence of operators. Lecture Notes in Mathematics, 1113:29–44, 1985. CODEN LN-MAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072269/.

Antosik:1985:Ib

[AS85f] Piotr Antosik and Charles Swartz. Imbedding c_{o} and l^{∞}. Lecture Notes in Mathematics, 1113:94–104, 1985. CODEN LN-MAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072274/.

Antosik:1985:Ia

[AS85g] Piotr Antosik and Charles Swartz. Introduction. Lecture Notes in Mathematics, 1113:1–6, 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072265/.

Antosik:1985:C

[AS85h] Piotr Antosik and Charles Swartz. K convergence. Lecture Notes in Mathematics, 1113:11–20, 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072267/.

Antosik:1985:MMA

[AS85i] Piotr Antosik and Charles Swartz. Matrix Methods in Analysis, volume 1113 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 114 pp. LCCN QA3 .L28 no. 1113. URL http:

//link.springer.com/book/10.1007/BFb0072264;   http://
www.springerlink.com/content/978-3-540-39266-8.

Antosik:1985:OPT

[AS85j] Piotr Antosik and Charles Swartz. Orlicz–Pettis Theorems. Lecture Notes in Mathematics, 1113:58–73, 1985. CODEN LN-MAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072271/.

Antosik:1985:SPL

[AS85k] Piotr Antosik and Charles Swartz. The Schur and Phillips lemma. Lecture Notes in Mathematics, 1113:74–84, 1985. CODEN LN-MAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072272/.

Antosik:1985:SLB

[AS85l] Piotr Antosik and Charles Swartz. The Schur lemma for bounded multiplier convergent series. Lecture Notes in Mathematics, 1113: 85–93, 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072273/.

Antosik:1985:UBP

[AS85m] Piotr Antosik and Charles Swartz. The uniform boundedness principle. Lecture Notes in Mathematics, 1113:21–28, 1985. CODEN LNMAA2. ISBN 3-540-15185-0 (print), 3-540-39266-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072268/.

Aseev:1988:QEP

[AS88] Vladislav V. Aseev and Anatoliĭ V. Syčev. Quasiconformal extension of plane quasi-Möbius embeddings. Lecture Notes in Mathematics, 1351:23–27, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081239/.

Accardi:1989:QST

[AS89a] L. Accardi and Kalyan Sinha. Quantum stop times. Lecture Notes in Mathematics, 1396:68–72, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083544/.

Assem:1989:API

[AS89b] Ibrahim Assem and Andrzej Skowroński. Algèbres pre-inclinées et catégories dérivées. (French) []. Lecture Notes in Mathematics, 1404:1–34, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084070/.

Asada:1985:CFS

[Asa85] Akira Asada. Curvature forms with singularities and non-integral characteristic classes. Lecture Notes in Mathematics, 1139:152– 168, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074582/.

Asada:1986:NAP

[Asa86] Akira Asada. Non abelian Poincarè lemma. Lecture Notes in Mathematics, 1209:37–65, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076619/.

Asada:1989:IFI

[Asa89] Akira Asada. Integrable forms on iterated loop spaces and higher dimensional non Abelian de Rham theory. (French) []. Lecture Notes in Mathematics, 1410:27–51, 1989. CODEN LN-MAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086409/.

Aschbacher:1986:FSG

[Asc86] Michael Aschbacher. Finite simple groups and their subgroups. Lecture Notes in Mathematics, 1185:1–57, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076170/.

Arad:1985:PPC

[ASH85] Z. Arad, J. Stavi, and M. Herzog. Powers and products of conjugacy classes in groups. Lecture Notes in Mathematics, 1112: 6–51, 1985. CODEN LNMAA2. ISBN 3-540-13916-8 (print), 3-540-39142-8 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0072285/;   http://link.springer.com/chapter/
10.1007/BFb0072286/.

Askey:1989:BIA

[Ask89] Richard Askey. Beta integrals and the associated orthogonal polynomials. Lecture Notes in Mathematics, 1395:84–121, 1989. CO-DEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086401/.

Asoh:1989:SAS

[Aso89] Tohl Asoh. Smooth SL(2, C) actions on the 3-sphere. Lecture Notes in Mathematics, 1375:6–15, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085595/.

Assadi:1985:LPT

[Ass85] Amir Assadi. Lower Postnikov terms of generalized CW complexes and semi-simple actions. Lecture Notes in Mathematics, 1167:1–23, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075213/.

Assadi:1986:HAC

[Ass86a] Amir H. Assadi. Homotopy actions and cohomology of finite groups. Lecture Notes in Mathematics, 1217:26–57, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072813/.

Assadi:1986:NLP

[Ass86b] Amir H. Assadi. Normally linear Poincaré complexes and equivariant splittings. Lecture Notes in Mathematics, 1217:58–78, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072814/.

Assadi:1989:FDS

[Ass89] Amir H. Assadi. “On finite domination and simple homotopy type of nonsimply-connected G-spaces”. Lecture Notes in Mathematics, 1375:16–32, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085596/.

Al-Salam:1988:GPS

[ASV88] W. A. Al-Salam and A. Verma. On the Geronimus polynomial sets. Lecture Notes in Mathematics, 1329:193–202, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083359/.

Acquistapace:1986:FSA

[AT86] Paolo Acquistapace and Brunello Terreni. On fundamental solutions for abstract parabolic equations. Lecture Notes in Mathematics, 1223:1–11, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099178/.

Atiyah:1985:CAM

[Ati85a] Michael Atiyah. Commentary on the article of Manin. Lecture Notes in Mathematics, 1111:103–109, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084586/. See [Man85].

Atiyah:1985:EDO

[Ati85b] Michael Atiyah. Eigenvalues of the Dirac operator. Lecture Notes in Mathematics, 1111:251–260, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084593/.

Attouch:1986:ECD

[Att86] H. Attouch. Epi-convergence and duality. Convergence of sequences of marginal and Lagrangians functions. Applications to homogenization problems in mechanics. Lecture Notes in Mathematics, 1190:21–56, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076701/.

Atzmon:1988:IRA

[Atz88] Aharon Atzmon. Irreducible representations of Abelian groups. Lecture Notes in Mathematics, 1359:83–92, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086589/.

Auslander:1986:ISE

[AU86] Maurice Auslander and Louise Unger. Isolated singularities and existence of almost split sequences. Lecture Notes in Mathematics, 1178:194–242, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075297/.

Aubin:1986:PYC

[Aub86] Thierry Aubin. Le problème de Yamabe concernant la courbure scalaire. (French) []. Lecture Notes in Mathematics, 1209: 66–72, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076620/.

Aubin:1989:MEK

[Aub89] Thierry Aubin. Métriques d’Einstein–Kähler. (French) []. Lecture Notes in Mathematics, 1410:52–59, 1989. CODEN LN-MAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086410/.

Audin:1987:CCL

[Aud87] Michèle Audin. Classes caractéristiques lagrangiennes. (French) []. Lecture Notes in Mathematics, 1298:1–16, 1987. CODEN LN-MAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082997/.

Aulbach:1985:CLS

[Aul85] Bernd Aulbach. Conditions for a limit set to be a singleton. Lecture Notes in Mathematics, 1151:35–41, 1985. CODEN LN-MAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074712/.

Aure:1989:SQT

[Aur89] Alf Bjørn Aure. Surfaces on quintic threefolds associated to the Horrocks–Mumford bundle. Lecture Notes in Mathematics, 1399: 1–9, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095964/.

Auslander:1985:FTI

[Aus85] M. Auslander. Finite type implies isolated singularity. Lecture Notes in Mathematics, 1142:1–4, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074789/.

Assadi:1985:SFG

[AV85] A. H. Assadi and P. Vogel. Semifree finite groups actions on compact manifolds. Lecture Notes in Mathematics, 1126:1–21, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074436/.

Alexander:1988:FIS

[AV88a] J. C. Alexander and Alberto Verjovsky. First integrals for singular holomorphic foliations with leaves of bounded volume. Lecture Notes in Mathematics, 1345:1–10, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081394/.

Awami:1988:FRN

[AV88b] Mohamed A. Awami and F. Van Oystaeyen. On filtered rings with Noetherian associated graded rings. Lecture Notes in Mathematics, 1328:8–27, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100910/.

Avramov:1986:GH

[Avr86] Luchezar L. Avramov. Golod homomorphisms. Lecture Notes in Mathematics, 1183:59–78, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075450/.

Avrin:1987:CPS

[Avr87] Joel D. Avrin. Convergence properties of strongly-damped semilinear wave equations. Lecture Notes in Mathematics, 1248:1– 4, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077410/.

Anderson:1988:IED

[AVV88] G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen. Inequalities for the extremal distortion function. Lecture Notes in Mathematics, 1351:1–11, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081237/.

Accardi:1985:QPA

[AvW85] Luigi Accardi and Wilhelm von Waldenfels, editors. Quantum Probability and Applications II: Proceedings of a Workshop held in Heidelberg, West Germany, October 1–5, 1984, volume 1136 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1136. URL http://link.springer.com/book/10.1007/
BFb0074453;   http://www.springerlink.com/content/978-
3-540-39570-6.

Arnold:1986:LES

[AW86a] L. Arnold and V. Wihstutz. Lyapunov exponents: a survey. Lecture Notes in Mathematics, 1186:1–26, 1986. CODEN LN-MAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076830/.

Arnold:1986:LEP

[AW86b] Ludwig Arnold and Volker Wihstutz, editors. Lyapunov Exponents: Proceedings of a Workshop held in Bremen, November 12–15, 1984, volume 1186 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1186. URL http:

//link.springer.com/book/10.1007/BFb0076829;   http://
www.springerlink.com/content/978-3-540-39795-3.

Accardi:1989:QRW

[AW89] L. Accardi and G. S. Watson. Quantum random walks. Lecture Notes in Mathematics, 1396:73–88, 1989. CODEN LN-MAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083545/.

Axelsson:1986:SEE

[Axe86] O. Axelsson. Stability and error estimates valid for infinite time, for strongly monotone and infinitely stiff evolution equations. Lecture Notes in Mathematics, 1192:275–284, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076081/.

Azema:1985:SPX

[AY85] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XIX 1983/84: Proceedings, volume 1123 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1123. URL http:

//link.springer.com/book/10.1007/BFb0075834;   http://
www.springerlink.com/content/978-3-540-39397-9.

Azema:1986:SPX

[AY86] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XX 1984/85: Proceedings, volume 1204 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1204. URL http:

//link.springer.com/book/10.1007/BFb0075705;   http://
www.springerlink.com/content/978-3-540-39860-8.

Azema:1987:ICH

[AY87] J. Azéma and M. Yor. Interprétation d’un calcul de H. Tanaka en théorie générale des processus. (French) []. Lecture Notes in Mathematics, 1247:262–269, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077640/.

Azema:1989:EMR

[AY89] J. Azéma and M. Yor. Étude d’une martingale remarquable. (French) []. Lecture Notes in Mathematics, 1372:88–130, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083962/.

Azema:1987:SPX

[AYM87] Jacques Azéma, Marc Yor, and Paul André Meyer. Séminaire de Probabilités XXI, volume 1247 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 302 (est.) pp. LCCN QA3

.L28 no. 1247. URL http://link.springer.com/book/10.1007/
BFb0077623;   http://www.springerlink.com/content/978-
3-540-47814-0.

Aznar:1987:SHM

[Azn87] V. Navarro Aznar. Sur les structures de Hodge mixtes associées aux cycles évanescents. (French) []. Lecture Notes in Mathematics, 1246:143–153, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077536/.

Aznar:1988:TAF

[Azn88] V. Navarro Aznar. Théorèmes d’annulation. (French) [Cancellation theorems]. Lecture Notes in Mathematics, 1335:133– 160, 1988. CODEN LNMAA2. ISBN 3-540-50023-5 (print), 3-540-69984-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085059/.

Bhargava:1985:TGR

[BA85] S. Bhargava and Chandrashekar Adiga. Two generalizations of Ramanujan’s continued fraction identities. Lecture Notes in Mathematics, 1122:56–62, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075750/.

Ben-Artzi:1989:NCR

[BA89] Matania Ben-Artzi. Numerical calculations of reacting flows. Lecture Notes in Mathematics, 1402:30–36, 1989. CODEN LN-MAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083866/.

Backelin:1986:RGH

[Bac86] Jörgen Backelin. On the rates of growth of the homologies of Veronese subrings. Lecture Notes in Mathematics, 1183:79– 100, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075451/.

Bachelot:1989:GEL

[Bac89] Alain Bachelot. Global existence of large amplitude solutions for Dirac–Klein–Gordon systems in Minkowski space. Lecture Notes in Mathematics, 1402:99–113, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083870/.

Badrikian:1985:RGF

[Bad85] Albert Badrikian. Réarrangement Gaussien de fonctions. (French) []. Lecture Notes in Mathematics, 1109:18–31, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101534/.

Baer:1986:HPW

[Bae86] Dagmar Baer. Homological properties of wild hereditary Artin algebras. Lecture Notes in Mathematics, 1177:1–12, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075255/.

Baernstein:1987:DST

[Bae87] Albert Baernstein II. Dubinin’s symmetrization theorem. Lecture Notes in Mathematics, 1275:23–30, 1987. CODEN LN-MAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078342/.

Baernstein:1989:STS

[Bae89] Albert Baernstein II. Some topics in symmetrization. Lecture Notes in Mathematics, 1384:111–123, 1989. CODEN LN-MAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086796/.

Bahri:1988:CPI

[Bah88] A. Bahri. Critical points at infinity in the variational calculus. Lecture Notes in Mathematics, 1324:1–29, 1988. CODEN LN-MAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100779/.

Baiocchi:1989:SLA

[Bai89] C. Baiocchi. Stability in linear abstract differential equations. Lecture Notes in Mathematics, 1386:1–21, 1989. CODEN LN-MAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089228/.

Bakry:1985:TRPa

[Bak85a] D. Bakry. Transformations de Riesz pour les semi-groupes symétriques. Première partie: Étude de la dimension 1. (French) []. Lecture Notes in Mathematics, 1123:130–144, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075843/.

Bakry:1985:TRPb

[Bak85b] D. Bakry. Transformations de Riesz pour les semi-groupes symétriques. seconde patrie: Étude sous la condition Γ_{2} ≥ 0. (French) []. Lecture Notes in Mathematics, 1123:145–174, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075844/.

Bakry:1985:RIL

[Bak85c] Dominique Bakry. Une remarque sur les inégalités de Littlewood– Paley sous l’hypothèse Γ_{2} ≥ 0. (French) []. Lecture Notes in Mathematics, 1123:175, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075845.

Baker:1987:CAI

[Bak87a] Andrew Baker. Combinatorial and arithmetic identities based on formal group laws. Lecture Notes in Mathematics, 1298:17– 34, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082998/.

Bakry:1987:ETR

[Bak87b] Dominique Bakry. Étude des transformations de Riesz dans les variétés riemanniennes à courbure de Ricci minorée. (French) []. Lecture Notes in Mathematics, 1247:137–172, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077631/.

Bakry:1988:PSH

[Bak88] Dominique Bakry. La propriété de sous-harmonicité des diffusions dans les variétés. (French) []. Lecture Notes in Mathematics, 1321:1–50, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084117/.

Bakry:1989:ICS

[Bak89] Dominique Bakry. Sur l’interpolation complexe des semigroupes de diffusion. (French) []. Lecture Notes in Mathematics, 1372: 1–20, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083956/.

Ballmann:1985:MNP

[Bal85] W. Ballmann. Manifolds of non positive curvature. Lecture Notes in Mathematics, 1111:261–268, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084594/.

Ballmann:1986:SMN

[Bal86a] Werner Ballmann. Structure of manifolds of nonpositive sectional curvature. Lecture Notes in Mathematics, 1201:1–13, 1986. CO-DEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075644/.

Balslev:1986:SOA

[Bal86b] Erik Balslev, editor. Schrödinger Operators, Aarhus 1985: Lectures given in Aarhus, October 2–4, 1985, volume 1218 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1218. URL http://link.springer.com/book/10.1007/
BFb0073041;   http://www.springerlink.com/content/978-
3-540-47119-6.

Balslev:1986:WOD

[Bal86c] Erik Balslev. Wave operators for dilation-analytic three-body Hamiltonians. Lecture Notes in Mathematics, 1218:39–60, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073043/.

Baldwin:1987:CT

[Bal87] John T. Baldwin. Classification theory: 1985. Lecture Notes in Mathematics, 1292:1–23, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082229/.

Ball:1988:SRG

[Bal88] Keith Ball. Some remarks on the geometry of convex sets. Lecture Notes in Mathematics, 1317:224–231, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081743/.

Ball:1989:VSC

[Bal89] Keith Ball. Volumes of sections of cubes and related problems. Lecture Notes in Mathematics, 1376:251–260, 1989. CODEN LN-MAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090058/.

Bank:1986:PEE

[Ban86] Randolph E. Bank. A-posteriori error estimates. Adaptive local mesh refinement and multigrid iteration. Lecture Notes in Mathematics, 1228:7–22, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072638/.

Barbu:1985:TOC

[Bar85a] Viorel Barbu. The time optimal control of variational inequalities. Dynamic programming and the maximum principle. Lecture Notes in Mathematics, 1119:1–19, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074777/.

Bartels:1985:QAV

[Bar85b] Dina Bartels. Quasihomogene affine Varietäten für SL(2, C). (German) []. Lecture Notes in Mathematics, 1146:1–105, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074535/.

Barth:1985:LES

[Bar85c] W. Barth. Lectures on K3- and Enriques surfaces. Lecture Notes in Mathematics, 1124:21–57, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074994/.

Barugola:1985:SPA

[Bar85d] A. Barugola. On some properties of an absorptive area and a chaotic area for an R^{2}-endomorphism. Lecture Notes in Mathematics, 1163:1–6, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076411/.

Barcilon:1986:IEP

[Bar86a] Victor Barcilon. Inverse eigenvalue problems. Lecture Notes in Mathematics, 1225:1–51, 1986. CODEN LNMAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic).  URL http://link.springer.com/
chapter/10.1007/BFb0072659/;   http://link.springer.com/
content/pdf/bfm:978-3-540-47353-4/1.pdf.

Barlet:1986:MP

[Bar86b]  D. Barlet.  Monodromy and poles of ∫_{X} |f|^{2λ}□.  Lec-

ture Notes in Mathematics, 1194:1–7, 1986. CODEN LN-MAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076990/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-39829-5/
1.pdf.

Barlet:1986:FTT

[Bar86c] Daniel Barlet. Fonctions de type trace réelles et formes hermitiennes horizontales. (French) []. Lecture Notes in Mathematics, 1188:232–244, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076825/.

Barlet:1986:CFH

[Bar86d] Daniel Barlet. Le calcul de la forme hermitienne canonique pour X^{a} + Y^{b} + Z^{c} = 0. (French) []. Lecture Notes in Mathematics, 1198:35–46, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077041/.

Bartusek:1986:POS

[Bar86e] M. Bartušek. On properties of oscillatory solutions of nonlinear differential equations of the n-th order. Lecture Notes in Mathematics, 1192:109–113, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076056/.

Barbasch:1987:URU

[Bar87a] Dan Barbasch. Unipotent representations and unitarity. Lecture Notes in Mathematics, 1243:73–85, 1987. CODEN LN-MAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073018/.

Barlet:1987:SHP

[Bar87b] D. Barlet. Symétrie de Hodge pour le polynôme de Bernstein– Sato. (French) []. Lecture Notes in Mathematics, 1295:1–10, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081974/.

Barlotti:1987:FSM

[Bar87c] Marco Barlotti. Faithful simple modules for the non-Abelian group of order pq. Lecture Notes in Mathematics, 1281:1–8, 1987. CO-DEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078684/.

Baron:1987:LPZ

[Bar87d] G. Baron. Ein Lineares Programm in der Zahlentheorie. (German) [A linear program in number theory]. Lecture Notes in Mathematics, 1262:1–13, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078591/.

Barrett:1987:BSB

[Bar87e] David E. Barrett. Boundary singularities of biholomorphic maps. Lecture Notes in Mathematics, 1268:24–28, 1987. CODEN LN-MAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097293/.

Barth:1987:PFT

[Bar87f] Thomas Barth. Sur les principes fondamentaux de la théorie du potentiel par rapport à un noyau. (French) []. Lecture Notes in Mathematics, 1235:1–14, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072745/.

Barchielli:1988:IOC

[Bar88a] Alberto Barchielli. Input and output channels in quantum systems and quantum stochastic differential equations. Lecture Notes in Mathematics, 1303:37–51, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078053/.

Barr:1988:CEU

[Bar88b] Michael Barr. On categories with effective unions. Lecture Notes in Mathematics, 1348:19–35, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081346/.

Barza:1988:DMK

[Bâr88c] Ilie Bârză. The dianalytic morphisms of the Klein bottles. Lecture Notes in Mathematics, 1351:38–51, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081241/.

Bardelli:1989:CGT

[Bar89a] Fabio Bardelli. Curves of genus three on a general Abelian threefold and the non-finite generation of the Griffiths group. Lecture Notes in Mathematics, 1399:10–26, 1989. CODEN LN-MAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0095965/.

Barros:1989:SFT

[Bar89b] Manuel Barros. Spherical finite type submanifolds. Applications. Lecture Notes in Mathematics, 1410:60–70, 1989. CODEN LN-MAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086411/.

Bartoszewicz:1989:SSC

[Bar89c] Artur Bartoszewicz. On some sufficient conditions for the existence of a consistent test in noncommutative statistics. Lecture Notes in Mathematics, 1391:16–22, 1989. CODEN LN-MAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083376/.

Bartsch:1989:BTM

[Bar89d] Thomas Bartsch. Bifurcation theory for metric parameter spaces. Lecture Notes in Mathematics, 1411:1–8, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086435/.

Bass:1987:IFB

[Bas87] Richard Bass. L_{p} inequalities for functionals of Brownian motion. Lecture Notes in Mathematics, 1247:206–217, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077635/.

Batchelor:1987:GMP

[Bat87] Marjorie Batchelor. Graded manifolds and pairs. Lecture Notes in Mathematics, 1251:65–72, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077315/.

Baues:1985:HCP

[Bau85] Hans Joachim Baues. On homotopy classification problems of J. H. C. Whitehead. Lecture Notes in Mathematics, 1172:17–55, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074422/.

Bauer:1987:SHT

[Bau87] Friedrich W. Bauer. Strong homology theories. Lecture Notes in Mathematics, 1283:3–29, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081414/.

Bauer:1988:HTM

[Bau88a] Stefan Bauer. The homotopy type of a 4-manifold with finite fundamental group. Lecture Notes in Mathematics, 1361:1–6, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083030/.

Baues:1988:CH

[Bau88b] Hans Joachim Baues. Combinatorial homotopy. Lecture Notes in Mathematics, 1318:31–50, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077793/.

Baumgartner:1989:RPO

[Bau89] James E. Baumgartner. Remarks on partition ordinals. Lecture Notes in Mathematics, 1401:5–17, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097328/.

Baxendale:1986:ABS

[Bax86] Peter H. Baxendale. Asymptotic behaviour of stochastic flows of diffeomorphisms. Lecture Notes in Mathematics, 1203:1–19, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076869/.

Bialynicki-Birula:1985:EUR

[BB85] Iwo Białynicki-Birula. Entropic uncertainty relations in quantum mechanics. Lecture Notes in Mathematics, 1136:90–103, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074463/.

Bedford:1986:BCP

[BB86] Eric Bedford and Steve Bell. Boundary continuity of proper holomorphic correspondences. Lecture Notes in Mathematics, 1198: 47–64, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077042/.

Baldwin:1987:CRL

[BB87] John T. Baldwin and Joel Berman. Concrete representations of lattices and the fundamental order. Lecture Notes in Mathematics, 1292:24–31, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082230/.

Berard:1988:NBS

[BB88a] P. H. Bérard and G. Besson. On the number of bound states and estimates on some geometric invariants. Lecture Notes in Mathematics, 1324:30–40, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100780/.

Bloom:1988:MSF

[BB88b] Lynette M. Bloom and Walter R. Bloom. Multipliers on spaces of functions with p-summable Fourier transforms. Lecture Notes in Mathematics, 1359:100–112, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086591/.

Boboc:1988:OCP

[BB88c] N. Boboc and Gh. Bucur. Order and convexity in potential theory. Lecture Notes in Mathematics, 1344:24–41, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103342/.

Buser:1988:RSL

[BBD88] Peter Buser, Marc Burger, and Jozef Dodziuk. Riemann surfaces of large genus and large λ_{1}. Lecture Notes in Mathematics, 1339:54–63, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083046/.

Bujalance:1989:GAN

[BBGM89] Emilio Bujalance, J. A. Bujalance, G. Gromadzki, and E. Martinez. The groups of automorphisms of non-orientable hyperelliptic Klein surfaces without boundary. Lecture Notes in Mathematics, 1398:43–51, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086238/.

Bernardi:1988:CHO

[BBP88] E. Bernardi, A. Bove, and C. Parenti. On a class of hyperbolic operators with double characteristics. Lecture Notes in Mathematics, 1340:23–30, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082882/.

Braaksma:1985:DSB

[BBT85] Boele L. J. Braaksma, Hendrik W. Broer, and Floris Takens, editors. Dynamical Systems and Bifurcations: Proceedings of a Workshop held in Groningen The Netherlands, April 16–20, 1984, volume 1125 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1125. URL http:

//link.springer.com/book/10.1007/BFb0075630;   http://
www.springerlink.com/content/978-3-540-39411-2.

Beyer:1985:FEE

[BC85] W. A. Beyer and P. J. Channell. A functional equation for the embedding of a homeomorphism of the interval into a flow. Lecture Notes in Mathematics, 1163:7–13, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076412/.

Barbosa:1986:CMS

[BC86a] J. Lucas M. Barbosa and A. Gervasio Colares. Complete minimal surfaces with finite total curvature. Lecture Notes in Mathematics, 1195:28–50, 1986. CODEN LNMAA2. ISBN 3-540-16491-X (print), 3-540-39830-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077108/.

Barbosa:1986:MS

[BC86b] J. Lucas M. Barbosa and A. Gervasio Colares. Minimal Surfaces in R^{3}, volume 1195 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16491-X (print), 3-540-39830-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 124 pp. LCCN QA3 .L28 no. 1195. URL http:

//link.springer.com/book/10.1007/BFb0077105;   http://
www.springerlink.com/content/978-3-540-39830-1.

Barbosa:1986:NCM

[BC86c] J. Lucas M. Barbosa and A. Gervasio Colares. Nonexistence of certain minimal surfaces. Lecture Notes in Mathematics, 1195: 98–113, 1986. CODEN LNMAA2. ISBN 3-540-16491-X (print), 3-540-39830-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077110/.

Barbosa:1986:PD

[BC86d] J. Lucas M. Barbosa and A. Gervasio Colares. A preliminary discussion. Lecture Notes in Mathematics, 1195:1–9, 1986. CODEN LNMAA2. ISBN 3-540-16491-X (print), 3-540-39830-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077106/.

Barbosa:1986:REC

[BC86e] J. Lucas M. Barbosa and A. Gervasio Colares. Recent examples of complete minimal surfaces. Lecture Notes in Mathematics, 1195:51–97, 1986. CODEN LNMAA2. ISBN 3-540-16491-X (print), 3-540-39830-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077109/.

Barbosa:1986:WRC

[BC86f] J. Lucas M. Barbosa and A. Gervasio Colares. The Weierstrass representation and the classical examples. Lecture Notes in Mathematics, 1195:10–27, 1986. CODEN LNMAA2. ISBN 3-540-16491-X (print), 3-540-39830-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077107/.

Bodigheimer:1988:RCC

[BC88] C.-F. Bödigheimer and F. R. Cohen. Rational cohomology of configuration spaces of surfaces. Lecture Notes in Mathematics, 1361:7–13, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083031/.

Brezzi:1986:SPP

[BCM86] Franco Brezzi, António Capelo, and Luisa Donatella Marini. Singular perturbation problems in semiconductor devices. Lecture Notes in Mathematics, 1230:191–198, 1986. CODEN LN-MAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072681/.

Baldi:1988:EVF

[BCM88] P. Baldi and M. Chaleyat-Maurel. An extension of Ventsel– Freidlin estimates. Lecture Notes in Mathematics, 1316:305– 327, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081937/.

Blanchard:1985:EQD

[BCSSC85] Ph. Blanchard, Ph. Combe, M. Sirugue, and M. Sirugue-Collin. Estimates of quantum deviations from classical mechanics using large deviation results. Lecture Notes in Mathematics, 1136:104– 111, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074464/.

Blanchard:1987:JPR

[BCSSC87] Ph. Blanchard, Ph. Combe, M. Sirugue, and M. Sirugue-Collin. Jump processes related to the two dimensional Dirac equation. Lecture Notes in Mathematics, 1250:1–13, 1987. CODEN LN-MAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077344/.

Bonneau:1986:DEL

[BCZ86] Pierre Bonneau, Anne Cumenge, and Ahmed Zériahi. Division dans les espaces de Lipschitz de fonctions holomorphes. (French) []. Lecture Notes in Mathematics, 1198:73–87, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077044/.

Berger:1987:TOF

[BCZ87] C. A. Berger, L. A. Coburn, and K. H. Zhu. Toeplitz operators and function theory in n-dimensions. Lecture Notes in Mathematics, 1256:28–35, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077736/.

Belopolskaya:1986:SED

[BD86a] Ya. I. Belopol’skaya and Yu. L. Daletskiĭ. Stochastic equations and differential geometry. Lecture Notes in Mathematics, 1214: 131–158, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075963/.

Burton:1986:UCD

[BD86b] T. A. Burton and D. P. Dwiggins. Uniqueness without continuous dependence. Lecture Notes in Mathematics, 1192:115–121, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076057/.

Branner:1988:SCP

[BD88] Bodil Branner and Adrien Douady. Surgery on complex polynomials. Lecture Notes in Mathematics, 1345:11–72, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081395/.

Beauville:1989:FTD

[BD89a] Arnaud Beauville and Olivier Debarre. Sur les fonctions thêta du second ordre. (French) []. Lecture Notes in Mathematics, 1399: 27–39, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095966/.

Bendersky:1989:CBC

[BD89b] Martin Bendersky and Donald M. Davis. On the complex bordism of classifying spaces. Lecture Notes in Mathematics, 1370: 53–56, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085217/.

Bisschop:1989:SDS

[BDD89] J. Bisschop, B. Dorhout, and A. Drud. Structural dependence and systems of equations. Lecture Notes in Mathematics, 1403: 209–224, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083466/.

Bratteli:1985:LDO

[BDE85] Ola Bratteli, Trond Digernes, and George A. Elliott. Locality and differential operators on C∗-algebras, II. Lecture Notes in Mathematics, 1132:46–83, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074879/.

Buendia:1988:DZP

[BDG88] E. Buendía, J. S. Dehesa, and F. J. Gálvez. The distribution of zeros of the polynomial eigenfunctions of ordinary differential operators of arbitrary order. Lecture Notes in Mathematics, 1329:222–235, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083361/.

Beck:1985:PBS

[BDH^{+}85] Anatole Beck, Richard Dudley, Marjorie Hahn, James Kuelbs, and Michael Marcus, editors. Probability in Banach Spaces V: Proceedings of the International Conference held in Medford, USA, July 16–27, 1984, volume 1153 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1153; QA3 .L28 no. 1153

QA273.43. URL http://link.springer.com/book/10.1007/
BFb0074941;   http://www.springerlink.com/content/978-
3-540-39645-1.

Bland:1987:CAG

[BDK87] J. Bland, T. Duchamp, and M. Kalka. A characterization of CP^{n} by its automorphism group. Lecture Notes in Mathematics, 1268:60–65, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097296/.

Brezinski:1985:POA

[BDM^{+}85] Claude Brezinski, André Draux, Alphonse P. Magnus, Pascal Maroni, and André Ronveaux, editors. Polynômes Orthogonaux et Applications: Proceedings of the Laguerre Symposium held at Bar-le-Duc, October 15–18, 1984, volume 1171 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1171. URL http://link.springer.com/book/10.1007/
BFb0076527;   http://www.springerlink.com/content/978-
3-540-39743-4.

Beals:1987:ISS

[BDT87a] R. Beals, Percy Deift, and Carlos Tomei. Inverse scattering for selfadjoint n^{th} order differential operators on the line. Lecture Notes in Mathematics, 1285:26–38, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080578/.

Blumlinger:1987:MSG

[BDT87b] M. Blümlinger, M. Drmota, and R. F. Tichy. Metrische Sätze der C-Gleichverteilung auf der Sphäre. (German) []. Lecture Notes in Mathematics, 1262:14–21, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078592/.

Bakry:1985:DH

[BÉ85] D. Bakry and Michel Émery. Diffusions hypercontractives. Lecture Notes in Mathematics, 1123:177–206, 1985. CODEN LN-MAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075847/.

Barbour:1986:TST

[BE86] A. D. Barbour and G. K. Eagleson. Tests for space–time clustering. Lecture Notes in Mathematics, 1212:42–51, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076236/.

Ballico:1987:BMR

[BE87] Edoardo Ballico and Philippe Ellia. Beyond the maximal rank conjecture for curves in IP^{3}. Lecture Notes in Mathematics, 1266: 1–23, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078176/.

Beattie:1986:SSB

[Bea86a] Margaret Beattie. The subgroup structure of the Brauer group of RG-dimodule algebras. Lecture Notes in Mathematics, 1197: 20–30, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076309/.

Beauville:1986:GMF

[Bea86b] Arnaud Beauville. Le groupe de monodromie des familles universelles d’hypersurfaces et d’intersections complètes. (French) [].

Lecture Notes in Mathematics, 1194:8–18, 1986. CODEN LN-MAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076991/.

Beals:1987:IRS

[Bea87] Michael Beals. Interaction of radially smooth nonlinear waves. Lecture Notes in Mathematics, 1256:1–27, 1987. CODEN LN-MAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077735/.

Beals:1988:NPC

[Bea88a] Richard Beals. Nonelliptic problems and complex analysis. Lecture Notes in Mathematics, 1306:23–37, 1988. CODEN LN-MAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082922/.

Beattie:1988:DTG

[Bea88b] Margaret Beattie. Duality theorems for group actions and gradings. Lecture Notes in Mathematics, 1328:28–32, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100911/.

Bebernes:1986:DBS

[Beb86] J. W. Bebernes. A description of blow-up for the solid fuel ignition model. Lecture Notes in Mathematics, 1192:193–196, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076068/.

Becker:1988:MCP

[Bec88] Howard S. Becker. More closure properties of pointclasses. Lecture Notes in Mathematics, 1333:31–36, 1988. CODEN LN-MAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084968/.

Buchweitz:1987:CMM

[BEH87] Ragnar-Olaf Buchweitz, David Eisenbud, and Jürgen Herzog. Cohen–Macaulay modules on quadrics. Lecture Notes in Mathematics, 1273:58–116, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078838/.

Beilinson:1987:HPB

[Beĭ87a] A. A. Beĭlinson. Height pairing between algebraic cycles. Lecture Notes in Mathematics, 1289:1–26, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078364/.

Beilinson:1987:HGP

[Beĭ87b] A. A. Beĭlinson. How to glue perverse sheaves. Lecture Notes in Mathematics, 1289:42–51, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078366/.

Beilinson:1987:DCP

[Beĭ87c] A. A. Beĭlinson. On the derived category of perverse sheaves. Lecture Notes in Mathematics, 1289:27–41, 1987. CODEN LN-MAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078365/.

daVeiga:1988:ILC

[Bei88] Hugo Beirão da Veiga. On the incompressible limit of the compressible Navier–Stokes equations. Lecture Notes in Mathematics, 1340:11–22, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082881/.

Bejancu:1985:SCS

[Bej85] Aurel Bejancu. A survey on CR-submanifolds of Kählerian manifolds. Lecture Notes in Mathematics, 1156:14–23, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075082/.

Bell:1985:HMS

[Bel85a] J. Bell. A hydrodynamical model of the sea hare’s propulsive mechanism. Lecture Notes in Mathematics, 1151:42–48, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074713/.

Bellissard:1985:SIQ

[Bel85b] J. Bellissard. Stability and instability in quantum mechanics. Lecture Notes in Mathematics, 1159:204–229, 1985. CODEN LN-MAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080334/.

Belbas:1987:NSC

[Bel87a] S. A. Belbas. Numerical solution of certain nonlinear parabolic partial differential equations. Lecture Notes in Mathematics, 1248: 5–14, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077411/.

Bell:1987:CFH

[Bel87b] S. Bell. Compactness of families of holomorphic mappings up to the boundary. Lecture Notes in Mathematics, 1268:29–42, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097294/.

Bell:1987:EBK

[Bel87c] S. Bell. Extendibility of the Bergman kernel function. Lecture Notes in Mathematics, 1276:33–41, 1987. CODEN LN-MAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078952/.

Bell:1989:SIP

[Bel89a] Murray G. Bell. Spaces of ideals of partial functions. Lecture Notes in Mathematics, 1401:1–4, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097327/.

Bellen:1989:PAS

[Bel89b] A. Bellen. Parallelism across the steps for difference and differential equations. Lecture Notes in Mathematics, 1386:22–35, 1989. CODEN LNMAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089229/.

Bemelmans:1988:FBP

[Bem88] Josef Bemelmans. Free boundary problems for the Navier–Stokes equations. Lecture Notes in Mathematics, 1357:96–115, 1988. CO-DEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082863/.

Beem:1985:TST

[BEMG85] John K. Beem, Paul E. Ehrlich, Steen Markvorsen, and Gregory J. Galloway. A Toponogov splitting theorem for Lorentzian manifolds. Lecture Notes in Mathematics, 1156:1–13, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075081/.

Bennett:1985:SCC

[Ben85a] M. K. Bennett. Separation conditions on convexity lattices. Lecture Notes in Mathematics, 1149:22–36, 1985. CODEN LN-MAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098453/.

Bensoussan:1985:SSP

[Ben85b] A. Bensoussan. Some singular perturbation problems arising in stochastic control. Lecture Notes in Mathematics, 1119:20– 31, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074778/.

Bentosela:1985:ESE

[Ben85c] F. Bentosela. Electrons of a solid in an external electric field. Lecture Notes in Mathematics, 1109:32–38, 1985. CODEN LN-MAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101535/.

Benson:1986:MFG

[Ben86a] D.-J. Benson. Modules for finite groups: Representation rings, quivers and varieties. Lecture Notes in Mathematics, 1178:1– 24, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075287/.

Bentosela:1986:RSD

[Ben86b] F. Bentosela. Resonance states in disordered systems. Lecture Notes in Mathematics, 1186:246–251, 1986. CODEN LN-MAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076844/.

Ben:1987:PTS

[Ben87] H. Ben Saad. Problème de type Skorohod. (French) []. Lecture Notes in Mathematics, 1235:15–38, 1987. CODEN LN-MAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072746/.

Bensoussan:1989:IFZ

[Ben89] Alain Bensoussan. On the integrated formulation of Zakai and Kushner equations. Lecture Notes in Mathematics, 1390:13– 23, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083932/.

Berard:1985:IIH

[Bér85a] Pierre H. Bérard. Isoperimetric inequalities, heat equation and geometric applications. Lecture Notes in Mathematics, 1156:24– 29, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075083/.

Bergeron:1985:ACM

[Ber85b] F. Bergeron. Une approche combinatoire de la méthode de Weisner. (French) []. Lecture Notes in Mathematics, 1171:111– 119, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076536/.

Berman:1985:LTS

[Ber85c] Simeon M. Berman. Limit theorems for sojourns of stochastic processes. Lecture Notes in Mathematics, 1153:40–71, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074945/.

Berndt:1985:RSR

[Ber85d] Bruce C. Berndt. Remarks on some of Ramanujan’s number theoretical discoveries found in his second notebook. Lecture Notes in Mathematics, 1122:47–55, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075749/.

Bertin:1985:ASN

[Ber85e] José Bertin. Automorphismes des surfaces non complètes, groupes fuchsiens et singularités quasihomogènes. (French) []. Lecture Notes in Mathematics, 1146:106–126, 1985. CODEN LN-MAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074536/.

Berard:1986:BSS

[Bér86a] Pierre H. Bérard. A brief survey of some recent developments in spectral geometry. Lecture Notes in Mathematics, 1207:144– 154, 1986. CODEN LNMAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076337/.

Berard:1986:GAI

[Bér86b] Pierre H. Bérard. Geometric applications of isoperimetric methods. Lecture Notes in Mathematics, 1207:116–143, 1986. CODEN LNMAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076336/.

Berard:1986:IM

[Bér86c] Pierre H. Bérard. Isoperimetric methods. Lecture Notes in Mathematics, 1207:78–97, 1986. CODEN LNMAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076334/.

Berard:1986:IMH

[Bér86d] Pierre H. Bérard. Isoperimetric methods and the heat equation. Lecture Notes in Mathematics, 1207:98–115, 1986. CODEN LN-MAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076335/.

Berard:1986:LRT

[Bér86e] Pierre H. Bérard. The Laplacian and related topics. Lecture Notes in Mathematics, 1207:44–77, 1986. CODEN LNMAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076333/.

Berard:1986:MPP

[Bér86f] Pierre H. Bérard. Motivations and the physical point of view. Lecture Notes in Mathematics, 1207:1–20, 1986. CODEN LN-MAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076331/.

Berard:1986:SGD

[Bér86g] Pierre H. Bérard. Spectral Geometry: Direct and Inverse Problems: With an Appendix by G. Besson, volume 1207 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 272 pp. LCCN QA3

.L28 no. 1207. URL http://link.springer.com/book/10.1007/
BFb0076330;   http://www.springerlink.com/content/978-
3-540-40958-8.

Berard:1986:TRG

[Bér86h] Pierre H. Bérard. Topics from Riemannian geometry. Lecture Notes in Mathematics, 1207:21–43, 1986. CODEN LN-MAA2. ISBN 3-540-16788-9 (print), 3-540-40958-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076332/.

Bergeron:1986:CRS

[Ber86i] F. Bergeron. Combinatorial representations of some Lie groups and Lie algebras. Lecture Notes in Mathematics, 1234:34–47, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072507/.

Bernardi:1986:SSN

[Ber86j] Marco Luigi Bernardi. On some singular nonlinear evolution equations. Lecture Notes in Mathematics, 1223:12–24, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099179/.

Bertero:1986:RML

[Ber86k] M. Bertero. Regularization methods for linear inverse problems. Lecture Notes in Mathematics, 1225:52–112, 1986. CODEN LN-MAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072660/.

Berenstein:1987:FW

[Ber87a] Carlos Berenstein. A final word. Lecture Notes in Mathematics, 1277:350, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078255.

Berger:1987:ESN

[Ber87b] M. S. Berger. The explicit solution of nonlinear ordinary and partial differential equations I. Conceptual ideas. Lecture Notes in Mathematics, 1248:15–27, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077412/.

Berndtsson:1987:CTI

[Ber87c] Bo Berndtsson. ∂_{b} and Carleson type inequalities. Lecture Notes in Mathematics, 1276:42–54, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078953/.

Bertoin:1987:TLI

[Ber87d] Jean Bertoin. Temps locaux et intégration stochastique pour les processus de Dirichlet. (French) []. Lecture Notes in Mathematics, 1247:191–205, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077634/.

Berg:1988:UMD

[Ber88a] Christian Berg. On the uniqueness of minimal definitizing polynomials for a sequence with finitely many negative squares. Lecture Notes in Mathematics, 1359:93–99, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086590/.

Bergh:1988:FBM

[Ber88b] Jöran Bergh. Functions of bounded mean oscillation and Hausdorff–Young type theorems. Lecture Notes in Mathematics, 1302:130–136, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078869/.

Berkes:1988:ADS

[Ber88c] István Berkes. On almost i.i.d. subsequences of the trigonometric system. Lecture Notes in Mathematics, 1332:54–63, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081611/.

Berndt:1989:RTP

[Ber89] Bruce C. Berndt. Ramanujan and the theory of prime numbers. Lecture Notes in Mathematics, 1395:122–139, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086402/.

Besson:1985:EQM

[Bes85] O. Besson. The entropy of quantum Markov states. Lecture Notes in Mathematics, 1136:81–89, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074462/.

Besson:1988:MEL

[Bes88] Gérard Besson. On the multiplicy of the eigenvalues of the Laplacian. Lecture Notes in Mathematics, 1339:32–53, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083045/.

Bouvier:1985:CPP

[BF85] Alain Bouvier and Marco Fontana. The catenarian property of the polynomial rings over a Prüfer domain. Lecture Notes in Mathematics, 1146:340–354, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074546/.

Benci:1988:BLT

[BF88a] Vieri Benci and Donato Fortunato. A “Birkhoff–Lewis” type result for non autonomous differential equations. Lecture Notes in Mathematics, 1324:85–96, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100784/.

Borisovich:1988:HMT

[BF88b] Yu. G. Borisovich and T. N. Fomenko. Homological methods in the theory of periodic and equivariant maps. Lecture Notes in Mathematics, 1334:21–41, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080421/.

Bolek:1986:MSD

[BFKP86] P. Bolek, J. Fořt, K. Kozel, and J. Polášek. Mathematical solution of direct and inverse problem for transonic cascade flows. Lecture Notes in Mathematics, 1192:361–366, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076094/.

Baratella:1988:BET

[BG88a] Paola Baratella and Luigi Gatteschi. The bounds for the error term of an asymptotic approximation of Jacobi polynomials. Lecture Notes in Mathematics, 1329:203–221, 1988. CODEN LN-MAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083360/.

Buser:1988:ANC

[BG88b] Peter Buser and Detlef Gromoll. On the almost negatively curved 3-sphere. Lecture Notes in Mathematics, 1339:78–85, 1988. CO-DEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0083048/.

Bergau:1989:NOO

[BG89a] Peter Bergau and D. Garbe. Non-orientable and orientable regular maps. Lecture Notes in Mathematics, 1398:29–42, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086237/.

Bielawski:1989:FPI

[BG89b] R. Bielawski and L. Górniewicz. A fixed point index approach to some differential equations. Lecture Notes in Mathematics, 1411: 9–14, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086436/.

Bourgain:1989:EBW

[BG89c] J. Bourgain and M. Gromov. Estimates of Bernstein widths of Sobolev spaces. Lecture Notes in Mathematics, 1376:176– 185, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090054/.

Byczkowski:1989:LKF

[BG89d] T. Byczkowski and P. Graczyk. Lévy–Kchinchine formula on vector spaces. Lecture Notes in Mathematics, 1391:23–32, 1989. CO-DEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0083377/.

Benkart:1989:ICH

[BGO^{+}89] G. M. Benkart, T. B. Gregory, J. M. Osborn, H. Strade, and R. L. Wilson. Isomorphism classes of Hamiltonian Lie algebras. Lecture Notes in Mathematics, 1373:42–57, 1989. CODEN LN-MAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088886/.

Bertrand:1985:PPQ

[BGR85] J. Bertrand, B. Gaveau, and G. Rideau. Poisson processes and quantum field theory: a model. Lecture Notes in Mathematics, 1136:74–80, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074461/.

Brown:1989:FPM

[BGS89] Robert F. Brown, Robert E. Greene, and Helga Schirmer. Fixed points of map extensions. Lecture Notes in Mathematics, 1411: 24–45, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086438/.

Borisovich:1986:GAS

[BGV86] Yuriǐ G. Borisovich, Yuriǐ E. Gliklikh, and A. M. Vershik. Global Analysis — Studies and Applications II, volume 1214 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 275 (est.) pp. LCCN QA3

.L28 no. 1214. URL http://link.springer.com/book/10.1007/
BFb0075955;   http://www.springerlink.com/content/978-
3-540-47084-7.

Bogvad:1986:CR

[BH86a] Rikard Bøgvad and Stephen Halperin. On a conjecture of roos. Lecture Notes in Mathematics, 1183:120–127, 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075453/.

Bouleau:1986:PAC

[BH86b] Nicolas Bouleau and Francis Hirsch. Propriétés d’absolue continuité dans les espaces de Dirichlet et application aux équations différentielles stochastiques. (French) []. Lecture Notes in Mathematics, 1204:131–161, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075717/.

Brown:1987:DIC

[BH87] R. C. Brown and Don Hinton. On the Dirichlet index conjecture. Lecture Notes in Mathematics, 1285:39–45, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080579/.

Berstein:1989:CWS

[BH89a] Israel Berstein and John R. Harper. Cogroups which are not suspensions. Lecture Notes in Mathematics, 1370:63–86, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085219/.

Bloom:1989:CPK

[BH89b] Walter R. Bloom and Herbert Heyer. Characterisation of potential kernels of transient convolution semigroups on a commutative hypergroup. Lecture Notes in Mathematics, 1379:21–35, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087842/.

Bouleau:1989:DRI

[BH89c] N. Bouleau and F. Hirsch. On the derivability, with respect to the initial data, of the solution of a stochastic differential equation with Lipschitz coefficients. Lecture Notes in Mathematics, 1393:

39–57, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085770/.

Bump:1989:SCR

[BH89d] Daniel Bump and Jeffrey Hoffstein. Some conjectured relationships between theta functions and Eisenstein series on the metaplectic group. Lecture Notes in Mathematics, 1383:1–11, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083566/.

Backelin:1988:MFH

[BHS88] Jörgen Backelin, Jürgen Herzog, and Herbert Sanders. Matrix factorizations of homogeneous polynomials. Lecture Notes in Mathematics, 1352:1–33, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082014/.

Burns:1989:SSF

[BHT89] John A. Burns, Terry L. Herdman, and Janos Turi. State-space formulation for functional differential equations of neutral-type. Lecture Notes in Mathematics, 1394:1–10, 1989. CODEN LN-MAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086747/.

Biane:1989:CAC

[Bia89] Philippe Biane. Comportement asymptotique de certaines fonctionnelles additives de plusieurs mouvements browniens. (French) []. Lecture Notes in Mathematics, 1372:198–233, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083974/.

Bigolin:1985:RIP

[Big85] Bruno Bigolin. Remarques sur les idéaux de polynômes. (French) []. Lecture Notes in Mathematics, 1165:14–29, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076143/.

Bihari:1985:PSS

[Bih85] I. Bihari. Periodic solutions of some second order nonlinear differential equations with periodic coefficients. Lecture Notes in Mathematics, 1151:49–59, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074714/.

Bingham:1986:ACL

[Bin86] Michael S. Bingham. On the assumptions of a central limit theorem for approximate martingale arrays on a group. Lecture Notes in Mathematics, 1210:1–12, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077167/.

Bingham:1989:FAP

[Bin89a] Michael S. Bingham. A Fourier-analytic proof that conditional probability distributions exist on a group. Lecture Notes in Mathematics, 1379:1–5, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087840/.

Bingham:1989:TTP

[Bin89b] N. H. Bingham. Tauberian theorems in probability theory. Lecture Notes in Mathematics, 1379:6–20, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087841/.

Biroli:1985:WOB

[Bir85] M. Biroli. Wiener obstacles for the biharmonic operator. Lecture Notes in Mathematics, 1151:60–67, 1985. CODEN LN-MAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074715/.

Bismut:1986:PG

[Bis86] Jean-Michel Bismut. Probability and geometry. Lecture Notes in Mathematics, 1206:1–60, 1986. CODEN LNMAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076299/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-40955-7/
1.pdf.

Baltus:1986:FBV

[BJ86] Christopher Baltus and William B. Jones. A family of best value regions for modified continued fractions. Lecture Notes in Mathematics, 1199:1–20, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075931/.

Barratt:1987:KIH

[BJM87] M. G. Barratt, J. D. S. Jones, and M. E. Mahowald. The Kervaire invariant and the Hopf invariant. Lecture Notes in Mathematics, 1286:135–173, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078740/.

Bjork:1989:ACN

[Bjö89a] Jan-Erik Björk. The Auslander condition on Noetherian rings. Lecture Notes in Mathematics, 1404:137–173, 1989. CODEN LN-MAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084075/.

Bourgeade:1989:PRH

[BJO89b] A. Bourgeade, H. Jourdren, and J. Ovadia. Problème de Riemann en hydrodynamique et applications. (French) []. Lecture Notes in Mathematics, 1402:37–55, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083867/.

Bano:1988:RFK

[BK88a] Shigetoshi Bando and Ryoichi Kobayashi. Ricci-flat Kähler metrics on affine algebraic manifolds. Lecture Notes in Mathematics, 1339:20–31, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083044/.

Boyle:1988:AMS

[BK88b] Mike Boyle and Wolfgang Krieger. Almost Markov and shift equivalent sofic systems. Lecture Notes in Mathematics, 1342: 33–93, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082823/.

Butkovic:1987:FAI

[BKK87] Davor Butković, Svetozar Kurepa, and Hrvoje Kraljević. Functional Analysis II, volume 1242 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 432 (est.) pp. LCCN QA3

.L28 no. 1242. URL http://link.springer.com/book/10.1007/
BFb0072440;   http://www.springerlink.com/content/978-
3-540-47876-8.

Butkovic:1987:AC

[BKS87] D. Butković, H. Kraljević, and N. Sarapa. On the almost convergence. Lecture Notes in Mathematics, 1242:396–417, 1987. CODEN LNMAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072446/.

Bourgain:1989:GFD

[BKT89] J. Bourgain, N. J. Kalton, and L. Tzafriri. Geometry of finite dimensional subspaces and quotients of L_{p}. Lecture Notes in Mathematics, 1376:138–175, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090053/.

Barchielli:1985:DOV

[BL85] A. Barchielli and G. Lupieri. Dilations of operation valued stochastic processes. Lecture Notes in Mathematics, 1136:57–66, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074459/.

Bouleau:1986:TLP

[BL86] N. Bouleau and D. Lamberton. Théorie de Littlewood–Paley– Stein et processus stables. (French) []. Lecture Notes in Mathematics, 1204:162–185, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075718/.

Bourgain:1988:PB

[BL88] J. Bourgain and J. Lindenstrauss. Projection bodies. Lecture Notes in Mathematics, 1317:250–270, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081746/.

Barchielli:1989:CSQ

[BL89a] A. Barchielli and G. Lupieri. Convolution semigroups in quantum probability and quantum stochastic calculus. Lecture Notes in Mathematics, 1396:107–127, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083548/.

Bourgain:1989:AES

[BL89b] J. Bourgain and J. Lindenstrauss. Almost Euclidean sections in spaces with a symmetric basis. Lecture Notes in Mathematics, 1376:278–288, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090062/.

Blackadar:1985:NST

[Bla85a] Bruce Blackadar. Noncommutative shape theory. Lecture Notes in Mathematics, 1132:38–45, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074878/.

Blass:1985:KDU

[Bla85b] Andreas Blass. Kleene degrees of ultrafilters. Lecture Notes in Mathematics, 1141:29–48, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076213/.

Bland:1987:IPO

[Bla87] J. S. Bland. The imbedding problem for open complex manifolds. Lecture Notes in Mathematics, 1268:43–59, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097295/.

Blass:1989:ASF

[Bla89] Andreas Blass. Applications of superperfect forcing and its relatives. Lecture Notes in Mathematics, 1401:18–40, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097329/.

Bleiler:1985:PTC

[Ble85a] Steven A. Bleiler. Prime tangles and composite knots. Lecture Notes in Mathematics, 1144:1–13, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075008/.

Bleistein:1985:IMR

[Ble85b] Norman Bleistein. Inverse methods for reflector imaging. Lecture Notes in Mathematics, 1151:68–76, 1985. CODEN LN-MAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074716/.

Blel:1986:SEA

[Ble86] M. Blel. Sur les singularités des ensembles analytiques. (French) []. Lecture Notes in Mathematics, 1198:65–72, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077043/.

Bliznyakov:1986:CII

[Bli86] N. M. Bliznyakov. Cauchy indices and the index of a singular point of a vector field. Lecture Notes in Mathematics, 1214:1–20, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075956/.

Bliznyakov:1988:TIE

[Bli88] N. M. Bliznyakov. Topological index estimates. Lecture Notes in Mathematics, 1334:180–198, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080429/.

Benachour:1989:ANC

[BLL89] Saïd Benachour, Alain Yves Le Roux, and Marie-Noëlle Le Roux. Approximation to nonlinear convection diffusion equations. Lecture Notes in Mathematics, 1402:1–11, 1989. CODEN LN-MAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083864/.

Bourgain:1988:MSS

[BLM88] J. Bourgain, J. Lindenstrauss, and V. D. Milman. Minkowski sums and symmetrizations. Lecture Notes in Mathematics, 1317: 44–66, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081735/.

Bourgain:1989:ERS

[BLM89] J. Bourgain, J. Lindenstrauss, and V. Milman. Estimates related to Steiner symmetrizations. Lecture Notes in Mathematics, 1376:264–273, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090060/.

Blomstrom:1985:SIP

[Blo85] Carol Blomstrom. Symmetric immersions in pseudo-Riemannian space forms. Lecture Notes in Mathematics, 1156:30–45, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075084/.

Bloom:1986:IMC

[Blo86] Walter R. Bloom. Idempotent measures on commutative hypergroups. Lecture Notes in Mathematics, 1210:13–23, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077168/.

Bourgeade:1987:ASG

[BLR87] A. Bourgeade, P. Le Floch, and P.-A. Raviart. Approximate solution of the generalized Riemann problem and applications. Lecture Notes in Mathematics, 1270:1–9, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078313/.

Blum:1985:ALE

[Blu85a] H. Blum. On the approximation of linear elliptic systems on polygonal domains. Lecture Notes in Mathematics, 1121:28–37, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076260/.

Blumenfeld:1985:RIP

[Blu85b] Matthias Blumenfeld. The regularity of interface-problems on corner-regions. Lecture Notes in Mathematics, 1121:38–54, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076261/.

Blumenthal:1985:CFM

[Blu85c] Robert A. Blumenthal. Connections on foliated manifolds. Lecture Notes in Mathematics, 1165:30–35, 1985. CODEN LN-MAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076144/.

Blumenthal:1986:MBM

[Blu86] R. A. Blumenthal. Mappings between manifolds with Cartan connections. Lecture Notes in Mathematics, 1209:94–99, 1986. CO-DEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076622/.

Biane:1987:PQR

[BLY87] Ph. Biane, J.-F. Le Gall, and M. Yor. Un processus qui ressemble au pont brownien. (French) []. Lecture Notes in Mathematics, 1247:270–275, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077641/.

Bulinskaya:1985:ABS

[BM85] E. V. Bulinskaya and S. A. Molčanov. Asymptotic behaviour of some random fields. Lecture Notes in Mathematics, 1155:21– 46, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074811/.

Bank:1986:CMG

[BM86] R. E. Bank and H. D. Mittelmann. Continuation and multigrid for nonlinear elliptic systems. Lecture Notes in Mathematics, 1228:23–37, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072639/.

Barlet:1987:DAT

[BM87] D. Barlet and H.-M. Maire. Développements asymptotiques, transformation de Mellin complexe et intégration sur les fibres. (French) []. Lecture Notes in Mathematics, 1295:11–23, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081975/.

Barreto:1988:END

[BM88] Antônio Sà Barreto and Richard B. Melrose. Examples of nondiscreteness for the interaction geometry of semilinear progressing waves in two space dimensions. Lecture Notes in Mathematics, 1324:186–196, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100792/.

Bachmuth:1989:IGA

[BM89a] Seymour Bachmuth and H. Y. Mochizuki. Infinite generation of automorphism groups. Lecture Notes in Mathematics, 1398:25– 28, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086236/.

Bolondi:1989:CLP

[BM89b] Giorgio Bolondi and Juan Carlos Migliore. Configurations of linear projective subvarieties. Lecture Notes in Mathematics, 1389: 19–31, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085921/.

Borkar:1989:LAS

[BM89c] Vivek S. Borkar and Sanjoy K. Mitter. Lattice approximation in the stochastic quantization of (ϕ^{4})_{2} fields. Lecture Notes in Mathematics, 1390:24–32, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083933/.

Boyer:1989:IH

[BM89d] Charles P. Boyer and Benjamin M. Mann. Instantons and homotopy. Lecture Notes in Mathematics, 1370:87–102, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085220/.

Bourgain:1988:GI

[BMMP88] J. Bourgain, M. Meyer, V. Milman, and A. Pajor. On a geometric inequality. Lecture Notes in Mathematics, 1317:271–282, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081747/.

Bruner:1986:RST

[BMMS86] Robert R. Bruner, J. Peter May, James E. McClure, and Mark Steinberger. H_{∞} Ring Spectra and their Applications, volume 1176 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 377 (est.) pp. LCCN QA3 .L28 no. 1176. URL http:

//link.springer.com/book/10.1007/BFb0075405;   http://
www.springerlink.com/content/978-3-540-39778-6.

Beilinson:1987:SVN

[BMS87] A. A. Beĭlinson, Yu. I. Manin, and V. V. Schechtman. Sheaves of the Virasoro and Neveu–Schwarz algebras. Lecture Notes in Mathematics, 1289:52–66, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078367/.

Bion-Nadal:1985:BBA

[BN85] J. Bion-Nadal. Banach bimodule associated to an action of a discrete group on a compact space. Lecture Notes in Mathematics, 1132:30–37, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074877/.

Brietzke:1988:MAM

[BN88] Eduardo Brietzke and Pedro Nowosad. Mathematical aspects of the minimum critical mass problem. Lecture Notes in Mathematics, 1324:53–64, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100782/.

Babuska:1985:FEM

[BO85] I. Babuška and J. E. Osborn. Finite element methods for the solution of problems with rough input data. Lecture Notes in Mathematics, 1121:1–18, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076258/.

Bodigheimer:1987:SSM

[Böd87] C.-F. Bödigheimer. Stable splittings of mapping spaces. Lecture Notes in Mathematics, 1286:174–187, 1987. CODEN LN-MAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078741/.

Bohme:1986:GLE

[Böh86] Olaf Böhme. On a generalization of the Lyapunov exponent of a second order linear stochastic differential equation. Lecture Notes in Mathematics, 1186:191–199, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076839/.

Bohnke:1988:ATB

[Boh88] Georges Bohnké. Algèbres de type Besov–Sobolev. (French) []. Lecture Notes in Mathematics, 1359:113–121, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086592/.

Bojarski:1988:RSI

[Boj88a] B. Bojarski. Remarks on Sobolev imbedding inequalities. Lecture Notes in Mathematics, 1351:52–68, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081242/.

Bojarski:1988:RLF

[Boj88b] Bogdan Bojarski. Remarks on local function spaces. Lecture Notes in Mathematics, 1302:137–152, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078870/.

Bokut:1988:SNR

[Bok88] L. A. Bokut’. Some new results in the combinatorial theory of rings and groups. Lecture Notes in Mathematics, 1352:34–43, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082015/.

Boldrini:1988:CSC

[Bol88] José Luiz Boldrini. Convergence of solutions of capillo-viscoelastic perturbations of the equations of elasticity. Lecture Notes in Mathematics, 1324:41–52, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100781/.

Bombieri:1987:TME

[Bom87] E. Bombieri. On the Thue–Mahler equation. Lecture Notes in Mathematics, 1290:213–243, 1987. CODEN LNMAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078711/.

Bongartz:1985:QFF

[Bon85] Klaus Bongartz. Quadratic forms and finite representation type. Lecture Notes in Mathematics, 1146:325–339, 1985. CODEN LN-MAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074545/.

Bony:1989:AMS

[Bon89] Jean-Michel Bony. Analyse microlocale et singularités non linéaires. (French) []. Lecture Notes in Mathematics, 1402:114– 127, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083871/.

Borell:1985:IBS

[Bor85] Christer Borell. Intrinsic bounds on some real-valued stationary random functions. Lecture Notes in Mathematics, 1153:72– 95, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074946/.

Borell:1986:BSA

[Bor86] Christer Borell. A brief survey of Antoine Ehrhard’s scientific work. Lecture Notes in Mathematics, 1193:1–3, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077095/.

Borel:1987:VTR

[Bor87a] Armand Borel. A vanishing theorem in relative Lie algebra cohomology. Lecture Notes in Mathematics, 1271:1–16, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079230/.

Borho:1987:NOP

[Bor87b] Walter Borho. Nilpotent orbits, primitive ideals, and characteristic classes. Lecture Notes in Mathematics, 1271:17–32, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079231/.

Borisovich:1988:POB

[Bor88a] A. Yu. Borisovich. Plateau operator and bifurcations of twodimensional minimal surfaces. Lecture Notes in Mathematics, 1334:1–20, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080420/.

Borisovich:1988:MAT

[Bor88b] Yu. G. Borisovich. Modern approach to the theory of topological characteristics of nonlinear operators I. Lecture Notes in Mathematics, 1334:199–220, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080430/.

Borisovich:1989:TCI

[Bor89] Yu. G. Borisovich. Topological characteristics of infinite- dimensional mappings and the theory of fixed points and coincidences. Lecture Notes in Mathematics, 1411:15–23, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086437/.

Bose:1986:ADB

[Bos86] A. Bose. Age distributions in birth and death processes. Lecture Notes in Mathematics, 1212:52–58, 1986. CODEN LN-MAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076237/.

Bourgain:1985:SRB

[Bou85a] J. Bourgain. Some remarks on the Banach space structure of the ball-algebras. Lecture Notes in Mathematics, 1166:4–10, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074686/.

Bourgain:1985:SAD

[Bou85b] J. Bourgain. Subspaces of L_{N}^{∞}, arithmetical diameter and Sidon sets. Lecture Notes in Mathematics, 1153:96–127, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074947/.

Bougerol:1986:OPM

[Bou86a] Philippe Bougerol. Oscillation des produits de matrices aléatoires dont l’exposant de Lyapounov est nul. (French) []. Lecture Notes in Mathematics, 1186:27–36, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076831/.

Bourguignon:1986:IIF

[Bou86b] Jean-Pierre Bourguignon. Invariants intégraux fonctionnels pour des équations aux derivées partielles d’origine géométrique. (French) []. Lecture Notes in Mathematics, 1209:100–108, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076623/.

Bouleau:1987:AVC

[Bou87a] Nicolas Bouleau. Autour de la variance comme forme de Dirichlet: filtrations et résolutions de l’identité contractions et BMO, espérances conditionnelles et principe complet du maximum. (French) []. Lecture Notes in Mathematics, 1235:39–53, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072747/.

Bouleau:1987:ELD

[Bou87b] Nicolas Bouleau. Énergie locale et densité de temps d’occupation. (French) []. Lecture Notes in Mathematics, 1235:54–56, 1987.

CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072748/.

Bourgain:1987:DCA

[Bou87c] J. Bourgain. A density condition for analyticity of the restriction algebra. Lecture Notes in Mathematics, 1267:151–156, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078142/.

Bourgain:1987:DFM

[Bou87d] J. Bourgain. On dimension free maximal inequalities for convex symmetric bodies in R^{\}. Lecture Notes in Mathematics, 1267: 168–176, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078144/.

Bourgain:1987:LPC

[Bou87e] J. Bourgain. On lattice packing of convex symmetric sets in R^{\}. Lecture Notes in Mathematics, 1267:5–12, 1987. CODEN LN-MAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078132/.

Bourgain:1987:REL

[Bou87f] J. Bourgain. Remarks on the extension of Lipschitz maps defined on discrete sets and uniform homeomorphisms. Lecture Notes in Mathematics, 1267:157–167, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078143/.

Bourguignon:1987:YMT

[Bou87g] Jean-Pierre Bourguignon. Yang–Mills theory: The differential geometric side. Lecture Notes in Mathematics, 1263:13–54, 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078609/.

Bourgain:1988:APE

[Bou88a] J. Bourgain. An approach to pointwise ergodic theorems. Lecture Notes in Mathematics, 1317:204–223, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081742/.

Bourgain:1988:FDH

[Bou88b] J. Bourgain. On finite dimensional homogeneous Banach spaces. Lecture Notes in Mathematics, 1317:232–238, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081744/.

Bourgain:1988:VVH

[Bou88c] J. Bourgain. Vector-valued Hausdorff–Young inequalities and applications. Lecture Notes in Mathematics, 1317:239–249, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081745/.

Bourn:1988:PNA

[Bou88d] Dominique Bourn. Pseudofunctors and non-abelian weak equivalences. Lecture Notes in Mathematics, 1348:55–71, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081348/.

Bourgain:1989:KRP

[Bou89a] J. Bourgain. On Kolmogorov’s rearrangement problem for orthogonal systems and Garsia’s conjecture. Lecture Notes in Mathematics, 1376:209–250, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090057/.

Bourgain:1989:BCL

[Bou89b] J. Bourgain. On the behavior of the constant in the Littlewood– Paley inequality. Lecture Notes in Mathematics, 1376:202–208, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090056/.

Boyle:1988:NSA

[Boy88] M. Boyle. Nasu’s simple automorphisms. Lecture Notes in Mathematics, 1342:23–32, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082822/.

Baronti:1988:NSC

[BP88a] Marco Baronti and Pier Luigi Papini. Nearby sets and centers. Lecture Notes in Mathematics, 1354:98–105, 1988. CODEN LN-MAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089585/.

Bondesson:1988:CM

[BP88b] Lennart Bondesson and Jaak Peetre. The classes V_{a} are monotone. Lecture Notes in Mathematics, 1302:153–157, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078871/.

Borceux:1988:DTB

[BP88c] Francis Borceux and Joan Wick Pelletier. Descent theory for Banach modules. Lecture Notes in Mathematics, 1348:36–54, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081347/.

Burghelea:1988:CHC

[BP88d] Dan Burghelea and Micheline Vigué Poirrier. Cyclic homology of commutative algebras I. Lecture Notes in Mathematics, 1318: 51–72, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077794/.

Billey:1987:DDE

[BPPS87] V. Billey, J. Périaux, P. Perrier, and B. Stoufflet. 2-D and 3-D Euler computations with finite element methods in aerodynamics. Lecture Notes in Mathematics, 1270:64–81, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078318/.

Bourgain:1989:DPE

[BPSTJ89] J. Bourgain, A. Pajor, S. J. Szarek, and N. Tomczak-Jaegermann. On the duality problem for entropy numbers of operators. Lecture Notes in Mathematics, 1376:50–63, 1989. CODEN LN-MAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090048/.

Burridge:1986:SRP

[BPW86] R. Burridge, G. Papanicolaou, and B. White. Statistics of reflected pulses. Lecture Notes in Mathematics, 1186:265–270, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076847/.

Balasubramanian:1985:HVT

[BR85a] R. Balasubramanian and K. Ramachandra. A hybrid version of a theorem of Ingham. Lecture Notes in Mathematics, 1122:38– 46, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075748/.

Bertrand:1985:IAE

[BR85b] J. Bertrand and G. Rideau. An intrinsic approach to the evolution of quantum observables in terms of stochastic processes on phase space. Lecture Notes in Mathematics, 1109:39–54, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101536/.

Bushnell:1985:SAM

[BR85c] Colin J. Bushnell and Irving Reiner. A survey of analytic methods in noncommutative number theory. Lecture Notes in Mathematics, 1142:50–87, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074792/.

Backelin:1986:WDY

[BR86] Jörgen Backelin and Jan-Erik Roos. When is the double Yoneda ext-algebra of a local Noetherian ring again Noetherian? Lecture Notes in Mathematics, 1183:101–119, 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075452/.

Bergelson:1988:WMA

[BR88] Vitaly Bergelson and Daniel J. Rudolph. Weakly mixing actions of F^{∞} have infinite subgroup actions which are Bernoulli. Lecture Notes in Mathematics, 1342:7–22, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082821/.

Berrick:1989:EMI

[BR89] A. J. Berrick and Duane Randall. On equivariant maps and immersions of real projective spaces. Lecture Notes in Mathematics, 1370:57–62, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085218/.

Braams:1986:MEC

[Bra86a] B. J. Braams. Magnetohydrodynamic equilibrium calculations using multigrid. Lecture Notes in Mathematics, 1228:38–51, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072640/.

Braess:1986:CMM

[Bra86b] Dietrich Braess. On the combination of the multigrid method and conjugate gradients. Lecture Notes in Mathematics, 1228: 52–64, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072641/.

Brailov:1986:CIN

[Bra86c] A. V. Brailov. Complete integrability with noncommuting integrals of certain Euler equations. Lecture Notes in Mathematics, 1214:21–39, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075957/.

Bramble:1986:CPE

[Bra86d] J. H. Bramble. The construction of preconditioners for elliptic problems by substructuring. Lecture Notes in Mathematics, 1230:158–166, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072678/.

Brandt:1986:MLA

[BRA86e] Achi Brandt, Dorit Ron, and Daniel J. Amit. Multi-level approaches to discrete-state and stochastic problems. Lecture Notes in Mathematics, 1228:65–98, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072642/.

Bratteli:1986:AR

[Bra86f] Ola Bratteli. Additional remarks. Lecture Notes in Mathematics, 1229:261–264, 1986. CODEN LNMAA2. ISBN 3-540-17199-1 (print), 3-540-47376-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098822/.

Bratteli:1986:DDG

[Bra86g] Ola Bratteli. Derivations, Dissipations and Group Actions on C∗- algebras, volume 1229 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17199-1 (print), 3-540-47376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 277 pp. LCCN QA3 .L28 no. 1229. URL http:

//link.springer.com/book/10.1007/BFb0098817;   http://
www.springerlink.com/content/978-3-540-47376-3.

Bratteli:1986:D

[Bra86h] Ola Bratteli. Dissipations. Lecture Notes in Mathematics, 1229: 241–260, 1986. CODEN LNMAA2. ISBN 3-540-17199-1 (print), 3-540-47376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098821/.

Bratteli:1986:GTD

[Bra86i] Ola Bratteli. General theory of derivations. Lecture Notes in Mathematics, 1229:4–33, 1986. CODEN LNMAA2. ISBN 3-540-17199-1 (print), 3-540-47376-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098819/.

Bratteli:1986:IHR

[Bra86j] Ola Bratteli. Introduction with historical remarks. Lecture Notes in Mathematics, 1229:1–3, 1986. CODEN LNMAA2. ISBN 3-540-17199-1 (print), 3-540-47376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098818/.

Bratteli:1986:NV

[Bra86k] Ola Bratteli. Noncommutative vectorfields. Lecture Notes in Mathematics, 1229:34–240, 1986. CODEN LNMAA2. ISBN 3-540-17199-1 (print), 3-540-47376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098820/.

Braam:1987:QFT

[Bra87] Peter Braam. Quantum field theory: the bridge between mathematics and the physical world. Lecture Notes in Mathematics, 1263:1–12, 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078608/.

Bracalova:1988:GTT

[Bra88] M. A. Bracalova. A generalization of the Teichmüller theorem. Lecture Notes in Mathematics, 1351:69–77, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081243/.

Brezzi:1985:NMF

[Bre85] Franco Brezzi, editor. Numerical Methods in Fluid Dynamics: Lectures given at the 3rd 1983 Session of the Centro Internationale Matematico Estivo (C.I.M.E.) held at Como, Italy, July 7–15, 1983, volume 1127 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15225-3 (print), 3-540-39391-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1127. URL http:

//link.springer.com/book/10.1007/BFb0074528;   http://
www.springerlink.com/content/978-3-540-39391-7.

Brenner:1986:CCF

[Bre86a] Sheila Brenner. A combinatorial characterisation of finite Auslander–Reiten quivers. Lecture Notes in Mathematics, 1177: 13–49, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075256/.

Bressoud:1986:DIE

[Bre86b] D. M. Bressoud. Definite integral evaluation by enumeration, partial results in the MacDonald conjectures. Lecture Notes in Mathematics, 1234:48–57, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072508/.

Brezzi:1986:RRA

[Bre86c] F. Brezzi. Recent results in the approximation of free boundaries. Lecture Notes in Mathematics, 1192:285–289, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076082/.

Brennan:1987:FRD

[Bre87] James E. Brennan. Functions with rapidly decreasing negative Fourier coefficients. Lecture Notes in Mathematics, 1275: 31–43, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078343/.

Brezinski:1988:EEP

[Bre88] Claude Brezinski. Error estimate in Padé approximation. Lecture Notes in Mathematics, 1329:1–19, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083350/.

Bressoud:1989:LPR

[Bre89a] David M. Bressoud. Lattice paths and the Rogers–Ramanujan identities. Lecture Notes in Mathematics, 1395:140–172, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086403/.

Brezis:1989:VMS

[Bré89b] Haïm Brézis. S^{k}-valued maps with singularities. Lecture Notes in Mathematics, 1365:1–30, 1989. CODEN LNMAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089176/.

Brinkhuis:1985:NIB

[Bri85a] Jan Brinkhuis. Normal integral bases and embeddings of fields. Lecture Notes in Mathematics, 1142:5–17, 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074790/.

Britton:1985:TWF

[Bri85b] N. F. Britton. Travelling wave front solutions of a differentialdifference equation arising in the modelling of myelinated nerve axon. Lecture Notes in Mathematics, 1151:77–89, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074717/.

Brilla:1986:SAN

[Bri86] I. Brilla. Spectral analysis of non-self-adjoint elliptic operators. Lecture Notes in Mathematics, 1192:197–202, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076069/.

Brion:1987:IAM

[Bri87] Michel Brion. Sur l’image de l’application moment. (French) []. Lecture Notes in Mathematics, 1296:177–192, 1987. CODEN LN-MAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078526/.

Brown:1985:OSN

[Bro85] Kenneth A. Brown. Ore sets in Noetherian rings. Lecture Notes in Mathematics, 1146:355–366, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074547/.

Brooks:1986:CPS

[Bro86] Robert Brooks. Combinatorial problems in spectral geometry. Lecture Notes in Mathematics, 1201:14–32, 1986. CODEN LN-MAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075645/.

Brownawell:1987:LTDa

[Bro87] W. Dale Brownawell. Large transcendence degree revisited. Lecture Notes in Mathematics, 1290:149–173, 1987. CODEN LN-MAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078707/.

Brownawell:1989:ACC

[Bro89] W. Dale Brownawell. Applications of Cayley–Chow forms. Lecture Notes in Mathematics, 1380:1–18, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086542/.

Bruning:1985:SAS

[Brü85] Jochen Brüning. Spectral analysis of singular Sturm–Liouville problems with operator coefficients. Lecture Notes in Mathematics, 1121:55–68, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076262/.

Bruner:1986:ASS

[Bru86a] Robert R. Bruner. The Adams spectral sequence of H_{∞} ring spectra. Lecture Notes in Mathematics, 1176:169–214, 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075411/.

Bruner:1986:HGR

[Bru86b] Robert R. Bruner. The homotopy groups of H_{∞} ring spectra. Lecture Notes in Mathematics, 1176:129–168, 1986. CODEN LN-MAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075410/.

Bruner:1986:HTR

[Bru86c] Robert R. Bruner. The homotopy theory of H_{∞} ring spectra. Lecture Notes in Mathematics, 1176:88–128, 1986. CODEN LN-MAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075409/.

Bruna:1987:BAC

[Bru87a] Joaquim Bruna. Boundary absolute continuity of functions in the ball algebra. Lecture Notes in Mathematics, 1276:55–59, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078954/.

Bruning:1987:CCR

[Brü87b] E. Brüning. A constructive characterization of Radon probability measures on infinite dimensional spaces. Lecture Notes in Mathematics, 1250:14–30, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077345/.

Bruning:1987:ITR

[Brü87c] J. Brüning. Index theory for regular singular operators and applications. Lecture Notes in Mathematics, 1256:36–54, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077737/.

Bruzzo:1987:LGT

[Bru87d] Ugo Bruzzo. Lagrangian gauge theories on supermanifolds. Lecture Notes in Mathematics, 1251:91–113, 1987. CODEN LN-MAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077317/.

Brungs:1988:CRV

[Bru88] H.-H. Brungs. Chain rings and valuations. Lecture Notes in Mathematics, 1328:33–37, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100912/.

Bruna:1989:SMF

[Bru89] Joaquim Bruna. On the size the maximal function and the Hilbert transform. Lecture Notes in Mathematics, 1384:124–131, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086797/.

Bryant:1985:MHS

[Bry85] Robert L. Bryant. Metrics with holonomy G_{2} or spin (7). Lecture Notes in Mathematics, 1111:269–277, 1985. CODEN LN-MAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084595/.

Bryant:1987:MLS

[Bry87a] Robert L. Bryant. Minimal Lagrangian submanifolds of Kähler– Einstein manifolds. Lecture Notes in Mathematics, 1255:1–12, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077676/.

Brylinski:1987:SEH

[Bry87b] Jean-Luc Brylinski. Some examples of Hochschild and cyclic homology. Lecture Notes in Mathematics, 1271:33–72, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079232/.

Bryant:1989:RSV

[Bry89a] Peter Bryant. Representing the super Virasoro algebra by meromorphic vectorfields on the graded Riemann sphere. Lecture Notes in Mathematics, 1410:71–76, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086412/.

Brylinski:1989:SCM

[Bry89b] Ranee Kathryn Brylinski. Stable calculus of the mixed tensor character I. Lecture Notes in Mathematics, 1404:35–94, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084071/.

Brzezinski:1985:BSS

[Brz85] J. Brzezinski. Brauer–Severi schemes of orders. Lecture Notes in Mathematics, 1142:18–49, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074791/.

Bastero:1986:PBS

[BS86] Jesús Bastero and Miguel San Miguel, editors. Probability and Banach Spaces: Proceedings of a Conference held in Zaragoza, Spain June 17–21, 1985, volume 1221 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 0-387-17186-X (paperback), 3-540-17186-X (paperback), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1221. URL http://link.springer.com/book/10.1007/
BFb0099107;   http://www.springerlink.com/content/978-
3-540-47344-2.

Baldwin:1987:ELW

[BS87a] John T. Baldwin and Charles Steinhorn. Exercises on local weight. Lecture Notes in Mathematics, 1292:127–131, 1987. CODEN LN-MAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082235/.

Beals:1987:HEG

[BS87b] Richard Beals and Nancy K. Stanton. The heat equation and geometry for the ¯∂-Neumann problem. Lecture Notes in Mathematics, 1276:25–32, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078951/.

Berenstein:1987:RCS

[BS87c] Carlos A. Berenstein and Daniele C. Struppa. A remark on “Convolutors in spaces of holomorphic functions”. Lecture Notes in Mathematics, 1276:276–280, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078963/. See [MS87a].

Bokowski:1989:ACG

[BS89a] Jürgen Bokowski and Bernd Sturmfels. Algebraic criteria for geometric realizability. Lecture Notes in Mathematics, 1355:61– 86, 1989. CODEN LNMAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089257/.

Bokowski:1989:CAM

[BS89b] Jürgen Bokowski and Bernd Sturmfels. Combinatorial and algebraic methods. Lecture Notes in Mathematics, 1355:32–60, 1989. CODEN LNMAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089256/.

Bokowski:1989:GM

[BS89c] Jürgen Bokowski and Bernd Sturmfels. Geometric methods. Lecture Notes in Mathematics, 1355:87–101, 1989. CODEN LN-MAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089258/.

Bokowski:1989:MCA

[BS89d] Jürgen Bokowski and Bernd Sturmfels. Matroids and chirotopes as algebraic varieties. Lecture Notes in Mathematics, 1355:147– 157, 1989. CODEN LNMAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089262/.

Bokowski:1989:EA

[BS89e] Jürgen Bokowski and Bernd Sturmfels. On the existence of algorithms. Lecture Notes in Mathematics, 1355:18–31, 1989. CODEN LNMAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089255/.

Bokowski:1989:FPM

[BS89f] Jürgen Bokowski and Bernd Sturmfels. On the finding of polyheadral manifolds. Lecture Notes in Mathematics, 1355:133– 146, 1989. CODEN LNMAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089261/.

Bokowski:1989:P

[BS89g] Jürgen Bokowski and Bernd Sturmfels. Preliminaries. Lecture Notes in Mathematics, 1355:1–17, 1989. CODEN LNMAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089254/.

Bokowski:1989:PM

[BS89h] Jürgen Bokowski and Bernd Sturmfels. Preprocessing methods. Lecture Notes in Mathematics, 1355:115–132, 1989. CODEN LN-MAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089260/.

Bokowski:1989:RTR

[BS89i] Jürgen Bokowski and Bernd Sturmfels. Recent topological results. Lecture Notes in Mathematics, 1355:102–114, 1989. CODEN LN-MAA2. ISBN 3-540-50478-8 (print), 3-540-46013-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089259/.

Brown:1989:RHT

[BS89j] Edgar H. Brown, Jr. and Robert H. Szczarba. On real homotopy theory. Lecture Notes in Mathematics, 1370:103–116, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085221/.

Baldoni-Silva:1987:IUR

[BSK87a] M. W. Baldoni-Silva and A. W. Knapp. Irreducible unitary representations of some groups of real rank two. Lecture Notes in Mathematics, 1243:15–36, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073016/.

Baldoni-Silva:1987:VAC

[BSK87b] M. W. Baldoni-Silva and A. W. Knapp. Vogan’s algorithm for computing composition series. Lecture Notes in Mathematics, 1243:37–72, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073017/.

Beauquier:1986:PFW

[BT86a] Joffroy Beauquier and Loÿs Thimonier. Prefix-free words of length N over M letters: two-sided well-balanced parentheses and palindroms. Lecture Notes in Mathematics, 1234:27–33, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072506/.

Boggs:1986:ESU

[BT86b] Paul T. Boggs and Jon W. Tolle. An efficient strategy for utilizing a merit function in nonlinear programming algorithms. Lecture Notes in Mathematics, 1230:127–137, 1986. CODEN LN-MAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072676/.

Brandt:1986:MMN

[BT86c] A. Brandt and S. Ta’asan. Multigrid method for nearly singular and slightly indefinite problems. Lecture Notes in Mathematics, 1228:99–121, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072643/.

Bretscher:1986:TNR

[BT86d] Otto Bretscher and Gordana Todorov. On a theorem of Nazarova and Roiter. Lecture Notes in Mathematics, 1177:50–54, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075257/.

Bohmer:1987:SIB

[BT87a] W. Böhmer and G. Trautmann. Special instanton bundles and Poncelet curves. Lecture Notes in Mathematics, 1273:325–336, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078852/.

Bourgain:1987:CSW

[BT87b] J. Bourgain and L. Tzafriri. Complements of subspaces of ℓ_{p}^{n}; p ≥ 1 which are uniquely determined. Lecture Notes in Mathematics, 1267:39–52, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078135/.

Brownawell:1987:LTDb

[BT87c] W. Dale Brownawell and Robert Tubbs. Large transcendence degree revisited. Lecture Notes in Mathematics, 1290:175–188, 1987. CODEN LNMAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078708/.

Batchelor:1988:FHT

[BT88a] Andrew Batchelor and Aubrey Truman. On first hitting times in stochastic mechanics. Lecture Notes in Mathematics, 1325:1– 15, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077912/.

Blazquez:1988:HBB

[BT88b] M. Blázquez and E. Tuma. Hereroclinic bifurcation in Banach spaces. Lecture Notes in Mathematics, 1331:12–37, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083063/.

Buchta:1985:ZPO

[Buc85] Christian Buchta. Zufällige Polyeder — Eine Obersicht. (German) [Random polyhedra — a summary]. Lecture Notes in Mathematics, 1114:1–13, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101638/.

Buekenhout:1986:GFS

[Bue86] Francis Buekenhout. The geometry of the finite simple groups. Lecture Notes in Mathematics, 1181:1–78, 1986. CODEN LN-MAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075512/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-39801-1/
1.pdf.

Buechler:1987:CSW

[Bue87] Steven Buechler. The classification of small weakly minimal sets I. Lecture Notes in Mathematics, 1292:32–71, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082231/.

Buium:1986:C

[Bui86a] Alexandru Buium. Complements. Lecture Notes in Mathematics, 1226:121–139, 1986. CODEN LNMAA2. ISBN 3-540-17194-0 (print), 3-540-47354-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101626/.

Buium:1986:DDT

[Bui86b] Alexandru Buium. Differential descent theory. Lecture Notes in Mathematics, 1226:28–86, 1986. CODEN LNMAA2. ISBN 3-540-17194-0 (print), 3-540-47354-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101624/.

Buium:1986:DFF

[Bui86c] Alexandru Buium. Differential Function Fields and Moduli of Algebraic Varieties, volume 1226 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17194-0 (print), 3-540-47354-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 146 pp. LCCN QA3 .L28

no. 1226. URL http://link.springer.com/book/10.1007/
BFb0101622;   http://www.springerlink.com/content/978-
3-540-47354-1.

Buium:1986:NDG

[Bui86d] Alexandru Buium. Normality in differential Galois theory. Lecture Notes in Mathematics, 1226:87–120, 1986. CODEN LN-MAA2. ISBN 3-540-17194-0 (print), 3-540-47354-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101625/.

Buium:1986:P

[Bui86e] Alexandru Buium. Preliminaries. Lecture Notes in Mathematics, 1226:1–27, 1986. CODEN LNMAA2. ISBN 3-540-17194-0 (print), 3-540-47354-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101623/.

Bumby:1985:HDS

[Bum85] Richard T. Bumby. Hausdorff dimension of sets arising in number theory. Lecture Notes in Mathematics, 1135:1–8, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074599/.

Burkholder:1986:MFA

[Bur86] Donald L. Burkholder. Martingales and Fourier analysis in Banach spaces. Lecture Notes in Mathematics, 1206:61–108, 1986. CODEN LNMAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076300/.

Burstall:1987:TMH

[Bur87] Francis E. Burstall. Twistor methods for harmonic maps. Lecture Notes in Mathematics, 1263:55–96, 1987. CODEN LN-MAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078610/.

Burkholder:1989:DSH

[Bur89] Donald L. Burkholder. Differential subordination of harmonic functions and martingales. Lecture Notes in Mathematics, 1384: 1–23, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086792/.

Buser:1988:CGP

[Bus88a] Peter Buser. Cayley graphs and planar isospectral domains. Lecture Notes in Mathematics, 1339:64–77, 1988. CODEN LN-MAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083047/.

Busque:1988:DFC

[Bus88b] Claudi Busqué. Directly finite ℵ_{0}-complete regular rings are unitregular. Lecture Notes in Mathematics, 1328:38–49, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100913/.

Buttazzo:1988:RPC

[But88] Giuseppe Buttazzo. Relaxation problems in control theory. Lecture Notes in Mathematics, 1340:31–39, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082883/.

Buzasi:1988:RIG

[Buz88] K. Buzási. On representations of infinite groups. Lecture Notes in Mathematics, 1352:44–59, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082016/.

Bruns:1988:ASL

[BV88a] Winfried Bruns and Udo Vetter. Algebras with straightening law on posets of minors. Lecture Notes in Mathematics, 1327: 38–49, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080382/.

Bruns:1988:A

[BV88b] Winfried Bruns and Udo Vetter. Appendix. Lecture Notes in Mathematics, 1327:202–218, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080394/.

Bruns:1988:DR

[BV88c] Winfried Bruns and Udo Vetter. Derivations and rigidity. Lecture Notes in Mathematics, 1327:184–201, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080393/.

Bruns:1988:DCG

[BV88d] Winfried Bruns and Udo Vetter. The divisor class group and the canonical class. Lecture Notes in Mathematics, 1327:93–104, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080386/.

Bruns:1988:GM

[BV88e] Winfried Bruns and Udo Vetter. Generic modules. Lecture Notes in Mathematics, 1327:162–173, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080391/.

Bruns:1988:GPIb

[BV88f] Winfried Bruns and Udo Vetter. Generic points and invariant theory. Lecture Notes in Mathematics, 1327:73–92, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080385/.

Bruns:1988:GPIa

[BV88g] Winfried Bruns and Udo Vetter. Generically perfect ideals. Lecture Notes in Mathematics, 1327:27–37, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080381/.

Bruns:1988:IMM

[BV88h] Winfried Bruns and Udo Vetter. Ideals of maximal minors. Lecture Notes in Mathematics, 1327:10–26, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080380/.

Bruns:1988:INS

[BV88i] Winfried Bruns and Udo Vetter. Integrity and normality. The singular locus. Lecture Notes in Mathematics, 1327:64–72, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080384/.

Bruns:1988:MKD

[BV88j] Winfried Bruns and Udo Vetter. The module of Kähler differentials. Lecture Notes in Mathematics, 1327:174–183, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080392/.

Bruns:1988:PIM

[BV88k] Winfried Bruns and Udo Vetter. Powers of ideals of maximal minors. Lecture Notes in Mathematics, 1327:105–121, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080387/.

Bruns:1988:P

[BV88l] Winfried Bruns and Udo Vetter. Preliminaries. Lecture Notes in Mathematics, 1327:1–9, 1988. CODEN LNMAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080379/.

Bruns:1988:PD

[BV88m] Winfried Bruns and Udo Vetter. Primary decomposition. Lecture Notes in Mathematics, 1327:122–134, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080388/.

Bruns:1988:PRS

[BV88n] Winfried Bruns and Udo Vetter. Principal radical systems. Lecture Notes in Mathematics, 1327:153–161, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080390/.

Bruns:1988:RT

[BV88o] Winfried Bruns and Udo Vetter. Representation theory. Lecture Notes in Mathematics, 1327:135–152, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080389/.

Bruns:1988:SA

[BV88p] Winfried Bruns and Udo Vetter. The structure of an ASL. Lecture Notes in Mathematics, 1327:50–63, 1988. CODEN LN-MAA2. ISBN 3-540-19468-1 (print), 3-540-39274-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080383/.

Bueken:1989:ACM

[BV89] P. Bueken and L. Vanhecke. Algebraic characterizations by means of the curvature in contact geometry. Lecture Notes in Mathematics, 1410:77–86, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086413/.

Barnett:1985:DMD

[BW85a] C. Barnett and I. F. Wilde. The Doob–Meyer decomposition for the square of Itô–Clifford L^{2}-martingales. Lecture Notes in Mathematics, 1136:67–73, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074460/.

Beagles:1985:TRE

[BW85b] A. E. Beagles and J. R. Whiteman. Treatment of a re-entrant vertex in a three-dimensional Poisson problem. Lecture Notes in Mathematics, 1121:19–27, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076259/.

Birman:1985:FBC

[BW85c] Joan S. Birman and Bronisław Wajnryb. 3-fold branched coverings and the mapping class group of a surface. Lecture Notes in Mathematics, 1167:24–46, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075214/.

Billeke:1988:BCO

[BW88] J. Billeke and M. Wallace. Bifurcationss of codimension one singularities of tangent vector fields on Whitney’s umbrella. Lecture Notes in Mathematics, 1331:1–11, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083062/.

Bolton:1985:ISS

[BWW85] J. Bolton, T. J. Willmore, and L. M. Woodward. Immersions of surfaces into space forms. Lecture Notes in Mathematics, 1156: 46–58, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075085/.

Barros:1986:FTS

[ByC86] Manuel Barros and Bang yen Chen. Finite type spherical submanifolds. Lecture Notes in Mathematics, 1209:73–93, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076621/.

Caenepeel:1986:GVA

[Cae86] S. Caenepeel. A graded version of Artin’s refinement theorem. Lecture Notes in Mathematics, 1197:31–44, 1986. CODEN LN-MAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076310/.

Caenepeel:1988:CTP

[Cae88] S. Caenepeel. Cancellation theorems for projective graded modules. Lecture Notes in Mathematics, 1328:50–59, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100914/.

Caffarelli:1989:FBP

[Caf89] Luis A. Caffarelli. Free boundary problems a survey. Lecture Notes in Mathematics, 1365:31–61, 1989. CODEN LNMAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089177/.

Caicedo:1985:FIQ

[Cai85] Xavier Caicedo. Failure of interpolation for quantifiers of monadic type. Lecture Notes in Mathematics, 1130:1–12, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075304/.

Cambini:1986:NLS

[Cam86] A. Cambini. Non-linear separation theorems, duality and optimality conditions. Lecture Notes in Mathematics, 1190:57–93, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076702/.

Camacho:1988:QHV

[Cam88] M. I. Camacho. Quasi-homogeneous vector fields of degree 2 in R^{3}. Lecture Notes in Mathematics, 1331:38–46, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083064/.

Cantrell:1985:LFE

[Can85] J. C. Cantrell. Locally flat embeddings of three dimensional manifolds in four dimensional manifolds. Lecture Notes in Mathematics, 1167:47–49, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075215/.

Cano:1988:DSF

[Can88] F. Cano. Dicriticalness of a singular foliation. Lecture Notes in Mathematics, 1345:73–94, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081396/.

Cardassi:1985:NRA

[Car85a] Carmen Silvia Cardassi. Numerical radius attaining operators. Lecture Notes in Mathematics, 1166:11–14, 1985. CODEN LN-MAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074687/.

Carlson:1985:VM

[Car85b] Jon F. Carlson. The variety of a module. Lecture Notes in Mathematics, 1142:88–95, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074793/.

Carpenter:1985:NPI

[Car85c] Gail A. Carpenter. The non-parametric influence of light on mammalian circadian rhythms. Lecture Notes in Mathematics, 1151: 90–108, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074718/.

Carlen:1986:ESP

[Car86a] Eric A. Carlen. Existence and sample path properties of the diffusions in Nelson’s stochastic mechanics. Lecture Notes in Mathematics, 1158:25–51, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080208/.

Carlsson:1986:FAP

[Car86b] Gunnar Carlsson. Free (Z/2)^{k}-actions and a problem in commutative algebra. Lecture Notes in Mathematics, 1217:79–83, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072815/.

Carmona:1986:RSO

[Car86c] René Carmona. Random Schrödinger operators. Lecture Notes in Mathematics, 1180:1–124, 1986. CODEN LNMAA2. ISBN 3-540-16441-3 (print), 3-540-39781-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074918/.

Carnal:1986:VAL

[Car86d] H. Carnal. Les variables aléatoires de loi stable et leur représentation selon P. Levy. (French) []. Lecture Notes in Mathematics, 1210:24–33, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077169/.

Carroll:1986:STM

[Car86e] Robert Carroll. Some transmutation methods for canonical systems. Lecture Notes in Mathematics, 1223:25–36, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099180/.

Carverhill:1986:SLE

[Car86f] A. Carverhill. Survey: Lyapunov exponents for stochastic flows on manifolds. Lecture Notes in Mathematics, 1186:292–307, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076849/.

Carlen:1988:SMF

[Car88a] Eric A. Carlen. Stochastic mechanics of free scalar fields. Lecture Notes in Mathematics, 1325:40–60, 1988. CODEN LN-MAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077915/.

Carter:1988:SEI

[Car88b] J. Scott Carter. Surgering the equatorial immersion in law dimensions. Lecture Notes in Mathematics, 1350:144–166, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081473/.

Carbonne:1989:GCD

[Car89a] P. Carbonne. Groupe des classes de diviseurs des algèbres graduées normales. (French) []. Lecture Notes in Mathematics, 1404:174– 200, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084076/.

Carroll:1989:SRF

[Car89b] Robert Carroll. Some remarks on forced integrable systems. Lecture Notes in Mathematics, 1394:11–17, 1989. CODEN LN-MAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086748/.

Carter:1989:CTC

[Car89c] Brandon Carter. Covariant theory of conductivity in ideal fluid or solid media. Lecture Notes in Mathematics, 1385:1–64, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084028/.

Casati:1985:ELD

[Cas85a] Giulio Casati. Energy level distributions and chaos in quantum mechanics. Lecture Notes in Mathematics, 1139:86–107, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074578/.

Casati:1985:LCM

[Cas85b] Giulio Casati. Limitations for chaotic motion in quantum mechanics. Lecture Notes in Mathematics, 1136:126–135, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074466/.

Cash:1986:BER

[Cas86] J. R. Cash. A block 5(4) explicit Runge–Kutta formula with ‘free’ interpolation. Lecture Notes in Mathematics, 1230:208–220, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072683/.

Catanese:1985:GIS

[Cat85a] Frabrizio Catanese. Generic invertible sheaves of 2-torsion and generic invertible thetacharacteristics on nodal plane curves. Lecture Notes in Mathematics, 1124:58–70, 1985. CODEN LN-MAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074995/.

Cathala:1985:BBC

[Cat85b] J.-C. Cathala. On the bifurcation between a chaotic area of T^{K} and a chaotic area of T. Lecture Notes in Mathematics, 1163: 14–22, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076413/.

Catanese:1988:MAS

[Cat88] F. Catanese. Moduli of algebraic surfaces. Lecture Notes in Mathematics, 1337:1–83, 1988. CODEN LNMAA2. ISBN 3-540-50080-4 (print), 3-540-45920-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082806/.

Cauchon:1986:CMT

[Cau86a] Gérard Cauchon. Commutants des modules de type Dini sur les algèbres noetheriennes. (French) []. Lecture Notes in Mathematics, 1220:109–117, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099508/.

Cauchon:1986:ESC

[Cau86b] Gérard Cauchon. Exemples de sous-corps commutatifs maximaux dans D_{2}(k). (French) []. Lecture Notes in Mathematics, 1197:45– 50, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076311/.

Cauchon:1988:CAP

[Cau88] G. Cauchon. Centraliseurs dans les anneaux de polynômes différentiels formels et leurs corps de fractions. (French) []. Lecture Notes in Mathematics, 1328:60–68, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100915/.

Cavazzuti:1986:CET

[Cav86] E. Cavazzuti. Convergence of equilibria in the theory of games. Lecture Notes in Mathematics, 1190:95–130, 1986. CODEN LN-MAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076703/.

Casas-Alvero:1985:AGS

[CAWXD85] Eduard Casas-Alvero, Gerald Welters, and Sebastian Xambó-Descamps, editors. Algebraic Geometry Sitges (Barcelona) 1983: Proceedings of a Conference held in Sitges (Barcelona), Spain October 5–12, 1983, volume 1124 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1124. URL http:

//link.springer.com/book/10.1007/BFb0074992;   http://
www.springerlink.com/content/978-3-540-39643-7.

Casas-Alvero:1986:APG

[CAXD86a] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Action of the projective group and proper solutions of an enumerative problem. Lecture Notes in Mathematics, 1196:18–22, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076507/.

Casas-Alvero:1986:CC

[CAXD86b] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Complete conics. Lecture Notes in Mathematics, 1196:3–8, 1986. CO-DEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076503/.

Casas-Alvero:1986:CCN

[CAXD86c] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Computation of characteristic numbers. Lecture Notes in Mathematics, 1196:31–35, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076511/.

Casas-Alvero:1986:CIC

[CAXD86d] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Conditions imposed on conics. Lecture Notes in Mathematics, 1196: 14–18, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076506/.

Casas-Alvero:1986:Ca

[CAXD86e] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Conics. Lecture Notes in Mathematics, 1196:1–2, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076502.

Casas-Alvero:1986:CSF

[CAXD86f] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Conics satisfying five independent conditions. Lecture Notes in Mathematics, 1196:48–56, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076515/.

Casas-Alvero:1986:Cb

[CAXD86g] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Cycles on W and on B. Lecture Notes in Mathematics, 1196:56–61, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076516/.

Casas-Alvero:1986:ETC

[CAXD86h] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. The Enumerative Theory of Conics after Halphen, volume 1196 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 129 pp. LCCN QA3

.L28 no. 1196. URL http://link.springer.com/book/10.1007/
BFb0076501;   http://www.springerlink.com/content/978-
3-540-39832-5.

Casas-Alvero:1986:EDH

[CAXD86i] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Equations of the degeneration hypersurfaces. Lecture Notes in Mathematics, 1196:11–14, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076505/.

Casas-Alvero:1986:E

[CAXD86j] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Examples. Lecture Notes in Mathematics, 1196:35–41, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076512/.

Casas-Alvero:1986:EHO

[CAXD86k] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Examples of higher order conditions. Lecture Notes in Mathematics, 1196:80–91, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076520/.

Casas-Alvero:1986:GHF

[CAXD86l] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Generalization of Halphen’s formula to conditions of higher order. Lecture Notes in Mathematics, 1196:74–80, 1986. CODEN LN-MAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076519/.

Casas-Alvero:1986:GDC

[CAXD86m] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Genericity of the data of conditions. Lecture Notes in Mathematics, 1196: 22–23, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076508.

Casas-Alvero:1986:HFF

[CAXD86n] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Halphen’s first formula. Lecture Notes in Mathematics, 1196:25–30, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076510/.

Casas-Alvero:1986:IIN

[CAXD86o] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Improper intersection numbers. Lecture Notes in Mathematics, 1196:67– 73, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076518/.

Casas-Alvero:1986:LCC

[CAXD86p] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Local characters of a condition. Lecture Notes in Mathematics, 1196: 23–24, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076509.

Casas-Alvero:1986:MCN

[CAXD86q] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Multiplicity cycles and Noether’s formula. Lecture Notes in Mathematics, 1196:61–67, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076517/.

Casas-Alvero:1986:SECa

[CAXD86r] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Strict equivalence of conditions. Lecture Notes in Mathematics, 1196: 41–45, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076513/.

Casas-Alvero:1986:SECb

[CAXD86s] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Strict equivalence of conditions. Lecture Notes in Mathematics, 1196: 92–99, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076521/.

Casas-Alvero:1986:SC

[CAXD86t] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Systems of conics. Lecture Notes in Mathematics, 1196:8–11, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076504/.

Casas-Alvero:1986:SCC

[CAXD86u] Eduardo Casas-Alvero and Sebastian Xambó-Descamps. Systems of conics contained in S_{p,q}. Lecture Notes in Mathematics, 1196: 45–47, 1986. CODEN LNMAA2. ISBN 3-540-16495-2 (print), 3-540-39832-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076514/.

Cazacu:1988:CKC

[Caz88] Cabiria Andreian Cazacu. Complete Klein coverings of Klein surfaces of characteristic 0. I. Lecture Notes in Mathematics, 1351: 12–22, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081238/.

Chuaqui:1985:ATT

[CB85] Rolando Chuaqui and Leopoldo Bertossi. Approximation to truth and theory of errors. Lecture Notes in Mathematics, 1130:13– 31, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075305/.

Choquet-Bruhat:1987:MCS

[CB87] Yvonne Choquet-Bruhat. Mathematics for classical supergravities. Lecture Notes in Mathematics, 1251:73–90, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077316/.

Curras-Bosch:1989:COF

[CB89] Carlos Currás-Bosch. On codimension-one foliations. Lecture Notes in Mathematics, 1410:100–107, 1989. CODEN LN-MAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086416/.

Combes:1987:SPS

[CBD87] J.-M. Combes, Ph. Briet, and P. Duclos. Spectral properties of Schrödinger operators with trapping potentials in the semiclassical limit. Lecture Notes in Mathematics, 1285:55–72, 1987.

CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080581/.

Chudnovsky:1985:APAa

[CC85a] D. V. Chudnovsky and G. V. Chudnovsky. Applications of Padé approximations to Diophantine inequalities in values of Gfunctions. Lecture Notes in Mathematics, 1135:9–51, 1985. CO-DEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074600/.

Chudnovsky:1985:APAb

[CC85b] D. V. Chudnovsky and G. V. Chudnovsky. Applications of Padé approximations to the Grothendieck conjecture on linear differential equations. Lecture Notes in Mathematics, 1135:52–100, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074601/.

Chan:1987:UBG

[CC87a] Whei-Ching C. Chan and Shui-Nee Chow. Uniform boundness and generalized inverses in Liapunov–Schmidt method for subharmonics. Lecture Notes in Mathematics, 1248:28–39, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077413/.

Chudnovsky:1987:CAN

[CC87b] D. V. Chudnovsky and G. V. Chudnovsky. Computer assisted number theory with applications. Lecture Notes in Mathematics, 1240:1–68, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072972/.

Chudnovsky:1988:EFG

[CC88] D. V. Chudnovsky and G. V. Chudnovsky. Elliptic formal groups over Z and F_{p} in applications to number theory, computer science and topology. Lecture Notes in Mathematics, 1326:11–54, 1988. CODEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078037/.

Cecil:1989:DSL

[CC89a] Thomas E. Cecil and Shiing-Shen Chern. Dupin submanifolds in Lie sphere geometry. Lecture Notes in Mathematics, 1369:1– 48, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087525/.

Chudnovsky:1989:CPA

[CC89b] D. V. Chudnovsky and G. V. Chudnovsky. Computational problems in arithmetic of linear differential equations. Some Diophantine applications. Lecture Notes in Mathematics, 1383:12–49, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083567/.

Chudnovsky:1985:NTS

[CCCN85] David V. Chudnovsky, Gregory V. Chudnovsky, Harvey Cohn, and Melvyn B. Nathanson, editors. Number Theory: A Seminar held at the Graduate School and University Center of the City University of New York 1983–84, volume 1135 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1135. URL http:

//link.springer.com/book/10.1007/BFb0074598;   http://
www.springerlink.com/content/978-3-540-39535-5.

Chudnovsky:1987:NTS

[CCCN87] David V. Chudnovsky, Gregory V. Chudnovsky, Harvey Cohn, and Melvyn B. Nathanson, editors. Number Theory: A Seminar held at the Graduate School and University Center of the City University of New York 1984–85, volume 1240 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1240. URL http:

//link.springer.com/book/10.1007/BFb0072971;   http://
www.springerlink.com/content/978-3-540-47756-3.

Carverhill:1986:CES

[CCE86] A. P. Carverhill, M. J. Chappell, and K. D. Elworthy. Characteristic exponents for stochastic flows. Lecture Notes in Mathematics, 1158:52–80, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080209/.

Charzynski:1985:CKJ

[CCS85] Zygmunt Charzyński, Jacek Chadzyński, and Przemysław Skibiński. A contribution to Keller’s Jacobian conjecture. Lecture Notes in Mathematics, 1165:36–51, 1985. CODEN LN-MAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076145/.

Chatzidakis:1987:OTS

[CCS^{+}87] Z. Chatzidakis, G. Cherlin, S. Shelah, G. Srour, and C. Wood. Orthogonality of types in separably closed fields. Lecture Notes in Mathematics, 1292:72–88, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082232/.

Cosnard:1985:DA

[CD85] M. Cosnard and J. Demongeot. On the definitions of attractors. Lecture Notes in Mathematics, 1163:23–31, 1985. CODEN LN-MAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076414/.

Cordes:1987:LCA

[CD87] H. O. Cordes and S. H. Doong. The Laplace comparison algebra of spaces with conical and cylindrical ends. Lecture Notes in Mathematics, 1256:55–90, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077738/.

Conti:1986:ORF

[CDG86] Roberto Conti, Ennio De Giorgi, and Franco Giannessi, editors. Optimization and Related Fields: Proceedings of the “G. Stampacchia International School of Mathematics” held at Erice, Sicily September 17–30, 1984, volume 1190 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1190. URL http:

//link.springer.com/book/10.1007/BFb0076699;   http://
www.springerlink.com/content/978-3-540-39817-2.

Coifman:1989:CZO

[CDMS89] R. Coifman, G. David, Y. Meyer, and S. Semmes. θ-Calderón– Zygmund operators. Lecture Notes in Mathematics, 1384:132– 145, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086798/.

Carmona:1987:NCH

[CDV87] Jacques Carmona, Patrick Delorme, and Michèle Vergne, editors. Non-Commutative Harmonic Analysis and Lie Groups: Proceedings of the International Conference held in Marseille–Luminy, June 24–29, 1985, volume 1243 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1243. URL http:

//link.springer.com/book/10.1007/BFb0073014;   http://
www.springerlink.com/content/978-3-540-47775-4.

Chappell:1988:FND

[CE88] M. J. Chappell and K. D. Elworthy. Flows of Newtonian diffusions. Lecture Notes in Mathematics, 1325:61–75, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077916/.

Cecchini:1985:NCS

[Cec85] Carlo Cecchini. Non commutative L^{p} spaces and K.M.S. functions. Lecture Notes in Mathematics, 1136:136–142, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074467/.

Cecconi:1986:PP

[Cec86] J. P. Cecconi. On Pareto’s problems. Lecture Notes in Mathematics, 1190:131–161, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076704/.

Cecchini:1988:SNR

[Cec88] Carlo Cecchini. Some noncommutative Radon–Nikodym theorems for von Neumann algebras. Lecture Notes in Mathematics, 1303:52–68, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078054/.

Cecchini:1989:SCN

[Cec89] Carlo Cecchini. Stochastic couplings for von Neumann algebras. Lecture Notes in Mathematics, 1396:128–141, 1989. CODEN LN-MAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083549/.

Cedo:1988:RGA

[Ced88] Ferran Cedó. Regular group algebras whose maximal right and left quotient rings coincide. Lecture Notes in Mathematics, 1328: 69–74, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100916/.

Cegrell:1987:SFM

[Ceg87] Urban Cegrell. Subharmonic functions and minimal surfaces. Lecture Notes in Mathematics, 1276:60–66, 1987. CODEN LN-MAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078955/.

Cegla:1989:CPC

[Ceg89] Wojciech Cegła. The covering property in a causal logic. Lecture Notes in Mathematics, 1396:142–148, 1989. CODEN LN-MAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083550/.

Cobos:1988:HSS

[CF88] Fernando Cobos and Dicesar Lass Fernandez. Hardy–Sobolev spaces and Besov spaces with a function parameter. Lecture Notes in Mathematics, 1302:158–170, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078872/.

Caranti:1987:SEI

[CFdG87] A. Caranti, S. Franciosi, and F. de Giovanni. Some examples of infinite groups in which each element commutes with its endomorphic images. Lecture Notes in Mathematics, 1281:9–17, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078685/.

Carreras:1986:HND

[CFM86] F. J. Carreras, A. Ferrández, and V. Miquel. Hermitian natural differential operators. Lecture Notes in Mathematics, 1209: 109–132, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076624/.

Chinea:1986:EAC

[CG86a] D. Chinea and C. González. An example of an almost cosymplectic homogeneous manifold. Lecture Notes in Mathematics, 1209:133–142, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076625/.

Cox:1986:CCT

[CG86b] J. Theodore Cox and David Griffeath. Critical clustering in the two dimensional voter model. Lecture Notes in Mathematics, 1212: 59–68, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076238/.

Chobanyan:1989:PRS

[CG89] S. A. Chobanyan and G. J. Georgobiani. A problem on rearrangements of summands in normed spaces and Rademacher sums. Lecture Notes in Mathematics, 1391:33–46, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083378/.

Cegarra:1988:ESF

[CGC88] A. M. Cegarra, A. R. Garzón, and P. Carrasco. An exact sequence in the first variable for non-abelian cohomology in algebraic categories. A Mayer–Vietoris sequence for non-abelian cohomology of groups. Lecture Notes in Mathematics, 1348:72–86, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081349/.

Curtis:1987:CUA

[CGMM87] Edward B. Curtis, Paul Goerss, Mark Mahowald, and R. James Milgram. Calculations of unstable Adams E_{2} terms for spheres. Lecture Notes in Mathematics, 1286:208–266, 1987. CODEN LN-MAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078744/.

Cordes:1987:PDO

[CGW87] Heinz O. Cordes, Bernhard Gramsch, and Harold Widom, editors. Pseudo-Differential Operators: Proceedings of a Conference held in Oberwolfach, February 2–8, 1986, volume 1256 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LN-MAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1256. URL http://link.springer.com/book/10.1007/
BFb0077734;   http://www.springerlink.com/content/978-
3-540-47886-7.

Chern:1985:RMA

[CH85] S. S. Chern and R. S. Hamilton. On Riemannian metrics adapted to three-dimensional contact manifolds. Lecture Notes in Mathematics, 1111:279–308, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084596/.

Hsiang:1989:PPD

[cH89a] Wu chung Hsiang. A personal perspective of differentiable transformation groups. Lecture Notes in Mathematics, 1375:1–5, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085594/.

Coutinho:1989:DOS

[CH89b] S. C. Coutinho and M. P. Holland. Differential operators on smooth varieties. Lecture Notes in Mathematics, 1404:201–219, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084077/.

Chalmers:1985:APC

[Cha85a] Bruce L. Chalmers. Absolute projection constants via absolute minimal projections. Lecture Notes in Mathematics, 1166:15– 21, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074688/.

Chaperon:1985:EPC

[Cha85b] Marc Chaperon. An elementary proof of the Conley–Zehnder theorem in symplectic geometry. Lecture Notes in Mathematics, 1125:1–8, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075631/.

Chan:1986:EMA

[Cha86a] Tony F. Chan. An efficient modular algorithm for coupled nonlinear systems. Lecture Notes in Mathematics, 1230:73–85, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072672/.

Chang:1986:MPL

[Cha86b] Kung-Ching Chang. On the Mountain Pass Lemma. Lecture Notes in Mathematics, 1192:203–208, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076070/.

Chatterji:1986:MTA

[Cha86c] S. D. Chatterji. Martingale theory: an analytical formulation with some applications in analysis. Lecture Notes in Mathematics, 1206:109–166, 1986. CODEN LNMAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076301/.

Chaperon:1988:IMP

[Cha88] Marc Chaperon. Invariant manifolds and a preparation lemma for local holomorphic flows and actions. Lecture Notes in Mathematics, 1345:95–110, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081397/.

Cheney:1985:FLA

[Che85] E. W. Cheney. Five lectures on the algorithmic aspects of approximation theory. Lecture Notes in Mathematics, 1129:1–20, 1985. CODEN LNMAA2. ISBN 3-540-15234-2 (print), 3-540-39412-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075156/.

Cheeger:1986:VTP

[Che86a] Jeff Cheeger. A vanishing theorem for piecewise constant curvature spaces. Lecture Notes in Mathematics, 1201:33–40, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075646/.

Chen:1986:DS

[Che86b] Kuo-Tsai Chen. On differentiable spaces. Lecture Notes in Mathematics, 1174:38–42, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076932/.

Chen:1986:HLG

[Che86c] Yu Chen. Homomorphisms from linear groups over division rings to algebraic groups. Lecture Notes in Mathematics, 1185:231– 265, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076177/.

Chen:1986:PVS

[Che86d] Z. Chen. A prehomogeneous vector space of characteristic 3. Lecture Notes in Mathematics, 1185:266–276, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076178/.

Chen:1987:ELB

[Che87] Zhihua Chen. An estimate of the lower bound of Levi form and its applications. Lecture Notes in Mathematics, 1255:13–25, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077677/.

Chen:1988:SSF

[Che88a] Shuxing Chen. Smoothness of shock front solutions for systems of conservation laws. Lecture Notes in Mathematics, 1306:38– 60, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082923/.

Chen:1988:DMA

[Che88b] Yazhe Chen. On degenerate Monge–Ampère equations in convex domains. Lecture Notes in Mathematics, 1306:61–68, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082924/.

Chernjakov:1988:OEF

[Che88c] A. G. Chernjakov. Obstructions to the existence of free nonholonomic structure on a smooth manifold. Lecture Notes in Mathematics, 1346:139–156, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082776/.

Chen:1989:MCT

[Che89] Weihuan Chen. The mean curvatures on the tubular hypersurfaces in a space of constant curvature. Lecture Notes in Mathematics, 1369:49–62, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087526/.

Chiappinelli:1985:BNP

[Chi85] Raffaele Chiappinelli. On bifurcation for nondifferentiable perturbations of selfadjoint operators. Lecture Notes in Mathematics, 1151:109–114, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074719/.

Chaillou:1989:BNF

[CHM89] Paul Chaillou, Pierre Hansen, and Yvon Mahieu. Best network flow bounds for the quadratic knapsack problem. Lecture Notes in Mathematics, 1403:225–235, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083467/.

Choda:1985:SEF

[Cho85a] Marie Choda. Strong ergodicity and full II_{1}-factors. Lecture Notes in Mathematics, 1132:84–90, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074880/.

Chong:1985:RTS

[Cho85b] C. T. Chong. Recursion theory on strongly Σ_{2} inadmissible ordinals. Lecture Notes in Mathematics, 1141:49–64, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076214/.

Chollet:1987:PRS

[Cho87a] Anne-Marie Chollet. Propriétés de recouvrement des sousensembles de la frontière d’un domaine strictement pseudoconvexe. (French) []. Lecture Notes in Mathematics, 1276:67– 78, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078956/.

Chow:1987:EFA

[Cho87b] P. L. Chow. Expectation functionals associated with some stochastic evolution equations. Lecture Notes in Mathematics, 1236:40– 56, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072882/.

Cho:1989:GDF

[Cho89a] Jung Rae Cho. On n-groupoids defined on fields. Lecture Notes in Mathematics, 1398:73–81, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086241/.

Chong:1989:RES

[Cho89b] C. T. Chong. Recursively enumerable sets in models of Σ_{2} collection. Lecture Notes in Mathematics, 1388:1–15, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083664/.

Choquet:1989:LMB

[Cho89c] Gustave Choquet. L’oeuvre de Marcel Brelot (1903–1987). (French) []. Lecture Notes in Mathematics, 1393:1–23, 1989. CO-DEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085768/.

Christophersen:1987:MCO

[Chr87] Jan Arthur Christophersen. Monomial curves and obstructions on cyclic quotient singularities. Lecture Notes in Mathematics, 1273:117–133, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078839/.

Christ:1989:ESP

[Chr89] Mike Christ. On the ¯∂_{b} equation and Szegő projection on CR manifolds. Lecture Notes in Mathematics, 1384:146–158, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086799/.

Connes:1985:DSS

[CHS85] Alain Connes, Uffe Haagerup, and Erling Størmer. Diameters of state spaces of type III factors. Lecture Notes in Mathematics, 1132:91–116, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074881/.

Chu:1988:NEP

[Chu88a] Hsin Chu. A note on the existence of periodic solutions of a differential system. Lecture Notes in Mathematics, 1342:94–99, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082824/.

Chung:1988:PMP

[Chu88b] Kai-Lai Chung. Probability methods in potential theory. Lecture Notes in Mathematics, 1344:42–54, 1988. CODEN LN-MAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103343/.

Cibils:1986:HHA

[Cib86] Claude Cibils. Hochschild homology of an algebra whose quiver has no oriented cycles. Lecture Notes in Mathematics, 1177:55– 59, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075258/.

Ciesielski:1988:NPD

[Cie88] Z. Ciesielski. Nonparametric polynomial density estimation in the L^{P} norm. Lecture Notes in Mathematics, 1354:1–10, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089579/.

Ciliberto:1987:HFF

[Cil87] Ciro Ciliberto. Hilbert functions of finite sets of points and the genus of a curve in a projective space. Lecture Notes in Mathematics, 1266:24–73, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078177/.

Cushman:1985:EMM

[CK85] R. Cushman and H. Knörrer. The energy momentum mapping of the Lagrange top. Lecture Notes in Mathematics, 1139:12– 24, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074575/.

Culik:1988:SEF

[CK88] K. Culik II and J. Karhumäki. Systems of equations over a finitely generated free monoid having an effectively findable equivalent finite subsystem. Lecture Notes in Mathematics, 1320:18–27, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083421/.

Christodoulou:1989:NSM

[CK89] D. Christodoulou and S. Klainerman. The nonlinear stability of the Minkowski metric in general relativity. Lecture Notes in Mathematics, 1402:128–145, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083872/.

Cattani:1987:HTP

[CKGP87] Eduardo Cattani, Aroldo Kaplan, Francisco Guillén, and Fernando Puerta, editors. Hodge Theory: Proceedings of the U.S.– Spain Workshop held in Sant Cugat (Barcelona), Spain June 24–30, 1985, volume 1246 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1246. URL http:

//link.springer.com/book/10.1007/BFb0077524;   http://
www.springerlink.com/content/978-3-540-47794-5.

Cattani:1987:SRI

[CKS87a] Eduardo Cattani, Aroldo Kaplan, and Wilfried Schmid. Some remarks on L^{2} and intersection cohomologies. Lecture Notes in Mathematics, 1246:32–41, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077527/.

Cattani:1987:VPH

[CKS87b] Eduardo Cattani, Aroldo Kaplan, and Wilfried Schmid. Variations of polarized Hodge structure: Asymptotics and monodromy. Lecture Notes in Mathematics, 1246:16–31, 1987. CODEN LN-MAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077526/.

Contessa:1985:AAS

[CL85] Maria Contessa and Léonce Lesieur. D-anneaux et anneaux F - semi-parfaits. (French) []. Lecture Notes in Mathematics, 1146: 367–404, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074548/.

Colombeau:1987:NTE

[CL87] J.-F. Colombeau and A. Y. Le Roux. Numerical techniques in elastoplasticity. Lecture Notes in Mathematics, 1270:103– 114, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078320/.

Chen:1989:INE

[CL89a] H. H. Chen and J. E. Lin. On the integrability of nonlinear evolution equations. Lecture Notes in Mathematics, 1394:30–43, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086750/.

Cranston:1989:SBH

[CL89b] M. Cranston and Y. Le Jan. Simultaneous boundary hitting for a two point reflecting Brownian motion. Lecture Notes in Mathematics, 1372:234–238, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083975/.

Clarke:1986:TRT

[Cla86] F. H. Clarke. Tonelli’s regurarity theory in the calculus of variations: Recent progress. Lecture Notes in Mathematics, 1190: 163–179, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076705/.

Clerc:1987:BRM

[Cle87] Jean-Louis Clerc. Bochner–Riesz means of H^{P} functions (0 < p < 1) on compact Lie groups. Lecture Notes in Mathematics, 1243:86–107, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073019/.

Clerc:1988:FBP

[Cle88] J.-L. Clerc. Fonctions K de Bessel pour les algèbres de Jordan. (French) []. Lecture Notes in Mathematics, 1359:122–134, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086593/.

Clote:1985:ALB

[Clo85a] P. Clote. Applications of the low-basis theorem in arithmetic. Lecture Notes in Mathematics, 1141:65–88, 1985. CODEN LN-MAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076215/.

Clote:1985:PRA

[Clo85b] P. Clote. Partition relations in arithmetic. Lecture Notes in Mathematics, 1130:32–68, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075306/.

Casagrande:1985:AET

[CLS85] F. Casagrande, L. A. Lugiato, and G. Strini. Adiabatic elimination technique for quantum dissipative systems. Lecture Notes in Mathematics, 1136:112–125, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074465/.

Chaleyat-Maurel:1987:CNF

[CM87a] Mireille Chaleyat-Maurel. Continuity in nonlinear filtering: some different approaches. Lecture Notes in Mathematics, 1236:25– 39, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072881/.

Chattot:1987:BSE

[CM87b] Jean-Jacques Chattot and Sylvie Malet. A “box-scheme” for the Euler equations. Lecture Notes in Mathematics, 1270:82– 102, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078319/.

Cima:1987:CTM

[CM87c] Professor Joseph A. Cima and Professor Thomas H. MacGregor. Cauchy Transforms of measures and univalent functions. Lecture Notes in Mathematics, 1275:78–88, 1987. CODEN LN-MAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078345/.

Cohen:1987:S

[CM87d] Fred Cohen and Mark Mahowald. The splitting of ω^{2}S^{2n+1}. Lecture Notes in Mathematics, 1286:188–192, 1987. CODEN LN-MAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078742/.

Cachafeiro:1988:ARL

[CM88a] M. A. Cachafeiro and F. Marcellán. Asymptotics for the ratio of the leading coefficients of orthogonal polynomials associated with a jump modification. Lecture Notes in Mathematics, 1354:111–117, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089587/.

Cachafeiro:1988:OPJ

[CM88b] M. A. Cachafeiro and F. Marcellán. Orthogonal polynomials and jump modifications. Lecture Notes in Mathematics, 1329:236– 240, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083362/.

Coven:1988:BMD

[CM88c] Ethan M. Coven and Irene Mulvey. The Barge–Martin decomposition theorem for pointwise nonwandering maps of the interval. Lecture Notes in Mathematics, 1342:100–104, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082825/.

Cervera:1989:SRN

[CM89a] Vicente Cervera and Francisca Mascaró. Some results on the normal subgroups of Diff^{Ω}(X×R^{+}, relX×[0]). Lecture Notes in Mathematics, 1410:87–93, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086414/.

Chow:1989:VIC

[CM89b] P. L. Chow and J.-L. Menaldi. Variational inequalities for the control of stochastic partial differential equations. Lecture Notes in Mathematics, 1390:42–52, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083935/.

Chaleyat-Maurel:1985:GGF

[CMJ85] Mireille Chaleyat-Maurel and Thierry Jeulin. Grossissement gaussien de la filtration brownienne. (French) []. Lecture Notes in Mathematics, 1118:59–109, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075771/.

Chaleyat-Maurel:1989:SDF

[CMM89] Mireille Chaleyat-Maurel and Dominique Michel. The support of the density of a filter in the uncorrelated case. Lecture Notes in Mathematics, 1390:33–41, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083934/.

Cervera:1989:VEN

[CMS89] Vicente Cervera, Francisca Mascaró, and Rafael Sivera. On volume elements on a non-compact manifold. Lecture Notes in Mathematics, 1410:94–99, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086415/.

Coleman:1986:RRF

[CO86] Bernard D. Coleman and David R. Owen. Recent research on the foundations of thermodynamics. Lecture Notes in Mathematics, 1174:83–99, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076936/.

Chiantini:1989:PSA

[CO89] Luca Chiantini and Ferruccio Orecchia. Plane sections of arithmetically normal curves in P^{3}. Lecture Notes in Mathematics, 1389:32–42, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085922/.

Coen:1987:SPC

[Coe87] S. Coen. Some properties of the canonical mapping of a complex space into its spectrum. Lecture Notes in Mathematics, 1276: 79–107, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078957/.

Cohn:1985:KPI

[Coh85] Harvey Cohn. Klein’s paradox, the icosahedron, and ring class fields. Lecture Notes in Mathematics, 1135:101–111, 1985. CO-DEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074602/.

Cohen:1986:PLC

[Coh86] Arjeh M. Cohen. Point-line characterizations of buildings. Lecture Notes in Mathematics, 1181:191–206, 1986. CODEN LN-MAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075515/.

Cohen:1987:CSA

[Coh87a] F. R. Cohen. A course in some aspects of classical homotopy theory. Lecture Notes in Mathematics, 1286:1–92, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078738/.

Cohen:1987:MFL

[Coh87b] Ralph L. Cohen. A model for the free loop space of a suspension. Lecture Notes in Mathematics, 1286:193–207, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078743/.

Cohn:1987:SDP

[Coh87c] Harvey Cohn. Successive diagonal projections of Hilbert modular functions. Lecture Notes in Mathematics, 1240:69–86, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072973/.

Cohen:1988:FFN

[Coh88a] Robert Cohen. A filtering formula for a non-linear system having a continuous observation, and a discrete observation at random times. Lecture Notes in Mathematics, 1316:352–371, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081940/.

Cohn:1988:VRF

[Coh88b] P. M. Cohn. Valuations on Ree fields. Lecture Notes in Mathematics, 1352:75–87, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082018/.

Cohn:1989:ITV

[Coh89] Harvey Cohn. Iteration of two-valued modular equations. Lecture Notes in Mathematics, 1383:50–58, 1989. CODEN LN-MAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083568/.

Collatz:1986:EAM

[Col86] L. Collatz. Einige Anwendungen der Mehrdimensionalen Approximationstheorie zur Lösungseinschließung bei Randwertaufgaben. (German) [Some applications of multidimensional approximation theory for enclosure-solution of boundary value problems]. Lecture Notes in Mathematics, 1192:367–372, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076095/.

Collet:1988:HDS

[Col88a] P. Collet. Hausdorff dimension of the singularities for invariant measures of expanding dynamical systems. Lecture Notes in Mathematics, 1331:47–58, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083065/.

Colley:1988:MPF

[Col88b] Susan Jane Colley. Multiple-point formulas and line complexes. Lecture Notes in Mathematics, 1311:9–22, 1988. CODEN LN-MAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082906/.

Colonius:1988:I

[Col88c] Fritz Colonius. Introduction. Lecture Notes in Mathematics, 1313: 1–7, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077932/.

Colonius:1988:LRM

[Col88d] Fritz Colonius. Local relaxed minima. Lecture Notes in Mathematics, 1313:86–103, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077937/.

Colonius:1988:OPC

[Col88e] Fritz Colonius. Optimal periodic control of ordinary differential equations. Lecture Notes in Mathematics, 1313:145–166, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077940/.

Colonius:1988:OT

[Col88f] Fritz Colonius. Optimization theory. Lecture Notes in Mathematics, 1313:8–30, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077933/.

Colonius:1988:RFD

[Col88g] Fritz Colonius. Retarded functional differential equations. Lecture Notes in Mathematics, 1313:31–47, 1988. CODEN LN-MAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077934/.

Colonius:1988:SLP

[Col88h] Fritz Colonius. A scenario for local properness. Lecture Notes in Mathematics, 1313:129–144, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077939/.

Colonius:1988:SLM

[Col88i] Fritz Colonius. Strong local minima. Lecture Notes in Mathematics, 1313:48–64, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077935/.

Colonius:1988:TLP

[Col88j] Fritz Colonius. Tests for local properness. Lecture Notes in Mathematics, 1313:104–128, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077938/.

Colonius:1988:WLM

[Col88k] Fritz Colonius. Weak local minima. Lecture Notes in Mathematics, 1313:65–85, 1988. CODEN LNMAA2. ISBN 3-540-19249-2 (print), 3-540-39170-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077936/.

Comer:1985:UAL

[Com85] Stephen D. Comer, editor. Universal Algebra and Lattice Theory: Proceedings of a Conference held at Charleston, July 11–14, 1984, volume 1149 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1149. URL http:

//link.springer.com/book/10.1007/BFb0098450;   http://
www.springerlink.com/content/978-3-540-39638-3.

Connes:1985:INC

[Con85a] A. Connes. Introduction to non commutative differential geometry. Lecture Notes in Mathematics, 1111:3–16, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084582/.

Conte:1985:FTS

[Con85b] A. Conte. Fano threefolds with singularities. Lecture Notes in Mathematics, 1124:71–78, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074996/.

Conti:1986:SCA

[Con86] Benedetto Conti. Simply connected algebras of tree-class A_{n} and D_{n}. Lecture Notes in Mathematics, 1177:60–90, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075259/.

Conforti:1989:FPG

[Con89] Michele Conforti. (K_{4}-e)-free perfect graphs and star cutsets. Lecture Notes in Mathematics, 1403:236–253, 1989. CODEN LN-MAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083468/.

Cooper:1986:SFT

[Coo86a] Jeffery Cooper. Scattering frequencies for time — periodic scattering problems. Lecture Notes in Mathematics, 1223:37–48, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099181/.

Cooper:1986:TAS

[Coo86b] S. Clement Cooper. On M-tables associated with strong moment problems. Lecture Notes in Mathematics, 1199:21–36, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075932/.

Cooper:1989:FSR

[Coo89] S. Clement Cooper. δ-Fraction solutions to Riccati equations. Lecture Notes in Mathematics, 1406:1–17, 1989. CODEN LN-MAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096162/.

CorradaC:1985:ASC

[Cor85] Manuel Corrada C. On the axiomatizability of sets in a class theory. Lecture Notes in Mathematics, 1130:69–75, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075307/.

Clenshaw:1989:LIA

[COT89] C. W. Clenshaw, F. W. J. Olver, and P. R. Turner. Level-index arithmetic: an introductory survey. Lecture Notes in Mathematics, 1397:95–168, 1989. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Couty:1989:FRA

[Cou89] Danielle Couty. Formes réduites des automorphismes analytiques de C^{n} à variété linéaire fixe et répulsive. (French) []. Lecture Notes in Mathematics, 1404:346–410, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084083/.

Clapp:1987:ICC

[CP87] Mónica Clapp and Dieter Puppe. Improving categorical coverings of spaces. Lecture Notes in Mathematics, 1283:38–42, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081416/.

Cenkl:1988:CN

[CP88] Bohumil Cenkl and Richard Porter. Cohomology of nilmanifolds. Lecture Notes in Mathematics, 1318:73–86, 1988. CODEN LN-MAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077795/.

Carr:1989:TST

[CP89a] Donna M. Carr and Donald H. Pelletier. Towards a structure theory for ideals on P_{κ}λ. Lecture Notes in Mathematics, 1401: 41–54, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097330/.

Chen:1989:DSP

[CP89b] Xiuxiong Chen and Chia-Kuei Peng. Deformation of surfaces preserving principal curvatures. Lecture Notes in Mathematics, 1369: 63–70, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087527/.

Cohen:1989:SRS

[CP89c] F. R. Cohen and F. P. Peterson. Some remarks on the space ℑJ. Lecture Notes in Mathematics, 1370:117–125, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085222/.

Chin:1985:SDO

[CQ85] Yuan-Shun Chin and Yuanxun Qin. On surfaces defined by ordinary differential equations: a new approach to Hilbert’s 16th Problem. Lecture Notes in Mathematics, 1151:115–131, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074720/.

Craig:1986:LID

[Cra86a] Walter Craig. The Lyapounov index, the density of states and their regularity for general stochastic potentials. Lecture Notes in Mathematics, 1186:252–257, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076845/.

Crauel:1986:LEI

[Cra86b] Hans Crauel. Lyapunov exponents and invariant measures of stochastic systems on manifolds. Lecture Notes in Mathematics, 1186:271–291, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076848/.

Crabb:1987:SS

[Cra87] M. C. Crabb. On the stable splitting of U(n) and ωU(n). Lecture Notes in Mathematics, 1298:35–53, 1987. CODEN LN-MAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082999/.

Cheney:1987:TDI

[CRD87] Margaret Cheney, James H. Rose, and Brian DeFacio. Threedimensional inverse scattering. Lecture Notes in Mathematics, 1285:46–54, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080580/.

Campbell:1989:COR

[CRT89] Colin M. Campbell, E. F. Robertson, and R. M. Thomas. On certain one-relator products of cyclic groups. Lecture Notes in Mathematics, 1398:52–64, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086239/.

Crumeyrolle:1985:CPG

[Cru85a] Albert Crumeyrolle. Constante de Planck et géométrie symplectique. (French) []. Lecture Notes in Mathematics, 1165:84– 107, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076147/.

Crumeyrolle:1985:CAL

[Cru85b] Albert Crumeyrolle. Constructions d’algèbres de Lie graduées orthosymplectiques et conformosymplectiques minkowskiennes. (French) []. Lecture Notes in Mathematics, 1165:52–83, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076146/.

Crumeyrolle:1987:PTP

[Cru87] A. Crumeyrolle. Le principe de trialité, le principe d’intéraction et les algèbres de Lie graduées orthosymplectiques. (French) []. Lecture Notes in Mathematics, 1251:49–64, 1987. CODEN LN-MAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077314/.

Campbell:1989:EPF

[CRW89] Colin M. Campbell, E. F. Robertson, and P. D. Williams. Efficient presentations for finite simple groups and related groups. Lecture Notes in Mathematics, 1398:65–72, 1989. CODEN LN-MAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086240/.

Cappel:1985:TG

[CS85a] Sylvain E. Cappell and Julius L. Shaneson. Torsion in L-groups. Lecture Notes in Mathematics, 1126:22–50, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074437/.

Cheeger:1985:DCG

[CS85b] Jeff Cheeger and James Simons. Differential characters and geometric invariants. Lecture Notes in Mathematics, 1167:50– 80, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075216/.

Chen:1986:SCW

[CS86] Benito Chen and P. G. Saffman. Stability of capillary waves on deep water. Lecture Notes in Mathematics, 1230:199–207, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072682/.

Carlson:1987:SVW

[CS87a] James A. Carlson and Carlos Simpson. Shimura varieties of weight two Hodge structures. Lecture Notes in Mathematics, 1246:1– 15, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077525/.

Cima:1987:PMB

[CS87b] Joseph A. Cima and Ted Suffridge. Proper mappings between balls in C^{n}. Lecture Notes in Mathematics, 1268:66–82, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097297/.

Cameron:1988:UFI

[CS88a] R. H. Cameron and D. A. Storvick. Unbounded Feynman integrable functionals defined in terms of analytic functions. Lecture Notes in Mathematics, 1351:78–92, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081244/.

Cotlar:1988:TLH

[CS88b] Mischa Cotlar and Cora Sadosky. Toeplitz liftings of Hankel forms. Lecture Notes in Mathematics, 1302:22–43, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078861/.

Cwikel:1988:EFT

[CS88c] Michael Cwikel and Yoram Sagher. An extension of Fourier type to quasi-Banach spaces. Lecture Notes in Mathematics, 1302:171– 176, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078873/.

Casazza:1989:BBS

[CS89a] Peter G. Casazza and Thaddeus J. Shura. Block basic sequences in Tsirelson’s space. Lecture Notes in Mathematics, 1363:19– 23, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085270/.

Casazza:1989:BLO

[CS89b] Peter G. Casazza and Thaddeus J. Shura. Bounded linear operators on T and the “blocking” principle. Lecture Notes in Mathematics, 1363:24–34, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085271/.

Casazza:1989:ETA

[CS89c] Peter G. Casazza and Thaddeus J. Shura. Embedding theorems about T and T. Lecture Notes in Mathematics, 1363:54–59, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085274/.

Casazza:1989:FJC

[CS89d] Peter G. Casazza and Thaddeus J. Shura. The Figiel–Johnson construction of Tsirelson’s space. Lecture Notes in Mathematics, 1363:8–18, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085269/.

Casazza:1989:IBS

[CS89e] Peter G. Casazza and Thaddeus J. Shura. Isomorphisms between subspaces of Tsirelson’s space which are spanned by subsequences of {t_{n}}_{n=1}^{∞}. Lecture Notes in Mathematics, 1363:60–66, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085275/.

Casazza:1989:MTS

[CS89f] Peter G. Casazza and Thaddeus J. Shura. Modified Tsirelson’s space: T_{M}. Lecture Notes in Mathematics, 1363:48–53, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085273/.

Casazza:1989:PUV

[CS89g] Peter G. Casazza and Thaddeus J. Shura. Permutations of the unit vector basis of Tsirelson’s space. Lecture Notes in Mathematics, 1363:67–81, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085276/.

Casazza:1989:PTC

[CS89h] Peter G. Casazza and Thaddeus J. Shura. Precursors of the Tsirelson construction. Lecture Notes in Mathematics, 1363:1– 7, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085268/.

Casazza:1989:SFC

[CS89i] Peter G. Casazza and Thaddeus J. Shura. Some final comments. Lecture Notes in Mathematics, 1363:119–120, 1989. CODEN LN-MAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085279.

Casazza:1989:SUV

[CS89j] Peter G. Casazza and Thaddeus J. Shura. Subsequences of the unit vector basis of Tsirelson’s space. Lecture Notes in Mathematics, 1363:35–47, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085272/.

Casazza:1989:UBC

[CS89k] Peter G. Casazza and Thaddeus J. Shura. Unconditional bases for complemented subspaces of Tsirelson’s space. Lecture Notes in Mathematics, 1363:82–94, 1989. CODEN LNMAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085277/.

Casazza:1989:VT

[CS89l] Peter G. Casazza and Thaddeus J. Shura. Variations on a theme. Lecture Notes in Mathematics, 1363:95–118, 1989. CODEN LN-MAA2. ISBN 3-540-50678-0 (print), 3-540-46069-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085278/.

Csorgo:1989:ESA

[Csö89] Sándor Csörgő. An extreme-sum approximation to infinitely divisible laws without a normal component. Lecture Notes in Mathematics, 1391:47–58, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083379/.

Carlson:1988:IMH

[CT88] James A. Carlson and Domingo Toledo. Integral manifolds, harmonic mappings, and the Abelian subspace problem. Lecture Notes in Mathematics, 1352:60–74, 1988. CODEN LN-MAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082017/.

Cohen:1989:LIT

[CT89] Ralph L. Cohen and Ulrike Tillmann. Lectures on immersion theory. Lecture Notes in Mathematics, 1369:71–124, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087528/.

Colliot-Thelene:1987:ASC

[CTKS87] Jean-Louis Colliot-Thélène, Dimitri Kanevsky, and Jean-Jacques Sansuc. Arithmétique des surfaces cubiques diagonales. (French) []. Lecture Notes in Mathematics, 1290:1–108, 1987. CODEN LNMAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078705/.

Curtis:1986:TTR

[Cur86] Charles W. Curtis. Topics in the theory of representations of finite groups. Lecture Notes in Mathematics, 1185:58–87, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076171/.

Cannarsa:1987:EUR

[CV87] Piermarco Cannarsa and Vincenzo Vespri. Existence and uniqueness results for a non linear stochastic partial differential equation. Lecture Notes in Mathematics, 1236:1–24, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072880/.

Cliff:1985:CPO

[CW85] G. H. Cliff and A. R. Weiss. Crossed product orders and wild ramification. Lecture Notes in Mathematics, 1142:96–104, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074794/.

Cerin:1987:BFP

[ČW87] Z. Čerin and T. Watanabe. Borsuk fixed point theorem for multivalued maps. Lecture Notes in Mathematics, 1283:30–37, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081415/.

Cazenave:1989:SRN

[CW89] Thierry Cazenave and Fred B. Weissler. Some remarks on the nonlinear Schrödinger equation in the critical case. Lecture Notes in Mathematics, 1394:18–29, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086749/.

Ye:1986:FCI

[cY86] Jia chen Ye. On the first Cartan invariant of the groups SL(3, p^{n}) and SU(3, p^{n}). Lecture Notes in Mathematics, 1185:388–400, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076185/.

Cermak:1986:FES

[ČZ86] L. Čermák and M. Zlámal. Finite element solution of a nonlinear diffusion problem with a moving boundary. Lecture Notes in Mathematics, 1192:291–294, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076083/.

DaPrato:1985:SRS

[Da 85] G. Da Prato. Some results on stationary Bellman equation in Hilbert spaces. Lecture Notes in Mathematics, 1119:32–51, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074779/.

DAntoni:1985:NPS

[D’A85a] C. D’Antoni. Normal product states and nuclearity: New aspects of algebraic quantum field theory. Lecture Notes in Mathematics, 1136:143–150, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074468/.

deAcosta:1985:LDS

[dA85b] A. de Acosta. On large deviations of sums of independent random vectors. Lecture Notes in Mathematics, 1153:1–14, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074942/.

DAngelo:1987:FTC

[D’A87a] John P. D’Angelo. Finite-type conditions for real hypersurfaces in C^{n}. Lecture Notes in Mathematics, 1268:83–102, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097298/.

DAngelo:1987:ICD

[D’A87b] John P. D’Angelo. Iterated commutators and derivatives of the Levi form. Lecture Notes in Mathematics, 1268:103–110, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097299/.

DalMaso:1986:CUC

[Dal86] G. Dal Maso. Convergence of unilateral convex sets. Lecture Notes in Mathematics, 1190:181–190, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076706/.

Daoudi:1989:EFG

[Dao89] A. Daoudi. Étude de fonctions de Green associées à des opérateurs elliptiques à coéfficients complexes. (French) []. Lecture Notes in Mathematics, 1393:58–71, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085771/.

Daubechies:1987:DSC

[Dau87] Ingrid Daubechies. Discrete sets of coherent states and their use in signal analysis. Lecture Notes in Mathematics, 1285:73–82, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080582/.

Dauge:1988:FSF

[Dau88a] Monique Dauge. Fredholm and semi-Fredholm results. Lecture Notes in Mathematics, 1341:26–57, 1988. CODEN LN-MAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086685/.

Dauge:1988:I

[Dau88b] Monique Dauge. Introduction. Lecture Notes in Mathematics, 1341:2–7, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086683/.

Dauge:1988:LO

[Dau88c] Monique Dauge. Laplace operator. Lecture Notes in Mathematics, 1341:153–170, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086689/.

Dauge:1988:Pa

[Dau88d] Monique Dauge. Preliminaries. Lecture Notes in Mathematics, 1341:8–24, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086684/.

Dauge:1988:Pb

[Dau88e] Monique Dauge. Proofs. Lecture Notes in Mathematics, 1341: 58–103, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086686/.

Dauge:1988:SAE

[Dau88f] Monique Dauge. Singularities along the edges. Lecture Notes in Mathematics, 1341:128–152, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086688/.

Dauge:1988:TDD

[Dau88g] Monique Dauge. Two-dimensional domains. Lecture Notes in Mathematics, 1341:104–127, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086687/.

Dauge:1988:VBVb

[Dau88h] Monique Dauge. Variational boundary value problems on polyhedral domains. Lecture Notes in Mathematics, 1341:186–212, 1988. CODEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086691/.

Dauge:1988:VBVa

[Dau88i] Monique Dauge. Variational boundary value problems on smooth domains. Lecture Notes in Mathematics, 1341:172–185, 1988. CO-DEN LNMAA2. ISBN 3-540-50169-X (print), 3-540-45942-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086690/.

Davie:1985:MNR

[Dav85a] A. M. Davie. Matrix norms related to Grothendieck’s inequality. Lecture Notes in Mathematics, 1166:22–26, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074689/.

Davis:1985:HDA

[Dav85b] James F. Davis. Higher diagonal approximations and skeletons of K(π, l)’s. Lecture Notes in Mathematics, 1126:51–61, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074438/.

Daverman:1987:ICA

[Dav87a] Robert J. Daverman. The intimate connections among decomposition theory, embedding theory, and manifold structure theory. Lecture Notes in Mathematics, 1283:43–47, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081417/.

Davies:1987:BMC

[Dav87b] Ian M. Davies. A “Brownian motion” with constant speed. Lecture Notes in Mathematics, 1250:31–35, 1987. CODEN LN-MAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077346/.

Davis:1987:BYI

[Dav87c] Burgess Davis. On the Barlow–Yor inequalities for local time. Lecture Notes in Mathematics, 1247:218–220, 1987. CODEN LN-MAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077636/.

Davis:1987:BAS

[Dav87d] Donald M. Davis. The Bo–Adams spectral sequence: Some calculations and a proof of its vanishing line. Lecture Notes in Mathematics, 1286:267–285, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078745/.

David:1989:SIS

[Dav89] Guy David. Singular integrals on surfaces. Lecture Notes in Mathematics, 1384:159–167, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086800/.

Dawson:1986:MVP

[Daw86a] Donald Dawson. Measure-valued processes construction, qualitative behavior and stochastic geometry. Lecture Notes in Mathematics, 1212:69–93, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076239/.

Dawson:1986:SEH

[Daw86b] Donald Dawson. Stochastic ensembles and hierarchies. Lecture Notes in Mathematics, 1203:20–37, 1986. CODEN LN-MAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076870/.

Day:1985:SIR

[Day85] Alan Day. Some independence results in the co-ordinization of Arguesian lattices. Lecture Notes in Mathematics, 1149:37–45, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098454/.

deBruin:1985:SPA

[dB85] Marcel G. de Bruin. Simultaneous Padé approximation and orthogonality. Lecture Notes in Mathematics, 1171:74–83, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076532/.

DeFacio:1987:IUS

[DB87] B. DeFacio and O. Brander. Information, uncertainty and the singular value decomposition of the filtered Fourier transformation. Lecture Notes in Mathematics, 1285:83–90, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080583/.

deBruin:1987:MGC

[dBJ87] Marcel G. de Bruin and Lisa Jacobsen. Modification of generalised continued fractions. I. Definition and application to the limitperiodic case. Lecture Notes in Mathematics, 1237:161–176, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072462/.

deChavez:1985:CMF

[dC85a] J. Ruíz de Chávez. Compensation multiplicative et ^{≪}produits de Wick^{≫}. (French) []. Lecture Notes in Mathematics, 1123:242– 247, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075853/.

deChavez:1985:EFP

[dC85b] J. Ruíz de Chávez. Espaces de Fock pour les processus de Wiener et de Poisson. (French) []. Lecture Notes in Mathematics, 1123: 230–241, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075852/.

deChavez:1985:ISM

[dC85c] J. Ruíz de Chávez. Sur les intégrales stochastiques multiples. (French) []. Lecture Notes in Mathematics, 1123:248–257, 1985.

CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075854/.

deChavez:1986:PCO

[dC86] J. Ruíz de Chávez. Sur la positivité de certains opérateurs. (French) []. Lecture Notes in Mathematics, 1204:338–340, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075725/.

deChavez:1988:RPD

[dC88] Juan Ruíz de Chávez. Une remarque sur les processus de Dirichlet forts. (French) []. Lecture Notes in Mathematics, 1321:85, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084121.

doCarmo:1989:HCM

[dC89] Manfredo P. do Carmo. Hypersurfaces of constant mean curvature. Lecture Notes in Mathematics, 1410:128–144, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086419/.

Dajczer:1985:ECN

[DD85] Marcos Dajczer and Peter Dombrowski. Examples of 1codimensional non totally geodesic isometric immersions of pseudo-Riemannian space forms with the same positive constant curvature and the same space-like rank. Lecture Notes in Mathematics, 1156:59–73, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075086/.

Doss:1986:EGD

[DD86] H. Doss and M. Dozzi. Estimations de grandes déviations pour les processus de diffusion à paramètre multidimensionnel. (French) []. Lecture Notes in Mathematics, 1204:68–80, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075713/.

Destuynder:1985:NMF

[DDL85] P. Destuynder, M. Djaoua, and S. Lescure. On a numerical method for fracture mechanics. Lecture Notes in Mathematics, 1121:69– 84, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076263/.

deLaPradelle:1989:PGO

[de 89] A. de La Pradelle. Potentiels de Green sur les ouverts fins. (French) []. Lecture Notes in Mathematics, 1393:202–215, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085781/.

Debiard:1987:SDH

[Deb87] Amédée Debiard. Système différentiel hypergéométrique et parties radiales des opérateurs invariants des espaces symétriques de type BC_{p}. (French) []. Lecture Notes in Mathematics, 1296:42– 124, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078523/.

Debarre:1989:DED

[Deb89] Olivier Debarre. Une démonstration élémentaire du théorème de Torelli pour les intersections de trois quadriques génériques de dimension impaire. (French) [An elementary proof of the theorem of Torelli to the intersections of three quadrics of generic odd dimension]. Lecture Notes in Mathematics, 1399:40–47, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095967/.

Dechevski:1988:MI

[Dec88] L. T. Dechevski. τ-Moduli and interpolation. Lecture Notes in Mathematics, 1302:177–190, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078874/.

Deheuvels:1989:AAC

[Deh89] René Deheuvels. Une application des algèbres de Clifford. (French) []. Lecture Notes in Mathematics, 1410:108–120, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086417/.

Dellacherie:1985:QRM

[Del85] C. Dellacherie. Quelques résultats sur les maisons de jeux analytiques. (French) []. Lecture Notes in Mathematics, 1123:222– 229, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075851/.

DellAntonio:1986:EFE

[Del86] G. F. Dell’Antonio. Electric field and effective dielectric constant in random media with non-linear response. Lecture Notes in Mathematics, 1158:81–103, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080210/.

Delaporte:1988:CCT

[Del88] Jacques Delaporte. Convoluteurs continus et topologie stricte. (French) []. Lecture Notes in Mathematics, 1359:135–141, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086594/.

Dellacherie:1989:CNP

[Del89] C. Dellacherie. Sur la caractérisation des noyaux potentiels. (French) []. Lecture Notes in Mathematics, 1393:78–95, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085773/.

Demailly:1986:IBK

[Dem86a] Jean-Pierre Demailly. Sur l’identité de Bochner–Kodaira–Nakano en géométrie hermitienne. (French) []. Lecture Notes in Mathematics, 1198:88–97, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077045/.

Demailly:1986:EFH

[Dem86b] Jean-Pierre Demailly. Un exemple de fibré holomorphe non de Stein à fibre C^{2} au-dessus du disque ou du plan. (French) []. Lecture Notes in Mathematics, 1198:98–104, 1986. CODEN LN-MAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077046/.

Demailly:1987:TAF

[Dem87a] Jean-Pierre Demailly. Sur les théorèmes d’annulation et de finitude de T. Ohsawa et O. Abdelkader. (French) []. Lecture Notes in Mathematics, 1295:48–58, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081977/.

Demailly:1987:PSC

[Dem87b] Jean-Pierre Demailly. Une preuve simple de la conjecture de Grauert–Riemenschneider. (French) []. Lecture Notes in Mathematics, 1295:24–47, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081976/.

Demailly:1988:VTT

[Dem88] Jean-Pierre Demailly. Vanishing theorems for tensor powers of a positive vector bundle. Lecture Notes in Mathematics, 1339: 86–105, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083049/.

Denkowska:1986:PTS

[Den86] Z. Denkowska. Une présentation de la théorie des sous-analytiques sans désingularisation. (French) []. Lecture Notes in Mathematics, 1198:105–116, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077047/.

Derighetti:1988:CPF

[Der88] Antoine Derighetti. Convoluteurs et projecteurs. (French) []. Lecture Notes in Mathematics, 1359:142–158, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086595/.

Derighetti:1989:PIC

[Der89] Antoine Derighetti. À propos de l’induction des convoluteurs. (French) []. Lecture Notes in Mathematics, 1379:50–63, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087844/.

Desrochers:1985:SDM

[Des85] Maryse Desrochers. Self-duality over the maximal order and torsion Galois modules. Lecture Notes in Mathematics, 1142:105– 111, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074795/.

Desarmenien:1986:FSS

[DF86a] Jacques Désarménien and Dominique Foata. Fonctions symétriques et séries hypergéométriques basiques multivariées, II. (French) []. Lecture Notes in Mathematics, 1234:68–90, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072510/.

Descloux:1986:ATM

[DF86b] J. Descloux and R. Ferro. Analysis of Thacker’s method for solving the linearized shallow water equations. Lecture Notes in Mathematics, 1192:295–302, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076084/.

Dolcetta:1985:RMM

[DFZ85] Italo Capuzzo Dolcetta, Wendell H. Fleming, and Tullio Zolezzi, editors. Recent Mathematical Methods in Dynamic Programming: Proceedings of the Conference held in Rome, Italy, March 26–28, 1984, volume 1119 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1119. URL http:

//link.springer.com/book/10.1007/BFb0074776;   http://
www.springerlink.com/content/978-3-540-39365-8.

deGraaf:1985:TSG

[dG85] J. de Graaf. Two spaces of generalized functions based on harmonic polynomials. Lecture Notes in Mathematics, 1171:164– 173, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076541/.

Dore:1986:SNA

[DG86a] Giovanni Dore and Davide Guidetti. On a singular nonautonomous equation in Banach spaces. Lecture Notes in Mathematics, 1223:92–101, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099185/.

Durr:1986:RCL

[DG86b] Detlef Dürr and Sheldon Goldstein. Remarks on the central limit theorem for weakly dependent random variables. Lecture Notes in Mathematics, 1158:104–118, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080211/.

Dawson:1989:GSS

[DG89] Donald A. Dawson and Luis G. Gorostiza. Generalized solutions of stochastic evolution equations. Lecture Notes in Mathematics, 1390:53–65, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083936/.

Dlab:1986:RTF

[DGM86a] Vlastimil Dlab, Peter Gabriel, and Gerhard Michler, editors. Representation Theory I Finite Dimensional Algebras: Proceedings of the Fourth International Conference on Representations of Algebras held in Ottawa, Canada, August 16–25, 1984, volume 1177 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1177. URL http://link.springer.com/book/10.1007/
BFb0075254;   http://www.springerlink.com/content/978-
3-540-39776-2.

Dlab:1986:RTI

[DGM86b] Vlastimil Dlab, Peter Gabriel, and Gerhard Michler, editors. Representation Theory II Groups and Orders: Proceedings of the Fourth International Conference on Representations of Algebras held in Ottawa, Canada, August 16–25, 1984, volume 1178 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1178. URL http://link.springer.com/book/10.1007/
BFb0075286;   http://www.springerlink.com/content/978-
3-540-39777-9.

Dahmen:1988:LSI

[DGM88] W. Dahmen, T. N. T. Goodman, and Charles A. Micchelli. Local spline interpolation schemes in one and several variables. Lecture Notes in Mathematics, 1354:11–24, 1988. CODEN LN-MAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089580/.

DeBruin:1987:SBZ

[DGR87] M. G. De Bruin, J. Gilewicz, and H.-J. Runckel. A survey of bounds for the zeros of analytic functions obtained by continued fraction methods. Lecture Notes in Mathematics, 1237:1–23, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072450/.

Diamond:1985:CS

[DH85a] H. Diamond and H. Halberstam. The combinatorial sieve. Lecture Notes in Mathematics, 1122:63–73, 1985. CODEN LN-MAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075751/.

Doebner:1985:DGM

[DH85b] Heinz-Dietrich Doebner and Jörg-Dieter Hennig, editors. Differential Geometric Methods in Mathematical Physics: Proceedings of an International Conference Held at the Technical University of Clausthal, FRG, August 30–September 2, 1983, volume 1139 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1139. URL http://link.springer.com/book/10.1007/
BFb0074572;   http://www.springerlink.com/content/978-
3-540-39585-0.

DeMeyer:1986:SRW

[DH86] F. R. DeMeyer and D. Hardy. Semigroup rings which are separable algebras. Lecture Notes in Mathematics, 1197:51–59, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076312/.

Diaz:1988:GSV

[DH88] Steven Diaz and Joe Harris. Geometry of Severi varieties, II: Independence of divisor classes and examples. Lecture Notes in Mathematics, 1311:23–50, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082907/.

Davis:1989:AMS

[DH89] Michael W. Davis and Jean-Claude Hausmann. Aspherical manifolds without smooth or PL structure. Lecture Notes in Mathematics, 1370:135–142, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085224/.

Dhooghe:1985:CIS

[Dho85] P. F. Dhooghe. Completely integrable systems of KdV-type related to isospectral periodic regular difference operators. Lecture Notes in Mathematics, 1139:236–252, 1985. CODEN LN-MAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074587/.

Doebner:1985:WSM

[DHS85] H.-D. Doebner, J. D. Hennig, and I. E. Segal. The work of Steven M. Paneitz. Lecture Notes in Mathematics, 1139:1–5, 1985. CO-DEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074573/.

DiPrisco:1985:MML

[Di 85] Carlos Augusto Di Prisco, editor. Methods in Mathematical Logic: Proceedings of the 6th Latin American Symposium on Mathematical Logic held in Caracas, Venezuela August 1–6, 1983, volume 1130 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1130. URL http:

//link.springer.com/book/10.1007/BFb0075303;   http://
www.springerlink.com/content/978-3-540-39414-3.

DiBlasio:1986:CSP

[Di 86] G. Di Blasio. On a class of semilinear parabolic equations in L^{1}. Lecture Notes in Mathematics, 1223:74–91, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099184/.

Diaconis:1988:ANC

[Dia88] Persi Diaconis. Applications of non-commutative Fourier analysis to probability problems. Lecture Notes in Mathematics, 1362: 51–100, 1988. CODEN LNMAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086179/.

Dickmann:1985:AMT

[Dic85] M. A. Dickmann. Applications of model theory to real algebraic geometry. Lecture Notes in Mathematics, 1130:76–150, 1985. CO-DEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075308/.

Dicks:1988:SRW

[Dic88] Warren Dicks. A survey of recent work on the cohomology of onerelator associative algebras. Lecture Notes in Mathematics, 1328: 75–81, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100917/.

Dieterich:1985:RTG

[Die85a] Ernst Dieterich. Representation types of group rings over complete discrete valuation rings II. Lecture Notes in Mathematics, 1142:112–125, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074796/.

Dieudonne:1985:FCP

[Die85b] J. Dieudonné. Fractions continuées et polynômes orthogonaux dans l’oeuvre de E. N. Laguerre. (French) []. Lecture Notes in Mathematics, 1171:1–15, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076528/.

Diers:1986:CBSa

[Die86a] Yves Diers. Categories of Boolean sheaves of algebras. Lecture Notes in Mathematics, 1187:19–47, 1986. CODEN LN-MAA2. ISBN 3-540-16459-6 (print), 3-540-39796-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075889/.

Diers:1986:CBSb

[Die86b] Yves Diers. Categories of Boolean sheaves of simple algebras. Lecture Notes in Mathematics, 1187:48–113, 1986. CODEN LN-MAA2. ISBN 3-540-16459-6 (print), 3-540-39796-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075890/.

Diers:1986:CBSc

[Die86c] Yves Diers. Categories of Boolean Sheaves of Simple Algebras, volume 1187 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16459-6 (print), 3-540-39796-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 168 pp. LCCN QA3 .L28 no. 1187. URL http:

//link.springer.com/book/10.1007/BFb0075887;   http://
www.springerlink.com/content/978-3-540-39796-0.

Diers:1986:CLB

[Die86d] Yves Diers. Categories of locally Boolean sheaves of simple algebras. Lecture Notes in Mathematics, 1187:114–160, 1986. CODEN LNMAA2. ISBN 3-540-16459-6 (print), 3-540-39796-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075891/.

Diers:1986:I

[Die86e] Yves Diers. Introduction. Lecture Notes in Mathematics, 1187: 1–18, 1986. CODEN LNMAA2. ISBN 3-540-16459-6 (print), 3-540-39796-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075888/.

Dieterich:1987:ARQ

[Die87] Ernst Dieterich. The Auslander–Reiten quiver of an isolated singularity. Lecture Notes in Mathematics, 1273:244–264, 1987. CO-DEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078848/.

Dierkes:1988:GMP

[Die88a] Ulrich Dierkes. A geometric maximum principle, Plateau’s problem for surfaces of prescribed mean curvature, and the two dimensional analogue of the catenary. Lecture Notes in Mathematics, 1357:116–141, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082864/.

Diers:1988:LHC

[Die88b] Yves Diers. Locally Hilbert categories. Lecture Notes in Mathematics, 1348:87–101, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081350/.

Din:1985:NLT

[Din85a] Allan M. Din. Non-linear techniques in two dimensional Grassmannian sigma models. Lecture Notes in Mathematics, 1139:253– 262, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074588/.

Dinculeanu:1985:CWC

[Din85b] Nicolae Dinculeanu. Characterization of weak compactness in function spaces by means of uniform convergence of extended operators. Lecture Notes in Mathematics, 1166:27–37, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074690/.

DiPerna:1987:NCS

[DiP87] Ronald J. DiPerna. Nonlinear conservative systems. Lecture Notes in Mathematics, 1285:99–109, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080585/.

Dixmier:1985:QRC

[Dix85] J. Dixmier. Quelques résultats et conjectures concernant les séries de Poincaré des invariants des formes binaires. (French) []. Lecture Notes in Mathematics, 1146:127–160, 1985. CODEN LN-MAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074537/.

Delfs:1985:BD

[DK85a] Hans Delfs and Manfred Knebusch. The basic definitions. Lecture Notes in Mathematics, 1173:1–86, 1985. CODEN LNMAA2. ISBN 3-540-16060-4 (print), 3-540-39744-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074552/.

Delfs:1985:CT

[DK85b] Hans Delfs and Manfred Knebusch. Completions and triangulations. Lecture Notes in Mathematics, 1173:87–225, 1985. CODEN LNMAA2. ISBN 3-540-16060-4 (print), 3-540-39744-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074553/.

Delfs:1985:H

[DK85c] Hans Delfs and Manfred Knebusch. Homotopies. Lecture Notes in Mathematics, 1173:226–294, 1985. CODEN LNMAA2. ISBN 3-540-16060-4 (print), 3-540-39744-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074554/.

Delfs:1985:LSS

[DK85d] Hans Delfs and Manfred Knebusch. Locally Semialgebraic Spaces, volume 1173 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16060-4 (print), 3-540-39744-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 329 pp. LCCN QA3 .L28 no. 1173. URL http:

//link.springer.com/book/10.1007/BFb0074551;   http://
www.springerlink.com/content/978-3-540-39744-1.

Demlova:1988:MCC

[DK88a] Marie Demlová and Václav Koubek. Minimal congruences and coextensions in semigroups. Lecture Notes in Mathematics, 1320: 28–83, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083422/.

Devaney:1988:DMC

[DK88b] Robert L. Devaney and Linda Keen. Dynamics of maps with constant Schwarzian derivative. Lecture Notes in Mathematics, 1351: 93–100, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081245/.

Devaney:1988:DT

[DK88c] Robert L. Devaney and Linda Keen. Dynamics of tangent. Lecture Notes in Mathematics, 1342:105–111, 1988. CODEN LN-MAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082826/.

Dwyer:1989:HSC

[DK89] W. G. Dwyer and D. M. Kan. Homology with simplicial coefficients. Lecture Notes in Mathematics, 1370:143–149, 1989. CO-DEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085225/.

DeVore:1988:PBT

[DKL88] R. A. DeVore, H. Kierstead, and G. G. Lorentz. A proof of Borsuk’s theorem. Lecture Notes in Mathematics, 1332:195–202, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081621/.

DaPrato:1986:PSL

[DL86a] Giuseppe Da Prato and Alessandra Lunardi. Periodic solutions for linear integrodifferential equations with infinite delay in Banach spaces. Lecture Notes in Mathematics, 1223:49–60, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099182/.

Doukhan:1986:IPE

[DL86b] P. Doukhan and J. R. León. Invariance principles for the empirical measure of a mixing sequence and for the local time of Markov processes. Lecture Notes in Mathematics, 1193:4–21, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077096/.

Daru:1987:ICS

[DL87] Virginie Daru and Alain Lerat. An implicit centered scheme which gives non-oscillatory steady shocks. Lecture Notes in Mathematics, 1270:115–127, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078321/.

Darling:1988:SEI

[DL88] R. W. R. Darling and Yves Le Jan. The statistical equilibrium of an isotropic stochastic flow with negative Lyapounov exponents is trivial. Lecture Notes in Mathematics, 1321:175–185, 1988.

CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084135/.

DeVore:1989:HOR

[DL89] Ronald A. DeVore and Bradley J. Lucier. High order regularity for solutions of the inviscid Burgers equation. Lecture Notes in Mathematics, 1402:147–154, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083873/.

Dijksma:1987:UCK

[DLdS87] Aad Dijksma, Heinz Langer, and Henk de Snoo. Unitary colligations in Krein spaces and their role in the extension theory of isometries and symmetric linear relations in Hilbert spaces. Lecture Notes in Mathematics, 1242:1–42, 1987. CODEN LN-MAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072441/.

delaHarpe:1985:RAD

[dlH85] Pierre de la Harpe. Reduced C∗-algebras of discrete groups which are simple with a unique trace. Lecture Notes in Mathematics, 1132:230–253, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074887/.

delaPradelle:1987:SRF

[dlP87a] A. de la Pradelle. Sur la subordination des résolvantes. (French) []. Lecture Notes in Mathematics, 1235:141–154, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072754/. See erra-

tum and addendum [dlP89].

delaPradelle:1987:RCP

[dlP87b] Arnaud de la Pradelle. Résolvantes complexes et potentiels intrinsèques. (French) []. Lecture Notes in Mathematics, 1235:155– 171, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072755/.

delaPradelle:1989:EAS

[dlP89] A. de la Pradelle. Erratum et addendum à ‘Sur la subordination des Résolvantes’ ‘Résolvantes complexes et potentiels intrinsèques’. (French) []. Lecture Notes in Mathematics, 1393:196– 201, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/

BFb0085780/. See [dlP87a].

dHumieres:1989:RFS

[dLQ89] D. d’Humières, P. Lallemand, and Y. H. Qian. Review of flow simulations using lattice gases. Lecture Notes in Mathematics, 1402: 56–68, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083868/.

Duflot:1985:PA

[DLS85] Jeanne Duflot, Peter S. Landweber, and Robert E. Stong. A problem of Adams on H ∗ (BG; Z_{p}). Lecture Notes in Mathematics, 1172:73–79, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074424/.

Dowbor:1986:GCA

[DLS86] Piotr Dowbor, Helmut Lenzing, and Andrzej Skowroński. Galois coverings of algebras by locally support-finite categories. Lecture Notes in Mathematics, 1177:91–93, 1986. CODEN LN-MAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075260/.

DeMeyer:1985:SG

[DM85a] Frank De Meyer and Richard Mollin. The Schur group. Lecture Notes in Mathematics, 1142:205–210, 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074802/.

DiPrisco:1985:S

[DM85b] Carlos A. Di Prisco and Wiktor Marek. On the space (λ)^{κ}. Lecture Notes in Mathematics, 1130:151–156, 1985. CODEN LN-MAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075309/.

Dietzfelbinger:1985:SRR

[DM85c] Martin Dietzfelbinger and Wolfgang Maass. Strong reducibilities in α- and β-recursion theory. Lecture Notes in Mathematics, 1141:89–120, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076216/.

Dani:1986:PSF

[DM86] S. G. Dani and M. McCrudden. Parabolic subgroups and factor compactness of measures on semisimple Lie groups. Lecture Notes in Mathematics, 1210:34–40, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077170/.

Davis:1987:SMA

[DM87a] M. H. A. Davis and Guillermo L. Gómez M. The semi-martingale approach to the optimal resource allocation in the controlled labour-surplus economy. Lecture Notes in Mathematics, 1250: 36–74, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077347/.

Dwyer:1987:HTC

[DM87b] W. G. Dwyer and G. Mislin. On the homotopy type of the components of map_{∗}(BS^{3}, BS^{3}). Lecture Notes in Mathematics, 1298: 82–89, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083001/.

deMedrano:1988:SSL

[dM88] Santiago López de Medrano. The space of Siegel leaves of a holomorphic vector field. Lecture Notes in Mathematics, 1345:233– 245, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081404/.

Dani:1989:EID

[DM89a] S. G. Dani and M. McCrudden. Embedding infinitely divisible probabilities on the affine group. Lecture Notes in Mathematics, 1379:36–49, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087843/.

Davis:1989:NSR

[DM89b] Donald M. Davis and Mark Mahowald. A new spectrum related to 7-connected cobordism. Lecture Notes in Mathematics, 1370:126–134, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085223/.

Devinatz:1987:SON

[DMR87] Allen Devinatz, Richard Moeckel, and Peter Rejto. On Schrödinger operators with von Neumann–Wigner type potentials. Lecture Notes in Mathematics, 1285:91–98, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080584/.

Dwyer:1987:HU

[DMW87] William G. Dwyer, Haynes R. Miller, and Clarence W. Wilkerson. The homotopic uniqueness of BS^{3}. Lecture Notes in Mathematics, 1298:90–105, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083002/.

DOttaviano:1985:MET

[D’O85] Ítala M. L. D’Ottaviano. The model extension theorems for L_{3}- theories. Lecture Notes in Mathematics, 1130:157–173, 1985. CO-DEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075310/.

Dobrowolski:1985:FEM

[Dob85] Manfred Dobrowolski. On finite element methods for nonlinear elliptic problems on domains with corners. Lecture Notes in Mathematics, 1121:85–103, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076264/.

Dolecki:1986:CBN

[Dol86] S. Dolecki. Continuity of bilinear and non-bilinear polarities. Lecture Notes in Mathematics, 1190:191–213, 1986. CODEN LN-MAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076707/.

Donaldson:1985:MII

[Don85a] S. K. Donaldson. 4-manifolds with indefinite intersection form. Lecture Notes in Mathematics, 1111:309–320, 1985. CODEN LN-MAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084597/.

Donkin:1985:Eb

[Don85b] Stephen Donkin. E_{6}. Lecture Notes in Mathematics, 1140:133– 139, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074646/.

Donkin:1985:Ec

[Don85c] Stephen Donkin. E_{7}. Lecture Notes in Mathematics, 1140:140– 168, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074647/.

Donkin:1985:Ed

[Don85d] Stephen Donkin. E_{8}. Lecture Notes in Mathematics, 1140:169– 226, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074648/.

Donkin:1985:F

[Don85e] Stephen Donkin. F_{4}. Lecture Notes in Mathematics, 1140:89– 132, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074645/.

Donkin:1985:G

[Don85f] Stephen Donkin. G_{2}. Lecture Notes in Mathematics, 1140:86– 88, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074644/.

Donkin:1985:CG

[Don85g] Stephen Donkin. The classical groups. Lecture Notes in Mathematics, 1140:52–75, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074642/.

Donkin:1985:EA

[Don85h] Stephen Donkin. An example and applications. Lecture Notes in Mathematics, 1140:227–233, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074649/.

Donkin:1985:HA

[Don85i] Stephen Donkin. Homological algebra. Lecture Notes in Mathematics, 1140:7–20, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074639/.

Donkin:1985:HAR

[Don85j] Stephen Donkin. Homological algebra revisited. Lecture Notes in Mathematics, 1140:76–85, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074643/.

Donkin:1985:I

[Don85k] Stephen Donkin. Introduction. Lecture Notes in Mathematics, 1140:1–6, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074638/.

Donkin:1985:M

[Don85l] Stephen Donkin. Miscellany. Lecture Notes in Mathematics, 1140: 234–249, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074650/.

Donkin:1985:MHA

[Don85m] Stephen Donkin. More homological algebra. Lecture Notes in Mathematics, 1140:21–30, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074640/.

Donkin:1985:RRA

[Don85n] Stephen Donkin. Rational Representations of Algebraic Groups: Tensor Products and Filtrations, volume 1140 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 254 pp. LCCN QA3

.L28 no. 1140. URL http://link.springer.com/book/10.1007/
BFb0074637;   http://www.springerlink.com/content/978-
3-540-39589-8.

Donkin:1985:R

[Don85o] Stephen Donkin. Reductions. Lecture Notes in Mathematics, 1140: 31–51, 1985. CODEN LNMAA2. ISBN 3-540-15668-2 (print), 3-540-39589-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074641/.

Dong:1986:FST

[Don86a] Chongying Dong. On full subgroups of twisted groups. Lecture Notes in Mathematics, 1185:277–288, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076179/.

Donnelly:1986:PSH

[Don86b] Harold Donnelly. Positive solutions of the heat and eigenvalue equations on Riemannian manifolds. Lecture Notes in Mathematics, 1209:143–151, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076626/.

Donnelly:1986:DPP

[Don86c] Peter Donnelly. Dual processes in population genetics. Lecture Notes in Mathematics, 1212:94–105, 1986. CODEN LN-MAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076240/.

Donagi:1988:SP

[Don88a] Ron Donagi. The Schottky problem. Lecture Notes in Mathematics, 1337:84–137, 1988. CODEN LNMAA2. ISBN 3-540-50080-4 (print), 3-540-45920-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082807/.

Donnay:1988:GFT

[Don88b] Victor J. Donnay. Geodesic flow on the two-sphere part II: Ergodicity. Lecture Notes in Mathematics, 1342:112–153, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082827/.

Donnelly:1988:DER

[Don88c] Harold Donnelly. Decay of eigenfunctions on Riemannian manifolds. Lecture Notes in Mathematics, 1339:106–117, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083050/.

Donaldson:1989:CCY

[Don89a] S. K. Donaldson. Compactification and completion of Yang–Mills moduli spaces. Lecture Notes in Mathematics, 1410:145–160, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086420/.

Donnelly:1989:RDE

[Don89b] Harold Donnelly. Rapid decay of eigenfunctions. Lecture Notes in Mathematics, 1410:161–176, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086421/.

Doplicher:1985:PQF

[Dop85] S. Doplicher. Problems in quantum field theory and in operator algebras. Lecture Notes in Mathematics, 1132:117–128, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074882/.

Dow:1989:CSC

[Dow89] Alan Dow. Compact spaces of countable tightness in the Cohen model. Lecture Notes in Mathematics, 1401:55–67, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097331/.

Dimiev:1985:APF

[DP85] Stančo Dimiev and Tsvetan Pelov. Almost pluriharmonic functions and symplectic mappings on generalized Kähler manifolds. Lecture Notes in Mathematics, 1165:108–121, 1985. CODEN LN-MAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076148/.

dePena:1986:ZRA

[dP86] D. A. de Peña. Zero relation algebras with oriented cycles of noninvertible morphisms. Lecture Notes in Mathematics, 1177: 256–268, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075268/.

DeVore:1988:ISN

[DP88] R. A. DeVore and V. A. Popov. Interpolation spaces and nonlinear approximation. Lecture Notes in Mathematics, 1302:191– 205, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078875/.

delPino:1986:LPS

[dPR86] Guido del Pino and Rolando Rebolledo, editors. Lectures in Probability and Statistics: Lectures given at the Winter School in Probability and Statistics held in Santiago de Chile, volume 1215 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 0-387-16822-2, 3-540-16822-2 (print), 3-540-47092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1215; QA3 .L28 no. 1215 QA273.18. URL http:

//link.springer.com/book/10.1007/BFb0075871;   http://
www.springerlink.com/content/978-3-540-47092-2.

Dovermann:1986:AAG

[DR86] Karl Heinz Dovermann and Melvin Rothenberg. An algebraic approach to the generalized Whitehead group. Lecture Notes in Mathematics, 1217:92–114, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072817/.

Dauge:1987:WFC

[DR87] Monique Dauge and Didier Robert. Weyl’s formula for a class of pseudodifferential operators with negative order on L^{2}(R^{n}). Lecture Notes in Mathematics, 1256:91–122, 1987. CODEN LN-MAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077739/.

Dilworth:1988:SQB

[DR88a] S. J. Dilworth and T. J. Ransford. Spectra in quasi-Banach algebras. Lecture Notes in Mathematics, 1332:175–178, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081619/.

Druck:1988:CSF

[DR88b] Iole F. Druck and Gonzalo E. Reyes. Sur la convergence de suites de fonctions suivant des filtres. (French) []. Lecture Notes in Mathematics, 1348:102–106, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081351/.

Dlab:1989:FRI

[DR89a] Vlastimil Dlab and Claus Michael Ringel. Filtrations of right ideals related to projectivity of left ideals. Lecture Notes in Mathematics, 1404:95–107, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084072/.

Dovermann:1989:GWT

[DR89b] Karl Heinz Dovermann and Melvin Rothenberg. The generalized Whitehead torsion of a g fibre homotopy equivalence. Lecture Notes in Mathematics, 1375:60–88, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085600/.

Draux:1985:OPR

[Dra85] André Draux. Orthogonal polynomials with respect to a linear functional lacunary of order S + 1 in a non-commutative algebra. Lecture Notes in Mathematics, 1171:84–91, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076533/.

Dranishnikov:1987:IDC

[Dra87a] A. N. Dranishnikov. Infinite-dimensional compacta with finite cohomological dimension modulo p. Lecture Notes in Mathematics, 1283:60–64, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081419/.

Dranishnikov:1987:RC

[Dra87b] A. N. Dranishnikov. On resolutions of LC^{n}-compacta. Lecture Notes in Mathematics, 1283:48–59, 1987. CODEN LN-MAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081418/.

Drasin:1987:SRS

[Dra87c] D. Drasin. Some recent successes in value-distribution theory. Lecture Notes in Mathematics, 1277:1–14, 1987. CODEN LN-MAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078242/.

Draux:1987:TPP

[Dra87d] André Draux. Two-point Padé-type and Padé approximants. Lecture Notes in Mathematics, 1237:51–62, 1987. CODEN LN-MAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072454/.

Draux:1988:CPA

[Dra88] André Draux. Convergence of Padé approximants in a noncommutative algebra. Lecture Notes in Mathematics, 1354:118– 130, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089588/.

Drdzinski:1985:RMH

[Drd85] Andrzej Drdziński. Riemannian manifolds with harmonic curvature. Lecture Notes in Mathematics, 1156:74–85, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075087/.

Dress:1985:RPE

[Dre85] Andreas W. M. Dress. Regular polytopes and equivariant tessellations from a combinatorial point of view. Lecture Notes in Mathematics, 1172:56–72, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074423/.

Drewnowski:1986:DPI

[Dre86] L. Drewnowski. On the Dunford and Pettis integrals. Lecture Notes in Mathematics, 1221:1–15, 1986. CODEN LNMAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099108/.

Drezet:1987:GPV

[Dre87] Jean-Marc Drezet. Groupe de Picard des variétés de modules de faisceaux semi-stables sur P_{2}. (French) []. Lecture Notes in Mathematics, 1273:337–362, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078853/.

Drmota:1987:GGS

[Drm87] M. Drmota. Gleichmässig Gleichverteilte Und Schwach Gleichmässig Gleichverteilte Funktionen Modulo 1. (German) []. Lecture Notes in Mathematics, 1262:22–36, 1987. CODEN LN-MAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078593/.

Desch:1986:LSN

[DS86a] W. Desch and W. Schappacher. Linearized stability for nonlinear semigroups. Lecture Notes in Mathematics, 1223:61–73, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099183/.

Dunau:1986:CDT

[DS86b] Jean-Louis Dunau and Henri Sénateur. Une caractérisation du type de la loi de Cauchy–Heisenberg. (French) []. Lecture Notes in Mathematics, 1210:41–57, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077171/.

daSilva:1987:EES

[dS87] J. A. Pereira da Silva. Essential extensions of some infinite dimensional Lie algebras. Lecture Notes in Mathematics, 1251:265– 283, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077325/.

Dittrich:1989:MGV

[DS89] Christel Dittrich and Rolf Sulanke. Möbius geometry VI. Characterization of the homogeneous tori. Lecture Notes in Mathematics, 1410:121–127, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086418/.

deSainte-Catherine:1985:CII

[dSCV85] Myriam de Sainte-Catherine and Gérard Viennot. Combinatorial interpretation of integrals of products of Hermite, Laguerre and Tchebycheff polynomials. Lecture Notes in Mathematics, 1171: 120–128, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076537/.

deSainte-Catherine:1986:ECY

[dSCV86] Myriam de Sainte-Catherine and Gérard Viennot. Enumeration of certain Young tableaux with bounded height. Lecture Notes in Mathematics, 1234:58–67, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072509/.

Dunau:1989:CTS

[DST89] Jean-Louis Dunau, Henri Senateur, and Insa Toulouse. Characterization of the type of some generalizations of the Cauchy distribution. Lecture Notes in Mathematics, 1379:64–74, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087845/.

DaPrato:1987:SPD

[DT87] Giuseppe Da Prato and Luciano Tubaro, editors. Stochastic Partial Differential Equations and Applications: Proceedings of a Conference held in Trento, Italy, Sept. 30–Oct. 5, 1985, volume 1236 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1236. URL http:

//link.springer.com/book/10.1007/BFb0072879;   http://
www.springerlink.com/content/978-3-540-47408-1.

Durran:1988:PD

[DT88a] Richard Durran and Aubrey Truman. Planetesimal diffusions. Lecture Notes in Mathematics, 1325:76–88, 1988. CODEN LN-MAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077917/.

Durran:1988:BMH

[DT88b] Richard M. Durran and Aubrey Truman. Brownian motion on hypersurfaces and computer simulation. Lecture Notes in Mathematics, 1325:89–100, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077918/.

DiLena:1989:SFM

[DT89] G. Di Lena and D. Trigiante. On the spectrum of families of matrices with applications to stability problems. Lecture Notes in Mathematics, 1386:36–53, 1989. CODEN LNMAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089230/.

Dubuc:1985:FEC

[Dub85] Serge Dubuc. Functional equations connected with peculiar curves. Lecture Notes in Mathematics, 1163:33–40, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076415/.

Dubrovin:1988:TOR

[Dub88] B. A. Dubrovin. Theory of operators and real algebraic geometry. Lecture Notes in Mathematics, 1334:42–59, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080422/.

Dudley:1985:EWT

[Dud85] R. M. Dudley. An extended Wichura theorem, definitions of Donsker class, and weighted empirical distributions. Lecture Notes in Mathematics, 1153:141–178, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074949/.

Dumcke:1985:LDL

[Düm85a] R. Dümcke. The low density limit for n-level systems. Lecture Notes in Mathematics, 1136:151–161, 1985. CODEN LN-MAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074469/.

Dumortier:1985:SGV

[Dum85b] Freddy Dumortier. On the structure of germs of vector fields in R^{3} whose linear part generates rotations. Lecture Notes in Mathematics, 1125:15–46, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075633/.

Dunham:1987:ECA

[Dun87] Charles B. Dunham. Existence of Chebyshev approximations by transformations of powered rationals. Lecture Notes in Mathematics, 1237:63–67, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072455/.

Dunham:1988:SUU

[Dun88] Charles B. Dunham. Subsets of unicity in uniform approximation. Lecture Notes in Mathematics, 1354:131–135, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089589/.

Dupont:1988:DSR

[Dup88] Johan L. Dupont. A dual simplicial de Rham complex. Lecture Notes in Mathematics, 1318:87–91, 1988. CODEN LN-MAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077796/.

Durand:1985:SEP

[Dur85a] M. Durand. Singularities in elliptic problems. Lecture Notes in Mathematics, 1121:104–112, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076265/.

Durr:1985:DCL

[Dür85b] Detlef Dürr. Diffusions and central limit theorems. Lecture Notes in Mathematics, 1109:55–67, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101537/.

Durrett:1985:UMT

[Dur85c] R. Durrett. On the unboundedness of martingale transforms. Lecture Notes in Mathematics, 1123:27–36, 1985. CODEN LN-MAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075837/.

Dur:1986:A

[Dür86a] Arne Dür. Applications. Lecture Notes in Mathematics, 1202: 77–108, 1986. CODEN LNMAA2. ISBN 3-540-16771-4 (print), 3-540-39818-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077476/.

Dur:1986:CSI

[Dür86b] Arne Dür. Categorical structures and incidence algebras. Lecture Notes in Mathematics, 1202:1–34, 1986. CODEN LNMAA2. ISBN 3-540-16771-4 (print), 3-540-39818-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077473/.

Dur:1986:MFI

[Dür86c] Arne Dür. Möbius Functions, Incidence Algebras and Power Series Representations, volume 1202 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16771-4 (print), 3-540-39818-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 134 pp. LCCN QA3

.L28 no. 1202. URL http://link.springer.com/book/10.1007/
BFb0077472;   http://www.springerlink.com/content/978-
3-540-39818-9.

Dur:1986:PFE

[Dür86d] Arne Dür. Partitions, functors and exponential formulas. Lecture Notes in Mathematics, 1202:35–49, 1986. CODEN LN-MAA2. ISBN 3-540-16771-4 (print), 3-540-39818-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077474/.

Dur:1986:SCS

[Dür86e] Arne Dür. Sheaflike categorical structures and incidence bialgebras. Lecture Notes in Mathematics, 1202:50–76, 1986. CODEN LNMAA2. ISBN 3-540-16771-4 (print), 3-540-39818-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077475/.

Durrett:1986:SPP

[Dur86f] Richard Durrett. Some peculiar properties of a particle system with sexual reproduction. Lecture Notes in Mathematics, 1212: 106–111, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076241/.

Durfee:1987:ACL

[Dur87] Alan H. Durfee. The L-adic cohomology of links. Lecture Notes in Mathematics, 1246:42–48, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077528/.

Duskin:1988:ACC

[Dus88] John W. Duskin. The Azumaya complex of a commutative ring. Lecture Notes in Mathematics, 1348:107–117, 1988. CODEN LN-MAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081352/.

Duval:1987:DLL

[Duv87] C. Duval. The Dirac & Levy–Leblond equations and geometric quantization. Lecture Notes in Mathematics, 1251:205–221, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077322/.

Denert:1986:CMO

[DV86] Marleen Denert and Jan Van Geel. Classnumbers of maximal orders in central simple algebras over global function fields. Lecture Notes in Mathematics, 1197:60–72, 1986. CODEN LN-MAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076313/.

Delorme:1987:IMS

[DvdB87] Patrick Delorme and van den Ban. Injection de modules sphériques pour les espaces symétriques réductifs dans certaines représentations induites. (French) [Injection of spherical modules for reductive symmetric spaces into certain induced representations]. Lecture Notes in Mathematics, 1243:108–143, 1987. CO-DEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073020/.

Dvir:1985:CPP

[Dvi85] Yoav Dvir. Covering properties of permutation groups. Lecture Notes in Mathematics, 1112:197–221, 1985. CODEN LN-MAA2. ISBN 3-540-13916-8 (print), 3-540-39142-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072288/.

Dydak:1987:SLC

[DW87] Jerzy Dydak and John Walsh. Sheaves that are locally constant with applications to homology manifolds. Lecture Notes in Mathematics, 1283:65–87, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081420/.

Dwyer:1988:M

[DW88] William G. Dwyer and Clarence W. Wilkerson. Maps of BZ/pZ to BG. Lecture Notes in Mathematics, 1318:92–98, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077797/.

Dwinger:1985:UOC

[Dwi85] Philip Dwinger. Unary operations on completely distributive complete lattices. Lecture Notes in Mathematics, 1149:46–81, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098455/.

Dwilewicz:1987:BST

[Dwi87] Roman Dwilewicz. Bergman–Szegő type theory for CR structures. Lecture Notes in Mathematics, 1277:15–34, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078243/.

Dylawerski:1988:DMB

[Dyl88] Grzegorz Dylawerski. An S^{1}-degree and S^{1}-maps between representation spheres. Lecture Notes in Mathematics, 1361:14– 28, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083032/.

Dieckerhoff:1985:PEO

[DZ85] R. Dieckerhoff and E. Zehnder. An ‘a-priori’ estimate for oscillatory-equations. Lecture Notes in Mathematics, 1125:9– 14, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075632/.

Dwyer:1987:MBC

[DZ87] W. Dwyer and A. Zabrodsky. Maps between classifying spaces. Lecture Notes in Mathematics, 1298:106–119, 1987. CODEN LN-MAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083003/.

Daoudi:1989:ADM

[DZ89a] A. Daoudi and M. Zahid. Approximation discrète de mesures dans un espace harmonique. (French) []. Lecture Notes in Mathematics, 1393:72–77, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085772/.

Dembo:1989:RAS

[DZ89b] Amir Dembo and Ofer Zeitouni. On the relation of anticipative Stratonovich and symmetric integrals: a decomposition formula. Lecture Notes in Mathematics, 1390:66–76, 1989. CODEN LN-MAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083937/.

Dziuk:1988:FEB

[Dzi88] Gerhard Dziuk. Finite elements for the Beltrami operator on arbitrary surfaces. Lecture Notes in Mathematics, 1357:142–155, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082865/.

Ellis:1988:NSS

[EB88] David B. Ellis and Michael G. Branton. Non-self-similar attractors of hyperbolic iterated function systems. Lecture Notes in Mathematics, 1342:158–171, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082829/.

Eberlein:1985:SMN

[Ebe85] Patrick Eberlein. Structure of manifolds of nonpositive curvature. Lecture Notes in Mathematics, 1156:86–153, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075088/.

Eberlein:1986:SSN

[Ebe86] Patrick Eberlein. L-subgroups in spaces of nonpositive curvature. Lecture Notes in Mathematics, 1201:41–88, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075647/.

Ebeling:1987:CDD

[Ebe87a] Wolfgang Ebeling. Computation of Dynkin diagrams. Lecture Notes in Mathematics, 1293:21–66, 1987. CODEN LN-MAA2. ISBN 3-540-18686-7 (print), 3-540-48072-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078931/.

Ebeling:1987:DDS

[Ebe87b] Wolfgang Ebeling. Dynkin diagrams for some special singularities. Lecture Notes in Mathematics, 1293:67–116, 1987. CODEN LNMAA2. ISBN 3-540-18686-7 (print), 3-540-48072-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078932/.

Ebeling:1987:ICI

[Ebe87c] Wolfgang Ebeling. Invariants of complete intersections. Lecture Notes in Mathematics, 1293:1–20, 1987. CODEN LNMAA2. ISBN 3-540-18686-7 (print), 3-540-48072-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078930/.

Ebeling:1987:MLM

[Ebe87d] Wolfgang Ebeling. Milnor lattices and monodromy groups. Lecture Notes in Mathematics, 1293:117–130, 1987. CODEN LN-MAA2. ISBN 3-540-18686-7 (print), 3-540-48072-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078933/.

Ebeling:1987:MGS

[Ebe87e] Wolfgang Ebeling. Monodromy groups of symmetric vanishing lattices. Lecture Notes in Mathematics, 1293:131–144, 1987. CODEN LNMAA2. ISBN 3-540-18686-7 (print), 3-540-48072-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078934/.

Ecalle:1985:IAC

[Éca85] Jean Écalle. Iteration and analytic classification of local diffeomorphisms of C^{ν}. Lecture Notes in Mathematics, 1163:41–48, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076416/.

Eckmann:1987:NGA

[Eck87] Beno Eckmann. Nilpotent group action and Euler characteristic. Lecture Notes in Mathematics, 1298:120–123, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083004/.

Ede:1986:CSD

[Ede86] D. A. Ede. Computer simulation of developmental processes in biology: Models for the developing limb. Lecture Notes in Mathematics, 1212:112–123, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076242/.

Edgar:1985:CMC

[Edg85] G. A. Edgar. Complex martingale convergence. Lecture Notes in Mathematics, 1166:38–59, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074691/.

Ein:1988:VCQ

[EEK88] Lawrence Ein, David Eisenbud, and Sheldon Katz. Varieties cut out by quadrics: Scheme-theoretic versus homogeneous generation of ideals. Lecture Notes in Mathematics, 1311:51–70, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082908/.

Eells:1985:MBI

[Eel85] James Eells. Minimal branched immersions into three-manifolds. Lecture Notes in Mathematics, 1167:81–94, 1985. CODEN LN-MAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075217/.

Epstein:1985:LUG

[EEŚ85] Marcelo Epstein, Marek Elżanowski, and Jedrzej Śniatycki. Locality and uniformity in global elasticity. Lecture Notes in Mathematics, 1139:300–310, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074591/.

Eigen:1988:PSI

[EH88a] Stanley Eigen and Arshag Hajian. Poincaré sequences in infinite measure spaces and complementing subsets of the integers. Lecture Notes in Mathematics, 1342:154–157, 1988. CODEN LN-MAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082828/.

Evans:1988:MQD

[EH88b] M. Evans and R. L. Hudson. Multidimensional quantum diffusions. Lecture Notes in Mathematics, 1303:69–88, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078055/.

Ehrenpreis:1985:TNP

[Ehr85] Leon Ehrenpreis. Transcendental numbers and partial differential equations. Lecture Notes in Mathematics, 1135:112–125, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074603/.

Ein:1988:VTV

[Ein88] Lawrence Ein. Vanishing theorems for varieties of low codimension. Lecture Notes in Mathematics, 1311:71–75, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082909/.

Echterhoff:1988:CGE

[EK88a] Siegfried Echterhoff and Eberhard Kaniuth. Certain group extensions and twisted covariance algebras with generalized continuous trace. Lecture Notes in Mathematics, 1359:159–169, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086596/.

Essannouni:1988:SAR

[EK88b] Hassane Essannouni and Amin Kaidi. Semiprime alternative rings with A.C.C. Lecture Notes in Mathematics, 1328:82–93, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100918/.

Ekedahl:1986:TES

[Eke86a] Torsten Ekedahl. Two examples of smooth projective varieties with non-zero Massey products. Lecture Notes in Mathematics, 1183:128–132, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075454/.

Ekeland:1986:TRC

[Eke86b] I. Ekeland. Two results in convex analysis. Lecture Notes in Mathematics, 1190:215–228, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076708/.

Ekeland:1988:SVP

[Eke88] Ivar Ekeland. Some variational problems arising from mathematical economics. Lecture Notes in Mathematics, 1330:1–18, 1988. CODEN LNMAA2. ISBN 3-540-50003-0 (print), 3-540-45883-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078156/.

Ekstrom:1989:ACG

[Eks89] Eva Kristina Ekström. The Auslander condition on graded and filtered Noetherian rings. Lecture Notes in Mathematics, 1404: 220–245, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084078/.

ElKory:1987:GVF

[El 87] Mohamed El Kory. Généralisation de la variation d’une fonction et application. (French) []. Lecture Notes in Mathematics, 1235: 57–87, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072749/.

Elencwajg:1988:EC

[EL88] G. Elencwajg and P. Le Barz. Explicit computations in Hilb^{3}P^{2}. Lecture Notes in Mathematics, 1311:76–100, 1988. CODEN LN-MAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082910/.

Elbert:1985:NC

[Elb85] Á. Elbert. On a Nehari’s constant. Lecture Notes in Mathematics, 1151:142–147, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074722/.

Elduque:1989:LAS

[Eld89] Alberto Elduque. On Lie algebras with a subalgebra of codimension one. Lecture Notes in Mathematics, 1373:58–66, 1989. CODEN LNMAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088887/.

Elezovic:1987:NOC

[Ele87] Neven Elezović. p-nuclear operators and cylindrical measures on tensor products of Banach spaces. Lecture Notes in Mathematics, 1242:418–432, 1987. CODEN LNMAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072447/.

Eliashberg:1988:SDW

[Eli88] Ya. M. Eliashberg. The structure of 1-dimensional wave fronts, non-standard Legendrian loops and Bennequin’s theorem. Lecture Notes in Mathematics, 1346:7–12, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082768/.

Ellia:1989:LHF

[Ell89] Ph. Ellia. Sur les lacunes d’Halphen. (French) []. Lecture Notes in Mathematics, 1389:43–65, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085923/.

Elschner:1985:ASC

[Els85a] Johannes Elschner. Appendix. Suboptimal convergence of the Galerkin method with splines for elliptic pseudodifferential equations. Lecture Notes in Mathematics, 1128:187–190, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075283/.

Elschner:1985:DPO

[Els85b] Johannes Elschner. Degenerate pseudodifferential operators on a closed curve. Lecture Notes in Mathematics, 1128:123–163, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075281/.

Elschner:1985:EDO

[Els85c] Johannes Elschner. Elliptic differential operators in R^{n} degenerating at one point. Lecture Notes in Mathematics, 1128:92– 122, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075280/.

Elschner:1985:FEM

[Els85d] Johannes Elschner. A finite element method for pseudodifferential equations on a closed curve. Lecture Notes in Mathematics, 1128:164–186, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075282/.

Elschner:1985:FDO

[Els85e] Johannes Elschner. Fuchsian differential operators with one singular point. Lecture Notes in Mathematics, 1128:13–36, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075277/.

Elschner:1985:LODa

[Els85f] Johannes Elschner. Lieear ordinary differential operators with one singular point. Lecture Notes in Mathematics, 1128:37–71, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075278/.

Elschner:1985:LODb

[Els85g] Johannes Elschner. Linear ordinary differential operators with several singular points. Lecture Notes in Mathematics, 1128:72– 91, 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075279/.

Elschner:1985:SOD

[Els85h] Johannes Elschner. Singular Ordinary Differential Operators and Pseudodifferential Equations, volume 1128 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15194-X (print), 3-540-39294-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 190 (est.) pp. LCCN QA3

.L28 no. 1128. URL http://link.springer.com/book/10.1007/
BFb0075276;   http://www.springerlink.com/content/978-
3-540-39294-1.

Elworthy:1988:GAD

[Elw88a] David Elworthy. Geometric aspects of diffusions on manifolds. Lecture Notes in Mathematics, 1362:277–425, 1988. CODEN LN-MAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086183/.

Elworthy:1988:BMH

[Elw88b] K. D. Elworthy. Brownian motion and harmonic forms. Lecture Notes in Mathematics, 1316:288–304, 1988. CODEN LN-MAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081936/.

Etgens:1985:PPT

[EM85] Maria Etgens and Zenon Moszner. On pseudo-processes and their extensions. Lecture Notes in Mathematics, 1163:49–58, 1985. CO-DEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076417/.

England:1986:SSC

[EM86] Roland England and Robert M. M. Mattheij. Sequential step control for integration of two-point boundary value problems. Lecture Notes in Mathematics, 1230:221–234, 1986. CODEN LN-MAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072684/.

Edelman:1988:OPM

[EM88] Alan Edelman and Charles A. Micchelli. Orthogonal polynomials in monotone and convex interpolation. Lecture Notes in Mathematics, 1329:20–31, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083351/.

Embid:1987:IVP

[Emb87] Pedro Embid. The initial value problem for the nonlinear equations for zero Mach number combustion. Lecture Notes in Mathematics, 1285:110–118, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080586/.

Emery:1988:CCV

[Éme88] Michel Émery. En cherchant une caractérisation variationnelle des martingales. (French) []. Lecture Notes in Mathematics, 1321: 147–154, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084131/.

Emery:1989:AM

[Éme89] Michel Émery. On the Azéma martingales. Lecture Notes in Mathematics, 1372:66–87, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083961/.

Ebbinghaus:1985:RTW

[EMS85] Heinz-Dieter Ebbinghaus, Gert H. Müller, and Gerald E. Sacks, editors. Recursion Theory Week: Proceedings of a Conference held in Oberwolfach, West Germany April 15–21, 1984, volume 1141 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1141. URL http://link.springer.com/book/10.1007/
BFb0076211;   http://www.springerlink.com/content/978-
3-540-39596-6.

Engel:1986:SCS

[EN86] K.-J. Engel and R. Nagel. On the spectrum of certain systems of linear evolution equations. Lecture Notes in Mathematics, 1223: 102–109, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099186/.

Erdos:1987:PRM

[EN87] Paul Erdős and Melvyn B. Nathanson. Problems and results on minimal bases in additive number theory. Lecture Notes in Mathematics, 1240:87–96, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072974/.

Ellis:1988:ESE

[EN88] R. Ellis and M. Nerurkar. Enveloping semigroup in ergodic theory and a proof of Moore’s ergodicity theorem. Lecture Notes in Mathematics, 1342:172–179, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082830/.

English:1986:LEO

[Eng86] Harald English. Lyapunov exponents and one-dimensional alloys. Lecture Notes in Mathematics, 1186:242–245, 1986. CODEN LN-MAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076843/.

Engler:1987:ERS

[Eng87a] Hans Engler. Existence of radially symmetric solutions of strongly damped wave equations. Lecture Notes in Mathematics, 1248: 40–51, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077414/.

Engquist:1987:COS

[Eng87b] Björn Engquist. Computation of oscillatory solutions to partial differential equations. Lecture Notes in Mathematics, 1270:10– 22, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078314/.

Engler:1989:QHI

[Eng89] Hans Engler. On quasilinear hyperbolic integrodifferential equations in unbounded domains. Lecture Notes in Mathematics, 1394:

44–55, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086751/.

Enoki:1988:SNT

[Eno88] Ichiro Enoki. Stability and negativity for tangent sheaves of minimal Kähler spaces. Lecture Notes in Mathematics, 1339:118– 126, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083051/.

Enss:1985:QST

[Ens85] Volker Enss. Quantum scattering theory for two- and three-body systems with potentials of short and long range. Lecture Notes in Mathematics, 1159:39–176, 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080332/.

Enss:1986:IAO

[Ens86] Volker Enss. Introduction to asymptotic observables for multiparticle quantum scattering. Lecture Notes in Mathematics, 1218: 61–92, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073044/.

Engler:1987:SDS

[ENS87] H. Engler, F. Neubrander, and J. Sandefur. Strongly damped semilinear second order equations. Lecture Notes in Mathematics, 1248:52–62, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077415/.

Enss:1988:ATE

[Ens88] Volker Enss. Asymptotic time evolutions for strictly outgoing multiparticle quantum systems with long-range potentials. Lecture Notes in Mathematics, 1324:65–84, 1988. CODEN LN-MAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100783/.

Erdos:1987:NFW

[EP87] Paul Erdős and Carl Pomerance. On the number of false witnesses for a composite number. Lecture Notes in Mathematics, 1240:97–100, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072975/.

Ellingsrud:1987:VNQ

[EPS87] Geir Ellingsrud, Ragni Piene, and Stein Arild Strømme. On the variety of nets of quadrics defining twisted cubics. Lecture Notes in Mathematics, 1266:84–96, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078179/.

Erdos:1985:SMP

[Erd85] P. Erdős. On some of my problems in number theory I would most like to see solved. Lecture Notes in Mathematics, 1122:74–84, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075752/.

Erdos:1989:R

[Erd89] Paul Erdős. Ramanujan and I. Lecture Notes in Mathematics, 1395:1–17, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086396/.

Erkip:1987:NSB

[Erk87] A. K. Erkip. Normal solvability of boundary value problems in half space. Lecture Notes in Mathematics, 1256:123–134, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077740/.

Erkama:1988:RRM

[Erk88] Timo Erkama. Rational Riemann maps. Lecture Notes in Mathematics, 1351:101–109, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081246/.

Elman:1986:PIN

[ES86] Howard C. Elman and Roy L. Streit. Polynomial iteration for nonsymmetric indefinite linear systems. Lecture Notes in Mathematics, 1230:103–117, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072674/.

Ellia:1987:CPC

[ES87a] Ph. Ellia and I. Sols. On the cohomology of projective curves. Lecture Notes in Mathematics, 1266:74–83, 1987. CODEN LN-MAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078178/.

Ellingsrud:1987:RMS

[ES87b] G. Ellingsrud and S. A. Strømme. On the rationality of the moduli space for stable rank-2 vector bundles on P^{2}. Lecture Notes in Mathematics, 1273:363–371, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078854/.

Erdos:1989:PPW

[ESN89] P. Erdős, M. Szalay, and J. L. Nicolas. Partitions into parts which are unequal and large. Lecture Notes in Mathematics, 1380:19– 30, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086543/.

Espanol:1988:SLB

[Esp88] L. Español. The spectrum lattice of Baer rings and polynomials. Lecture Notes in Mathematics, 1348:118–124, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081353/.

Erdos:1985:PRA

[ESS85] P. Erdős, A. Sárközy, and V. T. Sós. Problems and result on additive properties of general sequences, IV. Lecture Notes in Mathematics, 1122:85–104, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075753/.

Essen:1987:HMH

[Ess87] Matts Essén. Harmonic majorization, harmonic measure and minimal thinness. Lecture Notes in Mathematics, 1275:89–112, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078346/.

Esterle:1986:RSC

[Est86] Jean Esterle. Real semigroups in commutative Banach algebras. Lecture Notes in Mathematics, 1221:16–35, 1986. CODEN LN-MAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099109/.

Esteban:1988:SRS

[Est88] Dionisio Pérez Esteban. Semi-rings and spectral spaces. Lecture Notes in Mathematics, 1328:219–226, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100927/.

Evans:1985:ATD

[Eva85a] David E. Evans. The C∗-algebras of the two-dimensional Ising model. Lecture Notes in Mathematics, 1136:162–176, 1985. CO-DEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074470/.

Evans:1985:QPS

[Eva85b] David E. Evans. Quasi-product states on C∗-algebras. Lecture Notes in Mathematics, 1132:129–151, 1985. CODEN LN-MAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074883/.

Evans:1987:NST

[Eva87] Lawrence C. Evans. Nonlinear semigroup theory and viscosity solutions of Hamilton–Jacobi PDE. Lecture Notes in Mathematics, 1248:63–77, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077416/.

Emery:1987:SPL

[ÉY87] Michel Émery and J. E. Yukich. A simple proof of the logarithmic Sobolev inequality on the circle. Lecture Notes in Mathematics, 1247:173–175, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077632/.

Frigerio:1989:UDN

[FA89] Alberto Frigerio and Carola Aratari. Unitary dilation of a nonlinear quantum Boltzmann equation. Lecture Notes in Mathematics, 1396:89–98, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083546/.

Fabes:1988:LPM

[Fab88] Eugene Fabes. Layer potential methods for boundary value problems on Lipschitz domains. Lecture Notes in Mathematics, 1344: 55–80, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103344/.

Faber:1989:SRC

[Fab89] Carel F. Faber. Some results on the codimension-two Chow group of the moduli space of stable curves. Lecture Notes in Mathematics, 1389:66–75, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085924/.

Factorovich:1989:LBS

[Fac89] I. Factorovich. Limiting behaviour of the sum of I.I.D. random variables and its terms of greatest moduli in the case of logarithmic type tall function. Lecture Notes in Mathematics, 1412: 22–32, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084162/.

Fadell:1989:TVF

[Fad89] E. Fadell. Two vignettes in fixed point theory. Lecture Notes in Mathematics, 1411:46–51, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086439/.

Fajardo:1985:CTG

[Faj85] Sergio Fajardo. Completeness theorems for the general theory of stochastic processes. Lecture Notes in Mathematics, 1130:174– 194, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075311/.

Faltings:1985:AKM

[Fal85] G. Faltings. Arithmetische Kompaktifizierung des Modulraums der abelschen Varietäten. (German) []. Lecture Notes in Mathematics, 1111:321–383, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084598/.

Falkowski:1986:LSK

[Fal86] B.-J. Falkowski. Levy–Schoenberg kernels on Riemannian symmetric spaces of noncompact type. Lecture Notes in Mathematics, 1210:58–67, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077172/.

Faltings:1987:ATS

[Fal87] Gerd Faltings. Arithmetic theory of Siegel modular forms. Lecture Notes in Mathematics, 1240:101–108, 1987. CODEN LN-MAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072976/.

Fannes:1988:AFT

[Fan88] M. Fannes. An application of de Finetti’s theorem. Lecture Notes in Mathematics, 1303:89–102, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078056/.

Fannes:1989:TSS

[Fan89] M. Fannes. Temperature states of spin-boson models. Lecture Notes in Mathematics, 1396:149–157, 1989. CODEN LN-MAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083551/.

Farjoun:1987:HHD

[Far87a] Emmanuel Dror Farjoun. Homotopy and homology of diagrams of spaces. Lecture Notes in Mathematics, 1286:93–134, 1987. CO-DEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078739/.

Farkas:1987:UCH

[Far87b] H. M. Farkas. Unramified coverings of hyperelliptic Riemann surfaces. Lecture Notes in Mathematics, 1275:113–130, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078347/.

Faraut:1988:FDB

[Far88] Jacques Faraut. Formule du binôme généralisée. (French) []. Lecture Notes in Mathematics, 1359:170–180, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086597/.

Farnsteiner:1989:LTM

[Far89] Rolf Farnsteiner. Lie theoretic methods in cohomology theory. Lecture Notes in Mathematics, 1373:93–110, 1989. CODEN LN-MAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088890/.

Fiedler:1986:CSR

[FB86] B. Fiedler and P. Brunovský. Connections in scalar reaction diffusion equations with Neumann boundary conditions. Lecture Notes in Mathematics, 1192:123–128, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076058/.

Feyel:1989:CNL

[Fd89] D. Feyel and A. de La Pradelle. Cônes et noyaux locaux. (French) [Cones and local kernels]. Lecture Notes in Mathematics, 1393: 96–116, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085774/.

Fischbacher:1986:ART

[FdlP86] U. Fischbacher and J. A. de la Peña. Algorithms in representation theory of algebras. Lecture Notes in Mathematics, 1177: 115–134, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075262/.

Fefferman:1988:SRD

[Fef88] Robert Fefferman. Some recent developments in Fourier analysis and H^{P} theory on product domains — II. Lecture Notes in Mathematics, 1302:44–51, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078862/.

Feistauer:1986:MNS

[Fei86a] M. Feistauer. Mathematical and numerical study of nonlinear problems in fluid mechanics. Lecture Notes in Mathematics, 1192: 3–16, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076047/.

Feit:1986:BCD

[Fei86b] Walter Feit. Blocks with cyclic defect groups for some sporadic groups. Lecture Notes in Mathematics, 1178:25–63, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075288/.

Feinsilver:1989:BF

[Fei89] Philip Feinsilver. Bernoulli fields. Lecture Notes in Mathematics, 1396:158–181, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083552/.

Felshtyn:1988:NZF

[Fel88] A. L. Fel’shtyn. New zeta functions for dynamical systems and Nielsen fixed point theory. Lecture Notes in Mathematics, 1346: 33–55, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082770/.

Fenske:1989:FIP

[Fen89] Christian C. Fenske. Fuller’s index for periodic solutions of functional differential equations. Lecture Notes in Mathematics, 1411: 69–74, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086441/.

Fernique:1985:CMG

[Fer85a] X. Fernique. Comparaison de mesures gaussiennes et de mesures produit dans les espaces de Fréchet séparables. (French) []. Lecture Notes in Mathematics, 1153:179–197, 1985. CODEN LN-MAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074950/.

Fernique:1985:LRG

[Fer85b] X. Fernique. Laws of random Gaussian functions, some inequalities. Lecture Notes in Mathematics, 1166:60–64, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074692/.

Ferreira:1985:OAS

[Fer85c] Arturo Vaz Ferreira. On Oka’s analytic set-valued functions and spectral theory. Lecture Notes in Mathematics, 1165:122– 135, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076149/.

Ferrier:1986:MPS

[Fer86] J.-P. Ferrier. H^{1} et multiplicités pour le spectre d’une algèbre. (French) []. Lecture Notes in Mathematics, 1198:117–132, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077048/.

Ferry:1987:EC

[Fer87] Steven C. Ferry. UV^{k}-equivalent compacta. Lecture Notes in Mathematics, 1283:88–114, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081421/.

Fernandez:1988:QKT

[Fer88a] J. L. Fernández. On qualitative Korovkin theorems with adistance. Lecture Notes in Mathematics, 1354:136–139, 1988. CO-DEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0089590/.

Fernandez:1988:PDU

[Fer88b] José L. Fernandez. Problems on distortion under conformal mappings. Lecture Notes in Mathematics, 1344:198, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103352.

Ferrand:1988:CQM

[Fer88c] Jacqueline Ferrand. A characterization of quasiconformal mappings by the behaviour of a function of three points. Lecture Notes in Mathematics, 1351:110–123, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081247/.

Fernique:1989:RFA

[Fer89] Xavier Fernique. Régularité de fonctions aléatoires Gaussiennes stationnaires à valeurs vectorielles. (French) []. Lecture Notes in Mathematics, 1391:66–73, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083381/.

Franks:1989:LFP

[FF89] John Franks and David Fried. The Lefschetz function of a point. Lecture Notes in Mathematics, 1411:83–87, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086443/.

Feigin:1988:MOC

[FFR88] B. L. Feĭgin, D. B. Fuchs, and V. S. Retakh. Massey operations in the cohomology of the infinite dimensional Lie algebra L_{1}. Lecture Notes in Mathematics, 1346:13–31, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082769/.

Fernandez:1986:IM

[FG86] Marisa Fernández and Alfred Gray. The Iwasawa manifold. Lecture Notes in Mathematics, 1209:157–159, 1986. CODEN LN-MAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076628/.

Feichtinger:1988:UAA

[FG88] Hans G. Feichtinger and Karlheinz Gröchenig. A unified approach to atomic decompositions via integrable group representations. Lecture Notes in Mathematics, 1302:52–73, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078863/.

Ferus:1985:GDG

[FGHS85] Dirk Ferus, Robert B. Gardner, Sigurdur Helgason, and Udo Simon, editors. Global Differential Geometry and Global Analysis 1984: Proceedings of a Conference held in Berlin, June 10–14, 1984, volume 1156 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1156. URL http:

//link.springer.com/book/10.1007/BFb0075080;   http://
www.springerlink.com/content/978-3-540-39698-7.

Froberg:1986:FFL

[FGL86] Ralf Fröberg, Tor Gulliksen, and Clas Löfwall. Flat families of local, Artinian algebras with an infinite number of Poincaré series. Lecture Notes in Mathematics, 1183:170–191, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075457/.

Fadell:1989:ITN

[FH89] E. Fadell and S. Y. Husseini. Index theory for noncompact group actions with applications to Borsuk–Ulam theorems. Lecture Notes in Mathematics, 1411:52–68, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086440/.

Frazier:1989:TTL

[FHJW89] M. Frazier, Y.-S. Han, B. Jawerth, and G. Weiss. The T 1 theorem for Triebel–Lizorkin spaces. Lecture Notes in Mathematics, 1384:168–181, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086801/.

Fernique:1986:GSA

[FHMM86] Xavier Fernique, Bernard Heinkel, Paul-André Meyer, and Michael B. Marcus, editors. Geometrical and Statistical Aspects of Probability in Banach Spaces: Actes des Journées SMF de Calcul des Probabilités dans les Espaces de Banach, organisées à Strasbourg les 19 et 20 juin 1985, volume 1193 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 0-387-16487-1, 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1193; QA3 .L28 no. 1193 QA273.43. URL http:

//link.springer.com/book/10.1007/BFb0077094;   http://
www.springerlink.com/content/978-3-540-39826-4.

Felix:1986:R

[FHTT86] Yves Félix, Stephen Halperin, Daniel Tanré, and Jean-Claude Thomas. The radical of π_{∗}(ωS) ⊗ Q. Lecture Notes in Mathematics, 1183:133–135, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075455/.

Fiedler:1988:AG

[Fie88a] Bernold Fiedler. Appendix on genericity. Lecture Notes in Mathematics, 1309:116–134, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082953/.

Fiedler:1988:A

[Fie88b] Bernold Fiedler. Applications. Lecture Notes in Mathematics, 1309:92–105, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082951/.

Fiedler:1988:D

[Fie88c] Bernold Fiedler. Discussion. Lecture Notes in Mathematics, 1309: 106–115, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082952/.

Fiedler:1988:GGTb

[Fie88d] Bernold Fiedler. General global theory. Lecture Notes in Mathematics, 1309:84–91, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082950/.

Fiedler:1988:GGTa

[Fie88e] Bernold Fiedler. Generic global theory. Lecture Notes in Mathematics, 1309:68–83, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082949/.

Fiedler:1988:GLT

[Fie88f] Bernold Fiedler. Generic local theory. Lecture Notes in Mathematics, 1309:48–67, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082948/.

Fiedler:1988:I

[Fie88g] Bernold Fiedler. Introduction. Lecture Notes in Mathematics, 1309:1–14, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082944/.

Fiedler:1988:MR

[Fie88h] Bernold Fiedler. Main results. Lecture Notes in Mathematics, 1309:15–26, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082945/.

Fiedler:1988:NSS

[Fie88i] Bernold Fiedler. No symmetry — a survey. Lecture Notes in Mathematics, 1309:27–34, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082946/.

Fiedler:1988:VS

[Fie88j] Bernold Fiedler. Virtual symmetry. Lecture Notes in Mathematics, 1309:35–47, 1988. CODEN LNMAA2. ISBN 3-540-19234-4 (print), 3-540-39150-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082947/.

Filippi:1985:BIE

[Fil85] Paul J. T. Filippi. Boundary integral equations for sound radiation by a harmonically vibrating baffled plate. Lecture Notes in Mathematics, 1121:113–122, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076266/.

Finkelshtein:1988:RNT

[Fin88a] Andrei Martínez Finkelshtein. On the m-th row of Newtontype (α, β)-PADF tables and singular points. Lecture Notes in Mathematics, 1354:178–187, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089595/.

Finn:1988:CPC

[Fin88b] Robert Finn. Comparison principles in capillarity. Lecture Notes in Mathematics, 1357:156–197, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082866/.

Finn:1988:IG

[Fin88c] Robert Finn. The inclination of an H-graph. Lecture Notes in Mathematics, 1340:40–60, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082884/.

Fischbacher:1986:RFA

[Fis86] U. Fischbacher. The representation-finite algebras with at most 3 simple modules. Lecture Notes in Mathematics, 1177:94–114, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075261/.

Fischer:1987:AST

[Fis87a] Jürgen Fischer. An Approach to the Selberg Trace Formula via the Selberg Zeta-Function, volume 1253 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-15208-3 (print), 3-540-39331-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 184 pp. LCCN QA3 .L28

no. 1253. URL http://link.springer.com/book/10.1007/
BFb0077696;   http://www.springerlink.com/content/978-
3-540-39331-3.

Fischer:1987:BF

[Fis87b] Jürgen Fischer. Basic facts. Lecture Notes in Mathematics, 1253: 14–39, 1987. CODEN LNMAA2. ISBN 3-540-15208-3 (print), 3-540-39331-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077698/.

Fischer:1987:EFA

[Fis87c] Jürgen Fischer. The entire function Ξ associated with the Selberg zeta-function. Lecture Notes in Mathematics, 1253:113–161, 1987. CODEN LNMAA2. ISBN 3-540-15208-3 (print), 3-540-39331-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077700/.

Fischer:1987:GST

[Fis87d] Jürgen Fischer. The general Selberg trace formula. Lecture Notes in Mathematics, 1253:162–175, 1987. CODEN LNMAA2. ISBN 3-540-15208-3 (print), 3-540-39331-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077701/.

Fischer:1987:I

[Fis87e] Jürgen Fischer. Introduction. Lecture Notes in Mathematics, 1253: 1–13, 1987. CODEN LNMAA2. ISBN 3-540-15208-3 (print), 3-540-39331-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077697/.

Fischer:1987:TIR

[Fis87f] Jürgen Fischer. The trace of the iterated resolvent kernel. Lecture Notes in Mathematics, 1253:40–112, 1987. CODEN LN-MAA2. ISBN 3-540-15208-3 (print), 3-540-39331-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077699/.

FitzGerald:1987:AWU

[Fit87a] Carl H. FitzGerald. Arcs on which a univalent function has small magnitude. Lecture Notes in Mathematics, 1275:131–148, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078348/.

Fitzgibbon:1987:LSC

[Fit87b] W. E. Fitzgibbon. Longtime solutions for a class of convection diffusion systems. Lecture Notes in Mathematics, 1285:119– 125, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080587/.

Fitzsimmons:1988:PTS

[Fit88] P. J. Fitzsimmons. Penetration times and Skorohod stopping. Lecture Notes in Mathematics, 1321:166–174, 1988. CODEN LN-MAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084134/.

Frazier:1988:TAD

[FJ88] Michael Frazier and Björn Jawerth. The ϕ-transform and applications to distribution spaces. Lecture Notes in Mathematics, 1302:223–246, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078877/.

Fulling:1987:CFS

[FK87] S. A. Fulling and G. Kennedy. A closed form for the symbol of the resolvent parametrix of an elliptic operator. Lecture Notes in Mathematics, 1285:126–133, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080588/.

Fleischer:1988:EMA

[FK88] Vladimir Fleischer and Ulrich Knauer. Endomorphism monoids of acts are wreath products of monoids with small categories. Lecture Notes in Mathematics, 1320:84–96, 1988. CODEN LN-MAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083423/.

Finashin:1988:NDH

[FKV88] S. M. Finashin, M. Kreck, and O. Ya. Viro. Non-diffeomorphic but homeomorphic knottings of surfaces in the 4-sphere. Lecture Notes in Mathematics, 1346:157–198, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082777/.

Finkelshtein:1986:FPE

[FKZ86] A. M. Finkel’shteĭn, V. Ya. Kreĭnovich, and R. R. Zapatrin. Fundamental physical equations uniquely determined by their symmetry groups. Lecture Notes in Mathematics, 1214:159–170, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075964/.

Fagnola:1986:RIM

[FL86a] F. Fagnola and G. Letta. Sur la représentation intégrale des martingales du processus de Poisson. (French) []. Lecture Notes in Mathematics, 1204:28–29, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075707.

Feinberg:1986:THB

[FL86b] Martin Feinberg and Richard Lavine. Thermodynamics and the Hahn–Banach theorem. Lecture Notes in Mathematics, 1174:115– 117, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076938/.

Fourati:1987:THC

[FL87] Sonia Fourati and Erik Lenglart. Tribus homogènes et commutation de projections. (French) []. Lecture Notes in Mathematics, 1247:276–288, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077642/.

Flandoli:1987:DBV

[Fla87] Franco Flandoli. Dirichlet boundary value problem and optimal control for a stochastic distributed parameter system. Lecture Notes in Mathematics, 1236:57–71, 1987. CODEN LN-MAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072883/.

Fleischer:1985:DKR

[Fle85a] W. Fleischer. Ein Diskrepanzbegriff Für Kompakte Räume. (German) []. Lecture Notes in Mathematics, 1114:14–15, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101639.

Fleming:1985:SCA

[Fle85b] Wendell H. Fleming. A stochastic control approach to some large deviations problems. Lecture Notes in Mathematics, 1119:52– 66, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074780/.

Futaki:1988:OCR

[FM88] Akito Futaki and Toshiki Mabuchi. An obstruction class and a representation of holomorphic automorphisms. Lecture Notes in Mathematics, 1339:127–141, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083052/.

Fitzgibbon:1989:PSS

[FM89] W. E. Fitzgibbon and J. J. Morgan. Positive solutions for semilinear elliptic systems. Lecture Notes in Mathematics, 1394:56– 67, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086752/.

Fukushima:1987:DFR

[FNT87] M. Fukushima, S. Nakao, and M. Takeda. On Dirichlet forms with random data — recurrence and homogenization. Lecture Notes in Mathematics, 1250:87–97, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077349/.

Favini:1986:DEB

[FO86] Angelo Favini and Enrico Obrecht, editors. Differential Equations in Banach Spaces: Proceedings of a Conference held in Bologna, July 2–5, 1985, volume 1223 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1223. URL http:

//link.springer.com/book/10.1007/BFb0099177;   http://
www.springerlink.com/content/978-3-540-47350-3.

Follmer:1986:TRW

[Föl86] H. Föllmer. Time reversal on Wiener space. Lecture Notes in Mathematics, 1158:119–129, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080212/.

Follmer:1988:RFD

[Föl88] Hans Föllmer. Random fields and diffusion processes. Lecture Notes in Mathematics, 1362:101–203, 1988. CODEN LN-MAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086180/.

Fomenko:1988:IPH

[Fom88a] A. T. Fomenko. Invariant portrait of Hamiltonian integrable in Liouville sense. Lecture Notes in Mathematics, 1346:57–76, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082771/.

Fomenko:1988:QGT

[Fom88b] A. T. Fomenko. Qualitative geometrical theory of integrable systems. Classification of isoenergetic surfaces and bifurcation of Liouville tori at the critical energy values. Lecture Notes in Mathematics, 1334:221–245, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080431/.

Ford:1985:TDL

[For85] G. W. Ford. Temperature-dependent lamb shift of a quantum oscillator. Lecture Notes in Mathematics, 1136:202–206, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074472/.

Formanek:1986:GRM

[For86] Edward Formanek. Generating the ring of matrix invariants. Lecture Notes in Mathematics, 1197:73–82, 1986. CODEN LN-MAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076314/.

Formanek:1987:IM

[For87] Edward Formanek. The invariants of n×n matrices. Lecture Notes in Mathematics, 1278:18–43, 1987. CODEN LNMAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078804/.

Ford:1988:QLE

[For88a] G. W. Ford. The quantum Langevin equation from the independent-oscillator model. Lecture Notes in Mathematics, 1303:103–106, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078057/.

Forsman:1988:ADH

[For88b] Klas Forsman. Atomic decompositions in Hardy spaces on bounded Lipschitz domains. Lecture Notes in Mathematics, 1302: 206–222, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078876/.

Fountain:1988:FRA

[Fou88] John Fountain. Free right h-adequate semigroups. Lecture Notes in Mathematics, 1320:97–120, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083424/.

Fleckinger-Pelle:1985:AES

[FP85] J. Fleckinger-Pelle. Asymptotics of eigenvalues for some ‘nondefinite’ elliptic problems. Lecture Notes in Mathematics, 1151: 148–156, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074723/.

Fasano:1986:NDP

[FP86] Antonio Fasano and Mario Primicerio, editors. Nonlinear Diffusion Problems: Lectures given at the 2nd 1985 Session of the Centro Internazionale Matermatico Estivo (C.I.M.E.) held at Montecatini Terme, Italy June 10–June 18, 1985, volume 1224 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17192-4 (print), 3-540-47352-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1224. URL http://link.springer.com/book/10.1007/
BFb0072686;   http://www.springerlink.com/content/978-
3-540-47352-7.

Fannes:1985:IDC

[FQ85] M. Fannes and J. Quaegebeur. Infinite divisibility and central limit theorems for completely positive mappings. Lecture Notes in Mathematics, 1136:177–201, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074471/.

Forg-Rob:1985:PTL

[FR85] Wolfgang Förg-Rob. The pilgerschritt transform in Lie algebras. Lecture Notes in Mathematics, 1163:59–71, 1985. CODEN LN-MAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076418/.

Falk:1986:RCF

[FR86] Richard S. Falk and Gerard R. Richter. Remarks on a continuous finite element scheme for hyperbolic equations. Lecture Notes in Mathematics, 1230:63–72, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072671/.

Franzoni:1986:AC

[Fra86] T. Franzoni. Abstract Γ-convergence. Lecture Notes in Mathematics, 1190:229–241, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076709/.

Franek:1989:SIO

[Fra89] Frantisek Franek. Saturated ideals obtained via restricted iterated collapse of huge cardinals. Lecture Notes in Mathematics, 1401:73–96, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097333/.

Freese:1985:CCC

[Fre85] Ralph Freese. Connected components of the covering relation in free lattices. Lecture Notes in Mathematics, 1149:82–93, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098456/.

Freiman:1987:WSI

[Fre87] Gregory A. Freiman. What is the structure of K if K + K is small? Lecture Notes in Mathematics, 1240:109–134, 1987. CO-DEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0072977/.

Frehse:1988:RDE

[Fre88] Jens Frehse. Remarks on diagonal elliptic systems. Lecture Notes in Mathematics, 1357:198–210, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082867/.

Frey:1989:LBS

[Fre89] Gerhard Frey. Links between solutions of A − B = C and elliptic curves. Lecture Notes in Mathematics, 1380:31–62, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086544/.

Friedman:1985:IPO

[Fri85a] Sy D. Friedman. An immune partition of the ordinals. Lecture Notes in Mathematics, 1141:141–147, 1985. CODEN LN-MAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076218/.

Frigerio:1985:CSQ

[Fri85b] Alberto Frigerio. Construction of stationary quantum Markov processes through quantum stochastic calculus. Lecture Notes in Mathematics, 1136:207–222, 1985. CODEN LNMAA2. ISBN 3-540-15661-5 (print), 3-540-39570-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074473/.

Friedman:1986:FBP

[Fri86] A. Friedman. Free boundary problems in fluid dynamics. Lecture Notes in Mathematics, 1192:17–22, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076048/.

Fried:1988:CC

[Fri88a] David Fried. Counting circles. Lecture Notes in Mathematics, 1342:196–215, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082832/.

Fried:1988:CRR

[Fri88b] David Fried. Cyclic resultants of reciprocal polynomials. Lecture Notes in Mathematics, 1345:124–128, 1988. CODEN LN-MAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081399/.

Frigerio:1988:QPP

[Fri88c] Alberto Frigerio. Quantum Poisson processes: Physical motivations and applications. Lecture Notes in Mathematics, 1303:107– 127, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078058/.

Frigerio:1989:SAQ

[Fri89] Alberto Frigerio. Some applications of quantum probability to stochastic differential equations in Hilbert space. Lecture Notes in Mathematics, 1390:77–90, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083938/.

Forg-Rob:1985:PIO

[FRN85] Wolfgang Förg-Rob and Norbert Netzer. Product-integration and one-parameter subgroups of linear Lie-groups. Lecture Notes in Mathematics, 1163:72–82, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076419/.

Frohlich:1985:TRL

[Frö85] A. Fröhlich. Tame representations, local principal orders and local Weil groups. Lecture Notes in Mathematics, 1142:126–149, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074797/.

Frolicher:1986:CCC

[Frö86] Alfred Frölicher. Cartesian closed categories and analysis of smooth maps. Lecture Notes in Mathematics, 1174:43–51, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076933/.

Fejer:1985:EEE

[FS85a] Peter A. Fejer and Richard A. Shore. Embeddings and extensions of embeddings in the r.e. tt and wtt-degrees. Lecture Notes in Mathematics, 1141:121–140, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076217/.

Frangos:1985:CDB

[FS85b] Nikos E. Frangos and Louis Sucheston. On convergence and demiconvergence of block martingales and submartingales. Lecture Notes in Mathematics, 1153:198–225, 1985. CODEN LN-MAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074951/.

Fornaess:1987:PFR

[FS87] John Erik Fornæss and Nessim Sibony. Plurisubharmonic functions on ring domains. Lecture Notes in Mathematics, 1268:111– 120, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097300/.

Flaminio:1988:RRT

[FS88a] L. Flaminio and R. J. Spatzier. Ratner’s rigidity theorem for geometrically finite Fuchsian groups. Lecture Notes in Mathematics, 1342:180–195, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082831/.

Freiman:1988:GST

[FS88b] Gregory A. Freiman and Boris M. Schein. Group and semigroup theoretic considerations inspired by inverse problems of the additive number theory. Lecture Notes in Mathematics, 1320:121– 140, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083425/.

Feinsilver:1989:OAP

[FS89a] Philip Feinsilver and René Schott. An operator approach to processes on Lie groups. Lecture Notes in Mathematics, 1391:59– 65, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083380/.

Feinsilver:1989:OSP

[FS89b] Philip Feinsilver and René Schott. Operators, stochastic processes, and Lie groups. Lecture Notes in Mathematics, 1379:75–85, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087846/.

Felix:1986:OHR

[FT86a] Y. Félix and J.-C. Thomas. Sur l’opération d’holonômte rationnelle. (French) []. Lecture Notes in Mathematics, 1183:136– 169, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075456/.

Ferguson:1986:PC

[FT86b] Pamela A. Ferguson and Alexandre Turull. Prime characters. Lecture Notes in Mathematics, 1178:64–93, 1986. CODEN LN-MAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075289/.

Feigin:1987:AT

[FT87a] B. L. Feĭgin and B. L. Tsygan. Additive K-theory. Lecture Notes in Mathematics, 1289:67–209, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078368/.

Feigin:1987:CHA

[FT87b] B. L. Feĭgin and B. L. Tsygan. Cyclic homology of algebras with quadratic relations, universal enveloping algebras and group algebras. Lecture Notes in Mathematics, 1289:210–239, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078369/.

Felix:1988:EPC

[FT88a] Y. Félix and J. C. Thomas. An Euler–Poincaré characteristic for 1-connected spaces with Noetherian rational cohomology. Lecture Notes in Mathematics, 1318:124–137, 1988. CODEN LN-MAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077799/.

Felix:1988:FAS

[FT88b] Yves Félix and Daniel Tanré. Formalité d’une application et suite spectrale d’Eilenberg–Moore. (French) []. Lecture Notes in Mathematics, 1318:99–123, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077798/.

Figari:1987:CLT

[FTO87] R. Figari, S. Teta, and E. Orlandi. A central limit theorem for the Laplacian in regions with many small holes. Lecture Notes in Mathematics, 1250:75–86, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077348/.

Fu:1988:IBP

[Fu88] Hong-Yuan Fu. Initial and boundary problems for the degenerate or singular system of the filtration type. Lecture Notes in Mathematics, 1306:69–83, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082925/.

Fuchs:1986:AMG

[Fuc86] László Fuchs. An adaptive multi-grid scheme for simulation of flows. Lecture Notes in Mathematics, 1228:122–134, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072644/.

Fuglede:1986:FFH

[Fug86] Bent Fuglede. Fonctions finement holomorphes de plusieures variables — un essai. (French) []. Lecture Notes in Mathematics, 1198:133–145, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077049/.

Fuglede:1988:FPT

[Fug88a] Bent Fuglede. Fine potential theory. Lecture Notes in Mathematics, 1344:81–97, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103345/.

Fuglede:1988:RRF

[Fug88b] Bent Fuglede. On the Riesz representation of finely superharmonic functions. Lecture Notes in Mathematics, 1344:199–201, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103353/.

Fujiwara:1987:RTK

[Fuj87] Daisuke Fujiwara. A remark on Taniguchi–Kumanogo theorem for product of Fourier integral operators. Lecture Notes in Mathematics, 1256:135–153, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077741/.

Fukaya:1986:CSR

[Fuk86a] Kenji Fukaya. On a compactification of the set of Riemannian manifolds with bounded curvatures and diameters. Lecture Notes in Mathematics, 1201:89–107, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075648/.

Fukushima:1986:SAM

[Fuk86b] M. Fukushima. A stochastic approach to the minimum principle for the complex Monge–Ampère operator. Lecture Notes in Mathematics, 1203:38–50, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076871/.

Fuka:1988:CUF

[Fuk88a] Jaroslav Fuka. On a class of univalent functions. Lecture Notes in Mathematics, 1351:124–139, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081248/.

Fukushima:1988:TCS

[Fuk88b] Masatoshi Fukushima. On two classes of smooth measures for symmetric Markov processes. Lecture Notes in Mathematics, 1322:

17–27, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077863/.

Funaki:1986:CSP

[Fun86] Tadahisa Funaki. Construction of stochastic processes associated with the Boltzmann equation and its applications. Lecture Notes in Mathematics, 1203:51–65, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076872/.

Funaki:1988:HLS

[Fun88] Tadahisa Funaki. The hydrodynamical limit for scalar Ginzburg– Landau model on R. Lecture Notes in Mathematics, 1322:28– 36, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077864/.

Futaki:1988:AAR

[Fut88a] Akito Futaki. Aubin’s approach and related results. Lecture Notes in Mathematics, 1314:113–132, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078093/.

Futaki:1988:CGK

[Fut88b] Akito Futaki. The character f and its generalization to Kählerian invariants. Lecture Notes in Mathematics, 1314:46–55, 1988. CO-DEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078088/.

Futaki:1988:CCI

[Fut88c] Akito Futaki. The character f as a classical invariant. Lecture Notes in Mathematics, 1314:68–86, 1988. CODEN LN-MAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078090/.

Futaki:1988:CMM

[Fut88d] Akito Futaki. The character f as a moment map. Lecture Notes in Mathematics, 1314:99–112, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078092/.

Futaki:1988:CO

[Fut88e] Akito Futaki. The character f as an obstruction. Lecture Notes in Mathematics, 1314:56–67, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078089/.

Futaki:1988:I

[Fut88f] Akito Futaki. Introduction. Lecture Notes in Mathematics, 1314: 1–6, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078085/.

Futaki:1988:KEM

[Fut88g] Akito Futaki. Kähler–Einstein metrics and extremal Kähler metrics. Lecture Notes in Mathematics, 1314:31–45, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078087/.

Futaki:1988:LGC

[Fut88h] Akito Futaki. Lifting f to a group character. Lecture Notes in Mathematics, 1314:87–98, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078091/.

Futaki:1988:P

[Fut88i] Akito Futaki. Preliminaries. Lecture Notes in Mathematics, 1314: 7–30, 1988. CODEN LNMAA2. ISBN 3-540-19250-6 (print), 3-540-39172-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078086/.

Fraley:1989:NSP

[FV89] C. Fraley and J.-P. Vial. Numerical study of projective methods for linear programming. Lecture Notes in Mathematics, 1405:25– 38, 1989. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Fournier:1989:SVM

[FW89] G. Fournier and M. Willem. Simple variational methods for unbounded potentials. Lecture Notes in Mathematics, 1411:75– 82, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086442/.

Floer:1985:FPR

[FZ85] A. Floer and E. Zehnder. Fixed point results for symplectic maps related to the Arnold-conjecture. Lecture Notes in Mathematics, 1125:47–63, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075634/.

Farjoun:1987:HSS

[FZ87] E. Dror Farjoun and A. Zabrodsky. The homotopy spectral sequence for equivariant function complexes. Lecture Notes in Mathematics, 1298:54–81, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083000/.

Goerisch:1986:CNM

[GA86] F. Goerisch and J. Albrecht. The convergence of a new method for calculating lower bounds to eigenvalues. Lecture Notes in Mathematics, 1192:303–308, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076085/.

Gabai:1985:IHM

[Gab85] David Gabai. An internal hierarchy for 3-manifolds. Lecture Notes in Mathematics, 1144:14–17, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075009/.

Gago:1988:MGD

[Gag88] Felipe Gago. Morse germs in S.D.G. Lecture Notes in Mathematics, 1348:125–129, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081354/.

Gainza:1988:DCP

[Gai88a] P. Pascual Gainza. Descente cubique pour la K-théorie des faisceaux cohérents et l’homologie de Chow. (French) []. Lecture Notes in Mathematics, 1335:161–188, 1988. CODEN LNMAA2. ISBN 3-540-50023-5 (print), 3-540-69984-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085060/.

Gainza:1988:TMF

[Gai88b] P. Pascual Gainza. Théorèmes sur la monodromie. (French) []. Lecture Notes in Mathematics, 1335:43–58, 1988. CODEN LN-MAA2. ISBN 3-540-50023-5 (print), 3-540-69984-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085056/.

Gajewski:1986:USS

[Gaj86] H. Gajewski. On uniqueness and stability of steady-state carrier distributions in semiconductors. Lecture Notes in Mathematics, 1192:209–214, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076071/.

Galligo:1985:CSH

[Gal85] André Galligo. Computations of some Hilbert functions related with Schubert calculus. Lecture Notes in Mathematics, 1124:79– 97, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074997/.

Gallardo:1986:EHT

[Gal86] Léonard Gallardo. Exemples d’hypergroupes transients. (French) [Examples of transient hypergroups]. Lecture Notes in Mathematics, 1210:68–76, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077173/.

Gallardo:1989:RAG

[Gal89] Léonard Gallardo. Rectification à l’article de L. Gallardo et O. Gebuhrer “Lois de probabilités infiniment divisibles sur les hypergroupes commutatifs discrets dénombrables (Probability Measures on Groups VII Lecture Notes 1064). (French) []. Lecture Notes in Mathematics, 1379:86–89, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087847/. See [GG84].

Gang:1985:ACB

[Gan85a] Xiao Gang. Applications canoniques et bicanoniques. (French) []. Lecture Notes in Mathematics, 1137:71–88, 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075356/.

Gang:1985:CFA

[Gan85b] Xiao Gang. Classification des fibrations avec q = b + 1. (French) []. Lecture Notes in Mathematics, 1137:34–59, 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075354/.

Gang:1985:FAP

[Gan85c] Xiao Gang. Les fibrations avec petits invariants numériques. (French) []. Lecture Notes in Mathematics, 1137:60–70, 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075355/.

Gang:1985:INF

[Gan85d] Xiao Gang. Les invariants numériques. (French) []. Lecture Notes in Mathematics, 1137:16–33, 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075353/.

Gang:1985:NFG

[Gan85e] Xiao Gang. Nombre de fibrations de genre 2 d’une surface. (French) []. Lecture Notes in Mathematics, 1137:89–96, 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075357/.

Gang:1985:PF

[Gan85f] Xiao Gang. Préliminaires. (French) []. Lecture Notes in Mathematics, 1137:1–15, 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075352/.

Gang:1985:SFC

[Gan85g] Xiao Gang. Surfaces fibrées en courbes de genre deux, volume 1137 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15662-3 (print), 3-540-39573-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 103 pp. LCCN QA3 .L28 no. 1137. URL http:

//link.springer.com/book/10.1007/BFb0075351;   http://
www.springerlink.com/content/978-3-540-39573-7.

Garling:1986:BMU

[Gar86a] D. J. H. Garling. Brownian motion and UMD-spaces. Lecture Notes in Mathematics, 1221:36–49, 1986. CODEN LN-MAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099110/.

Garsia:1986:ROY

[Gar86b] A. M. Garsia. Raising operators and Young’s rule. Lecture Notes in Mathematics, 1234:91–105, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072511/.

Gaarding:1987:BM

[Går87a] Lars Gårding. Back matter. Lecture Notes in Mathematics, 1241: ??, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47216-2/1.pdf.

Gaarding:1987:CVD

[Går87b] Lars Gårding. Changes of variables and duality for general oscillatory integrals. Lecture Notes in Mathematics, 1241:95–103, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073095/.

Gaarding:1987:FM

[Går87c] Lars Gårding. Front matter. Lecture Notes in Mathematics, 1241: ??, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-47216-2/1.pdf.

Gaarding:1987:GPF

[Går87d] Lars Gårding. A global parametrix for the fundamental solution of a first order hyperbolic pseudodifferential operator. Lecture Notes in Mathematics, 1241:85–94, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073094/.

Gaarding:1987:HJE

[Går87e] Lars Gårding. The Hamilton–Jacobi equation and symplectic geometry. Lecture Notes in Mathematics, 1241:73–84, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073093/.

Gaarding:1987:HOC

[Går87f] Lars Gårding. Hyperbolic operators with constant coefficients. Lecture Notes in Mathematics, 1241:10–33, 1987. CODEN LN-MAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073090/.

Gaarding:1987:PO

[Går87g] Lars Gårding. Pseudodifferential operators. Lecture Notes in Mathematics, 1241:51–72, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073092/.

Gaarding:1987:SDF

[Går87h] Lars Gårding. Sharp and diffuse fronts of paired oscillatory integrals. Lecture Notes in Mathematics, 1241:104–123, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073096/.

Gaarding:1987:SLWa

[Går87i] Lars Gårding. Singularities in linear wave propagation. Lecture Notes in Mathematics, 1241:1–9, 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073089/.

Gaarding:1987:SLWb

[Går87j] Lars Gårding. Singularities in Linear Wave Propagation, volume 1241 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 125 pp. LCCN QA3 .L28 no. 1241. URL http:

//link.springer.com/book/10.1007/BFb0073088;   http://
www.springerlink.com/content/978-3-540-47216-2.

Gaarding:1987:WFS

[Går87k] Lars Gårding. Wave front sets and oscillatory integrals. Lecture Notes in Mathematics, 1241:34–50, 1987. CODEN LN-MAA2. ISBN 3-540-18001-X (print), 3-540-47216-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073091/.

Garabedian:1988:MTV

[Gar88a] Paul R. Garabedian. On the mathematical theory of vortex sheets. Lecture Notes in Mathematics, 1340:61–69, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082885/.

Garofalo:1988:NEF

[Gar88b] Nicola Garofalo. New estimates of the fundamental solution and Wiener’s criterion for parabolic equations with variable coefficients. Lecture Notes in Mathematics, 1340:70–83, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082886/.

Garroni:1988:GFI

[Gar88c] Maria Giovanna Garroni. Green function and invariant density for an integro-differential operator. Lecture Notes in Mathematics, 1340:84–94, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082887/.

Gautschi:1985:SNA

[Gau85] Walter Gautschi. Some new applications of orthogonal polynomials. Lecture Notes in Mathematics, 1171:63–73, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076531/.

Gauthier:1987:UBT

[Gau87] Paul M. Gauthier. Uniform and better-than-uniform elliptic approximations on unbounded sets. Lecture Notes in Mathematics, 1275:149–157, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078349/.

Gouyou-Beauchamps:1986:CSD

[GB86] Dominique Gouyou-Beauchamps. Chemins sous-diagonaux et tableaux de Young. (French) []. Lecture Notes in Mathematics, 1234:112–125, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072513/.

Gu:1987:DGD

[GBB87] Chaohao Gu, Marcel Berger, and Robert L. Bryant, editors. Differential Geometry and Differential Equations: Proceedings of a Symposium, held in Shanghai, June 21 – July 6, 1985, volume 1255 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1255. URL http:

//link.springer.com/book/10.1007/BFb0077675;   http://
www.springerlink.com/content/978-3-540-47883-6.

Gear:1989:DOC

[Gea89] C. W. Gear. DAEs: ODEs with constraints and invariants. Lecture Notes in Mathematics, 1386:54–68, 1989. CODEN LN-MAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089231/.

Gebuhrer:1986:QPD

[Geb86] Olivier Gebuhrer. Quelques propriétés du noyau potentiel d’une Marche aléatoire sur les hypergroupes de type Kunze–Stein. (French) []. Lecture Notes in Mathematics, 1210:77–83, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077174/.

Geigle:1986:KDA

[Gei86] Werner Geigle. Krull dimension and Artin algebras. Lecture Notes in Mathematics, 1177:135–155, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075263/.

Gekeler:1986:C

[Gek86a] Ernst-Ulrich Gekeler. Complements. Lecture Notes in Mathematics, 1231:94–100, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072701/.

Gekeler:1986:DMC

[Gek86b] Ernst-Ulrich Gekeler. Drinfeld Modular Curves, volume 1231 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 107 pp. LCCN QA3 .L28 no. 1231. URL http:

//link.springer.com/book/10.1007/BFb0072692;   http://
www.springerlink.com/content/978-3-540-47386-2.

Gekeler:1986:DM

[Gek86c] Ernst-Ulrich Gekeler. Drinfeld modules. Lecture Notes in Mathematics, 1231:2–10, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072694/.

Gekeler:1986:DMR

[Gek86d] Ernst-Ulrich Gekeler. Drinfeld modules of rank 1. Lecture Notes in Mathematics, 1231:24–40, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072697/.

Gekeler:1986:EAC

[Gek86e] Ernst-Ulrich Gekeler. Expansions around cusps. Lecture Notes in Mathematics, 1231:58–78, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072699/.

Gekeler:1986:L

[Gek86f] Ernst-Ulrich Gekeler. Lattices. Lecture Notes in Mathematics, 1231:10–17, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072695/.

Gekeler:1986:MC

[Gek86g] Ernst-Ulrich Gekeler. Modular curves over C. Lecture Notes in Mathematics, 1231:40–58, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072698/.

Gekeler:1986:MFF

[Gek86h] Ernst-Ulrich Gekeler. Modular forms and functions. Lecture Notes in Mathematics, 1231:78–93, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072700/.

Gekeler:1986:N

[Gek86i] Ernst-Ulrich Gekeler. Notations. Lecture Notes in Mathematics, 1231:1–2, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072693.

Gekeler:1986:PZF

[Gek86j] Ernst-Ulrich Gekeler. Partial zeta functions. Lecture Notes in Mathematics, 1231:17–24, 1986. CODEN LNMAA2. ISBN 3-540-17201-7 (print), 3-540-47386-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072696/.

Gelman:1986:CCS

[Gel86] B. D. Gel’man. On certain classes of selections of many-valued mappings. Lecture Notes in Mathematics, 1214:63–84, 1986. CO-DEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075959/.

Gelman:1988:SSS

[Gel88] B. D. Gel’man. On the structure of the set of solutions for inclusions with multivalued operators. Lecture Notes in Mathematics, 1334:60–78, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080423/.

Gentili:1987:RCG

[Gen87] Graziano Gentili. Regular complex geodesics for the domain D_{n} = ((z_{1}, …, z_{n}) ∈ C^{n}: |z_{1}| + … + |z_{n}| < 1). Lecture Notes in Mathematics, 1277:35–45, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078244/.

Gerard:1986:COS

[Gér86a] R. Gérard. Une classe d’opèrateurs singuliers non linèaires à singularitè règulière. (French) []. Lecture Notes in Mathematics, 1198:146–162, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077050/.

Gerl:1986:SIR

[Ger86b] Peter Gerl. Sobolev inequalities and random walks. Lecture Notes in Mathematics, 1210:84–96, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077175/.

Gerl:1988:AGA

[Ger88a] Peter Gerl. Amenable groups and amenable graphs. Lecture Notes in Mathematics, 1359:181–190, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086598/.

Geronimo:1988:POR

[Ger88b] J. S. Geronimo. Polynomials orthogonal with respect to singular continuous measures. Lecture Notes in Mathematics, 1329: 32–45, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083352/.

Gershkovich:1988:NFD

[Ger88c] V. Ya. Gershkovich. On normal form of distribution jets. Lecture Notes in Mathematics, 1346:77–98, 1988. CODEN LN-MAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082772/.

Geroldinger:1989:FAI

[Ger89] Alfred Geroldinger. Factorizations of algebraic integers. Lecture Notes in Mathematics, 1380:63–74, 1989. CODEN LN-MAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086545/.

Gessel:1986:CTL

[Ges86a] Ira M. Gessel. Counting three-line latin rectangles. Lecture Notes in Mathematics, 1234:106–111, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072512/.

Gesztesy:1986:STO

[Ges86b] F. Gesztesy. Scattering theory for one-dimensional systems with nontrivial spatial asymptotics. Lecture Notes in Mathematics, 1218:93–122, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073045/.

Gali:1985:OCS

[GESEZ85] I. M. Gali, H. A. El-Saify, and S. A. El-Zahabi. Optimal control of a system governed by hyperbolic operator. Lecture Notes in Mathematics, 1151:157–163, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074724/.

Gustafsson:1987:FFB

[GF87] Bertil Gustafsson and Lars Ferm. Far field boundary conditions for steady state solutions to hyperbolic systems. Lecture Notes in Mathematics, 1270:238–252, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078332/.

Gallardo:1984:LPI

[GG84] Léonard Gallardo and Olivier Gebuhrer. Lois de probabilité infiniment divisibles sur les hypergroupes commutatifs, discrets, dénombrables. (French) []. Lecture Notes in Mathematics, 1064: 116–130, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/

BFb0073638/. See correction [Gal89].

Geymonat:1985:EEN

[GG85] G. Geymonat and P. Grisvard. Eigenfunction expansions for non self adjoint operators and separation of variables. Lecture Notes in Mathematics, 1121:123–136, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076267/.

GarciaHernandez:1986:HSE

[GG86] J. L. García Hernández and J. L. Gómez Pardo. Hereditary and semihereditary endomorphism rings. Lecture Notes in Mathematics, 1197:83–89, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076315/.

Gupta:1987:DSR

[GGL87] C. K. Gupta, N. D. Gupta, and F. Levin. On dimension subgroups relative to certain product ideals. Lecture Notes in Mathematics, 1281:31–35, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078688/.

Gine:1985:IDR

[GH85a] Evarist Giné and Marjorie G. Hahn. M-infinitely divisible random compact convex sets. Lecture Notes in Mathematics, 1153: 226–248, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074952/.

Gingold:1985:GSH

[GH85b] Harry Gingold and Po-Fang Hsieh. A global study of a Hamiltonian system with multi turning points. Lecture Notes in Mathematics, 1151:164–171, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074725/.

Ghys:1988:CAF

[Ghy88] Étienne Ghys. Codimension on Anosov flows and suspensions. Lecture Notes in Mathematics, 1331:59–72, 1988. CODEN LN-MAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083066/.

Ganter:1985:VLS

[GI85] Bernhard Ganter and Thomas Ihringer. Varieties with linear subalgebra geometries. Lecture Notes in Mathematics, 1149:94– 100, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098457/.

Giaquinta:1986:PRM

[Gia86] M. Giaquinta. Partial regularity of minimizers. Lecture Notes in Mathematics, 1192:215–219, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076072/.

Giambruno:1987:PII

[Gia87] A. Giambruno. Polynomial identities with involution and the hyperoctahedral group. Lecture Notes in Mathematics, 1281:18– 25, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078686/.

Giaquinta:1988:QGC

[Gia88a] Mariano Giaquinta. Quasiconvexity, growth conditions and partial regularity. Lecture Notes in Mathematics, 1357:211–237, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082868/.

Giaquinta:1988:SRR

[Gia88b] Mariano Giaquinta. Some remarks on the regularity of minimizers. Lecture Notes in Mathematics, 1340:95–103, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082888/.

Gilbert:1987:FGC

[Gil87] N. D. Gilbert. On the fundamental cat^{n}-group of an n-cube of spaces. Lecture Notes in Mathematics, 1298:124–139, 1987. CO-DEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0083005/.

Gilman:1988:PBL

[Gil88] Robert H. Gilman. Periodic behavior of linear automata. Lecture Notes in Mathematics, 1342:216–219, 1988. CODEN LN-MAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082833/.

Gilkey:1989:EIE

[Gil89a] Peter B. Gilkey. The eta invariant of even order operators. Lecture Notes in Mathematics, 1410:202–211, 1989. CODEN LN-MAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086423/.

Gillard:1989:EFM

[Gil89b] Roland Gillard. Étude d’une famille modulaire de variétés abéliennes. (French) []. Lecture Notes in Mathematics, 1380:75– 86, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086546/.

Giusti:1985:HMM

[Giu85] Enrico Giusti, editor. Harmonic Mappings and Minimal Immersions: Lectures given at the 1st 1984 Session of the Centro Internationale Matematico Estivo (C.I.M.E.) held at Montecatini, Italy, June 24–July 3, 1984, volume 1161 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-16040-X (print), 3-540-39716-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1161. URL http://link.springer.com/book/10.1007/
BFb0075135;   http://www.springerlink.com/content/978-
3-540-39716-8.

Giusti:1988:QFS

[Giu88] Enrico Giusti. Quadratic functionals with splitting coefficients. Lecture Notes in Mathematics, 1340:104–114, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082889/.

Guddat:1988:POP

[GJN88] Jürgen Guddat, Hubertus Th. Jongen, and Dieter Nowack. Parametric optimization: Pathfollowing with jumps. Lecture Notes in Mathematics, 1354:43–53, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089582/.

Greene:1987:BSM

[GK87a] Robert E. Greene and Steven G. Krantz. Biholomorphic self-maps of domains. Lecture Notes in Mathematics, 1276:136–207, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078959/.

Greene:1987:CCW

[GK87b] Robert E. Greene and Steven G. Krantz. Characterizations of certain weakly pseudoconvex domains with non-compact automorphism groups. Lecture Notes in Mathematics, 1268:121–157, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097301/.

Goralcik:1988:UE

[GK88a] P. Goralčík and V. Koubek. On universality of extensions. Lecture Notes in Mathematics, 1320:150–153, 1988. CODEN LN-MAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083427/.

Gutkin:1988:WMB

[GK88b] E. Gutkin and A. Katok. Weakly mixing billiards. Lecture Notes in Mathematics, 1345:163–176, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081401/.

Goecke:1989:EAP

[GKL89] O. Goecke, B. Korte, and L. Lovász. Examples and algorithmic properties of greedoids. Lecture Notes in Mathematics, 1403: 113–161, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083463/.

Grace:1985:SFR

[GL85] S. R. Grace and B. S. Lalli. Some further results on oscillatory behavior of solutions of nth order delay differential equations. Lecture Notes in Mathematics, 1151:172–182, 1985. CODEN LN-MAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074726/.

Geigle:1987:CWP

[GL87] Werner Geigle and Helmut Lenzing. A class of weighted projective curves arising in representation theory of finite dimensional algebras. Lecture Notes in Mathematics, 1273:265–297, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078849/.

Glass:1985:WIP

[Gla85] A. M. W. Glass. The word and isomorphism problems in universal algebra. Lecture Notes in Mathematics, 1149:123–128, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098459/.

Glaz:1989:CA

[Gla89a] Sarah Glaz. Coherent algebras. Lecture Notes in Mathematics, 1371:280–327, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084578/.

Glaz:1989:FC

[Gla89b] Sarah Glaz. Fundamental concepts. Lecture Notes in Mathematics, 1371:69–107, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084573/.

Glaz:1989:IC

[Gla89c] Sarah Glaz. Introduction to coherence. Lecture Notes in Mathematics, 1371:31–68, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084572/.

Glaz:1989:PCR

[Gla89d] Sarah Glaz. Particular coherent rings. Lecture Notes in Mathematics, 1371:190–227, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084576/.

Glaz:1989:PR

[Gla89e] Sarah Glaz. Polynomial rings. Lecture Notes in Mathematics, 1371:228–279, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084577/.

Glaz:1989:P

[Gla89f] Sarah Glaz. Preliminaries. Lecture Notes in Mathematics, 1371: 1–30, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084571/.

Glaz:1989:RCO

[Gla89g] Sarah Glaz. Ring constructions and overrings. Lecture Notes in Mathematics, 1371:149–189, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084575/.

Glaz:1989:RE

[Gla89h] Sarah Glaz. Ring extensions. Lecture Notes in Mathematics, 1371:108–148, 1989. CODEN LNMAA2. ISBN 3-540-51115-6 (print), 3-540-46159-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084574/.

Gomez:1986:GZR

[GLC86] S. Gömez, A. V. Lévy, and A. Calderon. A global zero residual least squares method. Lecture Notes in Mathematics, 1230:

1–10, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072667/.

Gliklikh:1986:RPT

[Gli86] Yu. E. Gliklikh. Riemannian parallel translation, the Itô integral, and stochastic equations on manifolds. Lecture Notes in Mathematics, 1214:171–195, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075965/.

Gliklikh:1988:GAN

[Gli88] Yu. E. Gliklikh. Geometrical aspects of Nelson’s stochastic quantization. Lecture Notes in Mathematics, 1334:246–268, 1988. CO-DEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080432/.

Garcia-Lazaro:1985:CFK

[GLM85] P. García-Lázaro and F. Marcellán. Christoffel formulas for nkernels associated to Jordan arcs. Lecture Notes in Mathematics, 1171:195–203, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076544/.

Glover:1987:CMS

[Glo87] Henry H. Glover. Coloring maps on surfaces. Lecture Notes in Mathematics, 1298:140–147, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083006/.

Grochenig:1986:UDS

[GLR86] K. Gröchenig, Viktor Losert, and H. Rindler. Uniform distribution in solvable groups. Lecture Notes in Mathematics, 1210: 97–107, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077176/.

Ghoussoub:1985:RNP

[GM85a] N. Ghoussoub and B. Maurey. On the Radon–Nikodým property in function spaces. Lecture Notes in Mathematics, 1166:65– 71, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074693/.

Glover:1985:SCM

[GM85b] H. Glover and G. Mislin. On the stable cohomology of the mapping class group. Lecture Notes in Mathematics, 1172:80–84, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074425/.

Gil-Medrano:1986:CSI

[GM86] Olga Gil-Medrano. Connected sums and the infimum of the Yamabe functional. Lecture Notes in Mathematics, 1209:160–167, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076629/.

Gehring:1987:DCG

[GM87a] F. W. Gehring and G. J. Martin. Discrete convergence groups. Lecture Notes in Mathematics, 1275:158–167, 1987. CODEN LN-MAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078350/.

Goresky:1987:TAT

[GM87b] M. Goresky and R. MacPherson. On the topology of algebraic torus actions. Lecture Notes in Mathematics, 1271:73–90, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079233/.

Ghilardi:1988:MTP

[GM88] S. Ghilardi and G. C. Meloni. Modal and tense predicate logic: Models in presheaves and categorical conceptualization. Lecture Notes in Mathematics, 1348:130–142, 1988. CODEN LN-MAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081355/.

Gomez:1989:SEC

[GM89] F. Gómez and M. C. Muñoz. Submanifolds with Euclidean complements. Lecture Notes in Mathematics, 1410:240–248, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086426/.

Gomez-Mont:1988:PCH

[GMM88] Xavier Gómez-Mont and Jesús Muciño. Persistent cycles for holomorphic foliations having a meromorphic first integral. Lecture Notes in Mathematics, 1345:129–162, 1988. CODEN LN-MAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081400/.

Gill:1986:CNA

[GMSW86] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Considerations of numerical analysis in a sequential quadratic programming method. Lecture Notes in Mathematics, 1230:46–62, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072670/.

Graves-Morris:1987:FAS

[GMW87] P. R. Graves-Morris and J. M. Wilkins. A fast algorithm to solve Kalman’s partial realisation problem for single input, multi-output systems. Lecture Notes in Mathematics, 1287:1–20, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078896/.

Greiner:1986:ST

[GN86] Günther Greiner and Rainer Nagel. Spectral theory. Lecture Notes in Mathematics, 1184:60–97, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074925/.

Godula:1988:IMS

[GN88] Janusz Godula and Maria Nowak. Integral means of α-strongly starlike functions. Lecture Notes in Mathematics, 1351:140– 145, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081249/.

Grebogi:1988:BSS

[GNOY88] Celso Grebogi, Helena E. Nusse, Edward Ott, and James A. Yorke. Basic sets: Sets that determine the dimension of basin boundaries. Lecture Notes in Mathematics, 1342:220–250, 1988. CODEN LN-MAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082834/.

Goberstein:1988:CS

[Gob88] Simon M. Goberstein. Correspondences of semigroups. Lecture Notes in Mathematics, 1320:141–149, 1988. CODEN LN-MAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083426/.

Godin:1987:ARW

[God87a] Paul Godin. On the analytic regularity of weak solutions of analytic systems of conservation laws with analytic data. Lecture Notes in Mathematics, 1256:154–168, 1987. CODEN LN-MAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077742/.

Godounov:1987:LCI

[God87b] S. K. Godounov. Lois de conservation et intégrales d’énergie des équations hyperboliques. (French) []. Lecture Notes in Mathematics, 1270:135–149, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078323/.

Goldman:1985:RFG

[Gol85] William M. Goldman. Representations of fundamental groups of surfaces. Lecture Notes in Mathematics, 1167:95–117, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075218/.

Goldberg:1986:IWM

[Gol86a] V. V. Goldberg. Isoclinic webs W (4, 2, 2) of maximum 2-rank. Lecture Notes in Mathematics, 1209:168–183, 1986. CODEN LN-MAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076630/.

Goldfarb:1986:EPA

[Gol86b] Donald Goldfarb. Efficient primal algorithms for strictly convex quadratic programs. Lecture Notes in Mathematics, 1230: 11–25, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072668/.

Goldstein:1987:EEN

[Gol87a] Jerome A. Goldstein. Evolution equations with nonlinear boundary conditions. Lecture Notes in Mathematics, 1248:78–84, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077417/.

Goldstern:1987:KVG

[Gol87b] Martin Goldstern. Eine Klasse Vollständig Gleichverteilter Folgen. (German) []. Lecture Notes in Mathematics, 1262:37–45, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078594/.

Goldstern:1987:VGD

[Gol87c] Martin Goldstern. Vollständige Gleichverteilung in Diskreten Räumen. (German) []. Lecture Notes in Mathematics, 1262:46– 49, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078595/.

Golod:1988:SBH

[Gol88] E. S. Golod. Standard bases and homology. Lecture Notes in Mathematics, 1352:88–95, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082019/.

Goldberg:1989:LCT

[Gol89] Vladislav V. Goldberg. On a linearizability condition for a threeweb on a two-dimensional manifold. Lecture Notes in Mathematics, 1410:223–239, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086425/.

Gonzalez:1988:ALF

[Gon88] Jorge Bustamante González. Approximation by Lipschitz functions and its application to boundary value of Cauchy-type integrals. Lecture Notes in Mathematics, 1354:106–110, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089586/.

Gootman:1985:AGA

[Goo85] Elliot C. Gootman. Abelian group actions on type I C∗-algebras. Lecture Notes in Mathematics, 1132:152–169, 1985. CODEN LN-MAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074884/.

Gorostiza:1986:ASG

[Gor86] Luis G. Gorostiza. Asymptotics and spatial growth of branching random fields. Lecture Notes in Mathematics, 1212:124–135, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076243/.

Gordon:1988:MIR

[Gor88] Y. Gordon. On Milman’s inequality and random subspaces which escape through a mesh in R^{n}. Lecture Notes in Mathematics, 1317: 84–106, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081737/.

Gossez:1988:NNF

[Gos88] Jean-Pierre Gossez. Nonresonance near the first eigenvalue of a second order elliptic problem. Lecture Notes in Mathematics, 1324:97–104, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100785/.

Goss:1989:RTT

[Gos89] David Goss. Report on transcendency in the theory of function fields. Lecture Notes in Mathematics, 1383:59–63, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083569/.

Gotz:1986:GDB

[Göt86] Thomas Götz. Generation-dependent branching processes with immigration: convergence of distributions. Lecture Notes in Mathematics, 1212:136–144, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076244/.

Gottlieb:1988:IPV

[Got88] Daniel H. Gottlieb. On the index of pullback vector fields. Lecture Notes in Mathematics, 1350:167–170, 1988. CODEN LN-MAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081474/.

Gouvea:1988:GR

[Gou88a] Fernando Quadros Gouvêa. Galois representations. Lecture Notes in Mathematics, 1304:71–116, 1988. CODEN LNMAA2. ISBN 3-540-18946-7 (print), 3-540-38854-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082114/.

Gouvea:1988:HO

[Gou88b] Fernando Quadros Gouvêa. The Hecke and U operators. Lecture Notes in Mathematics, 1304:30–70, 1988. CODEN LN-MAA2. ISBN 3-540-18946-7 (print), 3-540-38854-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082113/.

Gouvea:1988:AMF

[Gou88c] Fernando Quadros Gouvêa. p-adic modular forms. Lecture Notes in Mathematics, 1304:1–29, 1988. CODEN LNMAA2. ISBN 3-540-18946-7 (print), 3-540-38854-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082112/.

Glowinski:1985:FEL

[GP85] R. Glowinski and J. Périaux. Finite element, least squares and domains decomposition methods for the numerical solution of nonlinear problems in fluid dynamics. Lecture Notes in Mathematics, 1127:1–114, 1985. CODEN LNMAA2. ISBN 3-540-15225-3 (print), 3-540-39391-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074529/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-39391-7/
1.pdf.

Gruson:1987:SCC

[GP87a] Laurent Gruson and Christian Peskine. Space curves: Complete series and speciality. Lecture Notes in Mathematics, 1266:108– 123, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078181/.

Guillen:1987:HCA

[GP87b] F. Guillén and F. Puerta. Hyperrésolutions cubiques et applications à la théorie de Hodge–Deligne. (French) []. Lecture Notes in Mathematics, 1246:49–74, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077529/.

Garcia:1987:DGM

[GPR87] Pedro Luis García and Antonio Pérez-Rendón, editors. Differential Geometric Methods in Mathematical Physics: Proceedings of the 14th International Conference held in Salamanca, Spain, June 24–29, 1985, volume 1251 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1251. URL http:

//link.springer.com/book/10.1007/BFb0077312;   http://
www.springerlink.com/content/978-3-540-47854-6.

Gelbart:1987:F

[GPS87a] S. Gelbart and I. Piatetski-Shapiro. L-functions for G × GL(n). Lecture Notes in Mathematics, 1254:53–136, 1987. CODEN LN-MAA2. ISBN 3-540-17848-1 (print), 3-540-47880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078127/.

Gilewicz:1987:RAA

[GPS87b] Jacek Gilewicz, Maciej Pindor, and Wojciech Siemaszko, editors. Rational Approximation and its Applications in Mathematics and Physics: Proceedings, Łańcut 1985, volume 1237 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LN-MAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1237. URL http://link.springer.com/book/10.1007/
BFb0072449;   http://www.springerlink.com/content/978-
3-540-47412-8.

Gelbart:1987:ECA

[GPSR87] Stephen Gelbart, Ilya Piatetski-Shapiro, and Stephen Rallis. Explicit Constructions of Automorphic L-Functions, volume 1254 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LN-MAA2. ISBN 3-540-17848-1 (print), 3-540-47880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 152 pp. LCCN QA3

.L28 no. 1254. URL http://link.springer.com/book/10.1007/
BFb0078125;   http://www.springerlink.com/content/978-
3-540-47880-5.

Goldberg:1989:CCR

[GPT89] S. I. Goldberg, D. Perrone, and G. Tóth. Curvature of contact Riemannian three-manifolds with critical metrics. Lecture Notes in Mathematics, 1410:212–222, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086424/.

Gomez:1985:TES

[GQS85] C. Gomez, J.-P. Quadrat, and A. Sulem. Towards an expert system in stochastic control: Optimization in the class of local feedbacks. Lecture Notes in Mathematics, 1119:67–93, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074781/.

Galdi:1985:EMUa

[GR85a] Giovanni P. Galdi and Salvatore Rionero. Energy methods in unbounded domains: The case of a half space. Lecture Notes in Mathematics, 1134:14–35, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075382/.

Galdi:1985:EMUb

[GR85b] Giovanni P. Galdi and Salvatore Rionero. Energy methods in unbounded domains: The case of an exterior domain. Lecture Notes in Mathematics, 1134:36–48, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075383/.

Galdi:1985:ESS

[GR85c] Giovanni P. Galdi and Salvatore Rionero. Exchange of stabilities, symmetry and the connection between linear and nonlinear stability. Lecture Notes in Mathematics, 1134:55–77, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075385/.

Galdi:1985:ITS

[GR85d] Giovanni P. Galdi and Salvatore Rionero. Introductory topics on stability of viscous flows. Lecture Notes in Mathematics, 1134: 1–13, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075381/.

Galdi:1985:SSO

[GR85e] Giovanni P. Galdi and Salvatore Rionero. Skew-symmetric operators and nonlinear stabilization: The hydromagnetic Bénard problem. Lecture Notes in Mathematics, 1134:78–91, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075386/.

Galdi:1985:SPS

[GR85f] Giovanni P. Galdi and Salvatore Rionero. Some properties of steady solutions in exterior domains. Lecture Notes in Mathematics, 1134:49–54, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075384/.

Galdi:1985:WEMa

[GR85g] Giovanni P. Galdi and Salvatore Rionero. Weighted energy methods in elastostatics on unbounded domains. Lecture Notes in Mathematics, 1134:92–107, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075387/.

Galdi:1985:WEMc

[GR85h] Giovanni P. Galdi and Salvatore Rionero. Weighted Energy Methods in Fluid Dynamics and Elasticity, volume 1134 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 126 pp. LCCN QA3

.L28 no. 1134. URL http://link.springer.com/book/10.1007/
BFb0075380;   http://www.springerlink.com/content/978-
3-540-39523-2.

Galdi:1985:WEMb

[GR85i] Giovanni P. Galdi and Salvatore Rionero. Weighted energy methods in linear elastodynamics in unbounded domains. Lecture Notes in Mathematics, 1134:108–119, 1985. CODEN LNMAA2. ISBN 3-540-15645-3 (print), 3-540-39523-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075388/.

Grupp:1985:FLS

[GR85j] F. Grupp and H.-E. Richert. The functions of the linear sieve (summary). Lecture Notes in Mathematics, 1122:105–114, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075754/.

Guadalupe:1985:CAP

[GR85k] J. J. Guadalupe and M. L. Rezola. Closure of analytic polynomials in weighted Jordan curves. Lecture Notes in Mathematics, 1171:204–210, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076545/.

Goldstein:1986:SET

[GR86] Jerome A. Goldstein and Gisèle Ruíz Rieder. Some extensions of Thomas–Fermi theory. Lecture Notes in Mathematics, 1223: 110–121, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099187/.

Goldstein:1987:CPH

[GR87a] Jerome A. Goldstein and Gisèle Ruíz Rieder. The Coulomb potential in higher dimensions. Lecture Notes in Mathematics, 1285: 143–149, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080590/.

Graham:1987:NMS

[GR87b] R. Graham and D. Roekaerts. A Nicolai map for supersymmetric quantum mechanics on Riemannian manifolds. Lecture Notes in Mathematics, 1250:98–105, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077350/.

Goncar:1988:RRA

[GR88] A. A. Gončar and E. A. Rakhmanov. On the rate of rational approximation of analytic functions. Lecture Notes in Mathematics, 1354:25–42, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089581/.

Goldstein:1989:RRR

[GR89] Jerome A. Goldstein and Gisèle Ruíz Rieder. Recent rigorous results in Thomas–Fermi theory. Lecture Notes in Mathematics, 1394:68–82, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086753/.

Graffi:1985:SOL

[Gra85a] Sandro Graffi, editor. Schrödinger Operators: Lectures given at the 2nd 1984 Session of the Centro Internationale Matematico Estivo (C.I.M.E.) held at Como, Italy, Aug. 26–Sept. 4, 1984, volume 1159 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1159. URL http:

//link.springer.com/book/10.1007/BFb0080330;   http://
www.springerlink.com/content/978-3-540-39706-9.

Gray:1985:CTV

[Gra85b] Alfred Gray. Comparison theorems for volumes in surfaces. Lecture Notes in Mathematics, 1167:118–124, 1985. CODEN LN-MAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075219/.

Graffi:1986:CLC

[Gra86a] Sandro Graffi. Classical limit and canonical perturbation theory. Lecture Notes in Mathematics, 1218:123–135, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073046/.

Grauert:1986:CAA

[Gra86b] Hans Grauert, editor. Complex Analysis and Algebraic Geometry: Proceedings of a Conference held in Göttingen, June 25– July 2, 1985, volume 1194 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1194. URL http:

//link.springer.com/book/10.1007/BFb0076989;   http://
www.springerlink.com/content/978-3-540-39829-5.

Graham:1987:SBI

[Gra87] C. Robin Graham. Scalar boundary invariants and the Bergman kernel. Lecture Notes in Mathematics, 1276:108–135, 1987. CO-DEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078958/.

Gray:1989:DS

[Gra89] Brayton Gray. On the double suspension. Lecture Notes in Mathematics, 1370:150–162, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085226/.

Gregus:1986:CBV

[Gre86a] M. Greguš. On a certain boundary value problem of the third order. Lecture Notes in Mathematics, 1192:129–132, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076059/.

Greiner:1986:APSa

[Gre86b] Günther Greiner. Asymptotics of positive semigroups on C_{o}(X). Lecture Notes in Mathematics, 1184:204–232, 1986. CODEN LN-MAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074930/.

Greiner:1986:APSb

[Gre86c] Günther Greiner. Asymptotics of positive semigroups on Banach lattices. Lecture Notes in Mathematics, 1184:333–367, 1986. CO-DEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074934/.

Greiner:1986:STPa

[Gre86d] Günther Greiner. Spectral theory of positive semigroups on C_{o}(X). Lecture Notes in Mathematics, 1184:163–203, 1986. CO-DEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074929/.

Greiner:1986:STPb

[Gre86e] Günther Greiner. Spectral theory of positive semigroups on Banach lattices. Lecture Notes in Mathematics, 1184:292–332, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074933/.

Greven:1986:CIP

[Gre86f] Andreas Greven. On a class of infinite particle systems evolving in a random environment. Lecture Notes in Mathematics, 1212:145–164, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076245/.

Greene:1987:CDG

[Gre87] Robert E. Greene. Complex differential geometry. Lecture Notes in Mathematics, 1263:228–288, 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078614/.

Green:1989:RLS

[Gre89] Mark Green. Restrictions of linear series to hyperplanes, and some results of Macaulay and Gotzmann. Lecture Notes in Mathematics, 1389:76–86, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085925/.

Griffor:1985:ALD

[Gri85a] E. R. Griffor. An application of Π _{2}^{1}-logic to descriptive set theory. Lecture Notes in Mathematics, 1141:148–158, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076219/.

Grindrod:1985:MMN

[Gri85b] P. Grindrod. A model for a myelinated nerve axon. Lecture Notes in Mathematics, 1151:183–191, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074727/.

Griewank:1986:RCS

[Gri86a] Andreas Griewank. Rates of convergence for secant methods on nonlinear problems in Hilbert space. Lecture Notes in Mathematics, 1230:138–157, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072677/.

Grimmett:1986:PPD

[Gri86b] Geoffrey Grimmett. Percolation processes and dimensionality. Lecture Notes in Mathematics, 1212:165–174, 1986. CODEN LN-MAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076246/.

Grisvard:1986:ASS

[Gri86c] P. Grisvard. An approach to the singular solutions of elliptic problems via the theory of differential equations in Banach spaces. Lecture Notes in Mathematics, 1223:131–155, 1986. CODEN LN-MAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099189/.

Grove:1985:IIG

[Gro85] Karsten Grove. The isometry-invariant geodesics problem: Closed and open. Lecture Notes in Mathematics, 1167:125–140, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075220/.

Groetsch:1986:RLE

[Gro86a] C. W. Groetsch. Regularization with linear equality constraints. Lecture Notes in Mathematics, 1225:168–181, 1986. CODEN LN-MAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072663/.

Groh:1986:APS

[Gro86b] Ulrich Groh. Asymptotics of positive semigroups on c^{∗}-and w^{∗}- algebras. Lecture Notes in Mathematics, 1184:400–425, 1986. CO-DEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074938/.

Groh:1986:BRS

[Gro86c] Ulrich Groh. Basic results on semigroups and operator algebras. Lecture Notes in Mathematics, 1184:369–375, 1986. CODEN LN-MAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074935/.

Groh:1986:CPS

[Gro86d] Ulrich Groh. Characterization of positive semigroups on w^{∗}- algebras. Lecture Notes in Mathematics, 1184:376–378, 1986. CO-DEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074936/.

Groh:1986:STP

[Gro86e] Ulrich Groh. Spectral theory of positive semigroups on w^{∗}- algebras and their preduals. Lecture Notes in Mathematics, 1184: 379–399, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074937/.

Gromov:1986:LRM

[Gro86f] M. Gromov. Large Riemannian manifolds. Lecture Notes in Mathematics, 1201:108–121, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075649/.

Gross:1986:LGT

[Gro86g] Lconard Gross. Lattice gauge theory; heuristics and convergence. Lecture Notes in Mathematics, 1158:130–140, 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080213/.

Gromov:1987:MVI

[Gro87a] M. Gromov. Monotonicity of the volume of intersection of balls. Lecture Notes in Mathematics, 1267:1–4, 1987. CODEN LN-MAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078131/.

Gross:1987:AIE

[Gro87b] Fletcher Gross. Automorphisms of induced extensions. Lecture Notes in Mathematics, 1281:26–30, 1987. CODEN LN-MAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078687/.

Grosshans:1987:CIP

[Gro87c] Frank D. Grosshans. Constructing invariant polynomials via Tschirnhaus transformations. Lecture Notes in Mathematics, 1278:95–102, 1987. CODEN LNMAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078809/.

Grove:1987:MDG

[Gro87d] Karsten Grove. Metric differential geometri. Lecture Notes in Mathematics, 1263:171–227, 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078613/.

Gromov:1988:DNL

[Gro88] M. Gromov. Dimension, non-linear spectra and width. Lecture Notes in Mathematics, 1317:132–184, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081739/.

Guillot:1988:SCA

[GRT88] J.-C. Guillot, J. Ralston, and E. Trubowitz. Semi-classical approximations in solid state physics. Lecture Notes in Mathematics, 1324:263–269, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100798/.

Grubb:1986:TEE

[Gru86a] Gerd Grubb. Trace estimates for exterior boundary problems associated with the Schrödinger operator. Lecture Notes in Mathematics, 1218:136–150, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073047/.

Grunbaum:1986:SMP

[Grü86b] F. Alberto Grünbaum. Some mathematical problems motivated by medical imaging. Lecture Notes in Mathematics, 1225:113– 141, 1986. CODEN LNMAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072661/.

Grubb:1987:CPP

[Gru87] Gerd Grubb. Complex powers of pseudo-differential boundary value problems with the transmission property. Lecture Notes in Mathematics, 1256:169–191, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077743/.

Grunbaum:1988:DES

[Grü88a] F. Alberto Grünbaum. Differential equations in the spectral parameter and multiphase similarity solutions. Lecture Notes in Mathematics, 1324:105–117, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100786/.

Gruter:1988:MFG

[Grü88b] Michael Grüter. The monotonicity formula in geometric measure theory, and an application to a partially free boundary problem. Lecture Notes in Mathematics, 1357:238–255, 1988. CODEN LN-MAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082869/.

Gardner:1985:EOD

[GS85a] Robert B. Gardner and William F. Shadwick. Equivalence of one dimensional Lagrangian field theories in the plane I. Lecture Notes in Mathematics, 1156:154–179, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075089/.

Guo:1985:SPS

[GS85b] Ben Yu Guo and B. D. Sleeman. Spatial patterning of the spruce budworm in the presence of defoliation. Lecture Notes in Mathematics, 1151:192–203, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074728/.

Grimmer:1986:ESR

[GS86] Ronald Grimmer and Eugenio Sinestrari. The extent of spatial regularity for parabolic integrodifferential equations. Lecture Notes in Mathematics, 1223:122–130, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099188/.

Ghione:1987:GCL

[GS87] Franco Ghione and Gianni Sacchiero. Genre d’une courbe lisse tracée sur une variété réglée. (French) []. Lecture Notes in Mathematics, 1266:97–107, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078180/.

Gordon:1989:MPP

[GS89] Basil Gordon and Dale Sinor. Multiplicative properties of ηproducts. Lecture Notes in Mathematics, 1395:173–200, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086404/.

Goerss:1987:AIF

[GSZ87] P. Goerss, L. Smith, and S. Zarati. Sur les A-algèbres instables. (French) []. Lecture Notes in Mathematics, 1298:148–161, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083007/.

Garcia:1985:GC

[GT85a] Octavio C. García and Walter Taylor. Generalized commutativity. Lecture Notes in Mathematics, 1149:101–122, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098458/.

Gottlieb:1985:TSM

[GT85b] David Gottlieb and Eli Turkel. Topics in spectral methods. Lecture Notes in Mathematics, 1127:115–155, 1985. CODEN LN-MAA2. ISBN 3-540-15225-3 (print), 3-540-39391-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074530/.

Gu:1987:GSE

[Gu87a] Chaohao Gu. A global study of extremal surfaces in 3-dimensional Minkowski space. Lecture Notes in Mathematics, 1255:26–33, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077678/.

Guillemin:1987:SSP

[GU87b] V. Guillemin and A. Uribe. Some spectral properties of periodic potentials. Lecture Notes in Mathematics, 1256:192–213, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077744/.

Gudmundsdottir:1987:RTD

[Gud87] Gudrun Gudmundsdottir. A regularity theorem for differential equations of constant strength. Lecture Notes in Mathematics, 1285:150–154, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080591/.

Gudkov:1988:PRP

[Gud88] D. A. Gudkov. Plane real projective quartic curves. Lecture Notes in Mathematics, 1346:341–347, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082782/.

Guerre:1985:SS

[Gue85] Sylvie Guerre. Symmetric sequences in L^{p}, l ≤ p < +∞. Lecture Notes in Mathematics, 1166:72–73, 1985. CODEN LN-MAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074694.

Guerre:1986:AES

[Gue86] Sylvie Guerre. Almost exchangeable sequences in L^{q}, 1 ≤ q < 2. Lecture Notes in Mathematics, 1193:22–28, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077097/.

Guillen:1988:DCC

[Gui88a] F. Guillén. Descente cubique de la cohomologie de de Rham algébrique. (French) []. Lecture Notes in Mathematics, 1335:59– 86, 1988. CODEN LNMAA2. ISBN 3-540-50023-5 (print), 3-540-69984-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085057/.

Guillen:1988:HCF

[Gui88b] F. Guillén. Hyperrésolutions cubiques. (French) [Cubic hyperresolutions]. Lecture Notes in Mathematics, 1335:1–42, 1988. CODEN LNMAA2. ISBN 3-540-50023-5 (print), 3-540-69984-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085055/.

Gulliksen:1986:NIM

[Gul86a] Tor H. Gulliksen. A note on intersection multiplicities. Lecture Notes in Mathematics, 1183:192–194, 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075458/.

Gulliksen:1986:RPS

[Gul86b] Tor H. Gulliksen. Reducing the Poincaré series of local rings to the case of quadratic relations. Lecture Notes in Mathematics, 1183:195–198, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075459/.

Gulliver:1988:IPH

[Gul88a] Robert Gulliver. Isoperimetric problems having continua of solutions. Lecture Notes in Mathematics, 1357:256–263, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082870/.

Gulliver:1988:MSF

[Gul88b] Robert Gulliver. Minimal surfaces of finite index in manifolds of positive scalar curvature. Lecture Notes in Mathematics, 1340: 115–122, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082890/.

Gupta:1989:ACR

[Gup89a] C. Kanta Gupta. Automorphisms of certain relatively free solvable groups. Lecture Notes in Mathematics, 1398:82–85, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086242/.

Gupta:1989:HDS

[Gup89b] Narain Gupta. Higher dimension subgroups. Lecture Notes in Mathematics, 1398:86–93, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086243/.

Guralnick:1985:MUG

[Gur85] Robert M. Guralnick. Modules under ground ring extension. Lecture Notes in Mathematics, 1142:150–156, 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074798/.

Guralnick:1986:DFC

[Gur86] Robert M. Guralnick. The dimension of the first cohomology group. Lecture Notes in Mathematics, 1178:94–97, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075290/.

Gustafson:1987:IBV

[Gus87] Karl Gustafson. Intermittent bifurcation of vortex flows. Lecture Notes in Mathematics, 1285:155–161, 1987. CODEN LN-MAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080592/.

Gutzwiller:1987:MSC

[Gut87] Martin C. Gutzwiller. Mechanics on a surface of constant negative curvature. Lecture Notes in Mathematics, 1240:230–258, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072983/.

Gutierrez:1988:FSH

[Gut88] Carlos Gutiérrez. Foliations on surfaces having exceptional leaves. Lecture Notes in Mathematics, 1331:73–85, 1988. CODEN LN-MAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083067/.

Gomez:1987:CEG

[GV87] J. Gómez and F. Varela. Sur certaines expressions globales d’une forme volume. (French) []. Lecture Notes in Mathematics, 1251: 284–300, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077326/.

Glockner:1989:RNC

[GvW89] P. Glockner and W. von Waldenfels. The relations of the non-commutative coefficient algebra of the unitary group. Lecture Notes in Mathematics, 1396:182–220, 1989. CODEN LN-MAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083553/.

Grisvard:1985:SCM

[GWW85] Pierre Grisvard, Wolfgang L. Wendland, and John R. Whiteman, editors. Singularities and Constructive Methods for Their Treatment: Proceedings of the Conference held in Oberwolfach, West Germany, November 20–26, 1983, volume 1121 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1121. URL http://link.springer.com/book/10.1007/
BFb0076257;   http://www.springerlink.com/content/978-
3-540-39377-1.

Gyorgyi:1985:PMD

[Gyö85] Géza Györgyi. The perturbative method for discrete processes and its physical application. Lecture Notes in Mathematics, 1163: 83–94, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076420/.

Gyongy:1989:SSP

[Gyö89] I. Gyöngy. The stability of stochastic partial differential equations and applications. Theorems on supports. Lecture Notes in Mathematics, 1390:91–118, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083939/.

Gine:1986:LCL

[GZ86a] Evarist Giné and Joel Zinn. Lectures on the central limit theorem for empirical processes. Lecture Notes in Mathematics, 1221: 50–113, 1986. CODEN LNMAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099111/.

Gusein-Zade:1986:EAI

[GZ86b] S. M. Guseĭn-Zade. An equivariant analogue of the index of a gradient vector field. Lecture Notes in Mathematics, 1214:196– 210, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075966/.

Gill:1987:EFD

[GZ87a] Tepper L. Gill and W. W. Zachary. Existence and finitedimensionality of attractors for the Landau–Lifschitz equations.

Lecture Notes in Mathematics, 1285:134–142, 1987. CODEN LN-MAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080589/.

Gill:1987:NSP

[GZ87b] Tepper L. Gill and Woodford W. Zachary, editors. Nonlinear Semigroups, Partial Differential Equations and Attractors: Proceedings of a Symposium held in Washington, D.C., August 5–8, 1985, volume 1248 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1248. URL http:

//link.springer.com/book/10.1007/BFb0077409;   http://
www.springerlink.com/content/978-3-540-47791-4.

Gine:1989:MCL

[GZ89] Evarist Giné and Joel Zinn. L_{p} multipliers in the central limit theorem with p-stable limit. Lecture Notes in Mathematics, 1391: 74–81, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083382/.

Haagerup:1985:IDC

[Haa85] Uffe Haagerup. Injectivity and decomposition of completely bounded maps. Lecture Notes in Mathematics, 1132:170–222, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074885/.

Haas:1987:GMF

[Haa87] Andrew Haas. The geometry of Markoff forms. Lecture Notes in Mathematics, 1240:135–144, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072978/.

Haba:1987:SES

[Hab87] Z. Haba. Stochastic equations for some Euclidean fields. Lecture Notes in Mathematics, 1250:106–119, 1987. CODEN LN-MAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077351/.

Haba:1988:SSP

[Hab88] Z. Haba. Supersymmetry and stochastic processes. Lecture Notes in Mathematics, 1325:101–112, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077919/.

Hadjiev:1985:FPP

[Had85] Dimitar I. Hadjiev. The first passage problem for generalized Ornstein–Uhlenbeck-processes with non-positive jumps. Lecture Notes in Mathematics, 1123:80–90, 1985. CODEN LN-MAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075840/.

Hadeler:1986:BDP

[Had86] K. P. Hadeler. Birth and death processes with killing and applications to parasitic infections. Lecture Notes in Mathematics, 1212:175–186, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076247/.

Hahn:1985:OLF

[Hah85] Wolfgang Hahn. Über Orthogonalpolynome, die linearen Funktionalgleichungen genügen. (German) []. Lecture Notes in Mathematics, 1171:16–35, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076529/.

Hahn:1986:ATM

[Hah86] Alexander J. Hahn. Algebraic K-theory, Morita theory, and the classical groups. Lecture Notes in Mathematics, 1185:88– 117, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076172/.

Hahubia:1989:AEC

[Hah89] Ts. G. Hahubia. On the adaptive estimation of change points. Lecture Notes in Mathematics, 1412:33–49, 1989. CODEN LN-MAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084163/.

Haiman:1985:LLP

[Hai85] Mark Haiman. Linear lattice proof theory: an overview. Lecture Notes in Mathematics, 1149:129–141, 1985. CODEN LN-MAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098460/.

Hain:1987:HAM

[Hai87a] Richard M. Hain. Higher Albanese manifolds. Lecture Notes in Mathematics, 1246:84–91, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077531/.

Hain:1987:IIM

[Hai87b] Richard M. Hain. Iterated integrals and mixed Hodge structures on homotopy groups. Lecture Notes in Mathematics, 1246:75– 83, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077530/.

Hajarnavis:1987:HCM

[Haj87] C. R. Hajarnavis. Homological and Cohen–Macaulay properties in non-commutative Noetherian rings. Lecture Notes in Mathematics, 1296:235–246, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078530/.

Hajela:1989:ESF

[Haj89] D. Hajela. Exponential sums and faster than Nyquist signaling. Lecture Notes in Mathematics, 1383:64–75, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083570/.

Halperin:1986:RI

[Hal86] Stephen Halperin. The radical of π_{∗}(ωS) ⊗ Q, II. Lecture Notes in Mathematics, 1183:199–210, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075460/.

Hale:1987:ASS

[Hal87] Jack K. Hale. Asymptotically smooth semigroups and applications. Lecture Notes in Mathematics, 1248:85–93, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077418/.

Hall:1989:ISS

[Hal89] Mark E. Hall. An introduction to Schubert submodules. Lecture Notes in Mathematics, 1373:111–123, 1989. CODEN LN-MAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088891/.

Hanumanthachari:1985:ENT

[Han85] J. Hanumanthachari. An extended Nagell totient and related Ramanujan sums. Lecture Notes in Mathematics, 1122:115– 129, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075755/.

Handelman:1987:DN

[Han87a] David E. Handelman. Definitions and notation. Lecture Notes in Mathematics, 1282:6–13, 1987. CODEN LNMAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078911/.

Handelman:1987:FIS

[Han87b] David E. Handelman. Factoriality and integral simplicity. Lecture Notes in Mathematics, 1282:50–59, 1987. CODEN LN-MAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078916/.

Handelman:1987:ICC

[Han87c] David E. Handelman. Integral closure and Cohen–Macauleyness. Lecture Notes in Mathematics, 1282:28–40, 1987. CODEN LN-MAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078913/.

Handelman:1987:Ia

[Han87d] David E. Handelman. Introduction. Lecture Notes in Mathematics, 1282:1–5, 1987. CODEN LNMAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078910/.

Handelman:1987:Ib

[Han87e] David E. Handelman. Isomorphisms. Lecture Notes in Mathematics, 1282:67–88, 1987. CODEN LNMAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078918/.

Handelman:1987:MII

[Han87f] David E. Handelman. Meet-irreducibile ideals in R_{K}. Lecture Notes in Mathematics, 1282:60–66, 1987. CODEN LN-MAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078917/.

Handelman:1987:PMF

[Han87g] David E. Handelman. Projective R_{K}-modules are free. Lecture Notes in Mathematics, 1282:41–43, 1987. CODEN LN-MAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078914/.

Handelman:1987:RWP

[Han87h] David E. Handelman. A random walk problem. Lecture Notes in Mathematics, 1282:14–27, 1987. CODEN LNMAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078912/.

Handelman:1987:SI

[Han87i] David E. Handelman. States on ideals. Lecture Notes in Mathematics, 1282:44–49, 1987. CODEN LNMAA2. ISBN 3-540-18400-7 (print), 3-540-47951-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078915/.

Hansen:1987:DGP

[Han87j] Vagn Lundsgaard Hansen, editor. Differential Geometry: Proceedings of the Nordic Summer School held in Lyngby, Denmark Jul. 29–Aug. 9, 1985, volume 1263 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1263. URL http:

//link.springer.com/book/10.1007/BFb0078607;   http://
www.springerlink.com/content/978-3-540-47249-0.

Hansen:1988:BSN

[Han88] Wolfhard Hansen. Balayage spaces — a natural setting for potential theory. Lecture Notes in Mathematics, 1344:98–117, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103346/.

Hansen:1989:VPP

[Han89] Wolfhard Hansen. Valeurs propres pour l’opérateur de Schrödinger. (French) []. Lecture Notes in Mathematics, 1393:117–134, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085775/.

Happel:1986:DA

[Hap86] Dieter Happel. Dynkin algebras. Lecture Notes in Mathematics, 1220:1–14, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099501/.

Happel:1987:RC

[Hap87] Dieter Happel. Repetitive categories. Lecture Notes in Mathematics, 1273:298–317, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078850/.

Happel:1989:HCF

[Hap89] Dieter Happel. Hochschild cohomology of finite-dimensional algebras. Lecture Notes in Mathematics, 1404:108–126, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084073/.

Harris:1985:IIV

[Har85a] Joe Harris. An introduction to infinitesimal variations of Hodge structures. Lecture Notes in Mathematics, 1111:51–58, 1985. CO-DEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084584/.

Hartmann:1985:CS

[Har85b] F. Hartmann. Castigliano and Sobolev. Lecture Notes in Mathematics, 1121:137–152, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076268/.

Harris:1986:ISF

[Har86] T. E. Harris. Isotropic stochastic flows and a related property of non-random potential flows. Lecture Notes in Mathematics, 1203: 66–74, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076873/.

Harbater:1987:GCA

[Har87a] David Harbater. Galois coverings of the arithmetic line. Lecture Notes in Mathematics, 1240:165–195, 1987. CODEN LN-MAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072980/.

Harnik:1987:CLM

[Har87b] Victor Harnik. Countable or ω_{1}-like models of Presburger’s arithmetic. Lecture Notes in Mathematics, 1292:89–106, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082233/.

Hart:1987:EO

[Har87c] Bradd Hart. An exposition of OTOP. Lecture Notes in Mathematics, 1292:107–126, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082234/.

Harten:1987:PRE

[Har87d] Ami Harten. Preliminary results on the extension of Eno schemes to two-dimensional problems. Lecture Notes in Mathematics, 1270: 23–40, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078315/.

Hartley:1987:CLF

[Har87e] B. Hartley. Centralizers in locally finite groups. Lecture Notes in Mathematics, 1281:36–51, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078689/.

Haraux:1988:RRS

[Har88a] A. Haraux. Recent results on semi-linear hyperbolic problems in bounded domains. Lecture Notes in Mathematics, 1324:118– 126, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100787/.

Harbourne:1988:IBU

[Har88b] Brian Harbourne. Iterated blow-ups and moduli for rational surfaces. Lecture Notes in Mathematics, 1311:101–117, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082911/.

Harer:1988:CMS

[Har88c] John L. Harer. The cohomology of the moduli space of curves. Lecture Notes in Mathematics, 1337:138–221, 1988. CODEN LN-MAA2. ISBN 3-540-50080-4 (print), 3-540-45920-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082808/.

Haslinger:1985:NIG

[Has85] Doz. Dr. F. Haslinger. Newton’sche Interpolationspolynome Und Gleichverteilung. (German) [Newton’s interpolation polynomials and uniform distribution]. Lecture Notes in Mathematics, 1114: 16–18, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101640/.

Hattori:1986:ACA

[Hat86] Akio Hattori. Almost complex S^{1}-actions on cohomology complex projective spaces. Lecture Notes in Mathematics, 1217:115– 122, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072818/.

Hattori:1989:CAS

[Hat89] Akio Hattori. Circle actions on symplectic manifolds. Lecture Notes in Mathematics, 1375:89–97, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085601/.

Haussmann:1985:LZP

[Hau85] U. G. Haussmann. L’équation de Zakai et le problème séparé du contrôle optimal stochastique. (French) []. Lecture Notes in Mathematics, 1123:37–62, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075838/.

Hawkes:1987:SEP

[Haw87] Trevor Hawkes. Subgroup embedding properties. Lecture Notes in Mathematics, 1281:52–64, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078690/.

Hayman:1988:ATA

[Hay88] W. K. Hayman. Ahlfors’ theorem on asymptotic values and some developments from it. Lecture Notes in Mathematics, 1351:159– 168, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081251/.

Hazod:1989:DGC

[Haz89] W. Hazod. On the decomposability group of a convolution semigroup. Lecture Notes in Mathematics, 1391:99–111, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083384/.

Heath-Brown:1989:WIH

[HB89] D. R. Heath-Brown. Weyl’s inequality and Hua’s inequality. Lecture Notes in Mathematics, 1380:87–92, 1989. CODEN LN-MAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086547/.

He:1989:MSD

[He89] Yuanjiang He. MD-semigroups, decomposition of point processes, central limit theorems for certain T_{2}-semigroups. Lecture Notes in Mathematics, 1379:107–124, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087849/.

Hegenbarth:1988:TCS

[Heg88] Friedrich Hegenbarth. On the K-theory of the classifying spaces of the general linear groups over finite fields. Lecture Notes in Mathematics, 1350:259–265, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081479/.

Heinkel:1985:BLL

[Hei85] Bernard Heinkel. On Brunk’s law of large numbers in some type 2 spaces. Lecture Notes in Mathematics, 1153:249–257, 1985. CO-DEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074953/.

Heinkel:1986:AMI

[Hei86] Bernard Heinkel. An application of a martingale inequality of Dubins and Freedman to the law of large numbers in Banach spaces. Lecture Notes in Mathematics, 1193:29–43, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077098/.

Heineken:1987:SIG

[Hei87] Hermann Heineken. Soluble irreducible groups of automorphisms of certain groups of class two. Lecture Notes in Mathematics, 1281: 65–72, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078691/.

Heinonen:1988:NEM

[Hei88a] Juha Heinonen. Nonlinear elliptic measures. Lecture Notes in Mathematics, 1344:202–204, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103354/.

Heinz:1988:QMS

[Hei88b] Erhard Heinz. On quasi-minimal surfaces. Lecture Notes in Mathematics, 1340:139–145, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082892/.

Hejhal:1985:CAW

[Hej85] Dennis A. Hejhal. A classical approach to a well-known spectral correspondence on quaternion groups. Lecture Notes in Mathematics, 1135:127–196, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074604/.

Hejhal:1988:RGR

[Hej88] Dennis A. Hejhal. Regular b-groups and repeated Dehn twists. Lecture Notes in Mathematics, 1351:169–192, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081252/.

Helmberg:1985:IUU

[Hel85] Gilbert Helmberg. Itineraries under unimodal maps. Lecture Notes in Mathematics, 1163:95–100, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076421/.

Helson:1986:BM

[Hel86a] Henry Helson. Back matter. Lecture Notes in Mathematics, 1227: ??, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47366-4/1.pdf.

Helson:1986:BT

[Hel86b] Henry Helson. Bochner’s theorem. Lecture Notes in Mathematics, 1227:42–52, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101632/.

Helson:1986:CL

[Hel86c] Henry Helson. Cocycles on the line. Lecture Notes in Mathematics, 1227:76–101, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101634/.

Helson:1986:DC

[Hel86d] Henry Helson. Distribution of cocycles. Lecture Notes in Mathematics, 1227:53–75, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101633/.

Helson:1986:FM

[Hel86e] Henry Helson. Front matter. Lecture Notes in Mathematics, 1227: ??, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-47366-4/1.pdf.

Helson:1986:MSM

[Hel86f] Henry Helson. Multiplicity of spectral measures. Lecture Notes in Mathematics, 1227:1–22, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101630/.

Helson:1986:STa

[Hel86g] Henry Helson. The spectral theorem. Lecture Notes in Mathematics, 1227:23–41, 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101631/.

Helson:1986:STb

[Hel86h] Henry Helson. The Spectral Theorem, volume 1227 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17197-5 (print), 3-540-47366-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 104 pp. LCCN QA3

.L28 no. 1227. URL http://link.springer.com/book/10.1007/
BFb0101629;   http://www.springerlink.com/content/978-
3-540-47366-4.

Hellerstein:1987:RZD

[Hel87] Simon Hellerstein. Real zeros of derivatives of meromorphic functions and solutions of differential equations — a survey with problems and conjectures. Lecture Notes in Mathematics, 1275:197– 209, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078352/.

Helffer:1988:BM

[Hel88a] Bernard Helffer. Back matter. Lecture Notes in Mathematics, 1336:??, 1988. CODEN LNMAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-45913-2/1.pdf.

Helffer:1988:BKW

[Hel88b] Bernard Helffer. B.K.W. construction for a potential near the bottom in the case of non-degenerate minima. Lecture Notes in Mathematics, 1336:7–18, 1988. CODEN LNMAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078117/.

Helffer:1988:DE

[Hel88c] Bernard Helffer. The decay of the eigenfunctions. Lecture Notes in Mathematics, 1336:19–29, 1988. CODEN LNMAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078118/.

Helffer:1988:FM

[Hel88d] Bernard Helffer. Front matter. Lecture Notes in Mathematics, 1336:??, 1988. CODEN LNMAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-45913-2/1.pdf.

Helffer:1988:GSC

[Hel88e] Bernard Helffer. Generalities on semi-classical analysis. Lecture Notes in Mathematics, 1336:1–6, 1988. CODEN LNMAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078116/.

Helffer:1988:IRR

[Hel88f] Bernard Helffer. An introduction to recent results of Witten. Lecture Notes in Mathematics, 1336:61–70, 1988. CODEN LN-MAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078120/.

Helffer:1988:SOM

[Hel88g] Bernard Helffer. On Schrödinger operators with magnetic fields. Lecture Notes in Mathematics, 1336:79–99, 1988. CODEN LN-MAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078122/.

Helffer:1988:SOP

[Hel88h] Bernard Helffer. On Schrödinger operators with periodic electric potentials. Lecture Notes in Mathematics, 1336:71–78, 1988. CODEN LNMAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078121/.

Helffer:1988:SIB

[Hel88i] Bernard Helffer. Study of interaction between the wells. Lecture Notes in Mathematics, 1336:30–60, 1988. CODEN LN-MAA2. ISBN 3-540-50076-6 (print), 3-540-45913-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078119/.

Hellegouarch:1989:PDB

[Hel89] Yves Hellegouarch. Positive definite binary quadratic forms over k[X]. Lecture Notes in Mathematics, 1380:93–119, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086548/.

Hempel:1985:FGI

[Hem85] John Hempel. The finitely generated intersection property for Kleinian groups. Lecture Notes in Mathematics, 1144:18–24, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075010/.

Hemker:1986:DCH

[Hem86] P. W. Hemker. Defect correction and higher order schemes for the multi grid solution of the steady Euler equations. Lecture Notes in Mathematics, 1228:149–165, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072646/.

Hennig:1985:YMA

[Hen85] J. D. Hennig. Yang–Mills aspects of Poincaré gauge theories. Lecture Notes in Mathematics, 1139:169–188, 1985. CODEN LN-MAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074583/.

Hennart:1986:NMN

[Hen86a] J. P. Hennart. Nodal methods for the numerical solution of partial differential equations. Lecture Notes in Mathematics, 1230: 175–190, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072680/.

Hennart:1986:NAP

[Hen86b] Jean-Pierre Hennart, editor. Numerical Analysis: Proceedings of the Fourth IIMAS Workshop held at Guanajuato, Mexico, July 23–27, 1984, volume 1230 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1230. URL http:

//link.springer.com/book/10.1007/BFb0072666;   http://
www.springerlink.com/content/978-3-540-47379-4.

Hennequin:1986:EEP

[Hen86c] P. L. Hennequin, editor. École d’Été de Probabilités de Saint Flour XIV — 1984. (French) [], volume 1180 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16441-3 (print), 3-540-39781-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1180. URL http:

//link.springer.com/book/10.1007/BFb0074917;   http://
www.springerlink.com/content/978-3-540-39781-6.

Hensz:1989:OSS

[Hen89a] Ewa Hensz. Orthogonal series and strong laws of large numbers in von Neumann algebras. Lecture Notes in Mathematics, 1396:221–228, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083554/.

Hensz:1989:SLL

[Hen89b] Ewa Hensz. Strong laws of large numbers for orthogonal sequences in von Neumann algebras. Lecture Notes in Mathematics, 1391:112–124, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083385/.

Hernandez:1986:QMN

[Her86] Jesús Hernández. Qualitative methods for nonlinear diffusion equations. Lecture Notes in Mathematics, 1224:47–118, 1986. CODEN LNMAA2. ISBN 3-540-17192-4 (print), 3-540-47352-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072688/.

Hernandez:1987:PCP

[Her87a] L. J. Hernández. Proper cohomologies and the proper classification problem. Lecture Notes in Mathematics, 1298:171–191, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083009/.

Herzog:1987:LCM

[Her87b] J. Herzog. Linear Cohen–Macaulay modules on integral quadrics. Lecture Notes in Mathematics, 1296:214–227, 1987. CODEN LN-MAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078528/.

Hernandez:1988:CPR

[Her88a] J. L. García Hernández. Continuous and pf rings of quotients. Lecture Notes in Mathematics, 1328:94–105, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100919/.

Herrlich:1988:RPM

[Her88b] Horst Herrlich. On the representability of partial morphisms in top and in related constructs. Lecture Notes in Mathematics, 1348:143–153, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081356/.

Herron:1988:MBC

[Her88c] David A. Herron. Metric boundary conditions for plane domains. Lecture Notes in Mathematics, 1351:193–200, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081253/.

Hernandez:1989:ARS

[Her89] R. Hernandez. Appendix to “On regular and stable ruled surfaces in P^{3}”. Lecture Notes in Mathematics, 1389:16–18, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085920/.

Hesselink:1986:DFR

[Hes86] Wim H. Hesselink. Dimension formulas related to a tame quiver. Lecture Notes in Mathematics, 1220:15–24, 1986. CODEN LN-MAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099502/.

Heyer:1986:PMG

[Hey86] Herbert Heyer, editor. Probability Measures on Groups VIII: Proceedings of a Conference held in Oberwolfach, November 10–16, 1985, volume 1210 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 0-387-16806-0, 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1210. URL http:

//link.springer.com/book/10.1007/BFb0077166;   http://
www.springerlink.com/content/978-3-540-44852-5.

Hanin:1989:PUB

[HGYR89] L. G. Hanin, R. E. Goot, A. Yu. Yakovlev, and S. T. Rachev. Precise upper bounds for the functionals describing tumour treatment efficiency. Lecture Notes in Mathematics, 1412:50–67, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084164/.

Hartshorne:1985:SAS

[HH85] R. Hartshorne and A. Hirschowitz. Smoothing algebraic space curves. Lecture Notes in Mathematics, 1124:98–131, 1985. CO-DEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074998/.

Hirschowitz:1986:CFS

[HH86] A. Hirschowitz and K. Hulek. Complete families of stable vector bundles over P_{2}. Lecture Notes in Mathematics, 1194:19–33, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076992/.

Hida:1989:IDR

[Hid89] Takeyuki Hida. Infinite dimensional rotation group and unitary group. Lecture Notes in Mathematics, 1379:125–134, 1989. CO-DEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087850/.

Higgs:1985:IAL

[Hig85] Denis Higgs. Interpolation antichains in lattices. Lecture Notes in Mathematics, 1149:142–149, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098461/.

Higuchi:1987:PTD

[Hig87] Yasunari Higuchi. Percolation of the two-dimensional Ising model. Lecture Notes in Mathematics, 1250:120–127, 1987. CODEN LN-MAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077352/.

Hildebrandt:1985:HMR

[Hil85a] Stefan Hildebrandt. Harmonic mappings of Riemannian manifolds. Lecture Notes in Mathematics, 1161:1–117, 1985. CODEN LNMAA2. ISBN 3-540-16040-X (print), 3-540-39716-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075136/;   http:
//link.springer.com/content/pdf/bfm:978-3-540-39716-8/
1.pdf.

Hilsum:1985:SOL

[Hil85b] Michel Hilsum. Signature operator on Lipschitz manifolds and unbounded Kasparov bimodules. Lecture Notes in Mathematics, 1132:254–288, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074888/.

Hildebrand:1989:SNA

[Hil89] Adolf Hildebrand. Some new applications of the large sieve. Lecture Notes in Mathematics, 1383:76–88, 1989. CODEN LN-MAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083571/.

Hineva:1985:SSF

[Hin85] Stefana Hineva. Submanifolds and the second fundamental tensor. Lecture Notes in Mathematics, 1156:194–203, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075091/.

Hinz:1987:BES

[Hin87] Andreas M. Hinz. Behavior of eigenfunctions and the spectrum of Schrödinger operators. Lecture Notes in Mathematics, 1285: 175–179, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080595/.

Hindman:1989:URT

[Hin89] Neil Hindman. Ultrafilters and Ramsey theory — an update. Lecture Notes in Mathematics, 1401:97–118, 1989. CODEN LN-MAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097334/.

Hirsch:1985:ADT

[Hir85] Morris W. Hirsch. Attractors for discrete-time monotone dynamical systems in strongly ordered spaces. Lecture Notes in Mathematics, 1167:141–153, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075221/.

Hirschowitz:1987:SPM

[Hir87] André Hirschowitz. Sections planes et multisécantes pour les courbes gauches génériques principales. (French) []. Lecture Notes in Mathematics, 1266:124–155, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078182/.

Hirschowitz:1989:NFC

[Hir89a] A. Hirschowitz. Nouvelles fonctions constructibles et rang des images directes. (French) []. Lecture Notes in Mathematics, 1389: 103–111, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085928/.

Hirschowitz:1989:CDG

[Hir89b] André Hirschowitz. Cohérence et dualité sur le gros site de Zariski. (French) []. Lecture Notes in Mathematics, 1389:91–102, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085927/.

Hirschowitz:1989:RSH

[Hir89c] André Hirschowitz. La rationalité des schémas de Hilbert de courbes gauches rationnelles suivant Katsylo. (French) []. Lecture Notes in Mathematics, 1389:87–90, 1989. CODEN LN-MAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085926/.

Hoffmann-Jorgensen:1985:NSC

[HJ85] J. Hoffmann-Jørgensen. Necessary and sufficient condition for the uniform law of large numbers. Lecture Notes in Mathematics, 1153:258–272, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074954/.

Hastings:1987:CM

[HJ87a] Harold M. Hastings and Mahendra Jani. Covering maps in top^{N}. Lecture Notes in Mathematics, 1283:115–124, 1987. CODEN LN-MAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081422/.

Hoffmann-Jorgensen:1987:GMP

[HJ87b] J. Hoffmann-Jørgensen. The general marginal problem. Lecture Notes in Mathematics, 1242:77–367, 1987. CODEN LN-MAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072443/.

Hardie:1989:WPT

[HJ89] K. A. Hardie and A. V. Jansen. A Whitehead product for track groups. Lecture Notes in Mathematics, 1370:163–170, 1989. CO-DEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085227/.

Huang:1989:BPS

[HjJ89] Hai-Hua Huang and Bo ju Jiang. Braids and periodic solutions. Lecture Notes in Mathematics, 1411:107–123, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086445/.

Hasanis:1985:AGM

[HK85] Th. Hasanis and D. Koutroufiotis. Applications of the Gauss mapping for hypersurfaces of the sphere. Lecture Notes in Mathematics, 1156:180–193, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075090/.

Hardie:1987:HCH

[HK87a] K. A. Hardie and K. H. Kamps. The homotopy category of homotopy factorizations. Lecture Notes in Mathematics, 1298:162– 170, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083008/.

Hazareesingh:1987:SPI

[HK87b] L. Hazareesingh and D. Kannan. Stochastic product integration and stochastic equations. Lecture Notes in Mathematics, 1236: 72–120, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072884/.

Hammer:1989:BRD

[HK89] P. L. Hammer and B. Kalantari. A bound on the roof-duality gap. Lecture Notes in Mathematics, 1403:254–257, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083469/.

Hardt:1988:RAM

[HKL88] Robert Hardt, David Kinderlehrer, and Mitchell Luskin. Remarks about the mathematical theory of liquid crystals. Lecture Notes in Mathematics, 1340:123–138, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082891/.

Hahn:1989:ULI

[HKW89] M. G. Hahn, J. Kuelbs, and D. C. Weiner. A universal law of the iterated logarithm for trimmed and censored sums. Lecture Notes in Mathematics, 1391:82–98, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083383/.

Halperin:1986:HSC

[HL86] Stephen Halperin and Gerson Levin. High skeleta of CW complexes. Lecture Notes in Mathematics, 1183:211–217, 1986. CO-DEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075461/.

Halperin:1988:NCD

[HL88a] S. Halperin and J.-M. Lemaire. Notions of category in differential algebra. Lecture Notes in Mathematics, 1318:138–154, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077800/.

Herrera:1988:RAO

[HL88b] René Hernández Herrera and Guillermo López Lagomasino. Relative asymptotics of orthogonal polynomials with respect to varying measures II. Lecture Notes in Mathematics, 1354:140–154, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089591/.

Hlawka:1985:BLDa

[Hla85a] E. Hlawka. Bemerkung Zum Lemma Von Du Bois–Reymond. (German) []. Lecture Notes in Mathematics, 1114:26–29, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101642/.

Hlawka:1985:PIA

[Hla85b] E. Hlawka. Über Ein Produkt, Das In Der Interpolation Analytischer Funktionen Im Einheitskreis Auftritt. (German) []. Lecture Notes in Mathematics, 1114:19–25, 1985. CODEN LN-MAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101641/.

Hlawka:1985:BLDb

[Hla85c] Edmund Hlawka. Bemerkung Zum Lemma von Du Bois–Reymond II. (German) [Remark on the Du Bois–Reymond lemma II]. Lecture Notes in Mathematics, 1114:30–39, 1985. CODEN LN-MAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101643/.

Hlawka:1985:ZAW

[Hla85d] Edmund Hlawka, editor. Zahlentheoretische Analysis: Wiener Seminarberichte 1980–82, volume 1114 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1114. URL http:

//link.springer.com/book/10.1007/BFb0101637;   http://
www.springerlink.com/content/978-3-540-39281-1.

Hlawka:1987:DMV

[Hla87a] Edmund Hlawka. Über Die Direkten Methoden der Variationsrechnung Und Gleichverteilung. (German) []. Lecture Notes in Mathematics, 1262:50–85, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078596/.

Hlawka:1987:ZAI

[Hla87b] Edmund Hlawka, editor. Zahlentheoretische Analysis II: Wiener Seminarberichte 1984–86, volume 1262 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1262. URL http:

//link.springer.com/book/10.1007/BFb0078590;   http://
www.springerlink.com/content/978-3-540-47256-8.

Hilden:1985:UK

[HLM85a] Hugh M. Hilden, María Teresa Lozano, and José María Montesino. Universal knots. Lecture Notes in Mathematics, 1144:25–59, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075011/.

Hog:1985:PCM

[HLM85b] Cynthia Hog, Martin Lustig, and Wolfgang Metzler. Presentation classes, 3-manifolds and free products. Lecture Notes in Mathematics, 1167:154–167, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075222/.

Hilden:1988:UGB

[HLM88] Hugh M. Hilden, María Teresa Lozano, and José María Montesinos. On the universal group of the Borromean rings. Lecture Notes in Mathematics, 1350:1–13, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081465/.

Haydon:1985:LS

[HLR85] Richard Haydon, Mireille Lévy, and Yves Raynaud. On the local structure of L_{p}(X). Lecture Notes in Mathematics, 1166:74– 79, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074695/.

Hirsch:1987:STD

[HM87] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 8, volume 1235 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 209 pp. LCCN QA3 .L28

no. 1235. URL http://link.springer.com/book/10.1007/
BFb0072744;   http://www.springerlink.com/content/978-
3-540-47406-7.

Hinrichsen:1988:OPR

[HM88a] Diederich Hinrichsen and Matthias Motscha. Optimization problems in the robustness analysis of linear state space systems. Lecture Notes in Mathematics, 1354:54–78, 1988. CODEN LN-MAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089583/.

Hu:1988:IMS

[HM88b] Y. Z. Hu and P.-A. Meyer. Sur les intégrales multiples de Stratonovitch. (French) []. Lecture Notes in Mathematics, 1321: 72–81, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084119/.

Hu:1988:CWI

[HM88c] Yao Zhong Hu and Paul-André Meyer. Chaos de Wiener et intégrale de Feynman. (French) []. Lecture Notes in Mathematics, 1321:51–71, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084118/.

Hmissi:1987:RRA

[Hmi87] Mohamed Hmissi. Régularité relative au noyau de Poisson. (French) []. Lecture Notes in Mathematics, 1235:88–104, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072750/.

Hmissi:1989:SGD

[Hmi89] Mohamed Hmissi. Semi-groupes deterministes. Lecture Notes in Mathematics, 1393:135–144, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085776/.

Henle:1985:BE

[HMW85] J. M. Henle, A. R. D. Mathias, and W. Hugh Woodin. A barren extension. Lecture Notes in Mathematics, 1130:195–207, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075312/.

Hackbusch:1986:MMC

[HN86] W. Hackbusch and Z. P. Nowak. Multigrid methods for calculating the lifting potential incompressible flows around three-dimensional bodies. Lecture Notes in Mathematics, 1228:135–148, 1986. CO-DEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0072645/.

Hamachi:1988:TCB

[HN88] Toshihiro Hamachi and Masakazu Nasu. Topological conjugacy for 1-block factor maps of subshifts and sofic covers. Lecture Notes in Mathematics, 1342:251–260, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082835/.

Hazod:1989:CTT

[HN89] W. Hazod and S. Nobel. Convergence-of-types theorem for simply connected nilpotent Lie groups. Lecture Notes in Mathematics, 1379:90–106, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087848/.

Hanche-Olsen:1985:JAT

[HO85] Harald Hanche-Olsen. JB-algebras with tensor products are C∗- algebras. Lecture Notes in Mathematics, 1132:223–229, 1985. CO-DEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074886/.

Hodges:1989:TNR

[Hod89] Timothy J. Hodges. K-theory of Noetherian rings. Lecture Notes in Mathematics, 1404:246–268, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084079/.

Hoffman:1985:EMS

[Hof85] David Hoffman. Embedded minimal surfaces, computer graphics and elliptic functions. Lecture Notes in Mathematics, 1156: 204–215, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075092/.

Hoff:1987:TET

[Hof87] David Hoff. Two existence theorems for systems of conservation laws with dissipation. Lecture Notes in Mathematics, 1270:181– 194, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078328/.

Hognas:1989:NSA

[Hög89] Göran Högnäs. A note on the semigroup of analytic mappings with a common fixed point. Lecture Notes in Mathematics, 1379: 135–140, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087851/.

Hohlov:1987:ERC

[Hoh87] Yu. S. Hohlov. The estimation of the rate of convergence in the integral limit theorem in the Euclidean motion group. Lecture Notes in Mathematics, 1233:1–10, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072705/.

Holevo:1988:NGC

[Hol88] A. S. Holevo. A noncommutative generalization of conditionally positive definite functions. Lecture Notes in Mathematics, 1303:128–148, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078059/.

Holevo:1989:LTR

[Hol89a] A. S. Holevo. Limit theorems for repeated measurements and continuous measurement processes. Lecture Notes in Mathematics, 1396:229–255, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083555/.

Holm:1989:HTR

[Hol89b] Darryl D. Holm. Hamiltonian techniques for relativistic fluid dynamics and stability theory. Lecture Notes in Mathematics, 1385: 65–151, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084029/.

Holzapfel:1989:VFD

[Hol89c] R.-P. Holzapfel. Volumes of fundamental domains of Picard modular groups. Lecture Notes in Mathematics, 1399:60–88, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095969/.

Homer:1985:MPD

[Hom85] Steven Homer. Minimal polynomial degrees of nonrecursive sets. Lecture Notes in Mathematics, 1141:193–202, 1985. CODEN LN-MAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076221/.

Hoppe:1986:CAT

[Hop86] R. Hoppe. Constructive aspects in time optimal control. Lecture Notes in Mathematics, 1190:243–272, 1986. CODEN LN-MAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076710/.

Hopkins:1987:SRP

[Hop87] William E. Hopkins, Jr. Some remarks on a problem in stochastic optimal control. Lecture Notes in Mathematics, 1236:121– 130, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072885/.

Hopf:1989:DGL

[Hop89a] Heinz Hopf, editor. Differential Geometry in the Large: Seminar Lectures New York University 1946 and Stanford University 1956, volume 1000 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. CODEN LNMAA2. ISBN 3-540-51497-X (print), 3-540-39482-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

LCCN QA3 .L28 no. 1000. URL http://link.springer.com/
book/10.1007/3-540-39482-6;   http://www.springerlink.
com/content/978-3-540-39482-2.

Hopkins:1989:MAR

[Hop89b] Michael J. Hopkins. Minimal atlases of real projective spaces. Lecture Notes in Mathematics, 1370:171–192, 1989. CODEN LN-MAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085228/.

Hormander:1987:LCS

[Hör87] Lars Hörmander. The lifespan of classical solutions of non-linear hyperbolic equations. Lecture Notes in Mathematics, 1256:214– 280, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077745/.

Hovstad:1989:ICF

[Hov89] R. M. Hovstad. Irrational continued fractions. Lecture Notes in Mathematics, 1406:18–24, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096163/.

Hoyt:1985:SAI

[Hoy85] William L. Hoyt. On surfaces associated with an indefinite ternary lattice. Lecture Notes in Mathematics, 1135:197–210, 1985. CO-DEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074605/.

Hoyt:1987:NES

[Hoy87] William L. Hoyt. Notes on elliptic K3 surfaces. Lecture Notes in Mathematics, 1240:196–213, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072981/.

Hoyt:1989:EFK

[Hoy89] William L. Hoyt. Elliptic fiberings of Kummer surfaces. Lecture Notes in Mathematics, 1383:89–110, 1989. CODEN LN-MAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083572/.

Holz:1987:GCT

[HPS87a] Michael Holz, Klaus-Peter Podewski, and Karsten Steffens. General criteria and their applications. Lecture Notes in Mathematics, 1238:33–97, 1987. CODEN LNMAA2. ISBN 3-540-17221-1 (print), 3-540-47426-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072631/.

Holz:1987:ICF

[HPS87b] Michael Holz, Klaus-Peter Podewski, and Karsten Steffens. Injective Choice Functions, volume 1238 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17221-1 (print), 3-540-47426-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 183 pp. LCCN QA3 .L28

no. 1238. URL http://link.springer.com/book/10.1007/
BFb0072628;   http://www.springerlink.com/content/978-
3-540-47426-5.

Holz:1987:I

[HPS87c] Michael Holz, Klaus-Peter Podewski, and Karsten Steffens. Introduction. Lecture Notes in Mathematics, 1238:1–9, 1987. CODEN LNMAA2. ISBN 3-540-17221-1 (print), 3-540-47426-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072629/.

Holz:1987:MTM

[HPS87d] Michael Holz, Klaus-Peter Podewski, and Karsten Steffens. Miscellaneous theorems on marriages. Lecture Notes in Mathematics, 1238:158–174, 1987. CODEN LNMAA2. ISBN 3-540-17221-1 (print), 3-540-47426-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072633/.

Holz:1987:STF

[HPS87e] Michael Holz, Klaus-Peter Podewski, and Karsten Steffens. Set theoretic foundations. Lecture Notes in Mathematics, 1238:10– 32, 1987. CODEN LNMAA2. ISBN 3-540-17221-1 (print), 3-540-47426-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072630/.

Holz:1987:SPF

[HPS87f] Michael Holz, Klaus-Peter Podewski, and Karsten Steffens. Structural properties of families. Lecture Notes in Mathematics, 1238: 98–157, 1987. CODEN LNMAA2. ISBN 3-540-17221-1 (print), 3-540-47426-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072632/.

Heath:1989:NTN

[HPY89] Philip Heath, Renzo Piccinini, and Chengye You. Nielsen-type numbers for periodic points, I. Lecture Notes in Mathematics, 1411:88–106, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086444/.

Hodgson:1985:IIL

[HR85] Craig Hodgson and J. H. Rubinstein. Involutions and isotopies of lens spaces. Lecture Notes in Mathematics, 1144:60–96, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075012/.

Happel:1986:DCT

[HR86] Dieter Happel and Claus Michael Ringel. The derived category of a tubular algebra. Lecture Notes in Mathematics, 1177:156– 180, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075264/.

Hawkins:1988:ATF

[HR88a] Jane M. Hawkins and E. A. Robinson, Jr. Approximately transitive (2) flows and transformations have simple spectrum. Lecture Notes in Mathematics, 1342:261–280, 1988. CODEN LN-MAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082836/.

Holme:1988:EPV

[HR88b] Audun Holme and Joel Roberts. On the embeddings of projective varieties. Lecture Notes in Mathematics, 1311:118–146, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082912/.

Hriljac:1987:SFP

[Hri87] Paul Hriljac. Splitting fields of principal homogeneous spaces. Lecture Notes in Mathematics, 1240:214–229, 1987. CODEN LN-MAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072982/.

Hairer:1989:CIPa

[HRL89a] Ernst Hairer, Michel Roche, and Christian Lubich. Convergence for index 1 problems. Lecture Notes in Mathematics, 1409:23– 29, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093950/.

Hairer:1989:CIPb

[HRL89b] Ernst Hairer, Michel Roche, and Christian Lubich. Convergence for index 2 problems. Lecture Notes in Mathematics, 1409:30– 54, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093951/.

Hairer:1989:CIPc

[HRL89c] Ernst Hairer, Michel Roche, and Christian Lubich. Convergence for index 3 problems. Lecture Notes in Mathematics, 1409:71– 91, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093953/.

Hairer:1989:DDA

[HRL89d] Ernst Hairer, Michel Roche, and Christian Lubich. Description of differential-algebraic problems. Lecture Notes in Mathematics, 1409:1–13, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093948/.

Hairer:1989:EDA

[HRL89e] Ernst Hairer, Michel Roche, and Christian Lubich. Examples of differential-algebraic systems and their solution. Lecture Notes in Mathematics, 1409:106–123, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093956/.

Hairer:1989:LEE

[HRL89f] Ernst Hairer, Michel Roche, and Christian Lubich. Local error estimation. Lecture Notes in Mathematics, 1409:99–105, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093955/.

Hairer:1989:OCR

[HRL89g] Ernst Hairer, Michel Roche, and Christian Lubich. Order conditions of Runge–Kutta methods for index 2 systems. Lecture Notes in Mathematics, 1409:55–70, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093952/.

Hairer:1989:RKM

[HRL89h] Ernst Hairer, Michel Roche, and Christian Lubich. Runge–Kutta methods for differential-algebraic equations. Lecture Notes in Mathematics, 1409:14–22, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093949/.

Hairer:1989:SNS

[HRL89i] Ernst Hairer, Michel Roche, and Christian Lubich. Solution of nonlinear systems by simplified Newton. Lecture Notes in Mathematics, 1409:92–98, 1989. CODEN LNMAA2. ISBN 3-540-51860-6 (print), 3-540-46832-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093954/.

Hrushovski:1987:LMR

[Hru87] Ehud Hrushovski. Locally modular regular types. Lecture Notes in Mathematics, 1292:132–164, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082236/.

Harder:1985:SVH

[HS85] G. Harder and N. Schappacher. Special values of Hecke Lfunctions and abelian integrals. Lecture Notes in Mathematics, 1111:17–49, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084583/.

Hulek:1986:APC

[HS86] K. Hulek and S. A. Strømme. Appendix to the paper “Complete families of stable vector bundles over P”. Lecture Notes in Mathematics, 1194:34–40, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076993/.

Hajela:1987:MES

[HS87a] D. Hajela and B. Smith. On the maximum of an exponential sum of the Möbius function. Lecture Notes in Mathematics, 1240: 145–164, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072979/.

Haslinger:1987:RDW

[HS87b] F. Haslinger and M. Smejkal. Representation and duality in weighted Fréchet spaces of entire functions. Lecture Notes in Mathematics, 1275:168–196, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078351/.

Herzog:1987:GGI

[HS87c] Jürgen Herzog and Herbert Sanders. The Grothendieck group of invariant rings and of simple hypersurface singularities. Lecture Notes in Mathematics, 1273:134–149, 1987. CODEN LN-MAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078840/.

Hinich:1987:HLH

[HS87d] V. A. Hinich and V. V. Schechtman. On homotopy limit of homotopy algebras. Lecture Notes in Mathematics, 1289:240–264, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078370/.

Hinton:1987:ASS

[HS87e] D. B. Hinton and J. K. Shaw. Asymptotics of solutions and spectra of perturbed periodic Hamiltonian systems. Lecture Notes in Mathematics, 1285:169–174, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080594/.

Hawkins:1988:RRO

[HS88] Jane M. Hawkins and Cesar E. Silva. Remarks on recurrence and orbit equivalence of nonsingular endomorphisms. Lecture Notes in Mathematics, 1342:281–290, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082837/.

Hammer:1989:QFB

[HS89a] Peter L. Hammer and Bruno Simeone. Quadratic functions of binary variables. Lecture Notes in Mathematics, 1403:1–56, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083461/.

Hunt:1989:MCF

[HS89b] Fern Hunt and Francis Sullivan. Methods of computing fractal dimensions. Lecture Notes in Mathematics, 1394:83–95, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086754/.

Hermida:1988:SCS

[HSG88] José A. Hermida and Tomás Sánchez-Giralda. Some criteria for solvability of systems of linear equations over modules. Lecture Notes in Mathematics, 1328:122–134, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100921/.

Hsiang:1987:LTG

[Hsi87] Wu-Yi Hsiang. Lie transformation groups and differential geometry. Lecture Notes in Mathematics, 1255:34–52, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077679/.

Hirzebruch:1985:ABP

[HSS85] Friedrich Hirzebruch, Joachim Schwermer, and Silke Suter, editors. Arbeitstagung Bonn 1984: Proceedings of the meeting held by the Max-Planck-Institut für Mathematik, Bonn June 15–22, 1984, volume 1111 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1111. URL http:

//link.springer.com/book/10.1007/BFb0084581;   http://
www.springerlink.com/content/978-3-540-39298-9.

Hsiao:1985:IEF

[HSW85] George C. Hsiao, Ernst P. Stephan, and Wolfgang L. Wendland. An integral equation formulation for a boundary value problem of elasticity in the domain exterior to an arc. Lecture Notes in Mathematics, 1121:153–165, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076269/.

Hackbusch:1986:MMI

[HT86] Wolfgang Hackbusch and Ulrich Trottenberg, editors. Multigrid Methods II: Proceedings of the 2nd European Conference on Multigrid Methods held at Cologne, October 1–4, 1985, volume 1228 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1228. URL http://link.springer.com/book/10.1007/
BFb0072636;   http://www.springerlink.com/content/978-
3-540-47372-5.

Hayashi:1987:RSP

[HT87] Nakao Hayashi and Yoshio Tsutsumi. Remarks on the scattering problem for nonlinear Schrödinger equations. Lecture Notes in Mathematics, 1285:162–168, 1987. CODEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080593/.

Hu:1987:IPR

[Hu87] Yi Hu. The imbedding problem of Riemannian globally symmetric spaces of the compact type. Lecture Notes in Mathematics, 1255:53–66, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077680/.

Hu:1988:NED

[Hu88] Yao-Zhong Hu. Un nouvel exemple de distribution de Hida. (French) []. Lecture Notes in Mathematics, 1321:82–84, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084120/.

Huang:1986:CLS

[Hua86] Jianhua Huang. A characterization of the 2-local subgroups of Rudvalis group and Higman–Sims group. Lecture Notes in Mathematics, 1185:289–307, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076180/.

Huang:1988:IRS

[Hua88] Yumin Huang. On interior regularity of solutions of a class of hypoelliptic equations. Lecture Notes in Mathematics, 1306:84– 92, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082926/.

Huber:1986:TRF

[Hub86] Catherine Huber. Théorie de la robustesse. (French) [Robustness theory]. Lecture Notes in Mathematics, 1215:1–128, 1986. CODEN LNMAA2. ISBN 3-540-16822-2 (print), 3-540-47092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075872/.

Hubl:1989:DFH

[Hüb89a] Reinhold Hübl. Differential forms and Hochschild homology. Lecture Notes in Mathematics, 1368:28–46, 1989. CODEN LN-MAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098409/.

Hubl:1989:HHH

[Hüb89b] Reinhold Hübl. The Hochschild homology and the Hochschild cohomology of a topological algebra. Lecture Notes in Mathematics, 1368:8–27, 1989. CODEN LNMAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098408/.

Hubl:1989:I

[Hüb89c] Reinhold Hübl. Introduction. Lecture Notes in Mathematics, 1368: 1–7, 1989. CODEN LNMAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098407/.

Hubl:1989:TRH

[Hüb89d] Reinhold Hübl. The topological residue homomorphism. Lecture Notes in Mathematics, 1368:84–93, 1989. CODEN LN-MAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098413/.

Hubl:1989:TFR

[Hüb89e] Reinhold Hübl. Trace formulas for residues of differential forms. Lecture Notes in Mathematics, 1368:94–105, 1989. CODEN LN-MAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098414/.

Hubl:1989:TCI

[Hüb89f] Reinhold Hübl. Traces in complete intersections. Lecture Notes in Mathematics, 1368:70–83, 1989. CODEN LNMAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098412/.

Hubl:1989:THH

[Hüb89g] Reinhold Hübl. Traces in Hochschild homology. Lecture Notes in Mathematics, 1368:47–55, 1989. CODEN LNMAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098410/.

Hubl:1989:TDF

[Hüb89h] Reinhold Hübl. Traces of differential forms. Lecture Notes in Mathematics, 1368:56–69, 1989. CODEN LNMAA2. ISBN 3-540-50985-2 (print), 3-540-46125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098411/.

Hudry:1986:DAD

[Hud86] A. Hudry. Deux applications du principe d’additivité. (French) []. Lecture Notes in Mathematics, 1220:118–129, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099509/.

Hudson:1988:ATQ

[Hud88] R. L. Hudson. Algebraic theory of quantum diffusions. Lecture Notes in Mathematics, 1325:113–124, 1988. CODEN LN-MAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077920/.

Hudson:1989:QDA

[Hud89] R. L. Hudson. Quantum diffusions on the algebra of all bounded operators on a Hilbert space. Lecture Notes in Mathematics, 1396:256–269, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083556/.

Hulek:1985:CIC

[Hul85] Klaus Hulek. Complete intersection curves, the splitting of the normal bundle and the Veronese surface. Lecture Notes in Mathematics, 1124:132–145, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074999/.

Hurder:1988:ETF

[Hur88] Steven Hurder. Ergodic theory of foliations and a theorem of Sacksteder. Lecture Notes in Mathematics, 1342:291–328, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082838/.

Husch:1987:EC

[Hus87] L. S. Husch. Embedding of continua. Lecture Notes in Mathematics, 1283:125–130, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081423/.

Husseini:1989:CGA

[Hus89] S. Y. Husseini. Coincidences for Grassmannian and associated spaces. Lecture Notes in Mathematics, 1411:124–129, 1989. CO-DEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086446/.

Hausmann:1988:MCP

[HV88] Jean-Claude Hausmann and Pierre Vogel. Manifold categories of Poincaré duality spaces. Lecture Notes in Mathematics, 1350: 241–258, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081478/.

Hebisch:1988:AMI

[HvL88] U. Hebisch and L. C. A. van Leeuwen. On additively or multiplicatively idempotent semirings and partial orders. Lecture Notes in Mathematics, 1320:154–161, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083428/.

Hendriksen:1988:EIZ

[HvR88] E. Hendriksen and H. van Rossum. Electrostatic interpretation of zeros. Lecture Notes in Mathematics, 1329:241–250, 1988. CO-DEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0083363/.

Hamann:1988:ASE

[HW88] U. Hamann and G. Wildenhain. Approximation by solutions of elliptic boundary value problems in various function spaces. Lecture Notes in Mathematics, 1302:247–252, 1988. CODEN LN-MAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078878/.

Hinman:1985:PMO

[HZ85] Peter G. Hinman and Stathis Zachos. Probabilistic machines, oracles, and quantifiers. Lecture Notes in Mathematics, 1141:159– 192, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076220/.

Hain:1987:GUV

[HZ87a] Richard M. Hain and Steven Zucker. A guide to unipotent variations of mixed Hodge structure. Lecture Notes in Mathematics, 1246:92–106, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077532/.

Hain:1987:TMH

[HZ87b] Richard M. Hain and Steven Zucker. Truncations of mixed Hodge complexes. Lecture Notes in Mathematics, 1246:107–114, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077533/.

Iarrobino:1985:CAC

[Iar85] Anthony Iarrobino, Jr. Compressed algebras and components of the punctual Hilbert scheme. Lecture Notes in Mathematics, 1124:146–165, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075000/.

Ibragimov:1985:TLP

[Ibr85] I. A. Ibragimov. Théorèmes limites pour les marches aléatoires. (French) []. Lecture Notes in Mathematics, 1117:199–297, 1985. CODEN LNMAA2. ISBN 3-540-15203-2 (print), 3-540-39316-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099422/.

Ichihara:1986:EPS

[Ich86] Kanji Ichihara. Explosion problems for symmetric diffusion processes. Lecture Notes in Mathematics, 1203:75–89, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076874/.

Ichikawa:1987:SPS

[Ich87] Akira Ichikawa. The separation principle for stochastic differential equations with unbounded coefficients. Lecture Notes in Mathematics, 1236:164–171, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072888/.

Ida:1989:GSS

[Idà89] Monica Idà. Generating six skew lines in P^{3}. Lecture Notes in Mathematics, 1389:112–127, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085929/.

Iesan:1987:AM

[Ieş87a] Dorin Ieşan. Anisotropic materials. Lecture Notes in Mathematics, 1279:44–67, 1987. CODEN LNMAA2. ISBN 3-540-18361-2 (print), 3-540-47911-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078754/.

Iesan:1987:HM

[Ieş87b] Dorin Ieşan. Heterogeneous media. Lecture Notes in Mathematics, 1279:68–94, 1987. CODEN LNMAA2. ISBN 3-540-18361-2 (print), 3-540-47911-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078755/.

Iesan:1987:RSV

[Ieş87c] Dorin Ieşan. The relaxed Saint-Venant problem. Lecture Notes in Mathematics, 1279:1–30, 1987. CODEN LNMAA2. ISBN 3-540-18361-2 (print), 3-540-47911-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078752/.

Iesan:1987:SVP

[Ieş87d] Dorin Ieşan. Saint-Venant’s problem for Cosserat elastic bodies. Lecture Notes in Mathematics, 1279:95–147, 1987. CODEN LN-MAA2. ISBN 3-540-18361-2 (print), 3-540-47911-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078756/.

Iesan:1987:TLC

[Ieş87e] Dorin Ieşan. Theory of loaded cylinders: The problems of Almansi and Michell. Lecture Notes in Mathematics, 1279:31–43, 1987. CODEN LNMAA2. ISBN 3-540-18361-2 (print), 3-540-47911-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078753/.

Ito:1986:SPT

[IH86] Kiyosi Itô and Takeyuki Hida, editors. Stochastic Processes and Their Applications: Proceedings of the International Conference held in Nagoya, July 2–6, 1985, volume 1203 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1203. URL http://link.springer.com/book/10.1007/
BFb0076868;   http://www.springerlink.com/content/978-
3-540-39852-3.

Ikeda:1988:STA

[IK88] Nobuyuki Ikeda and Shigeo Kusuoka. Short time asymptotics for fundamental solutions of diffusion equations. Lecture Notes in Mathematics, 1322:37–49, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077865/.

Ikegami:1988:CHS

[Ike88] Teruo Ikegami. Compactifications of harmonic spaces and Hunt processes. Lecture Notes in Mathematics, 1351:201–213, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081254/.

Isbell:1988:RLG

[IKPR88] J. Isbell, I. Kříž, A. Pultr, and J. Rosický. Remarks on localic groups. Lecture Notes in Mathematics, 1348:154–172, 1988. CO-DEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081357/.

Ilinskii:1985:NMI

[Il’85] A. I. Il’inskiĭ. Normality of a multidimensional infinitely divisible distribution which coincides with a normal distribution in a cone. Lecture Notes in Mathematics, 1155:47–59, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074812/.

Illman:1986:PFE

[Ill86] Sören Illman. A product formula for equivariant Whitehead torsion and geometric applications. Lecture Notes in Mathematics, 1217:123–142, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072819/.

Illan:1988:RAF

[Ill88] J. Illán. On the rational approximation of H^{P} functions in the L_{P} (μ) metric. Lecture Notes in Mathematics, 1354:155–163, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089592/.

Illman:1989:ICR

[Ill89] Sören Illman. The isomorphism class of a representation of a compact Lie group is determined by the equivariant simple-homotopy type of the representation. Lecture Notes in Mathematics, 1375: 98–110, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085602/.

Ismail:1988:ADH

[ILV88] Mourad E. H. Ismail, Jean Letessier, and Galliano Valent. Associated dual Hahn polynomials. Lecture Notes in Mathematics, 1329:251–254, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083364/.

Iwase:1989:HHA

[IM89] Norio Iwase and Mamoru Mimura. Higher homotopy associativity. Lecture Notes in Mathematics, 1370:193–220, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085229/.

Imkeller:1988:BPD

[Imk88a] Peter Imkeller. Basics; processes depending on a parameter. Lecture Notes in Mathematics, 1308:28–65, 1988. CODEN LN-MAA2. ISBN 3-540-19233-6 (print), 3-540-39148-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078099/.

Imkeller:1988:I

[Imk88b] Peter Imkeller. Introduction. Lecture Notes in Mathematics, 1308: 1–22, 1988. CODEN LNMAA2. ISBN 3-540-19233-6 (print), 3-540-39148-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078097/.

Imkeller:1988:JMT

[Imk88c] Peter Imkeller. Jumps of martingales and their compensation. Lecture Notes in Mathematics, 1308:99–130, 1988. CODEN LN-MAA2. ISBN 3-540-19233-6 (print), 3-540-39148-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078101/.

Imkeller:1988:NC

[Imk88d] Peter Imkeller. Notations and conventions. Lecture Notes in Mathematics, 1308:23–27, 1988. CODEN LNMAA2. ISBN 3-540-19233-6 (print), 3-540-39148-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078098/.

Imkeller:1988:QVS

[Imk88e] Peter Imkeller. Quadratic variation and structure of martingales. Lecture Notes in Mathematics, 1308:131–168, 1988. CODEN LN-MAA2. ISBN 3-540-19233-6 (print), 3-540-39148-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078102/.

Imkeller:1988:TPP

[Imk88f] Peter Imkeller. Two-parameter processes. Lecture Notes in Mathematics, 1308:66–98, 1988. CODEN LNMAA2. ISBN 3-540-19233-6 (print), 3-540-39148-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078100/.

Ihrig:1988:DFN

[IMW88] Edwin Ihrig, Douglas McMahon, and Ta-Sun Wu. Distal flows of non Abelian groups with finite codimension. Lecture Notes in Mathematics, 1342:329–341, 1988. CODEN LNMAA2. ISBN 3-540-50174-6 (print), 3-540-45946-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082839/.

Iserles:1985:BOP

[IN85] A. Iserles and S. P. Nørsett. Bi-orthogonal polynomials. Lecture Notes in Mathematics, 1171:92–100, 1985. CODEN LN-MAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076534/.

Its:1986:ASDb

[IN86a] Alexander R. Its and Victor Yu. Novokshenov. Asymptotic solution to a direct problem of the monodromy theory for the system (1.26). Lecture Notes in Mathematics, 1191:80–86, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076668/.

Its:1986:ASDa

[IN86b] Alexander R. Its and Victor Yu. Novokshenov. Asymptotic solution to a direct problem of the monodromy theory for the system (1.9). Lecture Notes in Mathematics, 1191:62–79, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076667/.

Its:1986:DEI

[IN86c] Alexander R. Its and Victor Yu. Novokshenov. The dynamics of electromagnetic impulse in a long laser amplifier. Lecture Notes in Mathematics, 1191:205–211, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076675/.

Its:1986:I

[IN86d] Alexander R. Its and Victor Yu. Novokshenov. Introduction. Lecture Notes in Mathematics, 1191:1–14, 1986. CODEN LN-MAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076662/.

Its:1986:IPM

[IN86e] Alexander R. Its and Victor Yu. Novokshenov. Inverse problem of the monodromy theory for the systems (1.9) and (1.26). Asymptotic analysis of integral equations of the inverse problem. Lecture Notes in Mathematics, 1191:46–61, 1986. CODEN LN-MAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076666/.

Its:1986:IDM

[IN86f] Alexander R. Its and Victor Yu. Novokshenov. The Isomonodromic Deformation Method in the Theory of Painlevé Equations, volume 1191 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 313 pp. LCCN QA3 .L28 no. 1191. URL http:

//link.springer.com/book/10.1007/BFb0076661;   http://
www.springerlink.com/content/978-3-540-39823-3.

Its:1986:IDSb

[IN86g] Alexander R. Its and Victor Yu. Novokshenov. Isomonodromic deformations of systems (1.9) and (1.26) and Painlevé equations of II and III types. Lecture Notes in Mathematics, 1191:43–45, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076665/.

Its:1986:IDSa

[IN86h] Alexander R. Its and Victor Yu. Novokshenov. Isomonodromic deformations of systems of linear ordinary differential equations with rational coefficients. Lecture Notes in Mathematics, 1191: 38–42, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076664/.

Its:1986:LTA

[IN86i] Alexander R. Its and Victor Yu. Novokshenov. Large-time asymptotics of the solution of the Cauchy problem for MKdV equation. Lecture Notes in Mathematics, 1191:173–204, 1986. CODEN LN-MAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076674/.

Its:1986:MSPa

[IN86j] Alexander R. Its and Victor Yu. Novokshenov. The manifold of solutions of Painlevé II equation decreasing as χ → −∞. Parametrization of their asymptotics through the monodromy data. Ablowitz–Segur connection formulae for real-valued solutions decreasing exponentially as χ → +∞. Lecture Notes in Mathematics, 1191:87–93, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076669/.

Its:1986:MSPc

[IN86k] Alexander R. Its and Victor Yu. Novokshenov. The manifold of solutions to Painlevé II equation increasing as χ → +∞. The expression of their asymptotics through the monodromy data. The connection formulae for pure imaginary solutions. Lecture Notes in Mathematics, 1191:109–123, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076671/.

Its:1986:MSPb

[IN86l] Alexander R. Its and Victor Yu. Novokshenov. The manifold of solutions to Painlevé III equation. The connection formulae for the asymptotics of real-valued solutions to the Cauchy problem. Lecture Notes in Mathematics, 1191:94–108, 1986. CODEN LN-MAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076670/.

Its:1986:MDS

[IN86m] Alexander R. Its and Victor Yu. Novokshenov. Monodromy data for the systems of linear ordinary differential equations with rational coefficients. Lecture Notes in Mathematics, 1191:15–37, 1986.

CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076663/.

Its:1986:MPR

[IN86n] Alexander R. Its and Victor Yu. Novokshenov. The movable poles of real-valued solutions to Painlevé II equation and the eigenfunctions of anharmonic oscillator. Lecture Notes in Mathematics, 1191:124–146, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076672/.

Its:1986:MPS

[IN86o] Alexander R. Its and Victor Yu. Novokshenov. The movable poles of the solutions of Painlevé III equation and their connection with Mathieu functions. Lecture Notes in Mathematics, 1191:147–172, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076673/.

Its:1986:QMT

[IN86p] Alexander R. Its and Victor Yu. Novokshenov. Quasiclassical mode of the three-dimensional wave collapse. Lecture Notes in Mathematics, 1191:216–229, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076677/.

Its:1986:SLT

[IN86q] Alexander R. Its and Victor Yu. Novokshenov. The scaling limit in two-dimensional Ising model. Lecture Notes in Mathematics, 1191:212–215, 1986. CODEN LNMAA2. ISBN 3-540-16483-9 (print), 3-540-39823-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076676/.

Inoue:1989:EWT

[Ino89] Tsuyoshi Inoue. The equivariant Whitehead torsions of equivariant homotopy equivalences between the unit spheres of representations of cyclic groups. Lecture Notes in Mathematics, 1375: 111–125, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085603/.

Sugiyama:1988:LBE

[iS88] Ken ichi Sugiyama. Lower bounds of the essential spectrum of the Laplace–Beltrami operator and its application to complex geometry. Lecture Notes in Mathematics, 1339:226–247, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083058/.

Israel:1989:CFM

[Isr89] Werner Israel. Covariant fluid mechanics and thermodynamics: an introduction. Lecture Notes in Mathematics, 1385:152–210, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084030/.

Igusa:1986:NCF

[IT86] Kiyoshi Igusa and Gordana Todorov. A numerical characterization of finite Auslander–Reiten quivers. Lecture Notes in Mathematics, 1177:181–198, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075265/.

Ito:1988:MCS

[Itō88] Kiyosi Itō. Malliavin calculus on a Segal space. Lecture Notes in Mathematics, 1322:50–72, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077866/.

Ito:1989:DRA

[Ito89] Noboru Ito. Doubly regular asymmetric digraphs with rank 5 automorphism groups. Lecture Notes in Mathematics, 1398:94– 99, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086244/.

Ivanoff:1986:LTM

[Iva86] Gail Ivanoff. Limit theorems for multitype branching random walks. Lecture Notes in Mathematics, 1212:187–194, 1986. CO-DEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076248/.

Ivanov:1988:ATM

[Iva88] N. V. Ivanov. Automorphisms of Teichmüller modular groups. Lecture Notes in Mathematics, 1346:199–270, 1988. CODEN LN-MAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082778/.

Istad:1986:SNC

[IW86] Roy M. Istad and Haakon Waadeland. A strategy for numerical computation of limits regions. Lecture Notes in Mathematics, 1199:37–47, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075933/.

Iwasaki:1987:GHP

[Iwa87] Chisato Iwasaki. Gevrey-hypoellipticity and pseudo-differential operators on Gevrey class. Lecture Notes in Mathematics, 1256: 281–293, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077746/.

Jacod:1985:TLP

[Jac85a] J. Jacod. Théorèmes limite pour les processus. (French) []. Lecture Notes in Mathematics, 1117:298–406, 1985. CODEN LN-MAA2. ISBN 3-540-15203-2 (print), 3-540-39316-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099423/.

Jacod:1985:GFP

[Jac85b] Jean Jacod. Grossissement de filtration et processus d’Ornstein– Uhlenbeck généralisé. (French) []. Lecture Notes in Mathematics, 1118:36–44, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075769/.

Jacod:1985:GIH

[Jac85c] Jean Jacod. Grossissement initial, hypothèse (H′) et théorème de Girsanov. (French) []. Lecture Notes in Mathematics, 1118: 15–35, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075768/.

Jacobsen:1986:CLP

[Jac86a] Lisa Jacobsen. On the convergence of limit periodic continued fractions K(a_{n}/1), where a_{n} → −1/4. Part II. Lecture Notes in Mathematics, 1199:48–58, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075934/.

Jacobsen:1986:TSC

[Jac86b] Lisa Jacobsen. A theorem on simple convergence regions for continued fractions K(a_{n}/1). Lecture Notes in Mathematics, 1199: 59–66, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075935/.

Jacka:1987:MIM

[Jac87] S. D. Jacka. A maximal inequality for martingale local times. Lecture Notes in Mathematics, 1247:221–229, 1987. CODEN LN-MAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077637/.

Jackson:1988:APO

[Jac88a] Steve Jackson. Ad and the projective ordinals. Lecture Notes in Mathematics, 1333:117–220, 1988. CODEN LNMAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084974/.

Jackson:1988:FM

[Jac88b] Steve Jackson. Front matter. Lecture Notes in Mathematics, 1337: ??, 1988. CODEN LNMAA2. ISBN 3-540-50080-4 (print), 3-540-45920-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-45920-0/1.pdf.

Jacobowitz:1988:SHP

[Jac88c] Howard Jacobowitz. Systems of homogeneous partial differential equations with few solutions. Lecture Notes in Mathematics, 1324:127–136, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100788/.

Jacka:1989:NGL

[Jac89] S. D. Jacka. A note on the good lambda inequalities. Lecture Notes in Mathematics, 1372:57–65, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083960/.

Jagannathan:1985:PRF

[Jag85] R. Jagannathan. On projective representations of finite Abelian groups. Lecture Notes in Mathematics, 1122:130–139, 1985. CO-DEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075756/.

Jain:1985:ILD

[Jai85] Naresh C. Jain. An introduction to large deviations. Lecture Notes in Mathematics, 1153:273–296, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074955/.

Jajte:1985:AUC

[Jaj85a] Ryszard Jajte. Almost uniform convergence in von Neumann algebras. Lecture Notes in Mathematics, 1110:4–10, 1985. CODEN LNMAA2. ISBN 3-540-13915-X (print), 3-540-39139-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101455/.

Jajte:1985:CCE

[Jaj85b] Ryszard Jajte. Convergence of conditional expectations and martingales in von Neumann algebras. Lecture Notes in Mathematics, 1110:57–80, 1985. CODEN LNMAA2. ISBN 3-540-13915-X (print), 3-540-39139-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101457/.

Jajte:1985:ET

[Jaj85c] Ryszard Jajte. Ergodic theorems. Lecture Notes in Mathematics, 1110:11–56, 1985. CODEN LNMAA2. ISBN 3-540-13915-X (print), 3-540-39139-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101456/.

Jajte:1985:I

[Jaj85d] Ryszard Jajte. Introduction. Lecture Notes in Mathematics, 1110: 1–3, 1985. CODEN LNMAA2. ISBN 3-540-13915-X (print), 3-540-39139-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101454/.

Jajte:1985:SLL

[Jaj85e] Ryszard Jajte. Strong laws of large numbers in von Neumann algebras. Lecture Notes in Mathematics, 1110:81–105, 1985. CODEN LNMAA2. ISBN 3-540-13915-X (print), 3-540-39139-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101458/.

Jajte:1985:SLT

[Jaj85f] Ryszard Jajte. Strong Limit Theorems in Non-Commutative Probability, volume 1110 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-13915-X (print), 3-540-39139-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 152 pp. LCCN QA3 .L28 no. 1110. URL http:

//link.springer.com/book/10.1007/BFb0101453;   http://
www.springerlink.com/content/978-3-540-39139-5.

Jajte:1988:CSN

[Jaj88] Ryszard Jajte. Contraction semigroups in L^{2} over a von Neumann algebra. Lecture Notes in Mathematics, 1303:149–153, 1988.

CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078060/.

Jajte:1989:AFN

[Jaj89a] R. Jajte. Asymptotic formula for normal operators in noncommutative L_{2}-spaces. Lecture Notes in Mathematics, 1396:270– 278, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083557/.

Jajte:1989:SLL

[Jaj89b] R. Jajte. Strong laws of large numbers for several contractions in a von Neumann algebra. Lecture Notes in Mathematics, 1391: 125–139, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083386/.

Jameson:1985:TFC

[Jam85] Antony Jameson. Transonic flow calculations for aircraft. Lecture Notes in Mathematics, 1127:156–242, 1985. CODEN LN-MAA2. ISBN 3-540-15225-3 (print), 3-540-39391-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074531/.

Jameson:1986:MAC

[Jam86] Antony Jameson. Multigrid algorithms for compressible flow calculations. Lecture Notes in Mathematics, 1228:166–201, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072647/.

James:1989:TPP

[Jam89] I. M. James. Topology: past, present and future. Lecture Notes in Mathematics, 1370:1–9, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085213/.

Janssen:1986:ACS

[Jan86a] Arnold Janssen. Absolute continuity and singularity of distributions of dependent observations: Gaussian and exchangeable measures. Lecture Notes in Mathematics, 1210:108–124, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077177/.

Jantzen:1986:MRR

[Jan86b] J. C. Jantzen. Modular representations of reductive groups. Lecture Notes in Mathematics, 1185:118–154, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076173/.

Jantzen:1987:RLA

[Jan87] J. C. Jantzen. Restricted Lie algebra cohomology. Lecture Notes in Mathematics, 1271:91–108, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079234/.

Janson:1989:FD

[Jan89] Svante Janson. On functions with derivatives in H^{1}. Lecture Notes in Mathematics, 1384:193–201, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086803/.

Jarosz:1985:PMI

[Jar85a] Krzysztof Jarosz. I. Perturbations of multiplications and ontoisomorphisms. Lecture Notes in Mathematics, 1120:4–35, 1985. CODEN LNMAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076887/.

Jarosz:1985:II

[Jar85b] Krzysztof Jarosz. II. Into-isomorphisms. Lecture Notes in Mathematics, 1120:36–44, 1985. CODEN LNMAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076888/.

Jarosz:1985:IIS

[Jar85c] Krzysztof Jarosz. III. Isometries in semisimple, commutative Banach algebras. Lecture Notes in Mathematics, 1120:45–55, 1985. CODEN LNMAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076889/.

Jarosz:1985:IPO

[Jar85d] Krzysztof Jarosz. IV. Perturbations of operator algebras. Lecture Notes in Mathematics, 1120:56–76, 1985. CODEN LN-MAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076890/.

Jarosz:1985:PBA

[Jar85e] Krzysztof Jarosz. Perturbations of Banach Algebras, volume 1120 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 118 pp. LCCN QA3 .L28 no. 1120. URL http:

//link.springer.com/book/10.1007/BFb0076885;   http://
www.springerlink.com/content/978-3-540-39370-2.

Jarosz:1985:P

[Jar85f] Krzysztof Jarosz. Preliminaries. Lecture Notes in Mathematics, 1120:1–3, 1985. CODEN LNMAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076886/.

Jarosz:1985:VS

[Jar85g] Krzysztof Jarosz. V. Stability. Lecture Notes in Mathematics, 1120:77–110, 1985. CODEN LNMAA2. ISBN 3-540-15218-0 (print), 3-540-39370-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076891/.

Jarre:1989:MAC

[Jar89] F. Jarre. On the method of analytic centers for solving smooth convex programs. Lecture Notes in Mathematics, 1405:69–86, 1989. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Jaworowski:1986:BOF

[Jaw86] Jan Jaworowski. Balanced orbits for fibre preserving maps of S^{1} and S^{3} actions. Lecture Notes in Mathematics, 1217:143– 150, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072820/.

Jensen:1986:CMA

[Jen86] Arne Jensen. Commutator methods and asymptotic completeness for one-dimensional Stark effect Hamiltonians. Lecture Notes in Mathematics, 1218:151–166, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073048/.

Jenkins:1987:SEH

[Jen87] James A. Jenkins. Some estimates for harmonic measures. Lecture Notes in Mathematics, 1275:210–214, 1987. CODEN LN-MAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078353/.

Jeulin:1985:ATD

[Jeu85] T. Jeulin. Application de la théorie du grossissement à l’étude des temps locaux browniens. (French) []. Lecture Notes in Mathematics, 1118:197–304, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075775/.

Jezek:1985:SIS

[Jež85] Jaroslav Ježek. Subdirectly irreducible and simple Boolean algebras with endomorphisms. Lecture Notes in Mathematics, 1149: 150–162, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098462/.

Jiang:1989:SMB

[Jia89] Boju Jiang. Surface maps and braid equations, I. Lecture Notes in Mathematics, 1369:125–141, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087529/.

Jin:1988:KFL

[Jin88a] Gyo Taek Jin. On Kojima’s η-function of links. Lecture Notes in Mathematics, 1350:14–30, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081466/.

Jin:1988:CYP

[Jin88b] Zhiren Jin. A counterexample to the Yamabe problem for complete noncompact manifolds. Lecture Notes in Mathematics, 1306: 93–101, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082927/.

Jacobsen:1986:FRC

[JJW86] Lisa Jacobsen, William B. Jones, and Haakon Waadeland. Further results on the computation of incomplete gamma functions. Lecture Notes in Mathematics, 1199:67–89, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075936/.

Jacobsen:1987:CAC

[JJW87] L. Jacobsen, W. B. Jones, and H. Waadeland. Convergence acceleration for continued fractions K(a_{n}/1), where a_{n} → ∞. Lecture Notes in Mathematics, 1237:177–187, 1987. CODEN LN-MAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072463/.

Jona-Lasinio:1985:SRA

[JL85] G. Jona-Lasinio. Some recent applications of stochastic processes in quantum mechanics. Lecture Notes in Mathematics, 1159:230– 241, 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080335/.

Johnson:1987:LEF

[JLS87] William B. Johnson, Joram Lindenstrauss, and Gideon Schechtman. On Lipschitz embedding of finite metric spaces in low dimensional normed spaces. Lecture Notes in Mathematics, 1267: 177–184, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078145/.

Jarchow:1985:WCO

[JM85] Hans Jarchow and Urs Matter. On weakly compact operators on ℓ (k)-spaces. Lecture Notes in Mathematics, 1166:80–88, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074696/.

Jones:1986:SIG

[JM86a] A. Jones and G. O. Michler. On the structure of the integral Green ring of a cyclic group of order p^{2}. Lecture Notes in Mathematics, 1178:98–108, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075291/.

Journe:1986:MOB

[JM86b] J.-L. Journé and P.-A. Meyer. Une martingale d’opérateurs bornés, non représentable en intégrale stochastique. (French) []. Lecture Notes in Mathematics, 1204:313–316, 1986. CODEN LN-MAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/BFb0075720/. See re-

mark [Par86c].

Jackowski:1989:HAC

[JM89a] Stefan Jackowski and James E. McClure. Homotopy approximations for classifying spaces of compact Lie groups. Lecture Notes in Mathematics, 1370:221–234, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085230/.

Jaffard:1989:OF

[JM89b] S. Jaffard and Y. Meyer. Les ondelettes. (French) []. Lecture Notes in Mathematics, 1384:182–192, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086802/.

Janovsky:1986:BAS

[JMN86] V. Janovský, I. Marek, and J. Neuberg. Bifurcation analysis of stimulated Brillouin scattering. Lecture Notes in Mathematics, 1192:309–314, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076086/.

Jones:1986:SFP

[JNT86] W. B. Jones, O. Njåstad, and W. J. Thron. Schur fractions, Perron–Carathéodory fractions and Szegő polynomials, a survey. Lecture Notes in Mathematics, 1199:127–158, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075938/.

Jones:1987:PCC

[JNT87] William B. Jones, Olav Njåstad, and W. J. Thron. Perron– Carathéodory continued fractions. Lecture Notes in Mathematics, 1237:188–206, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072464/.

Jockusch:1985:GRE

[Joc85] Carl G. Jockusch, Jr. Genericity for recursively enumerable sets. Lecture Notes in Mathematics, 1141:203–232, 1985. CODEN LN-MAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076222/.

Johnson:1986:ECG

[Joh86] Russell A. Johnson. An example concerning the geometric significance of the rotation number — integrated density of states. Lecture Notes in Mathematics, 1186:216–226, 1986. CODEN LN-MAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076841/.

Johnson:1988:D

[Joh88a] B. E. Johnson. Derivations from L^{1}(G) into L^{1}(G) and L^{∞}(G). Lecture Notes in Mathematics, 1359:191–198, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086599/.

Johnson:1988:HBS

[Joh88b] William B. Johnson. Homogeneous Banach spaces. Lecture Notes in Mathematics, 1317:201–203, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081741/.

John:1989:SQL

[Joh89a] Fritz John. Solutions of quasi-linear wave equations with small initial data. Lecture Notes in Mathematics, 1402:155–184, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083874/.

Johnson:1989:BGP

[Joh89b] Ellis L. Johnson. On binary group problems having the Fulkerson property. Lecture Notes in Mathematics, 1403:57–112, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083462/.

Jondrup:1985:HPA

[Jøn85] S. Jøndrup. Hereditary P.I. algebras. Lecture Notes in Mathematics, 1146:405–420, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074549/.

Jones:1988:CSV

[Jon88] Peter R. Jones. Congruence semimodular varieties of semigroups. Lecture Notes in Mathematics, 1320:162–171, 1988. CODEN LN-MAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083429/.

Jones:1989:SFC

[Jon89] Peter W. Jones. Square functions, Cauchy integrals, analytic capacity, and harmonic measure. Lecture Notes in Mathematics, 1384:24–68, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086793/.

Joshi:1985:SRS

[Jos85a] V. S. Joshi. Sums of reciprocals of some non-vanishing multiplicative functions. Lecture Notes in Mathematics, 1122:140–146, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075757/.

Jost:1985:LHM

[Jos85b] Jürgen Jost. Lectures on harmonic maps. Lecture Notes in Mathematics, 1161:118–192, 1985. CODEN LNMAA2. ISBN 3-540-16040-X (print), 3-540-39716-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075137/.

Joseph:1987:CAC

[Jos87a] A. Joseph. Corrigenda and addenda to “On the cyclicity of vectors associated with Duflo involutions”. Lecture Notes in Mathematics, 1243:181–188, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073022/. See [Jos87b].

Joseph:1987:CVA

[Jos87b] Anthony Joseph. On the cyclicity of vectors associated with Duflo involutions. Lecture Notes in Mathematics, 1243:144– 180, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/

BFb0073021/. See corrigenda and addenda [Jos87a].

Jost:1988:HMA

[Jos88] Jürgen Jost. Harmonic maps — analytic theory and geometric significance. Lecture Notes in Mathematics, 1357:264–296, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082871/.

Joyal:1986:FAE

[Joy86] André Joyal. Foncteurs analytiques et espèces de structures. (French) []. Lecture Notes in Mathematics, 1234:126–159, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072514/.

Jozefiak:1988:SS

[Józ88] Tadeusz Józefiak. Semisimple superalgebras. Lecture Notes in Mathematics, 1352:96–113, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082020/.

Jackowski:1986:TGP

[JP86] Stefan Jackowski and Krzysztof Pawałowski, editors. Transformation Groups Poznań 1985: Proceedings of a Symposium held in Poznań, July 5–9, 1985, volume 1217 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1217. URL http:

//link.springer.com/book/10.1007/BFb0072810;   http://
www.springerlink.com/content/978-3-540-47097-7.

Qiu:1988:ASC

[jQxQ88] Qing jiu Qiu and Si xin Qian. Analysis of C^{∞}-singularities for a class of operators with varying multiple characteristics. Lecture Notes in Mathematics, 1306:141–148, 1988. CODEN LN-MAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082930/.

Johnstone:1988:WPH

[JS88] Peter Johnstone and Shu-Hao Sun. Weak products and Hausdorff locales. Lecture Notes in Mathematics, 1348:173–193, 1988.

CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081358/.

Jerison:1987:SSO

[JSC87] David Jerison and Antonio Sánchez-Calle. Subelliptic, second order differential operators. Lecture Notes in Mathematics, 1277: 46–77, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078245/.

Jacobsen:1986:OCR

[JT86] L. Jacobsen and W. J. Thron. Oval convergence regions and circular limit regions for continued fractions K(a_{n}/1). Lecture Notes in Mathematics, 1199:90–126, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075937/.

Jacobsen:1989:JWH

[JTW89] Lisa Jacobsen, W. J. Thron, and Haakon Waadeland. Julius Worpitzky, his contributions to the analytic theory of continued fractions and his times. Lecture Notes in Mathematics, 1406:25– 47, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096164/.

Jurek:1985:RIR

[Jur85] Zbigniew J. Jurek. Random integral representation for another class of limit laws. Lecture Notes in Mathematics, 1153:297– 309, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074956/.

Jurek:1989:LSA

[Jur89] Zbigniew J. Jurek. Linear supports and absolute continuity of certain random integrals. Lecture Notes in Mathematics, 1391: 140–147, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083387/.

Jutila:1989:MVE

[Jut89] M. Jutila. Mean value estimates for exponential sums. Lecture Notes in Mathematics, 1380:120–136, 1989. CODEN LN-MAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086549/.

Julg:1985:GAI

[JV85] Pierre Julg and Alain Valette. Group actions on Irees and Kamenability. Lecture Notes in Mathematics, 1132:289–296, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074889/.

Jones:1989:PFE

[JW89] William B. Jones and Nancy J. Wyshinski. Positive T -fraction expansions for a family of special functions. Lecture Notes in Mathematics, 1406:48–66, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096165/.

Jeulin:1985:GFE

[JY85] Th. Jeulin and M. Yor, editors. Grossissements de filtrations: exemples et applications: Séminaire de Calcul Stochastique 1982/83 Université Paris VI, volume 1118 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1118. URL http:

//link.springer.com/book/10.1007/BFb0075765;   http://
www.springerlink.com/content/978-3-540-39339-9.

Kacur:1986:MRE

[Kač86] J. Kačur. Method of Rothe in evolution equations. Lecture Notes in Mathematics, 1192:23–34, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076049/.

Kagan:1985:PSL

[Kag85] A. M. Kagan. Partial sufficiency of linear forms. Lecture Notes in Mathematics, 1155:60–80, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074813/.

Kagan:1989:MAC

[Kag89a] A. M. Kagan. A multivariate analog of the Cramér theorem on components of the Gaussian distributions. Lecture Notes in Mathematics, 1412:68–77, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084165/.

Kagan:1989:RLT

[Kag89b] F. M. Kagan. A refinement of Lukacs theorems. Lecture Notes in Mathematics, 1412:78–82, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084166/.

Kahane:1988:DMS

[Kah88] Jean-Pierre Kahane. Désintegration des mesures selon la dimension et chaos multiplicatif. (French) []. Lecture Notes in Mathematics, 1359:199–208, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086600/.

Kaise:1987:ADF

[Kai87a] Tetsuo Kaise. Algèbres de division et GL_{n}. (French) []. Lecture Notes in Mathematics, 1252:99–201, 1987. CODEN LN-MAA2. ISBN 3-540-17827-9 (print), 3-540-47871-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077392/.

Kaise:1987:RWF

[Kai87b] Tetsuo Kaise. Représentations de Weil de GL_{2}. (French) []. Lecture Notes in Mathematics, 1252:1–98, 1987. CODEN LNMAA2. ISBN 3-540-17827-9 (print), 3-540-47871-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077391/.

Kaise:1987:RWA

[Kai87c] Tetsuo Kaise. Représentations de Weil et GL_{2} Algèbres de division et GL_{n}: (Vers les corps de classes galoisiens I, II). (French) [Weil representations and rmGL_{2} division algebras and rmGL_{n}: (Toward the body of Galois classes I, II)], volume 1252 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17827-9 (print), 3-540-47871-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 203 pp. LCCN QA3

.L28 no. 1252. URL http://link.springer.com/book/10.1007/
BFb0077390;   http://www.springerlink.com/content/978-
3-540-47871-3.

Kakuda:1989:RFQ

[Kak89] Yuzuru Kakuda. The role of a filter quantifier in set theory. Lecture Notes in Mathematics, 1388:16–95, 1989. CODEN LN-MAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083665/.

Kalashnikov:1985:RCE

[Kal85] V. V. Kalashnikov. The rate of convergence of Erlang distribution to the degenerate one. Lecture Notes in Mathematics, 1155: 95–101, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074815/.

Kalton:1986:DPV

[Kal86a] N. J. Kalton. Differentiability properties of vector valued functions. Lecture Notes in Mathematics, 1221:141–181, 1986. CODEN LNMAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099113/.

Kalton:1986:SAV

[Kal86b] N. J. Kalton. Some applications of vector-valued analytic and harmonic functions. Lecture Notes in Mathematics, 1221:114– 140, 1986. CODEN LNMAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099112/.

Kamishima:1985:PAH

[Kam85] Yoshinobu Kamishima. Proper actions on homogeneous spaces. Lecture Notes in Mathematics, 1167:168–179, 1985. CODEN LN-MAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075223/.

Kamata:1989:CNU

[Kam89a] Masayoshi Kamata. On the characteristic numbers of unitary semi-free S^{1}-manifolds. Lecture Notes in Mathematics, 1375:126– 131, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085604/.

Kamishima:1989:CCA

[Kam89b] Yoshinobu Kamishima. Conformal circle actions on 3-manifolds. Lecture Notes in Mathematics, 1375:132–144, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085605/.

Kanev:1985:TDG

[Kan85] Vasil Kanev. Theta divisors of generalized Prym varieties I. Lecture Notes in Mathematics, 1124:166–215, 1985. CODEN LN-MAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075001/.

Kanai:1986:AIR

[Kan86a] Masahiko Kanai. Analytic inequalities, and rough isometries between non-compact Riemannian manifolds. Lecture Notes in Mathematics, 1201:122–137, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075650/.

Kaniuth:1986:EMP

[Kan86b] Eberhard Kaniuth. Ergodic and mixing properties of measures on locally compact groups. Lecture Notes in Mathematics, 1210: 125–129, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077178/.

Kantor:1986:GPS

[Kan86c] William M. Kantor. Generalized polygons, SCABs and GABs. Lecture Notes in Mathematics, 1181:79–158, 1986. CODEN LN-MAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075513/.

Kang:1987:PAD

[Kan87] Li Shan Kang. Parallel algorithms and domain decomposition. Lecture Notes in Mathematics, 1297:61–75, 1987. CODEN LN-MAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Kanai:1988:TEG

[Kan88a] Masahiko Kanai. Tensorial ergodicity of geodesic flows. Lecture Notes in Mathematics, 1339:142–157, 1988. CODEN LN-MAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083053/.

Kanda:1988:PRB

[Kan88b] Mamoru Kanda. Problems on a relation between measures and corresponding potentials. Lecture Notes in Mathematics, 1344: 205–206, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103355.

Kanenobu:1989:UDS

[Kan89] Taizo Kanenobu. Untwisted deform-spun knots: Examples of symmetry-spun 2-knots. Lecture Notes in Mathematics, 1375:145– 167, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085606/.

Karni:1985:CNG

[Kar85] Shimy Karni. Covering numbers of groups of small order and sporadic groups. Lecture Notes in Mathematics, 1112:52–196, 1985. CODEN LNMAA2. ISBN 3-540-13916-8 (print), 3-540-39142-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072287/.

Karoubi:1989:CHC

[Kar89] Max Karoubi. Cyclic homology and characteristic classes of bundles with additional structures. Lecture Notes in Mathematics, 1370:235–242, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085231/.

Kasahara:1986:EPS

[Kas86] Yuji Kasahara. Extremal process as a substitution for ‘one-sided stable process with index 0’. Lecture Notes in Mathematics, 1203: 90–100, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076875/.

Kashiwara:1987:PLV

[Kas87] Masaki Kashiwara. Poincaré lemma for a variation of polarized Hodge structure. Lecture Notes in Mathematics, 1246:115– 124, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077534/.

Kasue:1988:HFG

[Kas88] Atsushi Kasue. Harmonic functions with growth conditions on a manifold of asymptotically nonnegative curvature I. Lecture Notes in Mathematics, 1339:158–181, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083054/.

Katai:1985:RPK

[Kát85a] I. Kátai. A remark on a paper of K. Ramachandra. Lecture Notes in Mathematics, 1122:147–152, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075758/.

Katre:1985:JST

[Kat85b] S. A. Katre. Jacobsthal sums in terms of quadratic partitions of a prime. Lecture Notes in Mathematics, 1122:153–162, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075759/.

Katsuda:1986:PPL

[Kat86a] Atsushi Katsuda. A pinching problem for locally homogeneous spaces. Lecture Notes in Mathematics, 1201:150–155, 1986. CO-DEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075652/.

Katz:1986:NCA

[Kat86b] Gabriel Katz. Normal combinatorics of G-actions on manifolds. Lecture Notes in Mathematics, 1217:167–182, 1986. CODEN LN-MAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072822/.

Katsuda:1988:DTC

[Kat88a] Atsushi Katsuda. Density theorems for closed orbits. Lecture Notes in Mathematics, 1339:182–202, 1988. CODEN LN-MAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083055/.

Katz:1988:IMP

[Kat88b] Sheldon Katz. Iteration of multiple point formulas and applications to conics. Lecture Notes in Mathematics, 1311:147–155, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082913/.

Kato:1989:SAC

[Kat89a] Mitsuyoshi Kato. On some Abelian complex reflection groups. Lecture Notes in Mathematics, 1375:168–182, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085607/.

Katsutani:1989:SSM

[Kat89b] Hiroaki Katsutani. Syntactical simulation of many-valued logic. Lecture Notes in Mathematics, 1388:96–119, 1989. CODEN LN-MAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083666/.

Kaufman:1987:HMR

[Kau87] Robert Kaufman. Hausdorff measures and removable sets. Lecture Notes in Mathematics, 1275:215–222, 1987. CODEN LN-MAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078354/.

Kawohl:1985:I

[Kaw85a] Bernhard Kawohl. Introduction. Lecture Notes in Mathematics, 1150:1–6, 1985. CODEN LNMAA2. ISBN 3-540-15693-3 (print), 3-540-39639-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075061/.

Kawohl:1985:MP

[Kaw85b] Bernhard Kawohl. Maximum principles. Lecture Notes in Mathematics, 1150:100–122, 1985. CODEN LNMAA2. ISBN 3-540-15693-3 (print), 3-540-39639-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075063/.

Kawohl:1985:R

[Kaw85c] Bernhard Kawohl. Rearrangements. Lecture Notes in Mathematics, 1150:7–99, 1985. CODEN LNMAA2. ISBN 3-540-15693-3 (print), 3-540-39639-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075062/.

Kawohl:1985:RCL

[Kaw85d] Bernhard Kawohl. Rearrangements and Convexity of Level Sets in PDE, volume 1150 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15693-3 (print), 3-540-39639-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 136 pp. LCCN QA3 .L28 no. 1150. URL http:

//link.springer.com/book/10.1007/BFb0075060;   http://
www.springerlink.com/content/978-3-540-39639-0.

Kawakubo:1986:TIE

[Kaw86a] K. Kawakubo. Topological invariance of equivariant rational Pontrjagin classes. Lecture Notes in Mathematics, 1217:183–195, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072823/.

Kawamata:1986:MMP

[Kaw86b] Yujiro Kawamata. On the minimal model problem. Lecture Notes in Mathematics, 1194:41–55, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076994/.

Kawazoe:1987:AHS

[Kaw87] Takeshi Kawazoe. Atomic Hardy spaces on semisimple Lie groups. Lecture Notes in Mathematics, 1243:189–197, 1987. CODEN LN-MAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073023/.

Kawohl:1988:OPC

[Kaw88] Bernhard Kawohl. Open problems connected with level sets of harmonic functions. Lecture Notes in Mathematics, 1344:207– 210, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103356/.

Kawakubo:1989:CTD

[Kaw89] Katsuo Kawakubo. G-s-cobordism theorems do not hold in general for many compact Lie groups G. Lecture Notes in Mathematics, 1375:183–190, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085608/.

Kazdan:1987:PDE

[Kaz87] Jerry L. Kazdan. Partial differential equations in differential geometry. Lecture Notes in Mathematics, 1263:134–170, 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078612/.

Kazamaki:1989:RCM

[Kaz89] N. Kazamaki. A remark on the class of martingales with bounded quadratic variation. Lecture Notes in Mathematics, 1372:47– 51, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083958/.

Kania-Bartoszynska:1986:IH

[KB86] Joanna Kania-Bartoszynska. Involutions on 2-handlebodies. Lecture Notes in Mathematics, 1217:151–166, 1986. CODEN LN-MAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072821/.

Kechris:1985:SES

[Kec85] Alexander S. Kechris. Sets of everywhere singular functions. Lecture Notes in Mathematics, 1141:233–244, 1985. CODEN LN-MAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076223/.

Kechris:1988:AUE

[Kec88a] Alexander S. Kechris. “AD + uniformization” is equivalent to “half ad_{R}”. Lecture Notes in Mathematics, 1333:98–102, 1988. CODEN LNMAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084971/.

Kechris:1988:CTM

[Kec88b] Alexander S. Kechris. A coding theorem for measures. Lecture Notes in Mathematics, 1333:103–109, 1988. CODEN LN-MAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084972/.

Kechris:1988:SCR

[Kec88c] Alexander S. Kechris. Subsets of ℵ_{1} constructible from a real. Lecture Notes in Mathematics, 1333:110–116, 1988. CODEN LN-MAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084973/.

Keesling:1987:USS

[Kee87] James Keesling. Using solenoids in the study of the Stone– Čech compactification. Lecture Notes in Mathematics, 1283:131– 147, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081424/.

Keisler:1987:CES

[Kei87] H. Jerome Keisler. Choosing elements in a saturated model. Lecture Notes in Mathematics, 1292:165–181, 1987. CODEN LN-MAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082237/.

Kempf:1987:CI

[Kem87] George R. Kempf. Computing invariants. Lecture Notes in Mathematics, 1278:81–94, 1987. CODEN LNMAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078808/.

Kenig:1989:RTC

[Ken89] Carlos E. Kenig. Restriction theorems, Carleman estimates, uniform Sobolev inequalities and unique continuation. Lecture Notes in Mathematics, 1384:69–90, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086794/.

Kerner:1985:DGA

[Ker85] R. Kerner. Differential geometrical approach to the theory of amorphous solids. Lecture Notes in Mathematics, 1139:311– 327, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074592/.

Kerber:1986:EUF

[Ker86] Adalbert Kerber. Enumeration under finite group action: Symmetry classes of mappings. Lecture Notes in Mathematics, 1234: 160–176, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072515/.

Kerner:1987:CKK

[Ker87] Richard Kerner. Cosmology and Kaluza–Klein theories. Lecture Notes in Mathematics, 1251:150–174, 1987. CODEN LN-MAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077320/.

Kesten:1986:AFP

[Kes86] Harry Kesten. Aspects of first passage percolation. Lecture Notes in Mathematics, 1180:125–264, 1986. CODEN LNMAA2. ISBN 3-540-16441-3 (print), 3-540-39781-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074919/.

Keselman:1988:EBC

[Kes88a] D. G. Keselman. On the extremal boundary of convex compact measures which represent a non-regular point in Choquet simplex. Lecture Notes in Mathematics, 1344:211–213, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103357/.

Keselman:1988:PCH

[Kes88b] D. G. Keselman. The problem of construction of the harmonic space based on Choquet simplex. Lecture Notes in Mathematics, 1344:214–215, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103358.

Keselman:1988:PQI

[Kes88c] D. G. Keselman. The problem on quasi-interior in Choquet simplexes. Lecture Notes in Mathematics, 1344:216–219, 1988. CO-DEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0103359/.

Keyfitz:1987:SNH

[Key87] Barbara Lee Keyfitz. A survey of nonstrictly hyperbolic conservation laws. Lecture Notes in Mathematics, 1270:152–162, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078325/.

Kubicek:1986:NDS

[KH86] M. Kubíček and M. Holodniok. Nonlinear dynamics systemsbifurcations, continuation methods, periodic solutions. Lecture Notes in Mathematics, 1192:379–386, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076097/.

Khavinson:1985:NTO

[Kha85] D. Khavinson. A note on Toeplitz operators. Lecture Notes in Mathematics, 1166:89–94, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074697/.

Kharlamov:1988:NSD

[Kha88] V. M. Kharlamov. Nonamphicheiral surfaces of degree 4 in RP^{3}. Lecture Notes in Mathematics, 1346:349–356, 1988. CODEN LN-MAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082783/.

Khazal:1989:NMS

[Kha89] Reyadh R. Khazal. A note on maximal subgroups in finite groups. Lecture Notes in Mathematics, 1398:106–109, 1989. CODEN LN-MAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086246/.

Khimshiashvili:1986:TIL

[Khi86] G. N. Khimshiashvili. On the topology of invertible linear singular integral operators. Lecture Notes in Mathematics, 1214: 211–230, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075967/.

Kono:1987:SOC

[KI87] Akira Kono and Kiminao Ishitoya. Squaring operations in mod2 cohomology of quotients of compact Lie groups by maximal tori. Lecture Notes in Mathematics, 1298:192–206, 1987. CODEN LN-MAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083010/.

Kifer:1986:CER

[Kif86] Yuri Kifer. Characteristic exponents for random homeomorphisms of metric spaces. Lecture Notes in Mathematics, 1186:74–84, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076834/.

Kifer:1988:NIN

[Kif88a] Yuri Kifer. A note on integrability of C^{r}-norms of stochastic flows and applications. Lecture Notes in Mathematics, 1325:125– 131, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077921/.

Kifer:1988:NSM

[Kif88b] Yuri Kifer. A note on stochastic models with expanding transformations. Lecture Notes in Mathematics, 1325:132–140, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077922/.

Kikuchi:1989:BER

[Kik89] M. Kikuchi. The best estimation of a ratio inequality for continuous martingales. Lecture Notes in Mathematics, 1372:52–56, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083959/.

Kimura:1986:DMP

[Kim86] Motoo Kimura. Diffusion model of population genetics incorporating group selection, with special reference to an altruistic trait. Lecture Notes in Mathematics, 1203:101–118, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076876/.

Kirsch:1985:RSO

[Kir85a] Werner Kirsch. Random Schrödinger operators and the density of states. Lecture Notes in Mathematics, 1109:68–102, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101538/.

Kirschenhofer:1985:AUD

[Kir85b] P. Kirschenhofer. Asymptotische Untersuchungen Zur Durchschnittlichen Gestalt Gewisser Graphenklassen. (German) []. Lecture Notes in Mathematics, 1114:40–54, 1985. CODEN LN-MAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101644/.

Kirsch:1986:LES

[Kir86] W. Kirsch. Lyapunov exponents for Schrödinger operators with random, but deterministic potentials. Lecture Notes in Mathematics, 1186:227–241, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076842/.

Kirsch:1987:EDB

[Kir87a] Werner Kirsch. Estimates on the difference between succeeding eigenvalues and Lifshitz tails for random Schrödinger operators. Lecture Notes in Mathematics, 1250:138–151, 1987. CODEN LN-MAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077354/.

Kirschenhofer:1987:AUD

[Kir87b] Peter Kirschenhofer. Asymptotische Untersuchungen zur Durchschnittlichen Gestalt Gewisser Graphenklassen II. (German) [].

Lecture Notes in Mathematics, 1262:86–92, 1987. CODEN LN-MAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078597/.

Kirk:1988:LMF

[Kir88] Paul A. Kirk. Link maps in the four sphere. Lecture Notes in Mathematics, 1350:31–43, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081467/.

Kirby:1989:MMZ

[Kir89a] Robion C. Kirby. p_{1}(M) = 3σ(M), ω_{4}^{so} = Z and ω_{4}^{spin} = Z. Lecture Notes in Mathematics, 1374:57–58, 1989. CODEN LN-MAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089041.

Kirby:1989:TLC

[Kir89b] Robion C. Kirby. T_{Lie}^{3} and CP^{2} ♯^{9}(−CP^{2}). Lecture Notes in Mathematics, 1374:38–42, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089037/.

Kirby:1989:MD

[Kir89c] Robion C. Kirby. 3-manifolds; a digression. Lecture Notes in Mathematics, 1374:46–48, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089039/.

Kirby:1989:BMb

[Kir89d] Robion C. Kirby. Back matter. Lecture Notes in Mathematics, 1374:??, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-46171-5/1.pdf.

Kirby:1989:BMa

[Kir89e] Robion C. Kirby. Bounding 5-manifolds. Lecture Notes in Mathematics, 1374:49–56, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089040/.

Kirby:1989:CH

[Kir89f] Robion C. Kirby. Casson handles. Lecture Notes in Mathematics, 1374:72–85, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089044/.

Kirby:1989:CT

[Kir89g] Robion C. Kirby. Classification theorems. Lecture Notes in Mathematics, 1374:31–32, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089035.

Kirby:1989:E

[Kir89h] Robion C. Kirby. Exotic R^{4}’s. Lecture Notes in Mathematics, 1374:95–101, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089046/.

Kirby:1989:FW

[Kir89i] Robion C. Kirby. Freedman’s work. Lecture Notes in Mathematics, 1374:86–94, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089045/.

Kirby:1989:FM

[Kir89j] Robion C. Kirby. Front matter. Lecture Notes in Mathematics, 1374:??, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46171-5/1.pdf.

Kirby:1989:HFL

[Kir89k] Robion C. Kirby. Handlebodies and framed links. Lecture Notes in Mathematics, 1374:3–19, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089033/.

Kirby:1989:IM

[Kir89l] Robion C. Kirby. Immersing 4-manifolds in R^{6}. Lecture Notes in Mathematics, 1374:43–45, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089038/.

Kirby:1989:IF

[Kir89m] Robion C. Kirby. Intersection forms. Lecture Notes in Mathematics, 1374:20–30, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089034/.

Kirby:1989:I

[Kir89n] Robion C. Kirby. Introduction. Lecture Notes in Mathematics, 1374:1–2, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089032.

Kirby:1989:RT

[Kir89o] Robion C. Kirby. Rohlin’s theorem. Lecture Notes in Mathematics, 1374:64–71, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089043/.

Kirby:1989:SS

[Kir89p] Robion C. Kirby. Spin structures. Lecture Notes in Mathematics, 1374:33–37, 1989. CODEN LNMAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089036/.

Kirby:1989:WDC

[Kir89q] Robion C. Kirby. Wall’s diffeomorphisms and H-cobordism. Lecture Notes in Mathematics, 1374:59–63, 1989. CODEN LN-MAA2. ISBN 3-540-51148-2 (print), 3-540-46171-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089042/.

Kiss:1985:NVG

[Kis85] Emil W. Kiss. A note on varieties of graph algebras. Lecture Notes in Mathematics, 1149:163–166, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098463/.

Kisynski:1986:JPM

[Kis86] Jan Kisyński. On jumps of paths of Markov processes. Lecture Notes in Mathematics, 1210:130–145, 1986. CODEN LN-MAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077179/.

Kisynski:1989:LFI

[Kis89] Jan Kisyński. Localizations of Feller infinitesimal generators and uniqueness of corresponding killed processes. Lecture Notes in Mathematics, 1379:141–178, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087852/.

Kiwiel:1985:ASM

[Kiw85a] Krzysztof C. Kiwiel. Aggregate subgradient methods for unconstrained convex minimization. Lecture Notes in Mathematics, 1133:44–86, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074502/.

Kiwiel:1985:BM

[Kiw85b] Krzysztof C. Kiwiel. Bundle methods. Lecture Notes in Mathematics, 1133:299–344, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074507/.

Kiwiel:1985:FPM

[Kiw85c] Krzysztof C. Kiwiel. Feasible point methods for convex constrained minimization problems. Lecture Notes in Mathematics, 1133:190–228, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074505/.

Kiwiel:1985:F

[Kiw85d] Krzysztof C. Kiwiel. Fundamentals. Lecture Notes in Mathematics, 1133:1–43, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074501/.

Kiwiel:1985:MDN

[Kiw85e] Krzysztof C. Kiwiel. Methods of Descent for Nondifferentiable Optimization, volume 1133 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 362 pp. LCCN QA3 .L28

no. 1133. URL http://link.springer.com/book/10.1007/
BFb0074500;   http://www.springerlink.com/content/978-
3-540-39509-6.

Kiwiel:1985:MFD

[Kiw85f] Krzysztof C. Kiwiel. Methods of feasible directions for nonconvex constrained problems. Lecture Notes in Mathematics, 1133: 229–298, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074506/.

Kiwiel:1985:MSD

[Kiw85g] Krzysztof C. Kiwiel. Methods with subgradient deletion rules for unconstrained nonconvex minimization. Lecture Notes in Mathematics, 1133:139–189, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074504/.

Kiwiel:1985:MSL

[Kiw85h] Krzysztof C. Kiwiel. Methods with subgradient locality measures for minimizing nonconvex functions. Lecture Notes in Mathematics, 1133:87–138, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074503/.

Kiwiel:1985:NE

[Kiw85i] Krzysztof C. Kiwiel. Numerical examples. Lecture Notes in Mathematics, 1133:345–353, 1985. CODEN LNMAA2. ISBN 3-540-15642-9 (print), 3-540-39509-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074508/.

Kurzweil:1986:PIP

[KJ86] J. Kurzweil and J. Jarník. Perron integral, Perron product integral and ordinary linear differential equations. Lecture Notes in Mathematics, 1192:149–154, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076062/.

Klass:1985:LIL

[KK85] Michael Klass and James Kuelbs. The law of the iterated logarithm in the ℓ^{p} spaces. Lecture Notes in Mathematics, 1153: 310–322, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074957/.

Keyfitz:1989:VAS

[KK89a] Barbara Lee Keyfitz and Herbert C. Kranzer. A viscosity approximation to a system of conservation laws with no classical Riemann solution. Lecture Notes in Mathematics, 1402:185–197, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083875/.

Ku:1989:PNO

[KK89b] Hsu-Tung Ku and Mei-Chin Ku. The Pontrjagin numbers of an orbit map and generalized G-signature theorem. Lecture Notes in Mathematics, 1375:198–206, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085610/.

Kalton:1986:ELP

[KL86] N. J. Kalton and L. J. Lange. Equimodular limit periodic continued fractions. Lecture Notes in Mathematics, 1199:159–219, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075939/.

Knight:1987:SSC

[KL87a] Julia F. Knight and Alistair H. Lachlan. Shrinking, stretching, and codes for homogeneous structures. Lecture Notes in Mathematics, 1292:192–229, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082239/.

Knopfmacher:1987:AFC

[KL87b] Arnold Knopfmacher and D. S. Lubinsky. Analogues of Freud’s conjecture for Erdős type weights and related polynomial approximation problems. Lecture Notes in Mathematics, 1287:21– 69, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078897/.

Koski:1987:IDP

[KL87c] Timo Koski and Wilfried Loges. On identification for distributed parameter systems. Lecture Notes in Mathematics, 1250:152– 159, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077355/.

Kaiser:1988:EIM

[KL88] U. Kaiser and B. H. Li. Enumeration of immersions of m-manifolds in (2m − 2)-manifolds by the singularity method. Lecture Notes in Mathematics, 1350:171–187, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081475/.

Kleppe:1987:NRC

[Kle87] Jan O. Kleppe. Non-reduced components of the Hilbert scheme of smooth space curves. Lecture Notes in Mathematics, 1266:181– 207, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078184/.

Kleppe:1989:LFS

[Kle89] Jan O. Kleppe. Liaison of families of subschemes in P^{n}. Lecture Notes in Mathematics, 1389:128–173, 1989. CODEN LN-MAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085930/.

Kulkarni:1985:DSS

[KLR85] Ravi Kulkarni, Kyung Bai Lee, and Frank Raymond. Deformation spaces for Seifert manifolds. Lecture Notes in Mathematics, 1167:180–216, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075224/.

Kaspi:1986:PLT

[KM86a] Haya Kaspi and Bernard Maisonneuve. Predictable local times and exit systems. Lecture Notes in Mathematics, 1204:95–100, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075715/.

Kolp:1986:BCS

[KM86b] O. Kolp and H. Mierendorff. Bus coupled systems for multigrid algorithms. Lecture Notes in Mathematics, 1228:202–218, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072648/.

Konig:1987:CNC

[KM87] Hermann König and Vitali D. Milman. On the covering numbers of convex bodies. Lecture Notes in Mathematics, 1267:82–95, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078138/.

Kasahara:1988:WCF

[KM88] Yuji Kasahara and Makoto Maejima. Weak convergence of functionals of point processes on R^{d}. Lecture Notes in Mathematics, 1322:73–84, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077867/.

Klebanov:1987:SSL

[KMM87] L. B. Klebanov, J. A. Melamed, and C. M. Manija. υ_{p}-strictly stable laws and estimation of their parameters. Lecture Notes in Mathematics, 1233:23–31, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072707/.

Klebanov:1989:PRN

[KMR89] L. B. Klebanov, J. A. Melamed, and S. T. Rachev. On the products of a random number of random variables in connection with a problem from mathematical economics. Lecture Notes in Mathematics, 1412:103–109, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084168/.

Kaneko:1988:NAS

[KN88a] Hiroshi Kaneko and Shintaro Nakao. A note on approximation for stochastic differential equations. Lecture Notes in Mathematics, 1321:155–162, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084132/.

Kerscher:1988:PSC

[KN88b] Wolfgang Kerscher and Rainer Nagel. Positivity and stability for Cauchy problems with delay. Lecture Notes in Mathematics, 1324:216–235, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100794/.

Kawamoto:1989:ECL

[KN89] Naoki Kawamoto and N. Nomura. Extensions of the class of Lie algebras with the lattice of subideals. Lecture Notes in Mathematics, 1398:100–105, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086245/.

Knebusch:1989:BTW

[Kne89a] Manfred Knebusch. Basic theory of weakly semialgebraic spaces. Lecture Notes in Mathematics, 1367:1–105, 1989. CODEN LN-MAA2. ISBN 3-540-50815-5 (print), 3-540-46089-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084988/.

Knebusch:1989:HC

[Kne89b] Manfred Knebusch. Homology and cohomology. Lecture Notes in Mathematics, 1367:182–259, 1989. CODEN LNMAA2. ISBN 3-540-50815-5 (print), 3-540-46089-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084990/.

Knebusch:1989:PCH

[Kne89c] Manfred Knebusch. Patch complexes, and homotopies again. Lecture Notes in Mathematics, 1367:106–181, 1989. CODEN LN-MAA2. ISBN 3-540-50815-5 (print), 3-540-46089-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084989/.

Knebusch:1989:SS

[Kne89d] Manfred Knebusch. Simplicial spaces. Lecture Notes in Mathematics, 1367:260–351, 1989. CODEN LNMAA2. ISBN 3-540-50815-5 (print), 3-540-46089-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084991/.

Knight:1986:PRS

[Kni86] F. B. Knight. Poisson representation of strict regular step filtrations. Lecture Notes in Mathematics, 1204:1–27, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075706/.

Knight:1987:DMP

[Kni87] Julia F. Knight. Degrees of models with prescribed Scott set. Lecture Notes in Mathematics, 1292:182–191, 1987. CODEN LN-MAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082238/.

Knopp:1985:MID

[Kno85] Marvin I. Knopp. Modular integrals on Γ_{o}(N) and Dirichlet series with functional equations. Lecture Notes in Mathematics, 1135:211–224, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074606/.

Knorrer:1987:TMD

[Knö87] Horst Knörrer. Torsionsfreie Moduln bei Deformation von Kurvensingularitäten. (German) []. Lecture Notes in Mathematics, 1273:150–155, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078841/.

Knopp:1989:RDT

[Kno89] M. I. Knopp. Recent developments in the theory of rational period functions. Lecture Notes in Mathematics, 1383:111–122, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083573/.

Kobayashi:1987:WTP

[Kob87] Osamu Kobayashi. A Willmore type problem for S^{2} × S^{2}. Lecture Notes in Mathematics, 1255:67–72, 1987. CODEN LN-MAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077681/.

Kobayashi:1989:CBF

[Kob89] Satoshi Kobayashi. Consistency of Beeson’s formal system RPS and some related results. Lecture Notes in Mathematics, 1388: 120–140, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083667/.

Kock:1986:ISD

[Koc86] Anders Kock. Introduction to synthetic differential geometry, and a synthetic theory of dislocations. Lecture Notes in Mathematics, 1174:52–68, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076934/.

Kock:1988:GFB

[Koc88] Anders Kock. Generalized fibre bundles. Lecture Notes in Mathematics, 1348:194–207, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081359/.

Kodnar:1986:PEA

[Kod86] R. Kodnár. A posteriori estimations of approximate solutions for some types of boundary value problems. Lecture Notes in Mathematics, 1192:373–378, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076096/.

Koepf:1988:CFN

[Koe88] Wolfram Koepf. Convex functions and the Nehari univalence criterion. Lecture Notes in Mathematics, 1351:214–218, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081255/.

Kogan:1986:A

[Kog86a] Jacob Kogan. Appendix. Lecture Notes in Mathematics, 1216: 97–103, 1986. CODEN LNMAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077068/.

Kogan:1986:BEO

[Kog86b] Jacob Kogan. Bifurcation of Extremals in Optimal Control, volume 1216 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 106 pp. LCCN QA3 .L28 no. 1216. URL http:

//link.springer.com/book/10.1007/BFb0077060;   http://
www.springerlink.com/content/978-3-540-47065-6.

Kogan:1986:BPLb

[Kog86c] Jacob Kogan. Branching pairs in linear control problems. Lecture Notes in Mathematics, 1216:49–54, 1986. CODEN LN-MAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077063/.

Kogan:1986:BPLa

[Kog86d] Jacob Kogan. Branching points in linear control problems. Lecture Notes in Mathematics, 1216:40–48, 1986. CODEN LN-MAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077062/.

Kogan:1986:LSV

[Kog86e] Jacob Kogan. Linear systems with vector valued performance index. Lecture Notes in Mathematics, 1216:68–77, 1986. CODEN LNMAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077065/.

Kogan:1986:NC

[Kog86f] Jacob Kogan. The nonlinear case. Lecture Notes in Mathematics, 1216:55–67, 1986. CODEN LNMAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077064/.

Kogan:1986:NCS

[Kog86g] Jacob Kogan. Nonlinear control systems with vector cost. Lecture Notes in Mathematics, 1216:78–84, 1986. CODEN LN-MAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077066/.

Kogan:1986:OCP

[Kog86h] Jacob Kogan. Optimal control problems with constraints. Lecture Notes in Mathematics, 1216:85–96, 1986. CODEN LN-MAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077067/.

Kogan:1986:O

[Kog86i] Jacob Kogan. Overview. Lecture Notes in Mathematics, 1216: 1–39, 1986. CODEN LNMAA2. ISBN 3-540-16818-4 (print), 3-540-47065-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077061/.

Koicheva:1987:MMD

[Koi87] M. Koicheva. The method of metric distances in the problem of estimation of the deviation from the exponential distribution. Lecture Notes in Mathematics, 1233:32–35, 1987. CODEN LN-MAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072708/.

Kolchin:1986:DAG

[Kol86] E. R. Kolchin. Differential algebraic groups. Lecture Notes in Mathematics, 1185:155–174, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076174/.

Komiya:1989:CFP

[Kom89a] Katsuhiro Komiya. Congruences for fixed point indices of equivariant maps and iterated maps. Lecture Notes in Mathematics, 1411:130–136, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086447/.

Komiya:1989:CBR

[Kom89b] Katsuhiro Komiya. Congruences for the Burnside ring. Lecture Notes in Mathematics, 1375:191–197, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085609/.

Kon:1985:ICS

[Kon85a] Mark A. Kon. Integrability criteria and some applications. Lecture Notes in Mathematics, 1148:83–103, 1985. CODEN LN-MAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077162/.

Kon:1985:IIR

[Kon85b] Mark A. Kon. Integrals of independent random variables. Lecture Notes in Mathematics, 1148:16–27, 1985. CODEN LN-MAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077157/.

Kon:1985:I

[Kon85c] Mark A. Kon. Introduction. Lecture Notes in Mathematics, 1148: 1–9, 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077155/.

Kon:1985:JDA

[Kon85d] Mark A. Kon. Joint distributions and applications. Lecture Notes in Mathematics, 1148:104–112, 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077163/.

Kon:1985:LI

[Kon85e] Mark A. Kon. The Lebesgue integral. Lecture Notes in Mathematics, 1148:66–82, 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077161/.

Kon:1985:PA

[Kon85f] Mark A. Kon. Physical applications. Lecture Notes in Mathematics, 1148:51–65, 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077160/.

Kon:1985:PDQ

[Kon85g] Mark A. Kon. Probability Distributions in Quantum Statistical Mechanics, volume 1148 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 120 pp. LCCN QA3 .L28

no. 1148. URL http://link.springer.com/book/10.1007/
BFb0077154;   http://www.springerlink.com/content/978-
3-540-39637-6.

Kon:1985:SCLa

[Kon85h] Mark A. Kon. Singular central limit theorems, I. Lecture Notes in Mathematics, 1148:28–38, 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077158/.

Kon:1985:SCLb

[Kon85i] Mark A. Kon. Singular central limit theorems, II. Lecture Notes in Mathematics, 1148:39–50, 1985. CODEN LNMAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077159/.

Kon:1985:VFC

[Kon85j] Mark A. Kon. Value functions on a canonical ensemble. Lecture Notes in Mathematics, 1148:10–15, 1985. CODEN LN-MAA2. ISBN 3-540-15690-9 (print), 3-540-39637-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077156/.

Kono:1985:DPD

[Kôn85k] Norio Kôno. Démonstration probabiliste du théorème de d’Alembert. (French) []. Lecture Notes in Mathematics, 1123: 207–208, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075848.

Koornwinder:1985:SOP

[Koo85] T. H. Koornwinder. Special orthogonal polynomial systems mapped onto each other by the Fourier–Jacobi transform. Lecture Notes in Mathematics, 1171:174–183, 1985. CODEN LN-MAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076542/.

Koornwinder:1988:GTI

[Koo88] Tom H. Koornwinder. Group theoretic interpretations of Askey’s scheme of hypergeometric orthogonal polynomials. Lecture Notes in Mathematics, 1329:46–72, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083353/.

Kopetzky:1989:SRD

[Kop89] Hans G. Kopetzky. Some results on Diophantine approximation related to Dirichlet’s Theorem. Lecture Notes in Mathematics, 1380:137–149, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086550/.

Korevaar:1987:SCA

[Kor87a] J. Korevaar. Some C^{N} capacities and applications. Lecture Notes in Mathematics, 1276:208–228, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078960/.

Korezlioglu:1987:PTP

[Kör87b] H. Körezlioğlu. Passage from two-parameters to infinite dimension. Lecture Notes in Mathematics, 1236:131–153, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072886/.

Korolev:1987:ANA

[Kor87c] V. Yu. Korolev. The accuracy of the normal approximation to the distribution of the sum of a random number of independent random variables. Lecture Notes in Mathematics, 1233:36–40, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072709/.

Korchagin:1988:ICP

[Kor88] A. B. Korchagin. Isotopy classification of plane seventh degree curves with the only singular point Z_{15}. Lecture Notes in Mathematics, 1346:407–426, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082785/.

Korolev:1989:ADR

[Kor89] V. Yu. Korolev. The asymptotic distributions of random sums. Lecture Notes in Mathematics, 1412:110–123, 1989. CODEN LN-MAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084169/.

Koschorke:1985:HOH

[Kos85] Ulrich Koschorke. Higher order homotopy invariants for higher dimensional link maps. Lecture Notes in Mathematics, 1172:116– 129, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074427/.

Kosarew:1986:MHA

[Kos86a] Siegmund Kosarew. Modulräume holomorpher Abbildungen auf komplexen Mannigfaltigkeiten mit 1-konkavem Rand. (German) []. Lecture Notes in Mathematics, 1194:56–79, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076995/.

Koshitani:1986:GAF

[Kos86b] Shigeo Koshitani. On group algebras of finite groups. Lecture Notes in Mathematics, 1178:109–128, 1986. CODEN LN-MAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075292/.

Koschorke:1988:MPI

[Kos88a] Ulrich Koschorke. Multiple point invariants of link maps. Lecture Notes in Mathematics, 1350:44–86, 1988. CODEN LN-MAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081468/.

Koschorke:1988:SMI

[Kos88b] Ulrich Koschorke. The singularity method and immersions of mmanifolds into manifolds of dimensions 2m − 2, 2m − 3 and 2m − 4. Lecture Notes in Mathematics, 1350:188–212, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081476/.

Koslowski:1988:COP

[Kos88c] Jürgen Koslowski. Closure operators with prescribed properties. Lecture Notes in Mathematics, 1348:208–220, 1988. CODEN LN-MAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081360/.

Kotelenez:1989:SRD

[Kot89] Peter Kotelenez. A stochastic reaction-diffusion model. Lecture Notes in Mathematics, 1390:132–137, 1989. CODEN LN-MAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083941/.

Kovaceva:1987:BCR

[Kov87] R. K. Kovačeva. Best Chebyshev rational approximants and poles of functions. Lecture Notes in Mathematics, 1237:68–72, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072456/.

Kowalski:1985:ACO

[Kow85] Marek A. Kowalski. Algebraic characterization of orthogonality in the space of polynomials. Lecture Notes in Mathematics, 1171:101–110, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076535/.

Kozlowski:1985:ECS

[Koz85] Michael Kozlowski. Examples concerning the spectrum of a closed Riemannian manifold. Lecture Notes in Mathematics, 1156:219– 226, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075094/.

Kozniewski:1986:EAC

[Koź86] Tadeusz Koźniewski. On the existence of acyclic Γ complexes of the lowest possible dimension. Lecture Notes in Mathematics, 1217:196–209, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072824/.

Kuhnel:1985:TSS

[KP85] W. Kühnel and U. Pinkall. Tight smoothing of some polyhedral surfaces. Lecture Notes in Mathematics, 1156:227–239, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075095/.

Kaijser:1986:CN

[KP86a] Sten Kaijser and Joan Wick Pelletier. Categorical notions. Lecture Notes in Mathematics, 1208:44–72, 1986. CODEN LN-MAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077086/.

Kaijser:1986:CMC

[KP86b] Sten Kaijser and Joan Wick Pelletier. The classical methods from a categorical viewpoint. Lecture Notes in Mathematics, 1208:132– 159, 1986. CODEN LNMAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077091/.

Kaijser:1986:CM

[KP86c] Sten Kaijser and Joan Wick Pelletier. The complex method. Lecture Notes in Mathematics, 1208:33–43, 1986. CODEN LN-MAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077085/.

Kaijser:1986:D

[KP86d] Sten Kaijser and Joan Wick Pelletier. Duality. Lecture Notes in Mathematics, 1208:106–122, 1986. CODEN LNMAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077089/.

Kaijser:1986:FDD

[KP86e] Sten Kaijser and Joan Wick Pelletier. Finite dimensional Doolittle diagrams. Lecture Notes in Mathematics, 1208:73–92, 1986. CODEN LNMAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077087/.

Kaijser:1986:IFD

[KP86f] Sten Kaijser and Joan Wick Pelletier. Interpolation Functors and Duality, volume 1208 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 167 pp. LCCN QA3 .L28 no. 1208. URL http:

//link.springer.com/book/10.1007/BFb0077081;   http://
www.springerlink.com/content/978-3-540-47044-1.

Kaijser:1986:I

[KP86g] Sten Kaijser and Joan Wick Pelletier. Introduction. Lecture Notes in Mathematics, 1208:1–6, 1986. CODEN LNMAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077082/.

Kaijser:1986:KE

[KP86h] Sten Kaijser and Joan Wick Pelletier. Kan extensions. Lecture Notes in Mathematics, 1208:93–105, 1986. CODEN LN-MAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077088/.

Kaijser:1986:MAD

[KP86i] Sten Kaijser and Joan Wick Pelletier. More about duality. Lecture Notes in Mathematics, 1208:123–131, 1986. CODEN LN-MAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077090/.

Kaijser:1986:P

[KP86j] Sten Kaijser and Joan Wick Pelletier. Preliminaries. Lecture Notes in Mathematics, 1208:7–17, 1986. CODEN LNMAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077083/.

Kaijser:1986:RM

[KP86k] Sten Kaijser and Joan Wick Pelletier. The real method. Lecture Notes in Mathematics, 1208:18–32, 1986. CODEN LN-MAA2. ISBN 3-540-16790-0 (print), 3-540-47044-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077084/.

Kulkarni:1986:UGS

[KP86l] Ravi S. Kulkarni and Ulrich Pinkall. Uniformization of geometric structures with applications to conformal geometry. Lecture Notes in Mathematics, 1209:190–209, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076632/.

Kac:1987:GIT

[KP87a] Victor G. Kac and Dale H. Peterson. On geometric invariant theory for infinite-dimensional groups. Lecture Notes in Mathematics, 1271:109–142, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079235/.

Kirschenhofer:1987:AUO

[KP87b] P. Kirschenhofer and H. Prodinger. Asymptotische Untersuchungen ober Charakteristische Parameter von Suchbăumen (German) []. Lecture Notes in Mathematics, 1262:93–107, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078598/.

Kallianpur:1989:WCS

[KPA89] G. Kallianpur and V. Pérez-Abreu. Weak convergence of solutions of stochastic evolution equations on nuclear spaces. Lecture Notes in Mathematics, 1390:119–131, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083940/.

Kirschenhofer:1987:AGN

[KPS87] Peter Kirschenhofer, Helmut Prodinger, and Johannes Schoißengeier. Zur Auswertung Gewisser Numerischer Reihen mit Hilfe Modularer Funktionen. (German) []. Lecture Notes in Mathematics, 1262:108–110, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078599/.

Kirschenhofer:1985:OZN

[KPT85] P. Kirschenhofer, H. Prodinger, and R. F. Tichy. Ober Die Ziffernsumme Natürlicher Zahlen Und Verwandte Probleme. (German) [On the digit sum of natural numbers and related problems]. Lecture Notes in Mathematics, 1114:55–65, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101645/.

Kirschenhofer:1987:EFA

[KPT87] Peter Kirschenhofer, Helmut Prodinger, and Robert F. Tichy. Über Einige Funktionaldifferentialgleichungen Aus der Analyse Von Algorithmen. (German) []. Lecture Notes in Mathematics, 1262:111–123, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078600/.

Kalashnikov:1987:SPS

[KPZ87] Vladimir V. Kalashnikov, Boyan Penkov, and Vladimir M. Zolotarev, editors. Stability Problems for Stochastic Models: Proceedings of the 9th International Seminar held in Varna, Bulgaria, May 13–19, 1985, volume 1233 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1233. URL http:

//link.springer.com/book/10.1007/BFb0072704;   http://
www.springerlink.com/content/978-3-540-47394-7.

Khoai:1988:ANT

[KQ88] Hà Huy Khoái and My Vinh Quang. On p-adic Nevanlinna theory. Lecture Notes in Mathematics, 1351:146–158, 1988. CODEN LNMAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081250/.

Klebanov:1985:SLM

[KR85] L. B. Klebanov and S. T. Rachev. Stability of the lack-of-memory property in a finite number of points. Lecture Notes in Mathematics, 1155:131–143, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074817/.

Krantz:1987:RPF

[Kra87a] Steven G. Krantz. Recent progress and future directions in several complex variables. Lecture Notes in Mathematics, 1268:1– 23, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097292/.

Krause:1987:CIA

[Kra87b] Günter Krause. Chaînes d’idéaux annulateurs d’un anneau noethérien. (French) []. Lecture Notes in Mathematics, 1296:261– 275, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078532/.

Kral:1988:BRP

[Krá88a] Josef Král. Boundary regularity and potential-theoretic operators. Lecture Notes in Mathematics, 1344:220–222, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103360/.

Krall:1988:ROP

[Kra88b] Allan M. Krall. A review of orthogonal polynomials satisfying boundary value problems. Lecture Notes in Mathematics, 1329: 73–97, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083354/.

Krasilshchik:1988:SBC

[Kra88c] I. S. Krasil’shchik. Schouten bracket and canonical algebras. Lecture Notes in Mathematics, 1334:79–110, 1988. CODEN LN-MAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080424/.

Kral:1989:PNF

[Krá89] Josef Král. Problème de Neumann faible avec condition frontière dans L^{1}. (French) []. Lecture Notes in Mathematics, 1393:145– 160, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085777/.

Krbec:1986:NS

[Krb86] P. Krbec. On nonparasite solutions. Lecture Notes in Mathematics, 1192:133–139, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076060/.

Krenk:1985:SDE

[Kre85] Steen Krenk. Scalar decomposition of elastic waves scattered by planar cracks. Lecture Notes in Mathematics, 1121:166–180, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076270/.

Kreith:1986:UZB

[Kre86a] K. Kreith. Uniform zeros for beaded strings. Lecture Notes in Mathematics, 1192:141–148, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076061/.

Krejci:1986:PSP

[Kre86b] P. Krejčí. Periodic solutions of partial differential equations with hysteresis. Lecture Notes in Mathematics, 1192:221–226, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076073/.

Kreweras:1986:JDT

[Kre86c] Germain Kreweras. Joint distributions of three descriptive parameters of bridges. Lecture Notes in Mathematics, 1234:177– 191, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072516/.

Kree:1988:TDD

[Kré88] Paul Krée. La théorie des distributions en dimension quelconque et l’intégration stochastique. (French) []. Lecture Notes in Mathematics, 1316:170–233, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081932/.

Krieg:1985:EPS

[Kri85a] Aloys Krieg. Eisenstein- and Poincaré-series. Lecture Notes in Mathematics, 1143:136–181, 1985. CODEN LNMAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075951/.

Krieg:1985:IHM

[Kri85b] Aloys Krieg. Integral and Hermitian matrices. Lecture Notes in Mathematics, 1143:1–41, 1985. CODEN LNMAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075947/.

Krieg:1985:MFa

[Kri85c] Aloys Krieg. Modular forms. Lecture Notes in Mathematics, 1143: 71–98, 1985. CODEN LNMAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075949/.

Krieg:1985:MFH

[Kri85d] Aloys Krieg. Modular Forms on Half-Spaces of Quaternions, volume 1143 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 203 pp. LCCN QA3 .L28 no. 1143. URL http:

//link.springer.com/book/10.1007/BFb0075946;   http://
www.springerlink.com/content/978-3-540-39610-9.

Krieg:1985:MFb

[Kri85e] Aloys Krieg. Modular functions. Lecture Notes in Mathematics, 1143:182–195, 1985. CODEN LNMAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075952/.

Krieg:1985:MGF

[Kri85f] Aloys Krieg. Modular group and fundamental domain. Lecture Notes in Mathematics, 1143:42–70, 1985. CODEN LN-MAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075948/.

Krieg:1985:TS

[Kri85g] Aloys Krieg. Theta-series. Lecture Notes in Mathematics, 1143: 99–135, 1985. CODEN LNMAA2. ISBN 3-540-15679-8 (print), 3-540-39610-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075950/.

Krizek:1986:SRL

[Kří86] M. Křížek. Superconvergence results for linear triangular elements. Lecture Notes in Mathematics, 1192:315–320, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076087/.

Kamber:1986:ATS

[KRT86] F. W. Kamber, E. A. Ruh, and Ph. Tondeur. Almost transversally symmetric foliations. Lecture Notes in Mathematics, 1209: 184–189, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076631/.

Kruglov:1985:CLT

[Kru85] V. M. Kruglov. The central limit theorem and weak convergence of maxima of sums of independent random variables. Lecture Notes in Mathematics, 1155:144–179, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074818/.

Kruglov:1989:NDC

[Kru89] V. M. Kruglov. Normal and degenerate convergences of random sums. Lecture Notes in Mathematics, 1412:124–136, 1989. CO-DEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084170/.

Kalton:1985:BSP

[KS85a] Nigel J. Kalton and Elias Saab, editors. Banach Spaces: Proceedings of the Missouri Conference held in Columbia, USA, June 24–29, 1984, volume 1166 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1166. URL http:

//link.springer.com/book/10.1007/BFb0074684;   http://
www.springerlink.com/content/978-3-540-39736-6.

Kosmann-Schwarzbach:1985:MMF

[KS85b] Yvette Kosmann-Schwarzbach. On the momentum mapping in field theory. Lecture Notes in Mathematics, 1139:25–73, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074576/.

Kuznezova-Sholpo:1985:CMF

[KS85c] I. A. Kuznezova-Sholpo. Compound metrics with fixed minimal ones. Lecture Notes in Mathematics, 1155:180–189, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074819/.

Kasue:1986:GTC

[KS86a] Atsushi Kasue and Kunio Sugahara. Gap theorems for certain submanifolds of Euclidean space and hyperbolic space form II. Lecture Notes in Mathematics, 1201:138–149, 1986. CODEN LN-MAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075651/.

Katok:1986:AC

[KS86b] A. Katok and J.-M. Strelcyn. Absolute continuity. Lecture Notes in Mathematics, 1222:41–156, 1986. CODEN LNMAA2. ISBN 3-540-17190-8 (print), 3-540-47349-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099033/.

Katok:1986:EEA

[KS86c] A. Katok and J.-M. Strelcyn. The estimation of entropy from above through Lyapunov characteristic numbers. Lecture Notes in Mathematics, 1222:180–198, 1986. CODEN LNMAA2. ISBN 3-540-17190-8 (print), 3-540-47349-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099035/.

Katok:1986:EIM

[KS86d] A. Katok and J.-M. Strelcyn. Existence of invariant manifolds for smooth maps with singularities. Lecture Notes in Mathematics, 1222:1–40, 1986. CODEN LNMAA2. ISBN 3-540-17190-8 (print), 3-540-47349-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099032/.

Klingenberg:1986:RIR

[KS86e] W. Klingenberg and T. Sakai. Remarks on the injectivity radius estimate for almost 1/4-pinched manifolds. Lecture Notes in Mathematics, 1201:156–164, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075653/.

Koiso:1986:NHK

[KS86f] Norihito Koiso and Yusuke Sakane. Non-homogeneous Kähler– Einstein metrics on compact complex manifolds. Lecture Notes in Mathematics, 1201:165–179, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075654/.

Kuchminskaya:1987:RAI

[KS87] Khristina I. Kuchminskaya and Wojciech Siemaszko. Rational approximation and interpolation of functions by branched continued fractions. Lecture Notes in Mathematics, 1237:24–40, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072451/.

Katsura:1988:DLD

[KS88a] Masashi Katsura and H. J. Shyr. Decomposition of languages into disjunctive outfix codes. Lecture Notes in Mathematics, 1320: 172–175, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083430/.

Kleiman:1988:EGN

[KS88b] Steven L. Kleiman and Robert Speiser. Enumerative geometry of nodal plane cubics. Lecture Notes in Mathematics, 1311:156– 196, 1988. CODEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082914/.

Krasilnikov:1988:LFD

[KŠ88c] A. N. Krasil’nikov and A. L. Šmel’kin. On the laws of finite dimensional representations of solvable Lie algebras and groups. Lecture Notes in Mathematics, 1352:114–129, 1988. CODEN LN-MAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082021/.

Kumar:1989:DTR

[KS89] Ajay Kumar and Ajit Iqbal Singh. A dichotomy theorem for random walks on hypergroups. Lecture Notes in Mathematics, 1379:179–184, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087853/.

Katok:1986:IME

[KSLP86] Anatole Katok, Jean-Marie Strelcyn, François Ledrappier, and Feliks Przytycki. Invariant Manifolds, Entropy and Billiards; Smooth Maps with Singularities, volume 1222 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17190-8 (print), 3-540-47349-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 278 (est.) pp. LCCN QA3

.L28 no. 1222. URL http://link.springer.com/book/10.1007/
BFb0099031;   http://www.springerlink.com/content/978-
3-540-47349-7.

Kleiman:1987:SVS

[KSX87] S. Kleiman, S. A. Strømme, and S. Xambó. Sketch of a verification of Schubert’s number 5819539783680 of twisted cubics. Lecture Notes in Mathematics, 1266:156–180, 1987. CODEN LN-MAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078183/.

Kamber:1985:BPF

[KT85a] F. W. Kamber and Ph. Tondeur. The Bernstein problem for foliations. Lecture Notes in Mathematics, 1156:216–218, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075093/.

Kirschenhofer:1985:GDR

[KT85b] P. Kirschenhofer and R. F. Tichy. Gleichverteilung In Diskreten Räumen. (German) []. Lecture Notes in Mathematics, 1114:66– 76, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101646/.

Kruglov:1987:MPD

[KT87] V. M. Kruglov and A. N. Titov. Mixtures of probability distributions. Lecture Notes in Mathematics, 1233:41–56, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072710/.

Korezlioglu:1988:IFP

[KÜ88] H. Körezlioğlu and A. S. Üstünel. An Itô formula for processes with values in an abstract Wiener space. Lecture Notes in Mathematics, 1316:234–246, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081933/.

Kucera:1985:MCC

[Kuč85] Antonín Kučera. Measure, Π _{1}^{0}-classes and complete extensions of pa. Lecture Notes in Mathematics, 1141:245–259, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076224/.

Kucera:1986:SBP

[Kuč86a] M. Kučera. Stability and bifurcation problems for reactiondiffusion systems with unilateral conditions. Lecture Notes in Mathematics, 1192:227–234, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076074/.

Kuchment:1986:SSS

[Kuc86b] P. A. Kuchment. On the spectral synthesis in the spaces of solutions of invariant differential equations. Lecture Notes in Mathematics, 1214:85–100, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075960/.

Kuchminskaya:1987:AFT

[Kuc87] Kh. I. Kuchminskaya. On approximation of functions by twodimensional continued fractions. Lecture Notes in Mathematics, 1237:207–216, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072465/.

Kufner:1986:BVP

[Kuf86] A. Kufner. Boundary value problems in weighted spaces. Lecture Notes in Mathematics, 1192:35–48, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076050/.

Kuhn:1987:R

[Kuh87] Nicholas J. Kuhn. The rigidity of L(n). Lecture Notes in Mathematics, 1286:286–292, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078746/.

Kuhn:1989:MTI

[Kuh89] Nicholas J. Kuhn. Morava K-theories and infinite loop spaces. Lecture Notes in Mathematics, 1370:243–257, 1989. CODEN LN-MAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085232/.

Kumjian:1985:DAC

[Kum85] Alex Kumjian. Diagonals in algebras of continuous trace. Lecture Notes in Mathematics, 1132:297–311, 1985. CODEN LN-MAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074890/.

Kummerer:1988:STN

[Küm88] Burkhard Kümmerer. Survey on a theory of non-commutative stationary Markov processes. Lecture Notes in Mathematics, 1303: 154–182, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078061/.

Kung:1987:CFB

[Kun87] Joseph P. S. Kung. Canonical forms for binary forms of even degree. Lecture Notes in Mathematics, 1278:52–61, 1987. CODEN LNMAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078806/.

Kuo:1986:LOG

[Kuo86] Hui-Hsiung Kuo. On Laplacian operators of generalized Brownian functionals. Lecture Notes in Mathematics, 1203:119–128, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076877/.

Kuo:1987:HEF

[Kuo87] Hui-Hsiung Kuo. The heat equation and Fourier transforms of generalized Brownian functionals. Lecture Notes in Mathematics, 1236:154–163, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072887/.

Kuo:1988:BMD

[Kuo88] Hui-Hsiung Kuo. Brownian motion, diffusions and infinite dimensional calculus. Lecture Notes in Mathematics, 1316:130– 169, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081931/.

Kuo:1989:SPD

[Kuo89] Hui-Hsiung Kuo. Stochastic partial differential equations of generalized Brownian functionals. Lecture Notes in Mathematics, 1390:138–146, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083942/.

Kurepa:1987:QSF

[Kur87] Svetozar Kurepa. Quadratic and sesquilinear forms. Contributions to characterizations of inner product spaces. Lecture Notes in Mathematics, 1242:43–76, 1987. CODEN LNMAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072442/.

Kusuoka:1986:GMC

[Kus86a] Shigeo Kusuoka. The generalized Malliavin calculus based on Brownian sheet and Bismut’s expansion for large deviation. Lecture Notes in Mathematics, 1158:141–157, 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080214/.

Kusuoka:1986:PEF

[Kus86b] Shigeo Kusuoka. Precise estimates for the fundamental solutions to some degenerate elliptic differential equations. Lecture Notes in Mathematics, 1203:129–133, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076878/.

Kusuoka:1988:DTC

[Kus88] Shigeo Kusuoka. Degree theorem in certain Wiener Riemannian manifolds. Lecture Notes in Mathematics, 1322:93–108, 1988. CO-DEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077869/.

Kuulasmaa:1986:RRS

[Kuu86] Kari Kuulasmaa. On the reproduction rate of the spatial general epidemic. Lecture Notes in Mathematics, 1212:195–199, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076249/.

Kuusalo:1988:GFM

[Kuu88] Tapani Kuusalo. A Green formula with multiplicities. Lecture Notes in Mathematics, 1351:219–222, 1988. CODEN LN-MAA2. ISBN 3-540-50370-6 (print), 3-540-45992-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081256/.

Kalashnikov:1985:MEF

[KV85] V. V. Kalashnikov and S. Yu. Vsekhsvyat·skiĭ. Metric estimates of the first occurrence time in regenerative processes. Lecture Notes in Mathematics, 1155:102–130, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074816/.

Kharlamov:1988:EGR

[KV88] V. M. Kharlamov and O. Ya. Viro. Extensions of the Gudkov– Rohlin congruence. Lecture Notes in Mathematics, 1346:357– 406, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082784/.

Kalashnikov:1989:CRT

[KV89] V. V. Kalashnikov and S. Yu. Vsekhsvyat·skiĭ. On the connection of Renyi’s theorem and renewal theory. Lecture Notes in Mathematics, 1412:83–102, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084167/.

Kagan:1985:RDS

[KZ85a] A. M. Kagan and A. A. Zinger. A refinement of Darmois– Skitovitch and Heyde theorems. Lecture Notes in Mathematics, 1155:81–94, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074814/.

Kalashnikov:1985:SPS

[KZ85b] Vladimir V. Kalashnikov and Vladimir M. Zolotarev, editors. Stability Problems for Stochastic Models: Proceedings of the 8th International Seminar held in Uzhgorod, USSR, Sept. 23–29, 1984, volume 1155 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1155. URL http:

//link.springer.com/book/10.1007/BFb0074809;   http://
www.springerlink.com/content/978-3-540-39686-4.

Kagan:1987:CAT

[KZ87] A. M. Kagan and A. A. Zinger. Contribution to the analytic theory of linear forms of independent random variables. Lecture Notes in Mathematics, 1233:11–22, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072706/.

Labbas:1986:TPP

[Lab86a] Rabah Labbas. A two point problem for a second order abstract differential equation. Lecture Notes in Mathematics, 1223:156– 159, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099190/.

Labelle:1986:SNC

[Lab86b] Gilbert Labelle. Some new computational methods in the theory of species. Lecture Notes in Mathematics, 1234:192–209, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072517/.

Lirola:1988:CGE

[LAB88] A. Lirola, E. R. Aznar, and M. Bullejos. The cohomology groups of an epimorphism. Lecture Notes in Mathematics, 1348:230– 241, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081362/.

Lacroix:1986:CSP

[Lac86] Jean Lacroix. Computations of the sum of positive Lyapunov exponents for the Lloyd model in a strip. Lecture Notes in Mathematics, 1186:258–264, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076846/.

Laitinen:1986:UHT

[Lai86] Erkki Laitinen. Unstable homotopy theory of homotopy representations. Lecture Notes in Mathematics, 1217:210–248, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072825/.

Laine:1988:ANL

[Lai88] Ilpo Laine. Axiomatic non-linear potential theories. Lecture Notes in Mathematics, 1344:118–132, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103347/.

Lallement:1988:SAS

[Lal88] Gérard Lallement. Some algorithms for semigroups and monoids presented by a single relation. Lecture Notes in Mathematics, 1320:176–182, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083431/.

Lambek:1988:UAL

[Lam88] J. Lambek. On the unity of algebra and logic. Lecture Notes in Mathematics, 1348:221–229, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081361/.

Lance:1985:SPR

[Lan85a] E. Christopher Lance. Some problems and results on reflexive algebras. Lecture Notes in Mathematics, 1132:324–330, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074892/.

Lang:1985:VC

[Lan85b] Serge Lang. Vojta’s conjecture. Lecture Notes in Mathematics, 1111:407–419, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084600/.

Lange:1986:CFA

[Lan86] L. J. Lange. Continued fraction applications to zero location. Lecture Notes in Mathematics, 1199:220–262, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075940/.

Landis:1988:APT

[Lan88a] E. M. Landis. Application of the potential theory to the study of qualitative properties of solutions of the elliptic and parabolic equations. Lecture Notes in Mathematics, 1344:133–153, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103348/.

Landweber:1988:ECM

[Lan88b] Peter S. Landweber. Elliptic cohomology and modular forms. Lecture Notes in Mathematics, 1326:55–68, 1988. CODEN LN-MAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078038/.

Landweber:1988:EGI

[Lan88c] Peter S. Landweber. Elliptic genera: an introductory overview. Lecture Notes in Mathematics, 1326:1–10, 1988. CODEN LN-MAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078036/.

Landweber:1988:SEC

[Lan88d] Peter S. Landweber. Supersingular elliptic curves and congruences for Legendre polynomials. Lecture Notes in Mathematics, 1326:69–93, 1988. CODEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078039/.

Lapeyre:1988:MPS

[Lap88] Bernard Lapeyre. Majoration a priori des solutions d’équations différentielles stochastiques stables. (French) []. Lecture Notes in Mathematics, 1316:340–351, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081939/.

Larson:1988:SNH

[Lar88] David R. Larson. On similarity of nests in Hilbert space and in Banach spaces. Lecture Notes in Mathematics, 1332:179–194, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081620/.

Lascoux:1986:RDP

[Las86a] Alain Lascoux. La résultante de deux polynômes. (French) []. Lecture Notes in Mathematics, 1220:56–72, 1986. CODEN LN-MAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099505/.

Lasiecka:1986:SRR

[Las86b] Irena Lasiecka. “sharp” regularity results for mixed hyperbolic problems of second order. Lecture Notes in Mathematics, 1223: 160–175, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099191/.

Lascoux:1987:FCF

[Las87] Alain Lascoux. Forme canonique d’une forme binaire. (French) []. Lecture Notes in Mathematics, 1278:44–51, 1987. CODEN LNMAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078805/.

Lasser:1989:MSS

[Las89] R. Lasser. A modification of stationarity for stochastic processes induced by orthogonal polynomials. Lecture Notes in Mathematics, 1379:185–191, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087854/.

Laudal:1986:MMP

[Lau86] O. A. Laudal. Matric Massey products and formal moduli I. Lecture Notes in Mathematics, 1183:218–240, 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075462/.

Laubin:1987:PSA

[Lau87a] Pascal Laubin. Propagation of the second analytic wave front set along diffractive rays. Lecture Notes in Mathematics, 1256: 294–311, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077747/.

Laue:1987:AGW

[Lau87b] Hartmut Laue. On automorphism groups which normalize an Abelian normal subgroup. Lecture Notes in Mathematics, 1281: 73–84, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078692/.

Laville:1985:EDA

[Lav85] Guy Laville. Sur l’équation de Dirac avec champ électromagnétique quelconque. (French) []. Lecture Notes in Mathematics, 1165: 136–149, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076150/.

Laville:1987:CSF

[Lav87] G. Laville. Sur un calcul symbolique de Feynman. (French) [On a symbolic calculation by Feynman]. Lecture Notes in Mathematics, 1295:132–145, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081980/.

Lawrynowicz:1985:SDP

[Ław85] Julian Ławrynowicz, editor. Seminar on Deformations: Proceedings, Łódź–Warsaw 1982/84, volume 1165 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1165. URL http://link.springer.com/book/10.1007/
BFb0076141;   http://www.springerlink.com/content/978-
3-540-39734-2.

Lawvere:1986:I

[Law86] F. William Lawvere. Introduction. Lecture Notes in Mathematics, 1174:1–16, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076929/.

Lawler:1989:CSC

[Law89] Eugene L. Lawler. Combinatorial structures and combinatorial optimization. Lecture Notes in Mathematics, 1403:162–197, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083464/.

Lax:1987:SHD

[Lax87] Peter D. Lax. On symmetrizing hyperbolic differential equations. Lecture Notes in Mathematics, 1270:150–151, 1987. CODEN LN-MAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078324.

Lazer:1988:IMT

[Laz88] Alan Lazer. Introduction to multiplicity theory for boundary value problems with asymmetric nonlinearities. Lecture Notes in Mathematics, 1324:137–165, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100789/.

Loparo:1986:ASI

[LB86] K. A. Loparo and G. L. Blankenship. Almost sure instability of a class of linear stochastic systems with jump process coefficients. Lecture Notes in Mathematics, 1186:160–190, 1986. CODEN LN-MAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076838/.

Light:1985:VDS

[LC85a] William Allan Light and Elliott Ward Cheney. The L_{1}-version of the Diliberto–Straus algorithm. Lecture Notes in Mathematics, 1169:75–90, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075398/.

Light:1985:AG

[LC85b] William Allan Light and Elliott Ward Cheney. The algorithm of von Golitschek. Lecture Notes in Mathematics, 1169:67–74, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075397/.

Light:1985:AA

[LC85c] William Allan Light and Elliott Ward Cheney. The alternating algorithm. Lecture Notes in Mathematics, 1169:48–55, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075394/.

Light:1985:AMR

[LC85d] William Allan Light and Elliott Ward Cheney. Appendix on miscellaneous results in Banach spaces. Lecture Notes in Mathematics, 1169:126–133, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075402/.

Light:1985:ABI

[LC85e] William Allan Light and Elliott Ward Cheney. Appendix on the Bochner integral. Lecture Notes in Mathematics, 1169:113– 125, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075401/.

Light:1985:ATT

[LC85f] William Allan Light and Elliott Ward Cheney. Approximation Theory in Tensor Product Spaces, volume 1169 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 133 (est.) pp. LCCN QA3

.L28 no. 1169. URL http://link.springer.com/book/10.1007/
BFb0075391;   http://www.springerlink.com/content/978-
3-540-39741-0.

Light:1985:CPM

[LC85g] William Allan Light and Elliott Ward Cheney. Central proximity maps. Lecture Notes in Mathematics, 1169:56–59, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075395/.

Light:1985:DSA

[LC85h] William Allan Light and Elliott Ward Cheney. The Diliberto– Straus algorithm in C(S×T). Lecture Notes in Mathematics, 1169: 60–66, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075396/.

Light:1985:EPC

[LC85i] William Allan Light and Elliott Ward Cheney. Estimates of projection constants. Lecture Notes in Mathematics, 1169:91– 102, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075399/.

Light:1985:ITP

[LC85j] William Allan Light and Elliott Ward Cheney. An introduction to tensor product spaces. Lecture Notes in Mathematics, 1169: 1–34, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075392/.

Light:1985:MP

[LC85k] William Allan Light and Elliott Ward Cheney. Minimal projections. Lecture Notes in Mathematics, 1169:103–112, 1985. CODEN LNMAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075400/.

Light:1985:P

[LC85l] William Allan Light and Elliott Ward Cheney. Proximinality. Lecture Notes in Mathematics, 1169:35–47, 1985. CODEN LN-MAA2. ISBN 3-540-16057-4 (print), 3-540-39741-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075393/.

Lelong:1986:SAA

[LDS86] Pierre Lelong, Pierre Dolbeault, and Henri Skoda, editors. Séminaire d’Analyse: Années 1983/1984. (French) [], volume 1198 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1198. URL http:

//link.springer.com/book/10.1007/BFb0077039;   http://
www.springerlink.com/content/978-3-540-38729-9.

LeBruyn:1986:C

[Le 86a] Lieven Le Bruyn. Cohomology of IP_{(2)}^{m}. Lecture Notes in Mathematics, 1220:130–156, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099510/.

leBruyn:1986:PS

[Le 86b] Lieven Le Bruyn. The Poincaré series of Π_{m,2}. Lecture Notes in Mathematics, 1197:90–108, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076316/.

LeJan:1986:HDS

[Le 86c] Y. Le Jan. Haussdorf dimension for the statistical equilibrium of stochastics flows. Lecture Notes in Mathematics, 1158:201– 207, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080218/.

LeJan:1987:TLS

[Le 87] Y. Le Jan. Temps local et superchamp. (French) []. Lecture Notes in Mathematics, 1247:176–190, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077633/.

LeBruyn:1988:CGD

[Le 88a] Lieven Le Bruyn. Centers of generic division algebras and zetafunctions. Lecture Notes in Mathematics, 1328:135–164, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100922/.

LeGall:1988:FPP

[Le 88b] Jean-François Le Gall. Sur les fonctions polaires pour le mouvement brownien. (French) []. Lecture Notes in Mathematics, 1321:186–189, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084136/.

LeJan:1988:FSR

[Le 88c] Yves Le Jan. On the Fock space representation of occupations times for non reversible Markov processes. Lecture Notes in Mathematics, 1322:134–138, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077871/.

LeBruyn:1989:SES

[Le 89] Lieven Le Bruyn. Simultaneous equivalence of square matrices. Lecture Notes in Mathematics, 1404:127–136, 1989. CODEN LN-MAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084074/.

Lopez-Escobar:1985:PFC

[LE85] E. G. K. López-Escobar. Proof functional connectives. Lecture Notes in Mathematics, 1130:208–221, 1985. CODEN LN-MAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075313/.

Leandre:1987:DTP

[Léa87] Rémi Léandre. Densitè en temps petit d’un processus de sauts. (French) []. Lecture Notes in Mathematics, 1247:81–99, 1987. CO-DEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077628/.

Leandre:1988:AQG

[Léa88] Rémi Léandre. Applications quantitatives et géométriques du calcul de Malliavin. (French) []. Lecture Notes in Mathematics, 1322:109–133, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077870/.

Lech:1986:MCB

[Lec86a] Christer Lech. A method for constructing bad Noetherian local rings. Lecture Notes in Mathematics, 1183:241–247, 1986. CODEN LNMAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075463/.

Lech:1986:YAP

[Lec86b] Crister Lech. Yet another proof of a result by Ogoma. Lecture Notes in Mathematics, 1183:248–249, 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075464.

Ledoux:1986:SBC

[Led86a] Michel Ledoux. On the small balls condition in the central limit theorem in uniformly convex spaces. Lecture Notes in Mathematics, 1193:44–52, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077099/.

Ledrappier:1986:PES

[Led86b] F. Ledrappier. Positivity of the exponent for stationary sequences of matrices. Lecture Notes in Mathematics, 1186:56–73, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076833/.

Lehmann:1986:MMR

[Leh86] Daniel Lehmann. Modele minimal relatif des feuilletages. Lecture Notes in Mathematics, 1183:250–258, 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075465/.

Lenstra:1985:GTP

[Len85] H. W. Lenstra, Jr. Galois theory and primality testing. Lecture Notes in Mathematics, 1142:169–189, 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074800/.

Lenzing:1986:CSA

[Len86] Helmut Lenzing. Curve singularities arising from the representation theory of tame hereditary algebras. Lecture Notes in Mathematics, 1177:199–231, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075266/.

Lennox:1987:SGN

[Len87] John C. Lennox. Soluble groups with nilpotent-extensible subgroups. Lecture Notes in Mathematics, 1281:103–106, 1987. CO-DEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078694/.

LePage:1989:AMI

[LeP89a] Raoul LePage. Appendix multidimensional infinitely divisible variables and processes. Part I: Stable case. Lecture Notes in Mathematics, 1391:153–163, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083389/.

LePage:1989:CMC

[LeP89b] Raoul LePage. Conditional moments for coordinates of stable vectors. Lecture Notes in Mathematics, 1391:148–152, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083388/.

Lerman:1985:OCH

[Ler85] Manuel Lerman. On the ordering of classes in high/low hierarchies. Lecture Notes in Mathematics, 1141:260–270, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076225/.

Leroy:1986:DAA

[Ler86] A. Leroy. S-dérivations algébriques sur les anneaux premiers. (French) []. Lecture Notes in Mathematics, 1197:114–120, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076318/.

Lerat:1989:DSN

[Ler89] A. Lerat. Difference schemes for nonlinear hyperbolic systems — a general framework. Lecture Notes in Mathematics, 1402: 12–29, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083865/.

Lescot:1988:SPM

[Les88] Jack Lescot. Séries de Poincaré des modules multi-gradués sur les anneaux monômiaux. (French) []. Lecture Notes in Mathematics, 1318:155–161, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077801/.

Letac:1989:PCF

[Let89] Gérard Letac. Le problème de la classification des familles exponentielles naturelles de R^{d} ayant une fonction variance quadratique. (French) []. Lecture Notes in Mathematics, 1379:192– 216, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087855/.

Levasseur:1985:CBA

[Lev85a] Thierry Levasseur. Complexe bidualisant en algèbre non commutative. (French) []. Lecture Notes in Mathematics, 1146:270– 287, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074541/.

Levine:1985:CCP

[Lev85b] Harold Levine. Canonical coordinatized product neighborhoods (C^{2} pn). Lecture Notes in Mathematics, 1157:25–46, 1985. CO-DEN LNMAA2. ISBN 3-540-15995-9 (print), 3-540-39700-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075069/.

Levine:1985:CIS

[Lev85c] Harold Levine. Classifying Immersions into R^{4} over Stable Maps of 3-Manifolds into R^{2}, volume 1157 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15995-9 (print), 3-540-39700-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 163 pp. LCCN QA3 .L28

no. 1157. URL http://link.springer.com/book/10.1007/
BFb0075066;   http://www.springerlink.com/content/978-
3-540-39700-7.

Levine:1985:I

[Lev85d] Harold Levine. Introduction. Lecture Notes in Mathematics, 1157: 1–5, 1985. CODEN LNMAA2. ISBN 3-540-15995-9 (print), 3-540-39700-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075067/.

Levine:1985:LSM

[Lev85e] Harold Levine. Lifting stable maps of three manifolds into the plane to immersions in R^{4}. Lecture Notes in Mathematics, 1157: 47–128, 1985. CODEN LNMAA2. ISBN 3-540-15995-9 (print), 3-540-39700-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075070/.

Levine:1985:SSM

[Lev85f] Harold Levine. The singularities of stable maps of three manifolds into the plane. Lecture Notes in Mathematics, 1157:6–24, 1985. CODEN LNMAA2. ISBN 3-540-15995-9 (print), 3-540-39700-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075068/.

Levine:1985:LGH

[Lev85g] J. P. Levine. Lectures on groups of homotopy spheres. Lecture Notes in Mathematics, 1126:62–95, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074439/.

Levitt:1985:SRL

[Lev85h] Norman Levitt. Some remarks on local formulae for p_{1}. Lecture Notes in Mathematics, 1126:96–126, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074440/.

Levy:1985:EPL

[Lev85i] Y. Levy. Ergodic properties of the Lozi map. Lecture Notes in Mathematics, 1109:103–116, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101539/.

Levine:1988:IBL

[Lev88] J. P. Levine. The μ̄-invariants of based links. Lecture Notes in Mathematics, 1350:87–103, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081469/.

Levasseur:1989:ODS

[Lev89a] Thierry Levasseur. Operateurs differentiels sur les surfaces munies d’une bonne C*-action. (French) []. Lecture Notes in Mathematics, 1404:269–295, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084080/.

Levitt:1989:EPD

[Lev89b] Norman Levitt. Equivariant piecewise differentiable immersions. Lecture Notes in Mathematics, 1366:181–187, 1989. CODEN LN-MAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085004/.

Levitt:1989:FLP

[Lev89c] Norman Levitt. Formal links and the PL Grassmannian G_{n,k}. Lecture Notes in Mathematics, 1366:43–59, 1989. CODEN LN-MAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084997/.

Levitt:1989:GPS

[Lev89d] Norman Levitt. The Grassmannian for piecewise smooth immersions. Lecture Notes in Mathematics, 1366:116–160, 1989. CODEN LNMAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085002/.

Levitt:1989:ITS

[Lev89e] Norman Levitt. The immersion theorem for subcomplexes of G_{n,k}. Lecture Notes in Mathematics, 1366:70–86, 1989. CODEN LN-MAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084999/.

Levitt:1989:IER

[Lev89f] Norman Levitt. Immersions equivariant with respect to orthogonal actions on R^{n+k}. Lecture Notes in Mathematics, 1366:87– 100, 1989. CODEN LNMAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085000/.

Levitt:1989:ITM

[Lev89g] Norman Levitt. Immersions into triangulated manifolds (with R. Mladineo). Lecture Notes in Mathematics, 1366:101–115, 1989. CODEN LNMAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085001/.

Levitt:1989:I

[Lev89h] Norman Levitt. Introduction. Lecture Notes in Mathematics, 1366:1–10, 1989. CODEN LNMAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084995/.

Levitt:1989:LFC

[Lev89i] Norman Levitt. Local formulae for characteristic classes. Lecture Notes in Mathematics, 1366:11–42, 1989. CODEN LN-MAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084996/.

Levitt:1989:PDI

[Lev89j] Norman Levitt. Piecewise differentiable immersions into Riemannian manifolds. Lecture Notes in Mathematics, 1366:188– 197, 1989. CODEN LNMAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085005/.

Levitt:1989:SAS

[Lev89k] Norman Levitt. Some applications to smoothing theory. Lecture Notes in Mathematics, 1366:161–180, 1989. CODEN LN-MAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085003/.

Levitt:1989:SVC

[Lev89l] Norman Levitt. Some variations of the G_{n,k} construction. Lecture Notes in Mathematics, 1366:60–69, 1989. CODEN LN-MAA2. ISBN 3-540-50756-6 (print), 3-540-46078-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084998/.

Lewis:1986:EAB

[Lew86a] J. T. Lewis. An elementary approach to Brownian motion on manifolds. Lecture Notes in Mathematics, 1158:158–167, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080215/.

Lewis:1986:TS

[Lew86b] L. G. Lewis, Jr. Thom spectra. Lecture Notes in Mathematics, 1213:407–449, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075788/.

Lewis:1987:ASF

[Lew87] John L. Lewis. Approximations of Sobolev functions and related topics. Lecture Notes in Mathematics, 1275:223–234, 1987. CO-DEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078355/.

Lewis:1988:LDP

[Lew88a] J. T. Lewis. The large deviation principle in statistical mechanics: an expository account. Lecture Notes in Mathematics, 1325: 141–155, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077923/.

Lewis:1988:GEO

[Lew88b] L. Gaunce Lewis, Jr. The RO(G)-graded equivariant ordinary cohomology of complex projective spaces with linear Z/p actions. Lecture Notes in Mathematics, 1361:53–122, 1988. CODEN LN-MAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083034/.

Lex:1988:RAL

[Lex88] Wilfried Lex. Remarks on acts and the lattice of their torsion theories. Lecture Notes in Mathematics, 1320:183–193, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083432/.

Larrouturou:1989:EMC

[LF89] B. Larrouturou and L. Fézoui. On the equations of multicomponent perfect of real gas inviscid flow. Lecture Notes in Mathematics, 1402:69–98, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083869/.

Li:1986:CG

[Li86] Huiling Li. On 2-connectedness of geometries. Lecture Notes in Mathematics, 1185:308–313, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076181/.

Li:1988:FBP

[Li88] Huilai Li. Free boundary problems for degenerate parabolic equations. Lecture Notes in Mathematics, 1306:102–130, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082928/.

Li:1989:AMS

[Li89a] An-Min Li. Affine maximal surfaces and harmonic functions. Lecture Notes in Mathematics, 1369:142–151, 1989. CODEN LN-MAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087530/.

Li:1989:SFM

[Li89b] Chao-Qun Li. On solutions of frame mappings into manifolds. Lecture Notes in Mathematics, 1411:143–147, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086449/.

Lichnerowicz:1985:GDS

[Lic85] André Lichnerowicz. Geometry of dynamical systems with timedependent constraints and time-dependent Hamiltonians: an approach toward quantization. Lecture Notes in Mathematics, 1139: 122–135, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074580/.

Lichnerowicz:1986:RCQ

[Lic86] André Lichnerowicz. Représentation coadjointe quotient et espaces homogènes de contact. (French) []. Lecture Notes in Mathematics, 1209:210–221, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076633/.

Liess:1987:BRO

[Lie87] Otto Liess. Boundary regularity for one-sided solutions of linear partial differential equations with analytic coefficients. Lecture Notes in Mathematics, 1256:312–329, 1987. CODEN LN-MAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077748/.

Liess:1988:RSC

[Lie88] Otto Liess. Regularity of solutions of Cauchy problems with smooth Cauchy data. Lecture Notes in Mathematics, 1324:166– 177, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100790/.

Ligocka:1985:OPB

[Lig85a] Ewa Ligocka. An open problem on boundary behaviour of holomorphic mappings. Lecture Notes in Mathematics, 1165:196, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076153.

Ligocka:1985:RWB

[Lig85b] Ewa Ligocka. The regularity of the weighted Bergman projections. Lecture Notes in Mathematics, 1165:197–203, 1985. CODEN LN-MAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076154/.

Liggett:1986:NPS

[Lig86] Thomas M. Liggett. Nearest particle systems: Results and open problems. Lecture Notes in Mathematics, 1212:200–215, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076250/.

Lima:1985:PSP

[Lim85] R. Lima. Propriétés spectrales pour des hamiltoniens presquepériodiques. (French) []. Lecture Notes in Mathematics, 1109: 117–135, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101540/.

deLin:1986:QIK

[Lin86a] Cheng de Lin. Quand l’inégalité de Kunita–Watanabe est-elle une égalité?. (French) []. Lecture Notes in Mathematics, 1204: 40–47, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075710/.

Lindstrom:1986:NAP

[Lin86b] Tom Lindstrøm. Nonstandard analysis and perturbations of the Laplacian along Brownian paths. Lecture Notes in Mathematics, 1158:180–200, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080217/.

Lindblad:1988:DEQ

[Lin88a] Göran Lindblad. Dynamical entropy for quantum systems. Lecture Notes in Mathematics, 1303:183–191, 1988. CODEN LN-MAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078062/.

Lindenstrauss:1988:SUF

[Lin88b] J. Lindenstrauss. Some useful facts about Banach spaces. Lecture Notes in Mathematics, 1317:185–200, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081740/.

Lions:1985:OCV

[Lio85] P.-L. Lions. Optimal control and viscosity solutions. Lecture Notes in Mathematics, 1119:94–112, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074782/.

Lions:1986:DSU

[Lio86] J.-L. Lions. Distributed systems with uncomplete data and Lagrange multipliers. Lecture Notes in Mathematics, 1190:273– 305, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076711/.

Lions:1989:VSF

[Lio89] P.-L. Lions. Viscosity solutions of fully nonlinear second order equations and optimal stochastic control in infinite dimensions. Part II: Optimal control of Zakai’s equation. Lecture Notes in Mathematics, 1390:147–170, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083943/.

Lisica:1987:APD

[Lis87] Yu. T. Lisica. The Alexander–Pontryagin duality theorem for coherent homology and cohomology with coefficients in sheaves of modules. Lecture Notes in Mathematics, 1283:148–163, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081425/.

Lisitsky:1989:SRV

[Lis89] A. D. Lisitsky. Stable random vectors in Hilbert space. Lecture Notes in Mathematics, 1412:172–182, 1989. CODEN LN-MAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084172/.

Litherland:1985:STS

[Lit85] R. A. Litherland. Symmetries of twist-spun knots. Lecture Notes in Mathematics, 1144:97–107, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075013/.

Littlejohn:1988:OPS

[Lit88] L. L. Littlejohn. Orthogonal polynomial solutions to ordinary and partial differential equations. Lecture Notes in Mathematics, 1329:98–124, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083355/.

Liulevicius:1988:RLS

[Liu88] Arunas Liulevicius. A reciprocity law for symmetric products of Gsets. Lecture Notes in Mathematics, 1350:266–269, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081480/.

Livsic:1987:CNO

[Liv87] Moshe S. Livšic. Commuting nonselfadjoint operators and collective motions of systems. Lecture Notes in Mathematics, 1272: 1–38, 1987. CODEN LNMAA2. ISBN 3-540-18316-7 (print), 3-540-47877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078926/.

Ljapin:1988:SEP

[Lja88] E. S. Ljapin. Semigroup extensions of partial groupoids. Lecture Notes in Mathematics, 1320:205–217, 1988. CODEN LN-MAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083434/.

Lassner:1985:QAG

[LL85] G. Lassner and G. A. Lassner. Quasi-*-algebras and general Weyl quantization. Lecture Notes in Mathematics, 1139:108–121, 1985.

CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074579/.

Labelle:1986:CEP

[LL86] Gilbert Labelle and Pierre Leroux, editors. Combinatoire énumérative: Proceedings of the “Colloque de combinatoire énumérative”, held at Université du Québec à Montréal, May 28–June 1, 1985, volume 1234 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1234. URL http:

//link.springer.com/book/10.1007/BFb0072503;   http://
www.springerlink.com/content/978-3-540-47402-9.

Lewandowski:1989:RTC

[LL89] M. Lewandowski and W. Linde. Real translates of complex measures. Lecture Notes in Mathematics, 1391:164–177, 1989. CO-DEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0083390/.

Livingston:1985:AIS

[LM85] Charles Livingston and Paul Melvin. Abelian invariants of satellite knots. Lecture Notes in Mathematics, 1167:217–227, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075225/.

Lai:1986:VTI

[LM86a] King-Fai Lai and Ngaiming Mok. On a vanishing theorem on irreducible quotients of finite volume of polydiscs. Lecture Notes in Mathematics, 1198:163–171, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077051/.

Ledoux:1986:SRU

[LM86b] M. Ledoux and M. B. Marcus. Some remarks on the uniform convergence of Gaussian and Rademacher Fourier quadratic forms. Lecture Notes in Mathematics, 1193:53–72, 1986. CODEN LN-MAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077100/.

Lewis:1986:CUS

[LM86c] L. G. Lewis, Jr. and J. P. May. Change of universe, smash products, and change of groups. Lecture Notes in Mathematics, 1213: 54–116, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075781/.

Lewis:1986:EDT

[LM86d] L. G. Lewis, Jr. and J. P. May. Equivariant duality theory. Lecture Notes in Mathematics, 1213:117–174, 1986. CODEN LN-MAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075782/.

Lewis:1986:ESC

[LM86e] L. G. Lewis, Jr. and J. P. May. The equivariant stable category. Lecture Notes in Mathematics, 1213:6–53, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075780/.

Lewis:1986:ETS

[LM86f] L. G. Lewis, Jr. and J. P. May. Equivariant Thom spectra. Lecture Notes in Mathematics, 1213:450–474, 1986. CODEN LN-MAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075789/.

Lewis:1986:ET

[LM86g] L. G. Lewis, Jr. and J. P. May. Equivariant transfer. Lecture Notes in Mathematics, 1213:175–235, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075783/.

Longobardi:1987:NGC

[LM87] Patrizia Longobardi and Mercede Maj. On the nilpotence of groups with a certain lattice of normal subgroups. Lecture Notes in Mathematics, 1281:107–114, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078695/.

Lickorish:1988:EPL

[LM88a] W. B. R. Lickorish and K. C. Millett. An evaluation of the F - polynomial of a link. Lecture Notes in Mathematics, 1350:104– 108, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081470/.

Lindsay:1988:IKA

[LM88b] Martin Lindsay and Hans Maassen. An integral kernel approach to noise. Lecture Notes in Mathematics, 1303:192–208, 1988. CO-DEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078063/.

Leandre:1989:DDC

[LM89a] R. Léandre and P.-A. Meyer. Sur le développement d’une diffusion en chaos de Wiener. (French) []. Lecture Notes in Mathematics, 1372:161–164, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083968/.

Loridan:1989:RTL

[LM89b] P. Loridan and J. Morgan. ϵ-Regularized two-level optimization problems: approximation and existence results. Lecture Notes in Mathematics, 1405:99–113, 1989. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Lewis:1986:BRS

[LMM86] L. G. Lewis, Jr., J. P. May, and J. E. McClure. The Burnside ring and splittings in equivariant homology theory. Lecture Notes in Mathematics, 1213:236–298, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075784/.

Lewis:1986:ORS

[LMS86a] L. G. Lewis, Jr., J. P. May, and M. Steinberger. Operad ring spectra. Lecture Notes in Mathematics, 1213:350–383, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075786/.

Lewis:1986:THS

[LMS86b] L. G. Lewis, Jr., J. P. May, and M. Steinberger. Twisted half smash products and extended powers. Lecture Notes in Mathematics, 1213:299–349, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075785/.

Lewis:1986:ESH

[LMS86c] L. Gaunce Lewis, Jr., J. Peter May, and Mark Steinberger. Equivariant Stable Homotopy Theory, volume 1213 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 474 (est.) pp. LCCN QA3

.L28 no. 1213. URL http://link.springer.com/book/10.1007/
BFb0075778;   http://www.springerlink.com/content/978-
3-540-47077-9.

Lukes:1986:BBS

[LMZ86a] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Baire and Blumberg spaces. Lecture Notes in Mathematics, 1189:133–142, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075899/.

Lukes:1986:BOS

[LMZ86b] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Base operator spaces. Lecture Notes in Mathematics, 1189:5–37, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075896/.

Lukes:1986:BBF

[LMZ86c] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Boundary behaviour of functions. Lecture Notes in Mathematics, 1189:287– 317, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075904/.

Lukes:1986:CPF

[LMZ86d] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Connectivity properties of fine topologies. Lecture Notes in Mathematics, 1189:142– 147, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075900/.

Lukes:1986:DT

[LMZ86e] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Density topologies. Lecture Notes in Mathematics, 1189:147–229, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075901/.

Lukes:1986:E

[LMZ86f] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Errata. Lecture Notes in Mathematics, 1189:e1–e2, 1986. CODEN LN-MAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075910.

Lukes:1986:ET

[LMZ86g] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Extension theorems. Lecture Notes in Mathematics, 1189:259–287, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075903/.

Lukes:1986:FDP

[LMZ86h] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. The fine Dirichlet problem. Lecture Notes in Mathematics, 1189:388–407, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075908/.

Lukes:1986:FTP

[LMZ86i] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Fine topology in potential theory. Lecture Notes in Mathematics, 1189:317–347, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075905/.

Lukes:1986:FTM

[LMZ86j] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Fine Topology Methods in Real Analysis and Potential Theory, volume 1189 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 472 pp. LCCN QA3 .L28 no. 1189. URL http:

//link.springer.com/book/10.1007/BFb0075894;   http://
www.springerlink.com/content/978-3-540-39814-1.

Lukes:1986:FHF

[LMZ86k] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Finely hyperharmonic functions. Lecture Notes in Mathematics, 1189:365– 388, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075907/.

Lukes:1986:FEF

[LMZ86l] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Further examples of fine topologies. Lecture Notes in Mathematics, 1189:229– 259, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075902/.

Lukes:1986:GPF

[LMZ86m] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. General properties of fine topologies. Lecture Notes in Mathematics, 1189:38– 69, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075897/.

Lukes:1986:HS

[LMZ86n] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Harmonic spaces. Lecture Notes in Mathematics, 1189:347–365, 1986. CODEN LN-MAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075906/.

Lukes:1986:LMP

[LMZ86o] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. The Lusin– Menchoff property of fine topologies. Lecture Notes in Mathematics, 1189:70–132, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075898/.

Lukes:1986:P

[LMZ86p] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Prologue. Lecture Notes in Mathematics, 1189:1–4, 1986. CODEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075895/.

Lukes:1986:QTM

[LMZ86q] Jaroslav Lukeš, Jan Malý, and Luděk Zajíček. Quasi-topological methods. Lecture Notes in Mathematics, 1189:407–423, 1986. CO-DEN LNMAA2. ISBN 3-540-16474-X (print), 3-540-39814-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075909/.

Luo:1987:ABC

[LN87] L. S. Luo and J. Nuttall. Asymptotic behavior of the Christoffel function related to a certain unbounded set. Lecture Notes in Mathematics, 1287:105–116, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078900/.

Lin:1988:DCS

[LN88] Chang-Shou Lin and Wei-Ming Ni. On the diffusion coefficient of a semilinear Neumann problem. Lecture Notes in Mathematics, 1340:160–174, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082894/.

Leutbecher:1989:CEU

[LN89] Armin Leutbecher and Gerhard Niklasch. On cliques of exceptional units and Lenstra’s construction of Euclidean fields. Lecture Notes in Mathematics, 1380:150–178, 1989. CODEN LN-MAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086551/.

Liulevicius:1986:DOS

[LÖ86] Arunas Liulevicius and Murad Özaydın. Duality in orbit spaces. Lecture Notes in Mathematics, 1217:249–252, 1986. CODEN LN-MAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072826/.

Loeb:1986:PCO

[Loe86] Jean-Jacques Loeb. Pseudo-convexité des ouverts invariants et convexité géodésique dans certains espaces symétriques. (French) []. Lecture Notes in Mathematics, 1198:172–190, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077052/.

Loeser:1987:EIT

[Loe87] F. Loeser. Évaluation d’intégrales et théorie de Hodge. (French) []. Lecture Notes in Mathematics, 1246:125–142, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077535/.

Long:1985:CPA

[Lon85] D. D. Long. Constructing pseudo-Anosov maps. Lecture Notes in Mathematics, 1144:108–114, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075014/.

Longtin:1986:CNC

[Lon86] André Longtin. Une combinatoire non-commutative pour l’étude des nombres secants. (French) []. Lecture Notes in Mathematics, 1234:246–266, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072519/.

Longinetti:1988:SII

[Lon88] Marco Longinetti. Some isoperimetric inequalities for the level curves of capacity and Green’s functions on convex plane domains. Lecture Notes in Mathematics, 1340:175–192, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082895/.

Longobardi:1989:GPP

[Lon89] Patrizia Longobardi. On groups with a permutational property on commutators. Lecture Notes in Mathematics, 1398:110–116, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086247/.

Lopez:1988:STP

[Lóp88] G. López. Szegő’s Theorem for polynomials orthogonal with respect to varying measures. Lecture Notes in Mathematics, 1329: 255–260, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083365/.

Lorenz:1986:AA

[Lor86] Martin Lorenz. On affine algebras. Lecture Notes in Mathematics, 1197:121–126, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076319/.

Lorenz:1988:FRS

[Lor88] Martin Lorenz. Frobenius reciprocity and G_{0} of skew group rings. Lecture Notes in Mathematics, 1328:165–172, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100923/.

Losert:1985:GFF

[Los85a] Prof. Dr. Viktor Losert. Gleichverteilte Folgen Und Folgen, Für Die Fast Alle Teilfolgen Gleichverteilt Sind. (German) []. Lecture Notes in Mathematics, 1114:84–97, 1985. CODEN LN-MAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101648/.

Losert:1985:GFD

[Los85b] Prof. Dr. Viktor Losert. Gleichverteilung Von Folgen, Die Durch Additive Halbgruppen Definiert Sind. (German) []. Lecture Notes in Mathematics, 1114:77–83, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101647/.

Letta:1986:PAL

[LP86] Giorgio Letta and Maurizio Pratelli, editors. Probability and Analysis: Lectures given at the 1st 1985 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Varenna (Como), Italy May 31–June 8, 1985, volume 1206 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1206; QA3 .L28 no. 1206 QA273.18. URL http:

//link.springer.com/book/10.1007/BFb0076298;   http://
www.springerlink.com/content/978-3-540-40955-7.

LeBruyn:1987:ELS

[LP87a] Lieven Le Bruyn and Claudio Procesi. Etale local structure of matrix invariants and concomitants. Lecture Notes in Mathematics, 1271:143–175, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079236/.

Leinen:1987:ACG

[LP87b] Felix Leinen and Richard E. Phillips. Algebraically closed groups in locally finite group classes. Lecture Notes in Mathematics, 1281:

85–102, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078693/.

Laudal:1988:AIH

[LP88a] Olav Arnfinn Laudal and Gerhard Pfister. Applications to isolated hypersurface singularities. Lecture Notes in Mathematics, 1310: 61–71, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078944/.

Laudal:1988:AFM

[LP88b] Olav Arnfinn Laudal and Gerhard Pfister. Automorphisms of the formal moduli. Lecture Notes in Mathematics, 1310:15–31, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078942/.

Laudal:1988:GCL

[LP88c] Olav Arnfinn Laudal and Gerhard Pfister. The generic component of the local moduli suite. Lecture Notes in Mathematics, 1310:88–104, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078946/.

Laudal:1988:I

[LP88d] Olav Arnfinn Laudal and Gerhard Pfister. Introduction. Lecture Notes in Mathematics, 1310:1–7, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078940/.

Laudal:1988:KSM

[LP88e] Olav Arnfinn Laudal and Gerhard Pfister. The Kodaira–Spencer map and its kernel. Lecture Notes in Mathematics, 1310:32– 60, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078943/.

Laudal:1988:MSX

[LP88f] Olav Arnfinn Laudal and Gerhard Pfister. The moduli suite of x_{1}^{5}+ x_{2}^{11}. Lecture Notes in Mathematics, 1310:105–111, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078947/.

Laudal:1988:PCS

[LP88g] Olav Arnfinn Laudal and Gerhard Pfister. Plane curve singularities with k*-action. Lecture Notes in Mathematics, 1310:72– 87, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078945/.

Laudal:1988:PSF

[LP88h] Olav Arnfinn Laudal and Gerhard Pfister. The prorepresenting substratum of the formal moduli. Lecture Notes in Mathematics, 1310:8–14, 1988. CODEN LNMAA2. ISBN 3-540-19235-2 (print), 3-540-39153-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078941/.

Lawrynowicz:1985:HPE

[ŁR85] Julian Ławrynowicz and Jakub Rembieliński. Hurwitz pairs equipped with complex structures. Lecture Notes in Mathematics, 1165:184–195, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076152/.

Lopez:1988:RAO

[LR88a] G. López and E. A. Rakhmanov. Rational approximations, orthogonal polynomials and equilibrium distributions. Lecture Notes in Mathematics, 1329:125–157, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083356/.

Lorch:1988:SCH

[LR88b] Lee Lorch and Dennis Russell. On some contributions of Halasz to the Turan power-sum theory. Lecture Notes in Mathematics, 1354:169–177, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089594/.

Luck:1988:ST

[LR88c] W. Lück and A. Ranicki. Surgery transfer. Lecture Notes in Mathematics, 1361:167–246, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083036/.

Lee:1989:SMM

[LR89a] Kyung Bai Lee and Frank Raymond. Seifert manifolds modelled on principal bundles. Lecture Notes in Mathematics, 1375:207– 215, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085611/.

Levin:1989:NDT

[LR89b] V. L. Levin and S. T. Rachev. New duality theorems for marginal problems with some applications in stochastics. Lecture Notes in Mathematics, 1412:137–171, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084171/.

Liedl:1985:ITF

[LRT85] Roman Liedl, Ludwig Reich, and György Targonski, editors. Iteration Theory and its Functional Equations: Proceedings of the International Symposium held at Schloss Hofen (Lochau), Austria Sept. 28–Oct. 1, 1984, volume 1163 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1163. URL http:

//link.springer.com/book/10.1007/BFb0076410;   http://
www.springerlink.com/content/978-3-540-39749-6.

Lascoux:1985:INP

[LS85a] A. Lascoux and M.-P. Schützenberger. Interpolation de Newton à plusieurs variables. (French) []. Lecture Notes in Mathematics, 1146:161–175, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074538/.

LePage:1985:SRL

[LS85b] Raoul LePage and Bertram M. Schreiber. A square root law for diffusing particles. Lecture Notes in Mathematics, 1153:323– 328, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074958/.

Lawvere:1986:CCP

[LS86a] F. William Lawvere and Stephen H. Schanuel, editors. Categories in Continuum Physics: Lectures given at a Workshop held at SUNY, Buffalo 1982, volume 1174 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1174. URL http:

//link.springer.com/book/10.1007/BFb0076928;   http://
www.springerlink.com/content/978-3-540-39760-1.

Ledrappier:1986:EEB

[LS86b] F. Ledrappier and J.-M. Strelcyn. The estimation of entropy from below through Lyapunov characteristic exponents. Lecture Notes in Mathematics, 1222:157–179, 1986. CODEN LNMAA2. ISBN 3-540-17190-8 (print), 3-540-47349-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099034/.

Landweber:1987:DRI

[LS87a] Peter S. Landweber and Robert E. Stong. The depth of rings of invariants over finite fields. Lecture Notes in Mathematics, 1240:259–274, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072984/.

Lannes:1987:SI

[LS87b] Jean Lannes and Lionel Schwartz. On the structure of the muinjectives. Lecture Notes in Mathematics, 1298:207–214, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083011/.

Levin:1987:SEA

[LS87c] A. L. Levin and E. B. Saff. Some examples in approximation on the unit disk by reciprocals of polynomials. Lecture Notes in Mathematics, 1287:70–82, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078898/.

Shen:1987:BCY

[lS87d] Chun li Shen. C^{k}-bound of curvatures in Yang–Mills theory. Lecture Notes in Mathematics, 1255:88–95, 1987. CODEN LN-MAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077684/.

Lindenstrauss:1987:RBD

[LS87e] Joram Lindenstrauss and Andrzej Szankowski. The relation between the distance and the weak distance for spaces with a symmetric basis. Lecture Notes in Mathematics, 1267:21–38, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078134/.

Lubinsky:1987:SAE

[LS87f] D. S. Lubinsky and E. B. Saff. Strong asymptotics for L_{p} extremal polynomials (1 < p ≤ ∞) associated with weights on [−1, 1]. Lecture Notes in Mathematics, 1287:83–104, 1987. CODEN LN-MAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078899/.

Lacomba:1988:BTK

[LS88a] Ernesto A. Lacomba and Guillermo Sienra. Blow up techniques in the Kepler problem. Lecture Notes in Mathematics, 1345:177– 191, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081402/.

Laurence:1988:SRR

[LS88b] Peter Laurence and Edward Stredulinsky. A survey of recent regularity results for second order queer differential equations. Lecture Notes in Mathematics, 1340:146–159, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082893/.

Lubinsky:1988:ACWa

[LS88c] Doron S. Lubinsky and Edward B. Saff. Approximation by certain weighted polynomials, I. Lecture Notes in Mathematics, 1305: 80–90, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082424/.

Lubinsky:1988:ACWb

[LS88d] Doron S. Lubinsky and Edward B. Saff. Approximation by certain weighted polynomials, II. Lecture Notes in Mathematics, 1305: 91–110, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082425/.

Lubinsky:1988:BFB

[LS88e] Doron S. Lubinsky and Edward B. Saff. Bernstein’s formula and Bernstein extremal polynomials. Lecture Notes in Mathematics, 1305:111–119, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082426/.

Lubinsky:1988:COP

[LS88f] Doron S. Lubinsky and Edward B. Saff. The case p = 2: Orthonormal polynomials. Lecture Notes in Mathematics, 1305:136– 145, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082429/.

Lubinsky:1988:FP

[LS88g] Doron S. Lubinsky and Edward B. Saff. Further properties of U_{n,R}(x). Lecture Notes in Mathematics, 1305:67–72, 1988. CO-DEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082422/.

Lubinsky:1988:IFR

[LS88h] Doron S. Lubinsky and Edward B. Saff. Infinite-finite range inequalities and their sharpness. Lecture Notes in Mathematics, 1305:49–56, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082420/.

Lubinsky:1988:IE

[LS88i] Doron S. Lubinsky and Edward B. Saff. Integral equations. Lecture Notes in Mathematics, 1305:28–39, 1988. CODEN LN-MAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082418/.

Lubinsky:1988:I

[LS88j] Doron S. Lubinsky and Edward B. Saff. Introduction. Lecture Notes in Mathematics, 1305:1–11, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082414/.

Lubinsky:1988:LZE

[LS88k] Doron S. Lubinsky and Edward B. Saff. The largest zeros of extremal polynomials. Lecture Notes in Mathematics, 1305:57– 66, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082421/.

Lubinsky:1988:TRA

[LS88l] Doron S. Lubinsky and Edward B. Saff. N th root asymptotics for extremal polynomials. Lecture Notes in Mathematics, 1305: 73–79, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082423/.

Lubinsky:1988:NIN

[LS88m] Doron S. Lubinsky and Edward B. Saff. Notation and index of notation. Lecture Notes in Mathematics, 1305:12–16, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082415/.

Lubinsky:1988:PAP

[LS88n] Doron S. Lubinsky and Edward B. Saff. Polynomial approximation of potentials. Lecture Notes in Mathematics, 1305:40–48, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082419/.

Lubinsky:1988:PA

[LS88o] Doron S. Lubinsky and Edward B. Saff. Proof of the asymptotics for E_{np}(W). Lecture Notes in Mathematics, 1305:120–127, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082427/.

Lubinsky:1988:PAE

[LS88p] Doron S. Lubinsky and Edward B. Saff. Proof of the asymptotics for the L_{p} extremal polynomials. Lecture Notes in Mathematics, 1305:128–135, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082428/.

Lubinsky:1988:SMR

[LS88q] Doron S. Lubinsky and Edward B. Saff. Statement of main results. Lecture Notes in Mathematics, 1305:17–23, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082416/.

Lubinsky:1988:WPZ

[LS88r] Doron S. Lubinsky and Edward B. Saff. Weighted polynomials and zeros of extremal polynomials. Lecture Notes in Mathematics, 1305:24–27, 1988. CODEN LNMAA2. ISBN 3-540-18958-0 (print), 3-540-38857-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082417/.

Shen:1989:HMR

[lS89] Chun li Shen. On the holomorphic maps from Riemann surfaces to Grassmannians. Lecture Notes in Mathematics, 1369: 225–234, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087536/.

Louveau:1988:SBW

[LSR88] Alain Louveau and Jean Saint-Raymond. The strength of Borel Wadge determinacy. Lecture Notes in Mathematics, 1333:1– 30, 1988. CODEN LNMAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084967/.

Laurent-Thiebaut:1985:RVF

[LT85] Christine Laurent-Thiébaut. Régularisation sur une variété. (French) []. Lecture Notes in Mathematics, 1165:150–183, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076151/.

Laurent-Thiebaut:1987:TBM

[LT87]  Christine Laurent-Thiébaut.  Transformation de Bochner–

Martinelli dans une variété de Stein. (French) []. Lecture Notes in Mathematics, 1295:96–131, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081979/.

Lippert:1988:RSF

[LT88] D. Lippert and W. Thomas. Relativized star-free expressions, first-order logic, and a concatenation game. Lecture Notes in Mathematics, 1320:194–204, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083433/.

Li:1989:CIL

[LT89] Bang-He Li and Zizhou Tang. Codimension 1 and 2 immersions of lens spaces. Lecture Notes in Mathematics, 1369:152–163, 1989.

CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087531/.

Lubinsky:1985:EEF

[Lub85] D. S. Lubinsky. Even entire functions absolutely monotone in [0, ∞) and weights on the whole real line. Lecture Notes in Mathematics, 1171:221–229, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076547/.

Luck:1988:ED

[Lüc88] Wolfgang Lück. The equivariant degree. Lecture Notes in Mathematics, 1361:123–166, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083035/.

Luck:1989:MG

[Lüc89a] Wolfgang Lück. RΓ-modules and geometry. Lecture Notes in Mathematics, 1408:305–412, 1989. CODEN LNMAA2. ISBN 3-540-51846-0 (print), 3-540-46827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083684/.

Luck:1989:ADI

[Lüc89b] Wolfgang Lück. Algebraically defined invariants. Lecture Notes in Mathematics, 1408:156–304, 1989. CODEN LNMAA2. ISBN 3-540-51846-0 (print), 3-540-46827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083683/.

Luck:1989:BM

[Lüc89c] Wolfgang Lück. Back matter. Lecture Notes in Mathematics, 1408:??, 1989. CODEN LNMAA2. ISBN 3-540-51846-0 (print), 3-540-46827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-46827-1/1.pdf.

Luck:1989:FM

[Lüc89d] Wolfgang Lück. Front matter. Lecture Notes in Mathematics, 1408:??, 1989. CODEN LNMAA2. ISBN 3-540-51846-0 (print), 3-540-46827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46827-1/1.pdf.

Luck:1989:GDI

[Lüc89e] Wolfgang Lück. Geometrically defined invariants. Lecture Notes in Mathematics, 1408:1–155, 1989. CODEN LNMAA2. ISBN 3-540-51846-0 (print), 3-540-46827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083682/.

Luczak:1989:BSV

[Łuc89f] Andrzej Łuczak. Banach space-valued Gleason measures on von Neumann algebras. Lecture Notes in Mathematics, 1391:178– 190, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083391/.

Ludwig:1988:MDI

[Lud88] Jean Ludwig. Minimal C∗-dense ideals and algebraically irreducible representations of the Schwartz-algebra of a nilpotent Lie group. Lecture Notes in Mathematics, 1359:209–217, 1988. CO-DEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086601/.

Lumer:1987:PDM

[Lum87] G. Lumer. Principes du maximum paraboliques pour des domaines (x, t) non-cylindriques. (French) []. Lecture Notes in Mathematics, 1235:105–113, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072751/.

Lumer:1989:EDD

[Lum89] G. Lumer. Équations de diffusion dans des domaines (x, t) noncylindriques et semi-groupes ‘espace-temps’. (French) []. Lecture Notes in Mathematics, 1393:161–180, 1989. CODEN LN-MAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085778/.

Lunardi:1986:RFN

[Lun86a] Alessandra Lunardi. C^{∞} regularity for fully nonlinear abstract evolution equations. Lecture Notes in Mathematics, 1223:176– 185, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099192/.

Lundberg:1986:LIQ

[Lun86b] Lars-Erik Lundberg. Lorentz invariant quantum theory. Lecture Notes in Mathematics, 1218:167–178, 1986. CODEN LN-MAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073049/.

Lusztig:1987:FTS

[Lus87] George Lusztig. Fourier transforms on a semisimple Lie algebra over F_{q}. Lecture Notes in Mathematics, 1271:177–188, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079237/.

Lodkin:1985:AAA

[LV85] A. A. Lodkin and A. M. Vershik. Approximation for actions of amenable groups and transversal automorphisms. Lecture Notes in Mathematics, 1132:331–346, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074893/.

LeBruyn:1986:ED

[LV86a] Lieven Le Bruyn and Michel Van den Bergh. An explicit description of Π_{3,2}. Lecture Notes in Mathematics, 1197:109–113, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076317/.

Leroux:1986:CRS

[LV86b] Pierre Leroux and Gérard X. Viennot. Combinatorial resolution of systems of differential equations, I. Ordinary differential equations. Lecture Notes in Mathematics, 1234:210–245, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072518/.

Li:1989:SFR

[LV89] Huishi Li and Freddy Van Oystaeyen. Strongly filtered rings applied to Gabber’s integrability theorem and modules with regular singularities. Lecture Notes in Mathematics, 1404:296–321, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084081/.

Lee:1988:CSM

[LW88] Ronnie Lee and Steven H. Weintraub. On certain Siegel modular varieties of genus two and levels above two. Lecture Notes in Mathematics, 1361:29–52, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083033/.

Lee:1989:SMV

[LW89] Ronnie Lee and Steven H. Weintraub. The Siegel modular variety of degree two and level four: a report. Lecture Notes in Mathematics, 1399:89–102, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095970/.

Li:1989:TON

[LX89] Bang-He Li and Tao Xu. On third order nondegenerate immersions and maps of S^{1} in R^{2}. Lecture Notes in Mathematics, 1369:164–175, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087532/.

Yu:1987:IPD

[lY87] Yan lin Yu. On the infinitesimal parallel displacement. Lecture Notes in Mathematics, 1255:221–227, 1987. CODEN LN-MAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077693/.

Lychagin:1986:TAG

[Lyc86] V. V. Lychagin. Topological aspects of geometrical theory of differential equations. Lecture Notes in Mathematics, 1214:101– 116, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075961/.

Lafferty:1989:CAL

[LYZ89] John D. Lafferty, Yanlin Yu, and Weiping Zhang. Clifford asymptotics and the local Lefschetz index. Lecture Notes in Mathematics, 1411:137–142, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086448/.

Ma:1989:CSC

[Ma89] Zhisheng Ma. Complete surfaces in H^{3} with a constant principal curvature. Lecture Notes in Mathematics, 1369:176–182, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087533/.

Maagli:1987:RIP

[Maa87] Habib Maagli. Représentation intégrale des potentiels. (French) []. Lecture Notes in Mathematics, 1235:114–119, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072752/.

Maagli:1989:PSS

[Maa89] H. Maagli. Perturbation et subordination des semi-groupes. (French) []. Lecture Notes in Mathematics, 1393:181–195, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085779/.

Macdonald:1987:CDO

[Mac87] I. G. Macdonald. Commuting differential operators and zonal spherical functions. Lecture Notes in Mathematics, 1271:189– 200, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079238/.

MacLane:1988:DSK

[Mac88a] Saunders Mac Lane. Decker’s sharper Künneth formula. Lecture Notes in Mathematics, 1348:242–256, 1988. CODEN LN-MAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081363/.

Mackey:1988:IRA

[Mac88b] George W. Mackey. Induced representations and the applications of harmonic analysis. Lecture Notes in Mathematics, 1359:16– 51, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086586/.

Maddux:1985:FIR

[Mad85] Roger Maddux. Finite integral relation algebras. Lecture Notes in Mathematics, 1149:175–197, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098465/.

Madrecki:1989:LBP

[Mad89] Andrzej Madrecki. On the p-Lévy–Baxter property and its applications. Lecture Notes in Mathematics, 1391:191–199, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083392/.

Maejima:1987:SLT

[Mae87] Makoto Maejima. Some limit theorems for summability methods of I.I.D. random variables. Lecture Notes in Mathematics, 1233: 57–68, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072711/.

Magnusson:1986:SPP

[Mag86] Jón Magnússon. Singularités polaires et points de Weierstrass en codimension plus grande que un. (French) []. Lecture Notes in Mathematics, 1188:260–301, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076827/.

Magnus:1988:AAW

[Mag88] Alphonse P. Magnus. Associated Askey–Wilson polynomials as Laguerre–Hahn orthogonal polynomials. Lecture Notes in Mathematics, 1329:261–278, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083366/.

Mahowald:1987:TCS

[Mah87] Mark Mahowald. Thom complexes and the spectra Bo and Bu. Lecture Notes in Mathematics, 1286:293–297, 1987. CODEN LN-MAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078747/.

Maire:1988:NSC

[Mai88] H.-M. Maire. Necessary and sufficient condition for maximal hypoellipticity of ¯∂_{b}. Lecture Notes in Mathematics, 1324:178– 185, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100791/.

Makkai:1985:UCL

[Mak85] M. Makkai. Ultraproducts and categorical logic. Lecture Notes in Mathematics, 1130:222–309, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075314/.

Makovoz:1988:ANC

[Mak88] Yuly Makovoz. On approximation numbers of the canonical embedding ℓ_{p}^{m} → ℓ_{q}^{m}. Lecture Notes in Mathematics, 1332:145– 149, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081616/.

Malinovsky:1985:LTN

[Mal85a] V. K. Malinovsky. On limit theorems for the number of Markov renewals. Lecture Notes in Mathematics, 1155:190–222, 1985. CO-DEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074820/.

Malitz:1985:PTF

[Mal85b] Jerome I. Malitz. Problems in taxonomy, a floating log. Lecture Notes in Mathematics, 1130:310–316, 1985. CODEN LN-MAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075315/.

Malliavin:1985:SAP

[Mal85c] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil et Marie-Paule Malliavin: Proceedings, Paris 1983–1984 (36ème Année). (French) [], volume 1146 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1146. URL http:

//link.springer.com/book/10.1007/BFb0074534;   http://
www.springerlink.com/content/978-3-540-39628-4.

Malliavin:1986:IPP

[Mal86a] Marie-Paule Malliavin. Ideaux premiers purement codimensionels d’algèbres enveloppantes. (French) []. Lecture Notes in Mathematics, 1197:127–133, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076320/.

Malliavin:1986:SAP

[Mal86b] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil et Marie-Paule Malliavin: Proceedings, Paris 1985 (37ème Année). (French) [], volume 1220 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1220. URL http:

//link.springer.com/book/10.1007/BFb0099500;   http://
www.springerlink.com/content/978-3-540-47255-1.

Malliavin:1988:AHO

[Mal88] M.-P. Malliavin. Algèbre homologique et opérateurs différentiels. (French) []. Lecture Notes in Mathematics, 1328:173–186, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100924/.

Manin:1985:NDG

[Man85] Yu. I. Manin. New dimensions in geometry. Lecture Notes in Mathematics, 1111:59–101, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084585/. See commentary [Ati85a].

Mandel:1986:MIA

[Man86a] Jan Mandel. On multigrid and iterative aggregation methods for nonsymmetric problems. Lecture Notes in Mathematics, 1228:219– 231, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072649/.

Mangasarian:1986:SAP

[Man86b] O. L. Mangasarian. Some applications of penalty functions in mathematical programming. Lecture Notes in Mathematics, 1190: 307–329, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076712/.

Mane:1988:HDI

[Mañ88] Ricardo Mañé. The Hausdorff dimension of invariant probabilities of rational maps. Lecture Notes in Mathematics, 1331:86– 117, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083068/.

Maroni:1985:QED

[Mar85] Pascal Maroni. Sur quelques espaces de distributions qui sont des formes linéaires sur l’espace vectoriel des polynômes. (French) []. Lecture Notes in Mathematics, 1171:184–194, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076543/.

Marciniak:1986:CHI

[Mar86a] Zbigniew Marciniak. Cyclic homology and idempotents in group rings. Lecture Notes in Mathematics, 1217:253–257, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072827/.

Martensen:1986:RMN

[Mar86b] F. Martensen. The Rothe method for nonlinear hyperbolic problems. Lecture Notes in Mathematics, 1192:387–392, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076098/.

Maruyama:1986:SRS

[Mar86c] Masaki Maruyama. Stable rationality of some moduli spaces of vector bundles on P^{2}. Lecture Notes in Mathematics, 1194:80– 89, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076996/.

Martin:1987:AST

[Mar87] Robert H. Martin, Jr. Applications of semigroup theory to reaction-diffusion systems. Lecture Notes in Mathematics, 1248: 108–126, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077420/.

Marcati:1988:NQH

[Mar88a] Pierangelo Marcati. Nonhomogeneous quasilinear hyperbolic systems: Initial and boundary value problem. Lecture Notes in Mathematics, 1340:193–200, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082896/.

Margulis:1988:LGE

[Mar88b] G. A. Margulis. Lie groups and ergodic theory. Lecture Notes in Mathematics, 1352:130–146, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082022/.

Maroni:1988:CFL

[Mar88c] Pascal Maroni. Le calcul des formes linéaires et les polynômes orthogonaux semi-classiques. (French) []. Lecture Notes in Mathematics, 1329:279–290, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083367/.

Markus:1989:MCE

[Már89a] L. Márkus. The mean’s consistent estimation, in the case random processes, satisfying partial differential equations. Lecture Notes in Mathematics, 1412:183–193, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084173/.

Martens:1989:CS

[Mar89b] G. Martens. On curves on K3 surfaces. Lecture Notes in Mathematics, 1389:174–182, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085931/.

Martin:1989:ABS

[Mar89c] Robert H. Martin, Jr. Asymptotic behavior of solutions to quasimonotone parabolic systems. Lecture Notes in Mathematics, 1394: 96–116, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086755/.

Masque:1985:ACP

[Mas85] Jaime Muñoz Masqué. An axiomatic characterization of the Poincaré–Cartan form for second order variational problems. Lecture Notes in Mathematics, 1139:74–84, 1985. CODEN LN-MAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074577/.

Massart:1986:RCC

[Mas86a] Pascal Massart. Rates of convergence in the central limit theorem for empirical processes. Lecture Notes in Mathematics, 1193: 73–109, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077101/.

Masuda:1986:STS

[Mas86b] Mikiya Masuda. Z_{2} surgery theory and smooth involutions on homotopy complex projective spaces. Lecture Notes in Mathematics, 1217:258–289, 1986. CODEN LNMAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072828/.

Masser:1987:SVH

[Mas87] D. W. Masser. Small values of heights on families of Abelian varieties. Lecture Notes in Mathematics, 1290:109–148, 1987. CODEN LNMAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078706/.

Mascolo:1988:ERN

[Mas88a] Elvira Mascolo. Existence results for non convex problems of the calculus of variations. Lecture Notes in Mathematics, 1340:201– 207, 1988. CODEN LNMAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082897/.

Mason:1988:MNM

[Mas88b] J. C. Mason. The L_{p} minimality and near-minimality of orthogonal polynomial approximation and integration methods. Lecture Notes in Mathematics, 1329:291–299, 1988. CODEN LN-MAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083368/.

Masuda:1989:PFC

[Mas89] Mikiya Masuda. A product formula for connected sum. Lecture Notes in Mathematics, 1375:231–239, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085613/.

Matkowski:1985:CFE

[Mat85] Janusz Matkowski. Cauchy functional equation on a restricted domain and commuting functions. Lecture Notes in Mathematics, 1163:101–106, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076422/.

Mathiak:1986:PHS

[Mat86a] Karl Mathiak. Projective Hjelmslev spaces. Lecture Notes in Mathematics, 1175:75–93, 1986. CODEN LNMAA2. ISBN 3-540-16099-X (print), 3-540-39764-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074657/.

Mathiak:1986:VTA

[Mat86b] Karl Mathiak. Valuation topologies and the approximation theorem. Lecture Notes in Mathematics, 1175:30–50, 1986. CODEN LNMAA2. ISBN 3-540-16099-X (print), 3-540-39764-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074655/.

Mathiak:1986:VF

[Mat86c] Karl Mathiak. Valuations of fields. Lecture Notes in Mathematics, 1175:1–29, 1986. CODEN LNMAA2. ISBN 3-540-16099-X (print), 3-540-39764-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074654/.

Mathiak:1986:VSF

[Mat86d] Karl Mathiak. Valuations of Skew Fields and Projective Hjelmslev Spaces, volume 1175 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16099-X (print), 3-540-39764-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 116 pp. LCCN QA3 .L28 no. 1175. URL http:

//link.springer.com/book/10.1007/BFb0074653;   http://
www.springerlink.com/content/978-3-540-39764-9.

Mathiak:1986:VVS

[Mat86e] Karl Mathiak. Valued vector spaces. Lecture Notes in Mathematics, 1175:51–74, 1986. CODEN LNMAA2. ISBN 3-540-16099-X (print), 3-540-39764-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074656/.

Matzat:1987:UGG

[Mat87a] Bernd Heinrich Matzat. Das Umkehrproblem der Galoistheorie. (German) []. Lecture Notes in Mathematics, 1284:1–7, 1987. CO-DEN LNMAA2. ISBN 3-540-18444-9 (print), 3-540-47978-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098325/.

Matzat:1987:EHG

[Mat87b] Bernd Heinrich Matzat. Einbettungsprobleme Über Hilbertkörpern. (German) []. Lecture Notes in Mathematics, 1284:215–268, 1987. CODEN LNMAA2. ISBN 3-540-18444-9 (print), 3-540-47978-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098329/.

Matzat:1987:FGF

[Mat87c] Bernd Heinrich Matzat. Fundamentalgruppen. (German) [Fundamental groups]. Lecture Notes in Mathematics, 1284:9–55, 1987. CODEN LNMAA2. ISBN 3-540-18444-9 (print), 3-540-47978-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098326/.

Matzat:1987:KEG

[Mat87d] Bernd Heinrich Matzat. Klassenzahlen von Erzeugendensystemen. (German) []. Lecture Notes in Mathematics, 1284:57–129, 1987. CODEN LNMAA2. ISBN 3-540-18444-9 (print), 3-540-47978-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098327/.

Matzat:1987:TAG

[Mat87e] Bernd Heinrich Matzat. Topologische Automorphismen. (German) [Topological automorphisms]. Lecture Notes in Mathematics, 1284:131–214, 1987. CODEN LNMAA2. ISBN 3-540-18444-9 (print), 3-540-47978-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098328/.

Matet:1989:CSS

[Mat89a] Pierre Matet. Concerning stationary subsets of [λ]^{<κ}. Lecture Notes in Mathematics, 1401:119–127, 1989. CODEN LN-MAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097335/.

Matsuoka:1989:NPO

[Mat89b] Takashi Matsuoka. The number of periodic orbits of smooth maps. Lecture Notes in Mathematics, 1411:148–154, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086450/.

Mawhin:1986:CPT

[Maw86] J. Mawhin. Critical point theory and nonlinear differential equations. Lecture Notes in Mathematics, 1192:49–58, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076051/.

Mawanda:1988:CAZ

[Maw88] M.-M. Mawanda. On a categorical analysis of Zadeh generalized subsets of sets I. Lecture Notes in Mathematics, 1348:257– 269, 1988. CODEN LNMAA2. ISBN 3-540-50362-5 (print), 3-540-45985-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081364/.

May:1986:EPR

[May86a] J. P. May. Extended powers and H_{∞} ring spectra. Lecture Notes in Mathematics, 1176:1–20, 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075406/.

May:1986:MAS

[May86b] J. P. May. Miscellaneous applications in stable homotopy theory. Lecture Notes in Mathematics, 1176:21–55, 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075407/.

May:1986:PPI

[May86c] J. P. May. Preamble: a polemical introduction to the stable category. Lecture Notes in Mathematics, 1213:1–5, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075779/.

May:1986:RDS

[May86d] J. P. May. A remark on duality and the Segal conjecture. Lecture Notes in Mathematics, 1217:303–305, 1986. CODEN LN-MAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072830/.

Mazya:1986:BIE

[Maz86] V. G. Maz’ya. Boundary integral equations of elasticity in domains with piecewise smooth boundaries. Lecture Notes in Mathematics, 1192:235–242, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076075/.

Moussa:1985:SAP

[MB85] Pierre Moussa and Daniel Bessis. A solvable almost periodic Schrödinger operator. Lecture Notes in Mathematics, 1109:136– 147, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101541/.

McCrea:1987:PGT

[Mc 87] J. Dermott Mc Crea. Poincaré gauge theory of gravitation: Foundations, exact solutions and computer algebra. Lecture Notes in Mathematics, 1251:222–237, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077323/.

Mas-Colell:1988:FLD

[MC88] Andreu Mas-Colell. Four lectures on the differentiable approach to general equilibrium theory. Lecture Notes in Mathematics, 1330: 19–43, 1988. CODEN LNMAA2. ISBN 3-540-50003-0 (print), 3-540-45883-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078157/.

Marques:1989:AST

[MC89] Mauro Marques and Stamatis Cambanis. Admissible and singular translates of stable processes. Lecture Notes in Mathematics, 1391:239–257, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083394/.

McCleary:1985:CGS

[McC85a] John McCleary. Closed geodesics on Stiefel manifolds. Lecture Notes in Mathematics, 1172:157–162, 1985. CODEN LN-MAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074429/.

McConnell:1985:TFR

[McC85b] J. C. McConnell. The K-theory of filtered rings and skew Laurent extensions. Lecture Notes in Mathematics, 1146:288–298, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074542/.

McConnell:1985:SLR

[McC85c] Terry R. McConnell. A Skorohod-like representation in infinite dimensions. Lecture Notes in Mathematics, 1153:359–368, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074960/.

McCulloh:1985:SIM

[McC85d] Leon R. McCulloh. Stickelberger ideals, monoid rings, and Galois module structure. Lecture Notes in Mathematics, 1142:190– 204, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074801/.

McClure:1986:E

[McC86a] J. E. McClure. The modp k-theory of QX. Lecture Notes in Mathematics, 1176:291–377, 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075414/.

McClure:1986:RSS

[McC86b] J. E. McClure. H_{∞} ring spectra via space-level homotopy theory. Lecture Notes in Mathematics, 1176:215–248, 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075412/.

McClure:1986:POR

[McC86c] J. E. McClure. Power operations in H_{∞}^{d} ring theories. Lecture Notes in Mathematics, 1176:249–290, 1986. CODEN LN-MAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075413/.

McCabe:1989:CFA

[McC89] John H. McCabe. On continued fractions associated with polynomial type Padé approximants, with an application. Lecture Notes in Mathematics, 1406:67–75, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096166/.

McFadden:1988:ATP

[McF88] R. B. McFadden. Automated theorem proving applied to the theory of semigroups. Lecture Notes in Mathematics, 1320:235– 243, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083436/.

McGill:1988:ERB

[McG88a] Paul McGill. Excursions from random boundaries. Lecture Notes in Mathematics, 1325:156–161, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077924/.

McGregor:1988:SIE

[McG88b] Malcolm T. McGregor. A solution of the integral equation ∫_{0}^{∞} (sin(\frac{π}{4} + θ)e^{−θy} + sin(\frac{π}{4} − θ)e^{θy})π(x, y) dy = √2coshθx in convolution form and a problem in diffusion theory. Lecture Notes in Mathematics, 1325:162–166, 1988. CODEN LN-MAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077925/.

McNulty:1985:HCF

[McN85] George F. McNulty. How to construct finite algebras which are not finitely based. Lecture Notes in Mathematics, 1149:167–174, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098464/.

Miret:1989:GCC

[MD89] J. M. Miret and S. Xambó Descamps. Geometry of complete cuspidal plane cubics. Lecture Notes in Mathematics, 1389:195– 234, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085933/.

Mdzinarishvili:1987:EH

[Mdz87] L. D. Mdzinarishvili. On exact homology. Lecture Notes in Mathematics, 1283:164–182, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081426/.

Mehring:1985:CIR

[Meh85] Günther H. Mehring. On a criterion of iteration in rings of formal power series. Lecture Notes in Mathematics, 1163:107–114, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076423/.

Meister:1986:SSU

[Mei86a] E. Meister. Some solved and unsolved canonical problems of diffraction theory. Lecture Notes in Mathematics, 1192:393– 398, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076099/.

Meixner:1986:CCS

[Mei86b] Th. Meixner. Constructions of chamber systems of type M with transitive automorphism group. Lecture Notes in Mathematics, 1181:207–217, 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075516/.

Mei:1987:IFP

[Mei87] Xiang-Ming Mei. The integral formula of Pontrjagin characteristic forms. Lecture Notes in Mathematics, 1255:73–79, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077682/.

Meier:1988:ABS

[Mei88] Michael Meier. Asymptotic behavior of solutions of some quasilinear elliptic systems in exterior domains. Lecture Notes in Mathematics, 1357:297–316, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082872/.

Melamed:1989:LTS

[Mel89] J. A. Melamed. Limit theorems in the set up of summation of a random number of independent identically distributed random variables. Lecture Notes in Mathematics, 1412:194–228, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084174/.

Menal:1988:CMR

[Men88a] Pere Menal. Cancellation modules over regular rings. Lecture Notes in Mathematics, 1328:187–208, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100925/.

Menzala:1988:RIS

[Men88b] Gustavo Perla Menzala. On the resonances and the inverse scattering problem for perturbed wave equations. Lecture Notes in Mathematics, 1324:236–244, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100795/.

Mennicke:1989:FGS

[Men89] Jens L. Mennicke. On Fibonacci groups and some other groups. Lecture Notes in Mathematics, 1398:117–123, 1989. CODEN LN-MAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086248/.

Merklen:1986:AAW

[Mer86] Héctor A. Merklen. Artin algebras which are equivalent to a hereditary algebra modulo preprojectives. Lecture Notes in Mathematics, 1177:232–255, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075267/.

Meschiari:1985:CRC

[Mes85] Mauro Meschiari. A classification of real and complex finite dimensional J∗-algebras. Lecture Notes in Mathematics, 1164:1– 84, 1985. CODEN LNMAA2. ISBN 3-540-16048-5 (print), 3-540-39732-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081910/.

Metzger:1985:P

[Met85] Thomas A. Metzger. On the (F) property. Lecture Notes in Mathematics, 1166:95–98, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074698/.

Metivier:1986:SAA

[Mét86] Michel Métivier. On stochastic algorithms in adaptive filtering. Lecture Notes in Mathematics, 1203:134–156, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076879/.

Metivier:1987:WCM

[Mét87] Michel Métivier. Weak convergence of measure valued processes using Sobolev-imbedding techniques. Lecture Notes in Mathematics, 1236:172–183, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072889/.

Metivier:1989:OCO

[Mét89] G. Métivier. Ondes de choc, ondes de raréfaction et ondes soniques multidimensionnelles. (French) []. Lecture Notes in Mathematics, 1402:203–215, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083877/.

Meyer:1985:TLP

[Mey85a] P.-A. Meyer. Sur la théorie de Littlewood–Paley–Stein (d’après Coifman–Rochberg–Weiss et Cowling). (French) []. Lecture Notes in Mathematics, 1123:113–129, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075842/.

Meyer:1985:RTF

[Mey85b] P.-A. Meyer. Une remarque sur la topologie fine. (French) []. Lecture Notes in Mathematics, 1123:176, 1985. CODEN LN-MAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075846.

Meyer:1986:EPQ

[Mey86a] P.-A. Meyer. Éléments de probabilités quantiques. (French) []. Lecture Notes in Mathematics, 1204:186–312, 1986. CODEN LN-MAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075719/.

Meyer:1986:QRA

[Mey86b] P.-A. Meyer. Quelques remarques au sujet du calcul stochastique sur l’espace de Fock. (French) []. Lecture Notes in Mathematics, 1204:321–330, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075722/.

Meyer:1986:EOC

[Mey86c] P.-A. Meyer. Sur l’existence de l’opérateur carré du champ. (French) []. Lecture Notes in Mathematics, 1204:30–33, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075708/.

Meyer:1987:EPQ

[Mey87a] P.-A. Meyer. Éléments de probabilités quantiques. (French) []. Lecture Notes in Mathematics, 1247:33–80, 1987. CODEN LN-MAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077627/.

Meyer:1987:FSP

[Mey87b] P.-A. Meyer. Fock space and probability theory. Lecture Notes in Mathematics, 1250:160–170, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077356/.

Meyer:1988:NSC

[Mey88a] P.-A. Meyer. A note on shifts and cocycles. Lecture Notes in Mathematics, 1303:209–212, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078064/.

Meyer:1988:QHA

[Mey88b] P.-A. Meyer. Quasimartingales hilbertiennes d’après Enchev. (French) []. Lecture Notes in Mathematics, 1321:86–88, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084122/.

Meyer:1988:TBR

[Mey88c] P.-A. Meyer. Sur un théorème de B. Rajeev. (French) []. Lecture Notes in Mathematics, 1321:141–143, 1988. CODEN LN-MAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084129/.

Meyer:1988:SLM

[Mey88d] P.-A. Meyer. Une surmartingale limité de martingales continues. (French) []. Lecture Notes in Mathematics, 1321:138–140, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084128/.

Meyer:1988:RCW

[Mey88e] Paul-André Meyer. Une remarque sur les chaos de Wiener. (French) []. Lecture Notes in Mathematics, 1322:151–155, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077873/.

Meyer:1989:CSE

[Mey89a] P.-A. Meyer. Construction de solutions d’“équations de structure”. (French) []. Lecture Notes in Mathematics, 1372:142–145, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083965/.

Meyer:1989:EPQa

[Mey89b] P.-A. Meyer. Éléments de probabilités quantiques. X. (French) []. Lecture Notes in Mathematics, 1372:175–182, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083971/.

Meyer:1989:EPQb

[Mey89c] P.-A. Meyer. Éléments de probabilités quantiques XI. (French) []. Lecture Notes in Mathematics, 1372:183–185, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083972/.

Meyer:1989:ESM

[Mey89d] P.-A. Meyer. Équations de structure des martingales et probabilités quantiques. (French) []. Lecture Notes in Mathematics, 1372:139–141, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083964/.

Meyer:1989:CRC

[Mey89e] P.-A. Meyer. Un cas de représentation chaotique discrète. (French) []. Lecture Notes in Mathematics, 1372:146, 1989. CODEN LN-MAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083966.

Meziani:1986:PSC

[Mez86] A. Meziani. Perturbations de structures C.R.. (French) []. Lecture Notes in Mathematics, 1198:191–195, 1986. CODEN LN-MAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077053/.

Mingarelli:1988:LVO

[MH88a] Angelo B. Mingarelli and S. Gotskalk Halvorsen. Linear vector ordinary differential equations. Lecture Notes in Mathematics, 1338:49–78, 1988. CODEN LNMAA2. ISBN 3-540-50078-2 (print), 3-540-45918-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080639/.

Mingarelli:1988:NOD

[MH88b] Angelo B. Mingarelli and S. Gotskalk Halvorsen. Non-oscillation domains of differential equations with two parameters. Lecture Notes in Mathematics, 1338:i, 1988. CODEN LNMAA2. ISBN 3-540-50078-2 (print), 3-540-45918-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0120479.

Mingarelli:1988:SLO

[MH88c] Angelo B. Mingarelli and S. Gotskalk Halvorsen. Scalar linear ordinary differential equations. Lecture Notes in Mathematics, 1338:

1–48, 1988. CODEN LNMAA2. ISBN 3-540-50078-2 (print), 3-540-45918-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080638/.

Mingarelli:1988:SVS

[MH88d] Angelo B. Mingarelli and S. Gotskalk Halvorsen. Scalar Volterra– Stieltjes integral equations. Lecture Notes in Mathematics, 1338: 79–100, 1988. CODEN LNMAA2. ISBN 3-540-50078-2 (print), 3-540-45918-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080640/.

Mhaskar:1987:SDT

[Mha87] H. N. Mhaskar. Some discrepancy theorems. Lecture Notes in Mathematics, 1287:117–131, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078901/.

Michler:1986:BCC

[Mic86] Gerhard O. Michler. Brauer’s conjectures and the classification of finite simple groups. Lecture Notes in Mathematics, 1178:129– 142, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075293/.

Michor:1989:FM

[Mic89a] Peter W. Michor. Front matter. Lecture Notes in Mathematics, 1411:??, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46862-2/1.pdf.

Michor:1989:GDA

[Mic89b] Peter W. Michor. Graded derivations of the algebra of differential forms associated with a connection. Lecture Notes in Mathematics, 1410:249–261, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086427/.

Milgram:1985:CSE

[Mil85a] R. James Milgram. The Cappell–Shaneson example. Lecture Notes in Mathematics, 1126:159–165, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074442/.

Milgram:1985:ESF

[Mil85b] R. James Milgram. Evaluating the Swan finiteness obstruction for periodic groups. Lecture Notes in Mathematics, 1126:127– 158, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074441/.

Millett:1985:TSF

[Mil85c] Kenneth Millett. Transversality for surfaces in foliated 3manifolds. Lecture Notes in Mathematics, 1144:115–133, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075015/.

Milman:1985:RSP

[Mil85d] V. D. Milman. Random subspaces of proportional dimension of finite dimensional normed spaces: Approach through the isoperimetric inequality. Lecture Notes in Mathematics, 1166:106– 115, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074700/.

Milman:1985:VAI

[Mil85e] V. D. Milman. Volume approach and iteration procedures in local theory of normed spaces. Lecture Notes in Mathematics, 1166: 99–105, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074699/.

Milman:1987:DMI

[Mil87a] V. D. Milman. Diameter of a minimal invariant subset of equivariant Lipschitz actions on compact subsets of R^{k}. Lecture Notes in Mathematics, 1267:13–20, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078133/.

Milman:1987:SRU

[Mil87b] V. D. Milman. Some remarks on Urysohn’s inequality and volume ratio of cotype 2-spaces. Lecture Notes in Mathematics, 1267: 75–81, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078137/.

Miletta:1988:ABS

[Mil88a] Peter Miletta. Asymptotic behavior of solutions to abstract evolution equations. Lecture Notes in Mathematics, 1331:118–139, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083069/.

Milgram:1988:SRK

[Mil88b] R. J. Milgram. Some remarks on the Kirby–Siebenmann class. Lecture Notes in Mathematics, 1361:247–252, 1988. CODEN LN-MAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083037/.

Milman:1988:FOC

[Mil88c] V. D. Milman. A few observations on the connections between local theory and some other fields. Lecture Notes in Mathematics, 1317:283–289, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081748/.

Milman:1988:ISG

[Mil88d] V. D. Milman. Isomorphic symmetrization and geometric inequalities. Lecture Notes in Mathematics, 1317:107–131, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081738/.

Minda:1987:IHM

[Min87] David Minda. Inequalities for the hyperbolic metric and applications to geometric function theory. Lecture Notes in Mathematics, 1275:235–252, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078356/.

Mira:1985:RSB

[Mir85] Christian Mira. Rotation sequences and bifurcations structure of one-dimensional endomorphisms. Lecture Notes in Mathematics, 1163:115–124, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076424/.

Misiurewicz:1985:CAE

[Mis85] Michał Misiurewicz. Chaos almost everywhere. Lecture Notes in Mathematics, 1163:125–130, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076425/.

Mitter:1986:ETS

[Mit86] S. K. Mitter. Estimation theory and statistical physics. Lecture Notes in Mathematics, 1203:157–176, 1986. CODEN LN-MAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076880/.

Mitchell:1987:BFL

[Mit87] S. A. Mitchell. The Bott filtration of a loop group. Lecture Notes in Mathematics, 1298:215–226, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083012/.

Mizohata:1988:CPH

[Miz88] Sigeru Mizohata. On the Cauchy problem for hyperbolic equations in C^{∞} and Gevrey classes. Lecture Notes in Mathematics, 1324:197–215, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100793/.

Malliavin:1988:IPC

[MK88] Paul Malliavin and Y. Katznelson. Image des points critiques d’une application regulière. (French) []. Lecture Notes in Mathematics, 1322:85–92, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077868/.

Marcellan:1985:MRF

[MM85] F. Marcellán and L. Moral. Minimal recurrence formulas for orthogonal polynomials on Bernoulli’s lemniscate. Lecture Notes in Mathematics, 1171:211–220, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076546/.

Mazziotto:1986:PLS

[MM86] G. Mazziotto and A. Millet. Points, lignes et systèmes d’arrêt flous et problème d’arrêt optimal. (French) []. Lecture Notes in Mathematics, 1204:81–94, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075714/.

Maslov:1986:MED

[MMA86] V. P. Maslov, P. P. Mosolov, and M. M. Antsiferova. Model equation for dynamics of phase translation. Lecture Notes in Mathematics, 1214:231–248, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075968/.

Meise:1987:SSD

[MMT87] Reinhold Meise, Siegfried Momm, and B. Alan Taylor. Splitting of slowly decreasing ideals in weighted algebras of entire functions. Lecture Notes in Mathematics, 1276:229–252, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078961/.

Mate:1988:FSO

[MN88a] Attila Máté and Paul Nevai. Factorization of second order difference equations and its application to orthogonal polynomials. Lecture Notes in Mathematics, 1329:158–177, 1988. CODEN LN-MAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083357/.

McCoy:1988:CF

[MN88b] Robert A. McCoy and Ibula Ntantu. Cardinal functions. Lecture Notes in Mathematics, 1315:51–73, 1988. CODEN LN-MAA2. ISBN 3-540-19302-2 (print), 3-540-39181-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098394/.

McCoy:1988:COP

[MN88c] Robert A. McCoy and Ibula Ntantu. Completeness and other properties. Lecture Notes in Mathematics, 1315:74–105, 1988. CODEN LNMAA2. ISBN 3-540-19302-2 (print), 3-540-39181-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098395/.

McCoy:1988:CCS

[MN88d] Robert A. McCoy and Ibula Ntantu. Convergence and compact subsets. Lecture Notes in Mathematics, 1315:39–50, 1988. CODEN LNMAA2. ISBN 3-540-19302-2 (print), 3-540-39181-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098393/.

McCoy:1988:FST

[MN88e] Robert A. McCoy and Ibula Ntantu. Function space topologies. Lecture Notes in Mathematics, 1315:3–14, 1988. CODEN LN-MAA2. ISBN 3-540-19302-2 (print), 3-540-39181-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098391/.

McCoy:1988:I

[MN88f] Robert A. McCoy and Ibula Ntantu. Introduction. Lecture Notes in Mathematics, 1315:1, 1988. CODEN LNMAA2. ISBN 3-540-19302-2 (print), 3-540-39181-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098390.

McCoy:1988:NF

[MN88g] Robert A. McCoy and Ibula Ntantu. Natural functions. Lecture Notes in Mathematics, 1315:15–38, 1988. CODEN LN-MAA2. ISBN 3-540-19302-2 (print), 3-540-39181-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098392/.

Madlener:1988:GHF

[MO88] Klaus Madlener and Friedrich Otto. On groups having finite monadic Church–Rosser presentations. Lecture Notes in Mathematics, 1320:218–234, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083435/.

Modica:1986:SHE

[Mod86] L. Modica. Stochastic homogenization and ergodic theory. Lecture Notes in Mathematics, 1190:359–370, 1986. CODEN LN-MAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076714/.

Modugno:1987:SVV

[Mod87] Marco Modugno. Systems of vector valued forms on a fibred manifold and applications to gauge theories. Lecture Notes in Mathematics, 1251:238–264, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077324/.

Mohres:1987:TFN

[Möh87] Walter Möhres. Torsion-free nilpotent groups with bounded ranks of the Abelian subgroups. Lecture Notes in Mathematics, 1281: 115–117, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078696/.

Mok:1986:CKM

[Mok86] Ngaiming Mok. Compact Kähler manifolds of nonnegative holomorphic bisectional curvature. Lecture Notes in Mathematics, 1194:90–103, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076997/.

Montejano:1987:GCL

[Mon87] Luis Montejano. Geometric category and Lusternik–Schnirelmann category. Lecture Notes in Mathematics, 1283:183–192, 1987. CO-DEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081427/.

Moroz:1986:AAAb

[Mor86a] B. Z. Moroz. Analytic Arithmetic in Algebraic Number Fields, volume 1205 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16784-6 (print), 3-540-39996-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 177 pp. LCCN QA3 .L28 no. 1205. URL http:

//link.springer.com/book/10.1007/BFb0076403;   http://
www.springerlink.com/content/978-3-540-39996-4.

Moroz:1986:AAAa

[Mor86b] B. Z. Moroz. Analytical arithmetic in algebraic number fields. Lecture Notes in Mathematics, 1205:e1–e3, 1986. CODEN LN-MAA2. ISBN 3-540-16784-6 (print), 3-540-39996-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0120465/.

Moroz:1986:CB

[Mor86c] B. Z. Moroz. Classical background. Lecture Notes in Mathematics, 1205:1–77, 1986. CODEN LNMAA2. ISBN 3-540-16784-6 (print), 3-540-39996-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076404/.

Moroz:1986:IEN

[Mor86d] B. Z. Moroz. Ideals with equal norms and integral points on nonform varieties. Lecture Notes in Mathematics, 1205:141–167, 1986. CODEN LNMAA2. ISBN 3-540-16784-6 (print), 3-540-39996-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076406/.

Moroz:1986:RC

[Mor86e] B. Z. Moroz. Remarks and comments. Lecture Notes in Mathematics, 1205:168–170, 1986. CODEN LNMAA2. ISBN 3-540-16784-6 (print), 3-540-39996-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076407/.

Moroz:1986:SPF

[Mor86f] B. Z. Moroz. Scalar product of L-functions. Lecture Notes in Mathematics, 1205:78–140, 1986. CODEN LNMAA2. ISBN 3-540-16784-6 (print), 3-540-39996-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076405/.

Morava:1988:SWG

[Mor88] Jack Morava. Some Weil group representations motivated by algebraic topology. Lecture Notes in Mathematics, 1326:94–106, 1988. CODEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078040/.

Morgan:1989:GES

[Mor89] Jeff Morgan. Global existence for semilinear parabolic systems via Lyapunov type methods. Lecture Notes in Mathematics, 1394: 117–121, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086756/.

Mosco:1985:SSO

[Mos85] U. Mosco. On some stochastic optimal impulse control problems. Lecture Notes in Mathematics, 1119:139–151, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074784/.

Mosco:1988:WCV

[Mos88] Umberto Mosco. Wiener criteria and variational convergences. Lecture Notes in Mathematics, 1340:208–238, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082898/.

Moser:1989:MFT

[Mos89] Jürgen Moser. Minimal foliations on a torus. Lecture Notes in Mathematics, 1365:62–99, 1989. CODEN LNMAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089178/.

Moussaoui:1985:ASD

[Mou85] M.-A. Moussaoui. Sur l’approximation des solutions du problème de Dirichlet dans un ouvert avec coins. (French) []. Lecture Notes in Mathematics, 1121:199–206, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076272/.

Moussu:1988:DAA

[Mou88] R. Moussu. Développement asymptotique de l’application retour d’un polycycle. (French) []. Lecture Notes in Mathematics, 1331: 140–149, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083070/.

Marcus:1985:SPS

[MP85] Michael B. Marcus and Gilles Pisier. Stochastic processes with sample paths in exponential Orlicz spaces. Lecture Notes in Mathematics, 1153:329–358, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074959/.

Marino:1986:GGN

[MP86] V. Marino and A. Pràstaro. On a geometric generalization of the Noether theorem. Lecture Notes in Mathematics, 1209:222– 234, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076634/.

Milani:1988:DTT

[MP88] A. Milani and R. Picard. Decomposition theorems and their application to non-linear electro- and magneto-static boundary value problems. Lecture Notes in Mathematics, 1357:317–340, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082873/.

Meyer:1989:SI

[MP89a] Mathieu Meyer and Alain Pajor. On Santaló’s inequality. Lecture Notes in Mathematics, 1376:261–263, 1989. CODEN LN-MAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090059/.

Milman:1989:IPI

[MP89b] V. D. Milman and A. Pajor. Isotropic position and inertia ellipsoids and zonoids of the unit ball of a normed n-dimensional space. Lecture Notes in Mathematics, 1376:64–104, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090049/.

Micali:1986:GEC

[MPS86] Artibano Micali, Antonio Paques, and Andrzej Solecki. Sur le groupe des extensions cubiques. (French) []. Lecture Notes in Mathematics, 1197:134–148, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076321/.

Mallet-Paret:1987:PSA

[MPS87] John Mallet-Paret and George R. Sell. The principle of spatial averaging and inertial manifolds for reaction-diffusion equations. Lecture Notes in Mathematics, 1248:94–107, 1987. CODEN LN-MAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077419/.

Mason:1985:BEM

[MPST85] J. C. Mason, A. P. Parker, R. N. L. Smith, and R. M. Thompson. Boundary element methods for singularity problems in fracture mechanics. Lecture Notes in Mathematics, 1121:181–198, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076271/.

Madsen:1985:SLL

[MR85a] Ib Madsen and Martin Raussen. Smooth and locally linear g homotopy representations. Lecture Notes in Mathematics, 1172: 130–156, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074428/.

Menaldi:1985:SCP

[MR85b] José-Luis Menaldi and Maurice Robin. Some control problems of degenerate diffusions with unbounded cost. Lecture Notes in Mathematics, 1119:113–138, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074783/.

Micchelli:1985:LOR

[MR85c] C. A. Micchelli and T. J. Rivlin. Lectures on optimal recovery. Lecture Notes in Mathematics, 1129:21–93, 1985. CODEN LNMAA2. ISBN 3-540-15234-2 (print), 3-540-39412-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075157/.

Min-Oo:1986:CD

[MR86] Min-Oo Maung and Ernst A. Ruh. Curvature deformations. Lecture Notes in Mathematics, 1201:180–190, 1986. CODEN LN-MAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075655/.

Manevitz:1987:FFU

[MR87] Larry M. Manevitz and Louis H. Rowen. Freedom via forcing: Uniform construction of relatively free or generic structures. Lecture Notes in Mathematics, 1292:230–246, 1987. CODEN LN-MAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082240/.

Madsen:1989:EPI

[MR89a] I. Madsen and M. Rothenberg. Equivariant pseudo-isotopies and K_{−I}. Lecture Notes in Mathematics, 1375:216–230, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085612/.

Metivier:1989:ITP

[MR89b] Guy Métivier and Jeffrey Rauch. The interaction of two progressing waves. Lecture Notes in Mathematics, 1402:216–226, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083878/.

Miro-Roig:1989:UAB

[MR89c] Rosa M. Miró-Roig. Unobstructed arithmetically Buchsbaum curves. Lecture Notes in Mathematics, 1389:235–241, 1989. CO-DEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085934/.

Meleard:1989:GEC

[MRC89] Sylvie Méléard and Sylvie Roelly-Coppoletta. A generalized equation for a continuous measure branching process. Lecture Notes in Mathematics, 1390:171–185, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083944/.

McGill:1988:ELC

[MRR88] P. McGill, B. Rajeev, and B. V. Rao. Extending Lévy’s characterisation of Brownian motion. Lecture Notes in Mathematics, 1321:163–165, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084133/.

Morgan:1985:ICS

[MS85] John W. Morgan and Peter B. Shalen. An introduction to compactifying spaces of hyperbolic structures by actions on trees. Lecture Notes in Mathematics, 1167:228–240, 1985. CODEN LN-MAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075226/.

Matumoto:1986:PST

[MS86a] Takao Matumoto and Masahiro Shiota. Proper subanalytic transformation groups and unique triangulation of the orbit spaces. Lecture Notes in Mathematics, 1217:290–302, 1986. CODEN LN-MAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072829/.

May:1986:HAE

[MS86b] J. P. May and M. Steinberger. The homological analysis of extended powers. Lecture Notes in Mathematics, 1213:384–406, 1986. CODEN LNMAA2. ISBN 3-540-16820-6 (print), 3-540-47077-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075787/.

Milman:1986:AAL

[MS86c] Vitali D. Milman and Gideon Schechtman. Additional applications of Lévy families in the theory of finite dimensional normed spaces. Lecture Notes in Mathematics, 1200:60–68, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38822-7_10/.

Milman:1986:AESb

[MS86d] Vitali D. Milman and Gideon Schechtman. Almost Euclidean subspaces of ℓ_{p}^{n} spaces, of general n-dimensional normed spaces, and of quotient of n-dimensional spaces. Lecture Notes in Mathematics, 1200:19–26, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38822-7_5/.

Milman:1986:AESa

[MS86e] Vitali D. Milman and Gideon Schechtman. Almost Euclidean subspaces of a normed space. Lecture Notes in Mathematics, 1200: 12–18, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38822-7_4/.

Milman:1986:ATF

[MS86f] Vitali D. Milman and Gideon Schechtman. Asymptotic Theory of Finite Dimensional Normed Spaces, volume 1200 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 156 pp. LCCN

QA3 .L28 no. 1200. URL http://link.springer.com/book/10.
1007/978-3-540-38822-7;   http://www.springerlink.com/
content/978-3-540-38822-7.

Milman:1986:BM

[MS86g] Vitali D. Milman and Gideon Schechtman. Back matter. Lecture Notes in Mathematics, 1200:114–156, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-540-38822-7/1.pdf.

Milman:1986:E

[MS86h] Vitali D. Milman and Gideon Schechtman. Embedding ℓ_{p}^{m} into ℓ_{1}^{n}. Lecture Notes in Mathematics, 1200:42–50, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38822-7_8/.

Milman:1986:FDN

[MS86i] Vitali D. Milman and Gideon Schechtman. Finite dimensional normed spaces, preliminaries. Lecture Notes in Mathematics, 1200: 9–11, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38822-7_3/.

Milman:1986:FM

[MS86j] Vitali D. Milman and Gideon Schechtman. Front matter. Lecture Notes in Mathematics, 1200:I–X, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-38822-7/1.pdf.

Milman:1986:IIS

[MS86k] Vitali D. Milman and Gideon Schechtman. The isoperimetric inequality on S^{n−1} and some consequences. Lecture Notes in Mathematics, 1200:5–8, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38822-7_2/.

Milman:1986:KT

[MS86l] Vitali D. Milman and Gideon Schechtman. Krivine’s theorem. Lecture Notes in Mathematics, 1200:77–84, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38822-7_12/.

Milman:1986:LF

[MS86m] Vitali D. Milman and Gideon Schechtman. Lévy families. Lecture Notes in Mathematics, 1200:27–32, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38822-7_6/.

Milman:1986:M

[MS86n] Vitali D. Milman and Gideon Schechtman. Martingales. Lecture Notes in Mathematics, 1200:33–41, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38822-7_7/.

Milman:1986:MPT

[MS86o] Vitali D. Milman and Gideon Schechtman. The Maurey– Pisier Theorem. Lecture Notes in Mathematics, 1200:85–97, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38822-7_13/.

Milman:1986:P

[MS86p] Vitali D. Milman and Gideon Schechtman. Preliminaries. Lecture Notes in Mathematics, 1200:1–4, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38822-7_1/.

Milman:1986:PRE

[MS86q] Vitali D. Milman and Gideon Schechtman. Projections on random Euclidean subspaces of finite dimensional normed spaces. Lecture Notes in Mathematics, 1200:106–113, 1986. CODEN LN-MAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38822-7_15/.

Milman:1986:RP

[MS86r] Vitali D. Milman and Gideon Schechtman. The Rademacher projection. Lecture Notes in Mathematics, 1200:98–105, 1986. CO-DEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-540-38822-7_
14/.

Milman:1986:RTS

[MS86s] Vitali D. Milman and Gideon Schechtman. Ramsey’s theorem with some applications to normed spaces. Lecture Notes in Mathematics, 1200:69–76, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38822-7_11/.

Milman:1986:TCN

[MS86t] Vitali D. Milman and Gideon Schechtman. Type and cotype of normed spaces, and some simple relations with geometrical properties. Lecture Notes in Mathematics, 1200:51– 59, 1986. CODEN LNMAA2. ISBN 3-540-16769-2 (print), 3-540-38822-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38822-7_9/.

Morison:1986:OMP

[MS86u] K. R. Morison and R. W. H. Sargent. Optimization of multistage processes described by differential-algebraic equations. Lecture Notes in Mathematics, 1230:86–102, 1986. CODEN LN-MAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072673/.

Meril:1987:CSH

[MS87a] A. Meril and D. C. Struppa. Convolutors in spaces of holomorphic functions. Lecture Notes in Mathematics, 1276:253–275, 1987. CO-DEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078962/. See re-

mark [BS87c].

Meril:1987:PHE

[MS87b] A. Méril and D. C. Struppa. Phénomène de Hartogs et équations de convolution. (French) []. Lecture Notes in Mathematics, 1295: 146–156, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081981/.

Makagon:1989:NID

[MS89a] A. Makagon and H. Salehi. Notes on infinite dimensional stationary sequences. Lecture Notes in Mathematics, 1391:200–238, 1989. CODEN LNMAA2. ISBN 3-540-51548-8 (print), 3-540-48244-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083393/.

Mezzetti:1989:GHS

[MS89b] Emilia Mezzetti and Gianni Sacchiero. Gonality and Hilbert schemes of smooth curves. Lecture Notes in Mathematics, 1389: 183–194, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085932/.

Muller:1989:CSV

[MS89c] Paul F. X. Müller and Gideon Schechtman. On complemented subspaces of H^{1} and VMO. Lecture Notes in Mathematics, 1376: 113–125, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090051/.

Mahowald:1987:CIE

[MT87] Mark Mahowald and Robert Thompson. A commentary on the “Image of J in the EHP sequence”. Lecture Notes in Mathematics, 1286:298–304, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078748/.

Milman:1987:STI

[MTJ87] V. D. Milman and N. Tomczak-Jaegermann. Sudakov type inequalities for convex bodies in IR^{n}. Lecture Notes in Mathematics, 1267:113–121, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078140/.

Mui:1985:HOD

[Mùi85] Huỳnh Mùi. Homology operations derived from modular coinvariants. Lecture Notes in Mathematics, 1172:85–115, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074426/.

Mukherjea:1989:CPN

[Muk89] Arunava Mukherjea. Convolution products of non-identical distributions on a compact Abelian semigroup. Lecture Notes in Mathematics, 1379:217–241, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087856/.

Muldoon:1986:ZSS

[Mul86] M. E. Muldoon. On the zeros of some special functions: Differential equations and Nicholson-type formulas. Lecture Notes in Mathematics, 1192:155–160, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076063/.

Muller:1987:IGD

[Mül87a] Werner Müller. The L^{2}-index of generalized Dirac operators. Lecture Notes in Mathematics, 1244:84–99, 1987. CODEN LN-MAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077670/.

Muller:1987:CRO

[Mül87b] Werner Müller. Cusps of rank one. Lecture Notes in Mathematics, 1244:5–12, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077662/.

Muller:1987:EF

[Mül87c] Werner Müller. The Eisenstein functions. Lecture Notes in Mathematics, 1244:64–73, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077668/.

Muller:1987:HEC

[Mül87d] Werner Müller. The heat equation on the cusp. Lecture Notes in Mathematics, 1244:13–21, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077663/.

Muller:1987:HK

[Mül87e] Werner Müller. The heat kernel. Lecture Notes in Mathematics, 1244:60–63, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077667/.

Muller:1987:HC

[Mül87f] Werner Müller. The Hirzebruch conjecture. Lecture Notes in Mathematics, 1244:139–149, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077672/.

Muller:1987:MCRa

[Mül87g] Werner Müller. Manifolds with cusps of rank one. Lecture Notes in Mathematics, 1244:31–45, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077665/.

Muller:1987:MCRb

[Mül87h] Werner Müller. Manifolds with Cusps of Rank One: Spectral Theory and L^{2}-Index Theorem, volume 1244 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 158 pp. LCCN QA3

.L28 no. 1244. URL http://link.springer.com/book/10.1007/
BFb0077660;   http://www.springerlink.com/content/978-
3-540-47762-4.

Muller:1987:NLC

[Mül87i] Werner Müller. The Neumann Laplacian on the cusp. Lecture Notes in Mathematics, 1244:22–30, 1987. CODEN LN-MAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077664/.

Muller:1987:P

[Mül87j] Werner Müller. Preliminaries. Lecture Notes in Mathematics, 1244:1–4, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077661/.

Muller:1987:SRH

[Mül87k] Werner Müller. The spectral resolution of H. Lecture Notes in Mathematics, 1244:46–59, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077666/.

Muller:1987:SSF

[Mül87l] Werner Müller. The spectral shift function. Lecture Notes in Mathematics, 1244:74–83, 1987. CODEN LNMAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077669/.

Muller:1987:UCI

[Mül87m] Werner Müller. The unipotent contribution to the index. Lecture Notes in Mathematics, 1244:100–138, 1987. CODEN LN-MAA2. ISBN 3-540-17696-9 (print), 3-540-47762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077671/.

Muller:1988:IR

[Mül88] Werner Müller. L^{2}-index and resonances. Lecture Notes in Mathematics, 1339:203–211, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083056/.

Muller:1989:RET

[Mül89] Ingo Müller. Relativistic extended thermodynamics II. Lecture Notes in Mathematics, 1385:257–268, 1989. CODEN LN-MAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084033/.

Mumford:1988:SR

[Mum88a] David Mumford. Spec(R). Lecture Notes in Mathematics, 1358: 93–108, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21581-4_11/.

Mumford:1988:CP

[Mum88b] David Mumford. The category of preschemes. Lecture Notes in Mathematics, 1358:108–121, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_12/.

Mumford:1988:CS

[Mum88c] David Mumford. Closed subpreschemes. Lecture Notes in Mathematics, 1358:143–155, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_15/.

Mumford:1988:CM

[Mum88d] David Mumford. Coherent modules. Lecture Notes in Mathematics, 1358:205–215, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_20/.

Mumford:1988:CVa

[Mum88e] David Mumford. Complete varieties. Lecture Notes in Mathematics, 1358:75–80, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_9/.

Mumford:1988:CVb

[Mum88f] David Mumford. Complex varieties. Lecture Notes in Mathematics, 1358:80–89, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_10/.

Mumford:1988:DM

[Mum88g] David Mumford. Definition of a morphism: I. Lecture Notes in Mathematics, 1358:15–24, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_3/.

Mumford:1988:DPM

[Mum88h] David Mumford. Definition of prevarieties and morphism. Lecture Notes in Mathematics, 1358:35–45, 1988. CODEN LN-MAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21581-4_5/.

Mumford:1988:D

[Mum88i] David Mumford. Dimension. Lecture Notes in Mathematics, 1358: 56–67, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21581-4_7/.

Mumford:1988:FMb

[Mum88j] David Mumford. The fibres of a morphism. Lecture Notes in Mathematics, 1358:67–75, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_8/.

Mumford:1988:FD

[Mum88k] David Mumford. Fields of definition. Lecture Notes in Mathematics, 1358:131–142, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_14/.

Mumford:1988:FMa

[Mum88l] David Mumford. Front matter. Lecture Notes in Mathematics, 1358:1, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-21581-4/1/1.pdf.

Mumford:1988:FMc

[Mum88m] David Mumford. Front matter. Lecture Notes in Mathematics, 1358:91–92, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21581-4/2/1.pdf.

Mumford:1988:FMd

[Mum88n] David Mumford. Front matter. Lecture Notes in Mathematics, 1358:191, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21581-4/3/1.pdf.

Mumford:1988:FMe

[Mum88o] David Mumford. Front matter. Lecture Notes in Mathematics, 1358:N2–V, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21581-4/1.pdf.

Mumford:1988:FPP

[Mum88p] David Mumford. The functor of points of a prescheme. Lecture Notes in Mathematics, 1358:155–167, 1988. CODEN LN-MAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21581-4_16/.

Mumford:1988:IAS

[Mum88q] David Mumford. Irreducible algebraic sets. Lecture Notes in Mathematics, 1358:7–15, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_2/.

Mumford:1988:PHA

[Mum88r] David Mumford. Products and the Hausdorff axiom. Lecture Notes in Mathematics, 1358:46–55, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_6/.

Mumford:1988:PMF

[Mum88s] David Mumford. Proper morphisms and finite morphisms. Lecture Notes in Mathematics, 1358:168–176, 1988. CODEN LN-MAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21581-4_17/.

Mumford:1988:QCM

[Mum88t] David Mumford. Quasi-coherent modules. Lecture Notes in Mathematics, 1358:193–205, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_19/.

Mumford:1988:SAV

[Mum88u] David Mumford. Sheaves and affine varieties. Lecture Notes in Mathematics, 1358:24–35, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_4/.

Mumford:1988:SA

[Mum88v] David Mumford. Some algebra. Lecture Notes in Mathematics, 1358:2–7, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_1/.

Mumford:1988:S

[Mum88w] David Mumford. Specialization. Lecture Notes in Mathematics, 1358:177–189, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21581-4_18/.

Mumford:1988:VP

[Mum88x] David Mumford. Varieties are preschemes. Lecture Notes in Mathematics, 1358:121–131, 1988. CODEN LNMAA2. ISBN 3-540-50497-4 (print), 3-662-21581-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21581-4_13/.

Murasugi:1985:HL

[Mur85a] Kunio Murasugi. 2-heights of links. Lecture Notes in Mathematics, 1144:134–137, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075016/.

Murre:1985:AAT

[Mur85b] J. P. Murre. Applications of algebraic K-theory to the theory of algebraic cycles. Lecture Notes in Mathematics, 1124:216–261, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075002/.

Murty:1985:FCM

[Mur85c] V. Kumar Murty. Fourier coefficients of modular forms. Lecture Notes in Mathematics, 1122:163–172, 1985. CODEN LN-MAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075760/.

Muramatu:1987:ENP

[Mur87] Tosinobu Muramatu. Estimates for the norm of pseudo-differential operators by means of Besov spaces. Lecture Notes in Mathematics, 1256:330–349, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077749/.

Murai:1988:ACC

[Mur88a] Takafumi Murai. Analytic capacities of cranks. Lecture Notes in Mathematics, 1307:71–116, 1988. CODEN LNMAA2. ISBN 3-540-19091-0 (print), 3-540-39105-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078081/.

Murai:1988:CCP

[Mur88b] Takafumi Murai. The Calderón commutator (8 proofs of its boundedness). Lecture Notes in Mathematics, 1307:1–30, 1988. CODEN LNMAA2. ISBN 3-540-19091-0 (print), 3-540-39105-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078079/.

Murai:1988:RVM

[Mur88c] Takafumi Murai. A real variable method for the Cauchy transform on graphs. Lecture Notes in Mathematics, 1307:31–70, 1988. CODEN LNMAA2. ISBN 3-540-19091-0 (print), 3-540-39105-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078080/.

Murakami:1989:ESL

[Mur89] Shingo Murakami. Exceptional simple Lie groups and related topics in recent differential geometry. Lecture Notes in Mathematics, 1369:183–221, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087534/.

Muto:1986:FEL

[Mut86] Hideo Mutō. The first eigenvalue of the Laplacian of an isoparametric minimal hypersurface in a unit sphere. Lecture Notes in Mathematics, 1201:191–201, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075656/.

McBryan:1986:MMP

[MV86] Oliver A. McBryan and Eric F. Van de Velde. The multigrid method on parallel processors. Lecture Notes in Mathematics, 1228:232–260, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072650/.

Metivier:1988:WSS

[MV88] Michel Métivier and Michel Viot. On weak solutions of stochastic partial differential equations. Lecture Notes in Mathematics, 1322:139–150, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077872/.

Mortini:1988:SRP

[MvR88] Raymond Mortini and Michael von Renteln. Some remarkable properties of the Wiener algebra W^{+}. Lecture Notes in Mathematics, 1359:218–225, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086602/.

Moeglin:1987:CHI

[MVW87a] Colette Mœglin, Marie-France Vignéras, and Jean-Loup Waldspurger. Correspondance de Howe et induction. (French) []. Lecture Notes in Mathematics, 1291:51–77, 1987. CODEN LN-MAA2. ISBN 3-540-18699-9 (print), 3-540-48102-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082715/.

Moeglin:1987:EHF

[MVW87b] Colette Mœglin, Marie-France Vignéras, and Jean-Loup Waldspurger. Espaces hermitiens. (French) []. Lecture Notes in Mathematics, 1291:1–25, 1987. CODEN LNMAA2. ISBN 3-540-18699-9 (print), 3-540-48102-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082713/.

Moeglin:1987:PRD

[MVW87c] Colette Mœglin, Marie-France Vignéras, and Jean-Loup Waldspurger. Paires réductives duales non ramifiées. (French) []. Lecture Notes in Mathematics, 1291:99–125, 1987. CODEN LN-MAA2. ISBN 3-540-18699-9 (print), 3-540-48102-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082717/.

Moeglin:1987:RPR

[MVW87d] Colette Mœglin, Marie-France Vignéras, and Jean-Loup Waldspurger. Représentations de petit rang du groupe symplectique. (French) []. Lecture Notes in Mathematics, 1291:127–161, 1987. CODEN LNMAA2. ISBN 3-540-18699-9 (print), 3-540-48102-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082718/.

Moeglin:1987:RMC

[MVW87e] Colette Mœglin, Marie-France Vignéras, and Jean-Loup Waldspurger. Représentations métaplectiques et conjecture de Howe. (French) []. Lecture Notes in Mathematics, 1291:27–50, 1987. CODEN LNMAA2. ISBN 3-540-18699-9 (print), 3-540-48102-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082714/.

Moeglin:1987:CCC

[MVW87f] Colette Mœglin, Marie-France Vignéras, and Jean-Loup Waldspurger. Sur les classes de conjugaison dans certains groupes unitaires. (French) []. Lecture Notes in Mathematics, 1291:79– 97, 1987. CODEN LNMAA2. ISBN 3-540-18699-9 (print), 3-540-48102-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082716/.

Miele:1986:DPS

[MW86] A. Miele and T. Wang. Dual properties of sequential gradient — restoration algorithms for optimal control problems. Lecture Notes in Mathematics, 1190:331–357, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076713/.

Masser:1989:SEE

[MW89] D. W. Masser and G. Wüstholz. Some effective estimates for elliptic curves. Lecture Notes in Mathematics, 1399:103–109, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095971/.

Meyer:1987:PDE

[MY87a] P.-A. Meyer and J. A. Yan. À propos des distributions sur l’espace de Wiener. (French) []. Lecture Notes in Mathematics, 1247:8– 26, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077625/.

Mostow:1987:SSC

[MY87b] G. D. Mostow and Stephen S.-T. Yau. Some surfaces covered by the ball and a problem in finite groups. Lecture Notes in Mathematics, 1271:201–228, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079239/.

Meyer:1985:CPN

[MZ85] P.-A. Meyer and W. A. Zheng. Construction de processus de Nelson reversibles. (French) []. Lecture Notes in Mathematics, 1123: 12–26, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075836/.

Nagai:1986:SPS

[Nag86a] Hideo Nagai. Stopping problems of symmetric Markov processes and non-linear variational inequalites. Lecture Notes in Mathematics, 1158:208–215, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080219/.

Nagel:1986:OPS

[Nag86b] Rainer Nagel. One-parameter Semigroups of Positive Operators, volume 1184 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 425 (est.) pp. LCCN QA3 .L28 no. 1184. URL http:

//link.springer.com/book/10.1007/BFb0074922;   http://
www.springerlink.com/content/978-3-540-39791-5.

Nagylaki:1986:NMG

[Nag86c] Thomas Nagylaki. Neutral models of geographical variation. Lecture Notes in Mathematics, 1212:216–237, 1986. CODEN LN-MAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076251/.

Nagase:1987:SCP

[Nag87] M. Nagase. On sufficient conditions for pseudo-differential operators to be L^{p}-bounded. Lecture Notes in Mathematics, 1256: 350–359, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077750/.

Nagy:1988:SIW

[Nag88] Attila Nagy. Subdirectly irreducible we-2 semigroups with globally idempotent core. Lecture Notes in Mathematics, 1320:244– 250, 1988. CODEN LNMAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083437/.

Nakaoka:1989:PBU

[Nak89] Minoru Nakaoka. Parametrized Borsuk–Ulam theorems and characteristic polynomials. Lecture Notes in Mathematics, 1411:155– 170, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086451/.

Narasimhan:1987:SVB

[Nar87] M. S. Narasimhan. Survey of vector bundles on curves. Lecture Notes in Mathematics, 1273:1–8, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078835/.

Nation:1985:SVS

[Nat85a] J. B. Nation. Some varieties of semidistributive lattices. Lecture Notes in Mathematics, 1149:198–223, 1985. CODEN LN-MAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098466/.

Natsume:1985:A

[Nat85b] Toshikazu Natsume. Appendix. Lecture Notes in Mathematics, 1132:312–323, 1985. CODEN LNMAA2. ISBN 3-540-15643-7 (print), 3-540-39514-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074891/.

Natterer:1986:NTI

[Nat86] Frank Natterer. Numerical treatment of ill-posed problems. Lecture Notes in Mathematics, 1225:142–167, 1986. CODEN LN-MAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072662/.

Nathanson:1989:APC

[Nat89a] Melvyn B. Nathanson. Additive problems in combinatorial number theory. Lecture Notes in Mathematics, 1383:123–139, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083574/.

Nathanson:1989:SCF

[Nat89b] Melvyn B. Nathanson. Sumsets containing k-free integers. Lecture Notes in Mathematics, 1380:179–184, 1989. CODEN LN-MAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086552/.

Neeman:1985:SBN

[Ne’85] Yuval Ne’eman. Supermanifolds and Berezin’s new integral. Lecture Notes in Mathematics, 1139:189–198, 1985. CODEN LN-MAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074584/.

Neeman:1987:SHD

[Ne’87] Yuval Ne’eman. Strings: From hadron dual models to gravity, unification and the structure of space–time. Lecture Notes in Mathematics, 1251:175–204, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077321/.

Necas:1986:ECT

[Neč86] J. Nečas. Entropy compactification of the transonic flow. Lecture Notes in Mathematics, 1192:399–408, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076100/.

Nedelec:1986:MFE

[Néd86] J.-C. Nédélec. Mixed finite element in 3D in H(div) and H(curl). Lecture Notes in Mathematics, 1192:321–325, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076088/.

Neher:1987:CG

[Neh87a] Erhard Neher. Classification of grids. Lecture Notes in Mathematics, 1280:59–106, 1987. CODEN LNMAA2. ISBN 3-540-18362-0 (print), 3-540-47921-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078219/.

Neher:1987:C

[Neh87b] Erhard Neher. Classifications. Lecture Notes in Mathematics, 1280:136–183, 1987. CODEN LNMAA2. ISBN 3-540-18362-0 (print), 3-540-47921-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078221/.

Neher:1987:CT

[Neh87c] Erhard Neher. Coordinatization theorems. Lecture Notes in Mathematics, 1280:107–135, 1987. CODEN LNMAA2. ISBN 3-540-18362-0 (print), 3-540-47921-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078220/.

Neher:1987:SFC

[Neh87d] Erhard Neher. Special families of compatible tripotents. Lecture Notes in Mathematics, 1280:1–58, 1987. CODEN LNMAA2. ISBN 3-540-18362-0 (print), 3-540-47921-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078218/.

Neidinger:1985:FOT

[Nei85] Richard D. Neidinger. Factoring operators through hereditarilyℓ^{P} spaces. Lecture Notes in Mathematics, 1166:116–128, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074701/.

Neidinger:1988:CRI

[Nei88] R. D. Neidinger. Concepts in the real interpolation of Banach spaces. Lecture Notes in Mathematics, 1332:43–53, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081610/.

Nelson:1985:CD

[Nel85] Edward Nelson. Critical diffusions. Lecture Notes in Mathematics, 1123:1–11, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075835/.

Nelson:1988:SMR

[Nel88] Edward Nelson. Stochastic mechanics and random fields. Lecture Notes in Mathematics, 1362:427–459, 1988. CODEN LN-MAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086184/.

Nelson:1989:FM

[Nel89] Edward Nelson. Front matter. Lecture Notes in Mathematics, 1365:??, 1989. CODEN LNMAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46075-6/1.pdf.

Neto:1988:ASP

[Net88a] Alcides Lins Neto. Algebraic solutions of polynomial differential equations and foliations in dimension two. Lecture Notes in Mathematics, 1345:192–232, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081403/.

Netsvetaev:1988:IID

[Net88b] N. Yu. Netsvetaev. Incomplete intersections and degenerations of complete intersections. Lecture Notes in Mathematics, 1346: 427–432, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082786/.

Neubrander:1986:ASB

[Neu86a] Frank Neubrander. Asymptotics of semigroups on Banach spaces. Lecture Notes in Mathematics, 1184:98–116, 1986. CODEN LN-MAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074926/.

Neuman:1986:OLD

[Neu86b] F. Neuman. Ordinary linear differential equations — a survey of the global theory. Lecture Notes in Mathematics, 1192:59– 70, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076052/.

Neustupa:1986:GEW

[Neu86c] J. Neustupa. The global existence of weak solutions of the mollified system of equations of motion of viscous compressible fluid. Lecture Notes in Mathematics, 1192:409–414, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076101/.

Neumann:1989:PRT

[Neu89] Peter M. Neumann. Pathology in the representation theory of infinite soluble groups. Lecture Notes in Mathematics, 1398:124– 139, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086249/.

Nevzorov:1987:TCU

[Nev87] V. B. Nevzorov. Two characterizations using records. Lecture Notes in Mathematics, 1233:79–85, 1987. CODEN LN-MAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072713/.

Nevanlinna:1989:NPL

[Nev89] Olavi Nevanlinna. A note on Picard–Lindelöf iteration. Lecture Notes in Mathematics, 1386:97–102, 1989. CODEN LN-MAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089233/.

Naveira:1986:DGP

[NFM86] Antonio M. Naveira, Angel Ferrández, and Francisca Mascaró, editors. Differential Geometry Peñíscola 1985: Proceedings of the 2nd International Symposium held at Peñíscola, Spain, June 2–9, 1985, volume 1209 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1209. URL http:

//link.springer.com/book/10.1007/BFb0076616;   http://
www.springerlink.com/content/978-3-540-44844-0.

Nguyen:1985:SAG

[Ngu85] Dũng Xuân Nguyen. Self-adjointness for general elliptic operators with Sobolev-type coefficients. Lecture Notes in Mathematics, 1151:132–141, 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074721/.

Nicaise:1987:ASP

[Nic87] Serge Nicaise. Approche spectrale des problèmes de diffusion sur les réseaux. (French) []. Lecture Notes in Mathematics, 1235:120– 140, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072753/.

Nicolas:1989:ACN

[Nic89] J. L. Nicolas. Appendix: On composite numbers. Lecture Notes in Mathematics, 1395:18–20, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086397/.

Nikulin:1987:OSE

[Nik87] V. N. Nikulin. On orthogonal-series estimators for probability distributions. Lecture Notes in Mathematics, 1233:86–92, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072714/.

Nirenberg:1988:FNS

[Nir88] Louis Nirenberg. Fully nonlinear second order elliptic equations. Lecture Notes in Mathematics, 1340:239–247, 1988. CODEN LN-MAA2. ISBN 3-540-50119-3 (print), 3-540-45932-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082899/.

Nirenberg:1989:VMN

[Nir89] L. Nirenberg. Variational methods in nonlinear problems. Lecture Notes in Mathematics, 1365:100–119, 1989. CODEN LN-MAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089179/.

Nishikawa:1986:DRM

[Nis86] Seiki Nishikawa. On deformation of Riemannian metrics and manifolds with positive curvature operator. Lecture Notes in Mathematics, 1201:202–211, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075657/.

Nitsche:1986:FBP

[Nit86] J. A. Nitsche. Free boundary problems for Stokes’ flows and finite element methods. Lecture Notes in Mathematics, 1192:327– 332, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076089/.

Njaatad:1986:MPP

[Njå86] Olav Njåtad. A multi-point Padé approximation problem. Lecture Notes in Mathematics, 1199:263–268, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075941/.

Njastad:1988:ORF

[Njå88] Olav Njåstad. Orthogonal rational functions with poles in a finite subset of R. Lecture Notes in Mathematics, 1329:300–307, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083369/.

Njaastad:1989:MPA

[Njå89a] Olav Njåstad. Multipoint Padé approximants and related continued fractions. Lecture Notes in Mathematics, 1406:76–87, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096167/.

Njaastad:1989:SSR

[Njå89b] Olav Njåstad. A survey of some results on separate convergence of continued fractions. Lecture Notes in Mathematics, 1406:88– 115, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096168/.

Nocedal:1986:VCG

[Noc86] Jorge Nocedal. Viewing the conjugate gradient method as a trust region algorithm. Lecture Notes in Mathematics, 1230:118– 126, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072675/.

Noll:1986:CMG

[Nol86] Walter Noll. Continuum mechanics and geometric integration theory. Lecture Notes in Mathematics, 1174:17–29, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076930/.

Norguet:1986:FPV

[Nor86a] François Norguet, editor. Fonctions de Plusieurs Variables Complexes V: Séminaire François Norguet Octobre 1979–Juin 1985, volume 1188 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1188. URL http:

//link.springer.com/book/10.1007/BFb0076817;   http://
www.springerlink.com/content/978-3-540-39797-7.

Norris:1986:SMC

[Nor86b] James Norris. Simplified Malliavin calculus. Lecture Notes in Mathematics, 1204:101–130, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075716/.

Notbohm:1988:FPC

[Not88] Dietrich Notbohm. The fixed-point conjecture for p-toral groups. Lecture Notes in Mathematics, 1361:253–260, 1988. CODEN LN-MAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083038/.

Novak:1988:AEB

[Nov88a] Erich Novak. Average error bounds. Lecture Notes in Mathematics, 1349:66–89, 1988. CODEN LNMAA2. ISBN 3-540-50368-4 (print), 3-540-45987-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079796/.

Novak:1988:DEB

[Nov88b] Erich Novak. Deterministic error bounds. Lecture Notes in Mathematics, 1349:9–42, 1988. CODEN LNMAA2. ISBN 3-540-50368-4 (print), 3-540-45987-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079794/.

Novak:1988:EBM

[Nov88c] Erich Novak. Error bounds for Monte Carlo methods. Lecture Notes in Mathematics, 1349:43–65, 1988. CODEN LN-MAA2. ISBN 3-540-50368-4 (print), 3-540-45987-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079795/.

Novak:1988:I

[Nov88d] Erich Novak. Introduction. Lecture Notes in Mathematics, 1349: 1–8, 1988. CODEN LNMAA2. ISBN 3-540-50368-4 (print), 3-540-45987-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079793/.

Novikov:1988:ATH

[Nov88e] S. P. Novikov. Analytical theory of homotopy groups. Lecture Notes in Mathematics, 1346:99–112, 1988. CODEN LN-MAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082773/.

Nowak:1985:EBT

[Now85] Werner Georg Nowak. Einige Beiträge Zur Theorie Der Gitterpunkte. (German) []. Lecture Notes in Mathematics, 1114:98– 117, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101649/.

Nowak:1987:HFS

[Now87] Sławomir Nowak. On the homology of function spaces. Lecture Notes in Mathematics, 1283:193–201, 1987. CODEN LN-MAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081428/.

Nagy:1989:WHC

[NP89a] Zs. Nagy and S. Purisch. When hereditarily collectionwise Hausdorffness implies regularity. Lecture Notes in Mathematics, 1401: 128–134, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097336/.

Nguyen:1989:HSH

[NP89b] Sang Nguyen and Stefano Pallottino. Hyperpaths and shortest hyperpaths. Lecture Notes in Mathematics, 1403:258–271, 1989.

CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083470/.

Naidu:1985:MAS

[NR85a] Desineni S. Naidu and Ayalasomayajula K. Rao. Modelling and analysis of singularly perturbed difference equations in state space form. Lecture Notes in Mathematics, 1154:41–119, 1985. CODEN LNMAA2. ISBN 3-540-15981-9 (print), 3-540-39681-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074762/.

Naidu:1985:SPAa

[NR85b] Desineni S. Naidu and Ayalasomayajula K. Rao. Singular perturbation analysis of difference equations in classical form. Lecture Notes in Mathematics, 1154:1–40, 1985. CODEN LNMAA2. ISBN 3-540-15981-9 (print), 3-540-39681-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074761/.

Naidu:1985:SPAb

[NR85c] Desineni S. Naidu and Ayalasomayajula K. Rao. Singular Perturbation Analysis of Discrete Control Systems, volume 1154 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-15981-9 (print), 3-540-39681-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 195 pp. LCCN QA3

.L28 no. 1154. URL http://link.springer.com/book/10.1007/
BFb0074760;   http://www.springerlink.com/content/978-
3-540-39681-9.

Naidu:1985:SPN

[NR85d] Desineni S. Naidu and Ayalasomayajula K. Rao. Singularly perturbed nonlinear difference equations and closed-loop discrete optimal control problem. Lecture Notes in Mathematics, 1154:148– 188, 1985. CODEN LNMAA2. ISBN 3-540-15981-9 (print), 3-540-39681-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074764/.

Naidu:1985:TTS

[NR85e] Desineni S. Naidu and Ayalasomayajula K. Rao. Three-time-scale difference equations with application to open-loop optimal control problem. Lecture Notes in Mathematics, 1154:120–147, 1985.

CODEN LNMAA2. ISBN 3-540-15981-9 (print), 3-540-39681-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074763/.

Naveira:1985:GOE

[NR85f] A. M. Naveira and A. H. Rocamora. A geometrical obstruction to the existence of two totally umbilical complementary foliations in compact manifolds. Lecture Notes in Mathematics, 1139:263– 279, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074589/.

Nerode:1985:GOR

[NR85g] A. Nerode and J. B. Remmel. Generic objects in recursion theory. Lecture Notes in Mathematics, 1141:271–314, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076226/.

Neumann:1988:EMN

[NR88] Walter D. Neumann and Lee Rudolph. The enhanced Milnor number in higher dimensions. Lecture Notes in Mathematics, 1350:109–121, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081471/.

Naftalevich:1985:TPS

[NS85] A. Naftalevich and M. Schreiber. Trigonometric polynomials and sums of squares. Lecture Notes in Mathematics, 1135:225– 238, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074607/.

Nagel:1986:BRB

[NS86a] Rainer Nagel and Ulf Schlotterbeck. Basic results on Banach lattices and positive operators. Lecture Notes in Mathematics, 1184:233–246, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074931/.

Nagel:1986:BRSa

[NS86b] Rainer Nagel and Ulf Schlotterbeck. Basic results on semigroups on Banach spaces. Lecture Notes in Mathematics, 1184:1–24, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074923/.

Nagel:1986:BRSb

[NS86c] Rainer Nagel and Ulf Schlotterbeck. Basic results on spaces C_{o}(X). Lecture Notes in Mathematics, 1184:117–121, 1986. CO-DEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074927/.

Nagel:1986:CSB

[NS86d] Rainer Nagel and Ulf Schlotterbeck. Characterization of semigroups on Banach spaces. Lecture Notes in Mathematics, 1184: 25–59, 1986. CODEN LNMAA2. ISBN 3-540-16454-5 (print), 3-540-39791-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074924/.

Naveira:1986:IIG

[NS86e] A. M. Naveira and S. Segura. The isoperimetric inequality and the geodesic spheres. Some geometric consequences. Lecture Notes in Mathematics, 1209:235–242, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076635/.

Nagaev:1987:PMS

[NS87] A. V. Nagaev and S. M. Shcolnick. Properties of mode of spectral positive stable distributions. Lecture Notes in Mathematics, 1233:69–78, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072712/.

Nagaev:1989:SAP

[NS89a] A. V. Nagaev and S. M. Shkolnik. Some asymptotic properties of the stable laws. Lecture Notes in Mathematics, 1412:229– 238, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084175/.

Norsett:1989:APR

[NS89b] Syvert P. Nørsett and Harald H. Simonsen. Aspects of parallel Runge–Kutta methods. Lecture Notes in Mathematics, 1386: 103–117, 1989. CODEN LNMAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089234/.

Nualart:1989:MCD

[NÜ89] David Nualart and Ali Süleyman Üstünel. Mesures cylindriques et distributions sur l’espace de Wiener. (French) []. Lecture Notes in Mathematics, 1390:186–191, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083945/.

Nualart:1986:MCS

[Nua86] David Nualart. Malliavin calculus and stochastic integrals. Lecture Notes in Mathematics, 1221:182–194, 1986. CODEN LN-MAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099114/.

Nualart:1988:NSI

[Nua88] David Nualart. Nonclausal stochastic integrals and calculus. Lecture Notes in Mathematics, 1316:80–129, 1988. CODEN LN-MAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081930/.

Nualart:1989:RDC

[Nua89] David Nualart. Une remarque sur le développement en chaos d’une diffusion. (French) []. Lecture Notes in Mathematics, 1372:165– 168, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083969/.

Nummelin:1989:CWL

[Num89] E. Nummelin. A conditional weak law of large numbers. Lecture Notes in Mathematics, 1412:259–262, 1989. CODEN LN-MAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084177/.

Nikulin:1989:CSG

[NV89] M. S. Nikulin and V. G. Voinov. A chi-square goodness-of-fit test for exponential distributions of the first order. Lecture Notes in Mathematics, 1412:239–258, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084176/.

Nakata:1986:SSR

[NW86] Mie Nakata and Mary Fanett Wheeler. Some superconvergence results for mixed finite element methods for linear parabolic problems. Lecture Notes in Mathematics, 1230:167–174, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072679/.

Njaastad:1989:SRN

[NW89] Olav Njåstad and Haakon Waadeland. Some remarks on nearness problems for continued fraction expansions. Lecture Notes in Mathematics, 1406:116–123, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096169/.

Nyikos:1989:CCS

[Nyi89] Peter J. Nyikos. Classes of compact sequential spaces. Lecture Notes in Mathematics, 1401:135–159, 1989. CODEN LN-MAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097337/.

Neumann:1985:NIF

[NZ85] Walter D. Neumann and Don Zagier. A note on an invariant of Fintushel and Stern. Lecture Notes in Mathematics, 1167:241– 244, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075227/.

Nualart:1989:SSI

[NZ89] D. Nualart and M. Zakai. A summary of some identities of the Malliavin calculus. Lecture Notes in Mathematics, 1390:192– 196, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083946/.

Obata:1989:LLM

[Oba89] Nobuaki Obata. The Levy Laplacian and mean value theorem. Lecture Notes in Mathematics, 1379:242–253, 1989. CODEN LN-MAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087857/.

Obretenov:1985:SCB

[Obr85] Apostol Obretenov. Some characterization of bivariate distributions based on hazard rate properties. Lecture Notes in Mathematics, 1155:223–241, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074821/.

Obretenov:1989:RCE

[Obr89] A. Obretenov. On the rate of convergence for the extreme value in the case of IFR-distributions. Lecture Notes in Mathematics, 1412:263–269, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084178/.

Ochanine:1988:GEE

[Och88] Serge Ochanine. Genres elliptiques équivariants. (French) []. Lecture Notes in Mathematics, 1326:107–122, 1988. CODEN LN-MAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078041/.

Ocneanu:1985:ADA

[Ocn85a] Adrian Ocneanu. Actions of discrete amenable groups on von Neumann algebras, volume 1138 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 115 pp. LCCN QA3 .L28

no. 1138. URL http://link.springer.com/book/10.1007/
BFb0098579;   http://www.springerlink.com/content/978-
3-540-39579-9.

Ocneanu:1985:AG

[Ocn85b] Adrian Ocneanu. Amenable groups. Lecture Notes in Mathematics, 1138:16–22, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098583/.

Ocneanu:1985:CV

[Ocn85c] Adrian Ocneanu. Cohomology vanishing. Lecture Notes in Mathematics, 1138:59–77, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098587/.

Ocneanu:1985:I

[Ocn85d] Adrian Ocneanu. Introduction. Lecture Notes in Mathematics, 1138:1–5, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098580/.

Ocneanu:1985:IC

[Ocn85e] Adrian Ocneanu. Invariants and classification. Lecture Notes in Mathematics, 1138:8–16, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098582/.

Ocneanu:1985:MR

[Ocn85f] Adrian Ocneanu. Main results. Lecture Notes in Mathematics, 1138:5–7, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098581/.

Ocneanu:1985:MA

[Ocn85g] Adrian Ocneanu. The model action. Lecture Notes in Mathematics, 1138:23–31, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098584/.

Ocneanu:1985:MAI

[Ocn85h] Adrian Ocneanu. Model action isomorphism. Lecture Notes in Mathematics, 1138:95–111, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098589/.

Ocneanu:1985:MAS

[Ocn85i] Adrian Ocneanu. Model action splitting. Lecture Notes in Mathematics, 1138:77–95, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098588/.

Ocneanu:1985:RT

[Ocn85j] Adrian Ocneanu. The Rohlin theorem. Lecture Notes in Mathematics, 1138:41–59, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098586/.

Ocneanu:1985:UA

[Ocn85k] Adrian Ocneanu. Ultraproduct algebras. Lecture Notes in Mathematics, 1138:31–40, 1985. CODEN LNMAA2. ISBN 3-540-15663-1 (print), 3-540-39579-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098585/.

Ocone:1987:PDS

[Oco87] Daniel Ocone. Probability distributions of solutions to some stochastic partial differential equations. Lecture Notes in Mathematics, 1236:184–199, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072890/.

Ocone:1988:GSC

[Oco88] Daniel L. Ocone. A guide to the stochastic calculus of variations. Lecture Notes in Mathematics, 1316:1–79, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081929/.

Oberst:1985:AFG

[OD85a] Ulrich Oberst and Arne Dür. Actions of formal groups on formal schemes. Applications to control theory and combinatorics. Lecture Notes in Mathematics, 1146:214–269, 1985. CODEN LN-MAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074540/.

Oberst:1985:CCA

[OD85b] Ulrich Oberst and Arne Dür. A constructive characterization of all optimal linear codes. Lecture Notes in Mathematics, 1146: 176–213, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074539/.

Odifreddi:1985:SD

[Odi85] Piergiorgio Odifreddi. The structure of m-degrees. Lecture Notes in Mathematics, 1141:315–332, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076227/.

Ofman:1986:ITOc

[Ofm86a] Salomon Ofman. Injectivité de la transformation obtenue par intégration sur les cycles analytioues. (French) []. Lecture Notes in Mathematics, 1188:201–228, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076823/.

Ofman:1986:ITOa

[Ofm86b] Salomon Ofman. Injectivité de la transformation obtenue par intégration sur les cycles analytiques. (French) []. Lecture Notes in Mathematics, 1188:183–189, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076821/.

Ofman:1986:ITOb

[Ofm86c] Salomon Ofman. Injectivité de la transformation obtenue par intégration sur les cycles analytiques. (French) []. Lecture Notes in Mathematics, 1188:190–200, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076822/.

Ofman:1986:RDF

[Ofm86d] Salomon Ofman. Residu et dualité. (French) []. Lecture Notes in Mathematics, 1188:1–22, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076818/.

Ofman:1986:TRI

[Ofm86e] Salomon Ofman. Transformée de Radon et intégrale sur les cycles analytiques compacts de certains ouverts de l’espace projectif complexe. (French) []. Lecture Notes in Mathematics, 1188: 23–44, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076819/.

Oharu:1986:SEE

[Oha86] Shinnosuke Oharu. Semilinear evolution equations in Fréchet spaces. Lecture Notes in Mathematics, 1223:186–207, 1986. CO-DEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0099193/.

Ohtsuka:1988:WEL

[Oht88] Makoto Ohtsuka. Weighted extremal length and Beppo Levi functions. Lecture Notes in Mathematics, 1344:154–161, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103349/.

Okada:1985:TBE

[Oka85] Masami Okada. Transcendental Bézout estimate by the logarithmic function in C^{n}. Lecture Notes in Mathematics, 1165:204– 210, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076155/.

Okninski:1988:CMR

[Okn88a] Jan Okniński. Commutative monoid rings with Krull dimension. Lecture Notes in Mathematics, 1320:251–259, 1988. CODEN LN-MAA2. ISBN 3-540-19347-2 (print), 3-540-39225-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083438/.

Okninski:1988:NPS

[Okn88b] Jan Okniński. Noetherian property for semigroup rings. Lecture Notes in Mathematics, 1328:209–218, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100926/.

Okoneshnikov:1985:CIP

[Oko85] Yu. D. Okoneshnikov. Confidence intervals for the parameters of strongly stable laws. Lecture Notes in Mathematics, 1155:242– 255, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074822/.

Okonek:1986:CCC

[Oko86a] Christian Okonek. Concavity, convexity and complements in complex spaces. Lecture Notes in Mathematics, 1194:104–126, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076998/.

Okonek:1986:SHM

[Oko86b] Christian Okonek. Subvarieties in homogeneous manifolds. Lecture Notes in Mathematics, 1194:127–132, 1986. CODEN LN-MAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076999/.

Oliver:1985:PCG

[Oli85] Robert Oliver. Projective class groups of integral group rings: a survey. Lecture Notes in Mathematics, 1142:211–232, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074803/.

Ollagnier:1985:DSA

[Oll85a] Jean Moulin Ollagnier. Dynamical systems and amenable groups. Lecture Notes in Mathematics, 1115:15–34, 1985. CODEN LN-MAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101577/.

Ollagnier:1985:DSS

[Oll85b] Jean Moulin Ollagnier. Dynamical systems in statistical mechanics. Lecture Notes in Mathematics, 1115:105–116, 1985. CODEN LNMAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101582/.

Ollagnier:1985:EFV

[Oll85c] Jean Moulin Ollagnier. Entropy as a function and the variational principle. Lecture Notes in Mathematics, 1115:71–85, 1985. CODEN LNMAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101580/.

Ollagnier:1985:EAD

[Oll85d] Jean Moulin Ollagnier. Entropy of abstract dynamical systems. Lecture Notes in Mathematics, 1115:53–70, 1985. CODEN LN-MAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101579/.

Ollagnier:1985:ECA

[Oll85e] Jean Moulin Ollagnier. Equivalence of countable amenable groups. Lecture Notes in Mathematics, 1115:117–138, 1985. CODEN LN-MAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101583/.

Ollagnier:1985:ET

[Oll85f] Jean Moulin Ollagnier. Ergodic theorems. Lecture Notes in Mathematics, 1115:35–52, 1985. CODEN LNMAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101578/.

Ollagnier:1985:ETS

[Oll85g] Jean Moulin Ollagnier. Ergodic Theory and Statistical Mechanics, volume 1115 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 147 pp. LCCN QA3 .L28 no. 1115. URL http:

//link.springer.com/book/10.1007/BFb0101575;   http://
www.springerlink.com/content/978-3-540-39289-7.

Ollagnier:1985:PA

[Oll85h] Jean Moulin Ollagnier. Preliminary analysis. Lecture Notes in Mathematics, 1115:1–14, 1985. CODEN LNMAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101576/.

Ollagnier:1985:SML

[Oll85i] Jean Moulin Ollagnier. Statistical mechanics on a lattice. Lecture Notes in Mathematics, 1115:86–104, 1985. CODEN LN-MAA2. ISBN 3-540-15192-3 (print), 3-540-39289-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101581/.

Olver:1987:ITD

[Olv87] Peter J. Olver. Invariant theory and differential equations. Lecture Notes in Mathematics, 1278:62–80, 1987. CODEN LN-MAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078807/.

Omey:1989:RCE

[Ome89] E. Omey. On the rate of convergence in extreme value theory. Lecture Notes in Mathematics, 1412:270–279, 1989. CODEN LN-MAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084179/.

Ocone:1989:LAC

[OP89] Daniel Ocone and Étienne Pardoux. A Lie algebraic criterion for non-existence of finite dimensionally computable filters. Lecture Notes in Mathematics, 1390:197–204, 1989. CODEN LN-MAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083947/.

Obretenov:1987:EDB

[OR87a] A. Obretenov and S. Rachev. Estimates of the deviation between the exponential and new classes of bivariate distributions. Lecture Notes in Mathematics, 1233:93–102, 1987. CODEN LN-MAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072715/.

Oeljeklaus:1987:RRH

[OR87b] Karl Oeljeklaus and Wolfgang Richthofer. Recent results on homogeneous complex manifolds. Lecture Notes in Mathematics, 1277: 78–119, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078246/.

Orloff:1987:OIS

[Orl87] Jeremy Orloff. Orbital integrals on symmetric spaces. Lecture Notes in Mathematics, 1243:198–239, 1987. CODEN LN-MAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073024/.

Orobitg:1989:SSS

[Oro89] Joan Orobitg. Spectral synthesis in spaces of functions with derivatives in H^{1}. Lecture Notes in Mathematics, 1384:202–206, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086804/.

Okonek:1985:STG

[OS85] Christian Okonek and Heinz Spindler. Das Spektrum torsionsfreier Garben II. (German) []. Lecture Notes in Mathematics, 1165:211–234, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076156/.

Oshima:1987:ACT

[Ōsh87] Y. Ōshima. On absolute continuity of two symmetric diffusion processes. Lecture Notes in Mathematics, 1250:184–194, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077358/.

Ostrovskii:1985:GTC

[Ost85] I. V. Ostrovskiĭ. Generalization of the Titchmarsh convolution theorem and the complex-valued measures uniquely determined by their restrictions to a half-line. Lecture Notes in Mathematics, 1155:256–283, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074823/.

Oshima:1987:TSM

[ŌT87] Yōichi Ōshima and Masayoshi Takeda. On a transformation of symmetric Markov process and recurrence property. Lecture Notes in Mathematics, 1250:171–183, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077357/.

Ouadghiri:1988:MSH

[Oua88] A. Ouadghiri. Un modêle de Sullivan en homotopie modérée. (French) []. Lecture Notes in Mathematics, 1318:162–191, 1988.

CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077802/.

Omey:1987:DBD

[OW87] E. Omey and E. Willekens. On the difference between distributions of sums and maxima. Lecture Notes in Mathematics, 1233:103–113, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072716/.

Owen:1986:GLV

[Owe86] David R. Owen. Global and local versions of the second law of thermodynamics. Lecture Notes in Mathematics, 1174:100–114, 1986. CODEN LNMAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076937/.

Ozawa:1985:NTS

[Oza85] Tetsuya Ozawa. On the number of tritangencies of a surface in IR^{3}. Lecture Notes in Mathematics, 1156:240–253, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075096/.

Ozawa:1988:AGF

[Oza88] Shin Ozawa. Approximation of Green’s function in a region with many obstacles. Lecture Notes in Mathematics, 1339:212– 225, 1988. CODEN LNMAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083057/.

Paneitz:1985:IFD

[Pan85a] Stephen M. Paneitz. Indecomposable finite dimensional representations of the Poincaré group and associated fields. Lecture Notes in Mathematics, 1139:6–9, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074574/.

Pantcheva:1985:LTE

[Pan85b] E. Pantcheva. Limit theorems for extreme order statistics under nonlinear normalization. Lecture Notes in Mathematics, 1155: 284–309, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074824/.

Pansu:1986:QMM

[Pan86] Pierre Pansu. Quasiconformal mappings and manifolds of negative curvature. Lecture Notes in Mathematics, 1201:212–229, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075658/.

Papamichael:1985:TSO

[Pap85] N. Papamichael. The treatment of singularities in orthonormalization methods for numerical conformal mapping. Lecture Notes in Mathematics, 1121:207–227, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076273/.

Papanicolaou:1988:RPI

[Pap88a] George C. Papanicolaou. Remarks on the point interaction approximation. Lecture Notes in Mathematics, 1362:265–275, 1988. CODEN LNMAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086182/.

Papanicolaou:1988:WOD

[Pap88b] George C. Papanicolaou. Waves in one-dimensional random media. Lecture Notes in Mathematics, 1362:205–264, 1988. CODEN LNMAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086181/.

Pardoux:1986:GFR

[Par86a] É. Pardoux. Grossissement d’une filtration et retournement du temps d’une diffusion. (French) []. Lecture Notes in Mathematics, 1204:48–55, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075711/.

Parthasarathy:1986:QSC

[Par86b] K. R. Parthasarathy. Quantum stochastic calculus. Lecture Notes in Mathematics, 1203:177–196, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076881/.

Parthasarathy:1986:RPM

[Par86c] K. R. Parthasarathy. A remark on the paper ‘Une martingale d’opérateurs bornés, non représentable en intégrale stochastique’, by J. L. Journé and P.-A. Meyer. Lecture Notes in Mathematics, 1204:317–320, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075721/. See [JM86b].

Parthasarathy:1986:SAR

[Par86d] K. R. Parthasarathy. Some additional remarks on Fock space stochastic calculus. Lecture Notes in Mathematics, 1204:331– 333, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075723/.

Pardoux:1987:TSS

[Par87a] É. Pardoux. Two-sided stochastic calculus for spdes. Lecture Notes in Mathematics, 1236:200–207, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072891/.

Parusnikov:1987:CML

[Par87b] V. I. Parusnikov. On the convergence of the multidimensional limit-periodic continued fractions. Lecture Notes in Mathematics, 1237:217–227, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072466/.

Pardo:1988:RQE

[Par88a] J. L. Gómez Pardo. Rings of quotients of endomorphism rings. Lecture Notes in Mathematics, 1328:106–121, 1988. CODEN LN-MAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100920/.

Parthasarathy:1988:LMF

[Par88b] K. R. Parthasarathy. Local measures in Fock space stochastic calculus and a generalized Itô–Tanaka formula. Lecture Notes in Mathematics, 1303:213–231, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078065/.

Parthasarathy:1989:SCQ

[Par89] K. R. Parthasarathy. Some comments on quantum probability. Lecture Notes in Mathematics, 1372:169–174, 1989. CODEN LN-MAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083970/.

Pasini:1986:CFG

[Pas86] Antonio Pasini. Covers of finite geometries with non-spherical minimal circuit diagram. Lecture Notes in Mathematics, 1181: 218–241, 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075517/.

Passare:1987:CME

[Pas87a] Mikael Passare. Courants meromorphes et égalité de la valeur principale et de la partie finie. (French) []. Lecture Notes in Mathematics, 1295:157–166, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081982/.

Paszkowski:1987:QGR

[Pas87b] Stefan Paszkowski. Quelques generalisations de la representation de reels par des fractions continues. (French) []. Lecture Notes in Mathematics, 1237:228–238, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072467/.

Paszkiewicz:1989:CAT

[Pas89] Adam Paszkiewicz. Convergences in W ∗-algebras — their strange behaviour and tools for their investigation. Lecture Notes in Mathematics, 1396:279–294, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083558/.

Pati:1987:CAV

[Pat87] Vishwambhar Pati. L^{2}-cohomology of algebraic varieties in the Fubinu metric. Lecture Notes in Mathematics, 1246:154–164, 1987. CODEN LNMAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077537/.

Paterson:1988:CLC

[Pat88] Alan L. T. Paterson. The class of locally compact groups G for which C ∗ (G) is amenable. Lecture Notes in Mathematics, 1359:226–237, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086603/.

Paul:1986:CDA

[Pau86] T. Paul. A characterization of dilation-analytic operators. Lecture Notes in Mathematics, 1218:179–189, 1986. CODEN LN-MAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073050/.

Pavlista:1985:SEL

[Pav85] Targo Pavlista. Small eigenvalues of the Laplacian and examples. Lecture Notes in Mathematics, 1156:254–263, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075097/.

Pavel:1987:APD

[Pav87a] Nicolae H. Pavel. Applications to partial differential equations. Lecture Notes in Mathematics, 1260:152–233, 1987. CODEN LN-MAA2. ISBN 3-540-17974-7 (print), 3-540-47186-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077771/.

Pavel:1987:NEOa

[Pav87b] Nicolae H. Pavel. Nonlinear evolution operators. Lecture Notes in Mathematics, 1260:1–60, 1987. CODEN LNMAA2. ISBN 3-540-17974-7 (print), 3-540-47186-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077769/.

Pavel:1987:NEOb

[Pav87c] Nicolae H. Pavel. Nonlinear Evolution Operators and Semigroups: Applications to Partial Differential Equations, volume 1260 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LN-MAA2. ISBN 3-540-17974-7 (print), 3-540-47186-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 285 pp. LCCN QA3

.L28 no. 1260. URL http://link.springer.com/book/10.1007/
BFb0077768;   http://www.springerlink.com/content/978-
3-540-47186-8.

Pavel:1987:NS

[Pav87d] Nicolae H. Pavel. Nonlinear semigroups. Lecture Notes in Mathematics, 1260:61–151, 1987. CODEN LNMAA2. ISBN 3-540-17974-7 (print), 3-540-47186-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077770/.

Pawlucki:1985:QRB

[Paw85] Wiesław Pawłucki. Quasi-regular boundary and Stokes’ formula for a sub-analytic leaf. Lecture Notes in Mathematics, 1165:235– 252, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076157/.

Pedersen:1985:TAS

[Ped85] Erik Kjaer Pedersen. Topological H_{0} × H_{1}-actions on spheres and linking numbers. Lecture Notes in Mathematics, 1172:163– 166, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074430/.

Peetre:1988:MRL

[Pee88] Jaak Peetre. Marcel Riesz in Lund. Lecture Notes in Mathematics, 1302:1–10, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078859/.

Persson:1985:DSS

[Per85] Ulf Persson. Double sextics and singular K-3 surfaces. Lecture Notes in Mathematics, 1124:262–328, 1985. CODEN LN-MAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075003/.

Peternell:1986:RCM

[Pet86] Thomas Peternell. Rational curves in Moisězon 3-folds. Lecture Notes in Mathematics, 1194:133–144, 1986. CODEN LN-MAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077000/.

Petz:1988:CEQ

[Pet88] Dénes Petz. Conditional expectation in quantum probability. Lecture Notes in Mathematics, 1303:251–260, 1988. CODEN LN-MAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078067/.

Petho:1989:RBC

[Pet89a] Attila Pethő. On the representation of 1 by binary cubic forms with positive discriminant. Lecture Notes in Mathematics, 1380: 185–196, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086553/.

Petkovic:1989:BC

[Pet89b] Miodrag Petković. Basic concepts. Lecture Notes in Mathematics, 1387:10–30, 1989. CODEN LNMAA2. ISBN 3-540-51485-6 (print), 3-540-48174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083601/.

Petkovic:1989:BPP

[Pet89c] Miodrag Petković. Bell’s polynomials and parallel disk iterations. Lecture Notes in Mathematics, 1387:163–220, 1989. CODEN LN-MAA2. ISBN 3-540-51485-6 (print), 3-540-48174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083604/.

Petkovic:1989:CES

[Pet89d] Miodrag Petković. Computational efficiency of simultaneous methods. Lecture Notes in Mathematics, 1387:221–249, 1989. CODEN LNMAA2. ISBN 3-540-51485-6 (print), 3-540-48174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083605/.

Petkovic:1989:GRI

[Pet89e] Miodrag Petković. Generalized root iterations. Lecture Notes in Mathematics, 1387:69–162, 1989. CODEN LNMAA2. ISBN 3-540-51485-6 (print), 3-540-48174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083603/.

Petkovic:1989:I

[Pet89f] Miodrag Petković. Introduction. Lecture Notes in Mathematics, 1387:1–9, 1989. CODEN LNMAA2. ISBN 3-540-51485-6 (print), 3-540-48174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083600/.

Petkovic:1989:IMD

[Pet89g] Miodrag Petković. Iterative methods without derivatives. Lecture Notes in Mathematics, 1387:31–68, 1989. CODEN LN-MAA2. ISBN 3-540-51485-6 (print), 3-540-48174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083602/.

Petz:1989:PMM

[Pet89h] Dénes Petz. Positive mappings on matrix algebras. Lecture Notes in Mathematics, 1396:295–303, 1989. CODEN LNMAA2. ISBN 3-540-51613-1 (print), 3-540-46713-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083559/.

Peng:1989:RIP

[PH89] Chia-Kuei Peng and Zixin Hou. A remark on the isoparametric polynomials of degree 6. Lecture Notes in Mathematics, 1369: 222–224, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087535/.

Phelps:1989:ASR

[Phe89a] Robert R. Phelps. Asplund spaces, the Radon–Nikodým property and optimization. Lecture Notes in Mathematics, 1364:72– 89, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21569-2_5/.

Phelps:1989:BM

[Phe89b] Robert R. Phelps. Back matter. Lecture Notes in Mathematics, 1364:108–118, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21569-2/1.pdf.

Phelps:1989:CFR

[Phe89c] Robert R. Phelps. Convex functions on real Banach spaces. Lecture Notes in Mathematics, 1364:1–16, 1989. CODEN LN-MAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21569-2_1/.

Phelps:1989:FM

[Phe89d] Robert R. Phelps. Front matter. Lecture Notes in Mathematics, 1364:I–IX, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21569-2/1.pdf.

Phelps:1989:GDS

[Phe89e] Robert R. Phelps. Gâteaux differentiability spaces. Lecture Notes in Mathematics, 1364:90–96, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21569-2_6/.

Phelps:1989:GMO

[Phe89f] Robert R. Phelps. A generalization of monotone operators: Usco maps. Lecture Notes in Mathematics, 1364:97–103, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-662-21569-2_7/

.

Phelps:1989:LSC

[Phe89g] Robert R. Phelps. Lower semicontinuous convex functions. Lecture Notes in Mathematics, 1364:40–63, 1989. CODEN LN-MAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21569-2_3/.

Phelps:1989:MOS

[Phe89h] Robert R. Phelps. Monotone operators, subdifferentials and Asplund spaces. Lecture Notes in Mathematics, 1364:17–39, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21569-2_2/.

Phelps:1989:NR

[Phe89i] Robert R. Phelps. Notes and remarks. Lecture Notes in Mathematics, 1364:104–107, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21569-2_8/.

Phelps:1989:SVP

[Phe89j] Robert R. Phelps. A smooth variational principle and more about Asplund spaces. Lecture Notes in Mathematics, 1364:64– 71, 1989. CODEN LNMAA2. ISBN 3-540-50735-3 (print), 3-662-21569-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21569-2_4/.

Phillips:1987:AA

[Phi87a] N. Christopher Phillips. Af algebras. Lecture Notes in Mathematics, 1274:286–328, 1987. CODEN LNMAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078666/.

Phillips:1987:BPF

[Phi87b] N. Christopher Phillips. Basic properties of K-freeness. Lecture Notes in Mathematics, 1274:131–166, 1987. CODEN LN-MAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078661/.

Phillips:1987:ETC

[Phi87c] N. Christopher Phillips. Equivariant K-theory of c *-algebras. Lecture Notes in Mathematics, 1274:12–67, 1987. CODEN LN-MAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078659/.

Phillips:1987:ICC

[Phi87d] N. Christopher Phillips. Introduction: The commutative case. Lecture Notes in Mathematics, 1274:1–11, 1987. CODEN LN-MAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078658/.

Phillips:1987:IET

[Phi87e] N. Christopher Phillips. Introduction to equivariant KK-theory. Lecture Notes in Mathematics, 1274:68–130, 1987. CODEN LN-MAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078660/.

Phillips:1987:FSS

[Phi87f] N. Christopher Phillips. K-freeness, saturation, and the strong Connes spectrum. Lecture Notes in Mathematics, 1274:236– 274, 1987. CODEN LNMAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078664/.

Phillips:1987:S

[Phi87g] N. Christopher Phillips. Subgroups. Lecture Notes in Mathematics, 1274:167–188, 1987. CODEN LNMAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078662/.

Phillips:1987:TP

[Phi87h] N. Christopher Phillips. Tensor products. Lecture Notes in Mathematics, 1274:189–235, 1987. CODEN LNMAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078663/.

Phillips:1987:TA

[Phi87i] N. Christopher Phillips. Type I algebras. Lecture Notes in Mathematics, 1274:275–285, 1987. CODEN LNMAA2. ISBN 3-540-18277-2 (print), 3-540-47868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078665/.

Picco:1985:ABS

[Pic85] P. Picco. On the absence of breakdown of symmetry for the plane rotator model with long range unbounded random interaction. Lecture Notes in Mathematics, 1109:148–154, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101542/.

Picard:1986:CPS

[Pic86] Jean Picard. Une classe de processus stable par retournement du temps. (French) []. Lecture Notes in Mathematics, 1204:56– 67, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075712/.

Piccioni:1987:CID

[Pic87] M. Piccioni. Convergence of implicit discretization schemes for linear differential equations with application to filtering. Lecture Notes in Mathematics, 1236:208–229, 1987. CODEN LN-MAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072892/.

Picard:1988:ASD

[Pic88] Jean Picard. Approximation of stochastic differential equations and application of the stochastic calculus of variations to the rate of convergence. Lecture Notes in Mathematics, 1316:267– 287, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081935/.

Picard:1989:MCF

[Pic89a] Jean Picard. Martingales sur le cercle. (French) []. Lecture Notes in Mathematics, 1372:147–160, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083967/.

Piccioni:1989:GWT

[Pic89b] Mauro Piccioni. A generalization of Wahba’s theorem on the equivalence between spline smoothing and Bayesian estimation. Lecture Notes in Mathematics, 1390:205–220, 1989. CODEN LN-MAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083948/.

Piene:1985:PET

[Pie85] Ragni Piene. On the problem of enumerating twisted cubics. Lecture Notes in Mathematics, 1124:329–337, 1985. CODEN LN-MAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075004/.

Piedra:1988:SRI

[Pie88a] René Piedra. On simultaneous rational interpolants of type (α, β). Lecture Notes in Mathematics, 1354:188–198, 1988. CODEN LN-MAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089596/.

Pier:1988:SVE

[Pie88b] Jean-Paul Pier. Some views on the evolution of harmonic analysis. Lecture Notes in Mathematics, 1359:1–15, 1988. CODEN LNMAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086585/.

Pierce:1989:ARP

[Pie89a] John F. Pierce. The analysis of the reduced problem. Lecture Notes in Mathematics, 1377:117–139, 1989. CODEN LNMAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093786/.

Pierce:1989:CAP

[Pie89b] John F. Pierce. Conclusions and additional problems. Lecture Notes in Mathematics, 1377:163–166, 1989. CODEN LN-MAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093788/.

Pierce:1989:I

[Pie89c] John F. Pierce. Introduction. Lecture Notes in Mathematics, 1377: 1–6, 1989. CODEN LNMAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093781/.

Pierce:1989:RBPa

[Pie89d] John F. Pierce. The reduction of the bifurcation problem. Lecture Notes in Mathematics, 1377:90–116, 1989. CODEN LN-MAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093785/.

Pierce:1989:RBPb

[Pie89e] John F. Pierce. The results for the bifurcation problem. Lecture Notes in Mathematics, 1377:140–162, 1989. CODEN LN-MAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093787/.

Pierce:1989:REP

[Pie89f] John F. Pierce. The rod equilibrium problem. Lecture Notes in Mathematics, 1377:66–89, 1989. CODEN LNMAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093784/.

Pierce:1989:SC

[Pie89g] John F. Pierce. The spaces of configurations. Lecture Notes in Mathematics, 1377:7–31, 1989. CODEN LNMAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093782/.

Pierce:1989:SL

[Pie89h] John F. Pierce. The spaces of loads. Lecture Notes in Mathematics, 1377:32–65, 1989. CODEN LNMAA2. ISBN 3-540-51304-3 (print), 3-540-46190-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093783/.

Pike:1986:BM

[Pik86a] E. R. Pike. Back matter. Lecture Notes in Mathematics, 1225: 201–209, 1986. CODEN LNMAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47353-4/1.pdf.

Pike:1986:ISF

[Pik86b] E. R. Pike. On the inversion of some first-kind Fredholm equations occurring in optical applications. Lecture Notes in Mathematics, 1225:182–200, 1986. CODEN LNMAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072664/.

Pillay:1987:SST

[Pil87] Anand Pillay. Simple superstable theories. Lecture Notes in Mathematics, 1292:247–263, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082241/.

Pinsky:1986:MET

[Pin86] Mark A. Pinsky. Mean exit times and hitting probabilities of Brownian motion in geodesic balls and tubular neighborhoods. Lecture Notes in Mathematics, 1158:216–223, 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080220/.

Pirillo:1987:PPS

[Pir87] Giuseppe Pirillo. On permutation properties for semigroups. Lecture Notes in Mathematics, 1281:118–119, 1987. CODEN LN-MAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078697.

Pirola:1989:NSG

[Pir89] Gian Pietro Pirola. On the Néron–Severi groups of the surfaces of special divisors. Lecture Notes in Mathematics, 1389: 242–245, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085935/.

Pisier:1986:PMG

[Pis86] Gilles Pisier. Probabilistic methods in the geometry of Banach spaces. Lecture Notes in Mathematics, 1206:167–241, 1986. CO-DEN LNMAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076302/.

Pitman:1987:SE

[Pit87] Jim Pitman. Stationary excursions. Lecture Notes in Mathematics, 1247:289–302, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077643/.

Pixley:1985:PCF

[Pix85] A. F. Pixley. Principal congruence formulas in arithmetical varieties. Lecture Notes in Mathematics, 1149:238–254, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098468/.

Plesniak:1987:PCL

[Ple87] Wiesław Pleśniak. Polynomial condition of Leja. Lecture Notes in Mathematics, 1237:41–45, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072452/.

Plotkin:1988:GTD

[Plo88] B. I. Plotkin. Galois theory of databases. Lecture Notes in Mathematics, 1352:147–161, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082023/.

Palais:1988:ACP

[PlT88a] Richard S. Palais and Chuu liang Terng. Advanced critical point theory. Lecture Notes in Mathematics, 1353:223–232, 1988. CO-DEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087452/.

Palais:1988:BM

[PlT88b] Richard S. Palais and Chuu liang Terng. Back matter. Lecture Notes in Mathematics, 1353:??, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-45996-5/1.pdf.

Palais:1988:CV

[PlT88c] Richard S. Palais and Chuu liang Terng. The calculus of variations. Lecture Notes in Mathematics, 1353:233–259, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087453/.

Palais:1988:ECP

[PlT88d] Richard S. Palais and Chuu liang Terng. Elementary critical point theory. Lecture Notes in Mathematics, 1353:181–222, 1988. CO-DEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087451/.

Palais:1988:FP

[PlT88e] Richard S. Palais and Chuu liang Terng. Focal points. Lecture Notes in Mathematics, 1353:65–71, 1988. CODEN LN-MAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087446/.

Palais:1988:FM

[PlT88f] Richard S. Palais and Chuu liang Terng. Front matter. Lecture Notes in Mathematics, 1353:??, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-45996-5/1.pdf.

Palais:1988:IS

[PlT88g] Richard S. Palais and Chuu liang Terng. Isoparametric submanifolds. Lecture Notes in Mathematics, 1353:106–139, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087448/.

Palais:1988:LGS

[PlT88h] Richard S. Palais and Chuu liang Terng. Local geometry of submanifolds. Lecture Notes in Mathematics, 1353:25–42, 1988. CO-DEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087444/.

Palais:1988:P

[PlT88i] Richard S. Palais and Chuu liang Terng. Preliminaries. Lecture Notes in Mathematics, 1353:3–24, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087443/.

Palais:1988:PFS

[PlT88j] Richard S. Palais and Chuu liang Terng. Proper Fredholm submanifolds in Hilbert spaces. Lecture Notes in Mathematics, 1353: 140–149, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087449/.

Palais:1988:TIS

[PlT88k] Richard S. Palais and Chuu liang Terng. Topology of isoparametric submanifolds. Lecture Notes in Mathematics, 1353:150–177, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087450/.

Palais:1988:TG

[PlT88l] Richard S. Palais and Chuu liang Terng. Transformation groups. Lecture Notes in Mathematics, 1353:72–105, 1988. CODEN LN-MAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087447/.

Palais:1988:WST

[PlT88m] Richard S. Palais and Chuu liang Terng. Weingarten surfaces in three dimensional space forms. Lecture Notes in Mathematics, 1353:43–64, 1988. CODEN LNMAA2. ISBN 3-540-50399-4 (print), 3-540-45996-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087445/.

Plucinska:1985:CGD

[Plu85] Agnieszka Plucińska. A characterization of Gaussian distribution by the linearity of conditional moments. Lecture Notes in Mathematics, 1155:310–319, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074825/.

Plucinska:1989:SPS

[Plu89] A. Plucińska. Some properties of stochastic processes with linear regression. Lecture Notes in Mathematics, 1412:280–287, 1989. CODEN LNMAA2. ISBN 3-540-51948-3 (print), 3-540-46863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084180/.

Perez:1989:DIB

[PM89] P. L. García Pérez and J. Muñoz Masqué. Differential invariants on the bundles of g-structures. Lecture Notes in Mathematics, 1410:177–201, 1989. CODEN LNMAA2. ISBN 3-540-51885-1 (print), 3-540-46858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086422/.

Pohlers:1989:AOI

[Poh89a] Wolfram Pohlers. The autonomous ordinal of the infinitary system Z_{∞} and the limits of predicativity. Lecture Notes in Mathematics, 1407:77–108, 1989. CODEN LNMAA2. ISBN 3-540-51842-8 (print), 3-540-46825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46825-7_3/.

Pohlers:1989:BM

[Poh89b] Wolfram Pohlers. Back matter. Lecture Notes in Mathematics, 1407:187–213, 1989. CODEN LNMAA2. ISBN 3-540-51842-8 (print), 3-540-46825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-46825-7/1.pdf.

Pohlers:1989:FM

[Poh89c] Wolfram Pohlers. Front matter. Lecture Notes in Mathematics, 1407:I–VIII, 1989. CODEN LNMAA2. ISBN 3-540-51842-8 (print), 3-540-46825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-46825-7/1.pdf.

Pohlers:1989:I

[Poh89d] Wolfram Pohlers. Introduction. Lecture Notes in Mathematics, 1407:1–6, 1989. CODEN LNMAA2. ISBN 3-540-51842-8 (print), 3-540-46825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46825-7_1/.

Pohlers:1989:OAP

[Poh89e] Wolfram Pohlers. Ordinal analysis of pure number theory. Lecture Notes in Mathematics, 1407:7–76, 1989. CODEN LNMAA2. ISBN 3-540-51842-8 (print), 3-540-46825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46825-7_2/.

Pohlers:1989:OAF

[Poh89f] Wolfram Pohlers. Ordinal analysis of the formal theory for noniterated inductive definitions. Lecture Notes in Mathematics, 1407:

109–186, 1989. CODEN LNMAA2. ISBN 3-540-51842-8 (print), 3-540-46825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46825-7_4/.

Poly:1986:PBM

[Pol86] J.-B. Poly. Problème des bords: Maximale complexite forte. Lecture Notes in Mathematics, 1198:196–205, 1986. CODEN LN-MAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077054/.

Pomerance:1985:DRN

[Pom85] Carl Pomerance. On the distribution of round numbers. Lecture Notes in Mathematics, 1122:173–200, 1985. CODEN LN-MAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075761/.

Pommerening:1987:IUG

[Pom87] Klaus Pommerening. Invariants of unipotent groups. Lecture Notes in Mathematics, 1278:8–17, 1987. CODEN LN-MAA2. ISBN 3-540-18360-4 (print), 3-540-47908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078802/;   http:
//link.springer.com/chapter/10.1007/BFb0078803/.

Ponce:1988:IVP

[Pon88] Gustavo Ponce. The initial value problem for Euler and Navier– Stokes equations in L_{s}^{p}(R^{2}). Lecture Notes in Mathematics, 1324: 245–252, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100796/.

Popov:1987:IBE

[Pop87a] V. A. Popov. On the inequalities of Berry–Esseen and V. M. Zolotarev. Lecture Notes in Mathematics, 1233:114–124, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072717/.

Popov:1987:A

[Pop87b] V. L. Popov. On actions of G_{a} on A^{n}. Lecture Notes in Mathematics, 1271:237–242, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079241/.

Porru:1985:CMV

[Por85] Giovanni Porru. On the continuity of the minima of variational integrals in Orlicz–Sobolev spaces. Lecture Notes in Mathematics, 1165:253–260, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076158/.

Pourcin:1988:DSH

[Pou88] Geneviève Pourcin. Deformations of singular holomorphic foliations on reduced compact C-analytic spaces. Lecture Notes in Mathematics, 1345:246–255, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081405/.

Povoas:1986:SSH

[Pov86] M. Povoas. On some singular hyperbolic evolution equations in Hilbert spaces. Lecture Notes in Mathematics, 1223:208–215, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099194/.

Pozo:1988:TIO

[Poz88] M. A. Jiménez Pozo. On the trajectories of inclined oil-wells. Lecture Notes in Mathematics, 1354:164–168, 1988. CODEN LN-MAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089593/.

Palagallo-Price:1987:PPO

[PPP87] Judith Palagallo-Price and Thomas E. Price. Properties of projections obtained by averaging certain polynomial interpolants. Lecture Notes in Mathematics, 1287:132–145, 1987. CODEN LN-MAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078902/.

Palis:1985:TIN

[PR85a] J. Palis and R. Roussarie. Topological invariants as numbers of translation. Lecture Notes in Mathematics, 1125:64–86, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075635/.

Peitgen:1985:MSM

[PR85b] Heinz-Otto Peitgen and Peter H. Richter. The Mandelbrot set in a model for phase transitions. Lecture Notes in Mathematics, 1111:111–134, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084587/.

Pino:1985:DUE

[PR85c] Ramon Pino and Jean-Pierre Ressayre. Definable ultrafilters and elementary end extensions. Lecture Notes in Mathematics, 1130: 341–350, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075317/.

Park:1989:CHP

[PR89a] Jae Keol Park and Klaus W. Roggenkamp. A characterization of hereditary P.I. rings. Lecture Notes in Mathematics, 1398:140– 145, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086250/.

Poffald:1989:DI

[PR89b] Esteban I. Poffald and Simeon Reich. A difference inclusion. Lecture Notes in Mathematics, 1394:122–130, 1989. CODEN LN-MAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086757/.

Praagman:1985:ILA

[Pra85] C. Praagman. Iterations and logarithms of automorphisms of complete local rings. Lecture Notes in Mathematics, 1163:131– 134, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076426/.

Pratelli:1988:ISG

[Pra88] Maurizio Pratelli. Intégration stochastique et géométrie des espaces de Banach. (French) []. Lecture Notes in Mathematics, 1321:129–137, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084127/.

Prestel:1985:APF

[Pre85] Alexander Prestel. On the axiomatization of PRC-fields. Lecture Notes in Mathematics, 1130:351–359, 1985. CODEN LN-MAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075318/.

Presutti:1987:CPS

[Pre87] Errico Presutti. Collective phenomena in stochastic particle systems. Lecture Notes in Mathematics, 1250:195–232, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077359/.

Preston:1988:CCI

[Pre88a] Chris Preston. Countable closed invariant sets. Lecture Notes in Mathematics, 1347:103–111, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079778/.

Preston:1988:ERS

[Pre88b] Chris Preston. Examples of register-shifts. Lecture Notes in Mathematics, 1347:50–56, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079774/.

Preston:1988:E

[Pre88c] Chris Preston. Extensions. Lecture Notes in Mathematics, 1347: 112–125, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079779/.

Preston:1988:I

[Pre88d] Chris Preston. Introduction. Lecture Notes in Mathematics, 1347: 1–19, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079770/.

Preston:1988:MOT

[Pre88e] Chris Preston. Mappings with one turning point. Lecture Notes in Mathematics, 1347:135–152, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079781/.

Preston:1988:PMM

[Pre88f] Chris Preston. Piecewise monotone mappings. Lecture Notes in Mathematics, 1347:20–33, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079771/.

Preston:1988:PPT

[Pre88g] Chris Preston. A proof of Parry’s theorem (theorem 2.6). Lecture Notes in Mathematics, 1347:57–67, 1988. CODEN LN-MAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079775/.

Preston:1988:PT

[Pre88h] Chris Preston. Proof of theorems 2.4 and 2.5. Lecture Notes in Mathematics, 1347:34–40, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079772/.

Preston:1988:Ra

[Pre88i] Chris Preston. Reductions. Lecture Notes in Mathematics, 1347: 68–81, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079776/.

Preston:1988:Rb

[Pre88j] Chris Preston. Refinements. Lecture Notes in Mathematics, 1347: 126–134, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079780/.

Preston:1988:SH

[Pre88k] Chris Preston. Sinks and homtervals. Lecture Notes in Mathematics, 1347:41–49, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079773/.

Preston:1988:SMR

[Pre88l] Chris Preston. Some miscellaneous results from real analysis. Lecture Notes in Mathematics, 1347:153–159, 1988. CODEN LN-MAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079782/.

Preston:1988:SS

[Pre88m] Chris Preston. The structure of the set D(f). Lecture Notes in Mathematics, 1347:82–102, 1988. CODEN LNMAA2. ISBN 3-540-50329-3 (print), 3-540-45971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079777/.

Pride:1989:SPC

[Pri89] Stephen J. Pride. Some problems in combinatorial group theory. Lecture Notes in Mathematics, 1398:146–155, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086251/.

Prodinger:1985:BGP

[Pro85] Helmut Prodinger. Die Bestimmung Gewisser Parameter Bei Binären Bäumen Mit Hilfe Analytischer Methoden. (German) []. Lecture Notes in Mathematics, 1114:118–133, 1985. CODEN LN-MAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0101650/.

Prodinger:1987:LTF

[Pro87] Helmut Prodinger. Über Längste 1-Teilfolgen In 0–1-Folgen. (German) [On the longest 1-part sequence in 0–1-sequences]. Lecture Notes in Mathematics, 1262:124–133, 1987. CODEN LN-MAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078601/.

Protter:1989:CBE

[Pro89] Philip Protter. A connection between the expansion of filtrations and Girsanov’s theorem. Lecture Notes in Mathematics, 1390:221–224, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083949/.

Pruss:1986:PST

[Prü86] Jan Prüss. Periodic solutions of the thermostat problem. Lecture Notes in Mathematics, 1223:216–226, 1986. CODEN LN-MAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099195/.

Pigozzi:1985:HPC

[PS85] Don Pigozzi and J. Sichler. Homomorphisms of partial and of complete Steiner triple systems and quasigroups. Lecture Notes in Mathematics, 1149:224–237, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098467/.

Pan:1987:SSR

[PS87a] Yanglian Pan and Yibin Shen. Some stability results of harmonic map from a manifold with boundary. Lecture Notes in Mathematics, 1255:80–87, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077683/.

Piatetski-Shapiro:1987:ITK

[PS87b] I. Piatetski-Shapiro. Invariant theory and Kloosterman sums. Lecture Notes in Mathematics, 1271:229–236, 1987. CODEN LN-MAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079240/.

Pop-Stojanovic:1987:EMP

[PS87c] Z. R. Pop-Stojanović. Energy in Markov processes. Lecture Notes in Mathematics, 1242:374–395, 1987. CODEN LNMAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072445/.

Parthasarathy:1988:RCQ

[PS88a] K. R. Parthasarathy and Kalyan B. Sinha. Representation of a class of quantum martingales II. Lecture Notes in Mathematics, 1303:232–250, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078066/.

Pesin:1988:SMC

[PS88b] Ya. B. Pesin and Ya. G. Sinaĭ. On stable manifolds for class of two-dimensional diffeomorphisms. Lecture Notes in Mathematics, 1346:113–125, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082774/.

Pontier:1988:USF

[PS88c] Monique Pontier and Jacques Szpirglas. Uniqueness of the solutions of the filtering equation with observations on a Riemannian symmetric space. Lecture Notes in Mathematics, 1316:328– 339, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081938/.

Peters:1989:PSR

[PS89a] C. Peters and J. Stienstra. A pencil of K3-surfaces related to Apéry’s recurrence for ζ(3) and Fermi surfaces for potential zero.

Lecture Notes in Mathematics, 1399:110–127, 1989. CODEN LN-MAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0095972/.

Piotrowski:1989:CCG

[PS89b] Wiktor Piotrowski and Maciej M. Sysło. A characterization of centroidal graphs. Lecture Notes in Mathematics, 1403:272–281, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083471/.

Piatetski-Shapiro:1987:FCG

[PSR87] I. Piatetski-Shapiro and S. Rallis. L-functions for the classical groups. Lecture Notes in Mathematics, 1254:1–52, 1987. CODEN LNMAA2. ISBN 3-540-17848-1 (print), 3-540-47880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078126/.

Pontier:1986:TRP

[PSS86] M. Pontier, C. Stricker, and J. Szpirglas. Sur le théorème de représentation par rapport à l’innovation. (French) []. Lecture Notes in Mathematics, 1204:34–39, 1986. CODEN LN-MAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075709/.

Puerta:1988:AHC

[Pue88] F. Puerta. Applications des hyperrésolutions cubiques à la théorie de Hodge. (French) []. Lecture Notes in Mathematics, 1335:87– 132, 1988. CODEN LNMAA2. ISBN 3-540-50023-5 (print), 3-540-69984-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085058/.

Pugliese:1986:SQI

[Pug86] Andrea Pugliese. Some questions on the integrodifferential equation u′ = AK ∗ u + BM ∗ u. Lecture Notes in Mathematics, 1223:227–242, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099196/.

Puppe:1988:SCM

[Pup88] V. Puppe. Simply connected manifolds without S^{1}-symmetry. Lecture Notes in Mathematics, 1361:261–268, 1988. CODEN LN-MAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083039/.

Peterson:1987:RRP

[PV87] Dale Peterson and Michèle Vergne. Recurrence relations for Plancherel functions. Lecture Notes in Mathematics, 1243:240– 261, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073025/.

Palis:1988:CHD

[PV88] J. Palis and M. Viana. On the continuity of Hausdorff dimension and limit capacity for horseshoes. Lecture Notes in Mathematics, 1331:150–160, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083071/.

Paris:1985:CPB

[PW85a] J. Paris and A. Wilkie. Counting problems in bounded arithmetic. Lecture Notes in Mathematics, 1130:317–340, 1985. CODEN LN-MAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075316/.

Pedersen:1985:NDA

[PW85b] Erik K. Pedersen and Charles A. Weibel. A nonconnective delooping of algebraic K-theory. Lecture Notes in Mathematics, 1126:166–181, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074443/.

Pragacz:1986:CRD

[PW86] Piotr Pragacz and Jerzy Weyman. On the construction of resolutions of determinantal ideals: a survey. Lecture Notes in Mathematics, 1220:73–92, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099506/.

Pisier:1987:RSR

[PX87] Gilles Pisier and Quanhua Xu. Random series in the real interpolation spaces between the spaces v_{p}. Lecture Notes in Mathematics, 1267:185–209, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078146/.

Quaegebeur:1988:MQV

[Qua88a] Johan Quaegebeur. Mutual quadratic variation and Itô’s table in quantum stochastic calculus. Lecture Notes in Mathematics, 1303:261–274, 1988. CODEN LNMAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078068/.

Quasthoff:1988:MPA

[Qua88b] Uwe Quasthoff. On mixing properties of automorphisms of von Neumann algebras related to measure space transformations. Lecture Notes in Mathematics, 1303:275–284, 1988. CODEN LN-MAA2. ISBN 3-540-18919-X (print), 3-540-38846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078069/.

Queffelec:1987:A

[Que87a] Martine Queffélec. The algebra M(T). Lecture Notes in Mathematics, 1294:1–14, 1987. CODEN LNMAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081891/.

Queffelec:1987:BAS

[Que87b] Martine Queffélec. Bijective automata or substitutions. Lecture Notes in Mathematics, 1294:174–189, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081899/.

Queffelec:1987:DSAb

[Que87c] Martine Queffélec. Dynamical systems arising from substitutions. Lecture Notes in Mathematics, 1294:87–109, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081895/.

Queffelec:1987:DSAa

[Que87d] Martine Queffélec. Dynamical systems associated to sequences. Lecture Notes in Mathematics, 1294:60–86, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081894/.

Queffelec:1987:ESD

[Que87e] Martine Queffélec. Eigenvalues of substitution dynamical system. Lecture Notes in Mathematics, 1294:110–143, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081896/.

Queffelec:1987:MM

[Que87f] Martine Queffélec. Matrices of measures. Lecture Notes in Mathematics, 1294:144–159, 1987. CODEN LNMAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081897/.

Queffelec:1987:MRP

[Que87g] Martine Queffélec. Matrix Riesz products. Lecture Notes in Mathematics, 1294:160–173, 1987. CODEN LNMAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081898/.

Queffelec:1987:MST

[Que87h] Martine Queffélec. Maximal spectral type of general automata. Lecture Notes in Mathematics, 1294:190–209, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081900/.

Queffelec:1987:SMG

[Que87i] Martine Queffélec. Spectral multiplicity of general automata. Lecture Notes in Mathematics, 1294:210–224, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081901/.

Queffelec:1987:STD

[Que87j] Martine Queffélec. Spectral theory of dynamical systems. Lecture Notes in Mathematics, 1294:35–59, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081893/.

Queffelec:1987:STU

[Que87k] Martine Queffélec. Spectral theory of unitary operators. Lecture Notes in Mathematics, 1294:15–34, 1987. CODEN LN-MAA2. ISBN 3-540-18692-1 (print), 3-540-48088-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081892/.

Quinn:1985:GA

[Qui85a] Frank Quinn. Geometric algebra. Lecture Notes in Mathematics, 1126:182–198, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074444/.

Quinn:1985:HC

[Qui85b] Frank Quinn. Handlebodies and 2-complexes. Lecture Notes in Mathematics, 1167:245–259, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075228/.

Wang:1988:EMT

[qW88] Zhi qiang Wang. Equivariant Morse theory for isolated critical orbits and its applications to nonlinear problems. Lecture Notes in Mathematics, 1306:202–223, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082935/.

Wu:1989:JPC

[qW89] Ying qing Wu. Jones polynomial and the crossing number of links. Lecture Notes in Mathematics, 1369:286–288, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087541/.

Raby:1987:PCF

[Rab87] Gilles Raby. Paramétrix, cohomologie et formes méromorphes. (French) []. Lecture Notes in Mathematics, 1295:167–225, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081983/.

Rabinowitz:1988:PSP

[Rab88] Paul H. Rabinowitz. Periodic solutions of prescribed energy of Hamiltonian systems. Lecture Notes in Mathematics, 1324:253– 262, 1988. CODEN LNMAA2. ISBN 3-540-50111-8 (print), 3-540-45928-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100797/.

Rachev:1985:EFS

[Rac85] S. T. Rachev. Extreme functionals in the space of probability measures. Lecture Notes in Mathematics, 1155:320–348, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074826/.

Racke:1988:IBV

[Rac88] Reinhard Racke. Initial boundary value problems in thermoelasticity. Lecture Notes in Mathematics, 1357:341–358, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082874/.

Radu:1987:SFP

[Rad87] Viorel Radu. Some fixed point theorems probabilistic metric spaces. Lecture Notes in Mathematics, 1233:125–133, 1987. CO-DEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0072718/.

Rallis:1987:GT

[Ral87a] Stephen Rallis. Global theory. Lecture Notes in Mathematics, 1245:174–199, 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077901/.

Rallis:1987:IPF

[Ral87b] Stephen Rallis. Inner product formulae. Lecture Notes in Mathematics, 1245:49–86, 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077898/.

Rallis:1987:FOR

[Ral87c] Stephen Rallis. L-Functions and the Oscillator Representation, volume 1245 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 239 pp. LCCN QA3 .L28 no. 1245. URL http:

//link.springer.com/book/10.1007/BFb0077894;   http://
www.springerlink.com/content/978-3-540-47761-7.

Rallis:1987:LF

[Ral87d] Stephen Rallis. Local l-factors. Lecture Notes in Mathematics, 1245:128–173, 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077900/.

Rallis:1987:NP

[Ral87e] Stephen Rallis. Notation and preliminaries. Lecture Notes in Mathematics, 1245:1–9, 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077895/.

Rallis:1987:SFC

[Ral87f] Stephen Rallis. Siegel formula — compact case. Lecture Notes in Mathematics, 1245:87–127, 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077899/.

Rallis:1987:SFR

[Ral87g] Stephen Rallis. Siegel formula revisited. Lecture Notes in Mathematics, 1245:25–48, 1987. CODEN LNMAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077897/.

Rallis:1987:SES

[Ral87h] Stephen Rallis. Special Eisenstein series on orthogonal groups. Lecture Notes in Mathematics, 1245:10–24, 1987. CODEN LN-MAA2. ISBN 3-540-17694-2 (print), 3-540-47761-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077896/.

Ranicki:1985:ATT

[Ran85] Andrew Ranicki. The algebraic theory of torsion I. Foundations. Lecture Notes in Mathematics, 1126:199–237, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074445/.

Ranga:1986:FSM

[Ran86] A. Sri Ranga. Ĵ-fractions and strong moment problems. Lecture Notes in Mathematics, 1199:269–284, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075942/.

Range:1987:IRT

[Ran87] R. Michael Range. Integral representations in the theory of the ¯∂-Neumann problem. Lecture Notes in Mathematics, 1276:281– 290, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078964/.

Randall:1988:ITF

[Ran88] Duane Randall. On indices of tangent fields with finite singularities. Lecture Notes in Mathematics, 1350:213–240, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081477/.

Ran:1989:DM

[Ran89] Ziv Ran. Deformations of maps. Lecture Notes in Mathematics, 1389:246–253, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085936/.

Rao:1987:FVB

[Rao87] A. Prabhakar Rao. A family of vector bundles on P^{3}. Lecture Notes in Mathematics, 1266:208–231, 1987. CODEN LN-MAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078185/.

Rao:1989:BHP

[Rao89] M. M. Rao. Bimeasures and harmonizable processes (analysis, classification, and representation). Lecture Notes in Mathematics, 1379:254–298, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087858/.

Rascle:1987:TEG

[Ras87] Michel Rascle. Un théorème d’existence globale en élasticité non linéaire mono-dimensionnelle. (French) []. Lecture Notes in Mathematics, 1270:210–220, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078330/.

Rauszer:1985:FCI

[Rau85] Cecylia Rauszer. Formalizations of certain intermediate logics. Part I. Lecture Notes in Mathematics, 1130:360–384, 1985. CO-DEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075319/.

Raviart:1985:APM

[Rav85] P.-A. Raviart. An analysis of particle methods. Lecture Notes in Mathematics, 1127:243–324, 1985. CODEN LNMAA2. ISBN 3-540-15225-3 (print), 3-540-39391-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074532/.

Ravenel:1988:CCT

[Rav88] Douglas C. Ravenel. Complex cobordism theory for number theorists. Lecture Notes in Mathematics, 1326:123–133, 1988. CODEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078042/.

Rawnsley:1985:STS

[Raw85] John H. Rawnsley. F -structures, F -twistor spaces and harmonic maps. Lecture Notes in Mathematics, 1164:85–159, 1985. CODEN LNMAA2. ISBN 3-540-16048-5 (print), 3-540-39732-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081911/.

Rawnsley:1987:TM

[Raw87] John Rawnsley. Twistor methods. Lecture Notes in Mathematics, 1263:97–133, 1987. CODEN LNMAA2. ISBN 3-540-18012-5 (print), 3-540-47249-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078611/.

Raynaud:1988:EAS

[Ray88] Yves Raynaud. Extraction of almost symmetric subsequences in r.i. spaces. Lecture Notes in Mathematics, 1332:33–42, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081609/.

Reed:1987:RHS

[RB87] M. C. Reed and J. J. Blum. A reaction-hyperbolic system in physiology. Lecture Notes in Mathematics, 1248:134–141, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077422/.

Reinhardt:1985:TCS

[RC85] William N. Reinhardt and Rolando Chuaqui. Types in class set theories. Lecture Notes in Mathematics, 1130:385–394, 1985. CO-DEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075320/.

Randjbar-Daemi:1985:SCF

[RD85] S. Randjbar-Daemi. Spontaneous compactification and fermion chirality. Lecture Notes in Mathematics, 1139:199–213, 1985. CO-DEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074585/.

Rea:1986:HEP

[Rea86] Claudio Rea. Holomorphic extendability past boundary points and flat functions in C^{2} and C^{n}. Lecture Notes in Mathematics, 1198:206–213, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077055/.

Read:1988:ISP

[Rea88] C. J. Read. The invariant subspace problem on a class of nonreflexive Banach spaces, 1. Lecture Notes in Mathematics, 1317: 1–20, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081733/.

Reckziegel:1985:HLI

[Rec85] Helmut Reckziegel. Horizontal lifts of isometric immersions into the bundle space of a pseudo-Riemannian submersion. Lecture Notes in Mathematics, 1156:264–279, 1985. CODEN LN-MAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075098/.

Reczek:1987:HAM

[Rec87] K. Reczek. Hyperbolic approximation of meromorphic functions. Lecture Notes in Mathematics, 1237:73–78, 1987. CODEN LN-MAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072457/.

Rego:1987:DMC

[Reg87] C. J. Rego. Deformation of modules on curves and surfaces. Lecture Notes in Mathematics, 1273:157–167, 1987. CODEN LN-MAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078842/.

Regev:1988:CMA

[Reg88] Amitai Regev. On the codimensions of matrix algebras. Lecture Notes in Mathematics, 1352:162–172, 1988. CODEN LN-MAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082024/.

Reich:1985:DEA

[Rei85] L. Reich. On a differential equation arising in iteration theory in rings of formal power series in one variable. Lecture Notes in Mathematics, 1163:135–148, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076427/.

Reiten:1986:CMM

[Rei86] Idun Reiten. Cohen–Macaulay modules over isolated singularities. Lecture Notes in Mathematics, 1220:25–32, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099503/.

Reichel:1987:BCF

[Rei87a] Lothar Reichel. Boundary collocation in Fejér points for computing eigenvalues and eigenfunctions of the Laplacian. Lecture Notes in Mathematics, 1287:146–160, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078903/.

Reiten:1987:FDA

[Rei87b] Idun Reiten. Finite dimensional algebras and singularities. Lecture Notes in Mathematics, 1273:35–57, 1987. CODEN LN-MAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078837/.

Reiffen:1988:LI

[Rei88a] Hans-Jörg Reiffen. Leafspaces and integrability. Lecture Notes in Mathematics, 1345:271–293, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081407/.

Reisner:1988:TTL

[Rei88b] Shlomo Reisner. On two theorems of Lozanovskii concerning intermediate Banach lattices. Lecture Notes in Mathematics, 1317: 67–83, 1988. CODEN LNMAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081736/.

Reiter:1988:FM

[Rei88c] H. Reiter. Front matter. Lecture Notes in Mathematics, 1361: ??, 1988. CODEN LNMAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46036-7/1.pdf.

Reiter:1988:MGH

[Rei88d] H. Reiter. Metaplectic groups and harmonic analysis. Lecture Notes in Mathematics, 1359:238–241, 1988. CODEN LN-MAA2. ISBN 3-540-50524-5 (print), 3-540-46032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086604/.

Reiter:1989:MGS

[Rei89a] Hans Reiter. The metaplectic group and Segal continuity in the adelic case. Lecture Notes in Mathematics, 1382:98–116, 1989. CODEN LNMAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093692/.

Reiter:1989:MGL

[Rei89b] Hans Reiter. The metaplectic group (local case); Segal continuity. Lecture Notes in Mathematics, 1382:91–97, 1989. CODEN LNMAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093691/.

Reiter:1989:PHA

[Rei89c] Hans Reiter. Preliminaries from harmonic analysis and group theory. Lecture Notes in Mathematics, 1382:1–15, 1989. CODEN LNMAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093684/.

Reiter:1989:PCQ

[Rei89d] Hans Reiter. Properties of certain quadratic forms. Lecture Notes in Mathematics, 1382:70–80, 1989. CODEN LNMAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093689/.

Reiter:1989:SAS

[Rei89e] Hans Reiter. Segal algebras; the Segal algebra G^{1}(G). Lecture Notes in Mathematics, 1382:16–35, 1989. CODEN LN-MAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093685/.

Reiter:1989:VSQ

[Rei89f] Hans Reiter. Vector spaces and quadratic forms ever local fields. Lecture Notes in Mathematics, 1382:62–69, 1989. CODEN LN-MAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093688/.

Reiter:1989:WOV

[Rei89g] Hans Reiter. Weil operators for vector spaces over local fields. Lecture Notes in Mathematics, 1382:81–90, 1989. CODEN LN-MAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093690/.

Reiter:1989:WGO

[Rei89h] Hans Reiter. Weil’s group of operators and related groups. Lecture Notes in Mathematics, 1382:43–61, 1989. CODEN LN-MAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0093687/.

Reiter:1989:WT

[Rei89i] Hans Reiter. Weil’s theorem 6. Lecture Notes in Mathematics, 1382:117–118, 1989. CODEN LNMAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093693.

Reiter:1989:WUO

[Rei89j] Hans Reiter. Weil’s unitary operators and the Segal algebra G^{1}(G). Lecture Notes in Mathematics, 1382:36–42, 1989. CODEN LNMAA2. ISBN 3-540-51417-1 (print), 3-540-46210-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093686/.

Rektorys:1986:NTT

[Rek86] K. Rektorys. Numerical and theoretical treating of evolution problems by the method of discretization in time. Lecture Notes in Mathematics, 1192:71–84, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076053/.

Ressel:1989:CCM

[Res89] Paul Ressel. A conjecture concerning mixtures of characters from a given closed subsemigroup in the dual. Lecture Notes in Mathematics, 1379:299–309, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087859/.

Reula:1986:EAL

[Reu86a] Francisco Javier Echarte Reula. Étude des algèbres de Lie resolubles réelles qui admettent des idéaux unidimensionnels n’appartenant pas au centre. (French) []. Lecture Notes in Mathematics, 1209:152–156, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076627/.

Reutenauer:1986:TPB

[Reu86b] Christophe Reutenauer. Theorem of Poincarè–Birkhoff–Witt, logarithm and symmetric group representations of degrees equal to Stirling numbers. Lecture Notes in Mathematics, 1234:267– 284, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072520/.

Reyes:1986:SRV

[Rey86a] Gonzalo E. Reyes. Synthetic reasoning and variable sets. Lecture Notes in Mathematics, 1174:69–82, 1986. CODEN LN-MAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076935/.

Reynolds:1986:PIC

[Rey86b] William F. Reynolds. Principal indecomposable characters and normal subgroups. Lecture Notes in Mathematics, 1178:143– 151, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075294/.

Rhemtulla:1989:GME

[Rhe89] Akbar H. Rhemtulla. Groups with many elliptic subgroups. Lecture Notes in Mathematics, 1398:156–162, 1989. CODEN LN-MAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086252/.

Richter:1986:DRT

[Ric86] F. Richter. On the K-dimensional Radon-transform of rapidly decreasing functions. Lecture Notes in Mathematics, 1209:243– 258, 1986. CODEN LNMAA2. ISBN 3-540-16801-X (print), 3-540-44844-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076636/.

Richardson:1987:NSS

[Ric87] R. W. Richardson. Normality of G-stable subvarieties of a semisimple Lie algebra. Lecture Notes in Mathematics, 1271:243– 264, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079242/.

Riedel:1987:ABD

[Rie87] M. Riedel. The asymptotic bias in a deviation of a location model. Lecture Notes in Mathematics, 1233:134–144, 1987. CODEN LN-MAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072719/.

Riley:1989:GOH

[Ril89] Robert Riley. Growth of order of homology of cyclic branched covers of knots. Lecture Notes in Mathematics, 1383:140–145, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083575/.

Rippon:1987:SFU

[Rip87] P. J. Rippon. Subharmonic functions with unilateral growth conditions. Lecture Notes in Mathematics, 1275:253–265, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078357/.

Rivlin:1988:GQA

[Riv88] T. J. Rivlin. Gauss quadrature for analytic functions. Lecture Notes in Mathematics, 1329:178–192, 1988. CODEN LN-MAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083358/.

Ranicki:1985:AGT

[RLQ85] Andrew Ranicki, Norman Levitt, and Frank Quinn, editors. Algebraic and Geometric Topology: Proceedings of a Conference held at Rutgers University, New Brunswick, USA July 6–13, 1983, volume 1126 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1126. URL http:

//link.springer.com/book/10.1007/BFb0074435;   http://
www.springerlink.com/content/978-3-540-39413-6.

Ruiperez:1987:VBP

[RM87] Daniel Hernández Ruipérez and Jaime Muñoz Masqué. Variational Berezinian problems and their relationship with graded variational problems. Lecture Notes in Mathematics, 1251:137–149, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077319/.

Roach:1988:IIT

[Roa88] G. F. Roach. An introduction to iterative techniques for potential problems. Lecture Notes in Mathematics, 1344:162–180, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103350/.

Robinson:1986:CMF

[Rob86] Geoffrey R. Robinson. On Cartan matrices of finite groups with a certain factorization. Lecture Notes in Mathematics, 1178:152– 176, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075295/.

Robinson:1987:VTC

[Rob87] Derek J. S. Robinson. Vanishing theorems for cohomology of locally nilpotent groups. Lecture Notes in Mathematics, 1281:120– 129, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078698/.

Roberts:1988:ONR

[Rob88] Joel Roberts. Old and new results about the triangle varieties. Lecture Notes in Mathematics, 1311:197–219, 1988. CODEN LN-MAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082915/.

Robinson:1989:RCT

[Rob89] Derek J. S. Robinson. Reflections on the constructive theory of polycyclic groups. Lecture Notes in Mathematics, 1398:163– 169, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086253/.

Rochberg:1988:WCS

[Roc88] Richard Rochberg. The work of Coifman and Semmes on complex interpolation, several complex variables, and PDE’s. Lecture Notes in Mathematics, 1302:74–90, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078864/.

Rodin:1987:DRR

[Rod87a] Yuri L. Rodin. The Dolbeault and Riemann–Roch theorems. Lecture Notes in Mathematics, 1288:1–28, 1987. CODEN LN-MAA2. ISBN 3-540-18572-0 (print), 3-540-48018-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082104/.

Rodin:1987:LIE

[Rod87b] Yuri L. Rodin. Linear integral equations connected with generalized analytic functions. Lecture Notes in Mathematics, 1288: 29–62, 1987. CODEN LNMAA2. ISBN 3-540-18572-0 (print), 3-540-48018-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082105/.

Rodin:1987:NAG

[Rod87c] Yuri L. Rodin. Nonlinear aspects of generalized analytic function theory. Lecture Notes in Mathematics, 1288:80–85, 1987. CODEN LNMAA2. ISBN 3-540-18572-0 (print), 3-540-48018-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082107/.

Rodin:1987:RBP

[Rod87d] Yuri L. Rodin. The Riemann boundary problem. Lecture Notes in Mathematics, 1288:63–79, 1987. CODEN LNMAA2. ISBN 3-540-18572-0 (print), 3-540-48018-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082106/.

Rodin:1987:SGA

[Rod87e] Yuri L. Rodin. Some generalizations and applications. Lecture Notes in Mathematics, 1288:86–108, 1987. CODEN LN-MAA2. ISBN 3-540-18572-0 (print), 3-540-48018-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082108/.

Roe:1987:UDS

[Roe87] P. L. Roe. Upwind differencing schemes for hyperbolic conservation laws with source terms. Lecture Notes in Mathematics, 1270: 41–51, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078316/.

Rofman:1985:AHJ

[Rof85] Edmundo Rofman. Approximation of Hamilton–Jacobi–Bellman equation in deterministic control theory. An application to energy production systems. Lecture Notes in Mathematics, 1119: 152–189, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074785/.

Rogers:1985:SES

[Rog85] Alice Rogers. Off-shell extended supergravity in extended superspace. Lecture Notes in Mathematics, 1139:214–235, 1985. CO-DEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074586/.

Roggenkamp:1986:GOF

[Rog86] K. W. Roggenkamp. Gorenstein orders of finite representation type and bijective lattices. Lecture Notes in Mathematics, 1178: 243–271, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075298/.

Roggenkamp:1987:ASS

[Rog87] K. W. Roggenkamp. Almost split sequences for some nonclassical lattice categories. Lecture Notes in Mathematics, 1273: 318–324, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078851/.

Roggenkamp:1988:SNP

[Rog88] K. W. Roggenkamp. Some new progress on the isomorphismproblem for integral group rings. Lecture Notes in Mathematics, 1328:227–236, 1988. CODEN LNMAA2. ISBN 3-540-19474-6 (print), 3-540-39278-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100928/.

Rogers:1989:MPM

[Rog89] L. C. G. Rogers. Multiple points of Markov processes in a complete metric space. Lecture Notes in Mathematics, 1372:186–197, 1989. CODEN LNMAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083973/.

Roiter:1986:NR

[Roĭ86] A. V. Roĭter. The norm of a relation. Lecture Notes in Mathematics, 1177:269–271, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075269/.

Rolfsen:1985:KTM

[Rol85] Dale Rolfsen, editor. Knot Theory and Manifolds: Proceedings of a Conference held in Vancouver, Canada, June 2–4, 1983, volume 1144 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1144. URL http:

//link.springer.com/book/10.1007/BFb0075007;   http://
www.springerlink.com/content/978-3-540-39616-1.

Ronan:1986:CBN

[Ron86] M. A. Ronan. A construction of buildings with no rank 3 residues of spherical type. Lecture Notes in Mathematics, 1181: 242–248, 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075518/.

Roos:1986:FPV

[Roo86a] Guy Roos. Fonctions de plusieurs variables complexes et formules de représentation intégrale. (French) []. Lecture Notes in Mathematics, 1188:45–182, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076820/.

Roos:1986:AAT

[Roo86b] Jan-Erik Roos, editor. Algebra, Algebraic Topology and their Interactions: Proceedings of a Conference held in Stockholm, Aug. 3–13, 1983, and later developments, volume 1183 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-16453-7 (print), 3-540-39790-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1183. URL http://link.springer.com/book/10.1007/
BFb0075445;   http://www.springerlink.com/content/978-
3-540-39790-8.

Roos:1988:HFL

[Roo88] Jan-Erik Roos. Homology of free loop spaces, cyclic homology and non-rational Poincaré–Betti series in commutative algebra. Lecture Notes in Mathematics, 1352:173–189, 1988. CODEN LN-MAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082025/.

Roos:1989:ACM

[Roo89] C. Roos. An O(n^{3}L) approximate center method for linear programming. Lecture Notes in Mathematics, 1405:147–158, 1989. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Rosenthal:1985:LAB

[Ros85a] Haskell Rosenthal. The Lie algebra of a Banach space. Lecture Notes in Mathematics, 1166:129–157, 1985. CODEN LN-MAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074702/.

Rosser:1985:CPS

[Ros85b] J. Barkley Rosser. Calculation of potential in a sector. Lecture Notes in Mathematics, 1121:228–257, 1985. CODEN LN-MAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076274/.

Rossler:1985:LLA

[Rös85c] O. E. Rössler. Long line attractors. Lecture Notes in Mathematics, 1163:149–160, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076428/.

Rosati:1986:BGD

[Ros86a] Luigi A. Rosati, editor. Buildings and the Geometry of Diagrams: Lectures given at the 3rd 1984 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Como, Italy, August 26– September 4, 1984, volume 1181 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1181. URL http:

//link.springer.com/book/10.1007/BFb0075511;   http://
www.springerlink.com/content/978-3-540-39801-1.

Rosenkranz:1986:SMD

[Ros86b] Gerd Rosenkranz. Stochastic measure diffusions as models of growth and spread. Lecture Notes in Mathematics, 1212:238– 248, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076252/.

Rost:1986:BHL

[Ros86c] Hermann Rost. On the behavior of the hydrodynamical limit for stochastic particle systems. Lecture Notes in Mathematics, 1215:129–164, 1986. CODEN LNMAA2. ISBN 3-540-16822-2 (print), 3-540-47092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075873/.

Rossberg:1987:CDT

[Ros87a] Hans-Joachim Rossberg. Cramér’s decomposition theorem within the continuation of distribution functions. Lecture Notes in Mathematics, 1233:145–155, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072720/.

Rossmann:1987:SRC

[Ros87b] W. Rossmann. Springer representations and coherent continuation representations of Weyl groups. Lecture Notes in Mathematics, 1243:273–282, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073027/.

Rosenthal:1988:C

[Ros88a] Haskell Rosenthal. L^{1}-convexity. Lecture Notes in Mathematics, 1332:156–174, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081618/.

Rosenthal:1988:CRT

[Ros88b] Haskell Rosenthal. On the Choquet representation theorem. Lecture Notes in Mathematics, 1332:1–32, 1988. CODEN LN-MAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081608/.

Rosales:1989:DEW

[Ros89a] Rodolfo R. Rosales. Diffraction effects in weakly nonlinear detonation waves. Lecture Notes in Mathematics, 1402:227–239, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083879/.

Rosales:1989:FM

[Ros89b] Rodolfo R. Rosales. Front matter. Lecture Notes in Mathematics, 1403:??, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-46810-3/1.pdf.

Rosenberger:1989:MGS

[Ros89c] Gerhard Rosenberger. Minimal generating systems for plane discontinuous groups and an equation in free groups. Lecture Notes in Mathematics, 1398:170–186, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086254/.

Roth:1989:PDN

[Rot89] Jean-Pierre Roth. Le problème du nuage de glace. (French) []. Lecture Notes in Mathematics, 1393:216–225, 1989. CODEN LN-MAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085782/.

Roux:1987:AVD

[Rou87] Bernard Roux. Anneaux de valuation discrète complêts, scindes, non commutatifs, en caractéristique zéro. (French) []. Lecture Notes in Mathematics, 1296:276–311, 1987. CODEN LN-MAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078533/.

Roussarie:1988:NFC

[Rou88] R. Roussarie. A note on finite cyclicity property and Hilbert’s 16th Problem. Lecture Notes in Mathematics, 1331:161–168, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083072/.

Rowen:1989:PIA

[Row89] Louis Halle Rowen. Primitive ideals of algebras over uncountable fields. Lecture Notes in Mathematics, 1404:322–345, 1989. CODEN LNMAA2. ISBN 3-540-51812-6 (print), 3-540-46814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084082/.

Rozanov:1987:BPS

[Roz87] Yu. A. Rozanov. Boundary problems for stochastic partial differential equations. Lecture Notes in Mathematics, 1250:233– 268, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077360/.

Reiner:1985:OTA

[RR85] Irving Reiner and Klaus W. Roggenkamp, editors. Orders and their Applications: Proceedings of a Conference held in Oberwolfach, West Germany June 3–9, 1984, volume 1142 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1142. URL http://link.springer.com/book/10.1007/
BFb0074788;   http://www.springerlink.com/content/978-
3-540-39601-7.

Rauch:1987:UNW

[RR87] Jeffrey Rauch and Michael C. Reed. Ultrasingularities in nonlinear waves. Lecture Notes in Mathematics, 1248:127–133, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077421/.

Romanowska:1985:APG

[RS85a] Anna B. Romanowska and Jonathan D. H. Smith. From affine to projective geometry via convexity. Lecture Notes in Mathematics, 1149:255–269, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098469/.

Ruess:1985:WDP

[RS85b] W. M. Ruess and C. P. Stegall. Weak*-denting points in duals of operator spaces. Lecture Notes in Mathematics, 1166:158– 168, 1985. CODEN LNMAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074703/.

Rohlfs:1987:CMD

[RS87] Jürgen Rohlfs and Birgit Speh. A cohomological method for the determination of limit multiplicities. Lecture Notes in Mathematics, 1243:262–272, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073026/.

Reichard:1988:PSQ

[RS88] Dr. K. Reichard and Prof. Dr. K. Spallek. Product singularities and quotients. Lecture Notes in Mathematics, 1345:256–270, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081406/.

Rochberg:1989:AOT

[RT89] Richard Rochberg and Mitchell Taibleson. An averaging operator on a tree. Lecture Notes in Mathematics, 1384:207–213, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086805/.

Rochberg:1988:IQS

[RTVW88] Richard Rochberg, Anita Tabacco Vignati, Marco Vignati, and Guido Weiss. Interpolation of quasinormed spaces by the complex method. Lecture Notes in Mathematics, 1302:91–98, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078865/.

RubiodeFrancia:1986:MIT

[Rub86] José L. Rubio de Francia. Martingale and integral transforms of Banach space valued functions. Lecture Notes in Mathematics, 1221:195–222, 1986. CODEN LNMAA2. ISBN 3-540-17186-X (print), 3-540-47344-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099115/.

Rubin:1989:WKA

[Rub89] Karl Rubin. The work of Kolyvagin on the arithmetic of elliptic curves. Lecture Notes in Mathematics, 1399:128–136, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095973/.

Ruggeri:1989:RET

[Rug89] Tommaso Ruggeri. Relativistic extended thermodynamics: General assumptions and mathematical procedure. Lecture Notes in Mathematics, 1385:269–277, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084034/.

Rump:1985:MVR

[Rum85] Wolfgang Rump. Module valuations and representations of completely reducible orders. Lecture Notes in Mathematics, 1142:233– 238, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074804/.

Rump:1986:ERT

[Rum86] Wolfgang Rump. Enlacements and representation theory of completely reducible orders. Lecture Notes in Mathematics, 1178:272– 308, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075299/.

Rumely:1989:A

[Rum89a] Robert S. Rumely. Applications. Lecture Notes in Mathematics, 1378:373–422, 1989. CODEN LNMAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084532/.

Rumely:1989:CD

[Rum89b] Robert S. Rumely. The canonical distance. Lecture Notes in Mathematics, 1378:56–132, 1989. CODEN LNMAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084528/.

Rumely:1989:F

[Rum89c] Robert S. Rumely. Foundations. Lecture Notes in Mathematics, 1378:23–55, 1989. CODEN LNMAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084527/.

Rumely:1989:GCT

[Rum89d] Robert S. Rumely. Global capacity theory. Lecture Notes in Mathematics, 1378:320–372, 1989. CODEN LNMAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084531/.

Rumely:1989:LCTa

[Rum89e] Robert S. Rumely. Local capacity theory — Archimedean case. Lecture Notes in Mathematics, 1378:133–184, 1989. CODEN LN-MAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084529/.

Rumely:1989:LCTb

[Rum89f] Robert S. Rumely. Local capacity theory — nonarchimedean case. Lecture Notes in Mathematics, 1378:185–319, 1989. CODEN LN-MAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084530/.

Rumely:1989:P

[Rum89g] Robert S. Rumely. Preliminaries. Lecture Notes in Mathematics, 1378:1–22, 1989. CODEN LNMAA2. ISBN 3-540-51410-4 (print), 3-540-46209-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084526/.

Rodriguez:1989:SEG

[RV89] María J. Rodríguez and Luis Vázquez. Spectrum estimations for the generalized quantum Henon–Heiles system. Lecture Notes in Mathematics, 1394:131–135, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086758/.

Ros:1985:PCM

[RVV85] Antonio Ros, Paul Verheyen, and Leopold Verstraelen. Positively curved minimal submanifolds. Lecture Notes in Mathematics, 1156:280–295, 1985. CODEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075099/.

Rosinski:1985:MIR

[RW85] J. Rosiński and W. A. Woyczyński. Moment inequalities for real and vector p-stable stochastic integrals. Lecture Notes in Mathematics, 1153:369–386, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074961/.

Llompart:1988:CCG

[RX88] F. Rosselló Llompart and S. Xambó Descamps. Computing Chow groups. Lecture Notes in Mathematics, 1311:220–234, 1988. CO-DEN LNMAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082916/.

Rude:1986:TSM

[RZ86] U. Rüde and Chr. Zenger. On the treatment of singularities in the multigrid method. Lecture Notes in Mathematics, 1228:261– 271, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072651/.

Rychlik:1987:AMB

[RZ87] Tomasz Rychlik and Ryszard Zieliński. An asymptotically most bias-robust invariant estimator of location. Lecture Notes in Mathematics, 1233:156–171, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072721/.

Sabri:1985:PIC

[Sab85] A. Sabri. Properties of invariant curves near a known invariant curve. Lecture Notes in Mathematics, 1163:161–169, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076429/.

Sacks:1985:SOQ

[Sac85] Gerald E. Sacks. Some open questions in recursion theory. Lecture Notes in Mathematics, 1141:333–342, 1985. CODEN LN-MAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076228/.

Sadovsky:1986:BND

[Sad86] Z. Sadovský. Bifurcations near a double eigenvalue of the rectangular plate problem with a domain parameter. Lecture Notes in Mathematics, 1192:415–420, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076102/.

Saerens:1987:ITS

[Sae87] Rita Saerens. Interpolation theory in C^{n}: a survey. Lecture Notes in Mathematics, 1268:158–188, 1987. CODEN LNMAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097302/.

Saff:1988:PCB

[Saf88] E. B. Saff. A principle of contamination in best polynomial approximation. Lecture Notes in Mathematics, 1354:79–97, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089584/.

Saito:1986:AAF

[Sai86] Yoshimi Saitō. Asymptotic and approximate formulas in the inverse scattering problem for the Schrödinger operator. Lecture Notes in Mathematics, 1218:190–200, 1986. CODEN LN-MAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073051/.

Sakamoto:1986:HI

[Sak86] Kunio Sakamoto. Helical immersions. Lecture Notes in Mathematics, 1201:230–241, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075659/.

Salamon:1985:HHM

[Sal85a] Simon Salamon. Harmonic and holomorphic maps. Lecture Notes in Mathematics, 1164:161–224, 1985. CODEN LNMAA2. ISBN 3-540-16048-5 (print), 3-540-39732-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081912/.

Salberger:1985:GDC

[Sal85b] Per Salberger. Galois descent and class groups of orders. Lecture Notes in Mathematics, 1142:239–255, 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074805/.

Salminen:1985:LTD

[Sal85c] P. Salminen. On local times of a diffusion. Lecture Notes in Mathematics, 1123:63–79, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075839/.

Salazar:1989:LGF

[Sal89] Jorge Salazar. Lignes de Green et frontière de R. S. Martin en quelques cas particuliers. (French) []. Lecture Notes in Mathematics, 1393:226–242, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085783/.

Sampson:1985:HMK

[Sam85] J. H. Sampson. Harmonic maps in Kähler geometry. Lecture Notes in Mathematics, 1161:193–205, 1985. CODEN LNMAA2. ISBN 3-540-16040-X (print), 3-540-39716-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075138/.

Sami:1986:DCB

[Sam86] K. Sami. Sur les domaines cerclés bornés de l’espace C_{o} (in). (French) []. Lecture Notes in Mathematics, 1198:214–232, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077056/.

Sanchez:1985:EEK

[Sán85a] Norma Sánchez. Einstein equations without Killing vectors, non-linear sigma models and self-dual Yang–Mills theory. Lecture Notes in Mathematics, 1139:280–299, 1985. CODEN LN-MAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074590/.

Sandling:1985:IPG

[San85b] Robert Sandling. The isomorphism problem for group rings: a survey. Lecture Notes in Mathematics, 1142:256–288, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074806/.

SanMartin:1986:ANT

[San86] L. San Martin. Addendum a note on transitivity of coupled control systems. Lecture Notes in Mathematics, 1186:126–128, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076836/.

Sano:1987:DSI

[San87] Shigeru Sano. Distributions sphériques invariantes sur l’espace symétrique semi-simple et son c-dual. (French) []. Lecture Notes in Mathematics, 1243:283–309, 1987. CODEN LNMAA2. ISBN 3-540-17701-9 (print), 3-540-47775-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073028/.

Sapronov:1986:DSS

[Sap86] Yu. I. Sapronov. The destruction of spherical symmetry in nonlinear variational problems. Lecture Notes in Mathematics, 1214: 249–275, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075969/.

Sapronov:1988:MST

[Sap88] Yu. I. Sapronov. Multidimensional sleeping tops. Lecture Notes in Mathematics, 1334:111–128, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080425/.

Sarnak:1985:NPC

[Sar85] Peter Sarnak. On the number of points on certain curves and an uncertainty principle. Lecture Notes in Mathematics, 1135: 239–253, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074608/.

Sarkozy:1989:HPN

[Sár89] A. Sárközy. Hybrid problems in number theory. Lecture Notes in Mathematics, 1383:146–169, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083576/.

Sato:1986:TBC

[Sat86] Hajime Sato. On topological Blaschke conjecture III. Lecture Notes in Mathematics, 1201:242–253, 1986. CODEN LN-MAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075660/.

Sanchez-Bringas:1988:SSG

[SB88] Federico Sánchez-Bringas. Structural stability of germs of vector fields on surfaces with a simple singularity. Lecture Notes in Mathematics, 1345:294–301, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081408/.

Sanders:1985:AIG

[SC85] J. A. Sanders and R. Cushman. Abelian integrals and global Hopf bifurcations. Lecture Notes in Mathematics, 1125:87–98, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075636/.

Schachermayer:1985:SRC

[Sch85a] W. Schachermayer. Some remarks concerning the Krein–Milman and the Radon–Nikodým property of Banach spaces. Lecture Notes in Mathematics, 1166:169–176, 1985. CODEN LN-MAA2. ISBN 3-540-16051-5 (print), 3-540-39736-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074704/.

Scharlemann:1985:MCS

[Sch85b] Martin Scharlemann. 3-manifolds with H_{2}(A, ∂A) = 0 and a conjecture of Stallings. Lecture Notes in Mathematics, 1144:138– 145, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075017/.

Schatz:1985:IAE

[Sch85c] A. H. Schatz. An introduction to the analysis of the error in the finite element method for second-order elliptic boundary value problems. Lecture Notes in Mathematics, 1129:94–139, 1985. CODEN LNMAA2. ISBN 3-540-15234-2 (print), 3-540-39412-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075158/.

Schmid:1985:RDR

[Sch85d] Wilfried Schmid. Recent developments in representation theory. Lecture Notes in Mathematics, 1111:135–153, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084588/.

Schnabl:1985:OFG

[Sch85e] R. Schnabl. Ober Eine C_{∞}-funktion. (German) []. Lecture Notes in Mathematics, 1114:134–142, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101651/.

Schnack:1985:SCG

[Sch85f] E. Schnack. Singularities of cracks with generalized finite elements. Lecture Notes in Mathematics, 1121:258–277, 1985. CODEN LN-MAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076275/.

Schoissengeier:1985:NWG

[Sch85g] Doz. Dr. J. Schoißengeier. Der Numerische Wert Gewisser Reihen. (German) []. Lecture Notes in Mathematics, 1114:143–147, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101652/.

Schoissengeier:1985:ODF

[Sch85h] Doz. Dr. J. Schoißengeier. Ober Die Diskrepanz Der Folgen (nα). (German) []. Lecture Notes in Mathematics, 1114:148–153, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101653/.

Schwaiger:1985:NFS

[Sch85i] Jens Schwaiger. Normal forms for systems of formal power series commuting in pairs and iteration problems. Lecture Notes in Mathematics, 1163:170–176, 1985. CODEN LNMAA2. ISBN 3-540-16067-1 (print), 3-540-39749-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076430/.

Schwartz:1985:CDD

[Sch85j] Laurent Schwartz. Construction directe d’une diffusion sur une variété. (French) []. Lecture Notes in Mathematics, 1123:91– 112, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075841/.

Schwenk:1985:AES

[Sch85k] Angela Schwenk. Affinsphären mit ebenen Schattengrenzen. (German) []. Lecture Notes in Mathematics, 1156:296–315, 1985. CO-DEN LNMAA2. ISBN 3-540-15994-0 (print), 3-540-39698-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075100/.

Schachermayer:1986:SRI

[Sch86a] Walter Schachermayer. Some remarks on integral operators and equimeasurable sets. Lecture Notes in Mathematics, 1206:242– 258, 1986. CODEN LNMAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076303/.

Schanuel:1986:WLP

[Sch86b] Stephen H. Schanuel. What is the length of a potato? Lecture Notes in Mathematics, 1174:118–126, 1986. CODEN LN-MAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076939/.

Schmidt:1986:EMP

[Sch86c] J. W. Schmidt. Enclosing methods for perturbed boundary value problems in nonlinear difference equations. Lecture Notes in Mathematics, 1192:333–338, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076090/.

Schneider:1986:CSN

[Sch86d] Michael Schneider. Complex surfaces with negative tangent bundle. Lecture Notes in Mathematics, 1194:150–157, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077002/.

Schneider:1986:RSL

[Sch86e] W. R. Schneider. Rigorous scaling laws for Dyson measures. Lecture Notes in Mathematics, 1158:224–233, 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080221/.

Schofield:1986:ULH

[Sch86f] A. H. Schofield. Universal localisation for hereditary rings and quivers. Lecture Notes in Mathematics, 1197:149–164, 1986. CO-DEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076322/.

Schott:1986:RRW

[Sch86g] R. Schott. Recurrent random walks on homogeneous spaces. Lecture Notes in Mathematics, 1210:146–152, 1986. CODEN LN-MAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077180/.

Schurmann:1986:CLT

[Sch86h] Michael Schürmann. A central limit theorem for coalgebras. Lecture Notes in Mathematics, 1210:153–157, 1986. CODEN LN-MAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077181/.

Schuur:1986:AAS

[Sch86i] Peter Cornelis Schuur. Asymptotic Analysis of Soliton Problems: An Inverse Scattering Approach, volume 1232 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 180 pp. LCCN QA3

.L28 no. 1232. URL http://link.springer.com/book/10.1007/
BFb0073054;   http://www.springerlink.com/content/978-
3-540-47387-9.

Schuur:1986:AESa

[Sch86j] Peter Cornelis Schuur. Asymptotic estimates of solutions of the Korteweg–de Vries equation on right half lines slowly moving to the left. Lecture Notes in Mathematics, 1232:34–77, 1986. CODEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073057/.

Schuur:1986:AESb

[Sch86k] Peter Cornelis Schuur. Asymptotic estimates of solutions of the sine-Gordon equation on right half lines almost linearly moving leftward. Lecture Notes in Mathematics, 1232:136–150, 1986. CO-DEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073062/.

Schuur:1986:DES

[Sch86l] Peter Cornelis Schuur. Decomposition and estimates of solutions of the modified Korteweg–de Vries equation on right half lines slowly moving leftward. Lecture Notes in Mathematics, 1232:109– 123, 1986. CODEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073060/.

Schuur:1986:ESK

[Sch86m] Peter Cornelis Schuur. The emergence of solitons of the Korteweg– de Vries equation from arbitrary initial conditions. Lecture Notes in Mathematics, 1232:13–33, 1986. CODEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073056/.

Schuur:1986:I

[Sch86n] Peter Cornelis Schuur. Introduction. Lecture Notes in Mathematics, 1232:1–12, 1986. CODEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073055/.

Schuur:1986:MPSa

[Sch86o] Peter Cornelis Schuur. Multisoliton phase shifts for the Korteweg– de Vries equation in the case of a nonzero reflection coefficient. Lecture Notes in Mathematics, 1232:78–88, 1986. CODEN LN-MAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073058/.

Schuur:1986:MPSb

[Sch86p] Peter Cornelis Schuur. Multisoliton phase shifts for the modified Korteweg–de Vries equation in the case of a nonzero reflection coefficient. Lecture Notes in Mathematics, 1232:124–135, 1986. CODEN LNMAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073061/.

Schuur:1986:ACP

[Sch86q] Peter Cornelis Schuur. On the approximation of a complex potential in the Zakharov–Shabat system by its reflectionless part. Lecture Notes in Mathematics, 1232:151–172, 1986. CODEN LN-MAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073063/.

Schuur:1986:ARP

[Sch86r] Peter Cornelis Schuur. On the approximation of a real potential in the Zakharov–Shabat system by its reflectionless part. Lecture Notes in Mathematics, 1232:89–108, 1986. CODEN LN-MAA2. ISBN 3-540-17203-3 (print), 3-540-47387-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073059/.

Schwarz:1986:MSC

[Sch86s] Mariola B. Schwarz. Mean square convergence of weak martingales. Lecture Notes in Mathematics, 1193:110–112, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077102/.

Schachermayer:1987:TJB

[Sch87a] W. Schachermayer. On a theorem of J. Bourgain on finite dimensional decompositions and the Radon–Nikodym property. Lecture Notes in Mathematics, 1267:96–112, 1987. CODEN LN-MAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078139/.

Schappert:1987:CSU

[Sch87b] Albert Schappert. A characterisation of strictly unimodular plane curve singularities. Lecture Notes in Mathematics, 1273:168– 177, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078843/.

Schechter:1987:CPL

[Sch87c] Eric Schechter. Compact perturbations of linear m-dissipative operators which lack Gihman’s property. Lecture Notes in Mathematics, 1248:142–161, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077423/.

Schechtman:1987:ES

[Sch87d] Gideon Schechtman. Embedding X_{p}^{m} spaces into ℓ_{r}^{n}. Lecture Notes in Mathematics, 1267:53–74, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078136/.

Schechtman:1987:DCC

[Sch87e] V. V. Schechtman. On the delooping of Chern character and Adams operations. Lecture Notes in Mathematics, 1289:265– 319, 1987. CODEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078371/.

Schmidt:1987:UEA

[Sch87f] Roland Schmidt. Untergruppenverbände endlicher auflösbarer Gruppen. (German) []. Lecture Notes in Mathematics, 1281:130– 150, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078699/.

Schneider:1987:GOS

[Sch87g] W. R. Schneider. Generalized one-sided stable distributions. Lecture Notes in Mathematics, 1250:269–287, 1987. CODEN LN-MAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077361/.

Schober:1987:KC

[Sch87h] Glenn E. Schober. Kirwan’s conjecture. Lecture Notes in Mathematics, 1275:266–271, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078358/.

Schoissengeier:1987:EFG

[Sch87i]  Johannes Schoißengeier. Eine Explizite Formel für ∑_{n≤NB_{2}({nα})}.

(German) []. Lecture Notes in Mathematics, 1262:134–138, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078602/.

Schreyer:1987:FCC

[Sch87j] Frank-Olaf Schreyer. Finite and countable CM-representation type. Lecture Notes in Mathematics, 1273:9–34, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078836/.

Schrohe:1987:SWS

[Sch87k] Elmar Schrohe. Spaces of weighted symbols and weighted Sobolev spaces on manifolds. Lecture Notes in Mathematics, 1256:360– 377, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077751/.

Schulze:1987:MEP

[Sch87l] B.-W. Schulze. Mellin expansions of pseudo-differential operators and conormal asymptotics of solutions. Lecture Notes in Mathematics, 1256:378–401, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077752/.

Schachermayer:1988:ECS

[Sch88a] W. Schachermayer. An example concerning strong regularity and points of continuity in Banach spaces. Lecture Notes in Mathematics, 1332:64–79, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081612/.

Schappacher:1988:AIC

[Sch88b] Norbert Schappacher. Abelian integrals with complex multiplication. Lecture Notes in Mathematics, 1301:128–137, 1988. CODEN LNMAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082099/.

Schappacher:1988:AHC

[Sch88c] Norbert Schappacher. Algebraic Hecke characters. Lecture Notes in Mathematics, 1301:1–22, 1988. CODEN LNMAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082095/.

Schappacher:1988:EIG

[Sch88d] Norbert Schappacher. Elliptic integrals and the gamma function. Lecture Notes in Mathematics, 1301:117–127, 1988. CODEN LN-MAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082098/.

Schappacher:1988:MAH

[Sch88e] Norbert Schappacher. Motives for algebraic Hecke characters. Lecture Notes in Mathematics, 1301:23–80, 1988. CODEN LN-MAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082096/.

Schappacher:1988:MCM

[Sch88f] Norbert Schappacher. Motives of CM modular forms. Lecture Notes in Mathematics, 1301:138–147, 1988. CODEN LN-MAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082100/.

Schappacher:1988:PAH

[Sch88g] Norbert Schappacher. The periods of algebraic Hecke characters. Lecture Notes in Mathematics, 1301:81–116, 1988. CODEN LN-MAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082097/.

Scheerer:1988:RTH

[Sch88h] H. Scheerer. Report on tame homotopy theory via differential forms. Lecture Notes in Mathematics, 1318:192–207, 1988. CO-DEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077803/.

Scheinkman:1988:DGE

[Sch88i] José A. Scheinkman. Dynamic general equilibrium models — two examples. Lecture Notes in Mathematics, 1330:44–71, 1988. CODEN LNMAA2. ISBN 3-540-50003-0 (print), 3-540-45883-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078158/.

Schumacher:1988:JC

[Sch88j] Carol S. Schumacher. JH* had the C*PCP. Lecture Notes in Mathematics, 1332:150–155, 1988. CODEN LNMAA2. ISBN 3-540-50018-9 (print), 3-540-45892-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081617/.

Schwartz:1988:FNC

[Sch88k] Lionel Schwartz. La filtration nilpotente de la catégorie μ et la cohomologie des espaces de lacets. (French) []. Lecture Notes in Mathematics, 1318:208–218, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077804/.

Schaumloffel:1989:WNS

[Sch89a] Kay-Uwe Schaumlöffel. White noise in space and time as the time-derivative of a cylindrical Wiener process. Lecture Notes in Mathematics, 1390:225–229, 1989. CODEN LNMAA2. ISBN 3-540-51510-0 (print), 3-540-48200-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083950/.

Schechter:1989:SLE

[Sch89b] Eric Schechter. A survey of local existence theories for abstract nonlinear initial value problems. Lecture Notes in Mathematics, 1394:136–184, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086759/.

Schechtman:1989:RCD

[Sch89c] Gideon Schechtman. A remark concerning the dependence on ϵ in Dvoretzky’s theorem. Lecture Notes in Mathematics, 1376:274– 277, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090061/.

Schlindwein:1989:SNS

[Sch89d] Chaz Schlindwein. Special non-special N_{1}-trees. Lecture Notes in Mathematics, 1401:160–166, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097338/.

Schoen:1989:CCB

[Sch89e] Chad Schoen. Cyclic covers of P^{v} branched along v + 2 hyperplanes and the generalized Hodge Conjecture for certain Abelian varieties. Lecture Notes in Mathematics, 1399:137–154, 1989. CO-DEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0095974/.

Schoen:1989:VTT

[Sch89f] Richard M. Schoen. Variational theory for the total scalar curvature functional for Riemannian metrics and related topics. Lecture Notes in Mathematics, 1365:120–154, 1989. CODEN LN-MAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089180/.

Schreyer:1989:GCG

[Sch89g] Frank-Olaf Schreyer. Green’s conjecture for general P -gonal curves of large genus. Lecture Notes in Mathematics, 1389:254– 260, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085937/.

Scoppola:1987:ENF

[Sco87] Carlo Maria Scoppola. An example of a nonabelian Frobenius– Wielandt complement. Lecture Notes in Mathematics, 1281:151– 157, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078700/.

Seda:1986:SBV

[Šed86] V. Šeda. Surjectivity and boundary value problems. Lecture Notes in Mathematics, 1192:161–170, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076064/.

Segal:1985:LG

[Seg85a] G. B. Segal. Loop groups. Lecture Notes in Mathematics, 1111: 155–168, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084589/.

Segal:1985:RAQ

[Seg85b] I. E. Segal. Regularity aspects of the quantized perturbative Smatrix in 4-dimensional space–time. Lecture Notes in Mathematics, 1139:136–149, 1985. CODEN LNMAA2. ISBN 3-540-15666-6 (print), 3-540-39585-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074581/.

Seidman:1987:TCL

[Sei87] Thomas I. Seidman. Two compactness lemmas. Lecture Notes in Mathematics, 1248:162–168, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077424/.

Seidman:1989:SSP

[Sei89a] Thomas I. Seidman. Switching systems and periodicity. Lecture Notes in Mathematics, 1394:199–210, 1989. CODEN LN-MAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086761/.

Seidman:1989:TSP

[Sei89b] Thomas I. Seidman. The transient semiconductor problem with generation terms, II. Lecture Notes in Mathematics, 1394:185– 198, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086760/.

Seligman:1986:SLM

[Sel86] George B. Seligman. Spin-like modules for certain infinitedimensional Lie algebras. Lecture Notes in Mathematics, 1220: 33–55, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099504/.

Selmi:1987:CCC

[Sel87] Mohamed Selmi. Critèare de comparaison de certains noyaux de Green. (French) []. Lecture Notes in Mathematics, 1235:172– 194, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072756/.

Selick:1988:MC

[Sel88a] Paul Selick. Moore conjectures. Lecture Notes in Mathematics, 1318:219–227, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077805/.

Seligman:1988:C

[Sel88b] George B. Seligman. Complements. Lecture Notes in Mathematics, 1300:179–185, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079304/.

Seligman:1988:CEA

[Sel88c] George B. Seligman. Construction of exceptional algebras from quadratic forms. Lecture Notes in Mathematics, 1300:52–83, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079300/.

Seligman:1988:CMS

[Sel88d] George B. Seligman. Construction of modules for the superexceptional algebras of rank one. Lecture Notes in Mathematics, 1300:141–178, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079303/.

Seligman:1988:CMP

[Sel88e] George B. Seligman. Construction of modules with prescribed relative highest weights, for the isotropic algebras of chapter 3. Lecture Notes in Mathematics, 1300:37–51, 1988. CODEN LN-MAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079299/.

Seligman:1988:GC

[Sel88f] George B. Seligman. General considerations. Lecture Notes in Mathematics, 1300:12–20, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079297/.

Seligman:1988:IE

[Sel88g] George B. Seligman. An introductory example: sl(n, D). Lecture Notes in Mathematics, 1300:1–11, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079296/.

Seligman:1988:IAM

[Sel88h] George B. Seligman. Involutorial algebras and modules for their skew elements. Lecture Notes in Mathematics, 1300:21–36, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079298/.

Seligman:1988:NRE

[Sel88i] George B. Seligman. Non-reduced exceptional algebras with a one-dimensional root space. Lecture Notes in Mathematics, 1300: 115–140, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079302/.

Seligman:1988:REA

[Sel88j] George B. Seligman. Representations of exceptional algebras constructed from quadratic forms. Lecture Notes in Mathematics, 1300:84–114, 1988. CODEN LNMAA2. ISBN 3-540-18973-4 (print), 3-540-38864-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079301/.

Seligman:1989:KMM

[Sel89a] George B. Seligman. Kac–Moody modules and generalized Clifford algebras. Lecture Notes in Mathematics, 1373:124–143, 1989. CODEN LNMAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088892/.

Seligman:1989:WL

[Sel89b] George B. Seligman. Workshop lectures. Lecture Notes in Mathematics, 1373:144–145, 1989. CODEN LNMAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088893.

Seneor:1986:AFR

[Sén86] R. Sénéor. Asymptotic freedom: a rigorous approach. Lecture Notes in Mathematics, 1158:234–244, 1986. CODEN LN-MAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080222/.

Serre:1989:ALM

[Ser89] Jean-Pierre Serre, editor. Algèbre locale, multiplicités; cours au Collège de France, 1957–1958. (French) [Local algebra, multiplicities; courses at the Collège of France, 1957–1958], volume 11 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. CODEN LNMAA2. ISBN 0-387-07028-1, 3-540-07028-1. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 11. URL

http://link.springer.com/book/10.1007/BFb0087610.

Sevryuk:1986:BTR

[Sev86a] M. B. Sevryuk. The behaviour of trajectories of reversible vectorfields near a symmetric cycle. Lecture Notes in Mathematics, 1211:159–170, 1986. CODEN LNMAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075881/.

Sevryuk:1986:CI

[Sev86b] M. B. Sevryuk. Chapter introduction. Lecture Notes in Mathematics, 1211:1–21, 1986. CODEN LNMAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075878/.

Sevryuk:1986:FSC

[Sev86c] M. B. Sevryuk. Families of symmetric cycles near an equilibrium of a reversible vectorfield. Lecture Notes in Mathematics, 1211:228–292, 1986. CODEN LNMAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075884/.

Sevryuk:1986:KTP

[Sev86d] M. B. Sevryuk. Kolmogorov tori of perturbations of integrable reversible diffeomorphisms and vectorfields. Lecture Notes in Mathematics, 1211:22–91, 1986. CODEN LNMAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075879/.

Sevryuk:1986:NAR

[Sev86e] M. B. Sevryuk. Non-autonomous reversible differential equations. Lecture Notes in Mathematics, 1211:171–180, 1986. CODEN LN-MAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075882/.

Sevryuk:1986:NFR

[Sev86f] M. B. Sevryuk. Normal forms for reversible diffeomorphisms and vectorfields near an equilibrium and their Kolmogorov tori. Lecture Notes in Mathematics, 1211:92–158, 1986. CODEN LN-MAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075880/.

Sevryuk:1986:SRZ

[Sev86g] M. B. Sevryuk. Structure of resonant zones of reversible diffeomorphisms and vectorfields. Lecture Notes in Mathematics, 1211:181–227, 1986. CODEN LNMAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075883/.

Sevryuk:1986:RS

[Sev86h] Mikhail B. Sevryuk. Reversible Systems, volume 1211 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16819-2 (print), 3-540-44865-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 319 pp. LCCN QA3

.L28 no. 1211. URL http://link.springer.com/book/10.1007/
BFb0075877;   http://www.springerlink.com/content/978-
3-540-44865-5.

Sewell:1987:QFG

[Sew87] Geoffrey L. Sewell. Quantum fields, gravitation and thermodynamics. Lecture Notes in Mathematics, 1250:288–297, 1987. CO-DEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077362/.

Schroder:1986:CSM

[SH86] W. Schröder and D. Hänel. A comparison of several Mg-methods for the solution of the time-dependent Navier–Stokes equations. Lecture Notes in Mathematics, 1228:272–284, 1986. CODEN LN-MAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072652/.

Shaginyan:1987:PAV

[Sha87] A. A. Shaginyan. On potential approximation of vector fields. Lecture Notes in Mathematics, 1275:272–279, 1987. CODEN LN-MAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078359/.

Shampine:1989:TPO

[Sha89] L. F. Shampine. Tolerance proportionality in ODE codes. Lecture Notes in Mathematics, 1386:118–136, 1989. CODEN LN-MAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089235/.

Shcolnick:1985:RIS

[Shc85] S. M. Shcolnick. On the ratio of independent stable random variables. Lecture Notes in Mathematics, 1155:349–354, 1985. CO-DEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074827/.

Shelah:1986:ACT

[She86a] Saharon Shelah. Around Classification Theory of Models, volume 1182 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 279 pp. LCCN QA3 .L28 no. 1182. URL http:

//link.springer.com/book/10.1007/BFb0098503;   http://
www.springerlink.com/content/978-3-540-39788-5.

Shelah:1986:CPI

[She86b] Saharon Shelah. Classification over a predicate II. Lecture Notes in Mathematics, 1182:47–90, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098505/.

Shelah:1986:CGQ

[She86c] Saharon Shelah. Classifying generalized quantifiers. Lecture Notes in Mathematics, 1182:1–46, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098504/.

Shelah:1986:ECS

[She86d] Saharon Shelah. The existence of coding sets. Lecture Notes in Mathematics, 1182:188–202, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098510/.

Shelah:1986:EER

[She86e] Saharon Shelah. Existence of endo-rigid Boolean algebras. Lecture Notes in Mathematics, 1182:91–119, 1986. CODEN LN-MAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098506/.

Shelah:1986:MLH

[She86f] Saharon Shelah. Monadic logic: Hanf numbers. Lecture Notes in Mathematics, 1182:203–223, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098511/.

Shelah:1986:MSC

[She86g] Saharon Shelah. More on stationary coding. Lecture Notes in Mathematics, 1182:224–246, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098512/.

Shelah:1986:NSU

[She86h] Saharon Shelah. Non standard uniserial module over a uniserial domain exists. Lecture Notes in Mathematics, 1182:135–150, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098508/.

Shelah:1986:NFA

[She86i] Saharon Shelah. A note on κ-freeness of abelian groups. Lecture Notes in Mathematics, 1182:260–268, 1986. CODEN LN-MAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098514/.

Shelah:1986:CTM

[She86j] Saharon Shelah. On countable theories with models — homogeneous models only. Lecture Notes in Mathematics, 1182:269– 271, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098515/.

Shelah:1986:DSF

[She86k] Saharon Shelah. On decomposable sentences for finite models. Lecture Notes in Mathematics, 1182:272–275, 1986. CODEN LN-MAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098516/.

Shelah:1986:NIB

[She86l] Saharon Shelah. On normal ideals and Boolean algebras. Lecture Notes in Mathematics, 1182:247–259, 1986. CODEN LN-MAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098513/.

Shelah:1986:SP

[She86m] Saharon Shelah. On the no(M) for M of singular power. Lecture Notes in Mathematics, 1182:120–134, 1986. CODEN LN-MAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098507/.

Shelah:1986:RS

[She86n] Saharon Shelah. Remarks on squares. Lecture Notes in Mathematics, 1182:276–279, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098517/.

Shelah:1986:RNI

[She86o] Saharon Shelah. Remarks on the numbers of ideals of Boolean algebra and open sets of a topology. Lecture Notes in Mathematics, 1182:151–187, 1986. CODEN LNMAA2. ISBN 3-540-16448-0 (print), 3-540-39788-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098509/.

Shekhtman:1987:GRP

[She87a] Boris Shekhtman. On the geometry of real polynomials. Lecture Notes in Mathematics, 1287:161–175, 1987. CODEN LN-MAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078904/.

Shelah:1987:ACT

[She87b] S. Shelah. On almost categorical theories. Lecture Notes in Mathematics, 1292:498–500, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082244/.

Shelah:1987:CNE

[She87c] Saharon Shelah. Classification of non elementary classes II abstract elementary classes. Lecture Notes in Mathematics, 1292: 419–497, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082243/.

Shelah:1987:UC

[She87d] Saharon Shelah. Universal classes. Lecture Notes in Mathematics, 1292:264–418, 1987. CODEN LNMAA2. ISBN 3-540-18674-3 (print), 3-540-48049-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082242/.

Sher:1987:CTS

[She87e] R. B. Sher. Complement theorems in shape theory, II. Lecture Notes in Mathematics, 1283:212–220, 1987. CODEN LN-MAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081430/.

Shelah:1989:CPP

[She89a] Saharon Shelah. Consistency of positive partition theorems for graphs and models. Lecture Notes in Mathematics, 1401:167– 193, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097339/.

Shen:1989:STR

[She89b] Yibing Shen. Stability of totally real minimal submanifolds. Lecture Notes in Mathematics, 1369:235–242, 1989. CODEN LN-MAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087537/.

Shinoda:1985:ATO

[Shi85] Juichi Shinoda. Absolute type 2 objects. Lecture Notes in Mathematics, 1141:343–356, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076229/.

Shi:1986:REC

[Shi86a] Jian-Yi Shi. σ^{−1}(λ) is an rl-equivalence class of A_{n}. Lecture Notes in Mathematics, 1179:235, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074983.

Shi:1986:AST

[Shi86b] Jian-Yi Shi. Admissible sign types of rank n. Lecture Notes in Mathematics, 1179:99–113, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074975/.

Shi:1986:ADA

[Shi86c] Jian-Yi Shi. The algebraic descriptions of the affine Weyl group A_{n} of type Ã_{n−1}, n > 2. Lecture Notes in Mathematics, 1179:

66–76, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074972/.

Shi:1986:AKL

[Shi86d] Jian-Yi Shi. Applications of Kazhdan–Lusztig theory. Lecture Notes in Mathematics, 1179:32–59, 1986. CODEN LN-MAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074970/.

Shi:1986:CGH

[Shi86e] Jian-Yi Shi. Coxeter groups, Hecke algebras and their representations. Lecture Notes in Mathematics, 1179:1–31, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074969/.

Shi:1986:GIK

[Shi86f] Jian-Yi Shi. Geometric interpretations of the Kazhdan–Lusztig polynomials. Lecture Notes in Mathematics, 1179:60–65, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074971/.

Shi:1986:GDA

[Shi86g] Jian-Yi Shi. A geometrical description of the affine Weyl group A_{n}. Lecture Notes in Mathematics, 1179:84–98, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074974/.

Shi:1986:IAS

[Shi86h] Jian-Yi Shi. The inserting algorithm on the set C. Lecture Notes in Mathematics, 1179:273–284, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074988/.

Shi:1986:ISO

[Shi86i] Jian-Yi Shi. Iterated star operations and interchanging operations on blocks. Lecture Notes in Mathematics, 1179:114–128, 1986.

CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074976/.

Shi:1986:KLC

[Shi86j] Jian-Yi Shi. The Kazhdan–Lusztig Cells in Certain Affine Weyl Groups, volume 1179 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 307 pp. LCCN QA3 .L28 no. 1179. URL http:

//link.springer.com/book/10.1007/BFb0074968;   http://
www.springerlink.com/content/978-3-540-39780-9.

Shi:1986:LRC

[Shi86k] Jian-Yi Shi. The left and right cells in σ^{−1}(λ). Lecture Notes in Mathematics, 1179:222–234, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074982/.

Shi:1986:LCC

[Shi86l] Jian-Yi Shi. Left cells are characterized by the generalized right τ-invariant. Lecture Notes in Mathematics, 1179:236–242, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074984/.

Shi:1986:OSA

[Shi86m] Jian-Yi Shi. The orbit space Ã_{n} of the affine Weyl group A_{n}. Lecture Notes in Mathematics, 1179:152–165, 1986. CODEN LN-MAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074979/.

Shi:1986:PAE

[Shi86n] Jian-Yi Shi. The partition of n associated with an element of the affine Weyl group A_{n}. Lecture Notes in Mathematics, 1179: 77–83, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074973/.

Shi:1986:ROL

[Shi86o] Jian-Yi Shi. Raising operations on layers. Lecture Notes in Mathematics, 1179:202–221, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074981/.

Shi:1986:SBE

[Shi86p] Jian-Yi Shi. The sequence ξ(w, k) beginning with an element of N_{λ}. Lecture Notes in Mathematics, 1179:166–201, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074980/.

Shi:1986:SNE

[Shi86q] Jian-Yi Shi. The set N_{λ} of normalized elements of σ^{−1}(λ). Lecture Notes in Mathematics, 1179:146–151, 1986. CODEN LN-MAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074978/.

Shi:1986:SPC

[Shi86r] Jian-Yi Shi. Some properties of cells and other equivalence classes of A_{n}. Lecture Notes in Mathematics, 1179:247–255, 1986. CO-DEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074986/.

Shi:1986:SSK

[Shi86s] Jian-Yi Shi. Some special kinds of sign types. Lecture Notes in Mathematics, 1179:256–272, 1986. CODEN LNMAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074987/.

Shi:1986:SAW

[Shi86t] Jian-Yi Shi. The subset σ^{−1}(λ) of the affine Weyl group A_{n}. Lecture Notes in Mathematics, 1179:129–145, 1986. CODEN LN-MAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074977/.

Shi:1986:TSC

[Shi86u] Jian-Yi Shi. The two-sided cells of the affine Weyl group A_{n}. Lecture Notes in Mathematics, 1179:243–246, 1986. CODEN LN-MAA2. ISBN 3-540-16439-1 (print), 3-540-39780-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074985/.

Shi:1986:NBG

[Shi86v] Shengming Shi. On the number of p-blocks with given defect groups and some applications of the p-power homomorphism. Lecture Notes in Mathematics, 1185:314–339, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076182/.

Shiohama:1986:CPB

[Shi86w] Katsuhiro Shiohama. Critical points of Busemann functions on complete open surfaces. Lecture Notes in Mathematics, 1201:254– 265, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075661/.

Shiota:1987:NM

[Shi87a] Masahiro Shiota. C^{0} Nash manifolds. Lecture Notes in Mathematics, 1269:175–188, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078577/.

Shiota:1987:ANMb

[Shi87b] Masahiro Shiota. Affine C^{ω} Nash manifolds. Lecture Notes in Mathematics, 1269:189–215, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078578/.

Shiota:1987:ANMa

[Shi87c] Masahiro Shiota. Affine C^{r} Nash manifolds. Lecture Notes in Mathematics, 1269:142–154, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078575/.

Shiota:1987:AT

[Shi87d] Masahiro Shiota. Approximation theorem. Lecture Notes in Mathematics, 1269:79–141, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078574/.

Shiota:1987:I

[Shi87e] Masahiro Shiota. Introduction. Lecture Notes in Mathematics, 1269:1–5, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078572/.

Shiota:1987:NNM

[Shi87f] Masahiro Shiota. Nonaffine C^{ω} Nash manifolds. Lecture Notes in Mathematics, 1269:155–174, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078576/.

Shiota:1987:P

[Shi87g] Masahiro Shiota. Preliminaries. Lecture Notes in Mathematics, 1269:6–78, 1987. CODEN LNMAA2. ISBN 3-540-18102-4 (print), 3-540-47763-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078573/.

Shirokov:1988:CIS

[Shi88a] Nikolai A. Shirokov. Closed ideals in the space X_{pq}^{μ}(ω, ℓ). Lecture Notes in Mathematics, 1312:174–205, 1988. CODEN LN-MAA2. ISBN 3-540-19255-7 (print), 3-540-39175-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082816/.

Shirokov:1988:P

[Shi88b] Nikolai A. Shirokov. The (F)-property. Lecture Notes in Mathematics, 1312:25–53, 1988. CODEN LNMAA2. ISBN 3-540-19255-7 (print), 3-540-39175-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082813/.

Shirokov:1988:I

[Shi88c] Nikolai A. Shirokov. Introduction. Lecture Notes in Mathematics, 1312:1–19, 1988. CODEN LNMAA2. ISBN 3-540-19255-7 (print), 3-540-39175-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082811/.

Shirokov:1988:MAF

[Shi88d] Nikolai A. Shirokov. Moduli of analytic functions smooth up to the boundary. Lecture Notes in Mathematics, 1312:54–135, 1988. CODEN LNMAA2. ISBN 3-540-19255-7 (print), 3-540-39175-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082814/.

Shirokov:1988:N

[Shi88e] Nikolai A. Shirokov. Notations. Lecture Notes in Mathematics, 1312:21–24, 1988. CODEN LNMAA2. ISBN 3-540-19255-7 (print), 3-540-39175-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082812/.

Shirokov:1988:ZTM

[Shi88f] Nikolai A. Shirokov. Zeros and their multiplicities. Lecture Notes in Mathematics, 1312:136–173, 1988. CODEN LNMAA2. ISBN 3-540-19255-7 (print), 3-540-39175-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082815/.

Shimoda:1989:EPS

[Shi89] Mamoru Shimoda. Elementary properties of a system of fundamental sequences for Γ_{0}. Lecture Notes in Mathematics, 1388: 141–152, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083668/.

Shnirelman:1986:PSW

[Shn86] A. I. Shnirel’man. On the principle of the shortest way in the dynamics of systems with constraints. Lecture Notes in Mathematics, 1214:117–130, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075962/.

Showalter:1985:VTA

[Sho85] R. E. Showalter. Variational theory and approximation of boundary value problems. Lecture Notes in Mathematics, 1129:140– 179, 1985. CODEN LNMAA2. ISBN 3-540-15234-2 (print), 3-540-39412-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075159/.

Shokin:1987:CAS

[Sho87] Y. I. Shokin. A computer-aided system for investigation and construction of difference schemes of gas dynamics. Lecture Notes in Mathematics, 1270:128–134, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078322/.

Siegel:1985:EAH

[Sie85] Jerrold Siegel. Extrema associated with homotopy classes of maps. Lecture Notes in Mathematics, 1167:260–267, 1985. CODEN LN-MAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075229/.

Siganov:1987:SEC

[Šig87] I. Šiganov. On stability estimates of Cramér’s theorem. Lecture Notes in Mathematics, 1233:178–181, 1987. CODEN LN-MAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072723/.

Silva:1988:ASC

[Sil88a] Ana Maria F. Silva. Atiyah sequences and complete closed pseudogroups preserving a local parallelism. Lecture Notes in Mathematics, 1345:302–316, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081409/.

Rehermann:1988:GFB

[Sil88b] Celiar Silva Rehermann. Généralisation de formules de bornage de L-I et applications aux L^{p}. (French) []. Lecture Notes in Mathematics, 1354:199–208, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089597/.

Silhol:1989:ACR

[Sil89a] Robert Silhol. Algebraic cycles on real varieties and application to rational surfaces. Lecture Notes in Mathematics, 1392:56–74, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088818/.

Silhol:1989:DRA

[Sil89b] Robert Silhol. Digression on real abelian varieties and classification of real abelian surfaces. Lecture Notes in Mathematics, 1392: 75–94, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088819/.

Silhol:1989:GRR

[Sil89c] Robert Silhol. General results on real algebraic surfaces. Lecture Notes in Mathematics, 1392:27–55, 1989. CODEN LN-MAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088817/.

Silhol:1989:PRA

[Sil89d] Robert Silhol. Preliminaries on real algebraic varieties. Lecture Notes in Mathematics, 1392:1–26, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088816/.

Silhol:1989:RRSb

[Sil89e] Robert Silhol. Rational real surfaces. Lecture Notes in Mathematics, 1392:109–138, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088821/.

Silhol:1989:RS

[Sil89f] Robert Silhol. Real K_{3} surfaces. Lecture Notes in Mathematics, 1392:178–200, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088823/.

Silhol:1989:RES

[Sil89g] Robert Silhol. Real elliptic surfaces. Lecture Notes in Mathematics, 1392:139–177, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088822/.

Silhol:1989:RRSa

[Sil89h] Robert Silhol. Real ruled surfaces. Lecture Notes in Mathematics, 1392:95–108, 1989. CODEN LNMAA2. ISBN 3-540-51563-1 (print), 3-540-70649-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088820/.

Silverman:1989:IPC

[Sil89i] Joseph H. Silverman. Integral points on curves and surfaces. Lecture Notes in Mathematics, 1380:202–241, 1989. CODEN LN-MAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086555/.

Simon:1985:SAT

[Sim85a] Barry Simon. Some aspects of the theory of Schrödinger operators. Lecture Notes in Mathematics, 1159:177–203, 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080333/.

Simon:1985:BM

[Sim85b] Leon Simon. Back matter. Lecture Notes in Mathematics, 1161: 278–289, 1985. CODEN LNMAA2. ISBN 3-540-16040-X (print), 3-540-39716-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-39716-8/1.pdf.

Simon:1985:ISE

[Sim85c] Leon Simon. Isolated singularities of extrema of geometric variational problems. Lecture Notes in Mathematics, 1161:206–277, 1985. CODEN LNMAA2. ISBN 3-540-16040-X (print), 3-540-39716-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075139/.

Simpson:1985:RTA

[Sim85d] Stephen G. Simpson. Recursion theoretic aspects of the dual Ramsey theorem. Lecture Notes in Mathematics, 1141:357–371, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076230/.

Simader:1986:HRW

[Sim86] C. G. Simader. Higher regularity of weak solutions of strongly nonlinear elliptic equations. Lecture Notes in Mathematics, 1192: 243–246, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076076/.

Simis:1987:MBN

[Sim87] A. Simis. Multiplicities and Betti numbers of homogeneous ideals. Lecture Notes in Mathematics, 1266:232–250, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078186/.

Simon:1989:DPL

[Sim89] Udo Simon. Dirichlet problems and the Laplacian in affine hypersurface theory. Lecture Notes in Mathematics, 1369:243–260, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087538/.

Singer:1987:AHS

[Sin87] William M. Singer. On the Λ-algebra and the homology of symmetric groups. Lecture Notes in Mathematics, 1286:305–341, 1987. CODEN LNMAA2. ISBN 3-540-18481-3 (print), 3-540-47986-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078749/.

Sion:1986:BM

[Sio86a] Maurice Sion. Back matter. Lecture Notes in Mathematics, 1206: 281–289, 1986. CODEN LNMAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-40955-7/1.pdf.

Sion:1986:CML

[Sio86b] Maurice Sion. Cylinder measures, local bases and nuclearity. Lecture Notes in Mathematics, 1206:259–280, 1986. CODEN LN-MAA2. ISBN 3-540-16787-0 (print), 3-540-40955-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076304/.

Siu:1985:SRR

[Siu85a] Yum-Tong Siu. Some recent results in complex manifold theory related to vanishing theorems for the semipositive case. Lecture Notes in Mathematics, 1111:169–192, 1985. CODEN LN-MAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084590/.

Siu:1985:TOP

[Siu85b] Yum-Tong Siu. Ten open problems connected with Hermitian and Kähler manifolds. Lecture Notes in Mathematics, 1165:279– 280, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076160.

Siu:1986:NVD

[Siu86] Yum-Tong Siu. Nonequidimensional value distribution theory and subvariety extension. Lecture Notes in Mathematics, 1194:158– 174, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077003/.

Siu:1987:KEM

[Siu87] Yum-Tong Siu. Kähler–Einstein metrics for the case of positive first Chern class. Lecture Notes in Mathematics, 1277:120– 130, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078247/.

Sleeman:1985:OPD

[SJ85] Brian D. Sleeman and Richard J. Jarvis, editors. Ordinary and Partial Differential Equations: Proceedings of the Eighth Conference held at Dundee, Scotland, June 25–29, 1984, volume 1151 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15694-1 (print), 3-540-39640-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1151. URL http://link.springer.com/book/10.1007/
BFb0074708;   http://www.springerlink.com/content/978-
3-540-39640-6.

Stara:1986:SRR

[SJ86] J. Stará and O. John. Some regularity results for quasilinear parabolic systems. Lecture Notes in Mathematics, 1192:247– 250, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076077/.

Sjostrand:1987:SRG

[Sjö87] Johannes Sjöstrand. Semiclassical resonances generated by nondegenerate critical points. Lecture Notes in Mathematics, 1256: 402–429, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077753/.

Skibsted:1986:DEL

[Ski86] Erik Skibsted. α-decay and the exponential law. Lecture Notes in Mathematics, 1218:201–222, 1986. CODEN LNMAA2. ISBN 3-540-16826-5 (print), 3-540-47119-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073052/.

Skorobogatko:1987:BCF

[Sko87] V. Ya. Skorobogat’ko. Branched continued fractions and convergence acceleration problems. Lecture Notes in Mathematics, 1237:

46–50, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072453/.

Slaman:1985:RPM

[Sla85a] Theodore A. Slaman. Reflection and the priority method in E-recursion theory. Lecture Notes in Mathematics, 1141:372– 404, 1985. CODEN LNMAA2. ISBN 3-540-15673-9 (print), 3-540-39596-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076231/.

Slavova:1985:BEB

[Sla85b] V. V. Slavova. On the Berry–Esseen bound for Student’s statistic. Lecture Notes in Mathematics, 1155:355–390, 1985. CODEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074828/.

Slemrod:1987:ACP

[Sle87] M. Slemrod. Admissibility criteria for phase boundaries. Lecture Notes in Mathematics, 1270:163–171, 1987. CODEN LN-MAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078326/.

Smith:1985:EMS

[Smi85a] Justin R. Smith. Equivariant Moore spaces. Lecture Notes in Mathematics, 1126:238–270, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074446/.

Smith:1985:ATG

[Smi85b] Larry Smith, editor. Algebraic Topology Göttingen 1984: Proceedings of a Conference held in Göttingen, Nov. 9–15, 1984, volume 1172 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1172. URL http:

//link.springer.com/book/10.1007/BFb0074420;   http://
www.springerlink.com/content/978-3-540-39745-8.

Smitalova:1986:DMP

[Smí86a] K. Smítalová. Delay makes problems in population modelling. Lecture Notes in Mathematics, 1192:421–424, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076103/.

Smith:1986:DOC

[Smi86b] S. P. Smith. Differential operators on commutative algebras. Lecture Notes in Mathematics, 1197:165–177, 1986. CODEN LN-MAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076323/.

Smith:1986:DOA

[Smi86c] S. P. Smith. Differential operators on the affine and projective lines in characteristic p > 0. Lecture Notes in Mathematics, 1220:157–177, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099511/.

Smith:1987:CDO

[Smi87] S. P. Smith. Curves, differential operators and finite dimensional algebras. Lecture Notes in Mathematics, 1296:158–176, 1987. CO-DEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078525/.

Smiley:1989:BSG

[Smi89] Michael W. Smiley. Breathers for the sine-Gordon equation. Lecture Notes in Mathematics, 1394:211–217, 1989. CODEN LN-MAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086762/.

Solo:1986:TAT

[Sol86] Victor Solo. Topics in advanced time series analysis. Lecture Notes in Mathematics, 1215:165–328, 1986. CODEN LNMAA2. ISBN 3-540-16822-2 (print), 3-540-47092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075874/.

Sommese:1986:ATS

[Som86] Andrew John Sommese. On the adjunction theoretic structure of projective varieties. Lecture Notes in Mathematics, 1194:175– 213, 1986. CODEN LNMAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077004/.

Sondow:1985:TIP

[Son85] Jonathan D. Sondow. Triviality of the involution on SK_{1} for periodic groups. Lecture Notes in Mathematics, 1126:271–276, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074447/.

Schulze-Pillot:1989:LRB

[SP89] Rainer Schulze-Pillot. A linear relation between theta series of degree and weight 2. Lecture Notes in Mathematics, 1380:197– 201, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086554/.

Spekreijse:1986:SOA

[Spe86] S. P. Spekreijse. Second order accurate upwind solutions of the 2D steady Euler equations by the use of a defect correction method. Lecture Notes in Mathematics, 1228:285–300, 1986. CODEN LN-MAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072653/.

Speiser:1988:TTF

[Spe88] Robert Speiser. Transversality theorems for families of maps. Lecture Notes in Mathematics, 1311:235–252, 1988. CODEN LN-MAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082917/.

Springer:1985:MAF

[Spr85] T. A. Springer. Microlocalisation algébrique. (French) []. Lecture Notes in Mathematics, 1146:299–316, 1985. CODEN LN-MAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074543/.

Springer:1986:CCA

[Spr86] T. A. Springer. Conjugacy classes in algebraic groups. Lecture Notes in Mathematics, 1185:175–209, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076175/.

Srinivasan:1985:SRN

[Sri85] S. Srinivasan. Some remarks in number theory. Lecture Notes in Mathematics, 1122:201–205, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075762/.

Sundaresan:1985:BDG

[SS85a] Kondagunta Sundaresan and Srinivasa Swaminathan. Basic definitions and geometric properties. Lecture Notes in Mathematics, 1131:3–11, 1985. CODEN LNMAA2. ISBN 3-540-15237-7 (print), 3-540-39415-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075325/.

Sundaresan:1985:GNA

[SS85b] Kondagunta Sundaresan and Srinivasa Swaminathan. Geometry and Nonlinear Analysis in Banach Spaces, volume 1131 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15237-7 (print), 3-540-39415-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 96 (est.) pp. LCCN QA3

.L28 no. 1131. URL http://link.springer.com/book/10.1007/
BFb0075323;   http://www.springerlink.com/content/978-
3-540-39415-0.

Sundaresan:1985:I

[SS85c] Kondagunta Sundaresan and Srinivasa Swaminathan. Introduction. Lecture Notes in Mathematics, 1131:1–2, 1985. CODEN LNMAA2. ISBN 3-540-15237-7 (print), 3-540-39415-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075324.

Sundaresan:1985:SPU

[SS85d] Kondagunta Sundaresan and Srinivasa Swaminathan. Smooth partitions of unity on Banach spaces. Lecture Notes in Mathematics, 1131:37–66, 1985. CODEN LNMAA2. ISBN 3-540-15237-7 (print), 3-540-39415-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075327/.

Sundaresan:1985:SAB

[SS85e] Kondagunta Sundaresan and Srinivasa Swaminathan. Smoothness and approximation in Banach spaces. Lecture Notes in Mathematics, 1131:67–96, 1985. CODEN LNMAA2. ISBN 3-540-15237-7 (print), 3-540-39415-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075328/.

Sundaresan:1985:SCB

[SS85f] Kondagunta Sundaresan and Srinivasa Swaminathan. Smoothness classification of Banach spaces. Lecture Notes in Mathematics, 1131:12–36, 1985. CODEN LNMAA2. ISBN 3-540-15237-7 (print), 3-540-39415-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075326/.

Sato:1986:SFH

[SS86] Eiichi Sato and Heinz Spindler. On the structure of 4 folds with a hyperplane section which is a P^{1} bundle over a ruled surface. Lecture Notes in Mathematics, 1194:145–149, 1986. CODEN LN-MAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077001/.

Segal:1987:STC

[SS87] J. Segal and S. Spież. On shape theory with compact supports. Lecture Notes in Mathematics, 1283:202–211, 1987. CODEN LN-MAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081429/.

Slaman:1988:DFD

[SS88a] Theodore A. Slaman and John R. Steel. Definable functions on degrees. Lecture Notes in Mathematics, 1333:37–55, 1988. CODEN LNMAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084969/.

Smith:1988:DSH

[SS88b] Larry Smith and R. E. Stong. Dirichlet series and homology theory. Lecture Notes in Mathematics, 1326:134–149, 1988. CODEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078043/.

Sternin:1988:LRI

[SS88c] B. Yu. Sternin and V. E. Shatalov. Laplace–Radon integral operators and singularities of solutions of differential equations on complex manifolds. Lecture Notes in Mathematics, 1334:129– 156, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080426/.

Sternin:1988:SSD

[SS88d] B. Yu. Sternin and V. E. Shatalov. Singularities of solutions of differential equations on complex manifolds (characteristical case). Lecture Notes in Mathematics, 1334:269–293, 1988. CODEN LN-MAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080433/.

Shinoda:1989:CHT

[SS89a] Juichi Shinoda and Theodore A. Slaman. The continuum hypothesis and the theory of the Kleene degrees. Lecture Notes in Mathematics, 1388:153–177, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083669/.

Sitaramaiah:1989:SAF

[SS89b] V. Sitaramaiah and M. V. Subbarao. Some asymptotic formulae involving powers of arithmetic functions. Lecture Notes in Mathematics, 1395:201–234, 1989. CODEN LNMAA2. ISBN 3-540-51595-X (print), 3-540-46681-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086405/.

Shiohama:1986:CTR

[SSS86] Katsuhiro Shiohama, Takashi Sakai, and Toshikazu Sunada, editors. Curvature and Topology of Riemannian Manifolds: Proceedings of the 17th International Taniguchi Symposium held in Katata, Japan, Aug. 26–31, 1985, volume 1201 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1201. URL http://link.springer.com/book/10.1007/
BFb0075643;   http://www.springerlink.com/content/978-
3-540-38827-2.

Sotomayor:1988:VFN

[ST88] J. Sotomayor and M. A. Teixeira. Vector fields near the boundary of a 3-manifold. Lecture Notes in Mathematics, 1331:169– 195, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083073/.

Tai:1989:CSF

[sT89a] Hsin sheng Tai. A class of symmetric functions and Chern classes of projective varieties. Lecture Notes in Mathematics, 1369:261– 274, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087539/.

Stromberg:1989:AD

[ST89b] Jan-Olov Strömberg and Alberto Torchinsky. The atomic decomposition. Lecture Notes in Mathematics, 1381:111–121, 1989. CODEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091162/.

Stromberg:1989:BI

[ST89c] Jan-Olov Strömberg and Alberto Torchinsky. The basic inequality. Lecture Notes in Mathematics, 1381:122–133, 1989. CODEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091163/.

Stromberg:1989:CI

[ST89d] Jan-Olov Strömberg and Alberto Torchinsky. Complex interpolation. Lecture Notes in Mathematics, 1381:177–188, 1989. CODEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091166/.

Stromberg:1989:DW

[ST89e] Jan-Olov Strömberg and Alberto Torchinsky. Decomposition of weights. Lecture Notes in Mathematics, 1381:18–29, 1989. CO-DEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0091156/.

Stromberg:1989:DC

[ST89f] Jan-Olov Strömberg and Alberto Torchinsky. A dense class. Lecture Notes in Mathematics, 1381:103–110, 1989. CODEN LN-MAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091161/.

Stromberg:1989:D

[ST89g] Jan-Olov Strömberg and Alberto Torchinsky. Duality. Lecture Notes in Mathematics, 1381:134–149, 1989. CODEN LN-MAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091164/.

Stromberg:1989:ED

[ST89h] Jan-Olov Strömberg and Alberto Torchinsky. Extensions of distributions. Lecture Notes in Mathematics, 1381:60–84, 1989. CO-DEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0091159/.

Stromberg:1989:FUH

[ST89i] Jan-Olov Strömberg and Alberto Torchinsky. Functions in the upper half-space. Lecture Notes in Mathematics, 1381:48–59, 1989. CODEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091158/.

Stromberg:1989:HS

[ST89j] Jan-Olov Strömberg and Alberto Torchinsky. The Hardy spaces. Lecture Notes in Mathematics, 1381:85–102, 1989. CODEN LN-MAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091160/.

Stromberg:1989:SMF

[ST89k] Jan-Olov Strömberg and Alberto Torchinsky. Sharp maximal functions. Lecture Notes in Mathematics, 1381:30–47, 1989. CO-DEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0091157/.

Stromberg:1989:SIM

[ST89l] Jan-Olov Strömberg and Alberto Torchinsky. Singular integrals and multipliers. Lecture Notes in Mathematics, 1381:150– 176, 1989. CODEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091165/.

Stromberg:1989:W

[ST89m] Jan-Olov Strömberg and Alberto Torchinsky. Weights. Lecture Notes in Mathematics, 1381:1–17, 1989. CODEN LNMAA2. ISBN 3-540-51402-3 (print), 3-540-46207-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091155/.

Szarek:1989:IVS

[ST89n] S. J. Szarek and M. Talagrand. An “isomorphic” version of the Sauer–Shelah lemma and the Banach–Mazur distance to the cube. Lecture Notes in Mathematics, 1376:105–112, 1989. CODEN LN-MAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090050/.

Stankiewicz:1985:SOP

[Sta85] Jan Stankiewicz. Some open problems on holomorphic functions of one variable. Lecture Notes in Mathematics, 1165:281–282, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076161.

Stakgold:1986:RDP

[Sta86a] Ivar Stakgold. Reaction-diffusion problems in chemical engineering. Lecture Notes in Mathematics, 1224:119–152, 1986. CODEN LNMAA2. ISBN 3-540-17192-4 (print), 3-540-47352-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072689/.

Stanley:1986:BDC

[Sta86b] Richard P. Stanley. A baker’s dozen of conjectures concerning plane partitions. Lecture Notes in Mathematics, 1234:285– 293, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072521/.

Stafford:1987:GDS

[Sta87a] J. T. Stafford. Global dimension of semiprime Noetherian rings. Lecture Notes in Mathematics, 1296:247–260, 1987. CODEN LN-MAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078531/.

Stahl:1987:NTH

[Sta87b] H. R. Stahl. A note on a theorem by H. N. Mhaskar and E. B. Saff. Lecture Notes in Mathematics, 1287:176–179, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078905/.

Stahl:1987:EUR

[Sta87c] Herbert Stahl. Existence and uniqueness of rational interpolants with free and prescribed poles. Lecture Notes in Mathematics, 1287:180–208, 1987. CODEN LNMAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078906/.

Stahl:1987:TDA

[Sta87d] Herbert Stahl. Three different approaches to a proof of convergence for Padé approximants. Lecture Notes in Mathematics, 1237:79–124, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072458/.

Stasheff:1988:CP

[Sta88a] James Stasheff. Cohomological physics. Lecture Notes in Mathematics, 1318:228–237, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077806/.

Stasheff:1988:CHI

[Sta88b] James D. Stasheff. Constrained Hamiltonians an introduction to homological algebra in field theoretical physics. Lecture Notes in Mathematics, 1326:150–160, 1988. CODEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078044/.

Stein:1985:PTS

[Ste85a] E. Stein. The practical treatment of stress concentrations and singularities within finite element displacement algorithms. Lecture Notes in Mathematics, 1121:278–299, 1985. CODEN LN-MAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076276/.

Steinberg:1985:SEB

[Ste85b] J. Steinberg. Series expansions of biharmonic functions around a slit. Lecture Notes in Mathematics, 1121:300–307, 1985. CODEN LNMAA2. ISBN 3-540-15219-9 (print), 3-540-39377-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076277/.

Stern:1985:GEW

[Ste85c] Jacques Stern. Generic extensions which do not ADD random reals. Lecture Notes in Mathematics, 1130:395–407, 1985. CODEN LNMAA2. ISBN 3-540-15236-9 (print), 3-540-39414-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075321/.

Steinberger:1986:HOR

[Ste86a] Mark Steinberger. Homology operations for H_{∞} and H_{n} ring spectra. Lecture Notes in Mathematics, 1176:56–87, 1986. CODEN LNMAA2. ISBN 3-540-16434-0 (print), 3-540-39778-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075408/.

Stetter:1986:AIS

[Ste86b] H. J. Stetter. Algorithms for the inclusion of solutions of ordinary initial value problems. Lecture Notes in Mathematics, 1192: 85–94, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076054/.

Steenbrink:1987:SRA

[Ste87a] J. H. M. Steenbrink. Some remarks about the Hodge conjecture. Lecture Notes in Mathematics, 1246:165–175, 1987. CODEN LN-MAA2. ISBN 3-540-17743-4 (print), 3-540-47794-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077538/.

Stensones:1987:EHF

[Ste87b] Berit Stensønes. Extendability of holomorphic functions. Lecture Notes in Mathematics, 1268:189–195, 1987. CODEN LN-MAA2. ISBN 3-540-18094-X (print), 3-540-47752-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097303/.

Steel:1988:LG

[Ste88] John R. Steel. Long games. Lecture Notes in Mathematics, 1333: 56–97, 1988. CODEN LNMAA2. ISBN 3-540-50020-0 (print), 3-540-45896-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084970/.

Stein:1989:BCD

[Ste89a] Alan H. Stein. Binomial coefficients not divisible by a prime. Lecture Notes in Mathematics, 1383:170–177, 1989. CODEN LN-MAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083577/.

Sternfeld:1989:HPD

[Ste89b] Yaki Sternfeld. Hilbert’s 13th problem and dimension. Lecture Notes in Mathematics, 1376:1–49, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090047/.

Stolz:1985:KG

[Sto85a] Stephan Stolz. Die bP_{m}-Komponente von ∂: a_{m}^{⟨k⟩} → ⊖_{m−1}. (German) []. Lecture Notes in Mathematics, 1116:108–114, 1985. CO-DEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0101617/.

Stolz:1985:ASG

[Sto85b] Stephan Stolz. Die Adams-Spektralsequenz. (German) []. Lecture Notes in Mathematics, 1116:32–34, 1985. CODEN LN-MAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101609/.

Stolz:1985:BG

[Sto85c] Stephan Stolz. Die Bordismusgruppen A_{∗}^{⟨k⟩}. (German) []. Lecture Notes in Mathematics, 1116:1–9, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101605/.

Stolz:1985:KAG

[Sto85d] Stephan Stolz. Die Kohomologie von A[k]. (German) []. Lecture Notes in Mathematics, 1116:19–31, 1985. CODEN LN-MAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101608/.

Stolz:1985:KKG

[Sto85e] Stephan Stolz. Die Kokern J-Komponente von ∂: a _{m}^{⟨k⟩} → ⊖_{m−1}. (German) []. Lecture Notes in Mathematics, 1116:99–107, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101616/.

Stolz:1985:PTK

[Sto85f] Stephan Stolz. Die Pontrjagin–Thom Konstruktion. (German) [The Pontryagin–Thom construction]. Lecture Notes in Mathematics, 1116:10–13, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101606/.

Stolz:1985:DAS

[Sto85g] Stephan Stolz. Differentiale in der Adams-Spektralsequenz von A[k]. (German) []. Lecture Notes in Mathematics, 1116:87–98, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101615/.

Stolz:1985:TAS

[Sto85h] Stephan Stolz. Ein E_{1}-Term der Adams-Spektralsequenz von A[k]. (German) []. Lecture Notes in Mathematics, 1116:51–57, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101611/.

Stolz:1985:EMB

[Sto85i] Stephan Stolz. Eine elementare Methode zur Berechnung von Ext-Gruppen. (German) []. Lecture Notes in Mathematics, 1116: 35–50, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101610/.

Stolz:1985:GIE

[Sto85j] Stephan Stolz. Geometrische Invarianten für Elemente der Filtrierung null. (German) []. Lecture Notes in Mathematics, 1116: 58–67, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101612/.

Stolz:1985:GKG

[Sto85k] Stephan Stolz. Geometrische Konstruktionen. (German) [Geometrical constructions]. Lecture Notes in Mathematics, 1116:68– 80, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101613/.

Stolz:1985:HMI

[Sto85l] Stephan Stolz. Hochzusammenhängende Mannigfaltigkeiten und ihre Ränder. (German) [Highly-related manifolds and their boundaries], volume 1116 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 134 pp. LCCN QA3 .L28 no. 1116. URL http:

//link.springer.com/book/10.1007/BFb0101604;   http://
www.springerlink.com/content/978-3-540-39334-4.

Stolz:1985:KSA

[Sto85m] Stephan Stolz. Konstruktion des Spektrums A[k]. (German) []. Lecture Notes in Mathematics, 1116:14–18, 1985. CODEN LN-MAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101607/.

Stolz:1985:RA

[Sto85n] Stephan Stolz. Relationen in π_{∗} (a[k]). Lecture Notes in Mathematics, 1116:81–86, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101614/.

Stolz:1985:THM

[Sto85o] Stephan Stolz. Trägheitsgruppen hochzusammenhängender Mannigfaltigkeiten. (German) []. Lecture Notes in Mathematics, 1116: 119–130, 1985. CODEN LNMAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101619/.

Stolz:1985:DGH

[Sto85p] Stephan Stolz. Zur Diffeomorphieklassifikation geschlossener hochzusammenhängender Mannigfaltigkeiten. (German) []. Lecture Notes in Mathematics, 1116:115–118, 1985. CODEN LN-MAA2. ISBN 3-540-15209-1 (print), 3-540-39334-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101618/.

Stoll:1986:VDT

[Sto86] Wilhelm Stoll. Value distribution theory for moving targets. Lecture Notes in Mathematics, 1194:214–235, 1986. CODEN LN-MAA2. ISBN 3-540-16490-1 (print), 3-540-39829-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077005/.

Stoll:1987:SRR

[Sto87a] A. Stoll. Self-repellent random walks and polymer measures in two dimensions. Lecture Notes in Mathematics, 1250:298–318, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077363/.

Stoll:1987:ARN

[Sto87b] Wilhelm Stoll. Algebroid reduction of Nevanlinna theory. Lecture Notes in Mathematics, 1277:131–241, 1987. CODEN LN-MAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078248/.

Stonehewer:1987:SSF

[Sto87c] Stewart E. Stonehewer. Subnormal subgroups of factorised groups. Lecture Notes in Mathematics, 1281:158–175, 1987. CODEN LN-MAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078701/.

Strehl:1985:PHG

[Str85a] Volker Strehl. Polynômes d’Hermite généralisés et identités de Szegő — une version combinatoire. (French) []. Lecture Notes in Mathematics, 1171:129–138, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076538/.

Stricker:1985:RSC

[Str85b] C. Stricker. Une remarque sue une certaine classe de semimartingales. (French) []. Lecture Notes in Mathematics, 1123: 218–221, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075850/.

Stricker:1985:LSC

[Str85c] Christophe Stricker. Loi de semimartingales et critères de compacité. (French) []. Lecture Notes in Mathematics, 1123:209– 217, 1985. CODEN LNMAA2. ISBN 3-540-15230-X (print), 3-540-39397-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075849/.

Streater:1986:FSC

[Str86a] R. F. Streater. The fermion stochastic calculus I. Lecture Notes in Mathematics, 1158:245–257, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080223/.

Strehl:1986:CJC

[Str86b] Volker Strehl. Combinatorics of Jacobi configurations I: Complete oriented matchings. Lecture Notes in Mathematics, 1234: 294–307, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072522/.

Streit:1986:QTS

[Str86c] L. Streit. Quantum theory and stochastic processes — some contact points. Lecture Notes in Mathematics, 1203:197–213, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076882/.

Strelcyn:1986:PBS

[Str86d] J.-M. Strelcyn. Plane billiards as smooth dynamical systems with singularities. Lecture Notes in Mathematics, 1222:199–278, 1986. CODEN LNMAA2. ISBN 3-540-17190-8 (print), 3-540-47349-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099036/.

Stromme:1987:PRC

[Str87a] Stein Arild Strømme. On parametrized rational curves in Grassmann varieties. Lecture Notes in Mathematics, 1266:251–272, 1987. CODEN LNMAA2. ISBN 3-540-18020-6 (print), 3-540-47708-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078187/.

Stroock:1987:HCR

[Str87b] Daniel W. Stroock. Homogeneous chaos revisited. Lecture Notes in Mathematics, 1247:1–7, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077624/.

Stricker:1988:PCM

[Str88a] Christophe Stricker. À propos d’une conjecture de Meyer. (French) []. Lecture Notes in Mathematics, 1321:144–146, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084130/.

Struwe:1988:AVM

[Str88b] Michael Struwe. Applications of variational methods to problems in the geometry of surfaces. Lecture Notes in Mathematics, 1357:359–378, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082875/.

Strade:1989:ATR

[Str89a] Helmut Strade. The absolute toral rank of a Lie algebra. Lecture Notes in Mathematics, 1373:1–28, 1989. CODEN LNMAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088884/.

Strottman:1989:RHH

[Str89b] D. Strottman. Relativistic hydrodynamics and heavy ion reactions. Lecture Notes in Mathematics, 1385:278–289, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084035/.

Sturm:1987:CMF

[Stu87] Jacob Sturm. On the congruence of modular forms. Lecture Notes in Mathematics, 1240:275–280, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072985/.

Sylvester:1987:RIB

[SU87] John Sylvester and Gunther Uhlmann. Remarks on an inverse boundary value problem. Lecture Notes in Mathematics, 1256: 430–441, 1987. CODEN LNMAA2. ISBN 3-540-17856-2 (print), 3-540-47886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077754/.

Sullivan:1986:CCR

[Sul86] Wayne G. Sullivan. L^{2} convergence of certain random walks on Z^{d} and related diffusions. Lecture Notes in Mathematics, 1212: 249–257, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076253/.

Suljagic:1987:ISS

[Sul87] Salih Suljagić. Invariant subspaces of shifts in quaternionic Hilbert space. Lecture Notes in Mathematics, 1242:368–373, 1987. CO-DEN LNMAA2. ISBN 3-540-17833-3 (print), 3-540-47876-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0072444/.

Sunada:1986:FGS

[Sun86] Toshikazu Sunada. L-functions in geometry and some applications. Lecture Notes in Mathematics, 1201:266–284, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075662/.

Sunada:1987:NTA

[Sun87] Toshikazu Sunada. Number theoretic analogues in spectral geometry. Lecture Notes in Mathematics, 1255:96–108, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077685/.

Sunada:1988:FGL

[Sun88] Toshikazu Sunada. Fundamental groups and Laplacians. Lecture Notes in Mathematics, 1339:248–277, 1988. CODEN LN-MAA2. ISBN 3-540-50113-4 (print), 3-540-45930-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083059/.

Smith:1985:CLS

[SV85] Roy Smith and Robert Varley. Components of the locus of singular theta divisors of genus 5. Lecture Notes in Mathematics, 1124:338–416, 1985. CODEN LNMAA2. ISBN 3-540-15232-6 (print), 3-540-39643-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075005/.

Svec:1986:SPC

[Šve86] M. Švec. Some problems concerning the equivalences of two systems of differential equations. Lecture Notes in Mathematics, 1192:171–179, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076065/.

Sanchez-Valenzuela:1987:AGH

[SVS87] Oscar A. Sánchez-Valenzuela and Shlomo Sternberg. The automorphism group of a Hermitian superalgebra. Lecture Notes in Mathematics, 1251:1–48, 1987. CODEN LNMAA2. ISBN 3-540-17816-3 (print), 3-540-47854-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077313/.

Salinetti:1986:HCP

[SW86a] G. Salinetti and R. Wets. On the hypo-convergence of probability measures. Lecture Notes in Mathematics, 1190:371–395, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076715/.

Sorsdal:1986:CCC

[SW86b] Ellen Sørsdal and Haakon Waadeland. On the convergence of a certain class of continued fractions K(a_{n}/1) with a_{n} → ∞. Lecture Notes in Mathematics, 1199:285–293, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075943/.

Sasvari:1987:CDR

[SW87] Z. Sasvári and W. Wolf. Characterizing the distributions of the random vectors X_{1}, X_{2}, X_{3} by the distribution of the statistic (X_{1} − X_{3}, X_{2} − X_{3}). Lecture Notes in Mathematics, 1233:172– 177, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072722/.

Serconek:1989:FRS

[SW89a] Shirlei Serconek and Robert Lee Wilson. Forms of restricted simple Lie algebras. Lecture Notes in Mathematics, 1373:67–80, 1989. CODEN LNMAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088888/.

Slaman:1989:CFI

[SW89b] Theodore A. Slaman and W. Hugh Woodin. Σ_{1}-collection and the finite injury priority method. Lecture Notes in Mathematics, 1388:178–188, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083670/.

Sternfeld:1989:ATV

[SW89c] Yaki Sternfeld and Yitzhak Weit. An approximation theorem for vector valued functions. Lecture Notes in Mathematics, 1376:126– 137, 1989. CODEN LNMAA2. ISBN 3-540-51303-5 (print), 3-540-46189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090052/.

Sweedler:1985:IAT

[Swe85] M. Sweedler. Introduction to the algebraic theory of positive characteristic differential geometry. Lecture Notes in Mathematics, 1146:317–324, 1985. CODEN LNMAA2. ISBN 3-540-15686-0 (print), 3-540-39628-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074544/.

Sweedler:1989:PCC

[Swe89] Moss E. Sweedler. Positive characteristic calculus and icebergs. Lecture Notes in Mathematics, 1383:178–185, 1989. CODEN LN-MAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083578/.

Song:1987:IPP

[SY87] S. Song and M. Yor. Inégalités pour les processus self-similaires arrêtés à un temps quelconque. (French) []. Lecture Notes in Mathematics, 1247:230–245, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077638/.

Suwa:1989:ACA

[SY89] Noriyuki Suwa and Noriko Yui. Arithmetic of certain algebraic surfaces over finite fields. Lecture Notes in Mathematics, 1383: 186–256, 1989. CODEN LNMAA2. ISBN 3-540-51549-6 (print), 3-540-46640-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083579/.

Schwarz:1985:GPM

[SZ85] Binyamin Schwarz and Abraham Zaks. Geometries of the projective matrix space. Lecture Notes in Mathematics, 1165:261– 278, 1985. CODEN LNMAA2. ISBN 3-540-16050-7 (print), 3-540-39734-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076159/.

Steenbrink:1987:PCR

[SZ87] Joseph Steenbrink and Steven Zucker. Polar curves, resolution of singularities and the filtered mixed Hodge structure on the vanishing cohomology. Lecture Notes in Mathematics, 1273:178– 202, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078844/.

Szekely:1986:HMR

[Szé86] Gábor J. Székely. Haar measures in a representation and a decomposition problem. Lecture Notes in Mathematics, 1210:158– 162, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077182/.

Szeidl:1987:EMR

[Sze87] László Szeidl. On the estimation of moments of regenerative cycles in a general closed central-server queueing network. Lecture Notes in Mathematics, 1233:182–189, 1987. CODEN LNMAA2. ISBN 3-540-17204-1 (print), 3-540-47394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072724/.

Szpirglas:1986:QRS

[Szp86] Aviva Szpirglas. Quelques résultats sur le schéma de Hilbert de C^{P} des sous-ensembles analytiques compacts de dimension O de C^{P}. (French) []. Lecture Notes in Mathematics, 1188:229–231, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076824/.

Tabib:1986:AIE

[Tab86] Claudette Tabib. About the inequalities of Erdős and Moser on the largest transitive subtournament of a tournament. Lecture Notes in Mathematics, 1234:308–320, 1986. CODEN LNMAA2. ISBN 3-540-17207-6 (print), 3-540-47402-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072523/.

Tachikawa:1986:SAT

[Tac86] Hiroyuki Tachikawa. Selfinjective algebras and tilting theory. Lecture Notes in Mathematics, 1177:272–307, 1986. CODEN LN-MAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075270/.

Tadmor:1987:EDN

[Tad87] Eitan Tadmor. The entropy dissipation by numerical viscosity in nonlinear conservative difference schemes. Lecture Notes in Mathematics, 1270:52–63, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078317/.

Tahara:1986:FHP

[Tah86] Hidetoshi Tahara. On Fuchsian hyperbolic partial differential equations. Lecture Notes in Mathematics, 1223:243–253, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099197/.

Takens:1985:NDD

[Tak85] Floris Takens. On the numerical determination of the dimension of an attractor. Lecture Notes in Mathematics, 1125:99–106, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075637/.

Takens:1988:IGA

[Tak88a] Floris Takens. Intermittancy: Global aspects. Lecture Notes in Mathematics, 1331:213–239, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083075/.

Takens:1988:LCH

[Tak88b] Floris Takens. Limit capacity and Hausdorff dimension of dynamically defined Cantor sets. Lecture Notes in Mathematics, 1331:196–212, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083074/.

Talenti:1986:IPL

[Tal86a] Giorgio Talenti, editor. Inverse Problems: Lectures given at the 1st 1986 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Montecatini Terme, Italy, May 28– June 5, 1986, volume 1225 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17193-2 (print), 3-540-47353-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1225. URL http:

//link.springer.com/book/10.1007/BFb0072658;   http://
www.springerlink.com/content/978-3-540-47353-4.

Talenti:1986:RFP

[Tal86b] Giorgio Talenti. Rearrangements of functions and partial differential equations. Lecture Notes in Mathematics, 1224:153–178, 1986. CODEN LNMAA2. ISBN 3-540-17192-4 (print), 3-540-47352-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072690/.

Tall:1989:TPS

[Tal89] Franklin D. Tall. Topological problems for set-theorists. Lecture Notes in Mathematics, 1401:194–200, 1989. CODEN LN-MAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097340/.

Tanaka:1987:LDD

[Tan87] H. Tanaka. Limit distribution for 1-dimensional diffusion in a reflected Brownian medium. Lecture Notes in Mathematics, 1247: 246–261, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077639/.

Tan:1988:IPN

[Tan88a] Yongji Tan. An inverse problem for nonlocal elliptic bvp and resistivity identification. Lecture Notes in Mathematics, 1306:149– 159, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082931/.

Tanaka:1988:LTO

[Tan88b] Hiroshi Tanaka. Limit theorem for one-dimensional diffusion process in Brownian environment. Lecture Notes in Mathematics, 1322:156–172, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077874/.

Tautu:1986:SSPa

[Tău86a] P. Tăutu. Stochastic spatial processes in biology: a concise historical survey. Lecture Notes in Mathematics, 1212:1–41, 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076235/.

Tautu:1986:RFA

[Tău86b] Petre Tăutu. Random fields: Applications in cell biology. Lecture Notes in Mathematics, 1212:258–299, 1986. CODEN LN-MAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076254/.

Tautu:1986:SSPb

[Tau86c] Petre Tautu, editor. Stochastic Spatial Processes: Mathematical Theories and Biological Applications Proceedings of a Conference held in Heidelberg, September 10–14, 1984, volume 1212 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16803-6 (print), 3-540-47053-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 1212. URL http://link.springer.com/book/10.1007/
BFb0076234;   http://www.springerlink.com/content/978-
3-540-47053-3.

Taylor:1985:RGM

[Tay85] M. J. Taylor. Relative Galois module structure of rings of integers. Lecture Notes in Mathematics, 1142:289–306, 1985. CODEN LNMAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074807/.

Taylor:1986:UPM

[Tay86] S. James Taylor. The use of packing measure in the analysis of random sets. Lecture Notes in Mathematics, 1203:214–222, 1986. CODEN LNMAA2. ISBN 3-540-16773-0 (print), 3-540-39852-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076883/.

Taylor:1988:LER

[Tay88] Jean E. Taylor. Local ellipticity of F and regularity of F minimizing currents. Lecture Notes in Mathematics, 1306:160–173, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082932/.

Taqqu:1985:RKH

[TC85] Murad S. Taqqu and Claudia Czado. Reproducing kernel Hilbert space for some non-Gaussian processes. Lecture Notes in Mathematics, 1153:128–140, 1985. CODEN LNMAA2. ISBN 3-540-15704-2 (print), 3-540-39645-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074948/.

Chuang:1987:TAN

[tC87] Chi tai Chuang. On the theory and applications of Q_{m}-normal families of meromorphic functions. Lecture Notes in Mathematics, 1275:44–77, 1987. CODEN LNMAA2. ISBN 3-540-18356-6 (print), 3-540-47899-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078344/.

tomDieck:1989:LCR

[tD89a] Tammo tom Dieck. Linking in cyclic representation forms. Lecture Notes in Mathematics, 1375:42–47, 1989. CODEN LN-MAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085598/.

tomDieck:1989:MLR

[tD89b] Tammo tom Dieck. Modification of linking in representation forms. Lecture Notes in Mathematics, 1375:33–41, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085597/.

tomDieck:1985:VFD

[tDL85] Tammo tom Dieck and Peter Löffler. Verschlingung von Fixpunktmengen in Darstellungsformen. I. (German) []. Lecture Notes in Mathematics, 1172:167–187, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074431/.

tomDieck:1986:VFD

[tDL86] Tammo tom Dieck and Peter Löffler. Verschlingungszahlen von Fixpunktmengen in Darstellungsformen. II. (German) []. Lecture Notes in Mathematics, 1217:84–91, 1986. CODEN LN-MAA2. ISBN 3-540-16824-9 (print), 3-540-47097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072816/.

tomDieck:1989:AMP

[tDP89] Tammo tom Dieck and Ted Petrie. The Abhyankar–Moh problem in dimension 3. Lecture Notes in Mathematics, 1375:48–59, 1989. CODEN LNMAA2. ISBN 3-540-51218-7 (print), 3-540-46178-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085599/.

Thi:1986:MPV

[Thi86] – Dào Trong Thi. Multidimensional parametrized variational problems on Riemannian manifolds. Lecture Notes in Mathematics, 1214:40–62, 1986. CODEN LNMAA2. ISBN 3-540-16821-4 (print), 3-540-47084-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075958/.

Thomson:1985:CS

[Tho85a] Brian S. Thomson. Cluster sets. Lecture Notes in Mathematics, 1170:44–69, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074382/.

Thomson:1985:C

[Tho85b] Brian S. Thomson. Continuity. Lecture Notes in Mathematics, 1170:70–84, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074383/.

Thomson:1985:DYR

[Tho85c] Brian S. Thomson. The Denjoy–Young relations. Lecture Notes in Mathematics, 1170:166–182, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074387/.

Thomson:1985:LS

[Tho85d] Brian S. Thomson. Local systems. Lecture Notes in Mathematics, 1170:1–43, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074381/.

Thomson:1985:M

[Tho85e] Brian S. Thomson. Monotonicity. Lecture Notes in Mathematics, 1170:115–137, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074385/.

Thomson:1985:RF

[Tho85f] Brian S. Thomson. Real Functions, volume 1170 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 229 pp. LCCN QA3

.L28 no. 1170. URL http://link.springer.com/book/10.1007/
BFb0074380;   http://www.springerlink.com/content/978-
3-540-39742-7.

Thomson:1985:RAD

[Tho85g] Brian S. Thomson. Relations among derivates. Lecture Notes in Mathematics, 1170:138–165, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074386/.

Thomson:1985:VF

[Tho85h] Brian S. Thomson. Variation of a function. Lecture Notes in Mathematics, 1170:85–114, 1985. CODEN LNMAA2. ISBN 3-540-16058-2 (print), 3-540-39742-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074384/.

Thomee:1986:EEF

[Tho86a] V. Thomée. Error estimates for finite element methods for semilinear parabolic problems with nonsmooth data. Lecture Notes in Mathematics, 1192:339–344, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076091/.

Thompson:1986:SFG

[Tho86b] John G. Thompson. Some finite groups which appear as galL/K, where K ⊂ Q(μ_{n}). Lecture Notes in Mathematics, 1185:210– 230, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076176/.

Thomas:1987:TIS

[Tho87] Pascal J. Thomas. Tents and interpolating sequences in the unit ball. Lecture Notes in Mathematics, 1276:291–302, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078965/.

Thom:1988:BFF

[Tho88] R. Thom. Sur les bouts d’une feuille d’un feuilletage au voisinage d’un point singulier isolé. (French) []. Lecture Notes in Mathematics, 1345:317–321, 1988. CODEN LNMAA2. ISBN 3-540-50226-2 (print), 3-540-45957-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081410/.

Thron:1986:ATC

[Thr86] Wolfgang Joseph Thron, editor. Analytic Theory of Continued Fractions II: Proceedings of a Seminar–Workshop held in Pitlochry and Aviemore, Scotland June 13–29, 1985, volume 1199 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LN-MAA2. ISBN 0-387-16768-4 , 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1199; QA3 .L28 no. 1199 QA295. URL http:

//link.springer.com/book/10.1007/BFb0075930;   http://
www.springerlink.com/content/978-3-540-38817-3.

Thron:1989:CFI

[Thr89] W. J. Thron. Continued fraction identities derived from the invariance of the crossratio under l.f.t. Lecture Notes in Mathematics, 1406:124–134, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096170/.

Tichy:1985:BEL

[Tic85] Robert F. Tichy. Bemerkung Zu Einem Lemma Aus Der Variationsrechnung. (German) []. Lecture Notes in Mathematics, 1114: 154–157, 1985. CODEN LNMAA2. ISBN 3-540-15189-3 (print), 3-540-39281-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101654/.

Tichy:1987:EBS

[Tic87] Robert F. Tichy. Einige Bemerkungen über Stetige Funktionen auf Topologischen Gruppen. (German) []. Lecture Notes in Mathematics, 1262:139–143, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078603/.

Tietavainen:1988:IST

[Tie88] A. Tietäväinen. Incomplete sums and two applications of Deligne’s result. Lecture Notes in Mathematics, 1352:190–205, 1988. CO-DEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082026/.

Timmesfeld:1986:BM

[Tim86a] F. G. Timmesfeld. Back matter. Lecture Notes in Mathematics, 1181:270–279, 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-39801-1/1.pdf.

Timmesfeld:1986:TCS

[Tim86b] F. G. Timmesfeld. Tits chamber systems and finite group theory. Lecture Notes in Mathematics, 1181:249–269, 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075519/.

Timmermans:1988:SSB

[Tim88] C. A. Timmermans. On C_{0}-semigroups in a space of bounded continuous functions in the case of entrance or natural boundary points. Lecture Notes in Mathematics, 1354:209–216, 1988. CODEN LNMAA2. ISBN 3-540-50443-5 (print), 3-540-46005-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089598/.

Tirapegui:1988:NFD

[Tir88] Enrique Tirapegui. Normal forms for deterministic and stochastic systems. Lecture Notes in Mathematics, 1331:240–250, 1988. CODEN LNMAA2. ISBN 3-540-50016-2 (print), 3-540-45889-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083076/.

Tits:1985:GGF

[Tit85] Jacques Tits. Groups and group functors attached to Kac–Moody data. Lecture Notes in Mathematics, 1111:193–223, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084591/.

Tits:1986:ITA

[Tit86] Jacques Tits. Immeubles de type affine. (French) []. Lecture Notes in Mathematics, 1181:159–190, 1986. CODEN LNMAA2. ISBN 3-540-16466-9 (print), 3-540-39801-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075514/.

Tits:1987:UEP

[Tit87] Jacques Tits. Unipotent elements and parabolic subgroups of reductive groups. II. Lecture Notes in Mathematics, 1271:265– 284, 1987. CODEN LNMAA2. ISBN 3-540-18234-9 (print), 3-540-47834-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079243/.

Thorup:1988:CBF

[TK88] Anders Thorup and Steven Kleiman. Complete bilinear forms. Lecture Notes in Mathematics, 1311:253–320, 1988. CODEN LN-MAA2. ISBN 3-540-19236-0 (print), 3-540-39157-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082918/.

Truman:1986:SMG

[TL86] A. Truman and J. T. Lewis. The stochastic mechanics of the ground-state of the hydrogen atom. Lecture Notes in Mathematics, 1158:168–179, 1986. CODEN LNMAA2. ISBN 3-540-15998-3 (print), 3-540-39703-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080216/.

Li:1988:GPR

[tLcZ88] Ta tsien Li and Yan chun Zhao. Global perturbation of the Riemann problem for the system of one-dimensional isentropic flow. Lecture Notes in Mathematics, 1306:131–140, 1988. CODEN LN-MAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082929/.

Li:1989:GCS

[tLmC89] Ta tsien Li and Yun mei Chen. Global classical solutions to the Cauchy problem for nonlinear wave equations. Lecture Notes in Mathematics, 1402:198–202, 1989. CODEN LNMAA2. ISBN 3-540-51746-4 (print), 3-540-46800-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083876/.

Li:1987:GSS

[tLqL87] Ta tsien Li and Da qian Li. Global solutions to some free boundary problems for quasilinear hyperbolic systems and applications. Lecture Notes in Mathematics, 1270:195–209, 1987. CODEN LN-MAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078329/.

Tonev:1987:NRB

[Ton87a] T. V. Tonev. New relations between Sibony–Basener boundaries. Lecture Notes in Mathematics, 1277:256–262, 1987. CODEN LN-MAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078250/.

Tonev:1987:IDG

[Ton87b] Toma V. Tonev. An infinite-dimensional generalization of the Shilov boundary and infinite dimensional analytic structures in the spectrum of a uniform algebra. Lecture Notes in Mathematics, 1277:242–255, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078249/.

Tong:1989:EW

[Ton89] Deyu Tong. On an example of Weier. Lecture Notes in Mathematics, 1411:171–181, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086452/.

Torrecillas:1986:HRT

[Tor86] Blas Torrecillas. Height relative to a torsion theory. Lecture Notes in Mathematics, 1197:178–184, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076324/.

Torres:1987:DST

[Tor87a] Felipe Cano Torres. Desingularization Strategies for Three-Dimensional Vector Fields, volume 1259 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17944-5 (print), 3-540-47173-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 194 pp. LCCN QA3 .L28

no. 1259. URL http://link.springer.com/book/10.1007/
BFb0077952;   http://www.springerlink.com/content/978-
3-540-47173-8.

Torres:1987:PWS

[Tor87b] Felipe Cano Torres. A partial winning strategy. Lecture Notes in Mathematics, 1259:35–78, 1987. CODEN LNMAA2. ISBN 3-540-17944-5 (print), 3-540-47173-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077954/.

Torres:1987:RSV

[Tor87c] Felipe Cano Torres. Resolution statements for a vector field. Lecture Notes in Mathematics, 1259:1–34, 1987. CODEN LN-MAA2. ISBN 3-540-17944-5 (print), 3-540-47173-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077953/.

Torres:1987:STT

[Tor87d] Felipe Cano Torres. Standard transitions from type I. Lecture Notes in Mathematics, 1259:79–115, 1987. CODEN LN-MAA2. ISBN 3-540-17944-5 (print), 3-540-47173-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077955/.

Torres:1987:TTT

[Tor87e] Felipe Cano Torres. Types two and three. Lecture Notes in Mathematics, 1259:172–185, 1987. CODEN LNMAA2. ISBN 3-540-17944-5 (print), 3-540-47173-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077957/.

Torres:1987:WST

[Tor87f] Felipe Cano Torres. A winning strategy for type one. Lecture Notes in Mathematics, 1259:116–171, 1987. CODEN LNMAA2. ISBN 3-540-17944-5 (print), 3-540-47173-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077956/.

Tortrat:1987:API

[Tor87g] Pierre Tortrat. Aspects potentialistes de l’iteration des polynômes. (French) []. Lecture Notes in Mathematics, 1235:195–209, 1987. CODEN LNMAA2. ISBN 3-540-17210-6 (print), 3-540-47406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072757/.

Trench:1986:LPG

[Tre86] W. F. Trench. Linear perturbations of general disconjugate equations. Lecture Notes in Mathematics, 1192:181–186, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076066/.

Tretkoff:1987:PFD

[Tre87] Marvin Tretkoff. Picard–Fuchs differential equations for the quadratic periods of Abelian integrals of the first kind. Lecture Notes in Mathematics, 1277:263–289, 1987. CODEN LN-MAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078251/.

Triebel:1986:RDT

[Tri86] H. Triebel. Recent developments in the theory of function spaces. Lecture Notes in Mathematics, 1192:95–106, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076055/.

Tromba:1988:OPD

[Tro88] A. J. Tromba. Open problems in the degree theory for disc minimal surfaces spanning a curve in R^{3}. Lecture Notes in Mathematics, 1357:379–401, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082876/.

Tromba:1989:CVA

[Tro89a] A. J. Tromba. A classical variational approach to Teichmüller theory. Lecture Notes in Mathematics, 1365:155–185, 1989. CODEN LNMAA2. ISBN 3-540-50727-2 (print), 3-540-46075-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089181/.

Tromba:1989:FM

[Tro89b] A. J. Tromba. Front matter. Lecture Notes in Mathematics, 1369: ??, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46137-1/1.pdf.

Trudinger:1986:CBV

[Tru86] N. S. Trudinger. Classical boundary value problems for Monge– Ampère type equations. Lecture Notes in Mathematics, 1192:251– 258, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076078/.

Tschantz:1985:MCE

[Tsc85] Steven T. Tschantz. More conditions equivalent to congruence modularity. Lecture Notes in Mathematics, 1149:270–282, 1985. CODEN LNMAA2. ISBN 3-540-15691-7 (print), 3-540-39638-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098470/.

Tomi:1988:GPM

[TT88] Friedrich Tomi and A. J. Tromba. A geometric proof of the Mumford compactness theorem. Lecture Notes in Mathematics, 1306:174–181, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082933/.

Tuan:1986:GTB

[Tua86] Hsio-Fu Tuan, editor. Group Theory, Beijing 1984: Proceedings of an International Symposium held in Beijing, Aug. 27–September 8, 1984, volume 1185 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 0-387-16456-1, 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1185. URL http:

//link.springer.com/book/10.1007/BFb0076169;   http://
www.springerlink.com/content/978-3-540-39793-9.

Turner:1985:NAL

[Tur85] Peter R. Turner, editor. Numerical Analysis Lancaster 1984: Proceedings of the SERC Summer School held in Lancaster, England, Jul. 15–Aug. 3, 1984, volume 1129 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15234-2 (print), 3-540-39412-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1129. URL http:

//link.springer.com/book/10.1007/BFb0075155;   http://
www.springerlink.com/content/978-3-540-39412-9.

Turnwald:1987:EBD

[Tur87] Gerhard Turnwald. Einige Bemerkungen zur Diskreten Gleichverteilung. (German) []. Lecture Notes in Mathematics, 1262: 144–149, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078604/.

Turaev:1988:COM

[Tur88a] V. G. Turaev. Classification of oriented Montesinos links via spin structures. Lecture Notes in Mathematics, 1346:271–289, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082779/.

Turaev:1988:TTC

[Tur88b] V. G. Turaev. Towards the topological classification of geometric 3-manifolds. Lecture Notes in Mathematics, 1346:291–323, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082780/.

Turnwald:1989:WUD

[Tur89] Gerhard Turnwald. Weak uniform distribution of second-order linear recurring sequences. Lecture Notes in Mathematics, 1380: 242–253, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086556/.

Tvrdy:1986:OCS

[Tvr86] M. Tvrdý. On optimal control of systems with interface side conditions. Lecture Notes in Mathematics, 1192:187–190, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076067/.

Tachikawa:1986:ARF

[TW86] Hiroyuki Tachikawa and Takayoshi Wakamatsu. Applications of reflection functors for selfinjective algebras. Lecture Notes in Mathematics, 1177:308–327, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075271/.

Wu:1987:CTD

[tW87a] Wen tsun Wu. A constructive theory of differential algebraic geometry based on works of J. F. Ritt with particular applications to mechanical theorem-proving of differential geometries. Lecture Notes in Mathematics, 1255:173–189, 1987. CODEN LN-MAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077689/.

Wu:1987:MHS

[tW87b] Wen tsün Wu. I*-measure of a homogeneous space — the Cartan theorem. Lecture Notes in Mathematics, 1264:116–142, 1987.

CODEN LNMAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082002/.

Wu:1987:RHT

[tW87c] Wen tsün Wu. Rational Homotopy Type: A Constructive Study via the Theory of the I*-measure, volume 1264 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 219 pp. LCCN QA3

.L28 no. 1264. URL http://link.springer.com/book/10.1007/
BFb0081997;   http://www.springerlink.com/content/978-
3-540-39025-1.

Ulrich:1988:EVE

[Ulr88a] Hanno Ulrich. Equivariant vertical Euclidean neighbourhood retracts. Lecture Notes in Mathematics, 1343:21–59, 1988. CODEN LNMAA2. ISBN 3-540-50187-8 (print), 3-540-45950-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079802/.

Ulrich:1988:FPI

[Ulr88b] Hanno Ulrich. The fixed point index of equivariant vertical maps. Lecture Notes in Mathematics, 1343:61–136, 1988. CODEN LN-MAA2. ISBN 3-540-50187-8 (print), 3-540-45950-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079803/.

Ulrich:1988:I

[Ulr88c] Hanno Ulrich. Introduction. Lecture Notes in Mathematics, 1343: 1–6, 1988. CODEN LNMAA2. ISBN 3-540-50187-8 (print), 3-540-45950-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079800/.

Ulrich:1988:PGA

[Ulr88d] Hanno Ulrich. Preliminaries on group actions. Lecture Notes in Mathematics, 1343:7–20, 1988. CODEN LNMAA2. ISBN 3-540-50187-8 (print), 3-540-45950-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079801/.

Unger:1986:PCT

[Ung86] Luise Unger. Preinjective components of trees. Lecture Notes in Mathematics, 1177:328–339, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075272/.

Urakawa:1986:SHM

[Ura86] Hajime Urakawa. Stability of harmonic maps and eigenvalues of Laplacian. Lecture Notes in Mathematics, 1201:285–307, 1986. CODEN LNMAA2. ISBN 3-540-16770-6 (print), 3-540-38827-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075663/.

Urbanik:1986:CMM

[Urb86] K. Urbanik. Compactness, medians and moments. Lecture Notes in Mathematics, 1210:163–173, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077183/.

Ustunel:1987:SAM

[Üst87] A. S. Üstünel. Some applications of the Malliavin calculus to stochastic analysis. Lecture Notes in Mathematics, 1236:230– 238, 1987. CODEN LNMAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072893/.

Ustunel:1988:SCF

[Üst88] A. S. Üstünel. Some comments on the filtering of diffusions and the Malliavin calculus. Lecture Notes in Mathematics, 1316:247– 266, 1988. CODEN LNMAA2. ISBN 3-540-19315-4 (print), 3-540-39186-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081934/.

Uemura:1988:DPH

[UW88] Hideaki Uemura and Shinzo Watanabe. Diffusion processes and heat kernels on certain nilpotent groups. Lecture Notes in Mathematics, 1322:173–197, 1988. CODEN LNMAA2. ISBN 3-540-19352-9 (print), 3-540-39232-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077875/.

Valette:1986:PTA

[Val86a] J. Valette. Polynômes tordus et T -anneaux à gauche. (French) []. Lecture Notes in Mathematics, 1197:185–189, 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076325/.

Valli:1986:QPS

[Val86b] A. Valli. Qualitative properties of the solutions to the Navier– Stokes equations for compressible fluids. Lecture Notes in Mathematics, 1192:259–264, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076079/.

VanBuskirk:1985:PKP

[Van85] James M. Van Buskirk. Positive knots have positive Conway polynomials. Lecture Notes in Mathematics, 1144:146–159, 1985. CODEN LNMAA2. ISBN 3-540-15680-1 (print), 3-540-39616-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075018/.

VandenBergh:1986:AID

[Van86a] M. Van den Bergh. The algebraic index of a division algebra. Lecture Notes in Mathematics, 1197:190–206, 1986. CODEN LN-MAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076326/.

VanPraag:1986:LLC

[Van86b] Paul Van Praag. L’élimination linéaire dans les corps gauches. (French) []. Lecture Notes in Mathematics, 1220:93–108, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099507/.

VanAssche:1987:APR

[Van87a] Walter Van Assche. Asymptotically periodic recurrence coefficients. Lecture Notes in Mathematics, 1265:43–86, 1987. CODEN LNMAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081883/.

VanAssche:1987:AOP

[Van87b] Walter Van Assche. Asymptotics for Orthogonal Polynomials, volume 1265 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 0-387-18023-0 (paperback), 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 201 pp. LCCN QA3 .L28 no. 1265; QA3 .L28 no. 1265; QA404.5. URL http:

//link.springer.com/book/10.1007/BFb0081880;   http://
www.springerlink.com/content/978-3-540-47711-2.

VanAssche:1987:I

[Van87c] Walter Van Assche. Introduction. Lecture Notes in Mathematics, 1265:1–13, 1987. CODEN LNMAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081881/.

VanAssche:1987:OPC

[Van87d] Walter Van Assche. Orthogonal polynomials on a compact set. Lecture Notes in Mathematics, 1265:14–42, 1987. CODEN LN-MAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081882/.

VanAssche:1987:OPU

[Van87e] Walter Van Assche. Orthogonal polynomials on unbounded sets. Lecture Notes in Mathematics, 1265:105–137, 1987. CODEN LN-MAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081885/.

VanAssche:1987:PPA

[Van87f] Walter Van Assche. Probabilistic proofs of asymptotic formulas. Lecture Notes in Mathematics, 1265:87–104, 1987. CODEN LNMAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081884/.

VanAssche:1987:SA

[Van87g] Walter Van Assche. Some applications. Lecture Notes in Mathematics, 1265:164–174, 1987. CODEN LNMAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081887/.

VanAssche:1987:ZDC

[Van87h] Walter Van Assche. Zero distribution and consequences. Lecture Notes in Mathematics, 1265:138–163, 1987. CODEN LN-MAA2. ISBN 3-540-18023-0 (print), 3-540-47711-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081886/.

VandenBergh:1987:RRD

[Van87i] Michel Van den Bergh. Regular rings of dimension three. Lecture Notes in Mathematics, 1296:228–234, 1987. CODEN LN-MAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078529/.

Varouchas:1986:PCC

[Var86a] J. Varouchas. Propriétés cohomologiques d’une classe de variétés analytiques complexes compactes. (French) []. Lecture Notes in Mathematics, 1198:233–243, 1986. CODEN LNMAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077057/.

Varouchas:1986:IVK

[Var86b] J. Varouchas. Sur l’image d’une variété kählérienne compacte. (French) []. Lecture Notes in Mathematics, 1188:245–259, 1986. CODEN LNMAA2. ISBN 3-540-16460-X (print), 3-540-39797-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076826/.

Varadhan:1988:LDA

[Var88a] Srinivasa R. S. Varadhan. Large deviations and applications. Lecture Notes in Mathematics, 1362:1–49, 1988. CODEN LN-MAA2. ISBN 3-540-50549-0 (print), 3-540-46042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086178/.

Varchenko:1988:IPM

[Var88b] A. N. Varchenko. Image of period mapping for simple singularities. Lecture Notes in Mathematics, 1334:294–308, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080434/.

Varea:1989:SLS

[Var89] Vicente R. Varea. The subalgebra lattice of a supersolvable Lie algebra. Lecture Notes in Mathematics, 1373:81–92, 1989. CODEN LNMAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088889/.

Vassilevski:1986:MMS

[Vas86] Panayot Vassilevski. Multigrid method in subspace and domain partitioning in the discrete solution of elliptic problems. Lecture Notes in Mathematics, 1228:301–314, 1986. CODEN LN-MAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072654/.

vandenBerg:1987:AL

[vdB87a] Imme van den Berg. Approximation lemmas. Lecture Notes in Mathematics, 1249:142–151, 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077582/.

vandenBerg:1987:AERb

[vdB87b] Imme van den Berg. Asymptotic expressions for the remainders

associated to expansions of type ∑_{n=0}^{∞} c_{n} \frac{z^{n}}{n!} , ∑_{n=0}^{∞} c_{n}z^{n}and ∑_{n=0c_{n}n!z^{n}}:

Critical regions, uniform behaviour. Lecture Notes in Mathematics, 1249:72–103, 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077580/.

vandenBerg:1987:AERa

[vdB87c] Imme van den Berg. Asymptotic expressions for the remainders

associated to expansions of type ∑_{n=0}^{∞} c_{n} \frac{z^{n}}{n!} , ∑_{n=0}^{∞} c_{n}z^{n}and ∑_{n=0c_{n}n!z^{n}},

where c_{n+p}/c_{n} → c. Lecture Notes in Mathematics, 1249:46– 71, 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077579/.

vandenBerg:1987:ES

[vdB87d] Imme van den Berg. External sets. Lecture Notes in Mathematics, 1249:106–141, 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077581/.

vandenBerg:1987:FEN

[vdB87e] Imme van den Berg. Four examples of nonstandard reasoning in asymptotics. Lecture Notes in Mathematics, 1249:2–43, 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077578/.

vandenBerg:1987:NAA

[vdB87f] Imme van den Berg. Nonstandard Asymptotic Analysis, volume 1249 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 187 pp. LCCN QA3 .L28 no. 1249. URL http:

//link.springer.com/book/10.1007/BFb0077577;   http://
www.springerlink.com/content/978-3-540-47810-2.

vandenBerg:1987:SE

[vdB87g] Imme van den Berg. Shadow expansions. Lecture Notes in Mathematics, 1249:152–176, 1987. CODEN LNMAA2. ISBN 3-540-17767-1 (print), 3-540-47810-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077583/.

vandenBerg:1988:LTS

[vdBL88] M. van den Berg and J. T. Lewis. Limit theorems for stochastic processes associated with a boson gas. Lecture Notes in Mathematics, 1325:16–23, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077913/.

vandenBerg:1988:LDB

[vdBLP88] M. van den Berg, J. T. Lewis, and J. V. Pulé. Large deviations and the boson gas. Lecture Notes in Mathematics, 1325:24–39, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077914/.

vandenEssen:1987:MRS

[vdE87] Arno van den Essen. Modules with regular singularities over filtered rings and algebraic micro-localization. Lecture Notes in Mathematics, 1296:125–157, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078524/.

vanderGeer:1985:SP

[vdG85] Gerard van der Geer. The Schottky problem. Lecture Notes in Mathematics, 1111:385–406, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084599/.

vanderGeer:1989:RDA

[vdG89] Gerard van der Geer. On rank-1 degenerations of Abelian varieties. Lecture Notes in Mathematics, 1389:261–269, 1989. CODEN LNMAA2. ISBN 3-540-51509-7 (print), 3-540-48188-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085938/.

vanderKallen:1985:MST

[vdK85] Wilberd van der Kallen. The Merkurjev–Suslin theorem. Lecture Notes in Mathematics, 1142:157–168, 1985. CODEN LN-MAA2. ISBN 3-540-15674-7 (print), 3-540-39601-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074799/.

vanderMeer:1985:FPN

[vdM85a] Jan-Cees van der Meer. Fibration preserving normal forms for energy-momentum maps. Lecture Notes in Mathematics, 1160: 46–65, 1985. CODEN LNMAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080361/.

vanderMeer:1985:HHBa

[vdM85b] Jan-Cees van der Meer. The Hamiltonian Hopf bifurcation. Lecture Notes in Mathematics, 1160:66–83, 1985. CODEN LN-MAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080362/.

vanderMeer:1985:HHBb

[vdM85c] Jan-Cees van der Meer. The Hamiltonian Hopf Bifurcation, volume 1160 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 115 pp. LCCN QA3 .L28 no. 1160. URL http:

//link.springer.com/book/10.1007/BFb0080357;   http://
www.springerlink.com/content/978-3-540-39710-6.

vanderMeer:1985:I

[vdM85d] Jan-Cees van der Meer. Introduction. Lecture Notes in Mathematics, 1160:1–3, 1985. CODEN LNMAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080358/.

vanderMeer:1985:NSR

[vdM85e] Jan-Cees van der Meer. Nonintegrable systems at resonance. Lecture Notes in Mathematics, 1160:84–94, 1985. CODEN LN-MAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080363/.

vanderMeer:1985:NFH

[vdM85f] Jan-Cees van der Meer. Normal forms for Hamiltonian functions. Lecture Notes in Mathematics, 1160:19–45, 1985. CODEN LNMAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080360/.

vanderMeer:1985:P

[vdM85g] Jan-Cees van der Meer. Preliminaries. Lecture Notes in Mathematics, 1160:4–18, 1985. CODEN LNMAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080359/.

vanderMeer:1985:RPT

[vdM85h] Jan-Cees van der Meer. The restricted problem of three bodies. Lecture Notes in Mathematics, 1160:95–103, 1985. CODEN LNMAA2. ISBN 3-540-16037-X (print), 3-540-39710-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080364/.

vanderWees:1986:FME

[vdW86] A. J. van der Wees. Fast multigrid employing ILU/SIP smoothing: a robust fast solver for 3D transonic potential flow. Lecture Notes in Mathematics, 1228:315–331, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072655/.

vanEijndhoven:1985:ACL

[vEdG85a] S. J. L. van Eijndhoven and J. de Graaf. Algebras of continuous linear mappings on analyticity spaces and trajectory spaces. Lecture Notes in Mathematics, 1162:133–208, 1985. CODEN LN-MAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081154/.

vanEijndhoven:1985:AST

[vEdG85b] S. J. L. van Eijndhoven and J. de Graaf. Analyticity spaces, trajectory spaces and linear mappings between them. Lecture Notes in Mathematics, 1162:10–44, 1985. CODEN LNMAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081151/.

vanEijndhoven:1985:CST

[vEdG85c] S. J. L. van Eijndhoven and J. de Graaf. Compound spaces, tensor products and kernel theorems. Lecture Notes in Mathematics, 1162:77–132, 1985. CODEN LNMAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081153/.

vanEijndhoven:1985:IEA

[vEdG85d] S. J. L. van Eijndhoven and J. de Graaf. Illustrative examples of analyticity spaces. Lecture Notes in Mathematics, 1162:45– 76, 1985. CODEN LNMAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081152/.

vanEijndhoven:1985:MID

[vEdG85e] S. J. L. van Eijndhoven and J. de Graaf. A mathematical interpretation of Dirac’s formalism for quantum mechanics. Lecture Notes in Mathematics, 1162:209–251, 1985. CODEN LNMAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081155/.

vanEijndhoven:1985:P

[vEdG85f] S. J. L. van Eijndhoven and J. de Graaf. Prologue. Lecture Notes in Mathematics, 1162:1–9, 1985. CODEN LNMAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081150/.

vanEijndhoven:1985:TSG

[vEdG85g] S. J. L. van Eijndhoven and J. de Graaf. Trajectory Spaces, Generalized Functions and Unbounded Operators, volume 1162 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 251 (est.) pp. LCCN QA3

.L28 no. 1162. URL http://link.springer.com/book/10.1007/
BFb0081149;   http://www.springerlink.com/content/978-
3-540-39747-2.

Vegter:1985:GSG

[Veg85] Gert Vegter. Global stability of generic two-parameter families of gradients on three-manifolds. Lecture Notes in Mathematics, 1125:107–129, 1985. CODEN LNMAA2. ISBN 3-540-15233-4 (print), 3-540-39411-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075638/.

vanEngelen:1989:TRA

[vEKM89] Fons van Engelen, Kenneth Kunen, and Arnold W. Miller. Two remarks about analytic sets. Lecture Notes in Mathematics, 1401: 68–72, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097332/.

Vesentini:1985:GSL

[Ves85] Edoardo Vesentini, editor. Geometry Seminar “Luigi Bianchi” II- 1984: Lectures given at the Scuola Normale Superiore, volume 1164 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-16048-5 (print), 3-540-39732-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1164. URL http:

//link.springer.com/book/10.1007/BFb0081909;   http://
www.springerlink.com/content/978-3-540-39732-8.

Vesentini:1987:HFH

[Ves87] Edoardo Vesentini. Holomorphic families of holomorphic isometries. Lecture Notes in Mathematics, 1277:290–302, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078252/.

Vershik:1988:GNS

[VG88] A. M. Vershik and V. Ya. Gershkovich. The geometry of the nonholonomic sphere for three-dimensional Lie group. Lecture Notes in Mathematics, 1334:309–331, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080435/.

vanGeemen:1989:NQ

[vGW89] Bert van Geemen and Jürgen Werner. Nodal quintics in P^{4}. Lecture Notes in Mathematics, 1399:48–59, 1989. CODEN LN-MAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0095968/.

Vial:1989:UAP

[Via89] Jean-Philippe Vial. A unified approach to projective algorithms for linear programming. Lecture Notes in Mathematics, 1405:191– 220, 1989. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

Viennot:1985:CTG

[Vie85] Gérard Viennot. A combinatorial theory for general orthogonal polynomials with extensions and applications. Lecture Notes in Mathematics, 1171:139–157, 1985. CODEN LNMAA2. ISBN 3-540-16059-0 (print), 3-540-39743-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076539/.

Vigue:1986:QAG

[Vig86] Jean-Pierre Vigué. Quelques applications des geodesiques complexes aux points fixes d’applications holomorphes. (French) []. Lecture Notes in Mathematics, 1198:244–260, 1986. CODEN LN-MAA2. ISBN 3-540-16762-5 (print), 3-540-38729-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077058/.

Vigneras:1989:CMG

[Vig89] Marie-France Vignéras. Correspondance modulaire Galoisquaternions pour un corps p-adique. (French) []. Lecture Notes in Mathematics, 1380:254–266, 1989. CODEN LNMAA2. ISBN 3-540-51397-3 (print), 3-540-46205-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086557/.

Vinter:1985:DPO

[Vin85] R. B. Vinter. Dynamic programming for optimal control problems with terminal constraints. Lecture Notes in Mathematics, 1119:190–202, 1985. CODEN LNMAA2. ISBN 3-540-15217-2 (print), 3-540-39365-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074786/.

Viro:1988:SIC

[Vir88] O. Ya. Viro. Some integral calculus based on Euler characteristic. Lecture Notes in Mathematics, 1346:127–138, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082775/.

Verma:1985:BAT

[VJ85] A. Verma and V. K. Jain. Basic analogues of transformations of nearly-poised basic hypergeometric series. Lecture Notes in Mathematics, 1122:206–217, 1985. CODEN LNMAA2. ISBN 3-540-15222-9 (print), 3-540-39642-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075763/.

vanKampen:1987:HDS

[vK87] N. G. van Kampen. How do stochastic processes enter into physics? Lecture Notes in Mathematics, 1250:128–137, 1987. CODEN LNMAA2. ISBN 3-540-17797-3 (print), 3-540-47835-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077353/.

vanOystaeyen:1986:RTP

[vO86] Freddy M. J. van Oystaeyen, editor. Ring Theory: Proceedings of an International Conference held in Antwerp, April 1–5, 1985, volume 1197 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16496-0 (print), 3-540-39833-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1197. URL http:

//link.springer.com/book/10.1007/BFb0076306;   http://
www.springerlink.com/content/978-3-540-39833-2.

Vogell:1985:IAT

[Vog85] Wolrad Vogell. The involution in the algebraic K-theory of spaces. Lecture Notes in Mathematics, 1126:277–317, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074448/.

Vogl:1987:CAK

[Vog87a] Fritz Vogl. Über Das Cauchysche Anfangswertproblem für Eine Klasse Linearer Partieller Funktional-Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 1262:150–158, 1987. CODEN LNMAA2. ISBN 3-540-18015-X (print), 3-540-47256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078605/.

Vogt:1987:REW

[Vog87b] Andrew Vogt. The Riccati equation: When nonlinearity reduces to linearity. Lecture Notes in Mathematics, 1248:169–185, 1987. CODEN LNMAA2. ISBN 3-540-17741-8 (print), 3-540-47791-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077425/.

Vogel:1988:MAL

[Vog88] Pierre Vogel. 2 × 2- matrices and application to link theory. Lecture Notes in Mathematics, 1361:269–298, 1988. CODEN LN-MAA2. ISBN 3-540-50528-8 (print), 3-540-46036-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083040/.

Vogt:1989:RER

[Vog89] Andrew Vogt. The Riccati equation revisited. Lecture Notes in Mathematics, 1394:218–233, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086763/.

Vojta:1987:AH

[Voj87a] Paul Vojta. Approximation to hyperplanes. Lecture Notes in Mathematics, 1239:89–123, 1987. CODEN LNMAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072995/.

Vojta:1987:CMC

[Voj87b] Paul Vojta. Consequences of the main conjecture. Lecture Notes in Mathematics, 1239:46–56, 1987. CODEN LNMAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072993/.

Vojta:1987:CNT

[Voj87c] Paul Vojta. A correspondence with Nevanlinna theory. Lecture Notes in Mathematics, 1239:30–45, 1987. CODEN LN-MAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072992/.

Vojta:1987:DA

[Voj87d] Paul Vojta. Diophantine approximations. Lecture Notes in Mathematics, 1239:16–29, 1987. CODEN LNMAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072991/.

Vojta:1987:DAV

[Voj87e] Paul Vojta. Diophantine Approximations and Value Distribution Theory, volume 1239 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 132 pp. LCCN QA3 .L28 no. 1239. URL http:

//link.springer.com/book/10.1007/BFb0072989;   http://
www.springerlink.com/content/978-3-540-47452-4.

Vojta:1987:HIP

[Voj87f] Paul Vojta. Heights and integral points. Lecture Notes in Mathematics, 1239:1–15, 1987. CODEN LNMAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072990/.

Vojta:1987:RT

[Voj87g] Paul Vojta. The ramification term. Lecture Notes in Mathematics, 1239:57–88, 1987. CODEN LNMAA2. ISBN 3-540-17551-2 (print), 3-540-47452-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072994/.

Volkmer:1988:I

[Vol88a] Hans Volkmer. Introduction. Lecture Notes in Mathematics, 1356: 1–7, 1988. CODEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089296/.

Volkmer:1988:MEPb

[Vol88b] Hans Volkmer. Multiparameter eigenvalue problems for bounded operators. Lecture Notes in Mathematics, 1356:21–49, 1988. CO-DEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0089298/.

Volkmer:1988:MEPa

[Vol88c] Hans Volkmer. Multiparameter eigenvalue problems for Hermitian matrices. Lecture Notes in Mathematics, 1356:8–20, 1988. CODEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089297/.

Volkmer:1988:MEPc

[Vol88d] Hans Volkmer. Multiparameter eigenvalue problems for unbounded operators. Lecture Notes in Mathematics, 1356:50– 75, 1988. CODEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089299/.

Volkmer:1988:METb

[Vol88e] Hans Volkmer. Multiparameter expansion theorems for bounded operators. Lecture Notes in Mathematics, 1356:87–115, 1988. CO-DEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0089301/.

Volkmer:1988:METa

[Vol88f] Hans Volkmer. Multiparameter expansion theorems for Hermitian matrices. Lecture Notes in Mathematics, 1356:76–86, 1988. CODEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089300/.

Volkmer:1988:METc

[Vol88g] Hans Volkmer. Multiparameter expansion theorems for unbounded operators. Lecture Notes in Mathematics, 1356:116– 141, 1988. CODEN LNMAA2. ISBN 3-540-50479-6 (print), 3-540-46015-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089302/.

Vossieck:1986:IPC

[Vos86a] Dieter Vossieck. On indecomposables in preprojective components. Lecture Notes in Mathematics, 1177:340, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075273.

Vossieck:1986:RTF

[Vos86b] Dieter Vossieck. Representation theory I — finite dimensional algebras. Lecture Notes in Mathematics, 1177:341, 1986. CODEN LNMAA2. ISBN 3-540-16432-4 (print), 3-540-39776-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075274.

Vigue-Poirrier:1988:CHQ

[VP88] Micheline Vigué-Poirrier. Cyclic homology and Quillen homology of a commutative algebra. Lecture Notes in Mathematics, 1318:238–245, 1988. CODEN LNMAA2. ISBN 3-540-19340-5 (print), 3-540-39204-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077807/.

VanGeel:1987:MOC

[VS87a] Jan Van Geel and P. Salberger. Maximal orders over curves. Lecture Notes in Mathematics, 1296:193–213, 1987. CODEN LN-MAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078527/.

vanStraten:1987:BNM

[vS87b] D. van Straten. On the Betti numbers of the Milnor fibre of a certain class of hypersurface singularities. Lecture Notes in Mathematics, 1273:203–220, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078845/.

Vsekhsvyatskii:1985:SNR

[Vse85] S. Yu. Vsekhsvyat·skiĭ. Some new results in a theorem due to Renyi. Lecture Notes in Mathematics, 1155:391–400, 1985. CO-DEN LNMAA2. ISBN 3-540-15985-1 (print), 3-540-39686-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074829/.

VanderHouwen:1989:CSC

[VSP89] P. J. Van der Houwen, B. P. Sommeijer, and G. Pontrelli. A comparative study of Chebyshev acceleration and residue smoothing in the solution of nonlinear elliptic difference equations. Lecture Notes in Mathematics, 1386:69–96, 1989. CODEN LN-MAA2. ISBN 3-540-51478-3 (print), 3-540-48144-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089232/.

Vuorinen:1988:BB

[Vuo88a] Matti Vuorinen. Boundary behavior. Lecture Notes in Mathematics, 1319:173–192, 1988. CODEN LNMAA2. ISBN 3-540-19342-1 (print), 3-540-39207-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077908/.

Vuorinen:1988:CG

[Vuo88b] Matti Vuorinen. Conformal geometry. Lecture Notes in Mathematics, 1319:1–47, 1988. CODEN LNMAA2. ISBN 3-540-19342-1 (print), 3-540-39207-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077905/.

Vuorinen:1988:MC

[Vuo88c] Matti Vuorinen. Modulus and capacity. Lecture Notes in Mathematics, 1319:48–119, 1988. CODEN LNMAA2. ISBN 3-540-19342-1 (print), 3-540-39207-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077906/.

Vuorinen:1988:QM

[Vuo88d] Matti Vuorinen. Quasiregular mappings. Lecture Notes in Mathematics, 1319:120–172, 1988. CODEN LNMAA2. ISBN 3-540-19342-1 (print), 3-540-39207-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077907/.

vonWaldenfels:1985:SSQ

[vW85] W. von Waldenfels. Stratonovich solution of a quantum stochastic differential equation describing light emission and absorption. Lecture Notes in Mathematics, 1109:155–176, 1985. CODEN LN-MAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101543/.

vonWaldenfels:1986:NCA

[vW86] Wilhelm von Waldenfels. Non-commutative algebraic central limit theorems. Lecture Notes in Mathematics, 1210:174–202, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077184/.

vanWeert:1989:SPR

[vW89] Ch. G. van Weert. Some problems in relativistic hydrodynamics. Lecture Notes in Mathematics, 1385:290–300, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084036/.

Vosmansky:1986:EPI

[VZ86] Jaromír Vosmanský and Miloš Zlámal, editors. Equadiff 6: Proceedings of the International Conference on Differential Equations and their Applications held in Brno, Czechoslovakia, Aug. 26–30, 1985, volume 1192 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (paperback), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1192; QA299.6-433. URL http://

link.springer.com/book/10.1007/BFb0076046.

Waadeland:1986:NPD

[Waa86] Haakon Waadeland. A note on partial derivatives of continued fractions. Lecture Notes in Mathematics, 1199:294–299, 1986. CODEN LNMAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075944/.

Waadeland:1987:LPC

[Waa87] Haakon Waadeland. Local properties of continued fractions. Lecture Notes in Mathematics, 1237:239–250, 1987. CODEN LN-MAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072468/.

Waadeland:1989:BVW

[Waa89] Haakon Waadeland. Boundary versions of Worpitzky’s Theorem and of parabola theorems. Lecture Notes in Mathematics, 1406:135–142, 1989. CODEN LNMAA2. ISBN 3-540-51830-4 (print), 3-540-46820-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096171/.

Wagner:1987:TEL

[Wag87] David H. Wagner. The transformation from Eulerian to Lagrangian coordinates for solutions with discontinuities. Lecture Notes in Mathematics, 1270:172–180, 1987. CODEN LN-MAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078327/.

Wainger:1989:PHA

[Wai89] Stephen Wainger. Problems in harmonic analysis related to curves and surfaces with infinitely flat points. Lecture Notes in Mathematics, 1384:91–110, 1989. CODEN LNMAA2. ISBN 3-540-51460-0 (print), 3-540-48134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086795/.

Waksman:1987:HAM

[Wak87] Leonid Waksman. Harmonic analysis of multi-parameter semigroups of contractions. Lecture Notes in Mathematics, 1272:39– 114, 1987. CODEN LNMAA2. ISBN 3-540-18316-7 (print), 3-540-47877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078927/.

Waldhausen:1985:ATS

[Wal85a] Friedhelm Waldhausen. Algebraic K-theory of spaces. Lecture Notes in Mathematics, 1126:318–419, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074449/.

Wall:1985:GGS

[Wal85b] C. T. C. Wall. Geometries and geometric structures in real dimension 4 and complex dimension 2. Lecture Notes in Mathematics, 1167:268–292, 1985. CODEN LNMAA2. ISBN 3-540-16053-1 (print), 3-540-39738-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075230/.

Walsh:1986:ISP

[Wal86a] John B. Walsh. An introduction to stochastic partial differential equations. Lecture Notes in Mathematics, 1180:265–439, 1986. CODEN LNMAA2. ISBN 3-540-16441-3 (print), 3-540-39781-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074920/.

Walsh:1986:MMP

[Wal86b] John B. Walsh. Martingales with a multidimensional parameter and stochastic integrals in the plane. Lecture Notes in Mathematics, 1215:329–491, 1986. CODEN LNMAA2. ISBN 3-540-16822-2 (print), 3-540-47092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075875/.

Walters:1986:UER

[Wal86c] Peter Walters. Unique ergodicity and random matrix products. Lecture Notes in Mathematics, 1186:37–55, 1986. CODEN LN-MAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076832/.

Waldvoge:1987:ZFD

[Wal87] J. Waldvogel. Zero-free disks in families of analytic functions. Lecture Notes in Mathematics, 1287:209–228, 1987. CODEN LN-MAA2. ISBN 3-540-18500-3 (print), 3-540-47991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078907/.

Wallin:1988:NOF

[Wal88] Hans Wallin. New and old function spaces. Lecture Notes in Mathematics, 1302:99–114, 1988. CODEN LNMAA2. ISBN 3-540-18905-X (print), 3-540-38841-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078866/.

Wang:1986:EQE

[Wan86] E Fang Wang. The equitype and quasi-equitype decompositions of arbitrary permutations by permutations of order 2 or 3. Lecture Notes in Mathematics, 1185:340–360, 1986. CODEN LN-MAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076183/.

Wang:1987:GMS

[Wan87a] Chang-Ping Wang. On the Gauss map of submanifold in R_{n} and S_{n}. Lecture Notes in Mathematics, 1255:109–129, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077686/.

Wang:1987:NSE

[Wan87b] Jinghua Wang. A nonhomogeneous system of equations of nonisentropic gas dynamics. Lecture Notes in Mathematics, 1270: 221–237, 1987. CODEN LNMAA2. ISBN 3-540-18200-4 (print), 3-540-47805-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078331/.

Wang:1987:AFP

[Wan87c] Xiaolu Wang. The C∗-algebras of foliations of the plane. Lecture Notes in Mathematics, 1257:111–158, 1987. CODEN LN-MAA2. ISBN 3-540-17903-8 (print), 3-540-47915-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077990/.

Wang:1987:DT

[Wan87d] Xiaolu Wang. Distinguished trees. Lecture Notes in Mathematics, 1257:53–110, 1987. CODEN LNMAA2. ISBN 3-540-17903-8 (print), 3-540-47915-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077989/.

Wang:1987:FP

[Wan87e] Xiaolu Wang. Foliations of the plane. Lecture Notes in Mathematics, 1257:10–34, 1987. CODEN LNMAA2. ISBN 3-540-17903-8 (print), 3-540-47915-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077987/.

Wang:1987:I

[Wan87f] Xiaolu Wang. Introduction. Lecture Notes in Mathematics, 1257: 1–9, 1987. CODEN LNMAA2. ISBN 3-540-17903-8 (print), 3-540-47915-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077986/.

Wang:1987:CAF

[Wan87g] Xiaolu Wang. On the C∗-Algebras of Foliations in the Plane, volume 1257 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17903-8 (print), 3-540-47915-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 165 pp. LCCN QA3 .L28 no. 1257. URL http:

//link.springer.com/book/10.1007/BFb0077985;   http://
www.springerlink.com/content/978-3-540-47915-4.

Wang:1987:VTG

[Wan87h] Xiaolu Wang. Various trees and graphs. Lecture Notes in Mathematics, 1257:35–52, 1987. CODEN LNMAA2. ISBN 3-540-17903-8 (print), 3-540-47915-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077988/.

Wang:1988:HIF

[Wan88] Guanglie Wang. Harnack inequalities for functions in De Giorgi parabolic class. Lecture Notes in Mathematics, 1306:182–201, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082934/.

Wang:1989:EIC

[Wan89] Shicheng Wang. Essential invariant circles of surface automorphisms of finite order. Lecture Notes in Mathematics, 1369:275– 285, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087540/.

Warfield:1986:NLR

[War86] R. B. Warfield, Jr. Noncommutative localized rings. Lecture Notes in Mathematics, 1220:178–200, 1986. CODEN LNMAA2. ISBN 3-540-17185-1 (print), 3-540-47255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099512/.

Watanabe:1987:CAC

[Wat87] Tadashi Watanabe. The continuity axiom and the Čech homology. Lecture Notes in Mathematics, 1283:221–239, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081431/.

Watling:1988:FHK

[Wat88] Keith D. Watling. Formulæ for the heat kernel of an elliptic operator exhibiting small-time asymptotics. Lecture Notes in Mathematics, 1325:167–180, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077926/.

Weber:1988:TIP

[Web88] Claude Weber. A topological interpretation for the polar quotients of an algebraic plane curve singularity. Lecture Notes in Mathematics, 1350:122–143, 1988. CODEN LNMAA2. ISBN 3-540-50369-2 (print), 3-540-45990-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081472/.

Weinberger:1985:OFM

[Wei85a] Shmuel Weinberger. Oliver’s formula and Minkowski’s theorem. Lecture Notes in Mathematics, 1126:420–421, 1985. CODEN LN-MAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074450.

Weinberger:1985:SNC

[Wei85b] Shmuel Weinberger. Some nilpotent complexes. Lecture Notes in Mathematics, 1126:422–423, 1985. CODEN LNMAA2. ISBN 3-540-15235-0 (print), 3-540-39413-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074451.

Weissauer:1986:AKE

[Wei86a] Rainer Weissauer. Ableitungen der Klingenschen Eisensteinreihen. (German) []. Lecture Notes in Mathematics, 1219:80–83, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072243/.

Weissauer:1986:AFG

[Wei86b] Rainer Weissauer. Automorphe Formen. (German) [Automorphic forms]. Lecture Notes in Mathematics, 1219:45–50, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072239/.

Weissauer:1986:HDS

[Wei86c] Rainer Weissauer. Das holomorphe diskrete Spektrum von L^{2}(Γ_{n}/G). (German) []. Lecture Notes in Mathematics, 1219:108– 111, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072246/.

Weissauer:1986:GG

[Wei86d] Rainer Weissauer. Der Grenzefall k = n + j + 1/2. (German) []. Lecture Notes in Mathematics, 1219:97–107, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072245/.

Weissauer:1986:OG

[Wei86e] Rainer Weissauer. Der Operator M(ρ, s). (German) []. Lecture Notes in Mathematics, 1219:112–122, 1986. CODEN LN-MAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072247/.

Weissauer:1986:SEG

[Wei86f] Rainer Weissauer. Die Siegelschen Eisensteinreihen. (German) []. Lecture Notes in Mathematics, 1219:131–141, 1986. CODEN LN-MAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072249/.

Weissauer:1986:DGD

[Wei86g] Rainer Weissauer. Differentialoperatoren. (German) [Differential operators]. Lecture Notes in Mathematics, 1219:23–44, 1986. CO-DEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0072238/.

Weissauer:1986:EGE

[Wei86h] Rainer Weissauer. Eisensteinreihen. (German) [Eisenstein series]. Lecture Notes in Mathematics, 1219:59–67, 1986. CODEN LN-MAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072241/.

Weissauer:1986:EKT

[Wei86i] Rainer Weissauer. Eisensteinreihen vom Klingenschen Typ. (German) []. Lecture Notes in Mathematics, 1219:68–79, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072242/.

Weissauer:1986:HGH

[Wei86j] Rainer Weissauer. Hyperebenen. (German) [Hyperplanes]. Lecture Notes in Mathematics, 1219:51–58, 1986. CODEN LN-MAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072240/.

Weissauer:1986:I

[Wei86k] Rainer Weissauer. Introduction. Lecture Notes in Mathematics, 1219:1–6, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072236/.

Weissauer:1986:PEG

[Wei86l] Rainer Weissauer. Polstellen der Eisensteinreihen. (German) []. Lecture Notes in Mathematics, 1219:84–96, 1986. CODEN LN-MAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072244/.

Weissauer:1986:SLG

[Wei86m] Rainer Weissauer. Stabile Liftungen. (German) []. Lecture Notes in Mathematics, 1219:123–130, 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072248/.

Weissauer:1986:SMG

[Wei86n] Rainer Weissauer. Stabile Modulformen. (German) [Robust modular forms]. Lecture Notes in Mathematics, 1219:7–22, 1986. CO-DEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0072237/.

Weissauer:1986:SME

[Wei86o] Rainer Weissauer. Stabile Modulformen und Eisensteinreihen, volume 1219 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNMAA2. ISBN 3-540-17181-9 (print), 3-540-47248-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 147 pp. LCCN QA3 .L28 no. 1219. URL http:

//link.springer.com/book/10.1007/BFb0072235;   http://
www.springerlink.com/content/978-3-540-47248-3.

Weidmann:1987:SES

[Wei87a] Joachim Weidmann. L_{2}-solutions and essential spectrum. Lecture Notes in Mathematics, 1258:162–171, 1987. CODEN LN-MAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077975/.

Weidmann:1987:ASSa

[Wei87b] Joachim Weidmann. Appendix to section 1: The separation of the Dirac operator. Lecture Notes in Mathematics, 1258:16–22, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077963/.

Weidmann:1987:ASO

[Wei87c] Joachim Weidmann. Appendix to section 12: Operators with periodic coefficients on the half-line. Lecture Notes in Mathematics, 1258:191–193, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077977/.

Weidmann:1987:ASP

[Wei87d] Joachim Weidmann. Appendix to section 2: Proof of the Lagrange identity for n > 2. Lecture Notes in Mathematics, 1258: 35–40, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077965/.

Weidmann:1987:ASSb

[Wei87e] Joachim Weidmann. Appendix to section 6: Semi-boundedness of Sturm–Liouville type operators. Lecture Notes in Mathematics, 1258:104–109, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077970/.

Weidmann:1987:CSM

[Wei87f] Joachim Weidmann. Computation of the spectral matrix ϱ. Lecture Notes in Mathematics, 1258:140–149, 1987. CODEN LN-MAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077973/.

Weidmann:1987:DIS

[Wei87g] Joachim Weidmann. Deficiency indices and self-adjoint extensions of T_{0}. Lecture Notes in Mathematics, 1258:52–71, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077967/.

Weidmann:1987:DOP

[Wei87h] Joachim Weidmann. Differential operators with periodic coefficients. Lecture Notes in Mathematics, 1258:172–190, 1987. CO-DEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077976/.

Weidmann:1987:ESA

[Wei87i] Joachim Weidmann. Essential spectrum and absolutely continuous spectrum of Sturm–Liouville operators. Lecture Notes in Mathematics, 1258:227–241, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077980/.

Weidmann:1987:FSA

[Wei87j] Joachim Weidmann. Formally self-adjoint differential expressions. Lecture Notes in Mathematics, 1258:7–15, 1987. CODEN LN-MAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077962/.

Weidmann:1987:FPG

[Wei87k] Joachim Weidmann. Fundamental properties and general assumptions. Lecture Notes in Mathematics, 1258:23–35, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077964/.

Weidmann:1987:I

[Wei87l] Joachim Weidmann. Introduction. Lecture Notes in Mathematics, 1258:1–6, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077961/.

Weidmann:1987:LPL

[Wei87m] Joachim Weidmann. Limit point-limit circle criteria. Lecture Notes in Mathematics, 1258:88–103, 1987. CODEN LN-MAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077969/.

Weidmann:1987:MOM

[Wei87n] Joachim Weidmann. The minimal operator and the maximal operator. Lecture Notes in Mathematics, 1258:41–51, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077966/.

Weidmann:1987:OTR

[Wei87o] Joachim Weidmann. Oscillation theory for regular Sturm– Liouville operators. Lecture Notes in Mathematics, 1258:194– 212, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077978/.

Weidmann:1987:OTS

[Wei87p] Joachim Weidmann. Oscillation theory for singular Sturm– Liouville operators. Lecture Notes in Mathematics, 1258:213– 226, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077979/.

Weidmann:1987:RSA

[Wei87q] Joachim Weidmann. The resolvents of self-adjoint extensions of T_{0}. Lecture Notes in Mathematics, 1258:110–125, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077971/.

Weidmann:1987:SID

[Wei87r] Joachim Weidmann. The solutions of the inhomogeneous differential equation (τ − λ)u = f; Weyl’s alternative. Lecture Notes in Mathematics, 1258:72–87, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077968/.

Weidmann:1987:SPS

[Wei87s] Joachim Weidmann. Special properties of the spectral representation, spectral multiplicities. Lecture Notes in Mathematics, 1258: 150–161, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077974/.

Weidmann:1987:SRS

[Wei87t] Joachim Weidmann. The spectral representation of self-adjoint extensions of T_{0}. Lecture Notes in Mathematics, 1258:126–139, 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077972/.

Weidmann:1987:STO

[Wei87u] Joachim Weidmann. Spectral Theory of Ordinary Differential Operators, volume 1258 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNMAA2. ISBN 3-540-17902-X (print), 3-540-47912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 303 pp. LCCN QA3 .L28 no. 1258. URL http:

//link.springer.com/book/10.1007/BFb0077960;   http://
www.springerlink.com/content/978-3-540-47912-3.

Weissauer:1987:DSM

[Wei87v] R. Weissauer. Divisors of the Siegel modular variety. Lecture Notes in Mathematics, 1240:304–324, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072987/.

Weissauer:1989:CSM

[Wei89a] Rainer Weissauer. On the cohomology of Siegel modular threefolds. Lecture Notes in Mathematics, 1399:155–170, 1989. CODEN LNMAA2. ISBN 3-540-51729-4 (print), 3-540-46791-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095975/.

Weitzner:1989:RP

[Wei89b] Harold Weitzner. Relativistic plasmas. Lecture Notes in Mathematics, 1385:211–237, 1989. CODEN LNMAA2. ISBN 3-540-51466-X (print), 3-540-48142-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084031/.

Wente:1985:CSC

[Wen85] Henry C. Wente. A counterexample in 3-space to a conjecture of H. Hopf. Lecture Notes in Mathematics, 1111:421–429, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084601/.

Wermer:1987:BPH

[Wer87a] John Wermer. Balayage and polynomial hulls. Lecture Notes in Mathematics, 1276:303–312, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078966/.

Werner:1987:CPM

[Wer87b] Helmut Werner. On the continuity properties of the multivariate Padé-operator T_{m,n}. Lecture Notes in Mathematics, 1237:125– 133, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072459/.

Whiteman:1986:STT

[Whi86] J. R. Whiteman. Singularities in two- and three-dimensional elliptic problems and finite element methods for their treatment. Lecture Notes in Mathematics, 1192:345–352, 1986. CODEN LN-MAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076092/.

Whitehead:1989:WEH

[Whi89] George W. Whitehead. The work of Edgar H. Brown, Jr. in topology. Lecture Notes in Mathematics, 1370:10–14, 1989. CODEN LNMAA2. ISBN 3-540-51118-0 (print), 3-540-46160-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085214/.

Widom:1985:HC

[Wid85a] Harold Widom. f^{(k)}(σ; τ_{i}) in the Hermitian case. Lecture Notes in Mathematics, 1152:42–44, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075037/.

Widom:1985:AEP

[Wid85b] Harold Widom. Asymptotic Expansions for Pseudodifferential Operators on Bounded Domains, volume 1152 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 150 pp. LCCN QA3

.L28 no. 1152. URL http://link.springer.com/book/10.1007/
BFb0075033;   http://www.springerlink.com/content/978-
3-540-39641-3.

Widom:1985:DS

[Wid85c] Harold Widom. Derivation of the series. Lecture Notes in Mathematics, 1152:16–35, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075035/.

Widom:1985:I

[Wid85d] Harold Widom. Introduction. Lecture Notes in Mathematics, 1152: 1–15, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075034/.

Widom:1985:PHE

[Wid85e] Harold Widom. Proof of the heat expansion. Lecture Notes in Mathematics, 1152:102–147, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075041/.

Widom:1985:PSEa

[Wid85f] Harold Widom. Proof of the Szegő expansion in the nonself-adjoint case. Lecture Notes in Mathematics, 1152:65–94, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075039/.

Widom:1985:PSEb

[Wid85g] Harold Widom. Proof of the Szegő expansion in the self-adjoint case. Lecture Notes in Mathematics, 1152:95–101, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075040/.

Widom:1985:SHE

[Wid85h] Harold Widom. The Szegő and heat expansions. Lecture Notes in Mathematics, 1152:36–41, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075036/.

Widom:1985:TF

[Wid85i] Harold Widom. Trace formulas. Lecture Notes in Mathematics, 1152:45–64, 1985. CODEN LNMAA2. ISBN 3-540-15701-8 (print), 3-540-39641-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075038/.

Wiedemann:1986:GOC

[Wie86] Alfred Wiedemann. Gorenstein orders of class A_{n}. Lecture Notes in Mathematics, 1178:309–315, 1986. CODEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075300/.

Wiegerinck:1987:CFP

[Wie87] J. Wiegerinck. Convergence of formal power series and analytic extension. Lecture Notes in Mathematics, 1276:313–320, 1987. CODEN LNMAA2. ISBN 3-540-18357-4 (print), 3-540-47904-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078967/.

Wihstutz:1985:AEL

[Wih85] Volker Wihstutz. Analytic expansion of Lyapunov exponents associated to the Schrödinger operator. Lecture Notes in Mathematics, 1109:177–188, 1985. CODEN LNMAA2. ISBN 3-540-13914-1 (print), 3-540-39138-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101544/.

Wihstutz:1986:PDL

[Wih86] Volker Wihstutz. Parameter dependence of the Lyapunov exponent for linear stochastic systems. A survey. Lecture Notes in Mathematics, 1186:200–215, 1986. CODEN LNMAA2. ISBN 3-540-16458-8 (print), 3-540-39795-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076840/.

Willems:1986:GCP

[Wil86a] Wolfgang Willems. The generalized p′-core and the principal pblock. Lecture Notes in Mathematics, 1178:177–193, 1986. CO-DEN LNMAA2. ISBN 3-540-16433-2 (print), 3-540-39777-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075296/.

Williams:1986:SCP

[Wil86b] William O. Williams. Structure of continuum physics. Lecture Notes in Mathematics, 1174:30–37, 1986. CODEN LN-MAA2. ISBN 3-540-16096-5 (print), 3-540-39760-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076931/.

Wilson:1987:ECS

[Wil87] John S. Wilson. An embedding condition for subgroups of infinite groups. Lecture Notes in Mathematics, 1281:176–179, 1987. CODEN LNMAA2. ISBN 3-540-18399-X (print), 3-540-47948-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078702/.

Wildenhain:1988:PTM

[Wil88] G. Wildenhain. Potential theory methods for higher order elliptic equations. Lecture Notes in Mathematics, 1344:181–195, 1988. CODEN LNMAA2. ISBN 3-540-50210-6 (print), 3-540-45952-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103351/.

Wilson:1989:DFA

[Wil89] Robert Lee Wilson. Differential forms and the algebra W (m: n). Lecture Notes in Mathematics, 1373:29–41, 1989. CODEN LN-MAA2. ISBN 3-540-51147-4 (print), 3-540-46170-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088885/.

Winkler:1985:BCN

[Win85a] Gerhard Winkler. Basic concepts from noncompact Choquet theory. Lecture Notes in Mathematics, 1145:13–53, 1985. CODEN LNMAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075054/.

Winkler:1985:COS

[Win85b] Gerhard Winkler. Choquet Order and Simplices: with Applications in Probabilistic Models, volume 1145 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 143 pp. LCCN QA3

.L28 no. 1145. URL http://link.springer.com/book/10.1007/
BFb0075051;   http://www.springerlink.com/content/978-
3-540-39622-2.

Winkler:1985:FAC

[Win85c] Gerhard Winkler. Four aspects of Choquet order. Lecture Notes in Mathematics, 1145:54–74, 1985. CODEN LNMAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075055/.

Winkler:1985:I

[Win85d] Gerhard Winkler. Introduction. Lecture Notes in Mathematics, 1145:1–5, 1985. CODEN LNMAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075052/.

Winkler:1985:ILSa

[Win85e] Gerhard Winkler. Inverse limits of simplices. Lecture Notes in Mathematics, 1145:75–93, 1985. CODEN LNMAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075056/.

Winkler:1985:ILSb

[Win85f] Gerhard Winkler. Inverse limits of spaces of measures. Lecture Notes in Mathematics, 1145:94–122, 1985. CODEN LN-MAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075057/.

Winkler:1985:NDC

[Win85g] Gerhard Winkler. Notations, definitions and conventions. Lecture Notes in Mathematics, 1145:6–12, 1985. CODEN LNMAA2. ISBN 3-540-15683-6 (print), 3-540-39622-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075053/.

Winter:1989:TNS

[Win89] Pawel Winter. Topological network synthesis. Lecture Notes in Mathematics, 1403:282–303, 1989. CODEN LNMAA2. ISBN 3-540-51797-9 (print), 3-540-46810-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083472/.

Witsch:1988:MVF

[Wit88a] K. J. Witsch. On a modified version of the free geodetic boundaryvalue problem. Lecture Notes in Mathematics, 1357:402–423, 1988. CODEN LNMAA2. ISBN 3-540-50508-3 (print), 3-540-46024-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082877/.

Witten:1988:IDO

[Wit88b] Edward Witten. The index of the Dirac operator in loop space. Lecture Notes in Mathematics, 1326:161–181, 1988. CODEN LN-MAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078045/.

Watson:1986:LME

[WKK86] L. T. Watson, M. P. Kamat, and H. Y. Kwok. Location of multiple equilibrium configurations near limit points by a double dogleg strategy and tunnelling. Lecture Notes in Mathematics, 1230:26–45, 1986. CODEN LNMAA2. ISBN 3-540-17200-9 (print), 3-540-47379-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072669/.

Wodzicki:1987:NRC

[Wod87] Mariusz Wodzicki. Noncommutative residue. Chapter I. Fundamentals. Lecture Notes in Mathematics, 1289:320–399, 1987. CO-DEN LNMAA2. ISBN 3-540-18571-2 (print), 3-540-48016-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078372/.

Woess:1986:DMB

[Woe86] Wolfgang Woess. A description of the Martin boundary for nearest neighbour random walks on free products. Lecture Notes in Mathematics, 1210:203–215, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077185/.

Wojtkowiak:1987:M

[Woj87] Zdzisław Wojtkowiak. On maps from h\underline{o} lim F to Z. Lecture Notes in Mathematics, 1298:227–236, 1987. CODEN LNMAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083013/.

Wong:1987:CMA

[Won87] Pit-Mann Wong. Complex Monge–Ampère equation and related problems. Lecture Notes in Mathematics, 1277:303–330, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078253/.

Wood:1985:MRH

[Woo85] R. M. W. Wood. Modular representations of GL(n, F_{P}) and homotopy theory. Lecture Notes in Mathematics, 1172:188–203, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074432/.

Wood:1987:TCH

[Woo87a] J. C. Wood. Twistor constructions for harmonic maps. Lecture Notes in Mathematics, 1255:130–159, 1987. CODEN LN-MAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077687/.

Wood:1987:SAS

[Woo87b] R. M. W. Wood. Splitting Σ(CP^{∞} × … × CP^{∞}) and the action of Steenrod squares Sq^{i} on the polynomial ring F_{2}[x_{1}, …, x_{n}]. Lecture Notes in Mathematics, 1298:237–255, 1987. CODEN LN-MAA2. ISBN 3-540-18729-4 (print), 3-540-48122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083014/.

Wright:1989:CRU

[Wri89a] Steve Wright. Connes’ reduction of the uniqueness proof to the bicentralizer problem. Lecture Notes in Mathematics, 1413:9– 81, 1989. CODEN LNMAA2. ISBN 3-540-52130-5 (print), 3-540-46903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090180/.

Wright:1989:HSB

[Wri89b] Steve Wright. Haagerup’s solution of the bicentralizer problem. Lecture Notes in Mathematics, 1413:82–105, 1989. CODEN LN-MAA2. ISBN 3-540-52130-5 (print), 3-540-46903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090181/.

Wright:1989:I

[Wri89c] Steve Wright. Introduction. Lecture Notes in Mathematics, 1413: 1–8, 1989. CODEN LNMAA2. ISBN 3-540-52130-5 (print), 3-540-46903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090179/.

Wronicz:1987:MIG

[Wro87] Z. Wronicz. The Marchaud inequality for generalized moduli of smoothness. Lecture Notes in Mathematics, 1237:134–144, 1987. CODEN LNMAA2. ISBN 3-540-17212-2 (print), 3-540-47412-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072460/.

Wagstaff:1987:MFL

[WS87] Samuel S. Wagstaff, Jr. and J. W. Smith. Methods of factoring large integers. Lecture Notes in Mathematics, 1240:281–303, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072986/.

Wschebor:1985:ADT

[Wsc85a] Mario Wschebor. Approximation du temps local du processus de Wiener à d-paramètres (d ≥ 1). (French) []. Lecture Notes in Mathematics, 1147:88–106, 1985. CODEN LNMAA2. ISBN 3-540-15688-7 (print), 3-540-39633-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075077/.

Wschebor:1985:EEL

[Wsc85b] Mario Wschebor. Existence d’extrema locaux. (French) []. Lecture Notes in Mathematics, 1147:11–27, 1985. CODEN LN-MAA2. ISBN 3-540-15688-7 (print), 3-540-39633-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075075/.

Wschebor:1985:FRF

[Wsc85c] Mario Wschebor. Formules de Rice. (French) []. Lecture Notes in Mathematics, 1147:28–87, 1985. CODEN LNMAA2. ISBN 3-540-15688-7 (print), 3-540-39633-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075076/.

Wschebor:1985:PEB

[Wsc85d] Mario Wschebor. Périmètres d’ensembles boréliens dans R^{d}. (French) []. Lecture Notes in Mathematics, 1147:1–10, 1985. CO-DEN LNMAA2. ISBN 3-540-15688-7 (print), 3-540-39633-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075074/.

Wschebor:1985:SAM

[Wsc85e] Mario Wschebor. Surfaces Aléatoires: Mesure Géométrique des Ensembles de Niveau, volume 1147 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNMAA2. ISBN 3-540-15688-7 (print), 3-540-39633-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 111 pp. LCCN QA3

.L28 no. 1147. URL http://link.springer.com/book/10.1007/
BFb0075073;   http://www.springerlink.com/content/978-
3-540-39633-8.

Williams:1988:SMP

[WST88] David Williams, Nick Steele, and Aubrey Truman. Stochastic mechanics for a point source. Lecture Notes in Mathematics, 1325:181–187, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077927/.

Wen-tsun:1987:MH

[Wt87a] Wu Wen-tsün. I∗-measure and homotopy. Lecture Notes in Mathematics, 1264:93–115, 1987. CODEN LNMAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082001/.

Wen-tsun:1987:RST

[Wt87b] Wu Wen-tsün. The de Rham–Sullivan theorem and I∗-measure. Lecture Notes in Mathematics, 1264:58–92, 1987. CODEN LN-MAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082000/.

Wen-tsun:1987:DMM

[Wt87c] Wu Wen-tsün. Dga and minimal model. Lecture Notes in Mathematics, 1264:20–57, 1987. CODEN LNMAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081999/.

Wen-tsun:1987:ECA

[Wt87d] Wu Wen-tsün. Effective computation and axiomatic system of I∗-measure. Lecture Notes in Mathematics, 1264:143–177, 1987. CODEN LNMAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082003/.

Wen-tsun:1987:FCM

[Wt87e] Wu Wen-tsün. Fundamental concepts. Measure and calculability. Lecture Notes in Mathematics, 1264:1–19, 1987. CODEN LNMAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081998/.

Wen-tsun:1987:MCF

[Wt87f] Wu Wen-tsün. I*-measures connected with fibrations. Lecture Notes in Mathematics, 1264:178–213, 1987. CODEN LN-MAA2. ISBN 3-540-13611-8 (print), 3-540-39025-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082004/.

Wu:1986:EAV

[Wu86] Xiaolong Wu. On the extensions of Abelian varieties by affine group schemes. Lecture Notes in Mathematics, 1185:361–387, 1986. CODEN LNMAA2. ISBN 3-540-16456-1 (print), 3-540-39793-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076184/.

Wu:1987:TCH

[Wu87a] Chuanxi Wu. On two classes of hypersurfaces in a space of constant curvature. Lecture Notes in Mathematics, 1255:160– 172, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077688/.

Wu:1987:LT

[Wu87b] H. Wu. Liouville theorems. Lecture Notes in Mathematics, 1277: 331–349, 1987. CODEN LNMAA2. ISBN 3-540-18355-8 (print), 3-540-47893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078254/.

Wu:1987:COM

[Wu87c] Liming Wu. Construction de l’opérateur de Malliavin sur l’espace de Poisson. (French) []. Lecture Notes in Mathematics, 1247:100– 113, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077629/.

Wu:1987:ISE

[Wu87d] Liming Wu. Inégalité de Sobolev sur l’espace de Poisson. (French) []. Lecture Notes in Mathematics, 1247:114–136, 1987. CODEN LNMAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077630/.

Wu:1988:CDB

[Wu88a] Fontong Wu. A class diffractive boundary value problem with multiple characteristic. Lecture Notes in Mathematics, 1306:224– 239, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082936/.

Wu:1988:EUR

[Wu88b] Lancheng Wu. Existence, uniqueness and regularity of the minimizer of a certain functional. Lecture Notes in Mathematics, 1306:240–254, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082937/.

Wunram:1987:RMC

[Wun87] J. Wunram. Reflexive modules on cyclic quotient surface singularities. Lecture Notes in Mathematics, 1273:221–231, 1987. CODEN LNMAA2. ISBN 3-540-18263-2 (print), 3-540-47851-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078846/.

Wustholz:1987:NABa

[Wüs87a] G. Wüstholz. A new approach to Baker’s theorem on linear forms in logarithms I. Lecture Notes in Mathematics, 1290:189– 202, 1987. CODEN LNMAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078709/.

Wustholz:1987:NABb

[Wüs87b] G. Wüstholz. A new approach to Baker’s theorem on linear forms in logarithms II. Lecture Notes in Mathematics, 1290:203– 211, 1987. CODEN LNMAA2. ISBN 3-540-18597-6 (print), 3-540-48023-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078710/.

Xia:1987:RFG

[Xia87] C. Y. Xia. Remarks on the fundamental group of positively curved manifolds. Lecture Notes in Mathematics, 1255:190–197, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077690/.

Xiao:1989:CMS

[Xia89] Liang Xiao. On complete minimal surfaces with parallel and flat ends. Lecture Notes in Mathematics, 1369:289–294, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087542/.

Xin:1987:LTT

[Xin87] Y. L. Xin. Liouville type theorems and regularity of harmonic maps. Lecture Notes in Mathematics, 1255:198–208, 1987. CO-DEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077691/.

Xin:1989:RHM

[Xin89] Yuan-Long Xin. Regularity of harmonic maps into certain homogeneous spaces. Lecture Notes in Mathematics, 1369:295– 305, 1989. CODEN LNMAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087543/.

Xu:1987:CS

[Xu87a] Quanhua Xu. Cotype of the spaces (A_{0}, A_{1})_{θ1}. Lecture Notes in Mathematics, 1267:210–212, 1987. CODEN LNMAA2. ISBN 3-540-18103-2 (print), 3-540-47771-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078147/.

Xu:1987:ASY

[Xu87b] Youyu Xu. On absence of static Yang–Mills fields with variant mass. Lecture Notes in Mathematics, 1255:209–220, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077692/.

Yagasaki:1987:FST

[Yag87] Tatsuhiko Yagasaki. Fiber shape theory, shape fibrations and movability of maps. Lecture Notes in Mathematics, 1283:240– 252, 1987. CODEN LNMAA2. ISBN 3-540-18443-0 (print), 3-540-47975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081432/.

Yajima:1985:BM

[Yaj85a] Kenji Yajima. Back matter. Lecture Notes in Mathematics, 1159: 265–281, 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-39706-9/1.pdf.

Yajima:1985:QCL

[Yaj85b] Kenji Yajima. The quasi-classical limit of scattering amplitudefinite range potentials. Lecture Notes in Mathematics, 1159:242– 263, 1985. CODEN LNMAA2. ISBN 3-540-16035-3 (print), 3-540-39706-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080336/.

Yamakami:1989:CCL

[Yam89] Tomoyuki Yamakami. Computational complexity of languages counting random oracles. Lecture Notes in Mathematics, 1388: 189–202, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083671/.

Yan:1987:DDS

[Yan87] J. A. Yan. Développement des distributions suivant les chaos de Wiener et applications à l’analyse stochastique. (French) []. Lecture Notes in Mathematics, 1247:27–32, 1987. CODEN LN-MAA2. ISBN 3-540-17768-X (print), 3-540-47814-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077626/.

Yan:1988:EPQa

[Yan88a] Jia-An Yan. Éléments de probabilités quantiques. IX. Calculs antisymétriques et ^{≪}supersymétriques^{≫} en probabilités. (French) []. Lecture Notes in Mathematics, 1321:101–123, 1988. CODEN LN-MAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084125/.

Yan:1988:EPQb

[Yan88b] Jia-An Yan. Éléments de probabilités quantiques. X Calculs avec des noyaux discrets. (French) []. Lecture Notes in Mathematics, 1321:124–128, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084126/.

Yan:1988:FDT

[Yan88c] Jia-An Yan. A formula for densities of transition functions. Lecture Notes in Mathematics, 1321:92–100, 1988. CODEN LN-MAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084124/.

Yan:1988:PTS

[Yan88d] Jia-An Yan. A perturbation theorem for semigroups of linear operators. Lecture Notes in Mathematics, 1321:89–91, 1988. CODEN LNMAA2. ISBN 3-540-19351-0 (print), 3-540-39228-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084123/.

Yan:1988:ERS

[Yan88e] Ziqian Yan. Everywhere regularity for solutions to quasilinear elliptic systems of triangular form. Lecture Notes in Mathematics, 1306:255–261, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082938/.

Yang:1989:IDS

[Yan89] Wenmao Yang. On infinitesimal deformations of surfaces in E^{3}. Lecture Notes in Mathematics, 1369:306–321, 1989. CODEN LN-MAA2. ISBN 3-540-51037-0 (print), 3-540-46137-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087544/.

Yasue:1988:CSM

[Yas88] Kunio Yasue. Computer stochastic mechanics. Lecture Notes in Mathematics, 1325:188–204, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077928/.

Yasugi:1989:ICI

[Yas89] Mariko Yasugi. Infinitesimal calculus interpreted in infinitary logic. Lecture Notes in Mathematics, 1388:203–220, 1989. CODEN LNMAA2. ISBN 3-540-51527-5 (print), 3-540-48220-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083672/.

Yger:1987:FDP

[Yge87] Alain Yger. Formules de division et prolongement meromorphe. (French) []. Lecture Notes in Mathematics, 1295:226–283, 1987. CODEN LNMAA2. ISBN 3-540-18691-3 (print), 3-540-48084-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081984/.

Yu:1989:PSQ

[YM89] Qingyu Yu and Tian Ma. The positive solutions of quasilinear elliptic boundary value problem. Lecture Notes in Mathematics, 1411:182–188, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086453/.

Yoeurp:1985:TGG

[Yoe85] Chantha Yoeurp. Théorème de Girsanov généralisé et grossissement d’une filtration. (French) []. Lecture Notes in Mathematics, 1118:172–196, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075774/.

Yomdin:1988:ACF

[Yom88] Y. Yomdin. Approximational complexity of functions. Lecture Notes in Mathematics, 1317:21–43, 1988. CODEN LN-MAA2. ISBN 3-540-19353-7 (print), 3-540-39235-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081734/.

Yor:1985:GFA

[Yor85a] M. Yor. Grossissement de filtrations et absolue continuité de noyaux. (French) []. Lecture Notes in Mathematics, 1118:6–14, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075766/;
 http://link.springer.com/chapter/10.1007/BFb0075767/.

Yor:1985:IMCa

[Yor85b] M. Yor. Inégalités de martingales continues arrêtées à un temps quelconque. (French) []. Lecture Notes in Mathematics, 1118:110– 146, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075772/.

Yor:1985:EPG

[Yor85c] Marc Yor. Entropie d’une partition, et grossissement initial d’une filtration. (French) []. Lecture Notes in Mathematics, 1118:45– 58, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075770/.

Yor:1985:IMCb

[Yor85d] Marc Yor. Inégalités de martingales continues arrêtées à un temps quelconque: Le rôle de certains espaces BMO. (French) []. Lecture Notes in Mathematics, 1118:147–171, 1985. CODEN LN-MAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075773/.

Yi-ping:1989:EEP

[Yp89] Zhang Yi-ping. Ensembles équivalents à un point frontière dans un domaine lipschitzien. (French) []. Lecture Notes in Mathematics, 1393:256–265, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085785/.

Yui:1985:BGP

[Yui85] Noriko Yui. The Brauer group of the product of two curves over a finite field. Lecture Notes in Mathematics, 1135:254–283, 1985. CODEN LNMAA2. ISBN 3-540-15649-6 (print), 3-540-39535-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074609/.

Yui:1988:JQL

[Yui88] Noriko Yui. Jacobi quartics, Legendre polynomials and formal groups. Lecture Notes in Mathematics, 1326:182–215, 1988. CO-DEN LNMAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078046/.

Yukich:1986:MEC

[Yuk86] J. E. Yukich. Metric entropy and the central limit theorem in Banach spaces. Lecture Notes in Mathematics, 1193:113–128, 1986. CODEN LNMAA2. ISBN 3-540-16487-1 (print), 3-540-39826-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077103/.

Zabczyk:1987:EPI

[Zab87] J. Zabczyk. Exit problem for infinite dimensional systems. Lecture Notes in Mathematics, 1236:239–257, 1987. CODEN LN-MAA2. ISBN 3-540-17211-4 (print), 3-540-47408-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0072894/.

Zagier:1985:MPM

[Zag85] D. Zagier. Modular points, modular curves, modular surfaces and modular forms. Lecture Notes in Mathematics, 1111:225– 248, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084592/.

Zagier:1988:NLS

[Zag88] Don Zagier. Note on the Landweber–Stong elliptic genus. Lecture Notes in Mathematics, 1326:216–224, 1988. CODEN LN-MAA2. ISBN 3-540-19490-8 (print), 3-540-39300-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078047/.

Zahid:1989:CNS

[Zah89] M. Zahid. Comparaison de noyaux de Schrödinger. (French) [Comparison of Schrödinger nuclei]. Lecture Notes in Mathematics, 1393:243–255, 1989. CODEN LNMAA2. ISBN 3-540-51592-5 (print), 3-540-46675-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085784/.

Zalesskii:1988:EMC

[Zal88] A. E. Zalesskiĭ. Eigenvalues of matrices of complex representations of finite groups of Lie type. Lecture Notes in Mathematics, 1352:206–218, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082027/.

Zambrini:1988:NPA

[Zam88a] J.-C. Zambrini. New probabilistic approach to the classical heat equation. Lecture Notes in Mathematics, 1325:205–220, 1988. CODEN LNMAA2. ISBN 3-540-50015-4 (print), 3-540-45887-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077929/.

Zamir:1988:TNG

[Zam88b] Shmuel Zamir. Topics in noncooperative game theory. Lecture Notes in Mathematics, 1330:72–128, 1988. CODEN LN-MAA2. ISBN 3-540-50003-0 (print), 3-540-45883-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0078159/.

Zappa:1988:NFC

[Zap88] Guido Zappa. Normal fitting classes of groups and generalizations. Lecture Notes in Mathematics, 1352:219–226, 1988. CODEN LN-MAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082028/.

Zarati:1985:QPD

[Zar85] S. Zarati. Quelques propriétés du foncteur Hom_{u_{p}}(, H^{∗}V). (French) []. Lecture Notes in Mathematics, 1172:204–209, 1985. CODEN LNMAA2. ISBN 3-540-16061-2 (print), 3-540-39745-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074433/.

Zayed:1987:UMS

[Zay87] Maher Zayed. Ultraproduits et modules sans facteurs directs de type fini. (French) []. Lecture Notes in Mathematics, 1296:312– 324, 1987. CODEN LNMAA2. ISBN 3-540-18690-5 (print), 3-540-48081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078534/.

Zelmanov:1988:NNA

[Zel88] E. I. Zel’manov. On the nilpotency of nil algebras. Lecture Notes in Mathematics, 1352:227–240, 1988. CODEN LNMAA2. ISBN 3-540-50371-4 (print), 3-540-45994-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082029/.

Zenisek:1986:SNC

[Žen86] A. Ženíšek. Some new convergence results in finite element theories for elliptic problems. Lecture Notes in Mathematics, 1192: 353–358, 1986. CODEN LNMAA2. ISBN 3-540-16469-3 (print), 3-540-39807-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076093/.

Zeuner:1986:HH

[Zeu86] Hansmartin Zeuner. On hyperbolic hypergroups. Lecture Notes in Mathematics, 1210:216–224, 1986. CODEN LNMAA2. ISBN 3-540-16806-0 (print), 3-540-44852-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077186/.

Zhou:1988:IVP

[ZG88] Yulin Zhou and Boling Guo. Initial value problems for a nonlinear singular integral-differential equation of deep water. Lecture Notes in Mathematics, 1306:278–290, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082940/.

Zhang:1988:DPC

[Zha88] Ke-Wei Zhang. On the Dirichlet problem for a class of quasilinear elliptic systems of partial differential equations in divergence form. Lecture Notes in Mathematics, 1306:262–277, 1988. CODEN LNMAA2. ISBN 3-540-19097-X (print), 3-540-39107-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082939/.

Zhao:1989:BM

[Zha89a] Xuezhi Zhao. Back matter. Lecture Notes in Mathematics, 1411: ??, 1989. CODEN LNMAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-46862-2/1.pdf.

Zhao:1989:RNN

[Zha89b] Xuezhi Zhao. A relative Nielsen number for the complement. Lecture Notes in Mathematics, 1411:189–199, 1989. CODEN LN-MAA2. ISBN 3-540-51932-7 (print), 3-540-46862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086454/.

Zheng:1987:HKF

[Zhe87] Yong Fan Zheng. Harmonic and Killing forms on complete Riemannian manifolds. Lecture Notes in Mathematics, 1255:228– 243, 1987. CODEN LNMAA2. ISBN 3-540-17849-X (print), 3-540-47883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077694/.

Zinsmeister:1989:DAF

[Zin89] M. Zinsmeister. Les dérivations analytiques. (French) []. Lecture Notes in Mathematics, 1372:21–46, 1989. CODEN LN-MAA2. ISBN 3-540-51191-1 (print), 3-540-46176-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083957/.

Zheng:1986:CCD

[ZM86] W. A. Zheng and P.-A. Meyer. Sur la construction de certaines diffusions. (French) []. Lecture Notes in Mathematics, 1204:334– 337, 1986. CODEN LNMAA2. ISBN 3-540-16779-X (print), 3-540-39860-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075724/.

Zolezzi:1986:SAO

[Zol86] T. Zolezzi. Stability analysis in optimization. Lecture Notes in Mathematics, 1190:397–419, 1986. CODEN LNMAA2. ISBN 3-540-16476-6 (print), 3-540-39817-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076716/.

Zubr:1988:CSC

[Žub88] A. V. Žubr. Classification of simply-connected topological 6manifolds. Lecture Notes in Mathematics, 1346:325–339, 1988. CODEN LNMAA2. ISBN 3-540-50237-8 (print), 3-540-45958-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082781/.

Zvyagin:1988:NSC

[Zvy88] V. G. Zvyagin. On the number of solutions for certain boundaryvalue problems. Lecture Notes in Mathematics, 1334:157–172, 1988. CODEN LNMAA2. ISBN 3-540-50019-7 (print), 3-540-45894-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080427/.

Zwicker:1989:BSP

[Zwi89] William S. Zwicker. A beginning for structural properties of ideals on P_{κ}λ. Lecture Notes in Mathematics, 1401:201–217, 1989. CODEN LNMAA2. ISBN 3-540-51730-8 (print), 3-540-46795-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097341/.